A Complete Bibliography of Publications in The Journal of Supercomputing: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

19 April 2024 Version 3.28

Title word cross-reference

(K_{9} − C_{9})^{n} [2234]. (n, n) [2492]. 1 [2282]. 192 [712]. 2 [577]. 2n [704]. 3 [215, 257, 492, 623, 2876]. 8 [712]. ^{2} [2576]. ^{∘} [283]. _{1} [665]. _{2} [665]. A∗ [593, 1246]. α [602]. C [733, 1127, 2329]. D [326]. δ [422]. g [1309, 2048, 2498]. GF (p) [2576]. h [1299, 1632, 2098, 2633]. H_{∞} [764, 1571]. K [313, 728, 789, 981, 1168, 1211, 1250, 1311, 1312, 1385, 1454, 1660, 2172, 2295, 2366, 2381, 2467, 2636, 2814, 2870]. K_{4} [2633]. kα [1751]. l [1168]. L_{2} [1154]. LU [2540]. N [961]. N [1019, 1311, 1312, 1318, 1632, 2282, 2295, 2633]. O(log_{2} N) [1299]. P_{k} [1644]. Q [627, 1770, 1960, 2368, 2715, 2799, 2840]. QR [455]. R [31]. s^{+} [2078]. σ [1016]. t/s [1311, 1312]. τ [1494]. Z [1276, 1723].

-ANFIS [1016]. -ary [1311, 1312, 2295]. -axis [1723]. -based [450, 917]. -bit [704, 712]. -body [1318]. -connectivity [1644]. -Cover [2814]. -cube [1019, 1311, 1312, 1632]. -cubes [2295]. -D [577, 2876]. -diagnosability [1311, 1312]. -disjoint [789]. -diversity [1168]. -dominating [1211]. -edge-connected [2381]. -extra [1299, 1309, 1632, 2633]. -flexible [1660]. -good-neighbor [2048, 2498]. -learning [31, 2840]. -learning-based [1770, 2368, 2715]. -Lop [1494]. -maximal-clique [1751]. -mean [1127, 1454]. -means [728, 733, 981, 1168, 1250, 2329, 2467, 2636, 2870]. -means-based [2172]. -nearest [313]. -network [1960].

-numbers [1276]. -Probabilistic [602]. -Restricted [2098]. -rung [2799]. -shell [2366]. -structure [2098]. -substructure [2098]. -SVM [1154]. -th [2633]. -Transitive [422]. -version [961].

/M [665].

0/1 [1559].

1-bit [1726]. 10 [2261]. 14-dimensional [2697]. 19 [800, 973, 1136, 1314, 1325, 1423, 1493, 1872, 2039, 2113, 2175, 2203, 2393, 2508, 2570, 2618, 2708, 2762]. 1901 [123]. 1D [1579, 1831]. 1DCNN [1831].

2 [1586, 1949]. 2-D [1949]. 2-disjoint [1869]. 2-phase [1913]. 2-tuple [2799]. 2.0 [412]. 24-by-24 [2610]. 256 [1854]. 2D [1418, 2434, 2483, 2648, 2722, 2846]. 2DSPP [1502]. 2PBDC [285].

3 [1691]. 3-ary [1019]. 3-D [1691]. 3-disjoint [1587]. 32-bit [1760]. 3D [812, 1097, 1351, 1401, 1438, 1507, 1647, 1669, 1739, 1740, 1743, 1973, 2036, 2238, 2372, 2682, 2783]. 3D-based [1973, 2036]. 3D-CICCM [2783]. 3D-memories [1739].

4-bit [2627]. 4-digit [1544]. 4-Valued [1769]. 4.0 [2863]. 4D [1529, 2879]. 4D-ACSM [1529].

512 [91]. 5G [158, 740, 935, 1121, 1659, 1678, 1686, 1840, 1912, 2031, 2331, 2409]. 5GMAKA [1912].

6 [1015, 1856]. 6G [2398]. 6LoWPAN [2252]. 6LoWPAN-based [2252].

754 [2898].

802.11ah [2023].

A-CTO [1748]. A-NFVLearn [2389]. A3 [563]. A3-Storm [563]. AAR [2890]. abandonment [1697]. ABC [208, 1989]. ABC-BA [208, 1989]. ABE [1332, 2237, 2282, 2410]. ABGS [2075]. ability [547]. ablation [1465]. Abnormal [214, 646, 1551]. Absolute [839, 1255, 2257]. absolute-value [2257]. absorbing [389]. abstraction [302, 843, 1281]. abstractive [1164, 2742]. academic [375]. accelerate [170, 1501, 1774, 2541]. accelerated [431, 485, 793, 901, 1445, 1658, 1691, 1950, 2249, 2352, 2507]. Accelerating [239, 427, 574, 755, 1471, 2056, 2062, 2065]. Acceleration [413, 451, 498, 1063, 1577, 2161, 2321, 2444]. accelerator [1084, 1086, 1180, 1576, 2476, 2774, 2806]. accelerator-rich [1084]. accelerators [360, 509, 907, 1318, 1881, 1935, 2047, 2337]. accelerometers [840]. acceptance [1624]. Access [139, 304, 305, 307, 546, 1058, 1067, 1117, 1118, 1332, 1416, 1423, 1444, 1559, 1633, 1771, 1774, 1953, 2175, 2231, 2237, 2334, 2500, 2608, 2625]. accesses [1795]. accessibility [2764]. accident [2411, 2573, 2672]. accommodate [990]. according [791, 2239]. account [286]. accuracy [607, 780, 874, 1786, 1891, 2086, 2408, 2456, 2584, 2861]. accuracy-aware [2086]. Accurate [292, 718, 800, 1773, 1836, 2039, 2734, 2850]. ACEP [707]. achievement [1263, 2713]. Achieving [2525]. ACNN [1419]. ACNN-TL [1419]. ACO [1176, 1554, 1732, 2716, 2841]. ACO-based [1176]. ACO-list [1732]. acoustic [717, 2495]. acquisition [2726]. Across [185, 1060, 1279, 2467]. Across-camera [1060]. across-wind [185]. ACSM [1529]. action [94, 1059, 1725, 2091, 2517, 2696]. action-set [1725]. activation [2537, 2584]. Active [112, 441, 806, 1174, 1515, 2020, 2656]. active-lane [1515]. activities [985, 1391]. activity [93, 415, 606, 647, 711, 831, 852, 1105, 1120, 1253, 1297]. activity-based [831]. actor [2601]. actors [283]. ACTSSD [1174]. actuator [621, 1571, 2677]. acupuncture [1265]. acute [1146, 1462, 1464]. acyclic [1841]. ad [248, 346, 349, 363, 796, 929, 1661, 1666, 1998, 2290]. ADA [2298]. AdaBoost [493, 1672, 1716]. AdaInNet [1828]. Adaline [595]. adaptation [1378, 1535, 1952, 2058, 2298]. Adapted [346, 743]. Adapting [408]. adaption [76, 691, 1586]. Adaptive [64, 122, 174, 256, 337, 450, 499, 562, 592, 600, 621, 642, 662, 694, 707, 855, 872, 925, 1023, 1202, 1277, 1298, 1300, 1304, 1305, 1450– 1452, 1475, 1476, 1554, 1679, 1773, 1826, 1828, 2006, 2015, 2019, 2023, 2070, 2075, 2088, 2103, 2143, 2184, 2191, 2192, 2247, 2281, 2290, 2341, 2353, 2354, 2362, 2365, 2424, 2437, 2527, 2535, 2575, 2679, 2760, 2768, 2794, 2822, 2836, 2842, 2863, 2896]. Adaptively [2282, 2626]. ADAS [2001]. AdaXod [2437]. adder [96, 821, 1474, 2073, 2630]. adder-subtractor [821]. adders [724, 1347, 1544, 1718, 2853, 2898]. adders/subtractors [2898]. address [2190]. addressable [2427]. Addressing [1637, 2182]. ADELA [2436]. adjusted [2626]. Adjusting [258, 259, 1529, 1749, 1974]. adjustment [1859]. Admission [1631]. ADMM [829]. ADOL [527]. Adoption [587, 820, 850, 957, 1210, 1499]. ADS [1852]. ADS-B [1852]. Advanced [56, 107, 184, 828, 923, 1352, 1373, 1429, 1682, 1730, 2161, 2389, 2856, 2875, 2915]. Advancements [2862, 2913]. adversarial [359, 463, 1029, 1818, 1837, 1876, 1970, 2042, 2210, 2294, 2298, 2403, 2533, 2785]. advertisement [2345]. advertisements [1959]. AEGA [2277]. Aerial [157, 720, 1255, 1594, 2304, 2377, 2908]. aeroacoustics [1396]. AERQP [2679]. AES [1879, 2007, 2732]. Affective [2873]. affine [1961]. affinity [1596, 1674]. AFI [348]. after [919]. against [288, 365, 374, 452, 465, 1037, 1957, 2113, 2207, 2359, 2637]. age [315, 860, 1324, 1325, 2301]. age-based [2301]. age-related [1324]. age-structured [1325]. Agent [161, 194, 254, 343, 368, 1017, 1267, 1431, 1719, 1810, 2003, 2176, 2604, 2651, 2804]. Agent-based [194, 368, 2176, 2804]. agents [583, 629, 720]. agglomeration [1267]. agglomerative [773]. aggrandized [1748]. aggregate [602, 1660, 2335]. aggregates [13]. aggregation [202, 505, 719, 1071, 1109, 1119, 1266, 1558, 1684, 1871, 1906, 2167, 2307, 2542, 2895]. Aggressive [1973, 2036]. agile [841, 2432, 2455]. aging [592, 1530]. agreement [652, 740, 828, 864, 1035, 1096, 1200, 1240, 1375, 1496, 1525, 1677, 1912, 2207, 2664]. agricultural [617, 1092, 1243, 1448, 1590, 1939, 2233, 2795]. agriculture [553]. AGV [1172]. ahead [928, 946, 2405, 2711]. AHP [543, 929, 1238, 1397, 1778, 2455]. AHP-based [1778]. AHP-TOPSIS [1238, 1397]. AHP-TOPSIS-based [543, 929]. AI [218, 595, 661, 725, 848, 1404, 1555, 1999, 2456, 2542, 2734, 2845]. AI-based [595, 725, 2845]. AI-driven [2542]. AI-enabled [218, 1555]. aid [2537]. aided [100, 147, 413, 794, 1402, 2265, 2670, 2776]. AIEMLA [1039]. AIOps [2212]. AIoT [1183]. Air [102, 183, 313, 328, 687, 710, 741]. Airbnb [1024]. aircraft [2077]. ALADMM [829]. alerts [2792]. algae [2206]. Algorithm [22, 44, 95, 162, 227, 235, 278, 331, 391, 397, 408, 482, 506, 541, 550, 590, 593, 629, 632, 762, 771, 783, 835, 905, 923, 931, 966, 1001, 1020, 1027, 1044, 1106, 1133, 1149, 1172, 1182, 1210, 1212, 1217, 1218, 1235, 1246, 1282, 1320, 1372, 1377, 1423, 1454, 1466, 1482, 1502, 1503, 1627, 1645, 1653, 1661, 1684, 1688, 1748, 1765, 1766, 1823, 1908, 1923, 1928, 1957, 2019, 2025, 2026, 2040, 2049, 2104, 2119, 2286, 2327, 2342, 2373, 2402, 2459, 2467, 2481, 2546, 2565, 2602, 2604, 2628, 2643, 2669, 2675, 2702, 2727, 2731, 2737, 2783, 2794, 2797, 2813, 2857, 2882, 2920]. algorithm [5, 74, 108, 195, 199, 255, 280, 308, 319, 362, 366, 383, 472, 478, 479, 497, 543, 576, 599, 605, 636, 639, 649, 662, 703, 708, 733, 747, 755, 757, 761, 790, 803, 809, 867, 874, 890, 913, 914, 917, 929, 956, 964, 969, 980, 1048, 1051, 1080, 1101, 1147, 1190, 1204, 1238, 1250, 1269, 1277, 1429, 1435, 1463, 1476, 1492, 1534, 1580, 1582, 1672, 1704, 1729, 1735, 1799, 1878, 1890, 1900, 1971, 1991, 2004, 2030, 2130, 2153, 2175, 2236, 2247, 2253, 2270, 2283, 2362, 2413, 2437, 2444, 2470, 2488, 2497, 2504, 2562, 2564, 2596, 2606, 2651, 2687, 2689, 2718, 2802, 2846, 2861, 2872, 2876]. algorithm [34, 63, 68, 180, 197, 209, 213, 214, 242, 290, 330, 484, 526, 552, 556, 568, 631, 673, 754, 759, 769, 788, 829, 873, 901, 925, 933, 938–940, 947, 981, 1018, 1028, 1070, 1077, 1088, 1125, 1179, 1262, 1273, 1294, 1304, 1305, 1310, 1346, 1351, 1397, 1424, 1451, 1453, 1475, 1518, 1574, 1585, 1611, 1635, 1646, 1648, 1656, 1707, 1715, 1738, 1757, 1787, 1807, 1834, 1898, 1903, 1910, 1946, 2015, 2027, 2028, 2031, 2155, 2173, 2179, 2183, 2206, 2259, 2277, 2351, 2434, 2503, 2520, 2531, 2560, 2580, 2610, 2636, 2705, 2706, 2708, 2733, 2755, 2782, 2814, 2823, 2833, 2841, 2843, 2856]. algorithm [18, 24, 38, 43, 53, 151, 166, 206, 207, 260, 277, 281, 353, 354, 396, 454, 463, 495, 633, 685, 820, 857, 977, 991, 1019, 1031, 1085, 1219, 1223, 1259, 1299, 1315, 1316, 1389, 1428, 1488, 1490, 1511, 1512, 1573, 1596, 1638, 1658, 1674, 1708, 1759, 1835, 1863, 1866, 1888, 1994, 2050, 2055, 2078, 2092, 2108, 2149, 2204, 2244, 2258, 2325, 2355, 2382, 2461, 2471, 2499, 2512, 2513, 2525, 2535, 2586, 2616, 2631, 2698, 2716, 2724, 2787, 2840, 2880, 2881, 2900]. algorithm-based [43, 44, 108, 552, 1088, 1423, 2175, 2327]. algorithm-optimized [2689]. algorithmic [300]. Algorithms [25, 64, 67, 145, 239, 291, 299, 328, 364, 365, 372, 373, 379, 414, 492, 532, 533, 557, 569, 591, 726, 813, 886, 965, 967, 986, 1014, 1062, 1082, 1086, 1087, 1125, 1177, 1238, 1256, 1271, 1365, 1385, 1430, 1450, 1473, 1532, 1709, 1770, 1947, 2147, 2150, 2163, 2166, 2227, 2242, 2278, 2310, 2312, 2319, 2372, 2529, 2530]. aliasing [1868]. alignment [458, 472, 926, 1140, 1587, 1815, 2105, 2772]. alike [659]. all-optical [811, 1074]. alleviating [736]. alliance [1043, 2491]. Allocating [1080, 2800]. allocation [44, 61, 144, 194, 197, 349, 382, 513, 525, 592, 625, 1047, 1048, 1056, 1067, 1103, 1412, 1415, 1519, 1675, 1686, 1710, 1715, 1821, 1943, 2025, 2164, 2170, 2176, 2278, 2331, 2379, 2423, 2487, 2730, 2881]. allowlist [1662]. allreduce [829]. along [1131]. alternate [575]. alternating [1624, 2585]. alternative [377]. altitude [1594]. ALU [85, 511, 1033, 1111, 1112, 1760, 2089, 2448]. alumina [2738]. Alzheimer [2753]. Amazon [345, 2198]. ambient [1007, 1827]. Ameliorated [2405]. Amended [1106]. AMI [904, 2875]. AMO [1188]. among [1412, 1584, 1757, 1777]. amplifier [811]. amplifier-based [811]. amplifiers [1074]. AMT [1394]. Analysing [1114]. Analysis [13, 17, 32, 52, 59, 85, 91, 94, 111, 115, 123, 145, 156, 169, 184, 199, 200, 208, 221, 223, 245, 264, 327, 328, 379, 398, 425, 570, 595, 611, 617, 645, 665, 666, 701, 704, 741, 747, 777, 804, 815, 833, 834, 840, 897, 919, 945, 1009, 1013, 1081, 1129, 1131, 1224, 1232, 1265, 1284, 1319, 1341, 1362, 1448, 1510, 1521, 1529, 1532, 1560, 1577, 1583, 1600, 1602, 1617, 1619, 1632, 1667, 1804, 1877, 1885, 1939, 1977, 1987, 1989, 2081, 2084, 2110, 2220, 2227, 2264, 2327, 2346, 2468, 2487, 2505, 2506, 2516, 2611, 2723, 2726, 2732, 2761, 2845, 2851, 2898, 2915–2917]. analysis [49, 53–55, 102, 187, 209, 213, 215, 281, 377, 450, 451, 470, 539, 577, 578, 655, 671, 684, 746, 773, 805, 806, 820, 844, 857, 958, 1046, 1135, 1155, 1162, 1216, 1231, 1264, 1425, 1490, 1518, 1638, 1656, 1717, 1750, 1793, 1803, 1848, 1912, 1969, 2005, 2008, 2031, 2034, 2161, 2201, 2214, 2266, 2277, 2298, 2426, 2471, 2479, 2509, 2515, 2589, 2714, 2779, 2822, 2837, 2849, 2879, 2889]. analytic [65, 217, 1988]. Analytical [722, 893, 1024, 1153, 1158, 1498]. analytics [77, 78, 107, 357, 442, 751, 1064, 1166, 1217, 1218, 1270]. analyze [1455, 2911]. Analyzing [540, 549, 693, 1434, 2590]. ancilla [1089]. Android [499, 2727, 2835]. ANFIS [160, 450, 1016, 1692]. ANFIS-based [160]. Angle [1723]. angular [839]. ankle [840, 1192]. ankle-worn [840]. ANN [2430, 2659]. annealing [180, 977, 1076, 1235, 1635]. annotating [2769]. annotation [246, 247, 255, 860, 2385]. annulus [1743]. anomalies [1214, 1831, 2464]. Anomaly [220, 632, 1011, 1209, 1361, 1555, 1893, 2222, 2333, 2505, 2535, 2598, 2619, 2808, 2839, 2868, 2886]. AnonSURP [1369, 1370]. anonymity [836]. anonymization [2780]. anonymous [584, 1369, 1370, 1855, 2080, 2308, 2809]. ANOVA [2277]. ANP [311]. ANP-based [311]. answer [188, 2534]. answering [688, 1614, 2146, 2297]. answers [222]. ant [43, 231, 245, 362, 593, 782, 1048, 1381, 1578, 2055, 2823, 2843]. ant-based [1578]. antenna [223]. antennae [2166]. Anti [540, 688, 1564, 1868, 2674, 2739, 2792, 2834]. Anti-aliasing [1868]. anti-collusion [2834]. anti-counterfeiting [2739]. anti-interference [1564]. anti-money [2792]. Anti-negation [688]. anti-reversing [540]. anti-vax [2674]. anticipating [2087]. antipatterns [1286]. antlion [1269]. any [1625]. any-order [1625]. AOMDV [1692]. AOMDV-REPO-based [1692]. Apache [1168, 1246, 1258, 2200, 2697]. aperture [251]. API [2666]. app [554, 815]. appeal [1959]. appearance [1648]. Application [15, 48, 65, 101, 107, 144, 162, 176, 194, 243, 254, 312, 349, 475, 484, 499, 515, 623, 649, 701, 748, 764, 786, 833, 881, 924, 961, 1090, 1154, 1162, 1263, 1276, 1293, 1310, 1320, 1336, 1339, 1363, 1404, 1408, 1427, 1476, 1480, 1510, 1620, 1626, 1665, 1727, 1805, 1825, 1898, 1919, 2031, 2055, 2176, 2209, 2306, 2481, 2502, 2516, 2586, 2610, 2669, 2727, 2755, 2775, 2782, 2845, 2884]. Application-specific [786]. Applications [9, 109, 170, 361, 395, 417, 490, 504, 571, 657, 679, 726, 735, 744, 769, 860, 922, 979, 1023, 1045, 1069, 1124, 1150, 1158, 1216, 1280, 1340, 1439, 1446, 1460, 1610, 1618, 1628, 1629, 1653, 1654, 1750, 1761, 1828, 1873, 1882, 1891, 1892, 1902, 1920, 1960, 2032, 2056, 2068, 2123, 2141, 2150, 2166, 2275, 2300, 2312, 2431, 2514, 2627, 2640, 2650, 2651, 2687, 2712, 2735, 2764, 2879]. Applied [17, 74, 219, 575, 1358, 1573, 1941, 1945]. Applying [568, 594, 604, 611, 1044, 1125, 1192, 1645]. Approach [27, 81, 104, 107, 109, 143, 145, 190, 221, 244, 348, 360, 378, 386, 423, 517, 529, 549, 611, 635, 642, 670, 714, 722, 734, 749, 750, 841, 869, 874, 983, 1093, 1095, 1120, 1130, 1156, 1157, 1167, 1197, 1198, 1207, 1232, 1319, 1331, 1390, 1420, 1422, 1440, 1442, 1507, 1575, 1578, 1626, 1689, 1703, 1791, 1796, 1816, 1823, 1849, 1860, 1880, 1905, 1947, 1956, 1983, 2053, 2072, 2073, 2075, 2085, 2086, 2090, 2101, 2119, 2128, 2133, 2169, 2240, 2280, 2317, 2386, 2458, 2500, 2516, 2530, 2577, 2578, 2581, 2681, 2697, 2710, 2719, 2734, 2834, 2847, 2887]. approach [15, 40, 57, 82, 92, 161, 215, 229, 245, 271, 273, 433, 491, 494, 572, 773, 779, 785, 786, 801, 810, 856, 875, 877, 958, 985, 1039, 1062, 1213, 1223, 1241, 1264, 1302, 1342, 1349, 1447, 1556, 1605, 1642, 1664, 1719, 1746, 1810, 1817, 1841, 1843, 1911, 2018, 2038, 2066, 2092, 2099, 2202, 2265, 2273, 2290, 2364, 2374, 2417, 2472, 2487, 2513, 2533, 2553, 2592, 2599, 2629, 2645, 2647, 2656, 2776, 2779, 2799, 2906, 2908]. approach-based [2553, 2629]. approaches [67, 339, 514, 619, 677, 867, 906, 1248, 1539, 1698, 1706, 1815, 1889, 2057, 2384, 2426, 2468, 2703, 2713, 2793, 2907]. appropriate [2777]. approval [179]. Approximate [455, 1471, 1627, 1833, 1891, 2118, 2483, 2701, 2715, 2853]. approximation [885, 2606]. approximations [554]. Apriori [526, 1766, 2259]. APT [1373, 2359]. APT-Dt-KC [1373]. AQL [2715]. aqua [2665]. aquaculture [2228]. aquaponics [2519]. Aquila [1420]. Arabia [1325]. arbitrarily [206, 2105]. arbitrary [958, 2339]. arbitrary-oriented [2339]. arc [1102, 1623]. arc-consistency [1623]. Archimedes [1904, 2562]. Architecting [1924]. architectural [660, 744, 1098, 1788]. architecture [23, 145, 225, 345, 501, 537, 573, 578, 644, 704, 730, 851, 946, 1000, 1004, 1037, 1057, 1069, 1080, 1178, 1187, 1240, 1328, 1402, 1421, 1470, 1521, 1634, 1760, 1838, 1870, 1919, 1942, 1963, 2016, 2062, 2152, 2180, 2213, 2330, 2386, 2554, 2566, 2635, 2650, 2765, 2774, 2798, 2869, 2885]. architecture-based [345]. architectures [12, 85, 231, 437, 614, 1348, 1469, 1473, 1515, 1523, 2064, 2467, 2662, 2709, 2771, 2794]. archive [2481]. ARCSA [215]. ARdetector [2835]. area [410, 659, 926, 950, 1032, 1274, 1540, 1791, 1839, 1971, 2089, 2162, 2448, 2554, 2609, 2662, 2853]. area-efficient [1839, 2448, 2662]. areas [757, 1840]. ARGhost [2684]. ARIMA [604, 1666]. arithmetic [737, 1073, 1282, 1635, 1661, 1759, 1790, 1839, 2363, 2588, 2616, 2882]. arm [89, 746, 810, 2044, 2147]. army [1334]. arousal [1754]. ArRaNER [1668]. array [775, 2818]. arrays [1113, 1919, 2208, 2903]. arrhythmia [344, 351, 1560, 2086]. arrival [759]. art [573, 641, 1098, 2867, 2897]. arterial [919, 1465]. artery [1463]. artifact [1556]. Artificial [37, 185, 214, 245, 372, 478, 572, 591, 850, 1018, 1039, 1110, 1170, 1225, 1478, 1480, 1514, 1536, 1742, 1746, 1755, 1756, 1775, 1816, 1933, 2049, 2206, 2453, 2604, 2727, 2738, 2761, 2784, 2867, 2892, 2897]. ary [1019, 1311, 1312, 2295]. ASBA [1977]. ASIC [1152]. ASM [2486]. ASM-SDN [2486]. Aspect [1215, 1606, 1803, 1972, 1977, 2008, 2142, 2205, 2214, 2678, 2837, 2873, 2889, 2899]. aspect-based [1215, 1803, 1977, 2008, 2889]. Aspect-level [1606, 1972, 2142, 2205, 2837]. aspect-merged [2142]. aspects [2214]. asphalt [572]. assembler [2427]. assembly [2339]. assess [233]. assessing [2269, 2508]. assessment [17, 105, 115, 119, 176, 222, 560, 825, 1152, 1252, 1299, 1355, 1644, 1692, 1911, 2103, 2380, 2621, 2741]. assets [1391]. assigning [1529]. assignment [656, 778, 1082, 1086, 1584, 2331]. assimilation [1737]. assist [250]. assistance [1233]. assistant [2436]. assisted [158, 280, 647, 1200, 1328, 1496, 1798, 1850, 1875, 2330, 2518, 2764]. assisting [733]. association [1262, 1455, 2023, 2750, 2761]. association-based [2750]. assurance [2617]. assured [1388]. Asymmetric [186, 504, 600, 1720, 2180]. asynchronous [384, 829, 1270, 1394]. Ate [2662]. athlete [1263]. atmospheric [791]. atrophic [850]. atrophy [508]. atrous [2822]. Attack [32, 136, 211, 288, 346, 705, 989, 1037, 1226, 1601, 1607, 1685, 2007, 2026, 2169, 2359, 2380, 2574, 2732, 2832]. attack-tolerant [2732]. attacked [2533]. attackers [2026]. attacks [121, 268, 365, 465, 612, 709, 810, 1039, 1349, 1641, 1661, 1864, 1873, 2168, 2637, 2777]. Attention [60, 71, 646, 686, 1164, 1354, 1361, 1411, 1419, 1517, 1520, 1551, 1606, 1657, 1677, 1846, 1895, 1899, 1964, 1977, 2033, 2090, 2091, 2214, 2416, 2443, 2463, 2484, 2494, 2543, 2622, 2665, 2668, 2670, 2673, 2691, 2724, 2729, 2743, 2832, 2837, 2860, 2865, 2868, 2890, 2899, 2912]. Attention-based [1361, 1411, 1419, 1520, 1657, 2090, 2091, 2543, 2622, 2865]. attention-LSTM [2665]. attentional [1441, 2780]. attentions [2297]. attentive [738, 1972, 2394]. attenuation [1826]. attitudes [2536]. attractiveness [1763]. attractor [2879]. Attribute [38, 289, 304, 305, 308, 403, 695, 738, 833, 882, 1118, 1444, 2334, 2532, 2559, 2625, 2796]. Attribute-aware [695, 738]. attribute-based [289, 305, 308, 403, 1118, 1444, 2334, 2532, 2625]. attributes [318, 1391, 1643, 2411, 2573, 2672, 2731, 2831]. auction [388, 1995, 2308, 2834]. auction-based [1995]. Audio [458, 792, 1486, 2402, 2592]. audio-to-score [458]. Auditable [304, 1444]. Auditing [2, 279, 288, 715, 1007, 1736, 1862]. AUGChain [1301]. Augmentation [1734]. Augmented [48, 566, 1299, 1887, 1986, 2188, 2295]. Augmenting [2205]. Aurora [851]. auscultatory [344]. authentic [2081, 2388]. Authenticated [1045, 1200, 1496, 1526, 1896, 2080, 2447]. Authentication [45, 137, 243, 248, 374, 453, 565, 584, 652, 700, 796, 836, 845, 1035, 1068, 1075, 1091, 1096, 1240, 1274, 1287, 1301, 1449, 1525, 1569, 1574, 1650, 1651, 1655, 1663, 1677, 1701, 1794, 1848, 1884, 1912, 1982, 2023, 2052, 2081, 2207, 2241, 2252, 2292, 2491, 2518, 2571, 2605, 2607, 2608, 2612, 2634, 2658, 2664, 2676, 2739, 2804, 2805, 2809, 2811, 2821, 2864, 2883, 2914]. authentication-based [2739]. authorities [289]. authority [1117, 2334]. authorization [2804]. authors [407]. autism [619, 1478, 2854]. auto [142, 1280, 2113, 2271, 2724, 2839]. auto-encoder [2724]. auto-encoders [2271]. auto-labeling [2113]. auto-negotiation [142]. auto-tuning [1280]. autoencoder [635, 1493, 1613, 1787, 1932, 2418, 2808]. autoencoder-based [635, 1932]. autoencoders [1890, 2280]. Autofocus [926, 2920]. Automata [96, 127, 278, 547, 585, 747, 790, 865, 944, 950, 1033, 1111, 1112, 1131, 1157, 1303, 1335, 1718, 1725, 1726, 1839, 1879, 1919, 2124, 2217, 2387, 2630, 2816]. automata-based [1157]. Automated [290, 337, 496, 782, 1025, 1308, 1324, 1423, 2021, 2119, 2175, 2486, 2652]. Automatic [46, 222, 272, 344, 361, 527, 588, 716, 1151, 1201, 1231, 1331, 1372, 1440, 1676, 1694, 2051, 2212, 2261, 2496, 2530, 2570, 2686, 2792, 2905]. automatically [1072]. automation [219, 2384]. automaton [952, 1950]. automotive [2806]. Autonomic [537, 1207]. autonomous [61, 426, 583, 1031, 1143, 1344, 1535, 1703, 1728, 2129, 2258]. autoregression [2905]. autoscaling [2746]. Autotuning [502, 532]. Auxiliary [1099, 1164]. Availability [23, 212, 618, 881, 893, 1005, 1345, 1414, 1513, 2569]. Availability-aware [1005]. availability-constrained [212]. availability-scalability [23]. available [2170]. average [516]. average-utility [516]. Averse [1861]. avoidance [781, 1003, 1275, 1489, 1608, 2001, 2561]. AVR [712]. AVR-based [712]. AVS FD MVITS [2432]. AVX [91]. AVX-512 [91]. Aware [21, 30, 34, 88, 159, 228, 332, 376, 386, 406, 462, 465, 469, 475, 481, 515, 541, 562, 563, 575, 623, 714, 738, 742, 778, 779, 816, 922, 936, 968, 971, 972, 983, 993, 1005, 1009, 1061, 1067, 1070, 1080, 1088, 1095, 1097, 1183, 1206, 1224, 1227, 1268, 1339, 1351, 1437, 1489, 1540, 1549, 1554, 1581, 1654, 1670, 1675, 1690, 1704, 1753, 1774, 1785, 1845, 1889, 1928, 1996, 1997, 2085, 2086, 2111, 2128, 2183, 2195, 2215, 2226, 2240, 2251, 2254, 2283, 2312, 2318, 2341, 2342, 2361, 2431, 2441, 2442, 2470, 2475, 2488, 2553, 2556, 2594, 2725, 2797, 2800, 2904]. aware [19, 202, 354, 363, 695, 817, 1189, 1835, 1863, 1874, 1922, 2156, 2214, 2232, 2307, 2599, 2906]. Awareness [61, 292, 888, 1738, 1999, 2143]. AXI [2415]. AXI-interconnect [2415]. axial [1743]. axiom [1276]. axis [921, 1723].

B [1852, 2280]. B-VAE [2280]. BA

[208, 1989]. back [1923]. back-projection [1923]. backbones [2272, 2891]. background [201, 1775, 2075]. backpropagation [875]. backtracking [754]. bad [2186]. badger [2708]. bag [1382, 2438]. bag-of-tasks [1382, 2438]. balance [167, 2453, 2860]. Balanced [562, 674, 1029, 1036, 1309, 1667, 1821, 2011, 2088]. balancer [816, 1016, 1227, 1498, 1581, 1809, 2164]. Balancing [11, 207, 309, 494, 543, 629, 749, 826, 856, 920, 1000, 1077, 1182, 1195, 1213, 1217, 1218, 1238, 1300, 1807, 1905, 2055, 2124, 2280, 2527, 2535, 2557, 2614, 2748, 2820]. balancing-based [1195]. balls [1665, 2306]. band [343]. Bandwidth [382, 1205, 2025, 2164, 2312, 2375, 2424]. bandwidth-aware [2312]. bank [1495]. Barzilai [945, 2596]. base [68, 1398, 2275, 2368, 2376]. baseboard [2866]. Based [14, 22, 42, 81, 95, 221, 225, 244, 249, 278, 286, 307, 361, 380, 427, 445, 564, 583, 587, 626, 668, 670, 696, 784, 792, 825, 828, 923, 931, 991, 999, 1017, 1037, 1057, 1136, 1148, 1215, 1217, 1218, 1246, 1356, 1372, 1392, 1393, 1440, 1444, 1503, 1506, 1560, 1589, 1676, 1687, 1703, 1728, 1751, 1765, 1822, 1864, 1877, 1915, 1924, 1926, 1947, 1980, 2025, 2051, 2065, 2075, 2081, 2119, 2190, 2194, 2227, 2233, 2256, 2319, 2343, 2383, 2401, 2458, 2493, 2532, 2539, 2561, 2563, 2581, 2603, 2641, 2676, 2685, 2694, 2796, 2812, 2842, 2845]. based [33, 44, 72, 76, 107, 118, 159, 172, 251, 252, 294, 314, 323, 351, 391, 397, 493, 508, 541, 550, 586, 593, 595, 632, 635, 637, 652, 682, 705, 814, 864, 883, 892, 909, 934, 1096, 1102, 1133, 1167, 1174, 1183, 1205, 1208, 1236, 1260, 1285, 1320, 1468, 1505, 1514, 1553, 1575, 1657, 1692, 1722, 1808, 1831, 1885, 1889, 1894, 1904, 1934, 1940, 1973, 1979, 1981, 1983, 2036, 2049, 2054, 2109, 2193, 2225, 2240, 2248, 2297, 2327, 2391, 2395, 2425, 2428, 2491, 2528, 2536, 2537, 2549, 2568, 2582, 2587, 2601, 2675, 2697, 2703, 2731, 2737, 2780, 2783, 2790, 2813, 2830, 2867, 2882, 2895, 2904, 2910]. based [5, 26, 64, 73, 114, 126, 149, 153, 174, 199, 263, 289, 378, 453, 465, 482, 514, 518, 543, 549, 603, 619, 630, 642, 665, 690, 704, 720, 899, 904, 905, 1043, 1068, 1118, 1144, 1147, 1212, 1225, 1226, 1327, 1348, 1354, 1423, 1520, 1524, 1563, 1578, 1621, 1633, 1645, 1685, 1709, 1752, 1772, 1784, 1828, 1855, 1923, 1928, 1932, 1956, 1958, 1960, 1991, 1996, 2040, 2058, 2090, 2096, 2104, 2116, 2139, 2175, 2180, 2200, 2222, 2243, 2246, 2267, 2268, 2331, 2358, 2366, 2370, 2414, 2464, 2481, 2486, 2502, 2569, 2612, 2626, 2646, 2661, 2786, 2794, 2804, 2805, 2825, 2834, 2850, 2871, 2909]. based [4, 108, 131, 156, 160, 243, 256, 272, 311, 319, 322, 326, 340, 343, 346, 410, 414, 415, 420, 426, 496, 510, 512, 519, 561, 579, 605, 639, 651, 657, 658, 714, 725, 757, 760, 797, 809, 812, 815, 831, 847, 869, 888, 889, 913, 914, 929, 950, 962, 963, 969, 980, 1013, 1024, 1090, 1097, 1142, 1143, 1157, 1173, 1176, 1187, 1253, 1257, 1314, 1341, 1380, 1420, 1460, 1469, 1472, 1555, 1558, 1726, 1833, 1836, 1943, 1992, 2042, 2067, 2144, 2245, 2250, 2323, 2393, 2394, 2435, 2454, 2478, 2562, 2564, 2596, 2663, 2696, 2715, 2718, 2734, 2762, 2828, 2861, 2872, 2915]. based [3, 16, 60, 62, 104, 154, 173, 189, 192, 200, 201, 266, 276, 308, 333, 366, 371, 374, 399, 417, 440, 448, 464, 492, 546, 589, 600, 694, 711, 726, 756, 787, 793, 839, 855, 882, 895, 956, 964, 968, 1004, 1009, 1048, 1072, 1076, 1081, 1092, 1120, 1129, 1190, 1195, 1220, 1237, 1244, 1254, 1270, 1277, 1295, 1300, 1337, 1344, 1361, 1406, 1408, 1517, 1528, 1533, 1534, 1577, 1604, 1615, 1632, 1647, 1655, 1659, 1667, 1672, 1681, 1701, 1711, 1762, 1771, 1851, 1878, 1884, 1930, 1965, 2004, 2012, 2017, 2091, 2115, 2117, 2121, 2166, 2216, 2235, 2247, 2254, 2269, 2354, 2441, 2540, 2567, 2597, 2607, 2615, 2620, 2695, 2711, 2723, 2767, 2784, 2807, 2847, 2865, 2896]. based [10, 82, 137, 186, 216, 242, 290, 296, 305, 335, 345, 384, 458, 503, 515, 617, 712, 713, 715, 749, 750, 766, 816, 849, 878, 881, 917, 1011, 1062, 1093, 1116, 1125, 1198, 1227, 1238, 1241, 1250, 1301, 1302, 1373, 1381, 1390, 1400, 1438, 1450, 1516, 1525, 1540, 1545, 1564, 1570, 1574, 1608, 1617, 1649, 1662, 1719, 1739, 1770, 1827, 1844, 1897, 1900, 1977, 1997, 2007, 2026, 2072, 2079, 2145, 2236, 2238, 2255, 2320, 2337, 2368, 2427, 2434, 2439, 2452, 2480, 2519, 2541, 2551, 2638, 2648, 2654, 2658, 2681, 2722, 2766, 2815, 2848, 2864, 2888, 2889]. based [7, 34, 40, 86, 194, 197, 198, 203, 224, 304, 341, 347, 349, 368, 376, 389, 403, 447, 475, 526, 597, 686, 698, 854, 947, 992, 1046, 1075, 1077, 1088, 1119, 1161, 1163, 1166, 1228, 1229, 1268, 1274, 1304, 1305, 1355, 1360, 1364, 1384, 1397, 1419, 1447, 1571, 1598, 1616, 1641, 1663, 1679, 1702, 1716, 1758, 1787, 1795, 1811, 1814, 1838, 1853, 1873, 1935, 1949, 2027, 2038, 2044, 2069, 2136, 2157, 2158, 2170–2172, 2176, 2183, 2447, 2457, 2483, 2492, 2495, 2515, 2553, 2579, 2583, 2595, 2600, 2619, 2634, 2647, 2656, 2664, 2667, 2706, 2725, 2770, 2853, 2892, 2914]. based [51, 61, 70, 93, 106, 113, 142, 209, 214, 253, 271, 316, 330, 394, 435, 450, 471, 532, 533, 552, 556, 568, 607, 618, 739, 745, 781, 782, 811, 829, 845, 846, 872, 879, 911, 920, 933, 952, 1121, 1181, 1188, 1194, 1242, 1278, 1292, 1298, 1321, 1332, 1358, 1436, 1455, 1483, 1537, 1611, 1652, 1671, 1695, 1757, 1775, 1824, 1905, 1913, 1917, 1937, 1969, 2024, 2034, 2068, 2094, 2111, 2174, 2199, 2201, 2230, 2252, 2276, 2277, 2312, 2356, 2429, 2531, 2533, 2554, 2555, 2592, 2593, 2623, 2629, 2659, 2670, 2729, 2739, 2741, 2755, 2758, 2793, 2803, 2814, 2826, 2833, 2911]. based [11, 39, 43, 45, 53, 55, 102, 164, 180, 183, 229, 260, 300, 353, 454, 463, 466, 513, 530, 539, 572, 608, 634, 669, 687, 919, 927, 978, 1247, 1256, 1266, 1289, 1299, 1310, 1316, 1359, 1366, 1415, 1428, 1448, 1461, 1475, 1477, 1518, 1551, 1576, 1596, 1597, 1601, 1612, 1674, 1678, 1715, 1721, 1732, 1759, 1779, 1793, 1803, 1866, 1875, 1883, 1946, 1959, 1995, 2006, 2029, 2102, 2103, 2114, 2140, 2146, 2154, 2207, 2232, 2249, 2263, 2266, 2308, 2313, 2329, 2352, 2355, 2372, 2378, 2385, 2417, 2421, 2422, 2433, 2461, 2472, 2484, 2499, 2510, 2529, 2543, 2575, 2580, 2622, 2631, 2668, 2686, 2705, 2716, 2720, 2748, 2750, 2818, 2820, 2823, 2840, 2880, 2881]. based [38, 41, 75, 133, 254, 277, 292, 438, 495, 676, 717, 850, 857, 1019, 1085, 1171, 1172, 1180, 1279, 1342, 1411, 1446, 1488, 1638, 1693, 1767, 1778, 1782, 1792, 1801, 1810, 1812, 1819, 1847, 1925, 1938, 2008, 2160, 2301, 2305, 2334, 2375, 2382, 2400, 2465, 2471, 2512, 2517, 2535, 2625, 2640, 2666, 2679, 2724, 2743, 2809, 2831, 2868, 2900, 2912]. bases [1603]. basic [886]. basis [667, 2819, 2877]. Basketball [1081, 1514, 1742, 2696]. bat [1492, 1511, 1729, 1908]. batch [18, 635, 2122]. batched [1134, 1547, 2280]. batches [2316]. batteries [2367, 2734]. battery [1719, 1810, 2838]. Battle [2754, 2781]. battlefield [1999]. Bayes [725, 1553, 2031]. Bayesian [55, 384, 671, 863, 1048, 1135, 1607, 1940, 2387, 2577]. Bayesian-based [1940]. BB [715]. BB-tree [715]. BC [1673]. BCGAN [920]. BCN [1829]. beam [1462]. beamforming [2407]. Bearer [307]. bearing [2034]. Beck [866]. bee [245, 372, 478]. Beetle [1898, 2050, 2166]. Behavior [156, 570, 646, 669, 815, 887, 1058, 1249, 1257, 1455, 1628, 1629, 1632, 2009, 2201, 2259, 2318, 2358, 2442, 2479]. behavior-based [2358]. behavioral [1555, 2637]. behaviors [165, 788, 976, 1528, 1728, 2536]. behaviour [2222]. BejaGNN [2358]. belief [684, 1092, 1387, 1679, 2376, 2577]. benchmark [2144]. benchmarking [378, 615, 1280, 1970]. Benchmarks [2110]. benefits [833]. BER [218]. BeRAN [2398]. BERT [1292, 1977, 2214, 2261, 2426]. BERT-CNN [1292]. Best [33, 491, 500, 964, 1177, 1246, 2291]. best-first [1177]. best-fit [500]. best-fit-decreasing [33]. Best-tree [2291]. better [1184, 1185, 1418]. between [509, 720, 1648, 2188, 2240, 2319, 2336, 2387, 2415]. betweenness [1983]. bi [941, 1160, 1223, 1520, 1657, 1823, 1853, 2386, 2438]. bi-cluster [1160]. bi-layer [1853]. bi-level [2386, 2438]. Bi-LSTM [941, 1520, 1657]. bi-objective [1223, 1823]. bias [1713]. bibliographic [200]. bibliometric [777]. bibliometrics [169]. BID [2898]. Bidirectional [550, 593, 797, 1108, 1932, 2091, 2291, 2678].

Big [13, 56, 77, 78, 143, 169, 170, 201, 217, 221, 261, 370, 450, 571, 588, 589, 617, 751, 762, 773, 793, 800, 915, 942, 951, 1016, 1043, 1052, 1156, 1165, 1181, 1217, 1218, 1224, 1400, 1717, 1988, 2018, 2039, 2136, 2200, 2508, 2669, 2824, 2845, 2911, 2917]. big-data [793]. BiGAN [1872]. bigdata [285]. BiGRU [2668]. bike [2711]. bike-sharing [2711]. bilateral [2693, 2827]. bilinear [2297]. bilingual [73]. BiLSTM [2214, 2494, 2673, 2776]. bin [1821]. binarization [842]. Binary [156, 225, 377, 552, 739, 867, 917, 1073, 1121, 1201, 1303, 1407, 1488, 1635, 1908, 1984, 2034, 2035, 2151, 2155, 2208, 2265, 2281, 2544]. Binary-PSO-based [1121]. binding [1663]. bio [104, 948, 957, 1653, 2902]. bio-image [957]. bio-inspired [948, 1653, 2902]. biochemical [315]. biogas [2659]. biogeography [568, 1819, 2144]. biogeography-based [568, 1819, 2144]. biohashing [2805]. bioimaging [1192]. bioinformatic [990]. bioinspired [2702]. biological [186, 398, 536, 938, 1377, 1434, 1781]. biology [309, 1464]. biomedical [1635, 1668, 1672, 1772, 2296]. biometric [63, 224, 453, 1525, 2245]. biometric-based [453]. biotechnology [609]. bipartite [753, 1406]. bird [2790]. birds [625]. bit [138, 704, 712, 1726, 1760, 2448, 2627]. bit-level [138]. Bitcoin [2324]. bitmap [677]. bits [827]. bitwise [3, 276]. bivariate [2728]. Black [397, 489, 510]. blackheart [193]. blacklist [1601]. BLAST [590]. BLASTP [534]. blendshapes [2010]. blind [86]. BLIS [128]. Block [448, 512, 889, 964, 1285, 1364, 1441, 1545, 1609, 2312, 2560]. block-centric [2560]. Blockchain [267, 532, 651, 665, 822, 864, 878, 879, 881, 927, 980, 999, 1004, 1043, 1068, 1090, 1237, 1301, 1332, 1356, 1366, 1375, 1444, 1513, 1601, 1645, 1659, 1676, 1687, 1701, 1702, 1771, 1801, 1865, 1875, 1882, 1915, 1916, 2076, 2190, 2235, 2308, 2312, 2334, 2336, 2369, 2378, 2390, 2410, 2449, 2510, 2518, 2523, 2526, 2600, 2605, 2612, 2625, 2635, 2646, 2664, 2703, 2706, 2739, 2747, 2778, 2786, 2804, 2809, 2834, 2844, 2848, 2862, 2869, 2904, 2914]. Blockchain-assisted [2518]. Blockchain-based [651, 864, 878, 1068, 1301, 1332, 1356, 1601, 1645, 1659, 1676, 1701, 1702, 1771, 2190, 2235, 2312, 2510, 2612, 2646, 2664, 2809, 2834, 2914]. Blockchain-enabled [2369, 2390, 2844]. blockchains [1412, 2336, 2639, 2766]. blocking [696, 1494, 1587]. Blockly [848]. blocks [393, 2109]. blog [242]. blood [2582]. bloom [1950]. Bluelight [2391]. BOA [18]. body [342, 350, 606, 1274, 1318, 1434, 2238, 2609]. Boltzmann [982, 2284]. books [768]. Boolean [1937, 2532]. boost [334, 1807, 1885]. boosted [366]. boosting [1858, 1893, 1977, 2288]. booting [2866]. bootstrapped [2077]. bootstrapping [2919]. border [1407, 2151]. borne [1071]. Borwein [945, 2596]. boson [2178]. both [626, 2692]. botnet [499]. bound [2078, 2236, 2437]. boundary [633, 880, 920]. bounds [516, 683, 1806]. BOW [353]. bowerbird [1943]. box [24, 510, 1457]. box-counting [24]. boxes [2627]. BP [65, 113, 673, 803, 882, 1443]. BPNN [2893]. brain [508, 711, 807, 854, 1358, 1523, 1566, 1713, 1808, 2004, 2051, 2125, 2617, 2654]. branch [1241, 2102, 2236, 2462, 2593]. BRB [2078]. breadth [2029]. breadth-first [2029]. break [1206]. break-aware [1206]. breast [444, 559, 2705, 2727, 2828, 2884]. bridge [67, 891]. broad [2121, 2540]. broadcast [577, 1083, 1796]. broadcasting [732, 1920]. broker [2883]. brokering [12]. brown [115]. brown-shell [115]. Brute [2680]. Brute-force [2680]. BSP [299]. BSR [575]. BTDA [2491]. BTLBO [2872]. bubble [1813, 2338]. bubble-sort [2338]. budget [34, 1070]. budget-aware [1070]. budget-constrained [34]. buffer [228, 1290, 2288, 2375]. buffers [1542]. build [1079]. Building [49, 200, 607, 1017, 1025, 1455, 1615, 1917, 1967, 2037, 2455]. buildings [185, 767]. bundle [1133]. burn [2621]. burst [953, 2760]. Burstiness [30]. Burstiness-aware [30]. bursty [2879]. bus [1365]. business [178, 182, 189, 580, 1838, 2085, 2128, 2741]. butler [2009]. Butterfly [1201, 2342, 2421, 2780]. BW [1016]. bypassing [540, 1973, 2036].

C [298, 361, 1930]. C-DMR [1930]. C3D [1758, 1811]. CA [462, 1701]. CA-Dedupe [462]. cable [2339, 2562]. CAC [781, 1390]. cache [393, 608, 1189, 1195, 1234, 1275, 1930, 1973, 2036, 2164]. cache-based [1930]. caches [974]. Caching [395, 916, 1247, 2350, 2419, 2435]. Caching-based [1247]. calculating [1648]. Calculation [86, 1465, 1948, 2087]. calculator [2904]. call [1249]. calls [1323]. calorie [832]. CALYOLOv4 [2484]. CaM [1016]. CaM-BW [1016]. CAM5 [2444]. camera [868, 1060, 1669]. CAMID [1788]. CAMIRA [1489]. Camshift [1172]. can [1365, 1418]. cancellation [16, 1188]. cancer [329, 338, 444, 559, 684, 818, 877, 1465, 1507, 2642, 2705, 2727, 2737, 2828, 2884]. candidate [2712]. candidates [2072]. Canny [1476]. capabilities [990, 1232, 1730]. capability [122, 730, 1441, 2382]. capacities [1205]. capacities-based [1205]. capacitive [459]. capacity [487, 681, 736, 881, 935, 1540]. capacity-aware [1540]. capsized [2888]. capsule [1964, 2719, 2899]. capture [161, 1553, 2010]. capturing [367]. capuchin [2636, 2843]. Car [927, 1734, 1894, 2450]. car-hailing [1894]. Carbon [479, 2594]. carcinoma [37, 1210, 1461]. card [1096, 1595]. cardiac [344, 351, 1560]. cardiotocographic [52]. cardiovascular [912]. care [144, 213, 1274, 1386, 1967, 2037]. Carlo [320]. Carmel [954]. carotid [912]. carrier [1074]. carry [1544, 2630]. carry-save [2630]. cartesian [1469, 2633]. cascade [1847, 2648, 2722, 2740]. cascaded [2827]. case [13, 14, 155, 328, 441, 524, 609, 694, 1037, 1331, 1334, 1716, 1838, 2009, 2043, 2124, 2144, 2203, 2319, 2343]. case-based [14, 694]. Cassandra [590]. cassava [2591]. Cast [203]. casting [1950]. cat [221, 1846, 2531]. Catalyst [1196]. catastrophic [2365]. categorical [229]. categories [2828]. causal [206, 1292, 2094]. cause [1233]. caused [154]. CAVLC [1327]. CAVLCU [1327]. Cayley [560, 1931]. CBC [2504]. CBVoSD [1289]. CC [1705, 2072]. CC-IFIM [2072]. CC-RRTMG SW [1705]. CCECGP [2094]. CCRN [1943]. CDA [1690]. CDMA [16]. Cell [22, 186, 416, 869, 899, 1121, 1155, 1310, 1586, 1755, 2040, 2521, 2582, 2773, 2816]. cell-based [22, 869]. cellphone [2683]. cells [446, 1033]. Cellular [96, 127, 585, 747, 790, 862, 865, 869, 900, 950, 952, 1111, 1112, 1131, 1141, 1157, 1249, 1303, 1335, 1620, 1718, 1725, 1726, 1738, 1839, 1840, 1879, 1919, 1950, 2124, 2217, 2630, 2816]. cellular-level [1141]. CEMAR [1117]. center [35, 354, 394, 469, 600, 1176, 1546, 1643, 1797, 1918, 2152, 2392, 2509]. centers [82, 129, 130, 382, 514, 618, 648, 698, 798, 799, 862, 976, 993, 1028, 1030, 1051, 1093, 1382, 1589, 1821, 1863, 2432, 2455, 2578, 2606, 2619]. central [869, 2293]. centralized [1117, 2023, 2169]. centre [155, 933]. centres [507, 2087, 2800]. centric [936, 999, 1026, 1130, 1375, 2152, 2480, 2560]. Centroid [668]. Centroid-Based [668]. CEP [846]. Ceph [321]. cerberus [1896, 1901]. CERN [1065, 1066]. certain [1541, 2234]. certificate [2252]. certificate-based [2252]. Certificateless [828, 1862, 1954, 2412]. cervical [446, 684, 1265]. CF [2699]. CF-lines [2699]. CFD [126, 183, 2060, 2775]. CFF [2688]. CFIN [605]. CfoTS [336]. CGAN [920]. CGAN-based [920]. CGM [1256].

CGM-based [1256]. ChaCha20 [1221]. chain [177, 194, 572, 1373, 1771, 1911, 1953, 2100, 2176, 2239, 2491, 2639, 2739, 2805, 2834]. chaining [983, 2216]. chains [389, 889]. challenge [705]. challenge-response [705]. Challenges [6, 129, 130, 385, 486, 744, 1053, 1124, 1295, 1306, 2514, 2547, 2632, 2650, 2901]. Challenging [302, 1353, 1567]. change [175, 2192, 2874]. changes [1434]. changing [1264]. channel [71, 202, 562, 1010, 1152, 1188, 1551, 1602, 1827, 1899, 2033, 2130, 2307, 2321, 2386, 2600, 2723, 2837]. channel-wise [1551]. channels [1083, 2081]. chaos [547, 2512]. chaos-based [2512]. Chaotic [454, 828, 836, 867, 917, 986, 1240, 1452, 1525, 1729, 1866, 2015, 2412, 2434, 2470, 2579, 2581, 2658, 2805, 2819]. characterisation [1293]. Characteristic [2468, 2851]. characteristics [127, 964, 1452, 1492, 2391, 2874]. Characterization [52, 1731, 1869]. Characterizing [188, 1380]. charge [1575, 2367]. chart [1579]. Chebyshev [743, 828, 1240, 1525, 2412, 2537]. check [1195]. checking [422, 1535, 1806, 2119, 2459, 2482]. checkpoint [701, 735, 1628, 1629]. checkpoint/restart [735]. chemistry [1453]. chemoembolization [919, 1465]. chemotherapy [684]. CHESDA [719]. chest [2570]. Cheval [793]. chi [2168]. chi-square [2168]. chicken [1018, 1809]. child [1427, 1441]. children [1463, 2860]. chimp [2015, 2698]. China [2293, 2871]. Chinese [73, 104, 187, 324, 470, 1670]. chip [501, 696, 861, 1056, 1296, 1450, 1857, 2016, 2599, 2794, 2910]. Chipping [2833]. chips [541, 1626, 2187]. Cholesky [2541]. Chord [2766, 2784]. Chord-based [2784]. chromatic [641]. Chronic [2421]. churn [241, 375, 1202, 2028, 2173, 2351]. churn-resilient [375]. CICAPS [2001]. CICCM [2783]. cipher [1221, 1482]. ciphers [964]. ciphertext [2805]. circuit [1384, 1910, 2257, 2286, 2879]. circuit-based [1384]. circuits [692, 950, 1042, 1073, 1522, 1947, 2217, 2282]. circulant [1552]. circular [1981, 2798]. circumstances [1976]. CISTs [560]. cities [453, 984, 1468, 1669, 2411, 2573, 2672]. city [59, 162, 638, 1199, 1209, 1756]. CKHO [1608]. Clairvoyant [759]. class [17, 428, 429, 787, 819, 920, 1429, 1559, 1748, 1886, 2018, 2182, 2405, 2545, 2704]. class-specific [428, 429]. classes [2154]. classical [258, 259, 2322, 2538, 2817]. Classification [40, 47, 138, 152, 157, 217, 224, 242, 252, 272, 286, 340, 381, 385, 424, 434, 444, 446, 449, 460, 520, 551, 622, 630, 667, 733, 739, 743, 748, 797, 800, 807, 826, 855, 870, 877, 918, 923, 1009, 1154, 1215, 1257, 1295, 1419, 1478, 1537, 1539, 1560, 1566, 1579, 1594, 1606, 1630, 1672, 1734, 1775, 1820, 1852, 1892, 1940, 1972, 1988, 2018, 2033, 2038, 2039, 2054, 2058, 2079, 2099, 2136, 2142, 2198, 2205, 2222, 2227, 2230, 2250, 2363, 2421, 2435, 2456, 2537, 2582, 2584, 2591, 2642, 2652, 2654, 2656, 2661, 2702, 2717, 2719, 2737, 2762, 2790, 2824, 2899, 2901, 2905, 2917]. classification [57, 58, 1289, 1292, 1310, 1436, 1523, 1586, 1899, 2031, 2154, 2174, 2223, 2265, 2479, 2635, 2668, 2869]. classifications [1025, 2842]. classifier [338, 341, 352, 595, 725, 858, 1442, 1586]. classifiers [359, 645, 1537, 1787]. classifying [559, 2119, 2138]. clause [1233]. CLGR [1975]. CLGR-Net [1975]. click [2385, 2688]. click-through [2688]. client [41, 836, 2299, 2388, 2883]. client-server [2388]. clients [2387]. Clifford [2257]. Climate [2140, 2874]. climbing [337, 2180]. Clinical [355, 725, 930, 1465, 2261, 2847]. clique [1751, 2078]. clock [85, 530, 1634, 2089]. clonal [2499]. CloneCloud [726]. closed [926, 1911, 2072]. closed-loop [1911]. closest [739]. closures [422]. Cloud [4, 12, 81, 108, 208, 227, 262, 280, 287– 289, 309, 319, 380, 382, 420, 431, 474, 507, 515, 523, 549, 571, 580, 584, 610, 637, 651, 653, 742, 816, 823, 835, 889, 928, 932, 966, 969, 972, 1030, 1043, 1095, 1096, 1106, 1117, 1136, 1176, 1182, 1187, 1205, 1217, 1218, 1224, 1306, 1326, 1338, 1382, 1385, 1392, 1413, 1414, 1420, 1444, 1532, 1688, 1690, 1697, 1741, 1773, 1776, 1823, 1855, 1859, 1880, 1889, 1895, 1951, 1989, 2021, 2053, 2085, 2128, 2135, 2347, 2395, 2438, 2444, 2493, 2532, 2566, 2578, 2613, 2628, 2715, 2730, 2746, 2780, 2887, 2893]. cloud [6, 23, 30, 35, 82, 273, 285, 304, 305, 308, 369, 371, 383, 386, 389, 399, 440, 469, 486, 542, 568, 578, 589, 598, 617, 655, 698, 715, 734, 749, 750, 769, 779, 798, 799, 849, 856, 875, 916, 917, 922, 993, 996, 1021, 1028, 1051, 1077, 1080, 1125, 1140, 1163, 1181, 1190, 1197, 1199, 1227, 1238, 1242, 1276, 1332, 1345, 1346, 1388, 1426, 1499, 1535, 1581, 1608, 1631, 1655, 1663, 1707, 1731, 1789, 1809, 1821, 1841, 1884, 1992, 2014, 2063, 2069, 2094, 2107, 2157, 2206, 2237, 2241, 2314, 2440, 2480, 2497, 2553, 2595, 2606, 2619, 2667, 2689, 2841, 2843, 2901]. cloud [102, 134, 148, 265, 392, 656, 933, 946, 1159, 1321, 1366, 1389, 1416, 1446, 1584, 1732, 1835, 1863, 1875, 1954, 2103, 2360, 2392, 2764, 2822]. cloud-assisted [280, 1875, 2764]. cloud-based [102, 420, 440, 1163, 1181, 1187, 1217, 1218, 1889, 2157]. cloud-centric [2480]. cloud-edge [916]. cloud-fog [2085, 2128]. cloud-fog-IoT [2360]. cloud-healthcare [1655]. CloudBench [373]. cloudIoT [1698]. cloudlet [2472]. cloudlets [1095, 1340]. clouds [2, 339, 373, 652, 678, 1012, 1777, 2057, 2166]. clubbing [331]. Cluster [18, 70, 155, 207, 215, 229, 261, 362, 488, 756, 771, 826, 923, 1138, 1160, 1268, 1280, 1306, 1460, 1469, 1524, 1529, 1558, 2007, 2170, 2172, 2243, 2285, 2340, 2466, 2486, 2527, 2528, 2876]. cluster-based [70, 923, 1460, 1558, 2007, 2486]. cluster-tree-based [756]. clusterCL [504]. clustered [1485]. Clustering [53, 67, 95, 131, 157, 180, 216, 229, 242, 257, 269, 286, 325, 331, 343, 391, 398, 492, 493, 505, 549, 586, 681, 682, 693, 713, 728, 733, 773, 786, 825, 857, 908, 929, 1012, 1100, 1250, 1341, 1377, 1381, 1400, 1453, 1454, 1563, 1582, 1596, 1656, 1674, 1772, 1885, 1889, 1904, 1922, 1997, 2040, 2104, 2329, 2396, 2397, 2485, 2549, 2561, 2592, 2636, 2643, 2669, 2681, 2697, 2751, 2787, 2870, 2871, 2880]. Clustering-based [53, 492, 586, 1341, 2561, 2681]. clusters [26, 83, 95, 389, 456, 502, 504, 563, 582, 609, 644, 902, 1222, 1377, 1430, 1489, 1625, 1785, 2406, 2487, 2748, 2820, 2854, 2906]. cm [1465]. CMO [1372]. CMO-COOT [1372]. CMODLB [856]. CMOS [1040]. CMP [231]. CNN [470, 661, 941, 1044, 1046, 1292, 1308, 1344, 1423, 1426, 1647, 1694, 1895, 1926, 1935, 2175, 2244, 2321, 2386, 2403, 2433, 2450, 2494, 2584, 2668, 2673]. CNN-based [2668]. CNN-LSTM [1426]. CNN-RDM [1926]. CNN-VAE [2244]. CNN’s [686, 1180, 2828]. CNN’s-based [2828]. CNTFET [899, 2701]. CNTFET-based [899]. Co [172, 990, 1120, 1174, 2316, 2452]. co-design [2452]. Co-designing [990]. co-occurrence [172, 1120]. co-scheduling [2316]. co-training [1174]. coaching [1254]. coal [1572]. coalescing [1795]. coalition [1415, 2409]. CoAP [243]. CoAP-based [243]. coarse [126, 1109, 2587]. coarse- [2587]. coarse-grained [126]. Code [226, 282, 292, 300, 307, 648, 810, 950, 1015, 1055, 1331, 1417, 1681, 1884, 2219]. code-based [292]. code-reuse [810]. coded [218]. codes [89, 235, 564, 2261]. CoDFi [2615]. CoDFi-DL [2615]. coding [16, 42, 781, 1171, 1320, 1862, 2083, 2221, 2431]. coefficient [1076]. coefficients [1648]. coevolutionary [2353]. coexistence [2879]. CoFB [2316]. Cognitive [16, 17, 36, 44, 49, 314, 330, 583, 1527, 1686, 2417, 2564]. coherence [1610, 2671]. cohesion [439]. cohorts [778]. cold [1783, 2240]. cold-start [1783]. collaboration [153, 1445].

Collaborative [4, 198, 260, 340, 457, 531, 787, 975, 1022, 1100, 1300, 1975, 1992, 2094, 2287, 2463, 2615, 2736, 2766]. collaborative-based [4, 1992]. Collatz [624]. collected [1007]. collection [285, 379, 627, 1387, 1403, 1928]. collective [64, 68, 999, 2451]. college [1271, 1724, 2400]. collision [1003, 2001]. collusion [2834]. collusive [388]. COLMA [2512]. Colony [43, 231, 245, 362, 372, 478, 593, 782, 1048, 1294, 1381, 2055, 2823, 2843]. Color [203, 454, 673, 803, 1441, 1826, 1842, 2648, 2722, 2724]. Color-patterned [2724]. colorectal [877]. coloring [698, 1796, 2011]. column [1089]. column-wise [1089]. combat [1597]. combating [1746]. combination [306, 1164, 1191, 2329, 2696]. Combinatorial [339, 955, 1518]. Combined [133, 722, 1044, 1210, 1476, 1556, 1630, 1825, 2267, 2289, 2483, 2562]. combining [236, 1791, 2615, 2668, 2688, 2897]. command [1621, 1983]. command-and-control [1983]. commander [1999]. commensal [938]. Comment [279]. comments [2919]. commerce [2024]. commercial [2190]. committee [1916]. commodity [473]. Commonsense [2873]. communicating [382]. Communication [80, 123, 233, 248, 263, 307, 364, 420, 442, 475, 565, 720, 794, 829, 868, 1123, 1138, 1169, 1203, 1264, 1290, 1352, 1739, 1793, 1796, 1813, 1990, 2028, 2052, 2060, 2080, 2173, 2189, 2301, 2330, 2351, 2388, 2409, 2451, 2593, 2640, 2848, 2881, 2891, 2914]. communication-aware [475]. Communication-efficient [2301]. communication-intensive [364]. communications [732, 2791, 2875]. communities [753]. Community [25, 153, 403, 448, 605, 755, 938, 975, 1123, 1404, 1424, 1890, 2559, 2649, 2856]. community-based [605]. CoMoDa [1390]. CoMP [416]. compact [2428]. compaction [906]. companion [1561]. Comparative [146, 434, 655, 804, 805, 965, 1165, 1205, 1319, 1532, 1543, 1844, 1872, 2227, 2611]. comparator [704, 1726]. comparators [842]. Comparing [2187]. comparison [620, 645, 765, 1145, 1178, 1548, 1706, 2587]. compatibility [160, 1010]. compatible [72]. compensation [659, 949, 1826, 2267]. competing [1584]. competitive [319, 1711, 2258, 2698]. Compilation [361, 1094, 2178]. Compiler [1069, 2428]. Compiler-directed [1069]. complete [2208, 2633]. completely [1993]. completeness [641]. completion [1533]. Complex [251, 605, 707, 758, 915, 1118, 1281, 1405, 1578, 1890, 1943, 2138, 2303, 2366, 2644, 2810]. complex-valued [1943]. Complexity [52, 103, 155, 326, 327, 990, 1417, 1802]. compliance [2741]. complications [919]. Component [281, 945, 1242, 1341, 1440, 2329, 2516, 2585]. component-based [1242, 1440]. Component-wise [2329]. components [666]. composite [1197, 2107]. Composition [273, 758, 1142, 1145, 1549, 2017, 2061, 2183, 2435, 2872]. compounds [852]. Comprehensive [116, 389, 504, 823, 972, 1159, 1295, 1576, 2584, 2698, 2707, 2720]. compressible [940, 2525]. compressing [1791]. compression [72, 145, 323, 409, 434, 904, 945, 1171, 1285, 1556, 1671, 1860, 1925, 1927, 2109, 2116, 2199, 2211, 2289, 2414, 2429, 2623, 2755, 2798, 2865]. compression-based [2199]. compression-encryption [2109]. compressive [2109]. Compressor [724]. compressors [1743]. Computation [14, 158, 452, 508, 538, 678, 699, 714, 769, 800, 875, 915, 1008, 1023, 1047, 1091, 1102, 1122, 1144, 1175, 1385, 1446, 1512, 1622, 1822, 2039, 2060, 2076, 2118, 2199, 2264, 2289, 2312, 2323, 2419, 2477, 2524, 2597]. computational [273, 309, 601, 1389, 1396, 1400, 1519, 1616, 1778, 1945, 2917]. computations [522, 766, 1094, 1833, 2132, 2817]. compute [721, 1290]. computed [622, 1462, 1463, 1465]. Computer [46, 153, 179, 253, 316, 958, 1175, 1322, 1402, 1611, 1733, 1934, 1946, 2126, 2370, 2586, 2670, 2776, 2826]. Computer-aided [1402, 2670, 2776]. computerized [820]. computers [827, 2151]. Computing [4, 56, 62, 159, 185, 227, 262, 263, 280, 319, 371, 395, 398, 553, 579, 584, 589, 628, 637, 648, 675, 714, 718, 727, 752, 793, 825, 835, 848, 866, 924, 966, 972, 976, 994, 1026, 1041, 1043, 1047, 1065, 1066, 1071, 1095, 1096, 1106, 1107, 1122, 1144, 1172, 1182, 1183, 1207, 1224, 1283, 1306, 1328, 1339, 1340, 1385, 1392, 1395, 1414, 1423, 1439, 1498, 1519, 1532, 1563, 1633, 1676, 1683, 1699, 1753, 1776, 1778, 1825, 1880, 1895, 1909, 1951, 1992, 2041, 2086, 2116, 2135, 2175, 2178, 2240, 2409, 2423, 2435, 2493, 2532, 2628, 2641, 2692, 2703, 2797, 2811, 2886]. computing [15, 30, 36, 60, 68, 134, 158, 212, 282, 313, 383, 386, 389, 486, 487, 552, 568, 594, 608, 617, 644, 661, 697, 706, 734, 749, 750, 769, 779, 838, 856, 875, 893, 911, 917, 990, 1028, 1077, 1103, 1125, 1190, 1199, 1238, 1242, 1247, 1298, 1337, 1345, 1386, 1398, 1405, 1426, 1447, 1458, 1489, 1499, 1500, 1550, 1569, 1570, 1590, 1592, 1616– 1618, 1675, 1679, 1744, 1760, 1788, 1802, 1841, 1878, 1927, 1995, 2003, 2056, 2062, 2076, 2112, 2162, 2231, 2289, 2341, 2349, 2378, 2392, 2417, 2470, 2497, 2527, 2547, 2553, 2597, 2650, 2667, 2793, 2843, 2853, 2901, 2903, 2913]. computing [656, 1155, 1223, 1389, 1477, 1512, 1584, 1712, 2207, 2400, 2625, 2716, 2809]. computing-assisted [158]. computing-based [60, 159, 1144, 1477]. computing-enabled [2076, 2809]. concatenate [1846]. concave [2004]. Concentration [1632, 1921, 2633]. concept [324, 1202, 1276, 1962, 2842]. concept-sensitive [1202]. concepts [20, 1956]. conceptual [2310]. Concise [204]. concurrency [982]. concurrent [1078, 1435]. condition [789, 1170, 2295, 2720]. conditional [1029, 1060, 1837, 2344, 2498]. conditioned [871, 2297]. conditions [426, 1945, 2186, 2377]. cone [1462]. cone-beam [1462]. confidence [2584, 2617]. Confidentiality [304]. Confidentiality-preserving [304]. configurable [1348, 1919]. configuration [701, 847]. configurations [2500]. confined [852]. confirmation [2628]. conflict [770]. conformable [2658]. Conformance [1535, 2459]. confrontation [2017, 2651]. congenital [1214]. Congestion [683, 777, 1097, 1275, 1485, 1500, 1654, 1813, 2070, 2561, 2682, 2694]. congestion-aware [1097]. congestion-guided [2682]. conjugate [358]. connected [643, 1075, 1128, 1936, 2381, 2787, 2826]. connection [1004]. Connectivity [1264, 1299, 1632, 1644, 1673, 1887, 1931, 1986, 2098, 2295, 2368, 2585, 2601, 2633, 2704]. conquer [296, 1430, 2502]. conscious [333]. Consensus [849, 1133, 1237, 1645, 1687, 1903, 1913, 1916, 2523]. Consensus-based [849, 1133]. conservation [1859]. Considerations [987]. considered [105]. considering [4, 119, 135, 177, 425, 479, 636, 1027, 1058, 1232, 1333, 1512, 1540, 1571, 1992, 2192, 2325, 2411, 2460, 2573, 2672]. consistency [369, 422, 1623, 2094, 2569]. consolidating [1467]. consolidation [29, 82, 129, 130, 469, 503, 514, 520, 604, 749, 972, 1489, 1515, 1757, 2497]. consolidation-aware [1489]. consortium [2100, 2523]. constrained [34, 212, 482, 816, 986, 1021, 1227, 1415, 1457, 1690, 1693, 1696, 1984, 2057, 2316, 2789, 2841]. Constraint [583, 785, 1688, 2045]. constraints [525, 1378, 1549, 1791, 1922, 2379, 2559]. construct [1092, 2577]. constructed [2376]. Constructing [560, 883, 2344, 2640, 2878]. Construction [69, 321, 603, 673, 803, 1079, 1271, 1284, 1398, 1432, 1453, 1454, 1756, 1943, 2272, 2471, 2610, 2645, 2891]. construction-based [1943]. consumer [1959]. consumers [2259]. consumption [87, 184, 208, 557, 614, 746, 1530, 1534, 1609, 1959, 1989, 2197, 2259, 2269, 2300, 2569, 2609, 2819]. container [263, 689, 765, 1789, 1909, 2215, 2578, 2825]. Containerization [1124]. Containerized [698, 1012, 1302, 1809, 2166]. Containers [380, 1012, 1467, 1985]. containers-based [380]. content [25, 193, 271, 462, 666, 1036, 1116, 1435, 1697, 1853, 2119, 2427, 2770, 2821]. content-addressable [2427]. content-aware [462]. content-based [271, 1853]. contention [778, 1333, 1600, 1824]. contents [1251, 2911]. context [88, 159, 178, 406, 480, 936, 1009, 1145, 1289, 1338, 1586, 1849, 2205, 2296, 2657, 2725, 2744, 2753, 2784, 2892]. context-aware [88, 159, 406, 936, 1009, 2725]. context-driven [1849]. contextualized [1419]. contiguous [2522]. Continuous [175, 206, 404, 719, 1212, 1725, 1950, 2543, 2559, 2622, 2907]. contour [112, 1647, 2699]. contract [290, 980, 1090, 1375, 2076, 2384]. contract-based [290]. contract-guided [2076]. Contracts [298, 2287, 2625]. Contrast [754, 1194, 2371]. Contrast-enhanced [1194]. Contrastive [2555, 2757, 2759, 2912]. control [136, 174, 305, 337, 551, 554, 697, 764, 794, 852, 887, 1117, 1143, 1178, 1239, 1296, 1325, 1332, 1374, 1416, 1444, 1451, 1459, 1485, 1515, 1571, 1620, 1621, 1631, 1703, 1714, 1770, 1771, 1813, 1901, 1983, 2023, 2180, 2237, 2334, 2341, 2380, 2424, 2580, 2608, 2625, 2694, 2846]. control-based [2694]. controlled [1468]. Controller [474, 806, 854, 932, 992, 1032, 1050, 1269, 1358, 1421, 1540, 1845, 1971, 2192, 2300, 2375, 2748, 2802, 2820]. controller-based [992]. controllers [1008, 2866, 2902]. controls [1365]. Conv [1520]. Conv-LSTM [1520]. conventional [1282]. converged [265]. Convergence [264, 274, 624, 715, 2481, 2584, 2626]. conversational [2436]. conversion [215, 1814, 2533]. converter [806, 950]. converters [1073]. ConvNeXt [2854]. ConvNeXt-T [2854]. convolution [157, 663, 1297, 1354, 1525, 1585, 1868, 2027, 2147, 2507, 2822]. convolution-Chebyshev [1525]. Convolutional [164, 345, 415, 419, 444, 449, 461, 559, 635, 855, 1002, 1006, 1025, 1108, 1201, 1204, 1220, 1244, 1315, 1419, 1434, 1493, 1546, 1579, 1606, 1670, 1713, 1716, 1722, 1752, 1767, 1803, 1831, 1852, 1899, 2136, 2142, 2156, 2198, 2260, 2281, 2394, 2433, 2460, 2552, 2563, 2582, 2591, 2642, 2724, 2726, 2753, 2776, 2786, 2819, 2821]. convolutions [1128, 2181]. convolver [1167]. convolvers [2483]. cooling [1643, 1941, 2509]. cooperating [1176]. cooperation [2651]. Cooperative [44, 161, 314, 340, 395, 530, 1431, 1584, 1686, 1999, 2001, 2288, 2289, 2331, 2353, 2495, 2564, 2664, 2891]. Coordinate [307]. coordinated [2484, 2557]. coordination [931, 1748, 1886, 2488, 2562]. COOT [1372]. coplanar [821, 1839, 2448, 2816]. COPRAS [2356]. Cops [886]. copy [2451]. Cor [2099]. Cor-ENTC [2099]. CORD [2393]. CORD-19 [2393]. core [8, 85, 576, 592, 650, 699, 723, 776, 880, 907, 1069, 1177, 1236, 1380, 1394, 1430, 1445, 1602, 1854, 2064, 2180, 2181, 2199, 2213, 2283, 2415, 2507, 2876]. Cores [1330, 1333, 1769, 2122]. corner [1506]. Corona [1273]. coronary [1463]. corporate [405]. corpus [496, 2271]. Correcting [1417, 1818, 2627]. Correction [3, 42, 130, 140, 247, 259, 336, 392, 429, 799, 805, 914, 932, 963, 998, 1066, 1112, 1185, 1218, 1229, 1305, 1312, 1370, 1479, 1491, 1509, 1531, 1567, 1629, 1674, 1713, 1766, 1810–1812, 1942, 1986, 1987, 2036, 2037, 2083, 2084, 2128, 2173– 2175, 2351, 2352, 2439–2442, 2573, 2622, 2623, 2672, 2673, 2722, 2771, 2820, 2869, 2870, 2896]. Correlation [749, 900, 2099, 2469, 2682, 2915]. Correlation-based [2915]. correlations [326]. correlative [820]. CoSC [2287]. cosine [2270]. cosmological [1318]. Cost [97, 108, 233, 267, 309, 354, 410, 497, 506, 519, 541, 578, 580, 729, 781, 798, 799, 831, 1080, 1095, 1213, 1256, 1260, 1690, 1738, 1835, 1909, 1910, 1996, 2014, 2041, 2057, 2077, 2166, 2240, 2325, 2405, 2563, 2816]. Cost-aware [354, 541, 1690, 1996]. Cost-effective [1080, 1095, 1910, 2057]. Cost-efficient [519, 2166, 2816]. cost-saving [267]. costs [881, 1326]. count [2728]. counterfactual [2790]. counterfeiting [2739]. countering [179]. countermeasure [365]. counting [24]. counts [85]. Coupled [114, 1566]. coupling [525, 1140, 1419]. Course [611, 1488]. Cover [630, 2513, 2814]. coverability [789]. coverable [1869]. coverage [685, 1554, 1888, 1949, 2194, 2317, 2549]. covering [757]. covers [2338]. COVID [800, 973, 1136, 1314, 1325, 1423, 1493, 1872, 2039, 2113, 2175, 2203, 2508, 2570, 2618, 2708, 2762]. COVID-19 [800, 973, 1136, 1314, 1325, 1423, 1493, 1872, 2039, 2113, 2175, 2203, 2508, 2570, 2618, 2708, 2762]. COVID19 [1746]. CovidO [2618]. CovSumm [2393]. CP [1332, 2282, 2410]. CP-ABE-based [1332]. CP2EH [1159]. CPN [2269]. CPN-based [2269]. CPU [297, 509, 521, 964, 1038, 1138, 1290, 1377, 1625, 2130, 2312, 2316, 2375, 2467, 2709, 2771, 2794]. CPU-GPU [1290, 1377]. CPU/GPU [2709, 2771]. CQARPL [1654]. crawler [423]. crawling [109]. CRDT [1445]. creating [2240]. Creation [2114, 2897]. Credit [909, 1595]. cRedit-based [909]. credits [754]. crevice [1476]. criteria [327, 552, 1024, 1032, 1450, 1776, 2107, 2343]. criterion [2551]. CRITIC [2329]. critical [342, 544, 965, 1055, 1401, 1790, 1941, 2431, 2514, 2593, 2900]. Criticality [1288, 1774]. Criticality-aware [1774]. crop [154, 1836, 2233, 2624]. CropReco [2233]. Cross [251, 588, 1186, 1315, 1475, 1565, 1871, 1953, 1979, 2196, 2218, 2271, 2298, 2714, 2718, 2757, 2889]. Cross-chain [1953]. cross-corpus [2271]. cross-disciplinary [588]. cross-domain [1953, 2298, 2889]. cross-entropy [1475]. cross-entropy-based [2718]. cross-input [1186]. cross-modal [1315]. cross-platform [1565]. cross-prompt [2757]. Cross-range [251]. cross-silo [1871]. cross-technology [2196]. cross-variation [1979]. crossed [1644, 1920]. crossover [821]. crosstalk [528, 781]. crow [1738, 1890]. crowd [999, 1680, 1710, 2156, 2650]. Crowdsensing [1649, 1966, 2864]. Crowdsensing-based [2864]. crowdsourcing [1286]. CRState [735]. cruise [337]. crunch [736]. cryptanalysis [2007]. crypto [1634]. crypto-processor [1634]. cryptocurrency [2357]. cryptographic [374, 658, 2053, 2242, 2422]. cryptography [258, 259, 565, 1091, 2627, 2703]. CryptoQNRG [2242]. Cryptosystem [445, 2245, 2388]. CSCL [17]. CSL [35]. CSL-driven [35]. CSM [2119, 2576]. CSO [750, 1809]. CSO-based [750]. CSO-ILB [1809]. CSP [1918]. CSPL [17]. CT [818, 1493, 1746, 1872]. cTMvSDN [1191]. CTO [1748]. cube [421, 1019, 1311, 1312, 1342, 1632, 1644, 2309, 2640]. cube-based [1342, 2640]. cubes [13, 1299, 1887, 1920, 1986, 2208, 2295]. cubic [961]. Cuckoo [156, 271, 783, 1028, 1235, 1488, 2458, 2798]. CUDA [301, 534, 844, 1938, 2058, 2140, 2775]. CUDA-powered [844]. cultivation [1271]. culture [449]. cum [2393]. Cumulative [1711]. CUR [2127]. current [573, 931, 1178, 1624, 1748, 2373]. curse [13]. curve [565, 1091, 1511, 1634, 2158, 2576, 2880]. curves [445, 2714]. custom [1626]. customer [178, 241, 345, 2028, 2173, 2277, 2351, 2485]. customized [1913, 2415]. Customizing [2010]. cuts [203]. cutting [138]. CVCF [2192]. CVO [1273]. CXR [2762]. CXR-based [2762]. Cyber [119, 120, 136, 318, 465, 467, 539, 621, 727, 1107, 1334, 1605, 1607, 1952, 2103, 2235, 2767]. cyber-attack [1607]. Cyber-physical [119, 120, 136, 467, 621, 727, 1107, 1605, 1952, 2235, 2767]. Cyberattack [943, 1955]. cyberattacks [2845]. cybercrime [271]. cybersecurity [2222, 2231, 2555, 2611]. cyberthreats [452]. Cycle [280, 1578, 1634, 2150, 2768]. CycleGAN [1872]. CycleGAN-learned [1872]. cycles [1920]. cycling [562]. cysts [1820].

D [215, 257, 492, 577, 623, 1691, 1949, 2657, 2876]. D* [2620]. D-NoCs [623]. D2D [2041, 2409]. DAACS [1207]. DAG [1841]. DAGs [1481]. Daily [334, 2203, 2430]. damage [2677]. dance [1434, 1583]. dandelion [2705]. danmaku [876, 2919]. danmaku-video [2919]. dark [860]. DART [66]. Data [13, 22, 39, 56, 67, 78, 129, 130, 143, 155, 169, 190, 210, 217, 221, 279, 331, 402, 431, 443, 500, 507, 514, 571, 606, 708, 719, 762, 777, 791, 825, 862, 934, 951, 962, 963, 1015, 1030, 1043, 1052, 1195, 1217, 1218, 1224, 1382, 1444, 1529, 1545, 1554, 1589, 1682, 1684, 1737, 1752, 1766, 1785, 1797, 1852, 1928, 1953, 1976, 1988, 2025, 2040, 2043, 2101, 2122, 2139, 2182, 2190, 2200, 2336, 2390, 2397, 2401, 2491, 2502, 2505, 2508, 2510, 2566, 2591, 2643, 2653, 2677, 2694, 2697, 2700, 2747, 2780, 2796, 2808, 2824, 2834, 2842, 2845, 2895, 2917]. data [35, 59, 63, 82, 122, 124, 131, 145, 152, 177, 188, 200, 201, 226, 242, 246, 247, 266, 298, 305, 307, 308, 335, 370, 382, 406, 589, 596, 600, 617, 666, 698, 750, 773, 775, 793, 798, 799, 809, 849, 889, 898, 904, 935, 942, 943, 945, 993, 1016, 1029, 1051, 1083, 1091, 1093, 1118, 1156, 1160, 1173, 1176, 1187, 1226, 1257, 1265, 1271, 1275, 1352, 1381, 1400, 1519, 1546, 1558, 1594, 1617, 1630, 1635, 1649, 1672, 1681, 1727, 1762, 1805, 1821, 1884, 1920, 2116, 2255, 2260, 2289, 2323, 2335, 2344, 2432, 2439, 2487, 2522, 2538, 2578, 2606, 2681, 2708, 2800, 2848]. data [2, 9, 28, 49, 53, 58, 77, 113, 140, 151, 202, 206, 209, 213, 229, 248, 261, 347, 354, 357, 394, 396, 433, 442, 469, 477, 498, 504, 505, 618, 677, 697, 716, 732, 751, 774, 786, 800, 824, 846, 860, 915, 920, 933, 959, 965, 1028, 1069, 1119, 1135, 1165, 1179, 1181, 1202, 1259, 1279, 1316, 1359, 1366, 1387, 1388, 1409, 1416, 1641, 1643, 1652, 1695, 1780, 1787, 1801, 1856, 1863, 1875, 1906, 1927, 1955, 2018, 2039, 2100, 2136, 2146, 2171, 2219, 2230, 2259, 2307, 2314, 2315, 2334, 2356, 2392, 2418, 2440, 2455, 2471, 2509, 2543, 2557, 2610, 2619, 2622, 2664, 2764, 2798, 2860, 2911]. data [41, 166, 170, 277, 450, 587, 588, 859, 1717, 2046, 2152, 2479]. data-centre [155]. Data-flow [2401]. data-intensive [824]. data-parallel [504]. database [9, 402, 413, 1385, 1777]. databases [118, 188, 422, 1865, 2201, 2285]. datacenters [309]. dataflow [1180, 2452]. dataset [800, 1211, 1353, 1567, 1734, 2039, 2113, 2280, 2372, 2611, 2779]. datasets [1765, 1860, 1968, 2353, 2468, 2508]. DAVmS [779]. dawn [2302]. day [710, 2405]. day-ahead [2405]. DBAHHO [1679]. DBSCAN [762, 1759]. DBT [2102]. DCCP [1916]. DCN [2145]. DCNN [438, 1204, 1963]. DCNs [1974]. DCUs [2653]. DDBWS [1070]. DDoS [211, 346, 612, 989, 1022, 1226, 1349, 1680, 2169]. DDS [1896, 1901]. DDS-cerberus [1896, 1901]. de-identification [2042]. deadline [816, 1070, 1227, 1346, 1581, 1688, 1690, 2057, 2553, 2841]. deadline-aware [1227, 1581, 2553]. deadline-constrained [1690, 2057, 2841]. deadline-driven [1346]. deadlock [1010]. deadlock-free [1010]. deaf [2188]. deblurring [463]. debugging [582, 1132]. decade [169]. decay [531]. decaying [2203]. decentralised [1068]. Decentralized [442, 518, 525, 999, 1037, 1416, 2163, 2526]. decimal [2898]. Decision [37, 39, 107, 110, 111, 177, 198, 252, 254, 327, 355, 366, 381, 521, 742, 905, 1032, 1207, 1267, 1338, 1450, 1527, 1638, 1712, 1776, 1831, 1893, 1913, 1999, 2107, 2139, 2228, 2343, 2356, 2526, 2777]. decision-makers [1913]. Decision-making [198, 742, 1032, 1527, 1831, 1913, 2107, 2139, 2228, 2343, 2777]. decisions [1337]. decoders [127]. decoding [2218, 2538]. decomposition [164, 311, 561, 667, 898, 1341, 1762, 1894, 2127, 2247, 2276, 2323, 2366, 2540, 2541, 2855]. decomposition-based [561, 2247]. decouple [235]. decreasing [33, 1511]. Dedicated [2817]. dedupe [462, 942]. deduplication [462, 1486, 1608, 1855, 1875]. Deep [60, 272, 274, 430, 444, 446, 457, 518, 553, 559, 561, 606, 611, 630, 660, 694, 721, 741, 760, 791, 797, 814, 833, 870, 896, 923, 930, 934, 947, 1061, 1098, 1099, 1137, 1143, 1164, 1204, 1210, 1214, 1232, 1243, 1265, 1284, 1307, 1319, 1324, 1358, 1393, 1395, 1461, 1462, 1464, 1468, 1487, 1493, 1575, 1604, 1614, 1616, 1637, 1645, 1679, 1713, 1724, 1768, 1783, 1800, 1825, 1828, 1836, 1864, 1881, 1894, 1926, 1934, 1960, 1974, 1978, 1999, 2013, 2079, 2082, 2083, 2147, 2182, 2210, 2221, 2256, 2274, 2279, 2493, 2530, 2549, 2550, 2591, 2615, 2665, 2696, 2702, 2786, 2812, 2813, 2867, 2886]. deep [71, 157, 196, 384, 437, 440, 494, 597, 643, 645, 647, 663, 672, 713, 738, 766, 780, 807, 826, 846, 853, 879, 897, 906, 919, 943, 954, 985, 1018, 1023, 1046, 1092, 1166, 1175, 1192, 1194, 1206, 1255, 1266, 1271, 1278, 1329, 1355, 1365, 1367, 1368, 1386, 1387, 1430, 1448, 1463, 1491, 1507, 1523, 1528, 1548, 1550, 1565, 1583, 1585, 1608, 1613, 1617, 1665, 1714, 1775, 1780, 1814, 1820, 1933, 1939, 1941, 1955, 2003, 2027, 2044, 2063, 2136, 2145, 2184, 2306, 2337, 2374, 2389, 2418, 2430, 2437, 2515, 2533, 2551, 2567, 2583, 2645, 2652, 2654, 2725, 2760, 2765, 2774, 2847, 2849, 2874, 2884]. deep [164, 466, 477, 615, 669, 684, 818, 820, 850, 946, 1272, 1428, 1547, 1572, 1595, 1963, 2046, 2263, 2316]. deep-learning [643, 1324, 1393]. DeepFakE [561]. Deeply [1376]. defect [10, 1279, 2614, 2724]. defects [837, 1665, 1733, 2306]. defenses [268, 2103]. deficit [2860]. deficit/hyperactivity [2860]. Defined [19, 543, 1008, 1022, 1032, 1191, 1269, 1343, 1540, 1546, 1678, 1680, 1971, 2055, 2093, 2145, 2169, 2216, 2229, 2486]. defined-DCN [2145]. defining [2764]. deflection [1056]. deflection-routed [1056]. defogging [2729]. deformable [1944]. degeneration [1324]. degradation [1467]. degraded [292]. degree [1648]. dehazing [659, 2896]. Delay [395, 754, 1571, 1613, 1675, 1800, 2023, 2667, 2715, 2796]. Delay-discretization-based [1571]. delay-sensitive [395, 1675, 2715]. delegation [937]. deletion [1388]. delirium [2436]. delivery [782, 1697, 1873, 2600]. demand [31, 76, 108, 767, 1575, 1784, 1894, 2405, 2711]. DEMATEL [666]. demeanor [2471]. demixing [2322]. demographic [2615]. Dempster [617, 1831, 1938, 2255, 2439]. demultiplexer [585]. Deng [2572]. denial [136, 268, 365, 2637]. denial-of-sleep [365]. denoising [524, 1363, 1693, 1827, 1925, 2537, 2724]. Dense [1121, 1726, 2621, 2816]. densely [416, 1128]. Density [391, 637, 1493, 1922, 1971, 2104, 2787, 2880]. dental [566]. denylist [1662]. Dependability [79, 1499, 1643, 2449, 2509]. Dependable [1905, 1916]. Dependency [1010, 1166, 1761, 2008, 2873]. dependent [1283, 1574]. dependent-tasks [1283]. deployed [416, 2902]. Deploying [906, 2617]. Deployment [54, 385, 580, 598, 685, 765, 1121, 1302, 1340, 1369, 1370, 1454, 1618, 1840, 1849, 2027, 2270, 2457, 2760]. deployments [2594]. Depression [1231, 2847]. depth [2450, 2621]. deraining [2002, 2371, 2801]. Derivative [343]. Derivative-based [343]. dermoscopy [1694]. descending [1051]. Descent [595, 1819]. description [809, 1570, 1925, 2756, 2859]. descriptive [222, 1548]. descriptor [2125, 2511]. Design [49, 126, 205, 223, 253, 265, 311, 326, 336, 377, 386, 387, 445, 468, 511, 535, 585, 636, 692, 744, 747, 811, 821, 822, 899, 950, 959, 986, 1033, 1038, 1040, 1076, 1098, 1111, 1112, 1129, 1151, 1152, 1155, 1167, 1187, 1200, 1226, 1282, 1320, 1347, 1362, 1363, 1457, 1466, 1479, 1587, 1611, 1656, 1703, 1714, 1726, 1735, 1822, 1830, 1838, 1839, 1842, 1877, 1902, 1966, 1987, 2073, 2084, 2089, 2192, 2220, 2224, 2257, 2329, 2386, 2407, 2427, 2432, 2448, 2452, 2453, 2455, 2481, 2512, 2554, 2567, 2586, 2602, 2630, 2664, 2698, 2701, 2741, 2816, 2817, 2892]. design-building [2455]. Design-time [2741]. Designing [378, 418, 579, 611, 990, 1023, 1505, 1605, 2029, 2035, 2448]. designs [96, 536, 724, 1042, 1131, 2217]. despite [2456]. destination [514]. destructive [193]. detail [1826]. details [2010]. detect [810, 1391, 1613, 1665, 1733, 2306, 2629]. Detecting [298, 318, 370, 448, 670, 1146, 1214, 1349, 1661, 2119, 2201, 2374, 2637]. Detection [25, 27, 146, 204, 220, 250, 274, 286, 336, 344, 346, 388, 391, 419, 426, 446, 464, 480, 561, 612, 632, 660, 702, 705, 921, 975, 1128, 1174, 1210, 1214, 1220, 1225, 1286, 1313, 1324, 1336, 1393, 1395, 1401, 1402, 1423, 1493, 1503, 1506, 1522, 1555, 1604, 1685, 1706, 1728, 1752, 1808, 1830, 1864, 1872, 1890, 1893, 1962, 2026, 2075, 2086, 2125, 2159, 2169, 2175, 2182, 2222, 2229, 2256, 2279, 2280, 2333, 2339, 2358, 2359, 2369, 2383, 2446, 2505, 2506, 2549, 2565, 2587, 2598, 2644, 2665, 2675, 2717, 2740, 2808, 2825, 2828, 2832, 2838, 2839, 2845, 2850, 2854, 2861, 2886, 2918]. detection [54, 193, 203, 271, 277, 359, 499, 597, 640, 646, 647, 672, 686, 737, 755, 780, 820, 834, 943, 947, 953, 981, 1022, 1031, 1198, 1209, 1231, 1244, 1343, 1344, 1360, 1361, 1373, 1424, 1436, 1538, 1548, 1551, 1564, 1573, 1576, 1593, 1595, 1607, 1625, 1636, 1783, 1827, 1847, 1955, 1964, 2006, 2121, 2168, 2174, 2209, 2260, 2273, 2377, 2417, 2433, 2465, 2484, 2514, 2533, 2535, 2540, 2555, 2570, 2574, 2592, 2614, 2616, 2627, 2638, 2649, 2657, 2666, 2683, 2684, 2724, 2726, 2829, 2835, 2847, 2868]. detection-based [1604]. detector [1918, 2113, 2691, 2699]. determination [154, 390]. determine [306]. deterministic [211]. develop [1146]. developed [834, 1170]. Developing [189, 905]. Development [143, 200, 348, 384, 386, 648, 666, 979, 1048, 1254, 1441, 1495, 1510, 1948, 2165, 2343, 2415, 2502, 2510, 2537, 2807, 2867]. developmental [2907]. developments [134]. Device [148, 392, 438, 1441, 1954, 2547, 2826]. device-free [438]. Device-specific [2547]. devices [159, 237, 1232, 1297, 1302, 1457, 1677, 1804, 1948, 1970, 1982, 2056, 2165, 2241, 2415, 2777]. DEVS [1843]. dew [1200, 1496, 2207, 2233, 2409]. dew-assisted [1200, 1496]. dew-edge-based [2233]. DewGame [2409]. DFW [2770]. DFW-PP [2770]. DG [2284, 2643]. DG-means [2643]. DHkmeans [1168]. DHkmeans-ldiversity [1168]. DHSA [1871]. DHSVM [154]. diabetes [2583]. diabetic [912, 1194, 1548]. diabetics [217, 1988]. Diagnosability [1309, 1311, 1312, 2048, 2061, 2498, 2704]. diagnose [2860]. diagnosis [48, 341, 351, 449, 654, 662, 684, 866, 1099, 1210, 1386, 1461, 1462, 1507, 1700, 1844, 1963, 2034, 2153, 2466, 2677, 2727, 2753, 2768]. diagnosis/prediction [2727]. diagnostic [619, 930, 1402, 2670]. diagonal [448, 2506]. diagram [1360]. dialect [1565]. dialogue [2328, 2469, 2536]. DialogueINAB [2536]. diameter [1465]. Diamond [801, 865]. dicing [2833]. dictionary [2443]. DIESEL [766]. differences [1186]. different [329, 548, 597, 990, 1053, 1648, 2248, 2270]. Differential [24, 64, 260, 436, 439, 529, 594, 745, 964, 1040, 2007, 2133, 2326, 2353, 2475, 2575, 2690, 2732, 2831, 2855]. differentially [620, 2501]. differentiation [1779, 1812]. DiffPageRank [529]. diffusion [32, 790, 1596, 1674, 2849, 2853]. digit [1063, 1544, 2530]. Digital [45, 670, 768, 1166, 1184, 1185, 1476, 1804, 2134, 2821, 2902]. digitized [99]. digits [645]. dilated [419, 1128]. dilation [683, 2068]. Dimension [38, 1922]. Dimension-aware [1922]. dimensional [11, 335, 372, 466, 717, 801, 898, 940, 951, 1051, 1300, 1363, 1456, 1473, 1635, 1672, 1734, 1821, 2124, 2144, 2353, 2408, 2697]. Dimensionality [921, 1409, 1630, 2456, 2611]. dimensions [958, 1734]. diminished [2759]. Direct [101, 577, 633, 940, 1470, 1624, 2004]. direct-forcing [633]. directed [492, 1069, 1841, 1947]. direction [1062]. directional [2317, 2499, 2814]. directions [744, 777, 2749]. directories [2671]. disabilities [768]. disaggregation [635]. disassembly [1929]. disaster [79, 853, 1491, 2368, 2516, 2558]. disciplinary [588]. discontinuous [2296]. discovering [944, 1761]. Discovery [407, 586, 1000, 1041, 1262, 1698, 2137]. discrepancy [1586, 2271]. Discrete [627, 764, 956, 1193, 1235, 1450, 1473, 1929, 2115, 2183, 2476, 2565, 2690, 2802]. discretionary [1264]. discretization [733, 1571]. discriminant [94, 2589]. Discriminate [1680]. discrimination [37]. Disease [40, 450, 454, 662, 725, 866, 923, 1214, 1386, 1463, 2038, 2421, 2456, 2717, 2753]. diseases [48, 905, 2227, 2591, 2670, 2702]. disjoint [789, 1019, 1587, 1869, 1920, 1928, 2640]. disorder [619, 1478, 2860]. disorders [840]. dispatch [111]. displacement [874]. display [312, 868, 1254]. Disruption [2639]. Dissemination [22, 962, 963, 1936, 2067]. dissimilarity [220, 1926]. dissipation [704, 747, 1362, 1877]. distance [10, 292, 682, 722, 779, 926, 1255, 1506, 2205, 2247, 2309, 2453, 2787]. distance-analytical [722]. distance-based [2247]. distance-related [2205]. distances [225]. distillation [2815]. Distributed [11, 18, 23, 56, 61, 63, 66, 70, 73, 80, 95, 142, 167, 206, 210, 268, 274, 278, 292, 301, 378, 390, 421, 442, 468, 490, 518, 530, 550, 556, 589, 758, 761, 792, 809, 849, 866, 881, 896, 921, 939, 944, 945, 959, 992, 1008, 1017, 1023, 1062, 1147, 1165, 1168, 1237, 1266, 1355, 1377, 1408, 1415, 1421, 1424, 1512, 1516, 1568, 1605, 1619, 1637, 1691, 1728, 1757, 1785, 1796, 1801, 1828, 1850, 1914, 1955, 2059, 2088, 2131, 2132, 2162, 2184, 2285, 2294, 2392, 2414, 2502, 2505, 2560, 2628, 2643, 2709, 2771, 2782, 2800, 2803, 2814, 2824]. distributed-memory [1691]. distributing [862]. Distribution [199, 349, 375, 500, 509, 663, 668, 754, 897, 1134, 1435, 1589, 1681, 1748, 1896, 2154, 2325, 2424, 2562, 2718]. distributional [1960]. distributions [1562]. divergence [2462]. divergent [1515, 2695]. diversity [693, 809, 1097, 1168, 2626]. divide [296, 1430, 2502]. divide-and-conquer [296, 1430]. Division [307, 2392]. DKEMA [1574]. DL [18, 1864, 2418, 2615]. DL-ESD [1864]. DMCSC [1796]. DMR [1930]. DMZ [636]. DNA [620, 1222, 1822, 2105, 2349, 2581, 2610]. DNA-based [1822]. DNetUnet [661]. DNN [1232, 1993, 2429, 2623, 2774]. DNNs [1857]. dOCAL [301]. Docker [1530]. docking [86, 554]. document [224, 1166, 1656, 1975, 2074, 2393, 2503, 2515, 2710, 2744]. document-based [224]. document-level [1975]. documentation [2261]. documents [323, 1668]. DODAF [1924]. DoE [1093]. DoE-based [1093]. does [1747]. dolphins [187]. domain [103, 214, 300, 386, 496, 527, 1289, 1439, 1586, 1788, 1882, 1953, 2006, 2058, 2100, 2124, 2130, 2298, 2457, 2465, 2469, 2851, 2889]. domain-based [496]. domain-specific [300]. domains [106]. dominance [448]. dominance-based [448]. dominating [1075, 1211, 1928]. Dominator [49]. dominoes [865]. donation [2157]. dose [522, 1310]. dose-based [1310]. DOSP [1568]. Dot [127, 585, 747, 950, 1111, 1112, 1131, 1157, 1303, 1335, 1726, 1839, 1879, 1919, 2217, 2630, 2816]. dot2dot [1222]. DOTMIX [1115]. DOTMIX-Pro [1115]. dots [2588]. double [1051, 1601, 2832]. double-spending [1601]. doubly [1871]. downhole [2909]. downtime [50]. DPAHMA [2362]. DPD [2898]. DPDK [1138]. DQN [1101, 2568]. dragonfly [568, 2413]. draining [465]. DRAM [1013, 1139]. DRAM-PCM [1013, 1139]. drawbacks [367]. drift [790, 1962, 2842]. drift-diffusion-recombination [790]. drinking [219]. drive [814, 854]. Driven [35, 357, 446, 810, 962, 963, 1187, 1346, 1472, 1607, 1849, 1964, 2493, 2534, 2542, 2690, 2710, 2831]. Driver [1354]. drives [512]. Driving [426, 669, 1143, 1344, 2479, 2714, 2806, 2877]. DRL [2863]. drone [506, 1255]. drop [364]. dropping [74, 2683]. DRTDBS [778]. drug [782, 2137]. DRX [184]. drying [1682]. DS [1831]. DS-1DCNN [1831]. DSDN [992]. DSL [295]. Dt [1373]. DTC [854]. DTC-SVM [854]. DTLS [243]. DTMC [2746]. Dual [560, 814, 921, 938, 1323, 1463, 1657, 1859, 2008, 2102, 2152, 2232, 2253, 2362, 2415, 2416, 2649, 2663, 2668, 2671, 2678, 2743, 2862]. dual-attention [2743]. dual-axis [921]. dual-biological-community [938]. dual-branch [2102]. dual-centric [2152]. dual-CISTs [560]. dual-core [2415]. Dual-drive [814]. dual-enhanced [2678]. dual-grain [2671]. dual-hop [2663]. dual-level [2008]. dual-modality [2416]. dual-population [2362]. dual-rate [1859]. Dual-source [1463]. dual-stage [1657]. dual-threshold [2253]. dual-tone [1323]. duct [183]. Dueling [2568]. Dung [2050]. duplex [2564]. duplicate [1636]. duplication [471, 2209]. duplication-based [471]. during [304, 1132]. duty [562]. DV [639, 1034]. DV-Hop [1034]. DV-Hop-based [639]. DVFS [339, 968, 1283]. DVFS-enabled [339, 1283].

DWT [1371, 2533]. dyeing [1714].

Dynamic [55, 62, 76, 82, 110, 185, 192, 244, 312, 351, 354, 380, 395, 408, 448, 465, 493, 509, 514, 524, 531, 544, 568, 596, 677, 715, 741, 749, 774, 795, 816, 819, 835, 841, 884, 917, 970, 975, 976, 1005, 1010, 1070, 1102, 1115, 1125, 1133, 1139, 1217, 1218, 1224, 1227, 1259, 1277, 1350, 1374, 1375, 1416, 1445, 1536, 1574, 1581, 1600, 1610, 1738, 1741, 1770, 1805, 1825, 1835, 1845, 1859, 1961, 1964, 1998, 2014, 2092, 2123, 2139, 2141, 2162, 2283, 2304, 2324, 2361, 2378, 2406, 2423, 2438, 2478, 2491, 2517, 2520, 2553, 2568, 2601, 2721, 2750, 2770, 2813, 2881]. dynamic-multithreading [1115]. Dynamical [2879]. dynamically [306, 1275, 1529]. Dynamics [3, 276, 1383, 1807, 1911, 2604, 2856]. DYRE [970].

e-commerce [2024]. E-CropReco [2233]. e-Health [652, 1159]. E-healthcare [1328]. e-learner [1921]. e-learning [1184, 1185, 2188]. e-mail [57]. E-RAD [2026]. e-smile [178]. E2LG [946]. EAAM [515]. EABM [228]. eagle [2172]. EALU [1790]. Early [154, 163, 179, 1152, 1477, 1548, 1864, 2026, 2168, 2369, 2665, 2753, 2918]. early-warning [1477]. earthquake [1755, 2494, 2673]. earthwork [1934]. ease [236]. east [2293]. Ebola [2780]. ECC [137, 1794, 2241]. ECC-based [137]. ECC-reliant [2241]. ECDSA [2883]. ECG [63, 351, 2363]. ECK [1692]. ECK-ANFIS [1692]. ECN [1974]. ecological [1466, 2904]. Economic [503, 833, 2529]. ecosystem [1536, 2185, 2453]. EDF [556]. EDF-RM [556]. EDGAN [463]. Edge [158, 159, 263, 357, 369, 371, 395, 399, 553, 589, 608, 683, 784, 847, 893, 916, 994, 1047, 1122, 1126, 1130, 1172, 1213, 1247, 1298, 1299, 1337, 1340, 1342, 1343, 1389, 1395, 1398, 1412, 1423, 1439, 1446, 1465, 1477, 1512, 1539, 1562, 1569, 1590, 1592, 1618, 1632, 1633, 1658, 1675, 1679, 1712, 1825, 1878, 1909, 1920, 1927, 1970, 1978, 1983, 1995, 1996, 2000, 2041, 2051, 2076, 2086, 2094, 2160, 2175, 2231, 2233, 2256, 2289, 2313, 2341, 2342, 2378, 2381, 2400, 2423, 2435, 2458, 2493, 2547, 2553, 2585, 2625, 2633, 2640, 2641, 2715, 2760, 2797, 2800, 2809, 2844, 2865, 2886]. edge-centric [1130]. edge-cloud [371, 589, 1446, 2094, 2715]. edge-computing [553]. edge-connected [2381]. edge-connectivity [1299, 1632, 2633]. edge-disjoint [1920, 2640]. edge-fog-cloud [2553]. edge-Internet [357]. edge-linking [1983]. edge-region-based [2051]. edges [784, 2768]. Editor [84, 117, 141, 171, 191, 230, 240, 284, 293, 303, 310, 400]. Editorial [98]. education [207, 609, 2400]. EDVWDD [962, 963]. EEG [676, 2116, 2363, 2676]. Effect [26, 379, 960, 1265, 1609, 2310, 2320, 2584, 2602]. Effective [9, 27, 50, 99, 244, 246, 247, 315, 340, 355, 409, 411, 423, 510, 787, 834, 883, 988, 1080, 1095, 1285, 1386, 1388, 1691, 1806, 1853, 1872, 1888, 1910, 2014, 2057, 2126, 2579, 2583, 2777, 2859, 2896]. Effectiveness [100, 112, 187, 2103]. Effects [684, 711, 2738]. Efficacy [818, 1465]. efficiencies [483]. Efficiency [26, 85, 521, 532, 637, 656, 895, 927, 949, 971, 1007, 1096, 1389, 1537, 1559, 2127, 2361]. efficiency-aware [2361]. Efficient [6, 22, 28, 86, 158, 262, 288, 317, 338, 371, 393, 414, 437, 476, 482, 505, 514, 529, 549, 576, 648, 712, 724, 729, 740, 757, 771, 821, 825, 871, 889, 905, 928, 951, 997, 1006, 1015, 1035, 1042, 1050, 1055, 1073, 1078, 1083, 1128, 1130, 1131, 1167, 1200, 1327, 1347, 1405, 1524, 1589, 1610, 1623, 1651, 1660, 1671, 1753, 1791, 1821, 1877, 1891, 1957, 1981, 2016, 2025, 2047, 2082, 2086, 2116, 2169, 2181, 2217, 2257, 2274, 2284, 2390, 2392, 2402, 2415, 2446, 2458, 2469, 2524, 2539, 2541, 2628, 2641, 2662, 2680, 2730, 2793, 2811, 2863, 2864, 2875, 2895]. efficient [35, 96, 216, 242, 266, 267, 273, 350, 352, 362, 369, 389, 404, 447, 471, 516, 519, 528, 651, 698, 699, 708, 714, 719, 726, 756, 775, 776, 824, 826, 856, 867, 910, 950, 954, 1016, 1028, 1064, 1115, 1179, 1214, 1240, 1278, 1333, 1335, 1473, 1479, 1525, 1544, 1574, 1582, 1616, 1622, 1630, 1664, 1695, 1736, 1740, 1777, 1789, 1840, 1851, 1871, 1918, 1944, 2014, 2042, 2052, 2068, 2072, 2073, 2089, 2107, 2118, 2166, 2170, 2177, 2196, 2199, 2202, 2213, 2290, 2340, 2370, 2412, 2448, 2523, 2554, 2576, 2609, 2619, 2658, 2733, 2816, 2817, 2843, 2846, 2853, 2872, 2898]. efficient [110, 140, 206, 533, 706, 998, 1121, 1138, 1171, 1180, 1296, 1396, 1612, 1759, 1839, 1856, 1875, 1922, 1938, 2060, 2100, 2132, 2301, 2324, 2465, 2795]. efficiently [810]. effort [2110, 2420]. EG [601]. eggs [115]. EICA [226]. EICA-optimized [226]. EIDM [2279]. eigenvalue [599]. Eisenstein [1497]. Ejection [1768]. Elastic [23, 88, 387, 399, 575, 707, 736, 922, 1741, 1785, 1809, 2226]. elasticity [474, 932, 2041]. Elastodynamic [613]. elder [611]. election [333, 550, 1027]. electric [1011, 1365, 1575, 2325, 2604, 2734]. electrical [1093, 2187]. Electricity [320, 751, 941, 2819]. ElectricVIS [59]. electro [1033]. electro-spin [1033]. electrocardiogram [2717]. electrochemical [2250]. electromagnetic [576, 2876]. electron [1740]. electronic [326, 2532]. Electrooculogram [794]. Electrooculogram-aided [794]. element [264, 1948, 2064, 2213]. elementary [1947]. elements [633, 744, 2769]. elephant [1368]. ELF [444]. eliminating [139, 1813]. elimination [15, 66]. elite [2761]. ELK [321]. elliptic [565, 1634, 2576]. ELS [582]. embarrassingly [1150]. Embedded [457, 459, 592, 847, 1088, 1378, 1602, 1673, 1745, 1790, 1849, 1907, 2677]. Embedding [122, 275, 283, 683, 1073, 1113, 1659, 1749, 2111, 2208, 2386, 2515, 2638, 2666, 2679, 2768, 2854]. embedding-based [2111]. embeddings [428, 429, 2198, 2230, 2759]. EMCP [2751]. EMD [833]. emergencies [1825, 2013]. emergency [111, 694, 853, 1239, 1491, 2275, 2558]. Emergent [78]. emerging [36]. emission [479, 2594]. emission-aware [2594]. Emotion [670, 934, 1233, 1561, 1842, 2102, 2126, 2271, 2273, 2536]. emotion-cause [1233]. Emotional [854, 1358, 1780]. emotions [1307]. emphasis [1239]. Empirical [107, 144, 491, 731, 765, 834, 1762, 2103, 2127, 2466]. EMRs [104, 324]. Emulation [412, 540, 582]. emulator [2231]. enable [894, 1427]. Enabled [210, 218, 339, 644, 740, 962, 963, 1039, 1283, 1420, 1555, 1702, 1756, 1997, 2076, 2202, 2340, 2369, 2390, 2409, 2419, 2444, 2493, 2605, 2665, 2765, 2809, 2844, 2861, 2887]. enabler [2071]. enables [1740]. Enabling [910, 1183, 1840, 1902, 2383]. encoder [839, 2724, 2839]. encoders [2271]. encoding [225, 226, 1248, 2096, 2898]. encrypted [304, 391, 1385, 2315, 2676]. Encryption [45, 134, 159, 289, 304, 308, 574, 658, 708, 1079, 1118, 1413, 1457, 1730, 1801, 1958, 2074, 2109, 2171, 2299, 2314, 2348, 2402, 2434, 2440, 2501, 2504, 2579, 2610, 2648, 2667, 2722, 2848]. End [466, 606, 658, 680, 1143, 1982, 2646, 2667]. End-to-end [466, 606, 658, 680, 1143, 2646, 2667]. endometrial [1461]. endorsement [1371]. endoscopic [37]. energies [1832]. Energy [19, 22, 26, 35, 87, 147, 184, 208, 216, 228, 262, 333, 350, 386, 418, 469, 514, 515, 557, 562, 573, 585, 614, 623, 635, 637, 657, 704, 708, 714, 719, 726, 747, 756, 767, 771, 776, 895, 904, 910, 950, 968, 971, 972, 1030, 1129, 1147, 1224, 1295, 1362, 1365, 1380, 1422, 1454, 1479, 1524, 1549, 1582, 1609, 1610, 1653, 1675, 1682, 1719, 1753, 1777, 1810, 1859, 1877, 1928, 1989, 2016, 2086, 2087, 2112, 2147, 2192, 2196, 2217, 2269, 2290, 2300, 2325, 2342, 2414, 2458, 2470, 2500, 2524, 2569, 2578, 2597, 2609, 2752, 2773, 2797, 2800, 2895]. energy [34, 110, 167, 342, 349, 389, 447, 467, 532, 618, 698, 785, 885, 902, 949, 1005, 1028, 1121, 1189, 1252, 1268, 1333, 1389, 1437, 1537, 1616, 1695, 1696, 1839, 2068, 2202, 2340, 2405, 2836, 2843, 2906]. energy- [19].

Energy-aware

[34, 228, 386, 515, 562, 623, 972, 1005, 1224, 1268, 1549, 1675, 1928, 2342, 2470, 2797]. energy-conscious [333]. energy-efficiency [949]. Energy-efficient [22, 35, 110, 216, 262, 350, 389, 447, 514, 698, 708, 714, 719, 726, 756, 771, 776, 910, 1028, 1121, 1333, 1479, 1524, 1610, 1616, 1753, 1777, 2086, 2196, 2524, 2843, 2895]. energy-harvesting [1189]. energy-objective [2800]. energy-oriented [657]. Energy-saving [968, 1030, 1147, 2500, 2597]. engine [414, 1139, 1440, 1828, 1969, 2077]. engineering [231, 482, 841, 986, 1506, 2108, 2129, 2193, 2453, 2481, 2512, 2602, 2687, 2698, 2712]. engines [1165, 2451]. English [496, 1284, 1648]. enhance [237, 2053, 2280, 2908]. Enhanced [77, 243, 324, 333, 444, 463, 506, 541, 543, 546, 566, 616, 672, 782, 872, 923, 929, 999– 1001, 1097, 1194, 1238, 1253, 1492, 1536, 1593, 1834, 1970, 1998, 2000, 2008, 2026, 2142, 2191, 2236, 2247, 2277, 2318, 2327, 2377, 2442, 2503, 2531, 2534, 2615, 2634, 2678, 2698, 2702, 2757, 2772, 2773, 2856, 2861, 2873, 2883]. enhancement [23, 103, 319, 488, 643, 1826]. Enhancing [88, 288, 359, 423, 780, 784, 802, 1096, 1419, 1604, 1786, 2243, 2418, 2655, 2711, 2776]. Enriched [1017]. enrollment [978]. Ensemble [27, 93, 180, 229, 337, 632, 682, 686, 721, 834, 946, 985, 1251, 1259, 1289, 1703, 1746, 1787, 2079, 2101, 2111, 2197, 2222, 2397, 2420, 2582, 2631, 2917]. Ensemble-based [682, 2222]. ensembled [1393, 2099]. ensembles [464]. ensure [908]. ENTC [2099]. Enterprise [201, 1145, 1243, 1293, 1398, 1838]. enterprises [177, 194, 2176, 2228, 2893]. entire [2477]. entities [283]. Entity [60, 68, 324, 876, 942, 1668, 2296, 2503, 2548, 2700, 2744, 2859, 2919]. Entity-level [876]. entrance [1351]. entropies [1832]. entropies-graph [1832]. Entropy [52, 252, 667, 1171, 1304, 1305, 1475, 1621, 2356, 2572, 2718, 2831, 2880]. Entropy-driven [2831]. entry [2671]. enumeration [1751]. Environment [57, 66, 69, 123, 147, 262, 264, 270, 274, 285, 290, 347, 353, 369, 371, 384, 395, 398, 589, 598, 606, 655, 750, 815, 835, 845, 856, 916, 996, 1077, 1103, 1133, 1143, 1173, 1195, 1205, 1246, 1300, 1399, 1408, 1420, 1427, 1446, 1486, 1487, 1608, 1665, 1678, 1689, 1692, 1733, 1741, 1776, 1794, 1809, 1835, 1841, 1850, 1884, 1905, 1966, 2009, 2053, 2065, 2085, 2128, 2206, 2219, 2292, 2306, 2330, 2360, 2438, 2449, 2458, 2478, 2524, 2527, 2532, 2553, 2571, 2583, 2681, 2730, 2780, 2841, 2844, 2887, 2897, 2901]. environmental [1267, 2247]. environments [6, 280, 515, 610, 647, 852, 889, 936, 1017, 1070, 1084, 1095, 1153, 1287, 1345, 1431, 1569, 1619, 1633, 1676, 1745, 1802, 1941, 1996, 2162, 2240, 2304, 2412, 2613, 2658, 2811, 2888]. EO [2481]. EO-based [2481]. EP2LBS [1612]. EP4DDL [1637]. Epidemic [973, 2867]. Epileptic [164, 676, 1371, 2383]. epistasis [1625]. EPSILOD [2132]. EQS [778]. Equality [2106]. equalizer [1639]. equation [766, 2855]. equations [594, 907, 1110, 2179, 2264, 2326, 2475]. Equilibrium [474, 932]. equipment [1770, 1868, 1929, 2755]. equipped [1164]. era [120, 942, 2123, 2867]. erasure [292, 1417]. ERPS [107]. erroneous [1245]. Error [737, 1216, 1317, 1545, 1663, 1727, 1891, 2032, 2267, 2616, 2627, 2631, 2701]. error-based [1663]. error-correcting [2627]. error-resilient [1891, 2701]. errors [2087, 2627]. ESD [1864]. ESDedup [1875]. ESKF [1140]. ESRGANs [1970]. establish [1399]. Establishing [329]. Establishment [3, 172, 276, 996, 1287]. estate [980, 2190, 2287]. estimate [722]. Estimating [2300]. Estimation [225, 251, 270, 426, 506, 585, 597, 831, 832, 950, 1104, 1188, 1245, 1255, 1329, 1408, 1556, 1598, 1602, 1745, 1768, 1836, 1839, 1858, 1907, 1941, 2044, 2130, 2149, 2156, 2184, 2203, 2303, 2313, 2420, 2450, 2549, 2682, 2720, 2734, 2788]. estimator [1541]. ET-DART [66]. ETA [1437]. ETA-HP [1437]. ETAS [1258]. ETD [1580]. Ethereum [290, 881]. Ethereum-based [881]. EV [1451]. evaders [161]. evaluate [422, 2364]. Evaluating [601, 838, 1326, 1348, 1467]. Evaluation [4, 14, 41, 65, 79, 100, 112, 132, 162, 292, 373, 475, 553, 684, 768, 772, 818, 893, 994, 996, 1002, 1077, 1129, 1192, 1205, 1241, 1272, 1276, 1414, 1440, 1457, 1463, 1495, 1498, 1513, 1530, 1562, 1594, 1638, 1692, 1694, 1754, 1865, 1921, 1992, 1998, 2106, 2165, 2224, 2242, 2285, 2372, 2398, 2400, 2449, 2454, 2455, 2509, 2510, 2545, 2567, 2767, 2799, 2893]. evaluation-based [1692]. evaluations [535]. evapotranspiration [334]. Event [73, 357, 642, 758, 764, 962, 963, 1187, 1438, 1472, 1551, 1553, 2249, 2352, 2476, 2555, 2912]. event-based [642, 2912]. Event-Driven [357, 962, 963, 1187, 1472]. event-triggered [764]. events [707, 770, 819, 953, 1876]. Everything [2071]. evidence [523, 1938, 2255, 2293, 2439, 2471]. Evolution [64, 153, 165, 745, 795, 938, 1825, 2013, 2353, 2481, 2575, 2690, 2750, 2823]. Evolutionary [213, 280, 452, 734, 763, 1478, 1490, 2153, 2247, 2319, 2386, 2629, 2651, 2751]. evolving [939, 988, 2009]. Exact [910, 1152, 1644, 1660]. exaflops [2106]. examination [222, 1461]. examinations [52]. Examining [2375]. example [2287]. exascale [2087, 2123, 2302]. exascale-era [2123]. excavator [1934]. exchange [517, 796, 967, 1045, 1149, 1526, 1587, 1757, 1987, 2084, 2220, 2399, 2447]. exclusion [1062]. executing [364]. execution [296, 297, 367, 648, 824, 887, 903, 1078, 1823, 1993, 2124, 2462]. executions [2901]. executor [236]. Exemplar [131]. Exemplar-based [131]. exercise [322]. expanded [236]. expansion [106, 1372, 1719, 1810]. expectation [733]. expectation-maximization [733]. Experience [551, 1505]. experiences [628, 1754]. experiment [17, 205, 1331]. experimental [379, 958, 1165, 1842]. experiments [995]. Explainable [640, 1746, 1933, 2456, 2734]. explanatory [478]. explicit [594]. exploit [1455]. exploitation [128, 236, 1492, 2565]. Exploiting [324, 787, 1010, 1493, 1722, 2451]. Exploration [1076, 1492, 1572, 2270, 2565, 2584, 2689]. explorations [709]. Exploratory [13, 691]. explore [2761]. Exploring [535, 930, 1244, 1433, 1806, 1892, 2328, 2467, 2905]. exponent [2141]. exponential [1792]. exponentially [2203]. exports [1448]. exposing [28, 140]. expression [196, 718, 915, 1230, 1842, 2517, 2641, 2852]. expressions [737, 2616]. expressway [2878]. extended [66, 657, 1636, 1886, 2247, 2277, 2891]. extender [1283]. Extending [1494, 1797]. extensible [896, 1760]. extension [89, 558, 746, 886, 958]. extensions [91, 238, 1537, 2161]. extensive [1331]. External [8, 2416]. External-attention [2416]. extra [1299, 1309, 1587, 1632, 2295, 2585, 2633]. extractable [1663]. Extracting [2860, 2919]. extraction [10, 116, 150, 496, 718, 770, 844, 846, 1233, 1476, 1658, 1694, 1847, 1975, 2117, 2141, 2506, 2548, 2645, 2678, 2697, 2871, 2873]. extractive [2268]. extractor [2605]. Extreme [152, 505, 1368, 1607, 1977, 2034, 2044, 2421]. extremely [2020]. eye [535, 1175].

F [1119, 1325]. F-LEACH [1119]. FaaS [2240]. fabric [1513, 1865, 2724]. face [440, 1255, 1476, 1745, 2042, 2113, 2194, 2511, 2675, 2684, 2740]. Facial [196, 224, 718, 1230, 1585, 1647, 1842, 1896, 2010, 2641, 2852].

Facilitating [675, 2157]. facility [2144]. fact [2119]. fact-checking-based [2119]. factor [380, 652, 973, 1274, 1443, 1495, 2024, 2201, 2491, 2605, 2900]. factor-based [2201]. factories [2333, 2350]. factoring [1640]. factorization [390, 406, 455, 627, 1322, 1827, 2541]. Factorized [907]. factors [29, 634, 1208, 1638, 2580, 2874]. factory [2485]. fading [2364]. FAGWO [1460]. FAGWO-H [1460]. Fail [1573]. Fail-Lagging [1573]. Failure [170, 550, 600, 1345, 1540, 1573, 1728, 1782, 1957, 2131, 2466, 2779, 2900]. failure-recovery [2131]. failures [1027, 2590]. Fair [1333, 1549, 1993, 2308, 2379]. fairness [2454]. fake [561, 2465, 2629]. fall [611]. false [1003, 2792]. family [2061]. fan [2282]. fan-in [2282]. fan-out [2282]. FAR [170]. FAR-MR [170]. farm [2374]. farms [2624]. Farsi [645]. FASR [2609]. FASR-LED [2609]. Fast [3, 24, 51, 170, 199, 225, 276, 516, 576, 631, 723, 858, 910, 1026, 1156, 1519, 1596, 1642, 1647, 1674, 1691, 1781, 1904, 1918, 1946, 1947, 1982, 2054, 2097, 2109, 2236, 2264, 2461, 2481, 2626, 2876]. faster [1115, 1344, 2477, 2667, 2861]. FastNBL [3, 276]. FastUDP [723]. fatality [2203]. fatigue [245, 711, 2877]. Fault [336, 419, 511, 528, 654, 674, 804, 805, 837, 842, 1019, 1097, 1150, 1198, 1317, 1342, 1347, 1460, 1498, 1544, 1700, 1704, 1790, 1844, 1863, 1905, 1920, 1930, 1963, 2000, 2034, 2120, 2153, 2194, 2253, 2260, 2266, 2272, 2498, 2542, 2627, 2640, 2677, 2732, 2765, 2768]. Fault-Tolerant [511, 837, 842, 1019, 1097, 1347, 1460, 1498, 1544, 1704, 1790, 1905, 1920, 1930, 2000, 2194, 2253, 2542, 2627, 2640, 2768]. faults [1019, 2466]. Faulty [1908, 2633]. FBMC [1902]. FC [2811]. FCA [1516, 2226]. FCI [2595]. FCM [454, 1268]. FE [1918]. FE-CSP [1918]. feasibility [199]. Feasible [211]. Feature [10, 40, 150, 157, 166, 204, 220, 221, 245, 272, 324, 335, 353, 437, 495, 631, 718, 733, 748, 770, 786, 844, 867, 991, 1059, 1105, 1181, 1376, 1481, 1551, 1588, 1617, 1635, 1667, 1729, 1765, 1783, 1847, 1883, 1886, 2012, 2034, 2038, 2125, 2155, 2227, 2238, 2255, 2265, 2277, 2347, 2353, 2397, 2439, 2471, 2506, 2550, 2565, 2572, 2583, 2657, 2697, 2729, 2734, 2757, 2770, 2827, 2852, 2865, 2871, 2915]. feature-based [40, 2038, 2734]. features [196, 222, 275, 351, 390, 485, 841, 1233, 1284, 1564, 1721, 1872, 1874, 2245, 2263, 2496, 2567, 2570, 2686, 2688, 2699, 2777, 2860, 2894]. Fechner [2511]. federated [761, 996, 1314, 1871, 2003, 2360, 2457, 2488, 2501, 2844, 2847]. federation [1880, 1996, 2472]. Feed [2659]. Feed-forward [2659]. feedback [610, 703, 1664, 1819, 1913]. FERMAT [885, 958]. Fermatean [2516]. ferrocyanide [2250]. ferroelectric [2373]. fertilizing [205]. FET [2373]. fetus [348]. FFT [838]. FGFS [1481]. FGPGAN [2403]. Fi [2068, 2196]. fidelity [1359]. Field [381, 1060, 1104, 1110, 1401, 1456, 1713, 1879, 1919, 2576, 2604, 2661, 2706]. field-programmable [1919]. fighting [2113]. file [7, 215, 1608, 1930, 2065, 2778]. files [802, 1323]. fill [1290]. Filter [51, 335, 366, 393, 643, 1125, 1363, 1877, 1902, 2281, 2434]. filter-based [335]. filtering [260, 393, 531, 654, 787, 975, 1100, 1390, 2232, 2389, 2404, 2615, 2736, 2896]. filters [1418]. Finally [483]. finance [430, 1771]. Financial [209, 1292, 1477, 1579, 1709, 1940, 2503]. find [1222]. Finding [278, 401, 500, 605, 997, 998, 1148, 1646, 1947, 2464]. Fine [83, 1117, 1874, 1946, 2210, 2371, 2587, 2736, 2790]. Fine-grained [83, 1874, 1946, 2587, 2736, 2790]. fine-tuning [2210]. finer [2403]. finer-grained [2403]. finger [621, 1868]. finger-like [621]. fingerprint [2468]. fingerprinting [2068]. finish [1721]. finite [264, 2064, 2141, 2213]. fire [173, 2918]. firebug [1257, 1392]. firefighting [100]. firefly [319, 365, 873, 1147, 1212, 1304, 1305, 1475]. firemen [2043]. fires [149]. firewalls [2505]. fireworks [482, 1001, 1223]. firms [405]. First [628, 1177, 1995, 2029, 2313]. fish [1766]. Fisher [2040, 2434, 2551]. fit [33, 500, 1268, 1747]. Fit-FCM [1268]. Fitch [498]. fitness [745, 1336, 2453]. Fiuncho [1625]. fixed [575]. fixed-alternate [575]. Flame [1978]. Flash [512, 1680, 1735, 2421]. flash-based [512]. flat [2339]. flexibility [990]. Flexible [602, 1020, 1118, 1346, 1422, 1660, 2299, 2339, 2613, 2662, 2671, 2758]. FlexSched [1078]. flight [991, 1304, 1305, 2512]. Flink [753, 988]. flip [97, 326, 2217]. flip-flop [326, 2217]. flip-flops [97]. floating [2616, 2898]. floating-point [2616]. flood [2168]. Floor [930, 1173]. floorplanning [2910]. floorplans [2362]. flop [326, 2217]. flops [1, 97, 1531]. flow [49, 53, 497, 618, 633, 687, 697, 880, 1178, 1193, 1401, 1515, 1617, 1813, 1932, 2354, 2359, 2401, 2460, 2520, 2525]. flower [2183]. flows [407, 940, 1059, 2316]. FLSGD [1508, 1509]. fluctuation [371]. fluid [2141]. fluids [1110]. Flux [1451, 2087]. Flux-weakening [1451]. Fly [528, 2211]. flying [929]. FM [910]. FM-index [910]. FMDH [223]. FMM [1318]. FMM-PM [1318]. focus [1037]. focused [423, 1052, 1338]. focusing [2110]. Fog [31, 280, 487, 578, 697, 714, 734, 796, 893, 924, 1041, 1080, 1103, 1136, 1144, 1223, 1302, 1328, 1339, 1447, 1458, 1498, 1519, 1539, 1563, 1616, 1631, 1676, 1689, 1744, 1753, 1778, 1828, 1850, 1905, 1915, 1936, 2003, 2085, 2116, 2128, 2206, 2207, 2331, 2360, 2377, 2387, 2431, 2470, 2553, 2558, 2595, 2597, 2613, 2681, 2716, 2730, 2793, 2811, 2856]. fog-assisted [1850]. fog-cloud [734, 1080, 2206, 2730]. fog-cloud-based [2595]. fog-computing-assisted [1328]. Fog-inspired [2558]. fog-integrated [1631]. fog/edge [1539]. FOI [1949]. folded [1887, 1986, 2208]. follower [2438]. following [1098]. FONS [924]. food [832]. foot [1548]. football [1192, 1395, 1433, 1724]. footprint [2904]. foraging [81]. foraging-based [81]. force [492, 1598, 2680]. force-directed [492]. forcibly [2381]. forcing [633]. Forecast [320, 2197, 2259]. Forecasting [76, 133, 430, 491, 687, 710, 760, 767, 1011, 1341, 1399, 1784, 1894, 1917, 2097, 2200, 2291, 2367, 2389, 2631, 2645, 2674]. forecasts [1368, 1818, 2293]. foreground [146]. forensic [1804, 2134]. forensics [1334]. Forest [149, 173, 195, 338, 725, 1432, 1442, 2780, 2918]. forests [76]. foreword [1049]. FOREX [88]. forgery [2592]. forgetting [2365, 2574]. form [1487]. Formal [1570, 1728, 2081, 2746]. formal-based [1728]. formally [642, 1068]. format [90]. formation [1415, 2409]. forms [1098]. formulation [623]. forum [2474]. forum-oriented [2474]. forward [137, 2659]. forwarding [433, 2356, 2401, 2796]. fountain [1681]. Four [1274, 2408, 2838]. Four-dimensional [2408]. Four-factor [1274]. four-stage [2838]. Fourier [1557, 1691]. FP [1204, 1262]. FP-DCNN [1204]. FP-growth [1262]. FPGA [234, 341, 376, 465, 515, 544, 690, 885, 1063, 1076, 1167, 1564, 1576, 1577, 1634, 1935, 2249, 2352, 2427, 2483, 2576, 2662, 2902, 2909]. FPGA-based [376, 465, 515, 1076, 1167, 1576, 1935, 2427, 2483, 2909]. FPGA-oriented [1564]. FPGAs [1050, 1479, 2149, 2898]. FPSA [1995]. FPSA-SMS [1995]. fractal [38]. Fraction [1768]. fractional [1429, 2326, 2658, 2846, 2855]. fragmentation [801]. Frame [1071, 1610]. Framework [50, 125, 126, 152, 156, 255, 275, 297, 350, 381, 421, 471, 496, 527, 542, 586, 610, 617, 686, 690, 713, 731, 793, 822, 843, 846, 848, 851, 866, 896, 948, 996, 1006, 1023, 1061, 1128, 1153, 1228, 1229, 1281, 1297, 1314, 1343, 1361, 1392, 1423, 1516, 1538, 1607, 1659, 1676, 1694, 1737, 1744, 1797, 1798, 1808, 1824, 1836, 1924, 1936, 1940, 2044, 2069, 2079, 2086, 2112, 2126, 2134, 2175, 2200, 2222, 2226, 2233, 2235, 2242, 2256, 2319, 2341, 2375, 2396, 2411, 2452, 2558, 2561, 2573, 2582, 2617, 2619, 2641, 2652, 2665, 2672, 2742, 2775, 2812, 2835, 2848, 2854, 2889, 2911, 2919]. framework [615, 723, 879, 955, 1134, 1367, 1547, 1691, 1843, 1922, 2529, 2635, 2645, 2795, 2856, 2869]. frameworks [78, 797, 809, 2446, 2858]. fraud [388, 1595]. Fredkin [2035]. free [183, 298, 438, 639, 1010, 1245, 1508, 1509, 1708, 1903, 2006, 2048, 2263, 2531, 2555]. frequencies [85]. Frequency [256, 532, 614, 968, 1323, 1560, 1571, 1693, 2122, 2851]. frequency-aware [968]. frequency-constrained [1693]. Frequency-domain [2851]. Frequent [13, 118, 396, 1772, 2072, 2296]. freshmen [978]. freshness [115]. Fresnel [2783]. FRHO [2243]. friendly [896]. FRLLE [550]. frog [541]. Frontier [1481]. frost [183]. frost-free [183]. FSCN [2356]. FSO [1393]. FSO-LSTM [1393]. FT [1097, 1790]. FT-EALU [1790]. FT-PDC [1097]. FtCFt [2194]. FTL [1735]. fuel [1534]. Full [96, 613, 724, 802, 821, 1347, 1418, 1585, 1743, 2073, 2564, 2853]. full-adders [724, 1347]. full-annulus [1743]. Full-convolution [1585]. full-duplex [2564]. full-text [802]. fully [574, 643, 754, 1025, 1796, 1852, 1958, 2826]. function [368, 444, 478, 579, 672, 836, 1151, 1493, 1792, 2027, 2108, 2216, 2240, 2305, 2329, 2336, 2537, 2571, 2819, 2877]. function-based [579]. Functional [471, 1225, 2299, 2617]. functionalities [813]. functionality [1605]. functions [573, 983, 1074, 1126, 1258, 1865, 2335, 2584]. fundus [673, 803]. furnace [1102]. Further [1705]. fuse [2297]. fused [437, 1376, 2443, 2839]. fusing [358, 2699]. fusion [113, 196, 622, 686, 705, 791, 934, 1278, 1507, 1783, 1847, 2012, 2138, 2238, 2255, 2263, 2416, 2428, 2439, 2519, 2521, 2548, 2657, 2691, 2729, 2757, 2827]. fusion-splitting [2428]. Future [81, 385, 573, 710, 775, 798, 799, 899, 1062, 1539, 2099, 2411, 2573, 2672, 2674, 2720, 2749]. fuzzing [510]. Fuzzy [33, 53, 176, 216, 219, 254, 256, 338, 341, 348, 415, 450, 607, 621, 733, 742, 760, 895, 905, 1031, 1119, 1127, 1228, 1229, 1272, 1381, 1451, 1597, 1631, 1661, 1687, 1784, 1904, 2125, 2136, 2183, 2195, 2202, 2243, 2245, 2329, 2335, 2516, 2564, 2577, 2580, 2605, 2696, 2735, 2799, 2821, 2914]. Fuzzy-based [216, 1119, 1228, 1229, 1381, 1784]. Fuzzy-logic-based [33].

G [542]. G-TOPSIS [542]. g2o [1140]. GA [1454, 2524, 2716]. GA-SAMP-MWPSO [1454]. GAAE [1787]. Gabor [1125]. GAC [112]. gain [806, 2054, 2773]. Gait [664, 840, 1722, 2238, 2263, 2496, 2693, 2860]. gallbladder [1210]. gallery [641]. Galois [1879]. Game [142, 177, 280, 314, 734, 969, 1337, 1355, 1395, 1412, 1514, 1584, 1620, 1686, 1754, 2317, 2331, 2378, 2409, 2472, 2802, 2823]. game-based [314, 2378, 2472]. gamma [1987, 2084, 2220, 2896]. GAN [713, 946, 1960, 2042, 2619]. GAN-based [1960, 2619]. GAN-SOM [713]. gantry [136]. gap [567, 1412]. GAP4NSH [2216]. garbage [379, 1089, 1403]. GARUDA [220]. gas [1604, 2661]. gastritis [850]. gate [127, 511, 837, 1474, 1726, 1919, 2035, 2286, 2448, 2853]. gate-diffusion-input-based [2853]. gated [1215, 2506]. gates [1155, 1347, 2257, 2817]. gateway [1677]. gathering [859, 1554, 2694]. Gauss [43]. Gaussian [220, 542, 702, 1392, 1894, 2178, 2591, 2839, 2896]. GBDT [366]. GBDT-KF [366]. GBM [2655]. GBM-optimized [2655]. GBRM [1659]. GCC [2248]. gCFS [1993]. GCN [2156, 2552]. Gear [253]. Gegenbauer [743]. Gelly [753]. GEMM [699, 1547]. gender [93]. Gene [331, 733, 1937, 2040, 2737]. general [731, 1467]. generalization [2551]. Generalized [122, 599, 654, 739, 907, 956, 1082, 1313, 1948, 2223, 2577]. generate [2113]. generated [827, 1931]. generating [2714]. Generation [300, 421, 432, 616, 899, 1072, 1244, 1331, 1719, 1810, 1838, 2017, 2091, 2190, 2242, 2328, 2349, 2384, 2454, 2586, 2659, 2710, 2784]. generative [457, 463, 1029, 1731, 1818, 1837, 1876, 1970, 2017, 2042, 2294, 2403]. generator [952, 1151]. generators [1076]. Generic [1079, 1822, 2132, 2576]. Genetic [39, 108, 197, 209, 629, 631, 685, 726, 905, 923, 956, 1082, 1088, 1106, 1190, 1423, 1447, 1452, 1466, 1597, 1599, 1757, 1765, 1787, 1834, 1888, 1910, 1946, 2040, 2049, 2163, 2175, 2216, 2277, 2286, 2373, 2689, 2727, 2856]. genetic-based [1447]. GENOCOP [522]. genome [1603]. genomics [2697]. genre [1387]. geo [2426]. geo-spatial [2426]. Geocast [1014]. geodesic [112]. geodynamic [2586]. geometric [138, 1523, 1588, 1980, 2737]. geometry [955]. GeoSpark [1165]. gestational [315]. gesture [1799]. gestures [918]. get [1418]. GHDC [2152]. GHFCM [331]. Giza [1943]. glaucoma [981]. GLCM [981, 2572]. GLNET [2828]. Global [281, 453, 790, 1026, 1149, 1282, 1294, 1301, 1316, 1452, 1729, 1866, 1975, 1979, 1994, 2015, 2050, 2097, 2313, 2668, 2682, 2687, 2698, 2712, 2828, 2882, 2894, 2904]. Global-Heap [1026]. global-local [2828]. glow [771]. glow-worm [771]. GM [2340]. GM-WOA [2340]. GMRES [1317, 2525]. GNN [2762]. GNP [2738]. GNP-alumina [2738]. goal [181]. Goaling [1081]. Godiva [1857]. GoF [1830]. golden [2191]. golf [957]. good [2048, 2204, 2498]. Google [155, 423, 1615]. Goore [1620]. governance [405, 666]. government [2686]. GPGPU [970, 1216, 1795, 2032]. GPGPUs [476]. GPS [99, 1408]. GPS-restricted [1408]. GPU [18, 24, 26, 86, 297, 398, 485, 498, 521, 533, 574, 644, 649, 774, 792, 793, 847, 901, 907, 940, 964, 993, 1038, 1067, 1082, 1127, 1171, 1285, 1290, 1327, 1348, 1377, 1401, 1432, 1473, 1482, 1501, 1574, 1609, 1691, 1705, 1769, 1774, 1789, 1923, 1947, 1950, 1973, 2036, 2062, 2096, 2105, 2130, 2140, 2167, 2199, 2264, 2284, 2316, 2346, 2444, 2461, 2467, 2709, 2771, 2794, 2906, 2916]. GPU-accelerated [485]. GPU-aware [993]. GPU-based [26, 86, 533, 792, 793, 847, 964, 1327, 1574, 1923, 1947, 2096]. GPU-enabled [644, 2444]. GPU-oriented [774]. GPUs [132, 135, 522, 534, 613, 706, 722, 1006, 1078, 1134, 1407, 1471, 1479, 1559, 1610, 1993, 2029, 2110, 2132, 2321, 2462]. GPUs-RRTMG LW [706]. Gradient [334, 358, 366, 484, 595, 1266, 1550, 1858, 1893, 1977, 2211, 2920]. gradient-boosted [366]. grain [2671]. grained [83, 126, 1117, 1874, 1946, 2371, 2403, 2587, 2736, 2790]. Gram [87]. granular [2238]. granularity [107, 497, 1036, 2121, 2671, 2741]. granularity-based [107]. Graph [8, 15, 73, 80, 249, 295, 323, 347, 356, 376, 377, 403, 409, 414, 418, 421, 422, 492, 626, 693, 698, 787, 851, 855, 890, 1010, 1130, 1220, 1244, 1270, 1321, 1406, 1517, 1593, 1606, 1623, 1659, 1670, 1747, 1767, 1795, 1803, 1832, 1834, 1841, 1899, 1947, 1961, 1964, 2011, 2027, 2114, 2131, 2142, 2146, 2156, 2260, 2313, 2354, 2358, 2393, 2428, 2460, 2465, 2534, 2552, 2560, 2598, 2633, 2649, 2669, 2670, 2750, 2763, 2772, 2776, 2815, 2822, 2837, 2868, 2873]. graph-based [249, 787, 1321, 1795, 2465, 2670]. Graph-enhanced [2534]. graph-parallel [80]. Graphene [2554]. Graphene-based [2554]. GraphFrames [753]. graphic [755, 1639, 2567]. graphics [455, 532, 812, 1610]. graphite [46]. graphlet [2831]. GraphMap [356]. graphs [278, 544, 560, 567, 682, 683, 789, 939, 988, 1405, 1424, 1552, 1570, 1751, 1761, 1869, 1931, 2048, 2078, 2177, 2338, 2498, 2559, 2859]. Grasshopper [281]. gravitational [552, 925, 931, 1866]. gray [2035, 2502]. grayscale [2533]. greater [2152]. greedy [221, 2643]. Greek [2185]. Green [49, 200, 262, 339, 386, 884, 1466, 1590, 1857, 1959, 2892]. greenhouse [2165]. greening [1283]. Grey [47, 173, 956, 1077, 1484, 1866, 1929, 2034, 2193, 2265, 2737]. Grid [106, 147, 257, 872, 904, 1046, 1379, 1593, 1719, 1810, 1906, 2052, 2197, 2406, 2542, 2634, 2718, 2773, 2778, 2875]. Grid-based [106]. grids [467, 886, 1109, 2525]. grocery [1917]. groundwater [176]. group [279, 864, 1142, 1375, 1523, 1786, 1913, 2412, 2526, 2585, 2638, 2848, 2912]. group-based [1142]. grouped [161, 1297]. Grouping [307, 427, 1376, 2094]. Grouping-Based [307]. Grover [2030]. growth [348, 495, 1262]. GRU [168, 1895, 2091, 2291]. GS4 [626]. GSAGA [1707]. GT [2523]. GT-NRSM [2523]. Gtpsum [2742]. Guard [2774]. guidance [530, 2385, 2534, 2695]. guide [109, 1363]. Guided [1465, 1481, 1517, 2076, 2131, 2232, 2682, 2721, 2742, 2801]. guidelines [411]. GVLE [2096]. GWO [331].

H [1015, 1460]. H-Code [1015]. Hadoop [519, 533, 569, 730, 785, 804, 805, 2529]. Hadoop-based [2529]. hailing [1367, 1894]. half [2768]. Hamilton [1627]. Hamiltonian [1829, 1920, 2768]. Hamming [225]. hand [270, 1329, 2332]. handle [1805]. handling [212, 688, 989, 2249, 2352, 2842]. Handoff [1249]. handover [1300]. handwritten [645]. haplotyping [2857]. HAR [1253]. hardening [1348]. Hardware [78, 413, 465, 502, 509, 591, 894, 1114, 1363, 1522, 1537, 1727, 1881, 1891, 1907, 2047, 2249, 2302, 2352, 2452, 2731, 2732, 2806, 2817]. Hardware-accuracy [1891]. Hardware-aided [413]. Hardware-based [1537]. hardware/software [2731]. harnessing [1610]. Harris [1054, 1679, 1820, 1979, 2355]. harvested [2836]. harvesting [124, 349, 1189, 1976]. Hash [427, 836, 1315, 1884]. Hash-based [427, 1884]. Hash-tree [427]. hashing [1854]. hate [2506]. Hawks [1054, 1679, 1820, 1979, 2355]. haystack [401]. Hazel [995]. HBI [1905]. HBI-LB [1905]. HC [2412]. HCI [1183]. HCN [1829]. HDFS [23, 802]. HDNN [1565]. HDSAP [2283]. head [771, 1268, 1606, 2172, 2243, 2543, 2622, 2899]. header [2216]. healing [1728]. Health [47, 67, 144, 213, 289, 446, 539, 611, 652, 1107, 1159, 1274, 1386, 1400, 1702, 1882, 2119, 2664, 2874]. health-care [1274]. Healthcare [77, 243, 341, 357, 450, 595, 879, 948, 984, 1119, 1136, 1328, 1422, 1655, 1761, 1787, 1902, 1915, 1951, 2251, 2369, 2390, 2435, 2583, 2646, 2804, 2848, 2863, 2864]. Heap [1026]. hearing [1169, 1990]. hearing-impaired [1169, 1990]. Heart [48, 435, 830, 905, 923, 1214, 2086, 2227, 2456, 2717]. heart-related [2227]. heartbeat [344]. heat [608, 2067]. heat-based [608]. heavy [1654]. helicopter [198, 2305]. hello [368]. Helmet [1557]. hematological [1308]. HEp [1586]. HEp-2 [1586]. herbal [172]. herd [227, 857, 1368]. Hermitian [445, 599]. Heronian [2799]. hesitant [742]. Heterogeneity [114, 119, 236, 922, 994, 2283, 2341]. Heterogeneity-aware [922, 2283, 2341]. Heterogeneous [18, 20, 139, 145, 233, 234, 239, 292, 309, 333, 354, 367, 378, 383, 456, 473, 489, 490, 502, 504, 509, 519, 521, 524, 525, 563, 594, 601, 656, 678, 716, 730, 902, 911, 958, 971, 993, 1050, 1069, 1084, 1088, 1133, 1333, 1377, 1437, 1550, 1582, 1589, 1616, 1637, 1677, 1696, 1730, 1777, 1785, 1802, 1860, 1900, 1913, 2022, 2206, 2278, 2304, 2331, 2370, 2406, 2419, 2438, 2486, 2487, 2638, 2706, 2751, 2794, 2814, 2823, 2843, 2885, 2906]. HetNet [2720]. Heuristic [180, 623, 778, 785, 1295, 1316, 1382, 1532, 1578, 1776, 1791, 1805, 1821, 1834, 1841, 2050, 2227, 2236, 2240, 2312, 2578, 2857]. heuristic-based [2240]. heuristics [219, 339, 575, 835, 995, 1761, 2029]. heuristics-based [2029]. HEVC [2083, 2221]. Hexagonal [22, 2349]. hexapod [62]. HFBO [2421]. HFBO-KSELM [2421]. HGP4CNN [1006]. hibernation [2866]. hidden [28, 140, 714, 763, 1500, 1872, 2861]. hiding [122, 226, 1359]. Hierarchical [289, 314, 340, 391, 773, 829, 889, 908, 1156, 1168, 1266, 1414, 1472, 1485, 1606, 1650, 2033, 2146, 2208, 2229, 2406, 2616, 2763]. hierarchies [502]. Hierarchy [65, 1024, 1621, 1983]. hierarchy-betweenness [1983]. Hierarchy-entropy [1621]. High [4, 11, 62, 99, 124, 135, 192, 212, 229, 250, 300, 301, 335, 360, 401, 516, 590, 597, 648, 656, 681, 706, 729, 752, 793, 794, 806, 851, 853, 854, 858, 874, 880, 899, 923, 929, 976, 979, 1007, 1038, 1040, 1059, 1134, 1143, 1157, 1161, 1175, 1180, 1192, 1256, 1316, 1330, 1335, 1345, 1357, 1359, 1384, 1396, 1401, 1434, 1474, 1491, 1500, 1523, 1564, 1577, 1583, 1617, 1635, 1646, 1672, 1760, 1860, 1902, 1907, 1992, 2011, 2035, 2058, 2062, 2070, 2087, 2112, 2136, 2236, 2263, 2324, 2353, 2525, 2586, 2682, 2702, 2714, 2775, 2806, 2840, 2913]. high-accuracy [874]. High-capacity [681]. High-correlation [2682]. high-cost [729]. High-dimensional [11, 335, 1635, 1672, 2353]. High-efficient [706, 1180]. high-energy [2087]. High-fidelity [1359]. high-gain [806]. high-impedance [1335]. High-level [301, 853, 979, 1330, 1434, 1491, 1577, 1583, 1907, 2236]. high-order [1059, 1396, 2775]. High-performance [4, 62, 124, 135, 212, 229, 250, 300, 360, 597, 648, 752, 793, 794, 806, 851, 880, 923, 976, 1134, 1161, 1256, 1384, 1474, 1523, 1617, 1646, 1860, 1992, 2011, 2035, 2062, 2070, 2136, 2586]. high-power [1192]. high-precision [1401, 2913]. high-Q [1040]. high-quality [929, 1316, 2840]. high-recall [401]. high-resolution [2058]. High-speed [192, 899, 1143, 1760, 2702, 2714, 2806]. high-throughput [2324]. high-utility [1357]. Highly [606, 723, 740, 1430, 1470, 2096, 2213, 2404, 2853]. highly-scalable [2404]. highway [572]. Hilbert [1963]. hill [337, 2180]. hill-climbing [2180]. HIN-based [2232]. HiperView [976]. histogram [272]. histological [877]. histopathological [559, 2828]. historical [2799]. History [1013, 1291]. HMAC [1884]. HMAC-R [1884]. HNN [365]. hoc [248, 346, 349, 363, 796, 929, 1661, 1666, 1998, 2290]. HOG [196]. hole [397, 2549]. holistic [823]. home [156, 1037, 1053, 1404, 1427, 1484]. homes [2658]. homogeneous [54, 1380]. homology [2803, 2814]. homology-based [2803, 2814]. homomorphic [45, 574, 1079, 1871, 1958, 2348, 2501]. homotopy [1162]. honey [2708]. Hop [123, 330, 639, 1034, 2146, 2546, 2663]. Hopfield [365]. horizon [120, 2097]. horizontal [399, 1417, 2886]. horizontally [2323]. Horn [1671]. hospital [2516, 2664]. hospitalized [2436]. host [514]. hosted [1414, 2798]. hosting [2764]. hot [1351, 2139, 2599]. hot-module-aware [2599]. Householder [455]. hovering [2731]. HP [1437]. HPC [7, 21, 235, 554, 650, 691, 765, 786, 860, 958, 990, 1132, 1573, 1740, 1985, 2065, 2185, 2302, 2528, 2590]. HPC-systems [990]. Hpcfolder [1087]. HPDMS [124]. HPM [913, 914]. HPMC [1700]. HPT [368]. HRV [1560]. HS6 [1015]. HSAC [829]. HSAC-ALADMM [829]. HSGWO [2382]. HSVNN [1630]. HTML5 [2769]. HTTP [2637]. Human [63, 156, 273, 587, 606, 788, 985, 1059, 1104, 1105, 1120, 1149, 1175, 1253, 1297, 1307, 1329, 1603, 1624, 1638, 1763, 1868, 1965, 2009, 2126, 2225, 2238, 2263, 2621, 2788]. human-computer [2126]. human-inspired [1149]. human-land [273]. human-robot [1624, 1763, 1868, 2225]. hunting [1924].

HWF [973]. Hybrid

[12, 21, 126, 143, 208, 237, 245, 297, 372, 413, 478, 482, 497, 540, 718, 719, 797, 821, 838, 874, 905, 958, 1013, 1021, 1038, 1097, 1106, 1125, 1153, 1212, 1235, 1238, 1281, 1285, 1314, 1320, 1349, 1372, 1382, 1383, 1393, 1426, 1460, 1484, 1503, 1507, 1516, 1614, 1635, 1688, 1692, 1700, 1707, 1765, 1807, 1823, 1851, 1891, 1943, 1963, 1977, 1989, 2016, 2053, 2058, 2130, 2145, 2243, 2262, 2274, 2275, 2345, 2362, 2367, 2393, 2470, 2498, 2506, 2524, 2549, 2598, 2615, 2654, 2717, 2727, 2730, 2736, 2738, 2744, 2755, 2758, 2765, 2802, 2838, 2854, 2866, 2882, 2887]. hybrid [166, 226, 578, 717, 911, 985, 1082, 1139, 1181, 1219, 1223, 1316, 1379, 1397, 1483, 1630, 1843, 1888, 1917, 1994, 2108, 2155, 2288, 2305, 2325, 2340, 2357, 2421, 2490, 2636, 2647, 2841, 2852, 2881, 2903]. Hybrid-based [1314]. Hybridized [2413]. HybriDroid [834]. hydraulic [1941]. hydrodynamic [186]. hydropower [1929]. hygienic [219]. HyPar [1516, 2226]. HyPar-FCA [1516, 2226]. hyper [343, 1599, 2478, 2689]. hyper-parameters [1599, 2478]. hyper-spectral [343]. hyper-task [2689]. hyperactivity [2860]. hyperautomation [2874]. Hyperbolic [1622]. hyperchaotic [2879]. hypercube [377, 1309, 2000, 2098, 2768]. hypercubes [674, 683, 1887, 1986]. hyperelliptic [1091]. Hypergraph [305, 973, 1652]. Hypergraph-partitioning-based [1652]. hypergraphed [2544]. hypergraphic [2381]. hyperledger [1513, 1865]. hypernetwork [1983]. hyperparameter [559, 2184]. hyperparameters [2727]. hyperspectral [460, 485, 739, 870, 901, 1025, 2656]. hyperthreading [387]. hypervisor [50]. hypervolume [397]. hypotheses [1342]. hypothesis [1937]. hypoxia [2665]. hysteroscopic [1461].

I/O [723, 949, 995, 1628, 1629, 2065]. IaaS [108, 1688, 1690]. IAP [2140]. ICD [2261]. ICD-10 [2261]. ICE [2675]. ICE-YoloX [2675]. ICP [566]. ICS [2845]. ICS-IDS [2845]. ICT [691]. ID [1855]. ID-based [1855]. IDCOS [915]. Ideal [2712]. Identification [458, 664, 780, 852, 1186, 1486, 1536, 1615, 1763, 1836, 1908, 1965, 2042, 2191, 2274, 2327, 2413, 2624, 2849, 2917]. Identify [2366, 2482, 2838, 2845]. identifying [107, 1528, 2474]. Identity [1163, 1184, 1185, 1366, 2491]. Identity-based [1163, 1366]. IDPC [391]. IDPNet [2788]. IDS [1393, 2433, 2845]. IDTracS [2401]. IEEE [123, 2898]. IEEE-754 [2898]. IF [238]. IF-statements [238]. IFIM [2072]. Igniteg [644]. Igniteg-GPU [644]. IHP [524]. II [912, 1381, 1548, 1798, 2391, 2647]. II-based [2647]. III [522]. IIoT [510, 1442]. IIoTs [1831]. IIR [1536]. ILB [1809]. ill [871]. ill-conditioned [871]. illness [477]. illumination [146]. illumination-invariant [146]. ILP [623]. IM [2810]. IMA [2267]. Image [45, 71, 118, 157, 224, 255, 353, 437, 454, 460, 508, 524, 622, 643, 661, 673, 684, 739, 743, 783, 803, 842, 843, 870, 912, 925, 957, 1029, 1044, 1081, 1085, 1099, 1127, 1167, 1230, 1298, 1304, 1305, 1315, 1336, 1363, 1428, 1433, 1434, 1462, 1463, 1475, 1476, 1557, 1585, 1658, 1720, 1792, 1826, 1853, 1860, 1886, 1904, 1925, 1926, 1944, 1969, 2002, 2051, 2079, 2109, 2154, 2209, 2253, 2275, 2276, 2297, 2425, 2434, 2445, 2450, 2483, 2492, 2504, 2517, 2579, 2581, 2635, 2648, 2656, 2701, 2718, 2722, 2726, 2754, 2783, 2790, 2801, 2853, 2869, 2871, 2896]. Image-based [508]. image-text [1315]. imagery [58, 1025, 1521]. images [37, 251, 343, 384, 485, 559, 622, 801, 807, 820, 844, 877, 901, 930, 1007, 1308, 1407, 1506, 1548, 1566, 1694, 1713, 1746, 1858, 1872, 2058, 2074, 2125, 2151, 2492, 2533, 2537, 2570, 2596, 2624, 2705, 2884]. imaging [103, 818, 1141, 1277, 1723, 2004, 2303, 2355]. imbalance [1202, 1383, 1752, 2182, 2917]. imbalanced [493, 759, 1259, 1672, 1787, 1968, 2018, 2154, 2717]. immersed [633]. Immune [214, 1480, 1755]. Impact [618, 1036, 1261, 1495, 1667, 2248, 2319, 2849, 2874]. impaired [1169, 1990]. Impairment [332, 575]. Impairment-aware [332, 575]. impedance [1335]. imperfect [442]. imperialist [319, 2258]. implant [566]. implantation [1643]. Implementation [102, 265, 312, 322, 336, 387, 509, 511, 534, 591, 625, 649, 712, 747, 751, 804, 805, 812, 821, 822, 874, 959, 980, 1008, 1109, 1187, 1221, 1271, 1327, 1347, 1377, 1397, 1473, 1482, 1544, 1611, 1639, 1735, 1769, 1955, 2096, 2130, 2149, 2212, 2249, 2284, 2352, 2554, 2576, 2630, 2732, 2898, 2903]. implementations [838, 1317, 2179, 2348, 2745]. implemented [2415]. implementing [232, 1052, 1505]. implicit [1116, 1650, 1664, 1670, 2328]. implicitization [1892]. importance [1708]. impossible [2007]. impossible-differential [2007]. imprecise [1162]. imprint [150]. Improve [21, 249, 325, 517, 703, 924, 1067, 1403, 1664, 1749, 2028, 2173, 2189, 2283, 2351, 2830]. Improved [27, 47, 180, 218, 223, 260, 296, 308, 335, 366, 386, 391, 541, 663, 762, 782, 783, 833, 898, 991, 1054, 1075, 1091, 1101, 1172, 1188, 1190, 1237, 1279, 1344, 1381, 1428, 1431, 1488, 1511, 1528, 1534, 1578, 1608, 1758, 1799, 1811, 1883, 1893, 1908, 1929, 1987, 1993, 1994, 1997, 2018, 2031, 2055, 2077, 2084, 2092, 2113, 2117, 2155, 2180, 2183, 2186, 2193, 2204, 2220, 2226, 2230, 2342, 2382, 2425, 2499, 2529, 2531, 2565, 2581, 2603, 2620, 2675, 2705, 2718, 2740, 2805, 2872, 2920]. Improvement [39, 183, 224, 322, 425, 927, 1031, 1242, 1596, 1674, 1870, 1942, 2276]. Improving [20, 347, 369, 456, 547, 561, 648, 669, 752, 759, 949, 1164, 1191, 1203, 1282, 1286, 1510, 1537, 1558, 1559, 1761, 1795, 1926, 2127, 2135, 2137, 2261, 2408, 2424, 2459, 2538, 2551, 2775, 2837]. impulse [2191]. imputation [1135, 1649]. in-kernel [735]. in-memory [9, 644, 793]. in-place [1394]. incentive [1249, 1613, 1966]. including [1125]. inclusion [680]. incomplete [1211, 2475]. incompressible [1110]. Incorporate [1984]. Incorporating [452, 2683, 2738, 2859]. increase [970]. increasing [827]. Incremental [109, 181, 388, 677, 939, 2072, 2152, 2866]. indecomposable [13]. Independent [1125, 1341, 1410, 1497, 2034, 2824]. Index [189, 313, 710, 910, 1751, 2078, 2293, 2631, 2645, 2789, 2895]. Index-based [1751, 2895]. index-tracking [2645]. indexes [154, 677]. indexing [11, 106, 801, 951, 2539]. indicator [1443]. indicators [315, 850]. indices [1832]. indigenous [611]. individual [269, 556]. individuals [794, 1175]. Indoor [438, 546, 815, 1173, 1793, 2068, 2468]. induced [338, 1572, 2888]. induction [854]. inductive [2556]. inductor [806, 1040]. Industrial [336, 439, 621, 763, 1555, 1682, 1909, 1927, 1960, 1963, 2148, 2231, 2292, 2327, 2380, 2859]. industry [209, 703, 1527, 2005, 2028, 2173, 2224, 2334, 2343, 2351]. inertial [484, 814, 1353, 1567]. infarct [508]. infarctions [1146]. inference [256, 450, 1631, 1687, 1828, 2261, 2316, 2321, 2865]. InfiniBand [894]. infinite [2191, 2879]. Influence [178, 239, 249, 405, 1499, 1861, 2375, 2569, 2690, 2761, 2810]. Influences [961, 1477]. influencing [29]. influential [605, 944, 1148, 1646, 2366]. Influenza [477]. Influenza-like [477]. INFO [2718]. information [25, 105, 114, 149, 244, 252, 277, 313, 353, 622, 651, 888, 897, 1092, 1184, 1185, 1199, 1233, 1243, 1276, 1278, 1372, 1391, 1463, 1593, 1615, 1749, 1851, 2054, 2067, 2319, 2612, 2638, 2657, 2683, 2741, 2752, 2859, 2880, 2909]. information-sharing [2319]. informative [2801]. informed [2828]. Infrared [926, 1298, 1548]. infrastructure [4, 22, 370, 828, 1338, 1352, 1439, 1655, 1707, 1992, 2014, 2862, 2875, 2904]. infrastructure-as-a-service [1338]. infrastructures [79, 1093, 2069, 2634]. initial [2163]. initialization [325]. injection [28, 140, 1003, 1317]. injuries [154]. injury [957, 1192]. InKS [235]. Innate [1480]. innovation [201, 1243, 1398]. innovative [92, 1129]. input [511, 837, 1134, 1186, 2205, 2448, 2853]. inputs [1089]. insensitive [1886]. insertion [696, 1598]. insight [2711]. Insights [116, 2746]. inspection [2718, 2878]. inspired [552, 917, 948, 1149, 1273, 1653, 1755, 1817, 2558, 2602, 2902]. Instance [1105, 1597, 1897, 2063, 2886]. instance-level [2886]. instant [675, 1775]. institutional [691]. instruction [1, 1241, 1531, 2427]. instruction-level [1241]. Int [1881]. Int-Monitor [1881]. intact [193]. integer [627, 685, 1599]. integral [1272]. integrate [1140]. integrated [264, 373, 418, 496, 692, 1041, 1442, 1571, 1605, 1631, 1738, 2047, 2296, 2341, 2343, 2736, 2821, 2871, 2883]. integrates [1386]. Integrating [430, 502, 609, 1643, 2863]. Integration [128, 160, 190, 888, 1183, 1293, 1805, 2113, 2188]. integrative [2645]. integrity [1366, 1663, 2681]. Intel [615, 880, 2149]. intelligence [64, 209, 355, 374, 591, 850, 938, 999, 1007, 1039, 1110, 1170, 1225, 1232, 1382, 1400, 1514, 1709, 1742, 1746, 1756, 1775, 1816, 1933, 2195, 2738, 2784, 2867, 2892, 2897]. intelligence-based [64]. intelligence-enabled [1756]. Intelligent [74, 136, 162, 205, 244, 271, 368, 419, 595, 638, 697, 703, 727, 763, 794, 806, 808, 854, 879, 892, 936, 947, 989, 1009, 1043, 1064, 1081, 1107, 1142, 1198, 1208, 1213, 1220, 1247, 1252, 1291, 1336, 1358, 1397, 1433, 1436, 1443, 1461, 1546, 1607, 1645, 1712, 1714, 1782, 1804, 1858, 1886, 1997, 2009, 2017, 2068, 2174, 2244, 2256, 2344, 2396, 2417, 2480, 2500, 2580, 2620, 2717, 2725, 2773, 2878]. intensity [2075]. intensive [364, 824]. intent [1969]. intention [1959, 2815]. inter [6, 1670, 1809, 1813, 2219]. inter-cloud [6, 1809]. inter-region [2219]. inter-ring [1813]. inter-sentence [1670]. interaction [398, 876, 1076, 1404, 1624, 1775, 2126, 2536]. interaction-coefficient-generators [1076]. interactions [1326, 2225, 2919]. Interactive [42, 109, 459, 979, 1254, 2318, 2385, 2442, 2688, 2899]. interconnect [2415]. interconnected [23]. interconnection [1342, 1644, 1857, 1987, 2070, 2084, 2153, 2220]. Interest [42, 188, 2401, 2688, 2793]. interests [1148]. interface [1087, 1175, 1900]. Interference [16, 43, 139, 1188, 1564, 1793]. interferometers [811]. interferometry [874]. interleaved [806]. interlocutors [2536]. intermediate [1195]. internal [1490]. international [179, 430]. Internet [36, 57, 95, 100, 131, 147, 158, 220, 223, 243, 261, 357, 368, 374, 439, 443, 446, 506, 564, 586, 603, 606, 616, 619, 658, 662, 709, 763, 822, 879, 900, 921, 936, 944, 961, 1004, 1023, 1039, 1041, 1058, 1068, 1081, 1092, 1129, 1170, 1199, 1243, 1244, 1284, 1336, 1350, 1356, 1360, 1369, 1370, 1374, 1386, 1387, 1422, 1433, 1453, 1485, 1563, 1583, 1590, 1645, 1709, 1714, 1724, 1730, 1742, 1780, 1864, 1905, 1909, 1911, 1927, 1939, 1967, 2013, 2037, 2071, 2136, 2169, 2188, 2202, 2231, 2233, 2251, 2252, 2407, 2480, 2524, 2676, 2703, 2706, 2747, 2807, 2808, 2878, 2897, 2904]. Internet [151, 202, 354, 436, 495, 1572, 1717, 2307]. Internet-of-Things [936, 944, 1023, 1905]. Internet-of-Things-aided [147]. Internet-of-Vehicle [1170]. Internet-scale [354]. Interoperability [2885]. Interplanetary [2778]. interpolation [328]. interruptible [1403]. intersection [2001, 2299, 2399]. Interval [698, 789, 1869, 1892, 2077, 2376, 2516]. interval-valued [2516]. Intervention [163, 957]. interventional [818]. interventions [2043]. intima [912]. Intra [128, 2012, 2083, 2221, 2451]. intra-coding [2083, 2221]. intra-node [2451]. Intra-pulse [2012]. intra-routine [128]. introducing [691]. Introduction [2129]. introspection [2189, 2825]. Intrusion [27, 54, 277, 359, 464, 660, 921, 1052, 1393, 1503, 1538, 1706, 1752, 2121, 2182, 2229, 2279, 2280, 2433, 2540, 2565, 2845]. intrusive [2115, 2201, 2311]. intuitionistic [760, 1228, 1229]. intuitive [2225]. invalidity [610]. invariant [146, 739, 743]. inverse [251, 2406]. inversion [613, 2461, 2909]. inverted [1358, 2843]. inverter [1384]. Investigating [569, 837, 1609]. Investigation [163, 1384, 1706, 1739, 2134, 2248, 2310, 2471]. investigations [995]. investment [405, 833]. involving [1913]. IO [1403]. IoBT [2407]. IoE [1439]. IoHT [1468]. IoHT-based [1468]. IoMT [822, 1607, 1655, 1706, 2116, 2383, 2605]. IoMT-based [1655, 2383]. ion [1379, 2367, 2734, 2838]. IoP [1422]. IoRT [1967, 2037]. iOS [523, 2727]. iOS-Android [2727]. IoT [79, 99, 125, 159, 160, 187, 270, 280, 322, 351, 374, 445, 447, 515, 565, 589, 703, 734, 744, 758, 764, 771, 863, 864, 889, 892, 921, 984, 1017, 1037, 1045, 1058, 1091, 1119, 1136, 1144, 1145, 1182, 1200, 1208, 1226, 1232, 1399, 1420, 1457, 1460, 1484, 1487, 1496, 1503, 1529, 1549, 1618, 1651, 1653, 1654, 1677, 1686, 1702, 1733, 1770, 1804, 1808, 1828, 1840, 1873, 1893, 1915, 2023, 2026, 2196, 2200, 2269, 2279, 2292, 2331, 2375, 2409, 2412, 2414, 2435, 2458, 2493, 2547, 2583, 2617, 2634, 2646, 2660, 2665, 2730, 2764, 2804, 2844, 2858, 2883, 2886, 2887]. IoT [265, 268, 313, 328, 341, 638, 647, 845, 846, 872, 1053, 1252, 1397, 1427, 1665, 1695, 1788, 1850, 1954, 2068, 2183, 2241, 2256, 2258, 2306, 2360, 2489, 2595, 2625, 2739, 2751, 2777, 2809]. IoT-based [322, 341, 351, 872, 892, 1397, 2583, 2634, 2646, 2804]. IoT-Cloud [515]. IoT-cloud-enabled [2887]. IoT-edge [2458]. IoT-enabled [2493]. IoT-fog-based [1915]. IoT-fog-cloud-based [1136].

IoT-oriented [1252]. IOTA [2628]. IoTs [814]. IoV [596, 2546, 2713]. IP [1957]. IP-over-WDM [1957]. iPartition [2560]. IPFS [822, 1801, 2848, 2862]. IPR [989]. iris [112, 340, 2574, 2726]. Irregular [1795, 2105]. irregularity [1832]. irrigation [334]. ISAR [2303, 2920]. Island [176, 2319]. Island-based [2319]. isogenies [1526]. isogeny [1287]. isolated [1719, 1810]. isolation [1789, 1993, 2026]. Isometric [2473]. issue [2129]. issues [129, 130, 385, 394, 486, 1053]. ISSWOA [2108]. item [817, 1772]. item-set [1772]. items [787, 2224]. itemset [396, 516, 2072]. itemsets [1357, 1791]. Iteration [358, 1162, 1266]. Iteration-fusing [358]. iterative [174, 196, 278, 356, 379, 524, 1127, 1611, 2132]. IWT [1085].

jackal [2191]. Jacobi [1497, 1611]. jammer [2663]. Jarratt [2780]. Java [2358]. Jaya [986, 1193, 2270, 2413, 2782]. Jaya-sine [2270]. JCF [2587]. JCOGIN [955]. Jinsha [853, 1491]. Job [490, 589, 1054, 1235, 1489, 1721, 2458, 2528, 2590, 2594]. job-shop [1054]. jobs [513]. join [533, 677, 2088]. joining [1249]. joint [19, 71, 556, 866, 1519, 1592, 1652, 1927, 2086, 2350, 2443, 2587, 2663, 2785, 2801]. jointly [731]. JPEG [1171]. Just [241]. Just-in-time [241].

K-DBSCAN [762]. K-NN [882]. KaaS [481]. Kalman [366]. Kapur [2718]. Kashin [866]. Kawasaki [1463]. KAZE [485]. KB [1671]. KC [1373]. KCSS [689]. KD [1554]. KD-tree [1554]. KDB [858]. Kemeny [2167]. Kepler [498]. Kernel [40, 338, 504, 735, 1078, 2038, 2305, 2346, 2371, 2421, 2497, 2516, 2541, 2608]. kernel-induced [338]. kernels [233, 1609, 1807, 2082]. key [258, 259, 349, 596, 616, 652, 740, 796, 828, 864, 1035, 1045, 1096, 1163, 1200, 1208, 1240, 1287, 1375, 1496, 1525, 1526, 1574, 1677, 1855, 1912, 2007, 2053, 2148, 2207, 2242, 2251, 2388, 2399, 2447, 2664, 2667, 2811, 2875]. key-based [2207]. key-dependent [1574]. key-scheduling [2242]. keyphrase [2117]. keypoint [2592]. keypoint-based [2592]. keyword [2532]. KF [366]. KG [2465]. KG-MFEND [2465]. KG2Lib [1767]. Kidney [2421]. kill [1373]. kill-chain [1373]. killer [1310, 1755]. Kinematic [1374]. kinematics [2267]. KITTI [2372]. KMLOD [402]. KMS [2875]. KMS-AMI [2875]. knapsack [1559]. knee [866, 1425]. KNN [2566, 2603]. KNN-based [2603]. KNNGAN [1968]. Knowledge [68, 116, 165, 200, 418, 481, 823, 980, 1061, 1398, 1431, 1438, 1453, 1767, 1972, 2298, 2328, 2465, 2478, 2551, 2695, 2772, 2773, 2776, 2795, 2815, 2859, 2873]. knowledge-as-a-service [481]. knowledge-aware [1061]. Knowledge-based [1438, 2478]. knowledge-graph-based [1767]. Kohonen [67]. kolam [2648, 2722]. Korean [402, 768, 2867]. KPI [144]. krill [227]. Krylov [2477]. KSELM [2421]. KubeGPU [1789]. Kubernetes [689, 1543, 1562, 2166, 2215, 2594]. Kutta [2425].

L [80, 809, 2042]. L-diversity [809]. L-GAN [2042]. L-PowerGraph [80]. label [38, 94, 1251, 1588, 2539]. labeled [2700]. labeling [2113]. laceability [2000]. Ladybug [1898]. Lagging [1573]. Lagrangian [2141]. lake [751, 1955]. LAN [1040]. Land [58, 273, 630, 1757]. land-cover [630]. landmark [602, 2042]. landmark-based [2042]. landowners [1757]. landscape [745]. landslide [2101]. landslides [853, 1491]. lane [1201, 1264, 1515]. language [300, 768, 780, 1284, 1799, 1977, 2545, 2689, 2756]. language-oriented [2689]. LARA [1067].

Large [8, 23, 53, 68, 95, 179, 312, 315, 323, 388, 409, 422, 431, 442, 507, 548, 582, 656, 678, 681, 682, 809, 886, 908, 935, 1064, 1154, 1160, 1196, 1211, 1375, 1377, 1401, 1424, 1445, 1500, 1520, 1524, 1578, 1580, 1646, 1711, 1805, 1819, 2022, 2078, 2177, 2179, 2313, 2371, 2461, 2545, 2633, 2651]. large-capacity [935]. large-kernel [2371]. large-scale [8, 23, 53, 68, 95, 388, 442, 507, 548, 582, 656, 678, 908, 1064, 1154, 1160, 1196, 1377, 1401, 1445, 1500, 1520, 1524, 1578, 1580, 1646, 1711, 1819, 2022, 2177, 2633, 2651]. large-scaled [409]. laser [926]. LASG [2140]. LASG/IAP [2140]. last [393, 1161]. last-level [393]. Latency [135, 184, 488, 714, 1080, 1540, 1552, 1902, 2219, 2316, 2321]. latency-aware [714, 1080]. latency-based [1540]. latency-constrained [2316]. Latency-optimized [2321]. latent [390, 1493, 2154, 2779]. LATOC [1238]. lattice [982]. laundering [2792]. layer [114, 332, 511, 880, 918, 1063, 1237, 1560, 1725, 1853, 2187, 2218, 2271, 2610, 2826]. layer-cross [2218]. Layered [888]. layers [1053]. layout [535, 2234]. layouts [1987, 2084, 2220]. lazy [829]. LB [1905]. LBCNet [2827]. LBM [1456]. LBP [1564]. LBTM [1356]. LCM [2648, 2722]. LDB [1886]. LDCF [1454]. ldiversity [1168]. LDOS [1390]. LDOS-CoMoDa [1390]. LEA [1027, 2602]. Lea-TN [1027]. LEACH [1119]. leader [550, 1027]. leadership [2545]. leadership-class [2545]. leakage [1152, 2829]. leaping [541]. learned [1505, 1872]. learner [1528, 1921]. Learning [17, 31, 94, 121, 174, 274, 278, 315, 351, 420, 428, 434, 449, 480, 493, 527, 553, 606, 611, 612, 619, 629, 630, 662, 675, 682, 721, 791, 814, 833, 896, 930, 931, 944, 987, 1098, 1099, 1105, 1137, 1174, 1193, 1210, 1265, 1319, 1368, 1393, 1395, 1421, 1464, 1468, 1487, 1510, 1575, 1614, 1637, 1689, 1706, 1724, 1825, 1828, 1864, 1926, 1934, 1974, 2009, 2013, 2025, 2082, 2083, 2133, 2135, 2147, 2182, 2188, 2193, 2210, 2221, 2225, 2256, 2274, 2279, 2370, 2383, 2387, 2421, 2493, 2530, 2549, 2665, 2674, 2677, 2692, 2702, 2703, 2792, 2812, 2813, 2844, 2867, 2886]. learning [63, 152, 189, 201, 384, 430, 437, 440, 446, 464, 551, 569, 643, 645, 663, 694, 713, 741, 752, 761, 766, 797, 807, 853, 877, 912, 942, 943, 1041, 1057, 1125, 1143, 1184, 1185, 1198, 1206, 1243, 1254, 1271, 1284, 1302, 1314, 1324, 1349, 1431, 1463, 1491, 1507, 1528, 1536, 1538, 1604, 1607, 1608, 1617, 1624, 1647, 1664, 1711, 1725, 1770, 1820, 1836, 1844, 1871, 1881, 1939, 1941, 1978, 2004, 2020, 2030, 2063, 2079, 2115, 2121, 2131, 2145, 2184, 2197, 2250, 2273, 2294, 2318, 2345, 2368, 2404, 2430, 2442, 2468, 2469, 2488, 2540, 2615, 2651, 2654, 2656, 2663, 2711, 2715, 2774, 2807, 2847, 2884]. learning [34, 144, 206, 353, 477, 505, 547, 615, 647, 669, 672, 687, 728, 780, 782, 800, 817, 826, 846, 850, 879, 884, 897, 906, 919, 946, 947, 954, 965, 978, 985, 1005, 1018, 1023, 1039, 1175, 1192, 1194, 1239, 1255, 1259, 1266, 1278, 1281, 1329, 1355, 1358, 1365, 1386, 1391, 1419, 1427, 1428, 1436, 1448, 1453, 1461, 1523, 1547, 1548, 1583, 1585, 1616, 1663, 1665, 1714, 1759, 1775, 1780, 1814, 1933, 1955, 1963, 2003, 2024, 2027, 2034, 2039, 2044, 2174, 2209, 2306, 2360, 2374, 2418, 2433, 2443, 2447, 2457, 2533, 2555, 2583, 2645, 2652, 2659, 2690, 2725, 2757, 2759, 2760, 2840, 2849, 2870]. learning [429, 818, 820, 1462, 1568, 1595, 1642, 2316, 2357, 2479, 2501, 2512, 2535, 2721, 2912]. learning-based [189, 464, 766, 846, 931, 947, 1143, 1302, 1358, 1436, 1461, 1616, 1711, 1836, 1844, 1864, 2079, 2174, 2193, 2250, 2370, 2433, 2656, 2659, 2663, 2711, 2725, 2812]. learning-enabled [2665]. learning-guided [2131]. lectures [1065, 1066]. LED [2609]. ledger [518, 2766]. Leffler [1240]. left [906]. Legacy [282, 2056]. Legio [1150]. LEGION [255]. Leighton [2745]. leisure [1404]. Lempel [1322]. length [72, 1113, 2096]. lens [2004]. lesion [1463]. less [1351, 1824, 2464, 2571]. lessons [1505]. LETHOR [1351]. letters [1648]. leukemia [1464]. leukocyte [1308]. level [95, 138, 232, 301, 324, 374, 393, 513, 610, 702, 705, 723, 853, 876, 979, 986, 1141, 1241, 1330, 1348, 1350, 1434, 1491, 1532, 1577, 1583, 1606, 1667, 1686, 1907, 1962, 1972, 1975, 2008, 2051, 2142, 2205, 2214, 2236, 2262, 2268, 2386, 2438, 2657, 2837, 2886, 2916]. leveling [1403]. levelling [69]. levels [990, 2224]. leverage [2219]. leverages [2030]. Leveraging [481, 822, 1055, 1684, 1749, 1874, 2112]. Levy [991, 1304, 1305, 2512]. lexicon [1009]. Li [2838]. Li-ion [2838]. libraries [302]. library [838, 1236, 1767, 2189]. LibSVM [1577]. LICOM3 [2140]. LICOM3-CUDA [2140]. LiDAR [791, 1353, 1567, 2372]. life [570, 1941, 2218, 2490, 2734]. Lifelong [1664, 2574]. lifestyle [1253]. lifetime [333, 859, 1379, 1558, 1817, 1888, 2751]. lifting [2276]. light [169, 697, 1203, 1239, 1793, 1858, 2059, 2341, 2450, 2655, 2665, 2788, 2881]. light-weight [2788]. light-weighted [2341]. LightGBM [1011, 1483]. lighting [1443]. Lightweight [80, 190, 305, 534, 565, 584, 658, 676, 700, 836, 845, 883, 900, 1297, 1356, 1449, 1569, 1677, 1912, 2046, 2052, 2159, 2235, 2252, 2292, 2388, 2478, 2480, 2484, 2492, 2504, 2684, 2691, 2766, 2774, 2827, 2850, 2918]. Like [477, 621, 860]. likely [1146]. limited [1542, 2810]. line [123, 841, 1849, 1885, 2699]. Linear [94, 175, 206, 358, 685, 766, 871, 1013, 1113, 1297, 1566, 2179, 2208, 2332, 2589, 2789, 2910]. lines [2699]. Linguistic [1276, 2395, 2799]. link [15, 229, 1027, 1725, 2145, 2272, 2707, 2859]. link-based [229]. link-elimination [15]. link-fault-tolerant [2272]. linked [402]. linking [68, 523, 1983]. links [2345, 2633]. LINPACK [1038]. Linux [380]. list [118, 410, 516, 911, 1732]. list-based [410, 911]. listen [2564]. listen-and-talk-based [2564]. lists [3, 276]. Lite [2620]. literacy [611]. Literature [129, 130, 169, 591, 679, 965, 1306, 1643, 1668, 1815, 1882, 2017]. lithium [2367, 2734]. lithium-ion [2367, 2734]. live [283, 929, 1777]. liver [818, 1402, 1465, 2652]. living [647]. LMAEB [470]. LMAEB-CNN [470]. LOA [1021]. Load [11, 76, 207, 256, 309, 320, 371, 494, 509, 543, 550, 562, 629, 749, 816, 826, 856, 941, 1000, 1011, 1016, 1077, 1182, 1213, 1217, 1218, 1227, 1238, 1341, 1383, 1426, 1498, 1571, 1581, 1807, 1809, 1821, 1905, 2055, 2088, 2115, 2124, 2291, 2311, 2375, 2527, 2557, 2614, 2748, 2760, 2820, 2880]. Load-balanced [562, 2088]. load-balancer [816, 1581]. load-balancing [2614]. load-based [550]. LOADng [1397]. LOADng-AT [1397]. loads [1384]. Local [164, 244, 436, 739, 745, 975, 1133, 1508, 1509, 1593, 1975, 2205, 2301, 2501, 2511, 2668, 2752, 2787, 2828, 2894]. local-global [1975]. Locality [959, 1067]. Locality-aware [1067]. Localization [438, 546, 639, 1034, 1245, 1353, 1429, 1492, 1567, 1782, 1980, 2022, 2066, 2468, 2531, 2592, 2782]. locally [2332]. locate [1680]. located [2368, 2375]. locating [1152, 1563]. Location [304, 363, 438, 757, 997, 998, 1401, 1612, 1758, 1811, 2142, 2158, 2237, 2413, 2600, 2685]. Location-based [438, 1612, 2685]. lock [298]. lock-free [298]. LOCP [2321]. LODNU [2159]. log [321, 943, 1955, 2041, 2464, 2505, 2590]. log-log [2041]. logarithm [627]. logarithmic [1833]. logging [570]. Logic [33, 415, 607, 1074, 1155, 1335, 1411, 1661, 1790, 1822, 1839, 1879, 1910, 1919, 1937, 2217, 2554, 2577, 2588, 2817]. logic-based [607, 1937]. logical [1957]. logistic [1166, 2648, 2722, 2855]. logistics [1043, 1208, 1487, 2892]. logs [1721]. Long [135, 221, 435, 477, 760, 926, 1001, 1142, 1257, 1515, 1758, 1784, 1811, 1852, 1967, 2037, 2105, 2463, 2655, 2773, 2784]. Long- [2463]. long-latency [135]. long-term [1142, 1967, 2037]. long-vector [1515]. longest [1578]. longwave [706]. lookahead [1544]. loop [1911]. loops [2262]. Lop [1494]. LoRa [1684, 2418]. Loss [444, 672, 696, 1779, 1812, 1885, 2116, 2223, 2785]. Loss-based [1779, 1812]. lotus [2602]. Louvain [755, 1424]. Low [42, 97, 326, 383, 439, 488, 521, 553, 656, 696, 699, 709, 774, 781, 798, 799, 899, 954, 1138, 1348, 1533, 1552, 1594, 1902, 1959, 2020, 2165, 2285, 2422, 2689, 2853]. low-area [2853]. Low-cohesion [439]. Low-complexity [326]. low-cost [781, 798, 799]. Low-latency [488, 1552, 1902]. low-level [1348]. Low-overhead [2422]. low-power [383, 521, 774, 899, 2165, 2285]. low-rank [42, 699, 1533]. low-resource [1959, 2689]. Lower [683, 2667]. Lp [1533]. Lp-norm [1533]. LSB [2492]. LSH [11]. LSH-based [11]. LSSVM [2305]. LSTM [67, 646, 797, 918, 941, 946, 1046, 1108, 1320, 1367, 1393, 1426, 1483, 1520, 1657, 1932, 1963, 2138, 2430, 2665, 2696]. LSTM-CNN-grid [1046]. LSTM-RNN-based [797]. LSTM-SN [2138]. LSTM/GAN [946]. LSTMs [432]. LTE [410, 1840]. LTMA [2237]. LTS [2767]. Lucy [1917]. LuNA [887]. lung [329, 338, 820, 830, 1128, 2642]. Lustre [7, 995, 2065]. Lustre-file-system-based [7]. LWD [2909]. LWE [1663]. Lyapunov [2141]. lymph [686].

M [450, 665, 1660]. M-ANFIS [450]. M-tree [1660]. M/M [665]. MAA [2206]. MAC [872, 2564]. Mach [811]. Machine [29, 30, 33, 63, 121, 144, 152, 201, 208, 227, 265, 315, 351, 389, 414, 420, 434, 441, 464, 469, 499, 505, 569, 612, 619, 645, 742, 779, 800, 875, 912, 965, 978, 1025, 1028, 1039, 1190, 1314, 1349, 1368, 1391, 1421, 1427, 1436, 1510, 1538, 1568, 1607, 1642, 1682, 1703, 1706, 1844, 1858, 1863, 1959, 1989, 2024, 2025, 2030, 2034, 2039, 2133, 2135, 2174, 2223, 2250, 2254, 2322, 2357, 2361, 2370, 2383, 2421, 2441, 2445, 2468, 2479, 2497, 2606, 2659, 2674, 2677, 2689, 2703, 2792, 2807, 2824, 2825]. machine-learning-based [420]. machinery [654]. Machines [382, 397, 548, 604, 949, 960, 1080, 1467, 1566, 2106]. Macro [1970]. macular [1324]. MADM [1900]. MAERI [1086]. Magas [1270]. MAGDM [2799]. magnetic [818, 1192, 1713]. magnitude [1726]. Mahalanobis [10]. mail [57, 101]. main [1139, 1545]. maintenance [369, 775, 831, 1782, 1929, 1941, 2077]. major [2466]. majority [127, 511, 579, 837, 1347, 1686, 2331, 2448]. makers [1913]. makespan [1835]. making [198, 521, 742, 1032, 1267, 1527, 1776, 1831, 1913, 1999, 2107, 2139, 2228, 2343, 2526, 2777]. Malatya [2513]. Malaysia [710]. Malek [1908]. Malicious [391, 702, 1601, 1936]. malleability [26, 128]. Malware [540, 834, 1220, 2358, 2666, 2825]. mammalian [186]. mammographic [2079]. MAN [1846]. management [81, 147, 179, 228, 258, 259, 267, 321, 389, 399, 410, 481, 507, 515, 610, 651, 694, 841, 861, 873, 878, 889, 933, 990, 993, 996, 1103, 1189, 1191, 1209, 1300, 1350, 1356, 1422, 1535, 1539, 1601, 1633, 1638, 1659, 1676, 1678, 1738, 1744, 1778, 1789, 1821, 1824, 1850, 1855, 1936, 1996, 2053, 2070, 2143, 2148, 2212, 2251, 2288, 2359, 2390, 2405, 2452, 2493, 2518, 2739, 2811, 2836, 2866, 2875, 2909]. management-based [933]. manager [279, 2319]. MANET [1381, 1692, 2172]. maneuvering [2303]. manifold [2540]. Manipulating [1334]. manipulation [1545]. manufacture [1399]. manufacturing [194, 311, 318, 703, 863, 1213, 1247, 1555, 1645, 1844, 2176, 2224, 2838]. many [483, 576, 650, 699, 880, 907, 1069, 1236, 1394, 1854, 2064, 2213, 2283, 2661, 2876]. many-core [576, 650, 699, 880, 907, 1069, 1236, 1394, 1854, 2064, 2213, 2283, 2876]. many-field [2661]. manycore [360, 451, 1470, 2733]. map [217, 828, 1240, 1669, 1988, 2015, 2648, 2658, 2722]. map-based [828]. MAPA [2214]. MAPE [1952]. MAPE-K [1952]. Mapping [15, 475, 541, 623, 1086, 1140, 1353, 1567, 1598, 1957, 1979, 2185, 2336, 2599, 2911]. Mapping-based [2911]. MapReduce [170, 291, 370, 378, 529, 538, 558, 785, 866, 1250]. MapReduce-based [1250]. maps [761, 1525, 1647, 2412, 2579]. marine [2019]. maritime [263]. marked [1876]. marker [1464]. markers [1210]. market [88, 92, 1816, 1996, 2405, 2871]. marketing [1939]. markets [430]. marking [211, 346, 1201, 1974]. Markov [389, 398, 572, 714, 763, 869, 1041, 1191, 1377, 1411, 1500, 2239]. Markov-process [1191]. Marvell [650]. Mashing [1807]. mashup [2114]. mask [2113, 2675, 2684, 2740]. masks [1745]. mass [186]. mass-based [186]. Massive [794, 802, 1822, 2714, 2876]. massively [431, 812, 2060]. match [1537]. matches [2761]. Matching [75, 314, 910, 969, 1105, 1130, 1518, 1781, 2012, 2061, 2118, 2162, 2177, 2445, 2550]. material [897]. Math [768, 1236]. math-to-speech [768]. Mathematical [1072, 2398, 2487]. MATLAB [2357]. MATLAB/Simulink [2357]. matrices [891, 1120, 1471, 1541, 2461]. Matrix [90, 127, 347, 390, 461, 473, 653, 954, 1134, 1270, 1364, 1394, 1471, 1533, 1588, 1827, 1926, 1958, 2082, 2507, 2541]. matrix-based [1270]. Matrix-product [1364]. matrix-vector [461]. MAVs [1408]. maximal [865, 1751, 1791, 1883]. Maximization [239, 401, 733, 1861, 2690, 2810]. maximize [1817]. maximizing [1888]. maximum [859, 1031, 1051, 1113, 1490, 1586, 2078, 2271]. mayfly [1994, 2092, 2512]. MBitCuts [138]. MC [16]. MCB [2288]. McCulloch [607]. MCDM [655]. MCI [1814]. MCS [2631]. mDesk [50]. MDP [2746]. MDS [564]. MDSbSP [564]. Mean [5, 164, 643, 1127, 1454, 1991, 2006, 2271, 2799, 2889]. Mean-Shift [5]. means [728, 733, 981, 1168, 1250, 1298, 2028, 2172, 2173, 2329, 2351, 2453, 2467, 2636, 2643, 2870]. measurable [2734]. Measure [220, 1791, 2119, 2735]. measurement [676, 839, 906, 1071, 1126, 1506]. measurements [179]. measures [52, 107, 149]. Measuring [155, 912, 1934]. MEC [740, 1153, 1300, 1316, 1659, 2160, 2419]. MEC-enabled [740, 2419]. mechanical [1048, 2755]. mechanism [19, 80, 153, 184, 187, 211, 336, 425, 528, 600, 646, 745, 781, 937, 1117, 1164, 1237, 1249, 1283, 1394, 1551, 1650, 1710, 1753, 1895, 1953, 1966, 2004, 2033, 2190, 2204, 2324, 2344, 2414, 2494, 2523, 2583, 2670, 2673, 2729, 2818, 2823, 2834, 2874, 2897]. mechanisms [101, 988, 1789, 1996, 2510, 2713]. MECInOT [2231]. media [58, 189, 912, 916, 973, 1009, 1251, 1257, 1334, 1365, 1775, 2095, 2474, 2506, 2638, 2770, 2847, 2897, 2911]. median [1359, 1418, 1877]. Medical [111, 214, 347, 384, 402, 606, 661, 662, 801, 822, 844, 981, 1091, 1141, 1425, 1614, 1630, 1875, 2235, 2297, 2480, 2492, 2532, 2612, 2676, 2719, 2744, 2747, 2783]. medicine [2739]. meets [2167]. mellitus [912]. melliuts [2583]. membrane [1740, 2692]. memetic [2362]. memories [1739]. Memory [8, 9, 132, 139, 221, 345, 435, 468, 477, 644, 722, 760, 793, 910, 1001, 1013, 1026, 1067, 1069, 1114, 1139, 1257, 1270, 1303, 1480, 1494, 1545, 1559, 1691, 1735, 1774, 1784, 1795, 1806, 1824, 1852, 2199, 2204, 2427, 2655, 2709, 2771, 2773, 2784]. memory-access [1067]. memory-centric [1026]. memory-efficient [2199]. memory-fully [1852]. MEMS [186]. mental [2363]. merge [190]. merged [2142]. merging [2880]. Mesh [126, 264, 360, 1329, 1396, 2064, 2621]. mesh-oriented [126]. meshes [1113, 2326]. message [45, 1087, 1574, 1873, 1884, 1965, 2081, 2518]. messages [1854, 2424]. messaging [675]. MESSB [1663]. MESSB-LWE [1663]. meta [339, 1295, 1382, 2050, 2227, 2312, 2578]. meta-heuristic [1382, 2050, 2227, 2312, 2578]. meta-heuristic-based [1295]. meta-heuristics [339]. metabolic [587]. metadata [182, 810, 860, 2618, 2629]. metadata-driven [810]. metaheuristic [156, 271, 549, 1125, 2248, 2310, 2562, 2579, 2687, 2802]. metaheuristic-based [156, 549, 2248]. metaheuristics [1935]. Metapath [1517, 2232]. Metapath-guided [1517, 2232]. metastasis [686]. metaverse [1868, 2134, 2632, 2897]. meteorological [2659]. metering [828, 1352, 2875]. Method [32, 93, 103, 127, 224, 226, 261, 266, 267, 311, 317, 362, 391, 398, 403, 447, 493, 502, 512, 592, 604, 623, 640, 688, 720, 760, 815, 831, 844, 885, 952, 1013, 1072, 1140, 1149, 1275, 1298, 1321, 1360, 1456, 1460, 1467, 1485, 1528, 1548, 1553, 1563, 1564, 1570, 1602, 1621, 1636, 1640, 1750, 1752, 1758, 1811, 1820, 1826, 1827, 1885, 1894, 1929, 1930, 1962, 2023, 2049, 2054, 2117, 2163, 2177, 2180, 2209, 2225, 2262, 2267, 2303, 2339, 2346, 2347, 2356, 2359, 2366, 2391, 2402, 2457, 2464, 2495, 2596, 2609, 2636, 2644, 2647, 2669, 2699, 2714, 2747, 2790, 2843, 2855, 2859, 2871, 2896, 2918]. method [9, 61, 119, 133, 176, 183, 633, 646, 659, 774, 857, 981, 982, 1171, 1315, 1323, 1396, 1415, 1693, 1755, 1792, 1819, 1883, 1938, 2154, 2296, 2305, 2315, 2329, 2575, 2666, 2668, 2683, 2764, 2791]. methodology [325, 484, 894, 1214, 1338, 1415, 1822, 1945, 2554, 2738, 2769, 2862]. methods [134, 286, 344, 468, 524, 612, 622, 630, 655, 834, 937, 997, 998, 1109, 1162, 1193, 1405, 1421, 1548, 1667, 2071, 2101, 2103, 2250, 2311, 2326, 2343, 2383, 2507, 2707, 2810]. methods-based [2103]. methods-image [524]. methylated [620]. metric [682, 772, 1708, 2309]. Mexican [568]. MFCINet [2657]. MFEND [2465]. MFMDet [2740]. MFQ [2246]. MI [1667]. MI-based [1667]. Micali [2745]. Micro [696, 862, 1302, 1719, 1810, 2082, 2122, 2517, 2800]. Micro-batch [2122]. micro-data [862]. Micro-expression [2517]. micro-grid [1719, 1810]. Micro-kernels [2082]. micro-ring [696]. micro-services [1302]. microarray [166, 331, 1154, 1765, 2255, 2439]. microblog [470]. Microblogs [1116]. microchannel [186]. microcontrollers [459]. microgrid [290, 2192, 2562]. microphones [1745]. microphysics [2444]. microrheology [441]. microscopic [1098]. microseismic [1693]. microservice [2212]. microservices [2069]. microservices-based [2069]. microwave [1465]. Mid [941]. Mid-term [941]. middleware [1017, 1788]. Migrating [625, 1015, 2088]. migration [7, 208, 262, 411, 543, 726, 902, 1489, 1633, 1777, 1856, 1989, 1995, 2135, 2144, 2486]. military [1703, 1999]. Mille [793]. millimeter [1602]. million [1764]. MILP [2236]. MIMO [210, 1188, 1313, 1602, 1686, 1793, 2130]. MIMO-MISO [1686]. MIMO-OFDM [1188]. mine [1693]. miner [2914]. mineral [2334]. miners [1761]. MinHash [1636]. Mini [554]. Mini-App [554]. Minimal [960, 1883, 2718]. minimization [108, 342, 785, 1295, 1716, 2752]. minimized [1454]. minimum [495, 497, 730, 1256, 1475, 2208, 2513]. minimum-cost [497]. Mining [67, 118, 151, 155, 200, 269, 277, 353, 396, 495, 516, 753, 795, 800, 815, 819, 898, 965, 1226, 1265, 1271, 1279, 1357, 1409, 1412, 1448, 1572, 1671, 1717, 1744, 1766, 1772, 1773, 2039, 2072, 2101, 2201, 2259, 2471, 2485, 2529]. mining-induced [1572]. Minkowski [558]. MinLA [2234]. minus [199]. misbehavior [1343]. misclassification [2563, 2584]. misinformation [2119]. MISO [1686]. MISR [2318, 2442]. missing [328, 1649, 2543, 2622]. mission [43, 1597, 1941, 2304, 2431]. mission-critical [2431]. misuse [2374]. mitigate [1680]. Mitigating [2320, 2902]. mitigation [18, 1022, 1055, 1662, 2168, 2639]. Mittag [1240]. Mittag-Leffler [1240]. mix [1846]. mixed [45, 870, 1288, 1439, 1599, 2355]. mixed-criticality [1288]. mixed-integer [1599]. mixing [874, 2738]. mixture [637, 702]. ML [299, 2227]. ML-based [2227]. MLDDP [1214]. MLEsIDSs [464]. MLIR [1565]. MLTM [610]. MM* [1309]. mmWave [2130]. MN [1188]. MN-based [1188]. mnemonics [2427]. MO [196]. Mobi [2431]. Mobi-Sense [2431]. Mobile [4, 22, 41, 48, 124, 223, 313, 363, 404, 714, 726, 732, 756, 769, 869, 962, 963, 994, 999, 1047, 1083, 1122, 1163, 1219, 1287, 1297, 1301, 1337, 1340, 1374, 1472, 1477, 1479, 1512, 1554, 1569, 1618, 1649, 1666, 1675, 1679, 1966, 1976, 1982, 1992, 1995, 2041, 2076, 2092, 2237, 2289, 2290, 2317, 2388, 2413, 2423, 2454, 2650, 2715, 2797]. Mobility [410, 453, 1301, 1512, 1666, 2431, 2488, 2694]. mobility-aware [2431, 2488]. MOBRO [2781]. modal [1315, 2074]. modalities [2126]. modality [2416]. mode [663, 1571, 1638, 1762, 1859, 2053, 2192, 2504, 2867, 2894]. Model [31, 65, 88, 108, 114, 153, 172, 235, 249, 309, 329, 378, 380, 457, 479, 559, 606, 665, 714, 721, 763, 823, 833, 900, 903, 913, 914, 924, 934, 1038, 1041, 1044, 1060, 1072, 1168, 1184, 1185, 1251, 1253, 1325, 1350, 1406, 1439, 1519, 1614, 1628, 1629, 1645, 1723, 1728, 1754, 1797, 1818, 1825, 1851, 1884, 1893, 1908, 1926, 1932, 1945, 2020, 2033, 2041, 2049, 2067, 2086, 2135, 2144, 2197, 2212, 2239, 2261, 2263, 2269, 2274, 2275, 2279, 2286, 2300, 2319, 2393, 2395, 2420, 2494, 2591, 2673, 2692, 2702, 2711, 2717, 2727, 2762, 2773, 2790, 2804, 2828, 2832, 2838, 2844, 2858, 2893]. model [60, 142, 154, 173, 175, 187, 194, 236, 282, 296, 306, 334, 335, 389, 470, 487, 491, 503, 555, 607, 614, 618, 653, 663, 673, 834, 879, 902, 918, 920, 943, 956, 985, 989, 1011, 1048, 1050, 1161, 1162, 1181, 1188, 1233, 1276, 1284, 1345, 1368, 1373, 1374, 1398, 1426, 1436, 1455, 1456, 1458, 1483, 1500, 1608, 1617, 1624, 1666, 1668, 1734, 1746, 1814, 1881, 1917, 1948, 1977, 2017, 2024, 2094, 2107, 2115, 2131, 2136, 2153, 2174, 2176, 2230, 2236, 2318, 2354, 2408, 2418, 2429, 2442, 2445, 2453, 2508, 2542, 2548, 2583, 2600, 2623, 2629, 2670, 2706, 2725, 2764, 2766, 2847, 2899]. model [105, 110, 435, 466, 477, 634, 687, 706, 717, 1267, 1289, 1309, 1494, 1573, 1595, 1601, 1700, 1708, 1843, 1950, 2046, 2100, 2140, 2258, 2298, 2357, 2380, 2400, 2443, 2465, 2471, 2482, 2543, 2622, 2625, 2684, 2772, 2868, 2894, 2912]. Model-based [618, 714, 2263]. model-free [2263]. model-oriented [2298]. model-view-controller [2300]. Modeling [4, 123, 149, 179, 433, 461, 475, 621, 702, 722, 731, 871, 881, 1123, 1158, 1414, 1682, 1764, 1876, 1983, 1992, 2069, 2305, 2327, 2398, 2438, 2475, 2728, 2767, 2819, 2877]. modelling [192, 1498, 2154, 2231, 2459]. modelling-based [2154]. models [78, 107, 115, 143, 387, 537, 578, 587, 893, 1023, 1145, 1146, 1314, 1331, 1400, 1561, 1594, 1843, 2044, 2145, 2250, 2545, 2677, 2713, 2738, 2746]. Modern [85, 90, 501, 838, 1006, 1469]. Modified [280, 450, 672, 771, 884, 1402, 1442, 1566, 1631, 1808, 1949, 2172, 2325, 2546, 2572, 2581, 2596, 2648, 2722, 2735]. modular [660, 712, 1865]. modularization [20]. Modulation [1313, 1902, 2012]. module [2447, 2599, 2608, 2829, 2865]. modules [362, 1919]. MOEA [2626]. MOGW [748]. Moka [2298]. Moka-ADA [2298]. moldable [513]. molecular [3, 276, 554, 733, 1383, 1464, 1807]. momental [2437]. moments [743, 1557]. momentum [2371]. momentum-contrast [2371]. monaural [424]. money [2792]. monitor [638, 1881]. Monitoring [67, 106, 113, 187, 205, 263, 315, 740, 741, 853, 872, 976, 984, 1107, 1136, 1147, 1491, 1522, 1702, 1741, 2005, 2115, 2189, 2311, 2369, 2400, 2489, 2755, 2818, 2878, 2887]. monitoring-based [1147, 2818]. Monkeypox [2369]. monocular [1669, 2450]. monolithic [1973, 2036]. Monte [320]. MOONLIT [2371]. MOSFETs [2284]. Mosques [852]. moth [482]. motif [855, 1261]. Motion [5, 113, 225, 463, 1225, 1336, 1408, 1991, 2149, 2693]. motor [854]. mould [1282, 1715]. mountain [1246]. mounted [2368, 2691]. movement [333, 1703, 2044]. movements [1175, 1434, 1583]. moving [106, 251, 426, 668, 2075, 2789]. MPI [126, 358, 717, 1087, 1132, 1150, 1222, 1494, 2058, 2189, 2451, 2522, 2791]. MPI-dot2dot [1222]. MPI/OpenMP [126]. MPIO [2382]. MPSoC [601, 1180]. MPSoC-based [1180]. MQCA [2073]. MQTT [2883]. MR [170, 807, 1566, 2125, 2884]. MRAM [1545]. MRAN [2090]. MRI [451, 1523, 1814, 1858, 2051, 2596, 2654]. MRI-based [1814, 2654]. MRIoEF [1439]. MRIPN [2267]. MRTensorCube [406]. MSSAMTO [2546]. MSSAMTO-IoV [2546]. much [1872]. Multi [8, 22, 32, 33, 38, 123, 198, 204, 227, 250, 307, 330, 343, 372, 374, 398, 426, 481, 497, 562, 610, 630, 660, 702, 736, 835, 882, 928, 951, 986, 1017, 1020, 1024, 1032, 1106, 1117, 1133, 1156, 1177, 1248, 1251, 1278, 1330, 1350, 1378, 1379, 1421, 1423, 1443, 1445, 1478, 1501, 1502, 1559, 1586, 1588, 1602, 1633, 1712, 1715, 1719, 1720, 1730, 1796, 1798, 1950, 1956, 1962, 1974, 1993, 2064, 2074, 2085, 2097, 2122, 2175, 2180, 2181, 2187, 2195, 2233, 2265, 2271, 2303, 2323, 2343, 2481, 2530, 2543, 2546, 2568, 2613, 2693, 2700, 2705, 2740, 2781, 2796, 2813, 2822, 2910]. multi [12, 20, 34, 40, 51, 54, 71, 94, 148, 161, 273, 327, 376, 392, 450, 471, 504, 544, 552, 592, 625, 677, 695, 698, 763, 765, 776, 801, 888, 933, 1018, 1051, 1063, 1088, 1125, 1176, 1240, 1269, 1300, 1333, 1341, 1430, 1431, 1450, 1484, 1535, 1537, 1551, 1601, 1606, 1626, 1663, 1725, 1776, 1780, 1794, 1809, 1817, 1821, 1827, 1835, 1900, 2003, 2012, 2017, 2018, 2038, 2062, 2106, 2107, 2115, 2131, 2206, 2214, 2231, 2238, 2254, 2281, 2299, 2371, 2438, 2441, 2445, 2457, 2469, 2496, 2514, 2534, 2571, 2600, 2647, 2651, 2657, 2681, 2691, 2729, 2741, 2748, 2751, 2785, 2815, 2820, 2899]. multi [83, 110, 202, 254, 299, 466, 525, 530, 723, 1267, 1323, 1415, 1597, 1691, 1810, 1847, 1899, 2100, 2128, 2146, 2152, 2307, 2334, 2465, 2552, 2622, 2698, 2724]. multi-access [1423, 1633, 2175, 2231]. multi-agent [161, 254, 343, 1017, 1267, 1431, 2003, 2651]. Multi-agent-based [1719, 1810]. multi-answer [2534]. multi-application [1626]. multi-architectural [660]. multi-aspects [2214]. Multi-attribute [882]. multi-attribute-based [2796]. multi-authority [1117, 2334]. Multi-Bearer [307]. MULTI-BSP [299]. multi-channel [202, 562, 1827, 1899, 2307]. multi-class [1559, 2018]. multi-client [2299]. multi-cloud [12, 273, 933, 1535, 1809, 2438, 2613]. multi-coloring [1796]. multi-coloring-based [698]. Multi-constraints [1378]. multi-container [765]. Multi-context [1586]. multi-controller [1421, 2748, 2820]. Multi-core [8, 592, 723, 776, 1177, 1430, 1445, 1602, 2064, 2180, 2181]. multi-cores [1330, 1333, 2122]. Multi-criteria [327, 552, 1032, 1450, 1776, 2107, 2343]. multi-device [148, 392]. multi-dimensional [372, 801, 951, 1051, 1300, 1821]. multi-disease [450]. multi-DNN [1993]. multi-domain [2100, 2457, 2465, 2469]. multi-exaflops [2106]. multi-extractable [1663]. Multi-factor [1443]. multi-feature [2012]. multi-fine-grained [2371]. multi-fog [2613]. multi-follower [2438]. multi-FPGA [544]. multi-frequency [1323]. multi-functional [471]. multi-gait [2496]. Multi-GPU [398, 1501, 1691, 1950, 2062]. multi-granular [2238]. Multi-granularity [497, 2741]. multi-graph [2552]. Multi-grid [1379]. Multi-head [1606, 2543, 2622, 2899]. Multi-helicopter [198]. multi-hidden [763]. Multi-hop [123, 330, 2146, 2546]. multi-horizon [2097]. multi-instance [1597]. multi-intention [2815]. multi-kernel [40, 504, 2038]. Multi-label [38, 94, 1251, 1588]. Multi-layer [1063, 1725, 2187, 2271]. Multi-level [374, 610, 702, 986, 1350, 1962, 2657]. multi-malicious [1601]. multi-match [1537]. multi-mobile [22]. Multi-modal [2074]. multi-model [2445]. multi-munition [530]. multi-node [2131]. Multi-objective [33, 34, 227, 625, 677, 835, 1018, 1020, 1088, 1156, 1176, 1248, 1269, 1478, 1712, 1715, 1798, 1817, 1900, 1956, 2085, 2128, 2195, 2206, 2254, 2281, 2441, 2481, 2514, 2530, 2568, 2647, 2751, 2781, 2910]. multi-parametric [2233]. multi-party [2323, 2600]. Multi-period [736]. multi-person [250]. multi-port [2152]. multi-processor [552]. multi-programming [20]. multi-queue [1974]. multi-replica [2681]. multi-resource [83]. Multi-scale [71, 204, 1341, 1551, 1720, 1847, 2691, 2693, 2724, 2740, 2785, 2822]. Multi-sensor [54, 110, 1278]. Multi-sensor-based [426]. multi-server [1240, 1794, 2571]. multi-source [1780]. multi-sourced [888]. Multi-spectral [630]. multi-stage [2700, 2729]. multi-step-ahead [928]. multi-strategy [1502]. multi-strategy-enhanced [2698]. multi-target [51, 2303]. multi-task [376, 695, 2115]. Multi-threshold [2705]. multi-tier [1730]. multi-track [2017]. multi-tracker [1125]. multi-UAV [525, 1133]. multi-UAVs [1415]. multi-user [273, 481, 1712]. multi-verse [1106, 1125, 1484]. Multi-view [32, 466]. multi-wave [1725]. multi-workflow [1835]. multibusiness [405]. multicast [1435, 2337]. multichannel [545, 1639, 2541]. multiclass [966, 2535]. Multicore [451, 614, 974, 1158, 1161, 1178, 1222, 1322, 1639, 1690, 2011, 2047, 2151, 2164]. multidevice [2288]. multidimensional [602, 1906]. multifactor [41, 2608, 2914]. Multifactorial [2004]. multifeature [2521]. multifunctional [1141]. multigrid [1109]. multilayer [662, 821, 1214, 2708]. Multilevel [576, 783, 916, 925, 1304, 1305, 1475, 1860, 1884, 2297, 2425, 2428, 2739, 2754, 2876]. Multilingual [2506]. Multimedia [151, 253, 353, 1378, 1697]. Multimodal [430, 479, 938, 1173, 1231, 2102, 2255, 2439, 2629, 2695, 2783]. multiobjective [1732, 2247, 2695]. MultiOff [1618]. multiperspective [2459]. multiphase [633, 2422]. Multiple [175, 211, 289, 307, 472, 477, 495, 634, 662, 706, 837, 977, 1083, 1296, 1346, 1406, 1584, 1618, 1736, 1788, 1815, 1854, 1874, 1886, 1961, 1993, 2126, 2318, 2332, 2424, 2442, 2460, 2627, 2686, 2745]. multiple-domain [1788]. multiple-response [1961]. multiplexer [326, 536, 579, 1131, 1362]. multiplication [90, 712, 747, 954, 1134, 1471, 1958, 2082, 2576]. multiplicative [2049, 2395]. multiplier [1089, 1891]. multipliers [811, 2701]. multiply [2203]. multipole [576, 2876]. multipopulation [2695]. multiprocessing [234]. multiprocessor [471, 917, 971, 1288, 1380, 1908]. multiprocessors [810, 861, 2278]. multiprojection [468]. multipublisher [1966]. multiresolution [1925]. Multiresource [2379]. multiscale [946, 1518, 1764, 2511]. Multisensor [113]. Multisource [197]. multistage [1595]. Multistep [861, 946, 1894, 2711]. multistep-ahead [946, 2711]. multitask [479, 1675, 1966]. multithreaded [1069]. multithreading [1055, 1115]. multitype [308]. multiuser [197]. Multivariate [1409, 1657, 1852, 1895, 2389, 2598, 2619, 2868]. MultiView [1367]. multivocal [1882]. munition [530]. murmuration [2712]. music [424, 459, 545, 1387, 1775, 1780, 2017, 2322, 2784]. mutation [64, 745, 1001, 1193, 1392, 2861]. mutation-enabled [2861]. mutual [700, 1035, 1062, 1233, 1274, 1651, 1701, 1912, 2292, 2889]. MUX [2422]. MV32 [1726]. MVDLSTM [1367]. MWPSO [1454]. myeloid [1464]. myocardial [1146].

Nacc [2774]. Nacc-Guard [2774]. NADE [292]. Nadu [256]. NAEEMD [1693]. Naive [725, 2031]. name [324]. Named [60, 1275, 1668, 2296, 2356, 2401, 2503, 2700, 2744]. NAND [512, 579, 1074]. nanites [782]. nano [2588]. nano-scale [2588]. nanocomparator [1157]. nanocomputing [1362]. nanofluid [2738]. nanomagnetic [2554]. nanoscale [97, 704, 2283, 2284]. nanotechnology [579]. narrative [2674]. NAS [2110]. nasopharyngeal [37]. national [833, 2185]. Natural [283, 780, 982, 1310, 1596, 1674, 1755, 2104, 2492]. nature [636, 2602]. nature-inspired [2602]. Navier [1469]. Navigation [593, 944, 1411]. NB [2489]. NBA [1653]. NBSVM [2426]. ND [2104]. ND-S [2104]. NEAE [2808]. Near [910, 926, 974, 1471, 1636]. near-duplicate [1636]. Near-infrared [926]. Near-optimal [974]. near-sparse [1471]. nearest [313, 602, 1660, 2315]. NEC [851]. neck [1265]. needed [691]. needle [2758]. needles [401]. needs [2185]. needy [2157]. needy-students [2157]. NEEF [350]. negation [688]. negative [390, 961, 1827, 2759]. negotiation [142]. neighbor [3, 276, 313, 602, 739, 1596, 1660, 1674, 2048, 2104, 2315, 2498]. Neighborhood [589, 682, 1186, 1732, 2247, 2859]. Nek5000 [1196]. nephropathy [1194]. NER [104]. nested [2700]. NestMSA [472]. net [475, 1128, 1402, 1521, 1975, 2046]. netflow [1955]. NetGuru [17]. Network [17, 55, 65, 88, 114, 199, 204, 210, 258, 259, 365, 370, 398, 444, 449, 452, 488, 517, 541, 559, 632, 660, 662, 696, 702, 758, 795, 863, 934, 983, 1002, 1010, 1027, 1057, 1126, 1201, 1263, 1307, 1320, 1376, 1393, 1460, 1478, 1495, 1552, 1558, 1566, 1587, 1657, 1685, 1706, 1720, 1722, 1737, 1748, 1752, 1768, 1784, 1797, 1825, 1862, 1889, 1960, 1972, 1975, 2042, 2121, 2148, 2150, 2158, 2159, 2182, 2187, 2215, 2243, 2256, 2280, 2286, 2297, 2331, 2358, 2390, 2394, 2476, 2486, 2507, 2536, 2561, 2563, 2582, 2591, 2608, 2661, 2665, 2675, 2693, 2727, 2773, 2786, 2794, 2819, 2832, 2850]. network [37, 71, 92, 157, 185, 186, 216, 272, 275, 307, 333, 345, 350, 352, 410, 460, 497, 546, 561, 639, 643, 686, 713, 753, 803, 832, 847, 855, 875, 935, 943, 967, 997, 998, 1009, 1046, 1048, 1092, 1104, 1198, 1204, 1220, 1241, 1244, 1268, 1274, 1295, 1297, 1301, 1343, 1374, 1399, 1406, 1408, 1517, 1538, 1546, 1555, 1626, 1630, 1644, 1679, 1716, 1734, 1738, 1758, 1778, 1832, 1900, 1901, 1911, 1912, 1987, 2012, 2017, 2027, 2062, 2067, 2084, 2091, 2099, 2106, 2216, 2220, 2260, 2337, 2354, 2403, 2431, 2457, 2469, 2506, 2551, 2562, 2577, 2638, 2642, 2649, 2664, 2689, 2696, 2715, 2719, 2723, 2726, 2761, 2788, 2815]. network [93, 110, 113, 164, 165, 202, 248, 277, 320, 321, 342, 359, 435, 572, 577, 646, 669, 672, 673, 813, 1025, 1058, 1063, 1121, 1135, 1230, 1261, 1272, 1296, 1315, 1364, 1387, 1411, 1419, 1428, 1434, 1443, 1572, 1585, 1592, 1598, 1599, 1613, 1642, 1670, 1678, 1680, 1708, 1767, 1782, 1803, 1811, 1868, 1888, 1899, 1933, 1944, 1964, 2046, 2111, 2114, 2136, 2138, 2146, 2152, 2266, 2307, 2312, 2325, 2416, 2433, 2482, 2490, 2515, 2535, 2552, 2599, 2657, 2678, 2720, 2729, 2743, 2765, 2776, 2782, 2789, 2827, 2837, 2868, 2874, 2877, 2892, 2899]. network-based [93, 272, 546, 686, 934, 1241, 1393, 1555, 1679, 1722, 2661, 2868]. network-constrained [2789]. Network-enabled [210]. network-on-chip [2599, 2794]. network-on-chips [1626, 2187]. networking [19, 1191, 1275, 1678, 2356]. Networks [22, 44, 66, 95, 123, 184, 221, 314, 336, 375, 388, 408, 415, 453, 461, 518, 543, 575, 602, 658, 740, 862, 869, 899, 904, 962, 1006, 1008, 1035, 1037, 1108, 1147, 1148, 1164, 1215, 1249, 1352, 1377, 1393, 1421, 1472, 1497, 1504, 1520, 1524, 1558, 1578, 1582, 1620, 1621, 1660, 1661, 1684, 1686, 1687, 1796, 1818, 1828, 1829, 1831, 1852, 1864, 1876, 1921, 1957, 1970, 1971, 1980, 1998, 2016, 2049, 2061, 2116, 2133, 2142, 2194, 2200, 2229, 2294, 2317, 2345, 2366, 2401, 2414, 2454, 2463, 2530, 2549, 2564, 2598, 2601, 2607, 2660, 2663, 2685, 2692, 2749, 2796, 2810, 2845, 2895]. networks [70, 124, 219, 257, 271, 325, 332, 346, 349, 385, 404, 419, 424, 494, 501, 555, 560, 600, 605, 645, 671, 709, 736, 754, 756, 873, 878, 895, 908, 929, 1022, 1029, 1032, 1034, 1056, 1100, 1116, 1187, 1245, 1261, 1269, 1397, 1405, 1429, 1436, 1450, 1492, 1501, 1540, 1550, 1565, 1579, 1591, 1606, 1616, 1646, 1666, 1725, 1800, 1837, 1890, 1900, 1937, 1976, 2000, 2026, 2066, 2070, 2093, 2099, 2153, 2198, 2216, 2269, 2270, 2272, 2290, 2398, 2413, 2430, 2437, 2460, 2489, 2495, 2585, 2597, 2609, 2704, 2751, 2753, 2778, 2784, 2785, 2803, 2814, 2833, 2915]. networks [54, 133, 158, 363, 416, 463, 530, 681, 684, 796, 859, 963, 1019, 1296, 1311, 1312, 1342, 1410, 1673, 1700, 1708, 1713, 1817, 1888, 2055, 2120, 2156, 2174, 2293, 2392, 2419, 2640, 2763, 2821, 2836, 2881, 2891, 2912]. networks-on-chip [501, 1056, 1296, 2016]. Neural [37, 65, 92, 157, 185, 221, 272, 325, 345, 352, 365, 370, 415, 419, 424, 449, 460, 461, 494, 518, 546, 559, 561, 643, 645, 660, 662, 803, 832, 847, 875, 884, 934, 1002, 1006, 1009, 1029, 1046, 1164, 1201, 1204, 1215, 1241, 1244, 1263, 1297, 1307, 1320, 1374, 1399, 1419, 1478, 1501, 1546, 1550, 1565, 1566, 1579, 1598, 1630, 1685, 1716, 1722, 1734, 1737, 1752, 1768, 1828, 1831, 1837, 1911, 1921, 2049, 2062, 2150, 2198, 2286, 2337, 2354, 2358, 2437, 2475, 2506, 2507, 2515, 2530, 2536, 2551, 2563, 2582, 2591, 2642, 2649, 2661, 2723, 2726, 2727, 2761, 2773, 2784, 2786, 2815, 2919]. neural [113, 133, 164, 320, 435, 572, 669, 671, 673, 1063, 1230, 1272, 1315, 1364, 1434, 1443, 1572, 1599, 1713, 1800, 1868, 1899, 1948, 2046, 2136, 2293, 2433, 2490, 2635, 2753, 2763, 2776, 2821, 2833, 2869, 2874]. neuro [256, 450, 621, 2202]. neuro-fuzzy [256, 450, 621, 2202]. neurodegenerative [1175]. NeuroEvolution [2808]. neurological [840]. neuromarketing [2071]. Neuron [607, 1086, 2049]. neuropeptide [820]. neutron [2087, 2302]. neutrons [581]. neutrosophic [69, 149, 961, 1778]. new-generation [2454, 2586]. newborns [1585]. news [73, 561, 2465, 2629, 2686]. next [817, 1540, 2239, 2830]. next-item [817]. NextG [1902]. NFV [862, 1845]. NFVLearn [2389]. NIASHPT [368]. NIDS [1537]. NIR [193]. NIST [712]. Nizar [2687]. NKA [1310]. NLFSR [2504]. NLP [2587]. NLR [1161]. NLR-OP [1161]. NMF [1613, 2709, 2771]. NMR [2909]. nn [882, 1385, 1846]. NNAPI [903]. NNBlocks [848]. NNEF [903]. NOC [475, 1097, 1973, 2036, 2337, 2424, 2452]. NoC-based [2337, 2452]. NoCs [623, 1351, 1870, 1942]. Node [41, 54, 119, 505, 668, 685, 686, 712, 924, 1027, 1501, 1613, 1708, 1899, 1936, 1957, 2066, 2131, 2272, 2413, 2451, 2514, 2531, 2549, 2577, 2731, 2895]. node- [2272]. Node-importance [1708]. NodeRank [1148]. nodes [167, 292, 349, 605, 668, 754, 796, 893, 1148, 1258, 1492, 1646, 2170, 2270, 2366, 2745, 2883]. nodule [71, 1128, 2785]. nodules [46, 820]. Noise [28, 140, 1886, 2596]. Noise-insensitive [1886]. noisy [530, 827, 1904]. NOMA [2364]. Non [167, 193, 279, 340, 390, 496, 668, 691, 696, 814, 859, 906, 1031, 1110, 1584, 1587, 1667, 1792, 1827, 2088, 2115, 2262, 2311, 2326, 2362, 2495, 2522, 2842]. Non-blocking [696, 1587]. non-compaction [906]. non-contiguous [2522]. Non-cooperative [340, 1584, 2495]. Non-destructive [193]. non-English [496]. non-ICT [691]. non-inertial [814]. non-intrusive [2115, 2311]. non-manager [279]. non-maximum [1031]. non-MI-based [1667]. non-migrating [2088]. non-negative [390, 1827]. non-slicing [2362]. non-stationary [2842]. non-tree [859]. non-uniform [167, 668, 1792, 2262, 2326]. non-viscous [1110]. nondominated [2626]. nonferrous [1102]. Nonlinear [52, 1162, 1384, 2610, 2846, 2855]. nonlocal [1298]. nonnegative [2541]. Nonparametric [306]. nonstationary [1762]. nonuniform [952]. nonvolatile [1189]. norm [1533]. normal [1412]. normalization [635]. north [1161, 2293]. North-Last [1161]. northeast [2293]. northwest [2293]. NoSQL [188, 356, 411, 2285, 2569]. NoSQL-based [2569]. notch [2896]. Note [84, 117, 141, 171, 191, 230, 240, 284, 293, 303, 310, 400, 1988– 1992, 2038, 2039, 2176, 2306, 2307]. notes [2261]. notifications [2070]. Novel [96, 104, 135, 161, 287, 331, 350, 352, 365, 368, 398, 478, 511, 512, 516, 527, 686, 690, 704, 708, 724, 766, 828, 845, 877, 918, 938, 966, 969, 996, 1020, 1037, 1046, 1082, 1111, 1136, 1217, 1273, 1328, 1331, 1335, 1352, 1368, 1397, 1400, 1457, 1613, 1653, 1668, 1690, 1718, 1726, 1737, 1787, 1816, 1820, 1841, 1860, 1870, 1913, 1942, 2019, 2035, 2052, 2073, 2144, 2217, 2268, 2276, 2362, 2395, 2448, 2507, 2516, 2530, 2537, 2544, 2568, 2581–2583, 2612, 2648, 2660, 2662, 2687, 2692, 2697, 2708, 2722, 2737, 2780, 2786, 2832, 2882, 2887, 2893, 2904]. novel [354, 470, 672, 745, 977, 1112, 1209, 1218, 1486, 1537, 1736, 1803, 1839, 1919, 2155, 2356, 2357, 2417, 2490, 2580, 2629, 2655, 2666]. novo [1603]. NP [641]. NP-completeness [641]. NPI [168]. NPR [843]. NRSM [2523]. NRWalk2Vec [2638]. NRWalk2Vec-HIN [2638]. Nscale [1856]. NSGA [1798, 2647]. nuclear [1533]. NUMA [2320]. number [390, 432, 574, 649, 865, 886, 952, 1089, 1266, 1833, 2139, 2242, 2728, 2735, 2913]. number-based [1266]. numbers [548, 1276, 2310]. numerical [282, 372, 478, 482, 489, 781, 940, 1818, 2015, 2326, 2355, 2512, 2586]. Nussbaumer [649]. NVIDIA [954].

O} [723, 949, 995, 1113, 1628, 1629, 2065]. O}-shaped [1113]. OBD [2479]. Object [204, 480, 535, 672, 780, 1057, 1060, 1232, 1446, 1576, 1594, 1874, 2006, 2075, 2159, 2385, 2533, 2539, 2657, 2743, 2752, 2850, 2861]. object-oriented [1446]. Objective [33, 34, 227, 625, 677, 835, 1018, 1020, 1088, 1156, 1176, 1223, 1248, 1269, 1478, 1712, 1715, 1798, 1817, 1823, 1900, 1956, 2085, 2128, 2195, 2206, 2254, 2265, 2281, 2441, 2481, 2514, 2530, 2568, 2647, 2751, 2781, 2800, 2813, 2910]. objectives [232, 1730]. objects [106, 426, 466, 1057, 2789]. objects-Based [1057]. OBL [1212]. OBL-based [1212]. oblivious [1079]. obstacle [1031, 1949]. obstacle-based [1949]. occurrence [172, 1120]. ocean [2140]. OCR [1505]. OCR-Vx [1505]. octree [1469]. octree-based [1469]. OFDM [218, 1188, 1793]. off [573, 1683, 1802, 2240, 2834]. off-chain [2834]. off-peak [1683]. offloading [158, 714, 716, 734, 769, 847, 994, 1047, 1122, 1126, 1144, 1337, 1421, 1446, 1512, 1616, 1618, 1679, 1712, 1753, 1828, 1878, 1927, 2076, 2086, 2160, 2289, 2350, 2419, 2524, 2546, 2553, 2595, 2597, 2653, 2793]. offs [1891, 2147]. OFNE [275]. OFP [1345]. OFP-TM [1345]. OG-RADL [816]. OHUQI [1357]. oil [2829]. OKCM [752]. old [153]. older [2436]. olive [1836]. OM2M [680]. omni [1504]. omni-path [1504]. omnidirectional [1433]. OmpSs [358]. OMRES [1808]. on-chip [1450, 1857]. on-demand [31, 108]. On-Fly-TOD [528]. on-shelf [1357]. on-the-fly [2211]. One [787, 1473, 1482, 1747, 1786, 1982, 2294, 2385, 2448]. one-bit [2448]. one-class [787]. one-click [2385]. one-dimensional [1473]. One-shot [2294]. Online [193, 222, 250, 271, 388, 752, 1065, 1066, 1261, 1333, 1345, 1367, 1652, 1710, 1856, 1894, 2067, 2119, 2277, 2404, 2711, 2832, 2867, 2912]. ontology [527, 888, 1851, 2618]. ontology-based [888, 1851]. OP [1161].

OP2 [2775]. open [129, 130, 402, 477, 744, 1020, 2112, 2632]. OpenACC [2095]. OpenCL [301, 735, 812, 1456, 1658, 2149, 2903]. Openflow [2375]. OpenGL [412]. OpenMP [126, 361, 716, 717, 1158, 1330, 2058, 2095, 2653]. OpenMP-based [361]. OpenWhisk [1258]. Operation [198, 420, 653, 729, 2192]. operations [3, 135, 276, 573, 1934, 2558]. operator [64, 1001, 1476, 1542, 2183, 2247]. operators [1450]. Opinion [275, 345, 1825, 1956, 2443]. Opportunistic [503, 562, 2454]. opportunities [6, 744, 1114]. opposite [1536]. opposition [814, 931, 1125, 1759, 2425, 2512]. opposition-based [814, 1125, 1759, 2425, 2512]. oppositional [1420]. OPSO [1238]. Optical [316, 332, 575, 696, 736, 811, 839, 868, 1059, 1074, 1161, 1611, 1946, 1957, 2187, 2826]. Optimal [40, 81, 138, 212, 253, 330, 343, 488, 580, 585, 728, 842, 869, 879, 916, 931, 974, 997, 998, 1056, 1102, 1122, 1188, 1190, 1213, 1320, 1340, 1355, 1417, 1542, 1552, 1568, 1570, 1692, 1712, 1715, 1748, 1761, 1858, 1880, 2016, 2038, 2063, 2234, 2243, 2262, 2270, 2312, 2413, 2506, 2520, 2549, 2662, 2779, 2832, 2846, 2870]. Optimisation [25, 559, 2738]. Optimization [14, 44, 126, 149, 162, 174, 208, 221, 231, 235, 343, 362, 372, 397, 420, 478, 482, 608, 639, 748, 771, 814, 913, 914, 983, 1016, 1020, 1034, 1093, 1132, 1149, 1173, 1176, 1197, 1204, 1212, 1221, 1236, 1248, 1257, 1260, 1282, 1318, 1392, 1442, 1466, 1478, 1503, 1559, 1627, 1653, 1661, 1678, 1682, 1699, 1748, 1762, 1820, 1849, 1851, 1909, 1923, 1928, 1935, 1943, 1979, 1984, 1989, 2021, 2041, 2045, 2144, 2193, 2227, 2243, 2327, 2342, 2361, 2373, 2391, 2421, 2435, 2458, 2481, 2530, 2568, 2602, 2636, 2653, 2687, 2694, 2712, 2718, 2727, 2731, 2737, 2780, 2797, 2813, 2882, 2910]. optimization [43, 47, 63, 71, 168, 245, 281, 337, 369, 384, 454, 479, 499, 568, 625, 717, 732, 754, 782, 857, 887, 915, 959, 986, 994, 1018, 1021, 1048, 1069, 1077, 1125, 1194, 1201, 1227, 1273, 1277, 1294, 1316, 1365, 1368, 1381, 1429, 1452, 1492, 1547, 1550, 1599, 1607, 1635, 1679, 1711, 1755, 1759, 1798, 1817, 1819, 1824, 1826, 1866, 1878, 1898, 1900, 1949, 1994, 2004, 2015, 2050, 2092, 2108, 2155, 2172, 2184, 2191, 2199, 2254, 2270, 2281, 2325, 2350, 2355, 2400, 2405, 2441, 2453, 2512, 2520, 2531, 2586, 2614, 2631, 2647, 2690, 2695, 2698, 2699, 2705, 2755, 2765, 2791, 2798, 2843, 2846, 2906]. Optimization-based [221, 1227, 1257, 1381, 1762, 2172, 2421, 2458, 2631, 2780]. optimization-driven [1607]. optimization-enabled [2765]. optimization-support [499]. optimize [1841, 1945]. Optimized [12, 133, 226, 261, 320, 442, 586, 704, 950, 985, 1009, 1085, 1180, 1262, 1363, 1372, 1393, 1430, 1726, 1759, 1808, 1809, 1840, 1850, 1902, 2022, 2066, 2096, 2136, 2286, 2321, 2538, 2578, 2642, 2655, 2676, 2689, 2716, 2842, 2856, 2877]. optimizer [1016, 1054, 1106, 1125, 1282, 1291, 1420, 1452, 1484, 1536, 1711, 1866, 1904, 1929, 1997, 2050, 2193, 2265, 2405, 2425, 2502, 2712, 2754, 2781]. optimizer-based [1420, 1866, 1904]. optimizers [2579]. Optimizing [7, 327, 862, 1015, 1160, 1233, 1382, 1705, 1856]. optimum [2562]. OptiSembleForecasting [2631]. orbit [1249]. orchestration [18, 1960]. orchestrator [924]. order [317, 1059, 1261, 1367, 1396, 1429, 1625, 2171, 2750, 2775]. order-picking [317]. order-shuffling [2171]. ordering [122, 181]. orderly [1388]. orders [75, 2485]. ordinal [1984]. ordinary [2400, 2475]. organ [930]. organic [852, 2659]. organisms [867, 1212, 2204]. organized [761]. organizing [1279]. oriented [126, 656, 657, 774, 1252, 1446, 1564, 1803, 1909, 2298, 2339, 2474, 2689, 2765]. ORSCA [1482]. ORSCA-GPU [1482]. orthogonal [641, 1557]. orthogonalization [87]. orthopair [2799]. oscillations [2879]. OSGAN [2294]. osmotic [2553]. OSN [500]. OSNs [2629]. osteoarthritis [1425]. out-of-core [2199]. out-of-memory [2709, 2771]. outage [2364]. outdoor [1438]. outlier [640, 1593, 1754, 2389]. outlook [573]. output [2826]. outputs [1089]. outsourced [2323, 2566]. outsourcing [1641, 2171]. outspread [1136]. Ovarian [1820]. over-current [1748]. over-sampling [10, 920]. overall [816]. overhead [29, 526, 1467, 1901, 2422, 2562]. overhead/cable [2562]. overheads [520, 1806]. overlapped [415]. overlapping [882, 975, 2649]. oversampling [1968, 2104, 2154]. overview [1537, 1907, 2383]. overwriting [1388]. ozone [791].

P [884, 1672]. P-AdaBoost-PAUC [1672]. P2P [363, 375, 404, 2266]. P4 [2216]. P4-based [2216]. PAARes [2170]. paced [2917]. pack [2838]. package [831, 2240]. package-aware [2240]. packages [102]. Packet [138, 211, 252, 346, 368, 667, 723, 858, 1537, 1739, 2054, 2291, 2375, 2661, 2863]. packet-based [1739]. packet-in [2375]. packets [433]. packing [1821]. padding [1051]. page [2466, 2916]. page-level [2916]. PageRank [529, 2763]. paging [410]. pain [1265]. pair [1233]. paired [1869]. pairing [2662]. pairs [1206]. pancake [560]. pandemic [1273, 1325, 1746]. panel [1135]. panoramic [1353, 1567]. PanoVILD [1353, 1567]. paradigm [19, 1539, 2185, 2431]. paradigms [451]. paralinguistic [434]. Parallel [56, 60, 74, 88, 126, 145, 192, 233, 234, 295, 302, 364, 372, 384, 398, 441, 458, 474, 489, 497, 502, 504, 517, 522, 524, 538, 545, 582, 625, 629, 649, 675, 679, 701, 717, 729, 752, 776, 790, 812, 830, 866, 871, 874, 887, 913, 932, 951, 958, 964, 986, 1049, 1130, 1177, 1204, 1222, 1246, 1248, 1260, 1313, 1322, 1377, 1401, 1407, 1432, 1489, 1519, 1580, 1611, 1623, 1634, 1639, 1699, 1815, 1822, 1837, 1854, 1947, 2011, 2045, 2058, 2095, 2109, 2110, 2124, 2151, 2179, 2213, 2353, 2391, 2462, 2476, 2541, 2586, 2593, 2669, 2745]. parallel [39, 68, 80, 180, 282, 555, 633, 716, 811, 914, 915, 940, 1082, 1150, 1171, 1256, 1379, 1508, 1509, 1946, 1948, 2060, 2118, 2132, 2461, 2525, 2610, 2733, 2826]. parallel-hybrid [874]. parallel-streaming [1948]. Parallelisation [2228]. parallelism [232, 294, 716, 1241, 1330, 1542, 2083, 2137, 2221]. Parallelization [86, 361, 576, 761, 1006, 1383, 1938, 2876]. parallelize [1087, 2064]. Parallelized [583]. parallelizing [282]. paralyzed [1468]. parameter [251, 1132, 1568, 1721, 1817, 2327, 2420, 2758]. parameter-based [2758]. Parameters [7, 1556, 1599, 2303, 2460, 2478, 2637, 2659, 2666]. parametric [2233]. ParaView [1196]. parcel [1757]. PARCSIM [1699]. parental [1459]. parenthesizing [1256]. Pareto [397]. Parity [97, 1015]. Parity-preserving [97]. parks [892]. parsimony [498, 2857]. parsing [1166, 1965, 2008]. Parsisanj [1440]. partial [594, 1109, 1586, 1697, 2326]. participants [153]. Particle [51, 149, 174, 454, 499, 814, 955, 983, 1227, 1762, 2281, 2361, 2695, 2731, 2882]. particles [1764]. partition [640, 1792, 2278]. partitioned [1288, 2124, 2323]. partitioning [15, 266, 769, 959, 988, 1036, 1195, 1516, 1652, 1747, 2164, 2560, 2731]. parts [597, 1734]. party [1385, 1767, 2207, 2323, 2600]. passenger [687]. passing [1087, 1965]. passive [438, 1449]. password [1035, 2447, 2571]. password-authenticated [2447]. patch [267]. Path [278, 332, 479, 544, 583, 681, 782, 789, 1097, 1101, 1219, 1246, 1281, 1504, 1534, 1587, 1627, 1798, 1867, 1869, 2092, 2337, 2338, 2411, 2477, 2481, 2573, 2580, 2620, 2672, 2758, 2892, 2900]. path-based [583]. pathogen [1310]. pathological [2884]. paths [1019]. patient [603, 740, 984, 1548, 2261]. patients [840, 1146, 1194, 1422, 1425, 1464, 1468]. pattern [13, 118, 214, 269, 422, 495, 739, 795, 882, 1102, 1130, 1579, 1773, 1781, 2201, 2223, 2485, 2694, 2860]. pattern-based [739]. patterned [2724]. patterns [118, 302, 323, 425, 693, 759, 1529, 1830, 1952, 2905]. PAUC [1672]. pavement [572]. payment [670, 1163]. PC [2528]. PCA [427, 1390, 2028, 2173, 2323, 2351, 2631]. PCA-based [2631]. PCA-PSO-K [2028]. PCB [1733]. PCGC [2428]. PCM [1013, 1139]. PCMIgr [2054]. PCNN [2851]. PCP [2841]. PCP-ACO [2841]. PCR [800, 2039]. PCVM.ARIMA [604]. PDC [1097]. PE [2818]. peak [1, 767, 1531, 1683, 1922]. peaks [391, 2104, 2787, 2880]. pear [193]. Pearson [749]. pedestrian [274, 597, 1918, 2496, 2552]. peer [518, 2025]. peer-to-peer [518, 2025]. pelvic [930]. penalty [302]. pendulum [1358]. pentanary [1632]. People [768, 999, 1169, 1990]. People-centric [999]. PEPS [776]. Per-user [2608]. perceiving [2669]. perceptron [662, 1063, 1214, 2187, 2708]. percutaneous [2758]. Perfect [137]. PerficientCloudSim [678]. performability [1643]. Performance [4, 17, 21, 62, 87, 107, 123, 124, 135, 145, 168, 250, 360, 364, 378, 394, 425, 461, 487, 488, 491, 553, 560, 569, 571, 573, 601, 615, 648, 657, 679, 722, 746, 752, 754, 765, 792, 793, 816, 838, 851, 880, 902, 923, 950, 960, 976, 983, 994, 1002, 1008, 1038, 1157, 1158, 1160, 1178, 1189, 1236, 1241, 1253, 1280, 1380, 1399, 1467, 1474, 1490, 1495, 1500, 1513, 1562, 1594, 1602, 1617, 1739, 1860, 1912, 1945, 1987, 1992, 1993, 2011, 2062, 2065, 2069, 2084, 2106, 2110, 2147, 2160, 2187, 2220, 2264, 2266, 2283, 2310, 2428, 2454, 2467, 2548, 2569, 2767, 2775]. performance [212, 218, 229, 292, 300, 405, 475, 572, 578, 597, 671, 772, 794, 806, 854, 995, 1077, 1134, 1161, 1175, 1203, 1256, 1384, 1398, 1403, 1419, 1523, 1530, 1543, 1646, 1750, 1763, 1865, 2035, 2070, 2136, 2143, 2288, 2398, 2510, 2525, 2586, 2856]. performance- [657]. Performance-aware [983, 1189]. performance-based [816]. Performance-energy [2147]. Performance-enhanced [1253]. performances [1127, 2363]. performed [2263]. period [736, 2068]. periodic [2909]. peripheral [2415]. permissioned [881]. permissionless [1916]. permutation [967, 1193, 2402]. persistence [1777, 2392]. persistent [1373, 2001]. Person [65, 250, 664, 1186, 1965, 2274]. person-post [65]. persona [2328]. personal [148, 289, 357, 392, 2510]. personality [2239]. Personalized [1024, 1453, 2328, 2795, 2830]. persons [2188, 2436]. perspective [163, 1389, 2093, 2746]. PERT [1253]. PERT-HAR [1253]. perturbation [484, 898, 1162]. pervasive [727]. pesticides [2374]. PesViT [2374]. Petri [475]. Petri-net-based [475]. petroleum [2893]. petrophysical [2909]. PG [1215]. PG-RNN [1215]. phase [55, 685, 687, 874, 1132, 1456, 1913, 2359, 2848, 2920]. phased [2544]. phases [2455, 2852]. phases-workflows [2455]. phenomenon [2633]. Phi [615, 880]. Philippine [2172]. phonocardiogram [667]. photo [1897]. photographic [2786]. photonic [501, 692, 696]. photovoltaic [2325]. Physical [119, 120, 136, 332, 465, 467, 539, 621, 727, 1107, 1120, 1131, 1605, 1952, 2235, 2400, 2571, 2767, 2892]. physics [2112]. picking [317]. picture [2349]. piecewise [175]. PII [1391]. pilot [1334]. pinball [2223]. pinhole [1277, 2355]. pipeline [812, 2550]. pISRA [105]. pitch [424]. Pitts [607]. Pix2Pix [1501]. pixel [122, 2075]. PKEET [2344]. place [1394]. placement [30, 31, 227, 354, 373, 742, 862, 875, 924, 983, 1005, 1028, 1032, 1083, 1144, 1176, 1190, 1213, 1269, 1275, 1296, 1339, 1447, 1540, 1592, 1845, 1909, 1971, 2254, 2258, 2342, 2441, 2578, 2606, 2613, 2797, 2802]. placing [865]. plagiarism [2587]. plan [1173]. plane [1296, 1723]. planning [43, 181, 202, 334, 487, 736, 782, 881, 997, 998, 1101, 1177, 1219, 1534, 1719, 1724, 1798, 1810, 1929, 2092, 2144, 2258, 2307, 2411, 2554, 2573, 2580, 2620, 2672, 2758, 2892]. plans [1194]. plant [2529, 2702]. platelets [1764]. platform [17, 21, 102, 234, 264, 274, 574, 649, 680, 697, 751, 933, 935, 1090, 1258, 1513, 1530, 1560, 1565, 2157, 2188, 2190, 2404, 2613, 2689]. platforms [233, 367, 378, 553, 569, 922, 1115, 1283, 1288, 1437, 1584, 1634, 1805, 1823, 2011, 2212, 2506]. play [1441]. Player [1395, 1742, 2396]. PLC [192, 263]. PLC-based [263]. PLEACH [2857]. Pleasure [1754]. Pleasure-arousal-outlier [1754]. Plus [2836]. Plus-profile [2836]. PM [1318]. PMC [1309, 2153]. PMLAP [2769]. PMSM [1451]. pneumonia [2570, 2635, 2869]. POI [1406]. Point [431, 1058, 1140, 1195, 1401, 2077, 2204, 2500, 2616, 2793, 2822, 2898]. point-based [1195]. point-of-interest [2793]. points [175, 546]. Poisson [2284]. polarity [496]. police [1527]. policies [974, 2505]. policy [386, 869, 1118, 1339, 2021, 2690]. Pollard [627]. pollination [2183]. pollutant [710]. pollution [102]. polyglot [1777]. polygon [1867]. polyhedral [1094]. polymorphic [782]. Polynomial [956, 1792, 2589]. polynomials [1405, 1541, 2537]. POPS [1683]. popularity [1753, 2674, 2770]. popularity-aware [1753]. population [2362, 2874]. populations [2163]. porous [2095]. port [1056, 1587, 2152]. portability [2467]. portable [2082, 2181, 2525]. portfolio [417, 2645]. portraits [2795]. Pose [270, 918, 1104, 1329, 2044, 2156, 2683, 2788]. position [426, 1215, 1233, 1648, 2214, 2429, 2549, 2623]. position-based [2429, 2623]. position-gated [1215]. positioning [892, 1278, 2068, 2368]. positions [2368]. positive [800, 2039, 2792]. POSNet [2274]. possible [581]. possibly [691]. post [65, 1526, 2399, 2558, 2745, 2786, 2867]. post-disaster [2558]. post-epidemic [2867]. post-photographic [2786]. post-quantum [1526, 2399, 2745]. posture [1254]. potassium [2250]. potential [176, 1252, 2650]. Power [59, 123, 277, 290, 383, 416, 425, 465, 521, 637, 709, 746, 774, 873, 899, 1095, 1131, 1152, 1192, 1296, 1371, 1484, 1524, 1571, 1609, 1748, 1885, 1907, 2165, 2197, 2285, 2312, 2520, 2529, 2554, 2799, 2844, 2880]. power-aware [465, 1095]. power-line [123]. power-saving [425]. powered [844, 1479, 2739, 2836]. powerful [1052]. PowerGraph [80]. powerlist [296]. powerlist-based [296]. PP [2770]. PPF [2246]. PPP [153]. PPT [1158]. PPT-Multicore [1158]. PQM [342]. practical [1397, 2007, 2315]. practices [841]. pragmatic [2468]. Pre [16, 1324, 1977, 2225, 2297, 2545, 2838]. pre-coding [16]. pre-identify [2838]. pre-processing [2225]. Pre-trained [1324, 1977, 2297]. pre-training [2545]. precise [1486, 2364]. precision [99, 870, 954, 1151, 1401, 2913]. precomputing [1683]. Predator [433, 1162]. Predator-prey [433]. predators [2019]. predict [74, 195, 348, 863, 2043]. predicted [287]. Predicting [345, 571, 663, 978, 1136, 1251, 1263, 1319, 1426, 1575, 1721, 2013, 2239, 2345, 2590, 2708, 2738, 2884]. Prediction [10, 21, 67, 92, 164, 241, 315, 329, 334, 338, 366, 435, 477, 514, 553, 572, 634, 637, 679, 725, 757, 861, 876, 900, 905, 912, 928, 941, 946, 981, 1100, 1108, 1158, 1201, 1202, 1226, 1241, 1257, 1314, 1345, 1354, 1367, 1368, 1442, 1483, 1500, 1514, 1520, 1530, 1657, 1666, 1709, 1723, 1725, 1742, 1773, 1818, 1895, 1918, 1932, 1940, 1969, 2009, 2028, 2032, 2044, 2063, 2077, 2101, 2133, 2173, 2203, 2218, 2351, 2354, 2414, 2430, 2446, 2460, 2490, 2494, 2519, 2528, 2552, 2556, 2583, 2619, 2659, 2673, 2707, 2711, 2727, 2728, 2748, 2765, 2770, 2812, 2820, 2833, 2836, 2859, 2871, 2906]. prediction [9, 676, 1279, 1572, 1755, 2357, 2365, 2679, 2688]. Prediction-based [514, 757]. prediction-compression-based [2414]. predictions [686, 1325, 1597]. Predictive [604, 694, 776, 1146, 1258, 1762, 1941, 2619]. predictors [1577]. preemption [1850]. Preface [36, 98]. preference [75, 2720]. prefetch [376, 1139, 2180]. prefetch-aware [376]. prefetching [1114]. prefix [729]. pregnant [1328]. preloading [1477]. preparation [179]. preparedness [2516]. preprocessing [53, 1181]. presbyopia [1479]. prescription [172]. presence [1680]. Presentation [705, 711, 2574]. preservation [1075, 2456]. preserved [2607, 2658]. preserving [2, 97, 266, 279, 285, 304, 347, 843, 845, 898, 948, 1159, 1359, 1375, 1385, 1392, 1612, 1683, 1779, 1812, 1878, 1906, 1922, 1944, 2190, 2194, 2259, 2315, 2323, 2646, 2692, 2834]. prevent [2436]. prevention [611, 1052, 2844]. prey [433, 1162]. prey-predator [1162]. Price [1483, 1995, 2041, 2063, 2293, 2357, 2871]. pricing [108, 1337, 1710, 1880, 1996, 2472]. primary [818, 1465]. prime [712, 2576]. primitives [1494]. principal [281, 945, 2516]. principles [1834]. PriNergy [447]. printing [1714]. prior [2683]. priorities [1205]. prioritization [1716]. Prioritized [1951]. Priority [447, 556, 596, 778, 966, 1051, 1600, 1774]. Priority-based [447, 556]. Privacy [2, 105, 260, 266, 279, 285, 347, 436, 439, 822, 823, 845, 878, 883, 889, 898, 948, 999, 1007, 1064, 1075, 1159, 1168, 1375, 1385, 1392, 1612, 1615, 1683, 1766, 1779, 1812, 1878, 1906, 1922, 2111, 2190, 2259, 2315, 2323, 2344, 2510, 2607, 2625, 2634, 2646, 2658, 2795, 2831, 2834, 2904]. Privacy-aware [2111, 2904]. privacy-enhanced [2634]. privacy-preserved [2658].

Privacy-preserving

[2, 266, 279, 285, 347, 845, 898, 948, 1159, 1375, 1385, 1612, 1683, 1878, 1906, 1922, 2190, 2259, 2315, 2323, 2646, 2834]. privacy-protected [2795]. private [881, 1414, 2501]. Pro [1115]. Proactive [707, 861, 989, 1351, 1689, 2314, 2440]. Probabilistic [538, 602, 651, 682, 1524, 1830, 2317, 2720]. Probability [329, 347, 420, 1493, 1738, 2145, 2364, 2577]. problem [69, 86, 317, 354, 401, 441, 542, 624, 625, 627, 641, 785, 977, 1054, 1082, 1106, 1193, 1235, 1256, 1282, 1484, 1490, 1540, 1559, 1578, 1627, 1637, 1725, 1903, 1946, 1971, 2078, 2182, 2248, 2258, 2281, 2438, 2497, 2502, 2513, 2520, 2782, 2814, 2857, 2920]. problems [482, 497, 599, 610, 867, 884, 938, 986, 1072, 1177, 1212, 1597, 1898, 1984, 2015, 2108, 2144, 2193, 2355, 2406, 2481, 2514, 2695, 2712, 2824, 2846, 2876, 2882]. procedure [1240, 1525]. process [26, 65, 82, 87, 155, 182, 580, 657, 675, 703, 759, 819, 869, 1024, 1152, 1191, 1332, 1414, 1844, 1894, 2170, 2356, 2459, 2634, 2683, 2741, 2838, 2839]. process-based [82]. processes [587, 742, 1731, 2085, 2128, 2907]. processing [8, 210, 246, 247, 313, 356, 370, 404, 414, 490, 532, 589, 707, 732, 755, 780, 793, 812, 851, 910, 1156, 1165, 1167, 1187, 1199, 1205, 1230, 1303, 1316, 1336, 1379, 1385, 1476, 1519, 1542, 1785, 1873, 1906, 1926, 2097, 2122, 2131, 2225, 2275, 2375, 2483, 2560, 2669, 2685, 2701, 2853]. processing-in-memory [1303]. processing-near-memory [910]. processor [387, 513, 552, 573, 576, 650, 731, 1236, 1394, 1602, 1634, 1854, 1908, 2415]. processor-workload [731]. processors [90, 238, 455, 473, 699, 776, 798, 799, 880, 954, 974, 1012, 1161, 1177, 1178, 1189, 1241, 1833, 2147, 2164, 2181, 2283, 2733]. product [245, 461, 841, 1048, 1319, 1364, 1510, 1849, 1939, 1948, 2345, 2633, 2742]. production [1399]. productivity [521, 1495, 1590, 1934, 2764]. products [345]. profile [308, 473, 2836]. profiles [1158]. profiling [786, 2239]. Profit [212]. prognosis [1210]. program [735, 1625, 2405, 2593]. programmability [1178]. programmable [1919]. Programming [20, 126, 235, 295, 297, 299, 301, 448, 685, 717, 785, 874, 955, 1050, 1102, 1566, 1914, 2110, 2487, 2910]. programming-based [1102, 2910]. programs [238, 296, 582, 716, 887]. progression [981]. progressive [1183]. project [153, 179]. projection [1923, 2473]. projective [1570]. projects [69, 1279, 2343]. prolapses [930]. prolonged [2751]. prominently [655]. promotion [101]. prompt [2757]. Proof [980, 1131, 2600]. propagation [114, 717, 1216, 1596, 1674, 1727, 1964, 2408]. properties [1311, 1312, 1829, 2482]. property [855, 2336]. proposal [204, 672, 1848]. proposals [2091]. proposed [143, 224, 559, 2028, 2157, 2173, 2351, 2572]. prospects [1539]. protect [365]. protected [2795]. Protecting [2777]. protection [260, 332, 439, 560, 823, 883, 1064, 1766, 1927, 1930, 2344, 2422, 2747, 2783, 2821, 2831]. protein [47, 215, 498, 1248, 1815]. proteins [195, 721]. protocol [39, 41, 70, 159, 228, 272, 333, 342, 453, 510, 564, 616, 652, 658, 668, 756, 828, 845, 849, 864, 908, 1075, 1200, 1287, 1346, 1369, 1370, 1381, 1397, 1449, 1569, 1655, 1677, 1687, 1701, 1862, 1873, 1916, 1998, 2023, 2094, 2148, 2195, 2241, 2399, 2447, 2571, 2607, 2660, 2664, 2751, 2840, 2895]. protocols [700, 827, 1079, 1472, 1496, 1848, 2081, 2564, 2749]. prototype [1002]. Provable [1794]. Provably [796, 836, 1375, 1526, 1569, 1954, 2809]. Provenance [323]. providers [944, 1321, 2893]. Providing [1163, 1330, 1504]. provision [616, 894]. provisioning [50, 549, 629, 881, 1346, 1631, 1642, 1689, 1840, 2021]. proximal [484, 2021]. proxy [2314, 2440, 2848]. pruning [671, 1956, 2177, 2281, 2321, 2386, 2403, 2563]. PSCLS [1954]. Pseudo [432, 2006]. Pseudo-random [432]. pseudo-supervised [2006]. pseudorandom [952, 2242]. PSO [343, 1021, 1121, 1227, 1320, 1949, 2028, 2173, 2305, 2351, 2361, 2478, 2524, 2544, 2758]. PSO-K [2173, 2351]. PSO-RDAL [1227]. pub [1036]. pub/sub [1036]. public [2, 49, 279, 288, 715, 927, 1272, 1366, 1486, 1825, 1862, 2013, 2508, 2639, 2667, 2805]. Publisher [1567, 1987, 2084, 2441, 2442]. publishing [266, 347]. PUF [2605]. Pulmonary [71]. pulpitis [1462]. Pulse [1566, 2012]. Pulse-Coupled [1566]. Puncalc [294]. puncture [2758]. punished [2787]. puppetmasters [2474]. purchase [2287]. purchasing [1455]. pure [2857]. pursuit [161]. puzzle [2680]. PVO [1359]. PWM [1384]. PyDTNN [896]. Pyramid [204, 480, 2002, 2517, 2691, 2762]. pyramids [1410, 1943]. Python [2110, 2522].

Q [1040, 2690, 2765]. Q&A [1123].

Q-learning-driven [2690]. QBCSSA [917]. QCA [326, 511, 579, 821, 837, 1347, 1362, 1474, 1760, 1877, 1910, 2035, 2089, 2448]. QCA-based [1877]. QCA-multiplexer [579]. QFD [603]. QGA [2286]. QGA-QGCNN [2286]. QGCNN [2286]. QoE [1071, 2493]. QoE-driven [2493]. QoS [19, 280, 481, 894, 1071, 1498, 1549, 1654, 2183, 2679]. QoS-aware [19, 481, 1654, 2183]. QoS/QoE [1071]. QR [2323]. qTESLA [649]. quad [2309]. quad-cube [2309]. quadratic [1560]. quadric [2399]. Quadro [2009]. Quadro-W [2009]. qualitative [2250]. Quality [20, 54, 328, 469, 587, 741, 784, 929, 1316, 1504, 1620, 1802, 1841, 1863, 2230, 2243, 2400, 2435, 2508, 2655, 2840, 2891]. quality-based [784]. quality-of-service [1620]. quantification [46, 311, 520, 2210]. Quantifying [618, 1647, 1832, 1901]. quantitative [132, 1357, 1754, 2767]. quantization [1556]. Quantum [96, 97, 127, 258, 259, 552, 579, 585, 628, 747, 827, 842, 917, 950, 956, 1033, 1065, 1066, 1076, 1111, 1112, 1131, 1157, 1303, 1335, 1526, 1544, 1627, 1685, 1718, 1726, 1839, 1879, 1919, 2030, 2178, 2217, 2242, 2253, 2257, 2286, 2399, 2412, 2538, 2588, 2630, 2703, 2745, 2816, 2817]. quantum-dot [127, 747, 1131, 1303, 1335, 1726, 1839, 2217, 2630, 2816]. quantum-dots [2588]. Quantum-inspired [552, 917]. quantum-to-classical [2538]. Quasi [1420]. quaternary [1073]. Queries [422, 538, 558, 1211, 2335]. Query [39, 53, 246, 247, 327, 404, 732, 1016, 1160, 1260, 1372, 1385, 1472, 1906, 1969, 2566, 2685]. query-driven [1472]. question [188, 688, 2146, 2297, 2534]. question-driven [2534]. questions [1614]. Queue [1500, 1974]. queueing [665, 2266]. queuing [316, 1458]. quotient [1858]. QVCO [1040].

R [102, 1344, 1647, 1884]. R-CNN [1344, 1647]. R2T [992]. R2T-DSDN [992]. races [28, 140]. RAD [2026]. radar [251, 882, 2012]. RadHPO [2184]. radial [2819, 2877]. radiation [522, 1368, 2302]. radiative [706, 1705]. radical [201]. radio [16, 44, 314, 330, 2564]. radio-based [16]. RADL [816, 1581]. RAFALE [1642]. RAID [1015, 1856]. RAID-6 [1015, 1856]. railroad [2728]. railway [833, 2714]. railways [2647]. rainbow [860]. rainfall [1818]. RainFormer [2002]. rainwater [1466]. RAM [1131, 1686, 2816]. Ramanujan [567]. random [76, 195, 338, 432, 725, 790, 827, 1060, 1432, 1442, 1593, 1666, 2012, 2190, 2310, 2780]. randomized [1903, 2179]. randomness [1512]. Range [251, 404, 639, 1245, 2531]. range-free [639, 2531]. Rank [42, 542, 699, 1533, 2026]. ranked [1772]. ranking [73, 1276, 1321, 1708, 2167, 2268, 2575]. ranking-based [2575]. ransomware [2835].

RAPCHI [1655]. rapid [466]. rare [819]. rat [508, 1997]. ratchets [658]. rate [550, 863, 1071, 1545, 1580, 1610, 1745, 1859, 1932, 2203, 2637, 2659, 2688]. ratings [610, 1319]. ratio [2054, 2075, 2211, 2738]. raw [419]. ray [913, 914, 1746, 2570]. ray-tracing [913, 914]. RBF [1495]. rcnn [2621, 2740]. RDAL [1227]. RDBMS [411]. RDF [323, 1671, 1981, 2335]. RDM [1926]. Re [231, 664, 1186, 1965, 2080, 2274, 2314, 2440, 2561, 2848]. re-encryption [2314, 2440, 2848]. Re-engineering [231]. re-identification [664, 1186, 1965, 2274]. re-registration [2080]. re-routing [2561]. REACH [2864]. reactive [804, 805, 1397]. read [292]. readability [1164]. reading [768]. Ready [769]. Ready-time [769]. Real [14, 111, 190, 237, 261, 357, 465, 556, 589, 592, 598, 642, 669, 697, 703, 904, 953, 976, 980, 992, 1071, 1081, 1141, 1187, 1253, 1336, 1437, 1564, 1600, 1735, 1746, 1790, 1892, 1898, 1902, 1934, 1976, 2005, 2059, 2186, 2190, 2249, 2287, 2322, 2333, 2352, 2483, 2539, 2726, 2818, 2827, 2849, 2909]. Real-time [14, 111, 190, 237, 261, 357, 465, 556, 589, 592, 598, 642, 697, 703, 904, 953, 976, 992, 1071, 1081, 1141, 1187, 1253, 1336, 1437, 1564, 1600, 1735, 1746, 1790, 1902, 1934, 2005, 2059, 2186, 2249, 2322, 2333, 2352, 2483, 2539, 2726, 2818, 2827, 2849, 2909]. real-valued [1892]. real-world [1898, 1976]. reality [48, 283, 566, 1439, 1624, 1868, 2188]. reality-supported [48]. Realization [1074, 1303, 1335, 1879, 2879]. realizing [798, 799, 936]. reallocation [33]. reasoning [694, 770, 1975, 2297, 2556]. rebalancing [819]. rebar [1506]. REC [557]. REC-SSS [557]. recall [401]. Received [2158]. receiving [1323]. receptive [1104]. rechargeable [685]. Recognition [60, 63, 93, 94, 196, 224, 324, 340, 415, 437, 606, 622, 645, 664, 718, 743, 832, 882, 934, 985, 1059, 1063, 1081, 1105, 1120, 1139, 1169, 1230, 1253, 1255, 1292, 1297, 1387, 1433, 1464, 1557, 1561, 1615, 1668, 1670, 1722, 1868, 1896, 1969, 1978, 1990, 2012, 2074, 2102, 2126, 2186, 2238, 2263, 2271, 2296, 2496, 2503, 2511, 2517, 2530, 2536, 2641, 2693, 2696, 2744, 2852, 2861, 2894]. recognizer [2700]. Recognizing [788]. recombination [790]. recommendation [249, 260, 327, 414, 457, 531, 607, 695, 738, 817, 1024, 1057, 1061, 1142, 1251, 1390, 1406, 1517, 1664, 1767, 1786, 2024, 2090, 2111, 2114, 2232, 2233, 2318, 2396, 2442, 2463, 2655, 2679, 2750, 2795, 2815, 2912]. Recommendations [406, 1116, 1459, 1749]. recommender [1179, 2408, 2603, 2615, 2736, 2830]. recommending [182]. Reconfigurable [15, 237, 784, 970, 1167, 1378, 1522, 2337, 2818]. reconfiguration [199, 1010, 1621, 2818]. reconstructing [2543, 2622]. reconstruction [55, 466, 613, 622, 1669, 1720, 2238, 2473]. record [289, 2532]. recorded [2677]. records [304]. Recovery [79, 170, 435, 484, 600, 1195, 1728, 2007, 2087, 2131, 2601]. rectangle [1502]. rectangular [1792]. rectified [1800]. recurrent [424, 435, 460, 1215, 1263, 1598, 1921, 2209, 2506]. recursive [1410, 2704]. Red [489]. Reduce [217, 557, 1383, 1988, 2023, 2083, 2221]. reduced [526, 1089, 1634, 2007]. reduced-round [2007]. reducer [253]. Reducing [226, 410, 779, 904, 1352, 1603, 1806, 2116, 2462, 2609]. reduct [2456]. reduction [50, 103, 406, 528, 592, 759, 885, 921, 1213, 1409, 1545, 1556, 1630, 1649, 1852, 2236, 2456, 2596, 2611, 2863, 2886, 2916]. redundancy [656, 1428, 1883, 2863]. redundant [1055]. Reefer [263]. reevaluation [294]. reference [2885]. refining [403]. reflectors [455]. reflex [818]. refraction [1193]. refrigerators [183]. region [42, 204, 333, 346, 640, 672, 1925, 2051, 2219, 2671]. region-based [333, 1925]. region-partition [640]. Regional [711, 1216, 1715, 2293]. regions [620, 1055, 1351]. register [326, 1930]. registration [466, 485, 869, 1044, 1548, 1944, 2080]. regression [175, 334, 1649, 1716, 1894, 1961, 2420, 2464]. regression-based [1649, 2464]. regressive [1166]. regular [1700]. regularization [94, 2477]. regularized [2540]. regulated [275]. regulation [1267, 1937]. rehabilitation [957]. Reinforcement [31, 34, 629, 782, 1005, 1198, 1206, 1239, 1302, 1431, 1604, 1645, 1820, 1828, 2003, 2027, 2549, 2651, 2656, 2663, 2760, 2813]. related [1324, 2205, 2227]. relatedness [2692]. Relation [627, 1975, 2556, 2772]. Relational [222, 1411, 2285]. relations [1670, 2797]. relationship [2548, 2919]. relative [426, 2738, 2787]. relaxed [1130]. relay [121, 330, 931, 1122, 1748, 2562, 2663]. relay-remote [1122]. Relevance [401, 1588, 1883]. relevant [116]. Reliability [21, 143, 237, 517, 539, 577, 581, 908, 970, 971, 1077, 1299, 1339, 1342, 1405, 1632, 1782, 1823, 1870, 1902, 1942, 2295, 2376]. Reliability-aware [971]. Reliable [50, 992, 1008, 1435, 1493, 1840, 2184, 2360, 2531, 2739]. reliant [2241]. ReLU [2826]. remaining [1941, 2218, 2490, 2734]. remanufacturing [1911]. remixing [2322]. Remodulation [2890]. Remote [176, 437, 565, 630, 739, 740, 791, 853, 1071, 1122, 1491, 1525, 1720, 2058]. removal [1428, 2075]. removed [318]. rendering [431, 812]. renewable [1719, 1810]. renewal [258, 259]. ReNo [1870, 1942]. renovation [153]. Renyi [667, 1304, 1305, 2572]. RenyiBS [667]. reorder [1795]. repair [1011]. repairable [2266]. repeats [1222]. replacement [608, 974, 2049]. replanning [1133]. replay [288]. replica [2139, 2681]. replicas [1736]. Replication [287, 380, 750, 849, 933, 1055, 2502]. REPO [1692]. report [1065, 1066]. reports [2686, 2911]. repositories [244]. Representation [127, 220, 338, 340, 1020, 1057, 1419, 1557, 1749, 1780, 1886, 1937, 2244, 2548, 2679, 2779, 2913]. representation-based [340, 2679]. representational [1926]. Reproducibility [555]. reputation [909, 2518]. request [1403]. requests [2730]. requirement [2143]. requirements [367, 2650, 2713, 2885]. rerouting [506]. resampling [51, 2182]. rescue [198, 2558]. Research [129, 130, 142, 150, 153, 162, 167, 198, 201, 202, 207, 209, 254, 320, 385, 491, 593, 636, 777, 940, 1062, 1208, 1306, 1404, 1466, 1495, 1539, 1799, 2067, 2160, 2307, 2614, 2632, 2675, 2753, 2805, 2826]. researchers [691]. reservation [698]. reservoir [1074]. resident [106]. residual [71, 93, 2171, 2361]. resilience [484, 1317]. resiliency [1150]. resilient [375, 759, 1891, 2701]. resistant [1641, 2207]. ResNet [1808]. resolution [71, 942, 1298, 1376, 1720, 1792, 1970, 2058, 2592]. resonance [818, 1192, 1713]. resonator [696]. resonator-based [696]. Resource [35, 44, 50, 69, 81, 83, 144, 212, 349, 363, 371, 399, 481, 549, 563, 629, 698, 809, 816, 835, 928, 993, 1041, 1067, 1103, 1191, 1227, 1300, 1306, 1333, 1337, 1346, 1415, 1457, 1519, 1530, 1539, 1581, 1592, 1631, 1633, 1638, 1659, 1678, 1686, 1689, 1698, 1744, 1789, 1821, 1850, 1895, 1909, 1943, 1944, 1959, 1960, 1996, 2331, 2370, 2389, 2423, 2472, 2487, 2689, 2716, 2730, 2731, 2881]. resource- [1227]. resource-aware [563, 816]. resource-constrained [1415, 1457]. resource-efficient [1944, 2370]. resources [111, 197, 503, 580, 1389, 1412, 1715, 1880, 2021, 2162, 2170]. respiratory [1745]. RESPNet [1944]. Response [185, 316, 705, 779, 1708, 1961, 2191, 2302, 2738]. responsive [535]. restart [735]. Restricted [1408, 2098]. results [535, 1418]. Rethinking [1642]. retinal [673, 803, 2670]. Retraction [1988–1992, 2038, 2039, 2176, 2306, 2307]. retrial [1249]. retrieval [244, 375, 401, 440, 801, 909, 1315, 1372, 1851, 1853, 2532, 2539]. returned [2773]. returning [690]. reuse [810, 1158]. reusing [841]. Revealing [982, 1486]. reversal [542]. reverse [7]. Reversible [97, 226, 724, 950, 1033, 1042, 1089, 1111, 1112, 1155, 1359, 1839, 2217, 2588]. reversing [540]. review [29, 77, 116, 129, 130, 134, 291, 464, 537, 591, 619, 679, 965, 1014, 1062, 1137, 1289, 1306, 1539, 1643, 1698, 1783, 1815, 1882, 1897, 2071, 2277, 2547, 2703, 2749]. reviews [670, 2198]. Revisiting [859, 1094]. Revocable [1416, 2334]. revocation [1117, 1163, 1332, 2080]. Revolutionizing [2907]. reweighting [1105]. rewrite [190]. rewrite/merge [190]. RFE [2077]. RFID [564, 845, 883, 1369, 1370, 1449, 2607]. RFID-based [564, 883, 2607]. RGB [868, 2657]. RGB-D [2657]. RGBT [2416]. RI [2840]. RI-RPL [2840]. Riccati [599]. Riccati-type [599]. rich [1084]. ride [681, 1367]. ride-hailing [1367]. ride-sharing [681]. ridesharing [75]. ridge [595, 2477]. Ridge-Adaline [595]. right [2332]. right-hand [2332]. rigid [466]. rigorous [2455]. Rijndael [1884]. Rijndael-based [1884]. Ring [360, 550, 696, 1813, 2766]. Ring-mesh [360]. Rio [148, 392]. RISC [2415]. RISC-V [2415]. Risk [105, 179, 912, 1861, 1911, 2380, 2911]. Risk-Averse [1861]. risks [1477]. rival [631]. river [853, 1491, 1945]. RKMIS [2148]. RLWE [2667]. RLWE-based [2667]. RMFFOG [1744]. RMSRM [2818]. RNN [104, 797, 1215]. RNN-based [104]. road [602, 681, 1025, 1428, 1521, 1660, 2394, 2685, 2812]. roads [2430, 2561]. robber [886]. robin [1228, 1229]. robot [62, 192, 1219, 1374, 1433, 1468, 1624, 1763, 1842, 1868, 2092, 2225, 2267, 2878]. Robotic [1967, 2037, 2044, 2339, 2524]. robots [61, 1141, 1404, 1561, 1615, 1650, 1779, 1812, 2617]. Robust [25, 32, 136, 152, 175, 270, 372, 511, 616, 821, 1533, 1655, 1874, 1904, 2148, 2210, 2245, 2448, 2513, 2581, 2654, 2804, 2864, 2896]. Robustness [1242]. rockburst [1572]. role [417]. root [639]. rotary [136, 839]. rotating [654]. rotation [739, 898]. rotation-invariant [739]. rotational [377]. rough [38, 1455]. round [1228, 1229, 1482, 1982, 2007]. round-robin [1228, 1229]. rounding [2447]. rounding-based [2447]. routability [2682]. route [1000, 2660, 2728]. routed [332, 1056]. router [696, 1161]. routes [269]. routine [128]. routines [502]. Routing [19, 66, 70, 95, 216, 228, 447, 562, 575, 600, 668, 709, 756, 859, 884, 908, 967, 989, 1014, 1019, 1039, 1097, 1295, 1316, 1351, 1381, 1397, 1450, 1460, 1472, 1524, 1546, 1622, 1692, 1864, 1873, 1998, 2070, 2145, 2172, 2195, 2202, 2290, 2340, 2489, 2561, 2660, 2682, 2749, 2794, 2840, 2895]. routings [560]. RoV [1323]. rover [150]. Row [2146]. Row-based [2146]. royale [2754, 2781]. RPL [1037, 1654, 2026, 2840]. RPL-based [1037, 2026, 2840]. rRT [800, 2039]. rRT-PCR [800, 2039]. RRTMG LW [706]. RRTMG SW [1705]. RSSI [1980, 2066]. RSSI-based [1980]. RT [236]. RTFTL [1735]. RTL [509]. RTOS [2249, 2352]. rule [244, 594, 846, 865, 895, 905, 1017, 1671, 2243, 2376]. rule-based [244, 895, 1017, 2243]. rules [768, 884, 952, 1262, 1455, 1671, 2428, 2761]. Rumor [1964, 2849]. rung [2799]. Runge [2425]. runner [639]. runner-root [639]. running [62, 237, 1721]. Runtime [21, 642, 1330, 1505, 1522, 1696, 2123, 2320, 2528, 2825]. rural [2430].

s [1201, 2104, 2259, 2287, 2341, 2627]. S-Apriori [2259]. S-boxes [2627]. S-Edge [2341]. S-Shaped [1201]. SA [1977, 2156]. SA-ASBA [1977]. SA-GCN [2156]. Saber.PAKE [2447]. safe [337]. safety [348, 467, 1427, 2394]. SAGE [2211]. SaHNoC [2016]. sailfish [1452]. sale [2287]. sales [1917]. salesman [939, 1946, 2248]. salesmen [977]. Saliency [1164, 2567, 2743]. Saliency-based [2743]. salient [2657]. salp [917, 991, 1656, 2565, 2861]. SAMP [1454]. samplers [2178]. samples [620, 2759]. Sampling [10, 391, 825, 920, 1400, 1610, 2150, 2313, 2422, 2829]. Sampling-based [825, 1400]. sampling-WIoU [2829]. sand [2531]. sandstorm [2377]. sanitized [347]. SAR [1071]. sarcasm [1436, 2174]. Sarcastic [1257]. SARSA [629]. Satellite [58, 853, 1491, 1521, 2434, 2624]. satin [1943]. satisfaction [178, 583, 1168, 2045, 2597]. satisfaction-based [2597]. satisfied [75]. saturation [1571]. Saudi [1325]. sAuth [1650]. save [2630]. saving [267, 309, 416, 425, 968, 1030, 1147, 2414, 2500, 2597]. saw [2833]. SBLA [1716]. SBLA-based [1716]. SBOX [1879]. SBSN [1057]. SCADA [1558]. Scalability [23, 91, 908, 2152, 2862]. Scalable [42, 50, 89, 356, 360, 460, 473, 625, 633, 636, 653, 706, 723, 746, 775, 808, 1029, 1109, 1237, 1350, 1421, 1424, 1470, 1550, 1558, 1642, 1699, 1750, 1764, 1796, 1822, 2035, 2404, 2523, 2613, 2671, 2697, 2875]. scalar [1948, 2576]. scale [8, 23, 53, 68, 71, 95, 204, 354, 388, 442, 507, 548, 582, 656, 671, 678, 908, 1064, 1154, 1160, 1196, 1341, 1377, 1401, 1445, 1500, 1520, 1524, 1551, 1578, 1580, 1646, 1708, 1711, 1720, 1819, 1847, 1918, 2022, 2177, 2461, 2580, 2588, 2633, 2651, 2691, 2693, 2724, 2740, 2785, 2822]. scale-free [1708]. scaled [409]. Scaling [251, 371, 399, 532, 602, 922, 1015, 1856, 2211]. scan [1746, 2798]. scanned [2074]. scanning [818, 2068]. scans [1493]. scatter [1020]. scattering [2876]. Scenario [1281, 1425, 1712, 2368, 2558, 2756]. scenarios [2364, 2380, 2760, 2761]. Scene [1923, 2644]. scenes [283]. Schedulability [1600]. scheduler [490, 509, 519, 563, 776, 1228, 1229, 1437, 1785, 2370]. schedulers [1619, 2454]. Scheduling [6, 21, 31, 34, 35, 76, 135, 139, 234, 236, 237, 291, 309, 319, 369, 376, 383, 397, 441, 465, 476, 503, 544, 548, 589, 596, 629, 689, 730, 752, 785, 835, 875, 917, 966, 969, 971, 1012, 1020, 1021, 1054, 1070, 1078, 1080, 1084, 1088, 1095, 1106, 1125, 1156, 1193, 1205, 1224, 1235, 1258, 1283, 1293, 1306, 1346, 1382, 1420, 1484, 1489, 1532, 1553, 1688, 1690, 1696, 1704, 1707, 1776, 1805, 1823, 1834, 1841, 1951, 1993, 2045, 2057, 2085, 2128, 2166, 2196, 2215, 2240, 2242, 2246, 2283, 2304, 2361, 2406, 2438, 2458, 2470, 2568, 2594, 2614, 2813, 2843, 2863]. scheduling [83, 471, 513, 552, 556, 568, 779, 911, 933, 1223, 1247, 1288, 1333, 1389, 1403, 1481, 1488, 1592, 1652, 1732, 1802, 1835, 1850, 1863, 1888, 2047, 2206, 2316, 2544, 2647, 2716, 2803, 2841, 2900, 2901]. Schema [374, 1331, 2710]. scheme [2, 45, 72, 135, 137, 211, 243, 248, 257–259, 288, 289, 323, 363, 404, 412, 431, 524, 565, 584, 651, 726, 796, 836, 911, 921, 983, 1034, 1035, 1056, 1068, 1096, 1119, 1159, 1186, 1274, 1301, 1324, 1352, 1366, 1388, 1457, 1553, 1554, 1571, 1612, 1613, 1641, 1651, 1681, 1683, 1685, 1695, 1705, 1736, 1771, 1794, 1801, 1859, 1862, 1875, 1878, 1897, 1902, 1906, 1912, 1925, 1943, 1954, 2052, 2089, 2109, 2202, 2207, 2237, 2245, 2251, 2252, 2292, 2314, 2334, 2336, 2350, 2390, 2401, 2410, 2412, 2423, 2440, 2444, 2491, 2492, 2595, 2605, 2612, 2646, 2661, 2755, 2796, 2798, 2831, 2875]. schemes [134, 765, 1079, 1083, 1308, 1730, 2276, 2348, 2901]. Schmidt [87]. Schrödinger [2284]. Schur [2276]. science [636, 2043]. scientific [34, 212, 397, 548, 629, 657, 772, 1021, 1160, 1280, 1688, 1690, 1704, 1823]. scope [2333]. scoping [134]. score [458, 491, 705, 1742, 2040]. scorecard [160]. scoring [512, 2757]. scrambling [2648, 2722]. scratchpad [1069]. screen [711]. screening [672, 850, 1478, 1638]. ScrimpCo [473]. SDAM [722]. SDC [1055]. SDLER [942]. SDN [394, 612, 858, 989, 992, 1421, 1633, 1845, 2025, 2099, 2168, 2340, 2359, 2375, 2435, 2486, 2493, 2661, 2748, 2802, 2820]. SDN-based [394, 1633, 2025, 2375, 2435]. SDN-enabled [2340]. SDVN [1343]. SDVNs [2093]. seagull [1984]. sealed [1995]. search [156, 198, 271, 403, 474, 482, 552, 564, 583, 589, 602, 783, 802, 867, 925, 931, 932, 964, 1020, 1028, 1046, 1177, 1182, 1212, 1235, 1294, 1316, 1440, 1488, 1596, 1660, 1674, 1732, 1738, 1866, 1888, 1890, 1969, 2004, 2029, 2062, 2166, 2204, 2319, 2382, 2406, 2458, 2497, 2546, 2550, 2559, 2616, 2635, 2636, 2680, 2712, 2833, 2843, 2869]. search-based [589, 1046]. Searchable [134, 1413, 2532, 2667]. searches [313, 2728]. searching [363, 2315, 2427]. seasonal [2259]. SeburSum [2268]. second [1261]. second-order [1261]. secondary [195, 721]. Secrecy [137, 2251]. secret [1641, 1695, 2492, 2747]. section [36]. sectional [1636]. sector [241]. Secure [70, 453, 518, 584, 647, 652, 658, 672, 708, 715, 719, 720, 750, 796, 808, 836, 864, 879, 1026, 1045, 1091, 1369, 1370, 1375, 1385, 1388, 1526, 1569, 1608, 1695, 1771, 1855, 1871, 1873, 1875, 1906, 1954, 1958, 2052, 2100, 2171, 2172, 2241, 2282, 2290, 2292, 2314, 2334, 2390, 2402, 2440, 2480, 2518, 2566, 2595, 2607, 2608, 2628, 2634, 2663, 2774, 2809, 2848, 2875, 2914]. Secured [66, 1381, 1967, 2037, 2778]. securer [2667]. Securing [121, 374, 436, 452, 2026]. Security [41, 55, 105, 120, 288, 308, 347, 374, 443, 445, 486, 616, 680, 700, 709, 742, 822, 878, 1017, 1037, 1053, 1096, 1184, 1185, 1560, 1605, 1778, 1848, 1884, 1893, 1927, 1933, 2093, 2224, 2237, 2378, 2547, 2625, 2655, 2703, 2764, 2777, 2883, 2908]. security-aware [742]. seed [510, 2171]. seed-residual [2171]. seeding [2690]. Seeking [491]. segment [2699]. Segmentation [112, 150, 454, 661, 673, 783, 803, 901, 1127, 1137, 1308, 1462, 1476, 1521, 1734, 1740, 1768, 1816, 1897, 1904, 2051, 2225, 2253, 2624, 2652, 2705, 2718, 2743, 2785, 2827, 2890]. segmentation-based [1897]. segmented [338]. Segmenter [2075]. segregation [424]. SEIR [1325]. seismic [2260]. seizure [164, 676, 1371, 2383, 2446]. Selecting [952, 1338]. Selection [125, 166, 221, 327, 330, 335, 343, 386, 437, 510, 512, 514, 542, 631, 667, 677, 733, 745, 748, 771, 786, 867, 924, 991, 1014, 1058, 1122, 1181, 1268, 1281, 1502, 1588, 1592, 1617, 1635, 1667, 1729, 1765, 1816, 1886, 1900, 2034, 2040, 2086, 2155, 2172, 2227, 2243, 2247, 2255, 2265, 2277, 2301, 2343, 2347, 2353, 2396, 2439, 2499, 2565, 2575, 2583, 2660, 2663, 2720, 2721, 2737, 2803, 2852, 2914]. selection-based [510]. selective [180, 345, 632, 1055, 1104, 1259, 2367, 2823]. selector [1296]. Self [61, 64, 258, 259, 669, 761, 874, 1279, 1378, 1689, 1728, 1933, 1952, 2806, 2818, 2899, 2912, 2917]. self-adaptation [1378, 1952]. Self-adaptive [64]. self-adjusting [258, 259]. self-attention [2899, 2912]. self-awareness [61]. self-driving [669, 2806]. self-healing [1728]. Self-learning [1689, 1933]. self-mixing [874]. self-organized [761]. self-organizing [1279]. Self-paced [2917]. self-reconfiguration [2818]. self-recovery [1728]. selfish [754, 857, 1360, 1613]. Semantic [20, 57, 222, 246, 247, 255, 298, 413, 979, 1183, 1233, 1768, 1851, 1965, 2225, 2232, 2624, 2678, 2692, 2741, 2801, 2827, 2890]. semantic-aware [1183, 2232]. semantics [626, 2669, 2759]. Semi [496, 661, 670, 869, 1109, 1288, 1538, 1859, 2020, 2207, 2535]. semi-aggregation [1109]. semi-automated [496]. semi-Markov [869]. Semi-partitioned [1288]. semi-sleep [1859]. Semi-supervised [661, 670, 1538, 2020, 2535]. semi-trusted [2207]. semiconductor [703, 811, 863, 1074, 1844]. semiconductors [2639]. SEN [517]. Sense [2431]. sensing [176, 437, 630, 739, 763, 791, 794, 853, 874, 926, 999, 1071, 1107, 1491, 1677, 1710, 1720, 2058, 2109, 2564]. Sensitive [395, 1080, 1133, 1202, 1439, 1675, 2715]. sensitivity [2509]. Sensor [22, 54, 66, 70, 110, 124, 216, 219, 246, 247, 257, 270, 333, 342, 349, 350, 357, 385, 426, 555, 606, 639, 658, 668, 712, 756, 839, 859, 873, 895, 908, 962, 963, 1034, 1035, 1187, 1198, 1245, 1268, 1278, 1295, 1408, 1429, 1460, 1472, 1492, 1524, 1558, 1582, 1620, 1662, 1695, 1744, 1817, 1888, 1976, 1980, 2066, 2148, 2158, 2194, 2229, 2243, 2270, 2317, 2413, 2431, 2495, 2549, 2601, 2749, 2782, 2803, 2814, 2836, 2891, 2895, 2915]. sensor-based [1662, 1695]. sensor-fog [2431]. sensorless [854, 1451]. sensors [557, 1007, 1147]. sentence [1292, 1670, 2759]. sentiment [189, 221, 470, 670, 876, 1009, 1046, 1215, 1289, 1319, 1419, 1606, 1803, 1972, 1977, 2008, 2013, 2142, 2198, 2205, 2214, 2230, 2298, 2426, 2515, 2668, 2678, 2736, 2837, 2873, 2889, 2899]. sentimental [2506]. sentinet [1009]. sentinet-based [1009]. Seoul [2711]. separable [1418]. separation [186, 545, 830]. SeQual [2347]. Sequence [472, 817, 1244, 1364, 1709, 1758, 1811, 1815, 1929]. Sequence-aware [817]. sequence-structure [1244]. sequenced [1603]. sequences [215, 1781, 2105, 2381, 2464]. sequencing [1603]. Sequential [269, 1042, 1122, 1465, 2020, 2201, 2217, 2463, 2485, 2690, 2900]. series [133, 328, 366, 435, 760, 773, 797, 1065, 1066, 1409, 1579, 1588, 1657, 1892, 1917, 1940, 2097, 2161, 2357, 2397, 2475, 2543, 2598, 2619, 2622, 2708, 2711, 2868, 2871, 2887]. serological [850]. server [82, 129, 130, 207, 387, 487, 520, 826, 847, 972, 1213, 1240, 1426, 1568, 1794, 2241, 2342, 2361, 2388, 2571, 2797]. serverless [1562, 1982, 2112, 2219]. servers [1467, 1498, 2152, 2387]. Service [12, 108, 125, 136, 178, 232, 268, 273, 402, 469, 481, 523, 542, 603, 653, 656, 661, 784, 847, 944, 1004, 1024, 1047, 1061, 1070, 1142, 1144, 1145, 1243, 1251, 1272, 1286, 1321, 1338, 1414, 1447, 1477, 1504, 1549, 1592, 1618, 1620, 1650, 1689, 1840, 1841, 1863, 1995, 2107, 2114, 2216, 2243, 2258, 2316, 2432, 2449, 2455, 2613, 2637, 2655, 2679, 2725, 2760, 2765, 2807, 2872, 2893]. service-aware [469]. service-oriented [656, 2765]. service-providers [944]. services [99, 163, 280, 418, 438, 584, 758, 1197, 1302, 1338, 1517, 1563, 1612, 1642, 1697, 1838, 1889, 1967, 2037, 2157, 2183, 2209, 2715, 2795]. session [616, 2388, 2830]. session-based [2830]. Set [1, 38, 743, 961, 1075, 1455, 1531, 1725, 1772, 2020, 2051, 2204, 2268, 2299, 2395, 2415, 2516, 2814]. set-based [1075, 2268]. sets [1888, 1928, 2046]. sets-based [1928]. SEUs [2902]. several [2640]. severity [2621]. SFC [1005, 2457]. SFLA [557]. SG [2644]. SGD [1508, 1509, 2301]. SHA [1854]. SHA-256 [1854]. shader [1290]. shadow [203, 2075]. Shafer [617, 1831, 1938, 2255, 2439]. shallow [2430]. Shannon [2356]. shape [338, 2724]. shaped [186, 717, 1113, 1201, 2278]. Shapley [1589]. sharding [2523, 2706]. share [730, 2492]. share-based [2492]. shared [2, 94, 279, 332, 393, 974, 1270, 1333, 1486, 1494, 2665]. shares [2171]. sharing [177, 314, 681, 1346, 1641, 1695, 1789, 2100, 2314, 2319, 2336, 2390, 2440, 2492, 2664, 2711, 2747, 2773, 2848]. sharing-based [1641, 1695]. shedding [256]. shelf [1357]. shell [115, 2366]. shepherd [1503]. Shift [5, 326, 1991]. ship [1564, 2580, 2888]. Shipping [2219]. shop [1020, 1054, 1193, 1235]. short [221, 435, 477, 760, 1001, 1228, 1229, 1257, 1341, 1784, 1852, 2033, 2291, 2430, 2463, 2655, 2671, 2773, 2784]. short-entry [2671]. short-term [221, 435, 477, 760, 1001, 1228, 1229, 1257, 1341, 1784, 1852, 2291, 2430, 2463, 2655, 2773, 2784]. shortest [278, 681]. shortwave [1705]. shot [381, 1814, 2225, 2294, 2721]. shuffle [517, 526, 967, 1587, 1987, 2084, 2220]. shuffle-exchange [517, 967]. shuffled [541, 1503]. shuffling [2171, 2434]. Siamese [1585, 2692, 2723]. Side [918, 1152, 1749, 2162, 2405, 2723]. Side-channel [1152, 2723]. sided [75]. sides [2332]. sieving [627]. sifter [186]. sifter-shaped [186]. Sigma [2343]. Sign [1100, 1169, 1344, 1799, 1847, 1990, 2691]. signal [99, 223, 434, 435, 484, 551, 764, 1313, 1560, 1693, 1873, 2158]. signaling [2031]. signals [72, 93, 419, 2012, 2363, 2717]. signature [45, 1163, 1954, 2745]. signcryption [2412, 2532]. signed [2133]. significantly [2083, 2221]. signs [2186, 2644]. silo [1871]. SiMAIM [2474]. SiMaLSTM [2692]. SiMaLSTM-SNP [2692]. SIMD [87, 90, 238, 1833, 2332]. Similar [713, 2154, 2380]. Similarity [11, 57, 817, 1116, 1224, 1648, 1875, 2119, 2567, 2587, 2735]. SIMOF [2195]. Simple [489, 952, 1079, 1087, 1362, 1867, 2763]. simplex [497, 857]. Simplified [1566, 2080]. Simplifying [232, 2522]. SIMT [1318, 1481]. simulate [678]. simulated [180, 977, 1235, 1619, 1635]. Simulating [1843, 1950]. Simulation [3, 88, 100, 153, 179, 187, 253, 276, 353, 576, 633, 790, 1038, 1111, 1112, 1315, 1318, 1379, 1456, 1466, 1580, 1703, 1984, 2095, 2284, 2327, 2586, 2888]. Simulation-based [2888]. simulations [536, 555, 880, 1196, 1383, 1743, 1807, 2476, 2525]. Simulator [4, 1076, 1438, 1699, 1992]. Simulink [2357]. simultaneous [1353, 1567, 2332]. sine [2270]. single [236, 414, 511, 1233, 1540, 2002, 2040, 2474, 2492, 2514, 2801, 2896]. single- [2514]. single-cell [2040]. single-image [2801, 2896]. single-layer [511]. single-machine-based [414]. single-task [1233]. SINGLETON [658]. singular [898]. Sink [22, 333, 668, 756, 962, 963, 1472, 1554, 2694]. sink-based [1472]. Sink-Enabled [962, 963]. sinks [124, 1976]. SIoT [944]. SIP [137]. SIR [114, 1325]. SIR-F [1325]. site [2474]. sites [2799]. situ [1196]. situation [55, 888, 1999]. Six [2343]. size [186, 711, 945, 1609, 1747, 2312, 2865]. size- [186]. sizes [85]. skeleton [297, 1430, 2132, 2860]. skeletons [300]. SkePU [297]. sketches [1961]. Sketching [2477]. skewed [2557]. skin [1507, 2621]. skinny [455]. skyline [39, 53, 327, 538, 2183, 2685]. Sl4u [2756].

SLA [386, 1535, 1676]. Slack [1283]. SLAM [2372]. SLAs [1326]. SLDChOA [2698]. sleep [365, 1697, 1859]. SLI [1765]. sliced [1919]. slicing [1960, 2362, 2715]. slides [2828]. Sliding [415, 1011, 1396, 1571, 2842]. slime [1282, 1715]. SLO [2215]. SLO-aware [2215]. slope [1786]. slot [2469]. slotted [872]. slow [2637]. SlowTrack [2637]. SM [1744]. Small [393, 416, 582, 802, 1121, 2046, 2305, 2432, 2461, 2644]. smallholder [2624]. Smart [59, 121, 147, 156, 194, 290, 311, 318, 357, 418, 453, 467, 499, 510, 638, 720, 725, 744, 767, 872, 904, 981, 984, 1037, 1053, 1057, 1090, 1096, 1155, 1169, 1199, 1209, 1228, 1229, 1278, 1281, 1375, 1386, 1393, 1422, 1459, 1468, 1484, 1487, 1527, 1555, 1669, 1756, 1804, 1902, 1906, 1967, 1978, 1990, 2037, 2052, 2165, 2176, 2197, 2209, 2241, 2275, 2287, 2333, 2350, 2384, 2411, 2485, 2542, 2573, 2609, 2625, 2634, 2658, 2672, 2725, 2778, 2807, 2875]. Smartphone [557, 573, 1105, 2390]. smartphones [1253, 1662]. smartwatch [557]. SMC [2078]. SMC-BRB [2078]. SMIGNN [2815]. smile [178]. SMO [1451]. SMOaaS [653]. smoke [926, 1978, 2918]. Smooth [2223]. SMR [2288]. SMS [1995]. SMVQ [226]. SN [2138]. SNCL [2903]. snow [2377]. SNP [2692]. snubber [1384]. Sobel [1658]. SoC [792]. Social [58, 81, 178, 188, 189, 249, 271, 408, 457, 570, 605, 616, 795, 825, 944, 953, 973, 975, 1009, 1057, 1100, 1116, 1148, 1174, 1251, 1257, 1261, 1334, 1350, 1356, 1400, 1436, 1613, 1615, 1646, 1725, 1779, 1812, 1842, 2067, 2090, 2111, 2133, 2136, 2138, 2174, 2292, 2318, 2345, 2442, 2474, 2506, 2638, 2690, 2770, 2797, 2815, 2831, 2832, 2847, 2911, 2912]. social-aware [2318, 2442]. societal [1123]. sockpuppets [2474]. SoCs [521, 1281]. Soft [621, 1216, 1380, 2032]. soft-core-based [1380]. Softplus [2421]. Software [10, 19, 20, 143, 362, 502, 509, 537, 543, 831, 965, 1008, 1022, 1032, 1114, 1191, 1234, 1269, 1279, 1343, 1348, 1540, 1546, 1678, 1680, 1699, 1833, 1914, 1936, 1971, 2055, 2093, 2145, 2169, 2216, 2229, 2343, 2417, 2420, 2452, 2486, 2731, 2907]. software-based [1348, 1833]. Software-defect [1279]. software-defined [19, 543, 1008, 1022, 1032, 1191, 1343, 1540, 1546, 1678, 2055, 2093, 2216, 2229, 2486]. Softwarization [1591]. solar [1368, 2836]. solar-powered [2836]. solid [512]. solidification [1950]. solids [193]. soluble [193]. solution [79, 367, 554, 676, 884, 907, 940, 970, 1020, 1708, 2005, 2014, 2332, 2547, 2712]. solutions [394, 486, 784, 878, 1017, 1053, 1145, 2435, 2514]. solvation [86]. solve [627, 1597, 2513]. solver [126, 1469, 1470, 2213, 2236, 2284]. solvers [871, 2248]. solving [317, 441, 482, 594, 599, 610, 766, 986, 1072, 1106, 1212, 1235, 1578, 1627, 1946, 1984, 2179, 2258, 2281, 2326, 2355, 2461, 2520, 2782, 2814, 2855]. SOM [713]. SOM-similar [713]. some [560, 1673, 1834]. somewhere [1663]. SOR [489]. SORec [1057]. sort [2338]. sorting [186, 1981, 2626]. sorting-based [2626]. SOS [2595]. SOS-FCI [2595]. sound [435]. sounds [344, 830]. Source [211, 361, 545, 830, 1384, 1463, 1780, 2006, 2112]. Source-free [2006]. Source-to-source [361]. sourced [888]. sources [442, 477, 761, 1152, 1719, 1810]. south [2293]. southwest [2293]. sowing [205]. Space [11, 55, 138, 172, 203, 613, 845, 1036, 1076, 1493, 1981, 2118, 2680, 2728, 2779]. Space-efficient [2118]. Space/time [1981]. Space/time-efficient [1981]. spaces [891]. spacing [1506]. Spam [57, 286, 1783]. spammer [1174, 2638]. Spanning [1410, 1490, 1497, 1867, 2000, 2640]. spark [379, 396, 421, 526, 569, 590, 751, 753, 809, 951, 968, 1168, 1195, 1246, 1543, 1757, 1824, 2018, 2200, 2529, 2697]. Spark-based [526, 809, 1757, 2200, 2529, 2697]. sparrow [1182, 2546, 2833]. Sparse [42, 90, 358, 461, 606, 610, 654, 766, 829, 1100, 1470, 1471, 1649, 1932, 2078]. Spatial [111, 313, 792, 900, 1165, 1244, 1313, 2056, 2121, 2130, 2315, 2426, 2460, 2517, 2566, 2753, 2801]. Spatial-guided [2801]. Spatial-temporal [900, 2121, 2460]. spatially [2665]. spatio [2482, 2490]. spatio-temporal [2482, 2490]. Spatiotemporal [102, 495, 1922, 2354, 2496, 2552, 2693, 2842, 2894]. special [36, 627, 1677, 2129]. special- [627]. specialty [1453]. specific [300, 428, 429, 786, 2547]. specification [2264]. spectral [115, 216, 343, 344, 567, 630, 633, 637, 1169, 1596, 1674, 1769, 1990, 2591]. spectrogram [2592]. spectrograms [1307]. spectroscopy [193]. spectrum [314, 619, 1371, 1478, 2564]. speculative [294]. Speech [424, 768, 1183, 1307, 1615, 1801, 2271, 2506, 2530, 2805]. speech-based [1183]. speed [192, 223, 858, 899, 1143, 1520, 1760, 2462, 2584, 2702, 2714, 2806]. speeded [2245]. Speeding [960, 1640]. spending [1601]. spheres [1290]. spherical [798, 799]. spheroidal [46]. spider [81, 1817]. spiking [352, 884, 2530]. spin [1033]. spiral [2520]. Spline [2158]. split [1094, 1384, 2733]. splitting [2428]. SPMD [1750]. SPMV [766, 1285]. spondylosis [1265]. sports [5, 381, 957, 1263, 1272, 1991]. spot [454, 2063]. spread [173]. spreadsheets [294]. SPTRSV [2346]. square [865, 2168]. SR [2361]. SR-PSO [2361]. SRAM [1155]. SRFNet [1104]. SS [808]. SS-ITS [808]. SSD [788, 2288]. SSDs [462, 1403]. SSML [2769]. SSS [557]. SSSAL [2020]. ST [781, 1451, 2850]. ST-CAC [781]. ST-SMO [1451]. ST-YOLOX [2850]. Stability [325, 1162, 2604]. stabilized [945]. Stable [75, 333, 362, 907, 969, 1837, 2660]. Stack [321, 704, 722]. stack-type [704]. Stacked [942, 1890, 2256, 2677]. Stackelberg [142, 1337]. Stacking [2101, 2420]. staff [1999]. stage [317, 391, 449, 468, 1186, 1415, 1657, 1864, 2346, 2700, 2711, 2729, 2838]. stages [179, 800, 1152, 2039]. staging [2884]. standalone [2192]. Star [1540, 1887, 1931, 1986, 2120, 2338, 2498].

Star-structure [1887, 1986]. starling [2712]. start [1783, 2240]. state [512, 573, 977, 1865, 2367, 2469, 2600, 2680, 2728]. state-of-charge [2367]. state-of-the-art [573]. state-space [2728]. statements [238]. Static [494, 677, 1088, 2916]. station [1929, 2368, 2486]. stationary [2842]. statistical [63, 351, 611, 612, 1224, 1553, 2049, 2446, 2714, 2810]. statistical-based [2049]. Statistically [1663, 2265]. status [2375, 2489]. stealing [937]. Stealth [1355]. steel [236, 1665, 2293, 2306]. STEEL-RT [236]. steepest [1819]. steganography [2431, 2581]. stencil [89, 1094, 2132, 2199, 2264]. step [262, 524, 657, 928, 945]. step-size [945]. Stereovision [1598]. STFM [2706]. stiffness [1598]. stitching [1044]. STM [2552]. STM-GCN [2552]. STMAD [1662]. Stochastic [82, 278, 487, 578, 595, 945, 959, 1553, 1731, 2378, 2575, 2647]. Stock [92, 151, 430, 1149, 1483, 2871]. Stokes [1469]. stop [306]. storage [41, 90, 148, 244, 261, 287, 288, 292, 392, 443, 523, 701, 715, 751, 775, 1117, 1129, 1165, 1180, 1237, 1332, 1366, 1416, 1444, 1663, 1719, 1771, 1801, 1810, 1855, 1875, 1884, 2192, 2209, 2288, 2324, 2334, 2510, 2532, 2569, 2848]. storages [23]. stores [1981, 2710]. storing [2766]. storm [563, 2004]. straggler [18, 370, 507, 1637]. Strait [2714]. strategic [1058]. Strategies [1584, 1747, 2064]. strategy [14, 139, 287, 368, 369, 395, 515, 596, 618, 656, 689, 691, 707, 729, 745, 904, 915, 938, 968, 994, 1014, 1206, 1277, 1285, 1355, 1454, 1489, 1502, 1592, 1616, 1679, 1779, 1812, 1824, 1983, 1995, 2004, 2019, 2060, 2080, 2139, 2160, 2180, 2194, 2246, 2268, 2288, 2355, 2499, 2550, 2557, 2698, 2748, 2777, 2820, 2823, 2880]. stream [131, 226, 232, 246, 247, 346, 626, 916, 953, 959, 1209, 1221, 1330, 1482, 1542, 1785, 2088, 2122, 2832, 2849]. Streaming [396, 490, 922, 929, 982, 1948, 2025, 2059, 2200, 2404]. streams [476, 793, 1202, 1259, 1896, 2313, 2557, 2632, 2842]. streets [1203]. strength [223, 2158, 2242]. strength-based [2158]. stretching [918]. string [2118]. stroke [508]. Structural [20, 165, 539, 1107, 1206, 1832, 2752]. Structure [47, 195, 206, 305, 516, 626, 674, 721, 843, 855, 951, 1019, 1089, 1209, 1234, 1237, 1244, 1315, 1704, 1727, 1736, 1887, 1926, 1931, 1944, 1950, 1986, 2098, 2120, 2156, 2160, 2727, 2744, 2789, 2859]. structure-aware [1704, 2156]. structure-based [2160]. Structure-preserving [843, 1944]. Structured [126, 1325, 1564, 1712]. structures [298, 1948, 1964]. STT [1545]. STT-MRAM [1545]. students [74, 2157]. studies [434, 2071]. Study [26, 111, 144, 146, 178, 441, 489, 524, 609, 620, 671, 691, 765, 804, 805, 927, 935, 956, 1036, 1090, 1126, 1165, 1308, 1331, 1543, 1844, 1872, 2009, 2043, 2144, 2231, 2319, 2466, 2713, 2901]. Studying [1727]. stuffing [2492]. style [1671]. stylization [843]. sub [1036]. subband [1507]. subgraph [2177, 2556]. subjects [165, 912]. suboptimal [167]. subpopulations [2319]. subset [1486, 1883]. subsets [295]. subspace [94, 2477]. substance [1486]. substances [820]. substantial [2917]. substitute [2396]. substitution [2402]. Substructure [2098, 2120, 2556]. Substructure-aware [2556]. subsystems [1643]. subtraction [2075]. subtractor [96, 821]. subtractors [1718, 2898]. Success [1208, 1291]. Succinct [1322]. SUDV [1936]. sufficient [789]. suffix [1981]. suitable [2046]. suite [159, 1889]. sum [558]. summarization [73, 409, 626, 1164, 1772, 1956, 2268, 2393, 2443, 2534, 2686, 2742]. summarized [246, 247]. summary [2268]. summary-level [2268]. summation [1240]. Sunway [699, 1221, 1234, 1236, 1394, 1470, 1743, 2213, 2391, 2733]. SunwayURANS [1743]. super [71, 195, 661, 1298, 1376, 1644, 1720, 1792, 1970, 2381, 2592]. super-computing [661]. super-resolution [71, 1298, 1376, 1720, 1792, 1970, 2592]. super-secondary [195]. Superblock [1236]. Superblock-based [1236]. supercomputer [384, 581, 1221, 2087, 2391]. supercomputers [483, 979, 1160, 2545]. Supercomputing [156, 567, 609, 923, 1603, 1721, 1819, 1941, 2030]. superior [2643]. Superiorization [484]. supermarkets [101]. supernode [2903]. superpixel [901]. supersingular [1287, 1526]. supertask [1244]. Supervised [189, 359, 661, 670, 1361, 1538, 1886, 1972, 2006, 2020, 2535, 2890]. supervising [1934]. suppliers [2224]. supply [59, 177, 194, 219, 1771, 1911, 2176, 2639, 2739]. Support [37, 187, 297, 355, 499, 504, 588, 903, 905, 960, 1025, 1338, 1566, 1618, 1630, 1788, 1914, 2123, 2157, 2223, 2669]. supported [48, 1694]. supporting [190, 1391]. supports [495]. suppression [32, 1031, 2792]. surface [466, 1665, 2306, 2614, 2738]. surfaces [2399]. SURFBCS [2245]. surgery [566]. surgical [1598]. Surix [696]. surrounding [2744]. Surveillance [788, 1064, 1551, 1746, 2304, 2493]. Survey [8, 268, 385, 394, 436, 486, 507, 692, 813, 878, 972, 1030, 1144, 1295, 1329, 1413, 1472, 1576, 1717, 1985, 2017, 2093, 2143, 2348, 2384, 2514, 2707, 2858, 2901]. Survivable [1957, 2457]. survival [329]. Suspicious [647]. sustainability [1643]. sustainable [273, 744, 2324, 2838]. sustained [772]. SVD [1085, 1693]. SVM [40, 854, 882, 1154, 2038, 2478]. SVM-BP [882]. SVR [2077]. SW26010 [576, 1236, 1854, 2876]. Swarm [149, 174, 208, 221, 242, 355, 454, 499, 771, 814, 819, 917, 938, 983, 991, 1018, 1141, 1227, 1257, 1382, 1408, 1656, 1709, 1711, 1762, 1809, 1989, 1997, 2195, 2281, 2361, 2500, 2565, 2695, 2731, 2756, 2861, 2882]. Swarm-based [242]. swarm-intelligence [1382]. swarming [1604]. SWaT [2611]. SWEclat [396]. swept [594]. Swin [2850]. switch [543, 1633, 1870, 1942]. switched [806]. switches [2375]. Switching [530, 2564]. swParaFEM [2213]. swPTS [2733]. swSuperLU [1470]. SX [851]. SX-Aurora [851]. sybil [1037, 1661, 2832]. SYCL [2467]. Sylvester [907]. symbiosis [1282]. symbiotic [867, 1212, 2204]. symbolic [2264]. Symmetric [258, 259, 577, 1040, 2207, 2388, 2610]. symmetry [1647]. symptoms [1231]. SYN [2168]. synchronization [530, 676, 1508, 1509, 1568]. synchronized [2089]. synchronous [1796]. Synergistic [2188]. Syntactic [2296, 2678, 2757]. Syntax [1670]. Syntax-Aware [1670]. synthesis [692, 1570, 1577, 1731, 1799, 1910, 2236]. synthesizing [1626]. synthetic [251, 1977, 2225]. SysML [1843]. system [8, 14, 16, 59, 139, 173, 205, 256, 263, 264, 272, 311, 312, 322, 327, 340, 375, 413, 440, 446, 503, 531, 546, 582, 595, 603, 665, 688, 701, 792, 839, 887, 892, 895, 905, 909, 936, 943, 980, 988, 1003, 1013, 1052, 1092, 1129, 1132, 1133, 1136, 1142, 1183, 1198, 1226, 1243, 1254, 1271, 1328, 1356, 1393, 1402, 1414, 1449, 1466, 1505, 1514, 1615, 1647, 1687, 1703, 1722, 1833, 1859, 1908, 1915, 1924, 1941, 2001, 2041, 2053, 2065, 2121, 2148, 2192, 2212, 2229, 2235, 2289, 2327, 2333, 2367, 2369, 2486, 2539, 2540, 2603, 2612, 2615, 2689, 2726, 2736, 2805, 2807, 2878, 2902, 2910]. system [7, 37, 75, 113, 119, 148, 183, 207, 214, 218, 219, 222, 254, 265, 318, 321, 341, 355, 384, 392, 450, 458, 467, 525, 539, 545, 588, 608, 638, 772, 781, 786, 794, 802, 830, 897, 984, 1163, 1170, 1179, 1188, 1288, 1332, 1358, 1366, 1386, 1443, 1453, 1536, 1571, 1631, 1664, 1702, 1714, 1721, 1911, 1955, 1969, 1997, 2088, 2140, 2160, 2191, 2266, 2308, 2316, 2377, 2380, 2433, 2480, 2590, 2600, 2625, 2655, 2725, 2776, 2778, 2892]. system-based [214]. system-on-chip [2910]. systematic [29, 129, 130, 591, 679, 965, 1530, 1539, 1643, 1698, 1717, 1815]. Systems [15, 20, 30, 77, 136, 210, 273, 358, 376, 380, 447, 464, 465, 468, 490, 519, 523, 537, 544, 550, 556, 592, 594, 598, 621, 642, 727, 730, 752, 766, 808, 838, 871, 883, 884, 917, 947, 948, 958, 971, 1038, 1071, 1088, 1107, 1119, 1184, 1200, 1242, 1270, 1313, 1355, 1369, 1370, 1378, 1380, 1386, 1496, 1500, 1570, 1600, 1651, 1696, 1710, 1728, 1790, 1849, 1907, 1937, 1952, 2005, 2052, 2129–2131, 2143, 2179, 2182, 2279, 2320, 2332, 2344, 2370, 2408, 2466, 2493, 2502, 2509, 2560, 2565, 2569, 2646, 2767, 2792, 2804, 2830, 2843, 2845, 2885]. systems [120, 316, 471, 509, 552, 656, 723, 990, 1007, 1084, 1185, 1207, 1480, 1597, 1605, 1691, 1865, 1875, 2024, 2047, 2733, 2879].

T [2257, 2854, 2883]. T-ECDSA [2883].

t2fcatr [1381]. table

[66, 368, 459, 1013, 1791, 2359, 2577, 2761]. table-based [1013]. tabu [1888]. tabular [2146]. TACHD [1214]. TACTIRSO [1997]. tagged [1806]. TaihuLight [1221, 1234, 1743]. Tailoring [2124]. Tails [507]. Taiwan [328, 2714]. Taiwanese [611]. TAKM [2811]. TAKM-FC [2811]. TAL [869]. TAL-based [869]. talent [1271, 1638]. talk [2564]. Tall [185, 455]. Tall-and-skinny [455]. TAMER [592]. Tamil [256]. tandem [1222]. TAoD [17]. tap [1102]. TARA [2331]. Target [51, 327, 685, 757, 1395, 1518, 2303, 2484, 2495]. targeting [361]. targets [251]. Task [6, 61, 236, 294, 309, 319, 354, 376, 383, 525, 544, 556, 592, 625, 629, 656, 695, 726, 734, 752, 902, 911, 917, 959, 966, 971, 994, 1012, 1048, 1080, 1088, 1106, 1156, 1205, 1223, 1233, 1247, 1293, 1415, 1420, 1505, 1512, 1532, 1649, 1707, 1710, 1805, 1834, 1850, 1927, 2045, 2086, 2115, 2160, 2278, 2320, 2331, 2350, 2363, 2370, 2419, 2546, 2553, 2568, 2689, 2793, 2794, 2813, 2843, 2900, 2901]. task-based [294, 1505, 2320, 2794]. task-single [236]. tasks [212, 237, 370, 465, 509, 568, 816, 1125, 1182, 1227, 1283, 1382, 1581, 1584, 1652, 1712, 1951, 2047, 2162, 2193, 2438]. taxonomies [1124]. taxonomy [134, 507, 1730, 2438, 2749]. Taylor [1541]. TB [2794]. TB-TBP [2794]. TBP [2794]. TCA [2058]. TCA-based [2058]. TCAM [899, 1537]. TCAM-based [1537]. TCEM [342]. TCEM-PQM [342]. TCP [2168]. TCR [375]. TDMA [1191]. TDMBBO [2144]. teacher [163, 2006, 2889]. Teaching [628, 1065, 1066, 1914, 2400]. Team [1763, 2193]. teams [1763]. Technique [16, 66, 186, 225, 365, 409, 489, 705, 718, 801, 826, 898, 923, 928, 1016, 1097, 1103, 1175, 1245, 1278, 1371, 1529, 1558, 1762, 1766, 1778, 1781, 1841, 1853, 1951, 1968, 1997, 2051, 2064, 2116, 2158, 2340, 2389, 2515, 2539, 2581, 2654, 2658, 2786]. Techniques [46, 116, 144, 217, 262, 315, 385, 411, 540, 645, 669, 692, 721, 804, 805, 825, 889, 912, 972, 982, 1078, 1137, 1295, 1348, 1384, 1816, 1844, 1852, 1907, 1988, 2071, 2083, 2087, 2107, 2221, 2228, 2377, 2429, 2565, 2611, 2623]. technologies [78, 1030, 1107, 1124, 1969, 2188]. technology [5, 179, 202, 209, 263, 387, 518, 684, 703, 821, 822, 863, 879, 897, 957, 1004, 1033, 1099, 1208, 1210, 1226, 1230, 1265, 1276, 1335, 1433, 1434, 1474, 1624, 1726, 1742, 1760, 1882, 1915, 1991, 2009, 2089, 2099, 2196, 2307, 2635, 2685, 2689, 2701, 2786, 2816, 2869, 2897]. telecare [2612]. telecom [1202]. telecommunication [241]. TeleDAL [2464]. telehealth [2336]. Telemedical [1560]. TempChain [2336]. temperature [861, 1437, 2738]. temperature-aware [1437]. temperatures [553]. template [1072, 2225, 2464]. template-based [1072]. template-less [2464]. Temporal [99, 646, 693, 900, 1244, 1376, 1874, 2091, 2121, 2460, 2482, 2490, 2519, 2559, 2750, 2772]. temporal-aware [1874]. Temporal-order [2750]. temporal-spatial-based [1244]. tennis [2761]. Tensor [406, 561, 1769, 2097, 2742]. tent [1979]. TentISSA [2893]. TentISSA-BPNN [2893]. Terahertz [2861]. term [221, 435, 477, 760, 941, 1001, 1142, 1228, 1229, 1257, 1341, 1784, 1852, 1967, 2037, 2291, 2395, 2430, 2463, 2655, 2773, 2784]. Terminal [577, 1004]. terminals [2844]. terms [1284, 1993]. ternary [316, 724, 1303, 1611, 1718, 1946, 2826]. Test [264, 800, 1441, 1716, 1838, 2039, 2045, 2168, 2420, 2798]. testicular [1214]. testing [306, 960, 1033, 1716, 2462, 2593]. tetrolet [103]. Text [748, 802, 953, 1164, 1292, 1315, 1448, 1656, 1772, 1851, 2033, 2074, 2138, 2198, 2244, 2445, 2548, 2584, 2629, 2686, 2719, 2849]. text-based [1851]. Text-image [2445]. texts [1956]. Textual [800, 1968, 2039, 2146, 2273, 2911]. texture [2125, 2570, 2572]. textures [659]. TFAS [2605]. TFMD [1343]. TFMD-SDVN [1343]. th [2633]. theater [1387]. their [101, 558, 768, 886, 2735]. theme [892]. theorem [958]. theoretic [574, 649]. Theoretical [1, 123, 1531, 2317]. theory [142, 149, 173, 280, 547, 617, 885, 969, 1412, 1455, 1458, 1580, 1733, 1763, 1831, 1938, 2331, 2802]. theory-based [2331]. therapies [329]. therapy [818]. Thermal [581, 861, 1351, 1589, 2195, 2254, 2441, 2529, 2909]. Thermal-aware [1351, 2195, 2254, 2441]. thermal-efficient [1589]. THIC [1443]. thick [1178]. thickness [912]. thing [1360]. Things [36, 47, 57, 95, 100, 131, 147, 151, 202, 220, 243, 261, 357, 368, 374, 443, 446, 495, 586, 603, 606, 619, 658, 662, 709, 763, 822, 879, 900, 921, 936, 944, 1004, 1023, 1039, 1041, 1058, 1068, 1081, 1092, 1129, 1170, 1199, 1243, 1244, 1284, 1336, 1350, 1356, 1360, 1374, 1386, 1387, 1433, 1453, 1485, 1563, 1572, 1583, 1590, 1645, 1709, 1714, 1717, 1724, 1730, 1742, 1780, 1864, 1905, 1911, 1933, 1939, 1967, 2013, 2037, 2136, 2169, 2188, 2202, 2231, 2233, 2251, 2252, 2307, 2407, 2480, 2524, 2676, 2703, 2706, 2747, 2807, 2878, 2897, 2904]. Things-aided [100]. things-based [879, 2676]. Things-driven [446]. Thinking [1049]. third [1767, 2207]. third-party [1767]. Thomas [2733]. Thou [1417]. thrashing [2916]. thread [2452, 2916]. Threat [581, 672, 763, 1373, 1662, 1924, 2256]. Threats [120, 374, 486, 2231, 2844]. Three [391, 466, 511, 652, 717, 781, 837, 898, 918, 1193, 1317, 1342, 1456, 1544, 1734, 1920, 2029, 2144, 2448, 2832, 2848, 2852]. Three-dimensional [466, 717, 898, 1456, 2144]. three-factor-based [652]. three-input [511, 837, 2448]. three-layer [918]. three-phase [2848]. Three-phases [2852]. three-stage [391]. three-stream [2832]. three-valued [781]. three-way [2029]. Threshold [417, 890, 1476, 1974, 2253, 2290, 2314, 2440, 2705]. Threshold-based [417]. threshold-proxy [2314, 2440]. Thresholding [783, 925, 1304, 1305, 1475, 2125, 2311, 2425, 2754]. throttling [2916]. throughput [590, 2324, 2454]. thruster [1379]. THS [391]. THS-IDPC [391]. ThunderX2 [650]. thyroid [40, 662, 2038, 2785]. Tianhe [1002, 2586]. Tianhe-3 [1002]. tier [1730, 2160, 2557]. tight [1140]. tighter [516]. tile [699]. tiled [312]. tiling [613, 1094]. tilted [1723]. Time [7, 14, 31, 111, 177, 190, 237, 241, 261, 308, 316, 342, 357, 366, 435, 465, 524, 531, 554, 556, 589, 592, 598, 613, 642, 657, 697, 703, 760, 769, 773, 779, 797, 884, 904, 939, 953, 960, 976, 988, 992, 1067, 1071, 1080, 1081, 1133, 1141, 1187, 1253, 1336, 1409, 1437, 1439, 1560, 1564, 1579, 1600, 1657, 1721, 1735, 1746, 1758, 1790, 1811, 1823, 1892, 1902, 1917, 1934, 1940, 2005, 2024, 2059, 2083, 2085, 2097, 2128, 2141, 2161, 2186, 2189, 2221, 2237, 2240, 2322, 2333, 2379, 2397, 2475, 2483, 2539, 2598, 2619, 2708, 2711, 2726, 2741, 2827, 2849, 2871, 2887, 2909]. time [133, 328, 363, 656, 2249, 2352, 2357, 2543, 2622, 2818, 2868]. Time-aware [363, 2085, 2128, 2475]. Time-critical [342]. time-efficient [1981]. time-evolving [939, 988]. time-frequency [1560]. time-Sensitive [1080, 1133, 1439]. time-series [1409, 2711, 2887]. time-space [613]. time-step [524]. time-step-based [657]. time-to-solution [554]. Time-variant [308]. tiny [2806]. TL [1419, 2433]. TL-CNN-IDS [2433]. TM [1072, 1345]. TM-generation [1072]. TMIS [836, 2658]. TN [1027]. Tnseg [2785]. TOD [528]. Toeplitz [1541]. Token [1062]. Token-based [1062]. tolerance [674, 804, 805, 1342, 1345, 1863, 2120]. Tolerant [511, 754, 837, 842, 1019, 1097, 1347, 1460, 1498, 1544, 1613, 1704, 1790, 1905, 1920, 1930, 2000, 2194, 2253, 2272, 2542, 2627, 2640, 2732, 2768, 2796]. tomography [622, 820, 1462, 1463, 1465, 1740]. tomography-guided [1465]. tone [1323]. Tool [25, 233, 678, 766, 1087, 1108, 1222, 1459, 1598]. tools [979, 2909]. Top [1211, 1751]. Top- [1211]. Topic [407, 693, 1123, 1400, 2515, 2911]. TopicLPRank [2117]. TopicRank [2117]. TopLeaders [408]. topological [203]. topologies [488, 567, 1552, 1570, 2270]. topology [25, 563, 1316, 1542, 1889, 1957, 2106, 2194]. topology- [563]. topology-aware [1889]. topology-based [2194]. topper [1429, 1748, 2405]. TOPSIS [542, 543, 929, 1238, 1397]. tori [886]. Torus [377, 577, 1027]. Total [1495, 2596]. totalistic [952]. touch [459]. tourism [2807]. tourist [269]. tourists [269]. tournament [2575]. toy [1459]. traceability [1953]. traceable [2334, 2410]. traceback [211]. traces [1797, 2347]. tracing [913, 914]. track [2017]. tracker [1125]. Tracking [5, 51, 250, 410, 426, 535, 1060, 1172, 1253, 1278, 1395, 1407, 1518, 1585, 1799, 1874, 1991, 2151, 2377, 2401, 2416, 2469, 2495, 2521, 2539, 2645, 2752]. tracking-based [2401]. tracks [523]. trade [573, 1802, 1891, 2147, 2240]. trade-off [573, 1802, 2240]. trade-offs [1891, 2147]. trading [290, 927, 1149, 1206, 2491]. traditional [2659]. Traffic [391, 425, 551, 563, 634, 636, 687, 694, 697, 702, 736, 826, 872, 900, 904, 1003, 1170, 1239, 1300, 1344, 1352, 1442, 1520, 1617, 1654, 1680, 1785, 1845, 1847, 2059, 2099, 2186, 2341, 2354, 2430, 2460, 2482, 2493, 2561, 2644, 2691, 2808, 2812, 2878]. traffic- [563]. traffic-adaptive [872]. Traffic-aware [1785, 1845]. trained [1324, 1977, 2297]. Training [5, 18, 359, 384, 494, 518, 588, 741, 960, 1006, 1029, 1174, 1501, 1724, 1737, 1837, 1991, 2020, 2049, 2097, 2184, 2185, 2187, 2437, 2545]. trains [2714]. trait [2757]. Trajectory [266, 424, 1724, 2365, 2552]. TRAM [1103]. transaction [2324, 2628]. transactional [347, 1114]. transactions [2201]. transcatheter [919, 1465]. transcendental [1151]. transcoding [481]. transfer [446, 449, 551, 706, 807, 877, 912, 1069, 1079, 1105, 1254, 1419, 1431, 1647, 1695, 1705, 1814, 1963, 2004, 2067, 2079, 2190, 2273, 2404, 2433, 2522, 2653, 2695]. transform [103, 574, 649, 1341, 1473, 1691, 1963, 2115, 2276, 2291, 2538, 2610, 2783, 2855]. transform-based [2276]. transformation [891, 1843, 1860, 2551, 2593]. transformations [2737]. transformed [2130]. Transformer [2002, 2097, 2102, 2218, 2393, 2394, 2519, 2528, 2598, 2790, 2801, 2829, 2850]. transformer-based [2394, 2519]. transformer-cum-graph-based [2393]. transformers [1354]. Transforming [296]. transforms [620, 1769]. Transit [1813]. Transition [415, 977, 1166]. transition-based [1166]. Transitive [422]. Transitivity [891, 2309]. translation [716, 1959, 2445, 2689, 2776]. Transmission [99, 115, 202, 416, 764, 774, 935, 1091, 1381, 1770, 2307, 2418, 2791]. transmitted [2116]. transonic [1743]. transport [349, 479, 955, 1278, 1560]. transportation [337, 808, 833, 947, 1712, 1997, 2725, 2894]. transpose [1394]. transposition [1931]. trapezoidal [2735]. TRAPPY [1339]. travel [997, 998]. traveling [977, 1946, 2248]. travelling [939]. treatment [818, 866, 1141, 1194, 1265, 1462, 1714]. treatments [1425]. Tree [39, 49, 157, 252, 338, 366, 381, 427, 474, 667, 704, 715, 728, 756, 859, 932, 1490, 1554, 1638, 1660, 1893, 2278, 2291, 2539, 2870]. Tree-based [704, 2539]. tree-shaped [2278]. trees [640, 1410, 1497, 1622, 1836, 1931, 2208, 2234, 2640, 2797]. Trend [773, 1940]. trends [169, 777, 2129, 2411, 2573, 2672]. tri [1169, 1990, 2121]. tri-broad [2121]. tri-spectral [1169, 1990]. triangle [422, 2048, 2313]. triangle-free [2048]. Triangular [216]. tridiagonal [2733]. Trigger [2555]. Trigger-free [2555]. triggered [764, 1881]. trimming [1623]. trip [663]. Triple [1417, 2103, 2550, 2837]. Triple-channel [2837]. triple-erasure-correcting [1417]. triplet [2678, 2724, 2873]. tristate [72, 1335]. TRIT [1763]. TRNG [2422]. trojan [465, 1522, 1881]. trolls [2404]. truncated [1533]. truncation [2064]. trunk [1043]. Trust [249, 368, 374, 457, 610, 878, 996, 1179, 1268, 1343, 1350, 1356, 1692, 1873, 1997, 2133, 2201, 2290, 2387, 2390, 2408, 2706, 2713, 2844, 2858]. Trust- [1268]. Trust-based [374, 1873]. Trust-embedded [457]. Trusted [1687, 2207]. trustworthy [375, 2635, 2869]. truthfulness [1339]. TSANET [2894]. TSMSA [1502]. TSUBASA [851]. TTLA [2387]. tuberculosis [449]. tumor [1137, 1210, 1402, 1523, 1808, 2654]. tumors [807, 2125, 2652]. tumour [1099]. Tunable [554]. tuning [582, 657, 1280, 2210, 2420, 2478]. tunnel [1476, 2373]. tunneling [2373]. tuple [2799]. turbo [218]. turbo-coded [218]. turbulence [1196]. turbulent [880]. Turing [498]. turning [1161]. tutorial [1, 1531]. tutoring [588]. Twin [1566].

Twitter [286, 1231, 1617, 2674]. Two [75, 95, 114, 317, 449, 468, 513, 685, 687, 940, 1186, 1209, 1237, 1363, 1385, 1415, 1686, 1734, 1766, 1982, 2033, 2124, 2160, 2162, 2262, 2336, 2359, 2387, 2491, 2544, 2557, 2566, 2605, 2711, 2811, 2875]. Two-channel [2033]. two-cloud [2566]. two-dimensional [940, 1363, 1734, 2124]. Two-factor [2491]. two-fish [1766]. two-layer [114, 1237]. Two-level [95, 513, 1686, 2262]. two-party [1385]. Two-phase [685, 687, 2359]. two-side [2162]. two-sided [75]. Two-stage [317, 449, 468, 1186, 1415, 2711]. two-stream [1209]. two-tier [2160, 2557]. Two-way [2387, 2811, 2875]. twofold [2902]. Type [136, 599, 704, 832, 912, 1034, 1127, 1381, 1548, 2564, 2624, 2905]. type-2 [1127]. typhoon [638].

U [1128, 1402, 1521, 2724]. U-Net [1128, 1402, 1521]. U-shape [2724]. UAV [43, 248, 525, 690, 1071, 1133, 1268, 1446, 1798, 2159, 2330, 2368, 2489]. UAV-assisted [1798, 2330]. UAV-based [1268, 1446]. UAV-borne [1071]. UAV-mounted [2368]. UAVs [420, 720, 1415, 2677, 2908]. ubiquitous [1788, 2330, 2510]. UDP [723]. ulceration [1548]. uLog [1833]. ultra [1121, 1449, 1726, 1891, 2089, 2217, 2816]. ultra-area-efficient [2089]. ultra-dense [1121, 1726, 2816]. ultra-efficient [1891]. ultra-energy-efficient [2217]. ultra-lightweight [1449]. ultrafast [1074]. Ultralightweight [1369, 1370, 2607]. Ultrascale [601]. Ultrasound [103, 684, 912, 930, 1194, 1461]. UML [1331]. UMOTS [1088]. unannotated [496]. unbalanced [1134, 1430, 2395]. uncertain [152, 1751]. uncertainty [1088, 1945, 2184, 2210]. uncertainty-aware [1088]. unclonable [2571]. under-frequency [256]. underlay [2041]. Understanding [120, 1307]. underutilized [514]. Underwater [717, 798, 799, 1826, 2495, 2749]. unequal [70, 95]. Unified [132, 1342, 2364, 2652, 2798]. Uniform [167, 668, 1403, 1792, 2262, 2326, 2381]. uniformity [827]. uniformization [2262]. unifying [20]. uninteresting [787]. Union [2046]. Union-net [2046]. unique [2746]. unit [199, 519, 585, 755, 1790, 1795, 1839, 2415, 2462, 2517, 2588]. unit-based [519]. units [532, 926, 2097]. unity [1438]. univariate [797]. universal [326]. university [74]. unknown [272, 2283, 2368]. unlinkability [2081]. unmanned [720, 2304, 2305, 2756, 2908]. unpaired [789]. unprecedented [1349]. unrelated [441]. unstructured [215, 2123, 2525, 2548, 2775]. Unsupervised [51, 325, 728, 844, 1586, 2131, 2229, 2347, 2393, 2443, 2464, 2743, 2759, 2779, 2839, 2870]. unwrapping [874]. update [413, 858, 1417, 1936, 2019, 2499]. Updating [282, 1015, 2520]. upgrade [798, 799]. uplift [607]. uplink [1203]. upper [516, 2078]. UPR [1010]. uprighting [2888]. UPSRVNet [2607]. URANS [1743]. URAP [1449]. urban [56, 1203, 1267, 1466, 1998, 2330, 2561, 2892]. URLLC [1153]. URMP [2337]. usability [245, 768]. usage [751, 815, 1138, 2389]. USCDRP [70]. Use [14, 58, 557, 650, 853, 1037, 1491, 1734, 1880, 2021, 2083, 2221, 2821]. used [509, 655, 927, 1087, 1585, 2701]. useful [1941, 2218, 2490, 2734]. useless [393]. User [188, 273, 308, 433, 481, 565, 610, 651, 723, 753, 896, 936, 1257, 1301, 1389, 1390, 1512, 1525, 1584, 1712, 1749, 1794, 1969, 2080, 2201, 2222, 2597, 2608, 2688, 2720, 2728, 2795]. user-based [1390]. user-centric [936]. user-friendly [896]. user-level [723]. user-user [753]. users [1058, 2368]. uses [1806]. Using [25, 66, 89, 112, 121, 182, 195, 217, 221, 315, 348, 366, 449, 511, 558, 629, 660, 703, 724, 771, 784, 863, 871, 907, 923, 928, 931, 944, 1060, 1092, 1105, 1108, 1128, 1158, 1164, 1168, 1182, 1201, 1215, 1347, 1376, 1382, 1391, 1392, 1399, 1421, 1444, 1454, 1501, 1521, 1554, 1560, 1661, 1733, 1748, 1768, 1816, 1852, 1891, 1902, 1914, 1915, 1961, 1970, 1988, 2025, 2043, 2051, 2182, 2190, 2261, 2280, 2342, 2363, 2387, 2427, 2493, 2505, 2549, 2565, 2581, 2591, 2628, 2630, 2634, 2665, 2677, 2710, 2740, 2761, 2773, 2792, 2797, 2830, 2832, 2844, 2886, 2919]. using [88, 127, 143, 239, 263, 272, 326, 351, 356, 362, 365, 397, 415, 421, 432, 538, 553, 561, 579, 590, 595, 602, 639, 651, 666, 675, 682, 693, 720, 725, 742, 757, 783, 809, 836, 874, 880, 889, 910, 921, 934, 941, 951, 975, 980, 1024, 1041, 1087, 1109, 1169, 1191, 1224, 1286, 1303, 1307, 1308, 1324, 1325, 1371, 1395, 1402, 1407, 1439, 1442, 1588, 1660, 1669, 1686, 1692, 1720, 1724, 1766, 1818, 1855, 1860, 1893, 1895, 1908, 1921, 1943, 1960, 1990, 2239, 2264, 2336, 2345, 2373, 2409, 2456, 2481, 2524, 2538, 2566, 2718, 2752, 2810, 2854, 2861, 2863]. using [16, 37, 52, 157, 208, 245, 255, 267, 269, 270, 308, 333, 340, 345, 374, 419, 423, 424, 446, 557, 585, 622, 643, 718, 733, 741, 749, 752, 832, 834, 840, 853, 855, 887, 898, 943, 945, 1016, 1057, 1074, 1110, 1186, 1206, 1240, 1254, 1257, 1269, 1290, 1297, 1341, 1390, 1429, 1431, 1456, 1474, 1484, 1491, 1492, 1507, 1525, 1546, 1608, 1649, 1662, 1745, 1890, 1941, 1977, 1978, 1989, 2062, 2106, 2126, 2196, 2197, 2250, 2262, 2270, 2294, 2386, 2412, 2430, 2431, 2468, 2470, 2485, 2496, 2504, 2506, 2551, 2571, 2577, 2639, 2717, 2726, 2728, 2744, 2784, 2802, 2878]. using [10, 19, 46, 63, 71, 185, 186, 219, 222, 224, 304, 305, 334, 341, 344, 347, 352, 355, 451, 485, 544, 565, 640, 663, 753, 773, 788, 800, 801, 807, 821, 826, 868, 875, 884, 912, 973, 984, 985, 1005, 1011, 1050, 1075, 1077, 1091, 1100, 1166, 1175, 1179, 1231, 1255, 1268, 1355, 1365, 1368, 1384, 1438, 1500, 1548, 1598, 1630, 1656, 1664, 1716, 1814, 1820, 1821, 1841, 1845, 1879, 1910, 1911, 1935, 1949, 1976, 2007, 2039, 2070, 2089, 2099, 2107, 2125, 2158, 2198, 2257, 2260, 2337, 2405, 2448, 2492, 2515, 2588, 2600, 2636, 2817, 2843, 2914]. using [47, 57, 144, 196, 213, 215, 271, 321, 342, 359, 368, 396, 435, 477, 542, 547, 627, 646, 647, 669, 672, 685, 754, 780, 796, 806, 811, 818, 879, 918, 920, 959, 981, 1031, 1059, 1163, 1165, 1188, 1223, 1233, 1294, 1323, 1329, 1379, 1398, 1412, 1433, 1566, 1572, 1586, 1593, 1595, 1631, 1642, 1666, 1731, 1742, 1746, 1757, 1786, 1858, 1888, 1919, 1925, 1955, 1963, 2059, 2066, 2092, 2152, 2155, 2168, 2191, 2201, 2209, 2258, 2276, 2312, 2325, 2418, 2422, 2426, 2433, 2450, 2526, 2570, 2605, 2625, 2629, 2631, 2670, 2691, 2747, 2759, 2760, 2776, 2777, 2798, 2840, 2849, 2874, 2891]. using [1155, 2149, 2479]. utility [516, 1357, 2720]. Utilization [456, 513, 928, 1909, 2424, 2895, 2906]. utilization-based [513]. Utilization-prediction-aware [2906]. utilize [280, 1012]. utilizing [2063, 2273, 2635, 2869, 2907]. UVaFTLE [2141].

V [2415]. V2I [2080]. V2V [1264]. vacation [665, 2266]. Vaccine [2426]. VAE [2244, 2280]. VAL [935]. validation [143, 1897]. validity [2329]. VALKYRIE [1889]. value [122, 177, 850, 898, 930, 1192, 1589, 2257, 2420, 2833]. value-based [1589]. Valued [781, 1769, 1892, 1943, 2516]. values [328]. VANET [1014, 1681, 1848, 2330, 2410]. VANETs [777, 1075, 2518]. variability [841]. Variable [72, 614, 1151, 1732, 2096, 2865]. variable-length [72, 2096]. variable-size [2865]. variables [149]. variant [308]. variants [1115, 2788]. variation [1979, 2596]. Variational [1162, 1685, 2280, 2839]. various [1308, 1745, 2248, 2364]. varying [1036, 1609, 2332]. vascular [2670]. vasovagal [818]. vault [2245]. vax [2674]. VBSF [90]. vCPUs [548]. Vector [85, 87, 89, 90, 172, 461, 499, 746, 845, 851, 891, 960, 1025, 1515, 1566, 1630, 2034, 2179, 2223, 2824]. vector-space-based [845]. vectorization [2161, 2332]. Vectorizing [238, 1515]. vectors [1289]. VEF [1797]. vegetable [205]. vehicle [337, 349, 564, 884, 947, 1170, 1278, 1468, 1575, 1703, 1998, 2325, 2377, 2479, 2691]. Vehicles [121, 158, 436, 616, 669, 720, 1031, 1239, 1365, 1369, 1370, 1936, 2304, 2620, 2734, 2908]. vehicular [346, 349, 796, 878, 1343, 1661, 2003, 2093, 2401, 2419, 2561, 2793]. vein [1868]. velocity [1511]. ventilation [2005]. ventricular [906]. Verifiable [2410, 2532]. verification [624, 673, 803, 1281, 1366, 2158, 2344, 2503, 2681]. verified [1068]. verifying [642]. versa [1073, 2035]. verse [1106, 1125, 1484]. version [961, 1566, 1636, 2140]. versions [1418, 2248]. versus [785, 1802]. Vertex [295, 2309, 2513]. vertex-cover [2513]. vertical [399, 801]. vertically [2323]. very [2058]. vessel [338, 673, 2905]. vessels [803]. VGL [851]. via [137, 190, 419, 461, 681, 739, 1210, 1252, 1284, 1314, 1366, 1385, 1453, 1742, 1856, 1886, 2078, 2121, 2142, 2357, 2358, 2403, 2654, 2728, 2757, 2916]. viability [1003, 1763]. vibration [419, 2005]. vice [1073, 2035]. victim [512]. Video [5, 42, 51, 353, 440, 481, 770, 876, 929, 1044, 1064, 1071, 1209, 1361, 1376, 1390, 1395, 1556, 1585, 1758, 1775, 1811, 1991, 2743, 2761, 2861, 2919]. video-based [2861]. Video-SAR [1071]. videos [381, 788, 1551]. Vietnamese [73]. view [32, 466, 965, 2300, 2468]. viewpoint [42]. violation [386]. violations [298, 2059]. violence [1758, 1811]. VIoT [2607]. Virtual [22, 29, 30, 33, 208, 227, 265, 283, 382, 389, 397, 469, 546, 548, 604, 742, 779, 813, 875, 949, 962, 963, 1028, 1080, 1086, 1126, 1280, 1438, 1467, 1624, 1642, 1863, 1868, 1889, 1989, 2027, 2135, 2254, 2272, 2361, 2436, 2441, 2497, 2600, 2606, 2714, 2825, 2891]. virtualization [387, 1591]. virtualized [129, 130, 983, 1489]. virus [1273]. viscoacoustic [2264]. viscosity [2738]. viscous [1110]. visibility [2810]. Visible [1203, 1793, 2881]. Vision [46, 1172, 1353, 1433, 1567, 1703, 1934, 2159, 2297, 2790]. Vision-based [1172, 1934]. vision-conditioned [2297]. Visual [59, 157, 270, 711, 825, 934, 1064, 1098, 1135, 1400, 1411, 1523, 1614, 2297, 2567, 2683, 2721]. visual-guided [2721]. visualization [215, 492, 620, 943, 1090, 1196, 1434, 1914, 1955]. Visualized [200]. VLSI [225, 2362]. VM [373, 386, 514, 749, 1176, 1345, 1346]. VMD [2755]. VMs [1741]. VNF [2389]. voice [1323, 1615]. VoIP [137, 1591]. volatile [852]. volume [183, 508, 2430, 2798]. volumes [1290, 1805]. voluminous [1791]. Voronoi [1360]. voter [1335]. VR [1868]. Vulnerability [119, 1216, 1722, 1766, 2032, 2417]. Vx [1505].

W [2009]. WaaS [1346]. WaaS-cloud [1346]. wafer [2614]. Wailingding [176]. wait [1903]. wait-free [1903]. walk [790, 1593, 1666]. walking [985]. warehouse [2617]. warehouses [677]. warehousing [190]. warning [154, 1477]. warp [135, 2246]. Wasserstein [1837, 1876]. waste [2659]. wastewater [1714, 1932]. water [197, 219, 280, 659, 1715, 2665]. waterlogging [154]. watermarking [1085, 2276, 2676]. waterpixel [901]. wave [711, 1602, 1725, 2029, 2264, 2888]. wave-induced [2888]. wavefield [613]. waveform [613]. wavelength [332]. wavelength-routed [332]. wavelet [568, 620, 667, 1473, 1860, 1894, 2115, 2276, 2291, 2538]. wavelets [1507]. way [422, 2029, 2387, 2811, 2875]. WBAN [2863]. WBANs [2195]. WCP [104]. WCP-RNN [104]. WDM [1957]. weakening [1451]. Weakly [1361, 1972, 2700, 2890]. wear [1108, 1403]. wear-leveling [1403]. wearable [1479]. Weather [491, 1368, 1438, 1818, 2186]. web [109, 207, 242, 413, 487, 535, 570, 1061, 1142, 1286, 1498, 1517, 1817, 2114, 2655, 2679, 2872]. website [535]. websites [160]. wedge [717]. wedge-shaped [717]. weight [1511, 1524, 1692, 1803, 2601, 2665, 2788]. weight-based [1524, 2601]. weight-oriented [1803]. weighted [76, 306, 739, 753, 905, 1259, 1276, 1582, 1686, 1725, 1962, 2331, 2341]. weighting [651, 973, 1784, 2329, 2397, 2770]. weighting-based [2770]. weights [325]. Whale [1277, 1442, 1878, 1900, 1928, 2155, 2325, 2327, 2520]. Wheel [150, 962, 963, 1640]. Wheel-based [962, 963]. wheeled [1374]. wheezing [1827]. White [115, 187, 1457, 2582]. White-box [1457]. white-shell [115]. Wi [2068, 2196]. Wi-Fi [2196]. wide [926, 1032, 1540, 1971, 2162]. wide-area [1032, 1540, 2162]. widespread [691]. WiFi [2881]. wildfires [871]. wind [185, 1252, 1571]. wind-integrated [1571]. window [415, 1011, 2088, 2092, 2379, 2842]. window-based [415, 1011]. Windowing [558]. windows [884]. Winograd [2181]. WIoU [2829]. wirelength [683, 2208]. Wireless [22, 54, 66, 70, 110, 124, 184, 216, 219, 257, 307, 333, 342, 350, 385, 452, 555, 639, 668, 756, 764, 859, 868, 873, 895, 908, 962, 963, 984, 1034, 1035, 1040, 1083, 1141, 1147, 1187, 1198, 1232, 1245, 1249, 1268, 1274, 1295, 1408, 1429, 1460, 1472, 1492, 1524, 1558, 1582, 1620, 1686, 1817, 1888, 1900, 1976, 1980, 2066, 2158, 2194, 2229, 2243, 2270, 2272, 2398, 2413, 2549, 2601, 2609, 2663, 2749, 2782, 2803, 2836, 2891, 2895, 2915]. wise [897, 1089, 1551, 2075, 2329]. within [507, 1279, 2333, 2345, 2386, 2450, 2548]. without [1035, 1214, 1701, 2116, 2365, 2574]. WK [1410]. WK-recursive [1410]. WLANs [2500]. WMNs [19]. WNoC [528]. WOA [2340]. Wolf [47, 1077, 1484, 1866, 1929, 2034, 2193, 2265, 2502, 2737]. women [1328]. word [172, 428, 429, 891, 1072, 1419, 2198, 2230, 2239]. Word2Vec [987]. words [496, 688, 1648, 2296]. work [1696]. worker [1258, 2301]. workflow [6, 14, 548, 769, 875, 933, 969, 1070, 1205, 1619, 1776, 1835, 2057, 2206, 2432, 2470, 2800, 2841, 2901]. workflow-as-a-service [1070]. workflows [34, 155, 397, 629, 824, 990, 1012, 1021, 1224, 1346, 1688, 1690, 1704, 1732, 1802, 1823, 2455, 2487]. working [841, 882, 1102, 2266, 2375]. Workload [731, 928, 946, 1589, 1731, 1895, 2226, 2347, 2406, 2710]. workload-aware [2226]. workload-driven [2710]. workloads [82, 549, 650, 765, 1773, 1795, 1797, 2283]. world [1898, 1976]. worm [771]. worn [840]. write [2288]. WSMP [2246]. WSN [167, 202, 228, 365, 505, 719, 945, 1000, 1375, 1454, 1653, 1798, 1949, 1997, 2022, 2202, 2307, 2694]. WSN-enabled [1997]. WSNs [562, 685, 708, 873, 1045, 1554, 1928, 2340].

X [1746, 2570]. X-ray [1746, 2570]. XAI [2456]. XAI-reduct [2456]. Xeon [234, 615, 880]. XGBoost [2842, 2884]. XGBoost-based [2842]. Xilinx [601]. XMI [1331]. XML [732, 1083]. XNOR [1074]. XOR [1474]. XPath [1260].

Yang [2680]. YARN [785, 2544]. Yates [2434]. YCbCr [203]. years [891]. yield [863]. Yin [2680]. yoga [1254]. YOLO [1031, 1846, 2644, 2684, 2696]. YOLO-ARGhost [2684]. YOLO-SG [2644]. YOLOOD [2339]. YOLOv4 [2484, 2806]. YOLOv4-tiny [2806]. YOLOv5 [2113, 2186]. YOLOv5m [2861]. YOLOX [2675, 2850].

Z [1335]. Z-Voter [1335]. Zehnder [811]. ZEN [2719]. zero [980, 1814, 2225, 2721, 2844]. zero-shot [1814, 2225, 2721]. zerotrust [2844]. ZFS [2065]. ZFS-based [2065]. ZigBee [2196]. Ziv [1322]. Zone [973, 1014, 2089].

References

Dolbeau:2018:TPF

[1] Romain Dolbeau. Theoretical peak FLOPS per instruction set: a tutorial. The Journal of Supercomputing, 74(3): 1341–1377, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [1508].

Wu:2018:PPA

[2] Libing Wu, Jing Wang, Sherali Zeadally, and Debiao He. Privacypreserving auditing scheme for shared data in public clouds. The Journal of Supercomputing, 74(11):6156– 6183, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See comment [279].

Li:2019:CFF

[3] Kun Li, Shigang Li, Shan Huang, Yifeng Chen, and Yunquan Zhang. Correction to: FastNBL: fast neighbor lists establishment for molecular dynamics simulation based on bitwise operations. The Journal of Supercomputing, 75(12):8339–8340, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-02956-w.pdf. See [276].

Kim:2019:SCM

[4] Hyun-Woo Kim, Jungho Kang, and Young-Sik Jeong. Simulator considering modeling and performance evaluation for high-performance computing of collaborative-based mobile cloud infrastructure. The Journal of Supercomputing, 75(8):4459–4471, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See retraction note [1992].

Hui:2019:MVT

[5] Qiuli Hui. Motion video tracking technology in sports training based on mean-shift algorithm. The Journal of Supercomputing, 75(9):6021– 6037, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See retraction note [1991].

Masdari:2020:ETW

[6] Mohammad Masdari and Mehran Zangakani. Efficient task and workflow scheduling in inter-cloud environments: challenges and opportunities. The Journal of Supercomputing, 76(1):499– 535, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rybintsev:2020:OPL

[7] Vladimir O. Rybintsev. Optimizing the parameters of the Lustre-file-system-based HPC system for reverse time migration. The Journal of Supercomputing, 76(1):536–548, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Huang:2020:SEM

[8] Jianqiang Huang, Wei Qin, Xiaoying Wang, and Wenguang Chen. Survey of external memory large-scale graph processing on a multi-core system. The Journal of Supercomputing, 76(1):549– 579, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yun:2020:EDP

[9] Ji-Tae Yun, Su-Kyung Yoon, Jeong-Geun Kim, and Shin-Dug Kim. Effective data prediction method for inmemory database applications. The Journal of Supercomputing, 76(1):580– 601, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

NezhadShokouhi:2020:SDP

[10] Mohammad Mahdi NezhadShokouhi, Mohammad Ali Majidi, and Abbas Rasoolzadegan. Software defect prediction using over-sampling and feature extraction based on Mahalanobis distance. The Journal of Supercomputing, 76(1): 602–635, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2020:LBD

[11] Jiagao Wu, Lu Shen, and Linfeng Liu. LSH-based distributed similarity indexing with load balancing in highdimensional space. The Journal of Supercomputing, 76(1):636–665, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khan:2020:OHS

[12] Minhaj Ahmad Khan. Optimized hybrid service brokering for multi-cloud architectures. The Journal of Supercomputing, 76(1):666–687, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fadishei:2020:CIA

[13] Hamid Fadishei and Azadeh Soltani. The curse of indecomposable aggregates for big data exploratory analysis with a case for frequent pattern cubes. The Journal of Supercomputing, 76(1): 688–707, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahmad:2020:UCB

[14] Saima Gulzar Ahmad, Hikmat Ullah Khan, Samia Ijaz, and Ehsan Ullah Munir. Use case-based evaluation of workflow optimization strategy in real-time computation system. The Journal of Supercomputing, 76(1): 708–725, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mohtavipour:2020:LEP

[15] Seyed Mehdi Mohtavipour and Hadi Shahriar Shahhoseini. A link-elimination partitioning approach for application graph mapping in reconfigurable computing systems. The Journal of Supercomputing, 76(1):726–754, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumar:2020:ICC

[16] R. Dhilip Kumar and K. S. Vishvaksenan. Interference cancellation in cognitive radio-based MC–CDMA system using pre-coding technique. The Journal of Supercomputing, 76(1):1– 15, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2020:CLP

[17] Yung-Hui Chen, Wei-Chun Lee, Chun-Hsiung Tseng, Lawrence Y. Deng, Chuan-Yu Chang, and Long-He Lee. Cognitive learning performance assessment and analysis with CSCL applied on the NetGuru platform and CSPL applied on the TAoD platform for the network experiment class.

The Journal of Supercomputing, 76(1): 16–46, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2020:BBO

[18] Eunju Yang, Dong-Ki Kang, and Chan-Hyun Youn. BOA: batch orchestration algorithm for straggler mitigation of distributed DL training in heterogeneous GPU cluster. The Journal of Supercomputing, 76(1):47–67, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nassiri:2020:JEQ

[19] Mohammad Nassiri and Reza Mohammadi. A joint energy- and QoS-aware routing mechanism for WMNs using software-defined networking paradigm. The Journal of Supercomputing, 76(1): 68–86, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kargar:2020:IMQ

[20] Masoud Kargar, Ayaz Isazadeh, and Habib Izadkhah. Improving the modularization quality of heterogeneous multi-programming software systems by unifying structural and semantic concepts. The Journal of Supercomputing, 76(1):87–121, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Naghshnejad:2020:HSP

[21] Mina Naghshnejad and Mukesh Singhal. A hybrid scheduling platform: a runtime prediction reliability aware scheduling platform to improve HPC scheduling performance. The Journal of Supercomputing, 76(1):122–149, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Amini:2020:EED

[22] S. M. Amini, A. Karimi, and M. Esnaashari. Energy-efficient data dissemination algorithm based on virtual hexagonal cell-based infrastructure and multi-mobile sink for wireless sensor networks. The Journal of Supercomputing, 76(1):150–173, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Maghsoudloo:2020:EHI

[23] M. Maghsoudloo and N. Khoshavi. Elastic HDFS: interconnected distributed architecture for availabilityscalability enhancement of large-scale cloud storages. The Journal of Supercomputing, 76(1):174–203, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

deMiras:2020:FDB

[24] Juan Ruiz de Miras. Fast differential box-counting algorithm on GPU. The Journal of Supercomputing, 76(1): 204–225, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bhih:2020:OTR

[25] Amhmed Bhih, Princy Johnson, and Martin Randles. An optimisation tool for robust community detection algorithms using content and topology information. The Journal of Supercomputing, 76(1):226–254, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-03018-x.pdf.

Iserte:2020:SEP

[26] Sergio Iserte and Krzysztof Rojek. An study of the effect of process malleability in the energy efficiency on GPUbased clusters. The Journal of Supercomputing, 76(1):255–274, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumar:2020:IEA

[27] Gulshan Kumar. An improved ensemble approach for effective intrusion detection. The Journal of Supercomputing, 76(1):275–291, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2020:ENI

[28] Misun Yu, Yu-Seung Ma, and Doo-Hwan Bae. Efficient noise injection for exposing hidden data races. The Journal of Supercomputing, 76(1):292– 323, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [140].

Bermejo:2020:VMC

[29] Belen Bermejo and Carlos Juiz. Virtual machine consolidation: a systematic review of its overhead influencing factors. The Journal of Supercomputing, 76(1): 324–361, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rahmani:2020:BAV

[30] Somayeh Rahmani, Vahid Khajehvand, and Mohsen Torabian. Burstinessaware virtual machine placement in cloud computing systems. The Journal of Supercomputing, 76(1):362–387, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Farhat:2020:RRL

[31] Peter Farhat, Hani Sami, and Azzam Mourad. Reinforcement R-learning model for time scheduling of ondemand fog placement. The Journal of Supercomputing, 76(1):388–410, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:MVA

[32] Yan Li, Chun zi Wang, and Ying chao Li. Multi-view analysis method for robust suppression of attack diffusion. The Journal of Supercomputing, 76(1): 411–426, January 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Braiki:2020:FLB

[33] Khaoula Braiki and Habib Youssef. Fuzzy-logic-based multi-objective bestfit-decreasing virtual machine reallocation. The Journal of Supercomputing, 76(1):427–454, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Qin:2020:EAS

[34] Yao Qin, Hua Wang, Shanwen Yi, Xiaole Li, and Linbo Zhai. An energyaware scheduling algorithm for budgetconstrained scientific workflows based on multi-objective reinforcement learning. The Journal of Supercomputing, 76(1):455–480, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:CDE

[35] Hongjian Li, Yuyan Zhao, and Shuyong Fang. CSL-driven and energyefficient resource scheduling in cloud data center. The Journal of Supercomputing, 76(1):481–498, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ramirez-Gonzalez:2020:PSS

[36] Gustavo Ramirez-Gonzalez and Enas Adbulhay. Preface on special section on cognitive computing for emerging Internet of Things. The Journal of Supercomputing, 76(2):1082–1085, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-02944-0.pdf.

Mohammed:2020:DSS

[37] Mazin Abed Mohammed, Mohd Khanapi Abd Ghani, N. Arunkumar, Raed Ibraheem Hamed, Salama A. Mostafa, Mohamad Khir Abdullah, and M. A. Burhanuddin. Decision support system for nasopharyngeal carcinoma discrimination from endoscopic images using artificial neural network. The Journal of Supercomputing, 76(2): 1086–1104, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2020:MLA

[38] Zhibin Zhang and Deyu Li. Multilabel algorithm based on rough set of fractal dimension attribute. The Journal of Supercomputing, 76(2):1105– 1115, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zeng:2020:DSQ

[39] Yifu Zeng, Yantao Zhou, and Fei Zheng. Data skyline query protocol based on parallel genetic improvement decision tree. The Journal of Supercomputing, 76(2):1116–1127, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shankar:2020:OFB

[40] K. Shankar, S. K. Lakshmanaprabu, Deepak Gupta, Andino Maseleno, and Victor Hugo C. de Albuquerque. Optimal feature-based multi-kernel SVM approach for thyroid disease classification. The Journal of Supercomputing, 76(2):1128–1143, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See retraction note [2038].

Zhou:2020:MCD

[41] Qianming Zhou and Jin Xie. Mobile client data security storage protocol based on multifactor node evaluation. The Journal of Supercomputing, 76(2):1144–1158, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ang:2020:SLR

[42] Li Ang and Zheng Bao-Yu. Sparse low-rank correction scalable video coding based on interactive region of interest viewpoint. The Journal of Supercomputing, 76(2):1159–1169, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yan:2020:GIA

[43] Fanlei Yan. Gauss interference ant colony algorithm-based optimization of UAV mission planning. The Journal of Supercomputing, 76(2):1170– 1179, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bharathi:2020:OAB

[44] G. P. Bharathi and K. Meena Alias Jeyanthi. An optimization algorithmbased resource allocation for cooperative cognitive radio networks. The Journal of Supercomputing, 76(2): 1180–1200, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2020:ESM

[45] Jing Yang, Mingyu Fan, and Guangwei Wang. Encryption scheme with mixed homomorphic signature based on message authentication for digital image. The Journal of Supercomputing, 76(2):1201–1211, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pereira:2020:AQS

[46] Renato F. Pereira, Valberto E. R. da Silva Filho, Lorena B. Moura, N. Arun Kumar, Auzuir R. de Alexandria, and Victor Hugo C. de Albuquerque. Automatic quantification of spheroidal graphite nodules using computer vision techniques. The Journal of Supercomputing, 76(2):1212– 1225, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sharma:2020:HTC

[47] Prerna Sharma, Apoorva Gupta, Aastha Aggarwal, Deepak Gupta, Ashish Khanna, Aboul Ella Hassanien, and Victor Hugo C. de Albuquerque. The health of things for classification of protein structure using improved grey wolf optimization. The Journal of Supercomputing, 76(2):1226–1241, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hemanth:2020:ARS

[48] Jude D. Hemanth, Utku Kose, Omer Deperlioglu, and Victor Hugo C. de Albuquerque. An augmented realitysupported mobile application for diagnosis of heart diseases. The Journal of Supercomputing, 76(2):1242– 1267, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2020:DTD

[49] Mengrui Wang. Dominator tree data flow cognitive analysis for green public building design. The Journal of Supercomputing, 76(2):1268–1276, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sugumar:2020:MSR

[50] T. N. Sugumar and N. Rajam Ramasamy. mDesk: a scalable and reliable hypervisor framework for effective provisioning of resource and downtime reduction. The Journal of Supercomputing, 76(2):1277–1292, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tiancheng:2020:UVM

[51] Wang Tiancheng. Unsupervised video multi-target tracking based on fast resampling particle filter. The Journal of Supercomputing, 76(2):1293– 1304, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Marques:2020:NCC

[52] João Alexandre Lobo Marques, Paulo C. Cortez, João P. V. Madeiro, Victor Hugo C. de Albuquerque, Simon James Fong, and Fernando S. Schlindwein. Nonlinear characterization and complexity analysis of cardiotocographic examinations using entropy measures. The Journal of Supercomputing, 76(2): 1305–1320, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zeng:2020:FCB

[53] Yifu Zeng, Yantao Zhou, Xu Zhou, and Fei Zheng. Fuzzy clusteringbased skyline query preprocessing algorithm for large-scale flow data analysis. The Journal of Supercomputing, 76(2):1321–1330, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Weizheng:2020:QAM

[54] Li Weizheng and Tu Xiumei. Quality analysis of multi-sensor intrusion detection node deployment in homogeneous wireless sensor networks. The Journal of Supercomputing, 76(2):1331– 1341, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zaiyi:2020:NSS

[55] Pu Zaiyi. Network security situation analysis based on a dynamic Bayesian network and phase space reconstruction. The Journal of Supercomputing, 76(2):1342–1357, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:APD

[56] Anonymous. Advanced parallel and distributed computing for big urban data. The Journal of Supercomputing, 76(2):755, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-020-03179-0.pdf.

Venkatraman:2020:SMC

[57] S. Venkatraman, B. Surendiran, and P. Arun Raj Kumar. Spam e-mail classification for the Internet of Things environment using semantic similarity approach. The Journal of Supercomputing, 76(2):756–776, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ye:2020:LUC

[58] Yaqin Ye, Ying An, Bo Chen, JunJue Wang, and Yingqiang Zhong. Land use classification from social media data and satellite imagery. The Journal of Supercomputing, 76(2):777–792, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lu:2020:EVA

[59] Qiang Lu, Wenqiang Xu, Haibo Zhang, Qingpeng Tang, Jie Li, and Rui Fang.

ElectricVIS: visual analysis system for power supply data of smart city. The Journal of Supercomputing, 76(2):793– 813, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Liu:2020:PCB

[60] Xiaojun Liu, Ning Yang, Yu Jiang, Lichuan Gu, and Xianzhang Shi. A parallel computing-based deep attention model for named entity recognition. The Journal of Supercomputing, 76(2): 814–830, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2020:DTA

[61] Zaijun Wang, Jinlin Zhu, XiaoPing Guo, YunTing Ma, and Zifan Li. Distributed task allocation method based on self-awareness of autonomous robots. The Journal of Supercomputing, 76(2):831–843, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Leng:2020:DRH

[62] Xiaokun Leng, Songhao Piao, Lin Chang, Zhicheng He, and Zheng Zhu. Dynamic running hexapod robot based on high-performance computing. The Journal of Supercomputing, 76(2):844– 857, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Patro:2020:EDO

[63] Kiran Kumar Patro, Surya Prakasa Rao Reddi, S. K. Ebraheem Khalelulla, P. Rajesh Kumar, and K. Shankar. ECG data optimization for biometric human recognition using statistical distributed machine learning algorithm. The Journal of Supercomputing, 76(2): 858–875, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Feng:2020:SAC

[64] Jinhong Feng, Jundong Zhang, Chuan Wang, and Minyi Xu. Self-adaptive collective intelligence-based mutation operator for differential evolution algorithms. The Journal of Supercomputing, 76(2):876–896, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dai:2020:ABN

[65] Weihuang Dai and Ping Hu. Application of BP neural network in the analytic hierarchy process of person-post evaluation model. The Journal of Supercomputing, 76(2):897–914, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chinnasamy:2020:SDR

[66] N. V. Chinnasamy and A. Senthilkumar. Secured distributed routing technique using extended DART and table elimination (ET-DART) technique in wireless sensor networks environment. The Journal of Supercomputing, 76(2): 915–931, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Guo:2020:DMA

[67] Aiping Guo, Ajuan Jiang, Jie Lin, and Xiaoxiao Li. Data mining algorithms for bridge health monitoring: Kohonen clustering and LSTM prediction approaches. The Journal of Supercomputing, 76(2):932–947, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-03045-8.pdf.

Xia:2020:CEL

[68] Yingchun Xia, Xingyue Wang, Lichuan Gu, Qijuan Gao, Jun Jiao, and Chao Wang. A collective entity linking algorithm with parallel computing on largescale knowledge base. The Journal of Supercomputing, 76(2):948–963, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abdel-Basset:2020:RLP

[69] Mohamed Abdel-Basset, Mumtaz Ali, and Asma Atef. Resource levelling problem in construction projects under neutrosophic environment. The Journal of Supercomputing, 76(2):964– 988, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vijayalakshmi:2020:UUS

[70] V. Vijayalakshmi and A. Senthilkumar. USCDRP: unequal secure clusterbased distributed routing protocol for wireless sensor networks. The Journal of Supercomputing, 76(2):989–1004, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Qi:2020:PNI

[71] Yongjun Qi, Junhua Gu, Weixun Li, Zepei Tian, Yajuan Zhang, and Juanping Geng. Pulmonary nodule image super-resolution using multi-scale deep residual channel attention network with joint optimization. The Journal of Supercomputing, 76(2):1005– 1019, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2020:VLC

[72] Tian Chen, Yongsheng Zuo, Xin An, and Fuji Ren. A variable-length compatible compression scheme based on tristate signals. The Journal of Supercomputing, 76(2):1020–1033, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gao:2020:CVB

[73] Shengxiang Gao, Zhengtao Yu, Yunlong Li, Yusen Wang, and Yafei Zhang. Chinese–Vietnamese bilingual news event summarization based on distributed graph ranking. The Journal of Supercomputing, 76(2):1034– 1048, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2020:PIA

[74] Zne-Jung Lee and Chou-Yuan Lee. A parallel intelligent algorithm applied to predict students dropping out of university. The Journal of Supercomputing, 76(2):1049–1062, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2020:STS

[75] Rong Zhao, Maozhu Jin, Peiyu Ren, and Qian Zhang. Stable two-sided satisfied matching for ridesharing system based on preference orders. The Journal of Supercomputing, 76(2):1063– 1081, February 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2020:ASB

[76] Mincheng Chen, Jingling Yuan, Dongling Liu, and Tao Li. An adaption scheduling based on dynamic weighted random forests for load demand forecasting. The Journal of Supercomputing, 76(3):1735–1753, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shafqat:2020:BDA

[77] Sarah Shafqat, Saira Kishwer, Raihan Ur Rasool, Junaid Qadir, Tehmina Amjad, and Hafiz Farooq Ahmad. Big data analytics enhanced healthcare systems: a review. The Journal of Supercomputing, 76(3):1754–1799, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Groppe:2020:EMF

[78] Sven Groppe. Emergent models, frameworks, and hardware technologies for big data analytics. The Journal of Supercomputing, 76(3):1800–1827, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Andrade:2020:DED

[79] Ermeson Andrade and Bruno Nogueira. Dependability evaluation of a disaster recovery solution for IoT infrastructures. The Journal of Supercomputing, 76(3):1828–1849, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2020:PLD

[80] Yue Zhao, Kenji Yoshigoe, Mengjun Xie, Jiang Bian, and Ke Xiong. L-PowerGraph: a lightweight distributed graph-parallel communication mechanism. The Journal of Supercomputing, 76(3):1850–1879, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abrol:2020:SSF

[81] Preeti Abrol and Savita Gupta. Social spider foraging-based optimal resource management approach for future cloud. The Journal of Supercomputing, 76(3): 1880–1902, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Naeen:2020:SPB

[82] Hossein Monshizadeh Naeen, Esmaeil Zeinali, and Abolfazl Toroghi Haghighat. A stochastic process-based server consolidation approach for dynamic workloads in cloud data centers. The Journal of Supercomputing, 76(3): 1903–1930, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhou:2020:FGS

[83] Mosong Zhou, Xiaoshe Dong, Heng Chen, and Xingjun Zhang. Finegrained scheduling in multi-resource clusters. The Journal of Supercomputing, 76(3):1931–1958, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:ENa

[84] Anonymous. Editor’s note. The Journal of Supercomputing, 76(3):1959, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

com/content/pdf/10.1007/s11227-
020-03238-6.pdf.

Barredo:2020:EAM

[85] Adrian Barredo, Juan M. Cebrian, Mateo Valero, Marc Casas, and Miquel Moreto. Efficiency analysis of modern vector architectures: vector ALU sizes, core counts and clock frequencies. The Journal of Supercomputing, 76(3): 1960–1979, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Saadi:2020:EGB

[86] Hocine Saadi, Nadia Nouali Taboudjemat, Abdellatif Rahmoun, Baldomero imbernón, Horacio Pérez-Sánchez, and José M. Cecilia. Efficient GPUbased parallelization of solvation calculation for the blind docking problem. The Journal of Supercomputing, 76(3): 1980–1998, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jakobs:2020:PEC

[87] Thomas Jakobs, Billy Naumann, and Gudula Rünger. Performance and energy consumption of the SIMD Gram– Schmidt process for vector orthogonalization. The Journal of Supercomputing, 76(3):1999–2021, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Contreras:2020:ECA

[88] Antonio V. Contreras, Antonio Llanes, Francisco J. Herrera, Sergio Navarro, Jose J. López-Espín, and José M. Cecilia. Enhancing the context-aware FOREX market simulation using a parallel elastic network model. The Journal of Supercomputing, 76(3):2022– 2038, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Armejach:2020:UAS

[89] Adrià Armejach, Helena Caminal, Juan M. Cebrian, Rubén Langarita, Rekai González-Alberquilla, Chris Adeniyi-Jones, Mateo Valero, Marc Casas, and Miquel Moretó. Using Arm’s scalable vector extension on stencil codes. The Journal of Supercomputing, 76(3):2039–2062, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:VNS

[90] Yishui Li, Peizhen Xie, Xinhai Chen, Jie Liu, Bo Yang, Shengguo Li, Chunye Gong, Xinbiao Gan, and Han Xu. VBSF: a new storage format for SIMD sparse matrix–vector multiplication on modern processors. The Journal of Supercomputing, 76(3):2063–2081, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cebrian:2020:SAA

[91] Juan M. Cebrian, Lasse Natvig, and Magnus Jahre. Scalability analysis of AVX-512 extensions. The Journal of Supercomputing, 76(3):2082– 2097, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pang:2020:INN

[92] Xiongwen Pang, Yanqiang Zhou, Pan Wang, Weiwei Lin, and Victor Chang. An innovative neural network approach for stock market prediction. The Journal of Supercomputing, 76(3):2098– 2118, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tuncer:2020:ERN

[93] Turker Tuncer, Fatih Ertam, Sengul Dogan, Emrah Aydemir, and Paweł Pławiak. Ensemble residual networkbased gender and activity recognition method with signals. The Journal of Supercomputing, 76(3):2119–2138, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:LSS

[94] Jianxin Li, Minjie Liu, Dongliang Ma, Jinyu Huang, Min Ke, and Tao Zhang. Learning shared subspace regularization with linear discriminant analysis for multi-label action recognition. The Journal of Supercomputing, 76(3): 2139–2157, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Amini:2020:TLD

[95] S. M. Amini and A. Karimi. Twolevel distributed clustering routing algorithm based on unequal clusters for large-scale Internet of Things networks. The Journal of Supercomputing, 76(3): 2158–2190, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sadeghi:2020:NEF

[96] Mostafa Sadeghi, Keivan Navi, and Mehdi Dolatshahi. Novel efficient full adder and full subtractor designs in quantum cellular automata. The Journal of Supercomputing, 76(3):

2191–2205, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Noorallahzadeh:2020:PPR

[97] Mojtaba Noorallahzadeh and Mohammad Mosleh. Parity-preserving reversible flip-flops with low quantum cost in nanoscale. The Journal of Supercomputing, 76(3):2206–2238, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hung:2020:EP

[98] Jason C. Hung, Neil Y. Yen, and Francisco Isidro Massetto. Editorial preface. The Journal of Supercomputing, 76(3):1359–1364, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-020-03228-8.pdf.

Park:2020:EDG

[99] Bum-Soo Park and Seung-Hee Kim. An effective digitized GPS signal transmission for high temporal precision IoT services. The Journal of Supercomputing, 76(3):1365–1382, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2020:EEI

[100] Kuei Min Wang and Lin Hui. Effectiveness evaluation of Internet of Things-aided firefighting by simulation. The Journal of Supercomputing, 76(3): 1383–1397, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2020:DMP

[101] Shih-Jung Wu, Rui-Dong Chiang, and Terng-Fang Wu. Direct mail promotion mechanisms and their application in supermarkets. The Journal of Supercomputing, 76(3):1398–1415, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2020:ICB

[102] Chao-Tung Yang, Yu-Wei Chan, Jung-Chun Liu, and Ben-Shen Lou. An implementation of cloud-based platform with R packages for spatiotemporal analysis of air pollution. The Journal of Supercomputing, 76(3):1416– 1437, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hsia:2020:CRM

[103] Chih-Hsien Hsia, Jia-Hao Yang, and Jen-Shiun Chiang. Complexity reduction method for ultrasound imaging enhancement in tetrolet transform domain. The Journal of Supercomputing, 76(3):1438–1449, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:WRN

[104] Jianqiang Li, Shenhe Zhao, Jijiang Yang, Zhisheng Huang, Bo Liu, Shi Chen, Hui Pan, and Qing Wang. WCP-RNN: a novel RNN-based approach for Bio–NER in Chinese EMRs. The Journal of Supercomputing, 76(3): 1450–1467, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wei:2020:PPC

[105] Yu-Chih Wei, Wei-Chen Wu, Gu-Hsin Lai, and Ya-Chi Chu. pISRA: privacy considered information security risk assessment model. The Journal of Supercomputing, 76(3):1468–1481, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shen:2020:GBI

[106] Jun-Hong Shen, Ching-Ta Lu, Mu-Yen Chen, and Neil Y. Yen. Grid-based indexing with expansion of resident domains for monitoring moving objects. The Journal of Supercomputing, 76(3): 1482–1501, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2020:EDA

[107] You-Shyang Chen and Jackson White. Empirical decision analytics approach of advanced granularity-based models for identifying performance measures of ERPS application. The Journal of Supercomputing, 76(3):1502–1535, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kansal:2020:GAB

[108] Sahil Kansal, Harish Kumar, Sakshi Kaushal, and Arun Kumar Sangaiah. Genetic algorithm-based cost minimization pricing model for on-demand IaaS cloud service. The Journal of Supercomputing, 76(3):1536–1561, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Liu:2020:GII

[109] Chien-Hung Liu, Woei-Kae Chen, and Chi-Chia Sun. GUIDE: an interactive and incremental approach for crawling Web applications. The Journal of Supercomputing, 76(3):1562–1584, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2020:EED

[110] Xuhui Yang, Qingguo Zhou, Jinqiang Wang, Rui Zhou, and Kuan-Ching Li. An energy-efficient dynamic decision model for wireless multi-sensor network. The Journal of Supercomputing, 76(3):1585–1603, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kao:2020:SDD

[111] Jui-Hung Kao, Po-Huan Hsiao, Yen-Jen Oyang, Wei-Zen Sun, Chih-Hong Sun, Horng-Twu Liaw, and Shin-Wen Chang. The study for dispatch decision of medical emergency resources with real-time spatial analysis. The Journal of Supercomputing, 76(3):1604– 1627, March 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chang:2020:EEI

[112] Yuan-Tsung Chang, Timothy K. Shih, Yung-Hui Li, and W. G. C. W. Kumara. Effectiveness evaluation of iris segmentation by using geodesic active contour (GAC). The Journal of Supercomputing, 76(3):1628–1641, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2020:MDF

[113] Shuxin Wang. Multisensor data fusion of motion monitoring system based on BP neural network. The Journal of Supercomputing, 76(3):1642–1656, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fan:2020:TLS

[114] Tongrang Fan, Wanting Qin, Wenbin Zhao, Feng Wu, and Jianmin Wang. A two-layer SIR information propagation model with heterogeneity based on coupled network. The Journal of Supercomputing, 76(3):1657–1679, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cheng:2020:TSA

[115] Ching-Wei Cheng, Shan-Yu Jung, Chia-Chun Lai, Sheng-Yu Tsai, and Chien-Chung Jeng. Transmission spectral analysis models for the assessment of white-shell eggs and brown-shell eggs freshness. The Journal of Supercomputing, 76(3):1680–1694, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shahid:2020:IRK

[116] Abdul Shahid, Muhammad Tanvir Afzal, Moloud Abdar, Mohammad Ehsan Basiri, Xujuan Zhou, Neil Y. Yen, and Jia-Wei Chang. Insights into relevant knowledge extraction techniques: a comprehensive review. The Journal of Supercomputing, 76(3):1695–1733, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:ENb

[117] Anonymous. Editor’s note. The Journal of Supercomputing, 76(3):1734, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.

com/content/pdf/10.1007/s11227-
020-03239-5.pdf.

Chang:2020:MIF

[118] Ye-In Chang, Jun-Hong Shen, Chia-En Li, Zih-Siang Chen, and Ming-Hsuan Tu. Mining image frequent patterns based on a frequent pattern list in image databases. The Journal of Supercomputing, 76(4):2597–2621, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2020:VAM

[119] Bo Wang, Hengrui Ma, Xunting Wang, Guiping Deng, Yan Yang, and Shaohua Wan. Vulnerability assessment method for cyber-physical system considering node heterogeneity. The Journal of Supercomputing, 76(4):2622–2642, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Walker-Roberts:2020:THU

[120] Steven Walker-Roberts, Mohammad Hammoudeh, Omar Aldabbas, Mehmet Aydin, and Ali Dehghantanha. Threats on the horizon: understanding security threats in the era of cyber-physical systems. The Journal of Supercomputing, 76(4):2643–2664, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-03028-9.pdf.

Ahmad:2020:SSV

[121] Usman Ahmad, Hong Song, Awais Bilal, Mamoun Alazab, and Alireza Jolfaei. Securing smart vehicles from relay attacks using machine learning. The Journal of Supercomputing, 76(4):

2665–2682, April 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2020:GPV

[122] Chin-Feng Lee, Jau-Ji Shen, Somya Agrawal, Yu-Ju Tseng, and Yu-Chi Kao. A generalized pixel value ordering data hiding with adaptive embedding capability. The Journal of Supercomputing, 76(4):2683–2714, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hao:2020:TMP

[123] Sheng Hao and Hu yin Zhang. Theoretical modeling for performance analysis of IEEE 1901 power-line communication networks in the multi-hop environment. The Journal of Supercomputing, 76(4):2715–2747, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Najjar-Ghabel:2020:HHP

[124] Samad Najjar-Ghabel, Leili Farzinvash, and Seyed Naser Razavi. HPDMS: high-performance data harvesting in wireless sensor networks with mobile sinks. The Journal of Supercomputing, 76(4):2748–2776, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Baranwal:2020:FIS

[125] Gaurav Baranwal, Manisha Singh, and Deo Prakash Vidyarthi. A framework for IoT service selection. The Journal of Supercomputing, 76(4):2777–2814, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

He:2020:SMO

[126] Feng He, Xiaoshe Dong, Nianjun Zou, Weiguo Wu, and Xingjun Zhang. Structured mesh-oriented framework design and optimization for a coarsegrained parallel CFD solver based on hybrid MPI/OpenMP programming. The Journal of Supercomputing, 76(4): 2815–2841, April 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Deng:2020:MRM

[127] Feifei Deng, Guangjun Xie, Renjun Zhu, and Yongqiang Zhang. A matrix representation method for decoders using majority gate characteristics in quantum-dot cellular automata. The Journal of Supercomputing, 76(4): 2842–2859, April 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rodriguez-Sanchez:2020:IEI

[128] Rafael Rodríguez-Sánchez, Francisco D. Igual, and Enrique S. Quintana-Ortí. Integration and exploitation of intraroutine malleability in BLIS. The Journal of Supercomputing, 76(4):2860– 2875, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abadi:2020:CSC

[129] Reza Mohamadi Bahram Abadi, Amir Masoud Rahmani, and Sasan Hossein Alizadeh. Challenges of server consolidation in virtualized data centers and open research issues: a systematic literature review. The Journal of Supercomputing, 76(4):2876–2927, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [130].

Abadi:2020:CCS

[130] Reza Mohamadi Bahram Abadi, Amir Masoud Rahmani, and Sasan Hossein Alizadeh. Correction to: Challenges of server consolidation in virtualized data centers and open research issues: a systematic literature review. The Journal of Supercomputing, 76(4):2928, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-019-03105-z.pdf.

See [129].

Jiang:2020:EBD

[131] Yizhang Jiang, Anqi Bi, Kaijian Xia, Jing Xue, and Pengjiang Qian. Exemplar-based data stream clustering toward Internet of Things. The Journal of Supercomputing, 76(4):2929–2957, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2020:QEU

[132] Qi Yu, Bruce Childers, Libo Huang, Cheng Qian, and Zhiying Wang. A quantitative evaluation of unified memory in GPUs. The Journal of Supercomputing, 76(4):2958–2985, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2020:OTS

[133] Kaiyi Zhao, Li Li, Saihua Cai, and Ruizhi Sun. An optimized time series combined forecasting method based on neural networks. The Journal of Supercomputing, 76(4):2986–3012, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Varri:2020:SRS

[134] Umasankararao Varri, Syamkumar Pasupuleti, and K. V. Kadambari. A scoping review of searchable encryption schemes in cloud computing: taxonomy, methods, and recent developments. The Journal of Supercomputing, 76(4):3013–3042, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Do:2020:NWS

[135] Cong Thuan Do, Hong Jun Choi, Sung Woo Chung, and Cheol Hong Kim. A novel warp scheduling scheme considering long-latency operations for high-performance GPUs. The Journal of Supercomputing, 76(4):3043–3062, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Haghighi:2020:IRC

[136] Mohammad Sayad Haghighi, Faezeh Farivar, Alireza Jolfaei, and Mohammad Hesam Tadayon. Intelligent robust control for cyber-physical systems of rotary gantry type under denial of service attack. The Journal of Supercomputing, 76(4):3063–3085, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nikooghadam:2020:PFS

[137] Mahdi Nikooghadam and Haleh Amintoosi. Perfect forward secrecy via an ECC-based authentication scheme for SIP in VoIP. The Journal of Supercomputing, 76(4):3086–3104, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abbasi:2020:MOB

[138] Mahdi Abbasi, Saeideh Vesaghati Fazel, and Milad Rafiee. MBitCuts: optimal bit-level cutting in geometric space packet classification. The Journal of Supercomputing, 76(4):3105–3128, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fang:2020:MSS

[139] Juan Fang, Mengxuan Wang, and Zelin Wei. A memory scheduling strategy for eliminating memory access interference in heterogeneous system. The Journal of Supercomputing, 76 (4):3129–3154, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-019-03135-7.pdf.

Yu:2020:CEN

[140] Misun Yu, Yu-Seung Ma, and Doo-Hwan Bae. Correction to: Efficient noise injection for exposing hidden data races. The Journal of Supercomputing, 76(4):3155, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-019-03124-w.pdf.

See [28].

Anonymous:2020:ENc

[141] Anonymous. Editor’s note. The Journal of Supercomputing, 76(4):2239, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

com/content/pdf/10.1007/s11227-
020-03237-7.pdf.

Tian:2020:RDA

[142] Ran Tian, Shanwei Li, and Guoying Yang. Research on a distributed autonegotiation model based on Stackelberg game theory. The Journal of Supercomputing, 76(4):2240–2251, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Govindasamy:2020:DSR

[143] P. Govindasamy and R. Dillibabu. Development of software reliability models using a hybrid approach and validation of the proposed models using big data. The Journal of Supercomputing, 76(4):2252–2265, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Skylakha:2020:ESA

[144] S. Skylakha, P. Sakthivel, and K. S. Arunselvan. Empirical study on application of machine learning techniques for resource allocation in health care using KPI. The Journal of Supercomputing, 76(4):2266–2274, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mahammad:2020:PAD

[145] Farooq Sunar Mahammad and V. Madhu Viswanatham. Performance analysis of data compression algorithms for heterogeneous architecture through parallel approach. The Journal of Supercomputing, 76(4):2275–2288, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Karthikeyan:2020:CSI

[146] P. R. Karthikeyan, P. Sakthivel, and T. S. Karthik. Comparative study of illumination-invariant foreground detection. The Journal of Supercomputing, 76(4):2289–2301, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Amarnath:2020:ITA

[147] D. Amarnath and S. Sujatha. Internetof-Things-aided energy management in smart grid environment. The Journal of Supercomputing, 76(4):2302–2314, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2020:RPS

[148] Feng Wang, Yanjun Wu, and Feiyue Huang. Rio: a personal storage system in multi-device and cloud. The Journal of Supercomputing, 76(4):2315–2338, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [392].

Gafar:2020:MNV

[149] Mona Gamal Gafar, Mohamed Elhoseny, and M. Gunasekaran. Modeling neutrosophic variables based on particle swarm optimization and information theory measures for forest fires. The Journal of Supercomputing, 76(4): 2339–2356, April 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:RFE

[150] Nan Li, Hai bo Gao, Liang Ding, Feng tian Lv, Zhong yan Bi, and Yi da Wang. Research on feature extraction and segmentation of rover wheel imprint. The Journal of Supercomputing, 76(4):2357–2373, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2502-7.pdf.

Yang:2020:NAS

[151] Jinfei Yang, Jiajia Li, and Shouqiang Liu. A new algorithm of stock data mining in Internet of multimedia things. The Journal of Supercomputing, 76(4):2374–2389, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jing:2020:REL

[152] Shibo Jing and Liming Yang. A robust extreme learning machine framework for uncertain data classification. The Journal of Supercomputing, 76(4): 2390–2416, April 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Guo:2020:REP

[153] Bin Guo and Jiamin Li. Research on the evolution of participants collaboration mechanism in PPP model based on computer simulation: based on the old community renovation project. The Journal of Supercomputing, 76(4): 2417–2434, April 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ma:2020:EWI

[154] Yan Ma, Qinxue Xiong, Jianqiang Zhu, and Shunyao Jiang. Early warning indexes determination of the crop injuries caused by waterlogging based on DHSVM model. The Journal of Supercomputing, 76(4):2435–2448, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fernandez-Cerero:2020:MDC

[155] Damián Fernández-Cerero, Ángel Jesús Varela-Vaca, Alejandro Fernández-Montes, María Teresa Gómez-López, and José Antonio Alvárez-Bermejo. Measuring data-centre workflows complexity through process mining: the Google cluster case. The Journal of Supercomputing, 76(4):2449–2478, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kaur:2020:BCS

[156] Maninder Kaur, Gurpreet Kaur, Pradip Kumar Sharma, Alireza Jolfaei, and Dhananjay Singh. Binary cuckoo search metaheuristic-based supercomputing framework for human behavior analysis in smart home. The Journal of Supercomputing, 76(4):2479–2502, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lin:2020:CTA

[157] Chuen Horng Lin, Chia Ching Yu, Ting You Wang, and Tsung Yi Chen. Classification of the tree for aerial image using a deep convolution neural network and visual feature clustering. The Journal of Supercomputing, 76(4): 2503–2517, April 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wan:2020:ECO

[158] Shaohua Wan, Xiang Li, Yuan Xue, Wenmin Lin, and Xiaolong Xu. Efficient computation offloading for Internet of vehicles in edge computingassisted 5G networks. The Journal of Supercomputing, 76(4):2518–2547, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dar:2020:CAE

[159] Zaineb Dar, Adnan Ahmad, Farrukh Aslam Khan, Furkh Zeshan, Razi Iqbal, Hafiz Husnain Raza Sherazi, and Ali Kashif Bashir. A contextaware encryption protocol suite for edge computing-based IoT devices. The Journal of Supercomputing, 76(4): 2548–2567, April 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumar:2020:ABC

[160] Akshi Kumar and Anshika Arora. An ANFIS-based compatibility scorecard for IoT integration in websites. The Journal of Supercomputing, 76(4): 2568–2596, April 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Qadir:2020:NAM

[161] Muhammad Zuhair Qadir, Songhao Piao, Haiyang Jiang, and Mohammed El Habib Souidi. A novel approach for multi-agent cooperative pursuit to capture grouped evaders. The Journal of Supercomputing, 76(5):3416–3426, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Duan:2020:ROA

[162] Ke Duan. Research on optimization and application of evaluation algorithm for intelligent city. The Journal of Supercomputing, 76(5):3427–3439, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alzahrani:2020:ITP

[163] Sultan Saeed Alzahrani. Investigation of teachers’ perspective about early intervention services. The Journal of Supercomputing, 76(5):3440–3461, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2020:ESP

[164] Zuyi Yu, Weiwei Nie, Weidong Zhou, Fangzhou Xu, Shasha Yuan, Yan Leng, and Qi Yuan. Epileptic seizure prediction based on local mean decomposition and deep convolutional neural network. The Journal of Supercomputing, 76(5):3462–3476, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wei:2020:SEK

[165] Qifeng Wei, Xin Gu, and Yuan Yuan. On the structural evolution of the knowledge network and behaviors of the knowledge subjects. The Journal of Supercomputing, 76(5):3477–3493, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zheng:2020:HFS

[166] Yuefeng Zheng, Ying Li, Gang Wang, Yupeng Chen, Qian Xu, Jiahao Fan, and Xueting Cui. A hybrid feature selection algorithm for microarray data. The Journal of Supercomputing, 76(5): 3494–3526, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2020:RSE

[167] Ruiying Wang, Guoping He, Xiaoming Wu, Fuqiang Wang, and Yifan Hu. Research on suboptimal energy balance of non-uniform distributed nodes in WSN. The Journal of Supercomputing, 76(5):3527–3541, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Liu:2020:GON

[168] Wei Liu, Quan Wang, Yunlong Zhu, and Hanning Chen. GRU: optimization of NPI performance. The Journal of Supercomputing, 76(5):3542–3554, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahmad:2020:DBD

[169] Iftikhar Ahmad, Gulzar Ahmed, Syed Adeel Ali Shah, and Ejaz Ahmed. A decade of big data literature: analysis of trends in light of bibliometrics. The Journal of Supercomputing, 76(5): 3555–3571, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhu:2020:FRM

[170] Yongqing Zhu, Juniarto Samsudin, Renuga Kanagavelu, Weiwen Zhang, Long Wang, Theint Theint Aye, and Rick Siow Mong Goh. Fast recovery MapReduce (FAR-MR) to accelerate failure recovery in big data applications. The Journal of Supercomputing, 76(5):3572–3588, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:ENd

[171] Anonymous. Editor’s note. The Journal of Supercomputing, 76(5):3589, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
020-03277-z.pdf.

Chen:2020:EHP

[172] Shouqiang Chen, Yang Chen, Feng Yuan, and Xiaowei Chang. Establishment of herbal prescription vector space model based on word cooccurrence. The Journal of Supercomputing, 76(5):3590–3601, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lv:2020:FFS

[173] Chundong Lv, Jia Wang, and Fanfei Zhang. Forest fire spread model based on the grey system theory. The Journal of Supercomputing, 76(5): 3602–3614, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gu:2020:AIL

[174] Qun Gu and Xiaohong Hao. Adaptive iterative learning control based on particle swarm optimization. The Journal of Supercomputing, 76(5):3615–3622, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shi:2020:RCP

[175] Shurong Shi, Yi Li, and Chuang Wan. Robust continuous piecewise linear regression model with multiple change points. The Journal of Supercomputing, 76(5):3623–3645, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Xu:2020:ARS

[176] Haoli Xu, Daqing Wang, Zhengdong Deng, Zhibin Ding, Zhixin Liu, Guangyuan Wang, and Borui Ni. Application of remote sensing fuzzy assessment method in groundwater potential in Wailingding Island. The Journal of Supercomputing, 76(5): 3646–3658, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pan:2020:DGD

[177] Xinyu Pan, Jingzhong Ma, and Chengxia Wu. Decision game of data sharing in supply chain enterprises considering data value over time. The Journal of Supercomputing, 76(5): 3659–3672, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Weijun:2020:SSS

[178] He Weijun, Wang Hui, Lin Ling, Thomas Stephen Ramsey, and Huang Zhengwei. Study of e-smile service influence on customers’ satisfaction in social business context. The Journal of Supercomputing, 76(5):3673–3688, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2599-8.pdf.

Shijun:2020:RMC

[179] Song Shijun. Risk management and countering measurements by computer modeling and simulation technology in the approval and early preparation stages of a large international project. The Journal of Supercomputing, 76(5):

3689–3701, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2020:HPS

[180] Meihong Wu. Heuristic parallel selective ensemble algorithm based on clustering and improved simulated annealing. The Journal of Supercomputing, 76 (5):3702–3712, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Liang:2020:NGO

[181] Ruishi Liang, Mingzhi Mao, Hui Ma, and Huan Wang. A new goal ordering for incremental planning. The Journal of Supercomputing, 76(5):3713–3728, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:UMR

[182] Zhao Li, Jun Wu, Xiaofeng Zhang, Jingsha He, Peng Chen, and Keqing He. Using metadata for recommending business process. The Journal of Supercomputing, 76(5):3729–3748, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ye:2020:AVI

[183] Jianjun Ye, Xingyu Huang, Yanglin Cheng, Junda Shao, and Yuan Zhang. Air volume improvement in the duct system in frost-free refrigerators based on the CFD method. The Journal of Supercomputing, 76(5):3749–3764, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Arunsundar:2020:AEC

[184] B. Arunsundar, P. Sakthivel, and E. Natarajan. Analysis of energy consumption and latency in advanced wireless networks through DRX mechanism. The Journal of Supercomputing, 76(5):3765–3787, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nikose:2020:CDA

[185] T. J. Nikose and R. S. Sonparote. Computing dynamic across-wind response of tall buildings using artificial neural network. The Journal of Supercomputing, 76(5):3788–3813, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Panwala:2020:ASS

[186] Fenil. C. Panwala and R. Kumar. Asymmetric sifter-shaped microchannel network in biological MEMS for size- and mass-based mammalian cell sorting and separation using hydrodynamic technique. The Journal of Supercomputing, 76(5):3814–3846, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Peng:2020:EAI

[187] Qiao Ming Peng, Lin Hui, Kuei Min Wang, and Liang Cheng Chang. Effectiveness analysis of an IoT mechanism in support of monitoring Chinese white dolphins by simulation model. The Journal of Supercomputing, 76(5): 3847–3865, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2020:CUI

[188] Minchul Lee, Sieun Jeon, and Min Song. Characterizing user interest in NoSQL databases of social question and answer data. The Journal of Supercomputing, 76(5):3866–3881, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2020:DSL

[189] Hyeonseo Lee, Nakyeong Lee, Harim Seo, and Min Song. Developing a supervised learning-based social media business sentiment index. The Journal of Supercomputing, 76(5):3882–3897, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        018-02737-x.pdf.

Cuzzocrea:2020:RMA

[190] Alfredo Cuzzocrea, Nickerson Ferreira, and Pedro Furtado. A rewrite/merge approach for supporting real-time data warehousing via lightweight data integration. The Journal of Supercomputing, 76(5):3898–3922, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:ENe

[191] Anonymous. Editor’s note. The Journal of Supercomputing, 76(5):3157, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        020-03275-1.pdf.

Liu:2020:HSP

[192] Yongxun Liu, Jingyun Zhao, Yonggang Yao, Qiuxia Cao, and Jichao Cui. Highspeed parallel robot dynamic modelling based on PLC. The Journal of Supercomputing, 76(5):3158–3172, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sun:2020:NDD

[193] Xudong Sun, Ke Zhu, Xiaogang Jiang, and Yande Liu. Non-destructive detection of blackheart and soluble solids content of intact pear by online NIR spectroscopy. The Journal of Supercomputing, 76(5):3173–3187, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pu:2020:ABS

[194] Zhenzi Pu, Qijie Jiang, Hongzhu Yue, and Maria Tsaptsinos. Agent-based supply chain allocation model and its application in smart manufacturing enterprises. The Journal of Supercomputing, 76(5):3188–3198, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See retraction note [2176].

Hu:2020:URF

[195] Xiu zhen Hu, Hai xia Long, Chang jiang Ding, Su juan Gao, and Rui Hou. Using random forest algorithm to predict super-secondary structure in proteins. The Journal of Supercomputing, 76(5):3199–3210, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2020:FER

[196] Hao Wang, Senbing Wei, and Baofu Fang. Facial expression recognition using iterative fusion of MO–HOG and deep features. The Journal of Supercomputing, 76(5):3211–3221, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Qi:2020:MMW

[197] Shaoqun Qi, Luhe Wan, and Baoling Fu. Multisource and multiuser water resources allocation based on genetic algorithm. The Journal of Supercomputing, 76(5):3222–3230, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shao:2020:MHC

[198] Quan Shao, Meng Jia, Chenchen Xu, and Yan Zhu. Multi-helicopter collaborative search and rescue operation research based on decision-making. The Journal of Supercomputing, 76(5): 3231–3251, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

He:2020:FDN

[199] Yuqing He, Pingjiang Lu, Lujiu Deng, Jianguo Yin, Yuanyuan He, Zhenghua Zhang, and Hongbin He. Fast distribution network reconfiguration algorithm based on minus feasibility analysis unit. The Journal of Supercomputing, 76(5): 3252–3265, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Liu:2020:VAK

[200] Xiaojun Liu, Mengmeng Wang, and Hanliang Fu. Visualized analysis of knowledge development in green building based on bibliographic data mining. The Journal of Supercomputing, 76 (5):3266–3282, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:RER

[201] Huimin Li, Qinpu Zhang, and Zuolong Zheng. Research on enterprise radical innovation based on machine learning in big data background. The Journal of Supercomputing, 76(5):3283–3297, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yue:2020:RDA

[202] Hongzhu Yue, Qijie Jiang, Chuanbin Yin, and Jonny Wilson. Research on data aggregation and transmission planning with Internet of Things technology in WSN multi-channel aware network. The Journal of Supercomputing, 76(5):3298–3307, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See retraction note [2307].

Shao:2020:CSD

[203] Quan Shao, Chenchen Xu, Yu Zhou, and Hongji Dong. Cast shadow detection based on the YCbCr color space and topological cuts. The Journal of Supercomputing, 76(5):3308–3326, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fang:2020:CFP

[204] Baofu Fang and Lu Fang. Concise feature pyramid region proposal network for multi-scale object detection. The Journal of Supercomputing, 76(5): 3327–3337, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jin:2020:DEI

[205] Xin Jin, Kaixuan Zhao, Jiangtao Ji, Zhaomei Qiu, Zhitao He, and Hao Ma.

Design and experiment of intelligent monitoring system for vegetable fertilizing and sowing. The Journal of Supercomputing, 76(5):3338–3354, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2020:ECS

[206] Jing Yang, Na Li, Ning An, Yu Chen, and Gil Alterovitz. An efficient causal structure learning algorithm for linear arbitrarily distributed continuous data. The Journal of Supercomputing, 76(5): 3355–3363, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2020:RWS

[207] Anji Yu and Shimin Yang. Research on web server cluster load balancing algorithm in web education system. The Journal of Supercomputing, 76(5): 3364–3373, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Karthikeyan:2020:ECA

[208] K. Karthikeyan, R. Sunder, K. Shankar, S. K. Lakshmanaprabu, V. Vijayakumar, Mohamed Elhoseny, and Gunasekaran Manogaran. Energy consumption analysis of virtual machine migration in cloud using hybrid swarm optimization (ABC-BA). The Journal of Supercomputing, 76(5):3374–3390, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See retraction note [1989].

Wang:2020:RIA

[209] Xiaojuan Wang, Lanshan Gan, and Songlin Liu. Research on intelligence analysis technology of financial industry data based on genetic algorithm. The Journal of Supercomputing, 76(5): 3391–3401, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Feng:2020:NEM

[210] Jiao Feng, Wen Liu, Peng Li, and Johnson Zhang. Network-enabled MIMO systems with distributed data processing. The Journal of Supercomputing, 76 (5):3402–3415, May 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Suresh:2020:FDA

[211] S. Suresh and N. Sankar Ram. Feasible DDoS attack source traceback scheme by deterministic multiple packet marking mechanism. The Journal of Supercomputing, 76(6):4232–4246, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Reshmi:2020:PRA

[212] B. Reshmi and P. Poongodi. Profit and resource availability-constrained optimal handling of high-performance scientific computing tasks. The Journal of Supercomputing, 76(6):4247–4261, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Suresh:2020:HCD

[213] A. Suresh, R. Kumar, and R. Varatharajan. Health care data analysis using evolutionary algorithm. The Journal of Supercomputing, 76(6):4262–4271, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sharmila:2020:AIS

[214] L. Sharmila and U. Sakthi. An artificial immune system-based algorithm for abnormal pattern in medical domain. The Journal of Supercomputing, 76(6):4272–4286, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vignesh:2020:VCA

[215] U. Vignesh and R. Parvathi. 3D visualization and cluster analysis of unstructured protein sequences using ARCSA with a file conversion approach. The Journal of Supercomputing, 76(6):4287–4301, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nisha:2020:TFB

[216] U. Nilabar Nisha and A. Mahabub Basha. Triangular fuzzy-based spectral clustering for energy-efficient routing in wireless sensor network. The Journal of Supercomputing, 76(6):4302–4327, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

AlZubi:2020:BDA

[217] Ahmad Ali AlZubi. Big data analytic diabetics using map reduce and classification techniques. The Journal of Supercomputing, 76(6):4328–4337, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See retraction note [1988].

Subramanium:2020:AET

[218] Priti Subramanium and Rajeshree D. Raut. AI-enabled turbo-coded OFDM system for improved BER performance.

The Journal of Supercomputing, 76(6): 4338–4348, June 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Priya:2020:AFH

[219] S. Kavi Priya, G. Shenbagalakshmi, and T. Revathi. Applied fuzzy heuristics for automation of hygienic drinking water supply system using wireless sensor networks. The Journal of Supercomputing, 76(6):4349–4375, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Aljawarneh:2020:GGD

[220] Shadi A. Aljawarneh and Radhakrishna Vangipuram. GARUDA: Gaussian dissimilarity measure for feature representation and anomaly detection in Internet of things. The Journal of Supercomputing, 76(6):4376–4413, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alarifi:2020:BDA

[221] Abdulaziz Alarifi, Amr Tolba, Zafer Al-Makhadmeh, and Wael Said. A big data approach to sentiment analysis using greedy feature selection with cat swarm optimization-based long short-term memory neural networks. The Journal of Supercomputing, 76(6): 4414–4429, June 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nandini:2020:AAD

[222] V. Nandini and P. Uma Maheswari. Automatic assessment of descriptive answers in online examination system using semantic relational features.

The Journal of Supercomputing, 76(6): 4430–4448, June 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Saravanan:2020:AMI

[223] S. Saravanan and P. Sudhakar. Analysis of mobile Internet speed, signal strength and FMDH antenna design for improved Internet speed. The Journal of Supercomputing, 76(6):4449–4475, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Senthilkumar:2020:PMI

[224] Rajarathinam Senthilkumar and Ramasamy Kannan Gnanamurthy. A proposed method for the improvement in biometric facial image recognition using document-based classification. The Journal of Supercomputing, 76(6):4476–4494, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Arun:2020:VAB

[225] A. Arun and M. Devaraju. VLSI architecture of binary encoding technique for fast motion estimation based on Hamming distances. The Journal of Supercomputing, 76(6):4495–4507, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Prabha:2020:HEM

[226] K. R. Prabha and M. Jagadeeswari. A hybrid encoding method for reducing code stream in EICA-optimized SMVQ reversible data hiding. The Journal of Supercomputing, 76(6):4508–4524, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Baalamurugan:2020:MOK

[227] K. M. Baalamurugan and S. Vijay Bhanu. A multi-objective krill herd algorithm for virtual machine placement in cloud computing. The Journal of Supercomputing, 76(6):4525–4542, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jayarajan:2020:EAB

[228] P. Jayarajan, G. R. Kanagachidambaresan, T. V. P. Sundararajan, K. Sakthipandi, R. Maheswar, and A. Karthikeyan. An energy-aware buffer management (EABM) routing protocol for WSN. The Journal of Supercomputing, 76(6):4543–4555, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yuvaraj:2020:HPL

[229] N. Yuvaraj and C. Suresh Ghana Dhas. High-performance link-based cluster ensemble approach for categorical data clustering. The Journal of Supercomputing, 76(6):4556–4579, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:ENf

[230] Anonymous. Editor’s note. The Journal of Supercomputing, 76(6):4580, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
020-03278-y.pdf.

Cecilia:2020:REA

[231] José M. Cecilia and José M. García. Reengineering the ant colony optimization for CMP architectures. The Journal of Supercomputing, 76(6):4581–4602, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Griebler:2020:SIS

[232] Dalvan Griebler, Adriano Vogel, Daniele De Sensi, Marco Danelutto, and Luiz G. Fernandes. Simplifying and implementing service level objectives for stream parallelism. The Journal of Supercomputing, 76(6):4603– 4628, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rico-Gallego:2020:TAC

[233] Juan A. Rico-Gallego, Sergio Moreno-Álvarez, Juan C. Díaz-Martín, and Alexey L. Lastovetsky. A tool to assess the communication cost of parallel kernels on heterogeneous platforms. The Journal of Supercomputing, 76(6): 4629–4644, June 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rodriguez:2020:PMS

[234] Andrés Rodríguez, Angeles Navarro, Rafael Asenjo, Francisco Corbera, Rubén Gran, Darío Suárez, and Jose Nunez-Yanez. Parallel multiprocessing and scheduling on the heterogeneous Xeon+FPGA platform. The Journal of Supercomputing, 76(6):4645–4665, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ejjaaouani:2020:IPM

[235] Ksander Ejjaaouani, Olivier Aumage, Julien Bigot, Michel Méhrenberger, Hitoshi Murai, Masahiro Nakao, and Mitsuhisa Sato. InKS: a programming model to decouple algorithm from optimization in HPC codes. The Journal of Supercomputing, 76(6):4666–4681, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rey:2020:SRC

[236] Antón Rey, Francisco D. Igual, and Manuel Prieto-Matías. STEEL-RT: combining single task-single executor model and expanded scheduling to ease heterogeneity exploitation. The Journal of Supercomputing, 76(6):4682– 4700, June 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ghavidel:2020:HSE

[237] Abolfazl Ghavidel, Yasser Sedaghat, and Mahmoud Naghibzadeh. Hybrid scheduling to enhance reliability of real-time tasks running on reconfigurable devices. The Journal of Supercomputing, 76(6):4701–4730, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sun:2020:VPI

[238] Huihui Sun, Sergei Gorlatch, and Rongcai Zhao. Vectorizing programs with IF-statements for processors with SIMD extensions. The Journal of Supercomputing, 76(6):4731–4746, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Haque:2020:AIM

[239] Mridul Haque and Dip Sankar Banerjee. Accelerating influence maximization using heterogeneous algorithms. The Journal of Supercomputing, 76(6): 4747–4769, June 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:ENg

[240] Anonymous. Editor’s note. The Journal of Supercomputing, 76(6):3923, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        020-03274-2.pdf.

Amin:2020:JTC

[241] Adnan Amin, Feras Al-Obeidat, Babar Shah, May Al Tae, Changez Khan, Hamood Ur Rehman Durrani, and Sajid Anwar. Just-in-time customer churn prediction in the telecommunication sector. The Journal of Supercomputing, 76(6):3924–3948, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Neeba:2020:SBC

[242] E. A. Neeba, S. Koteeswaran, and N. Malarvizhi. Swarm-based clustering algorithm for efficient web blog and data classification. The Journal of Supercomputing, 76(6):3949–3962, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumar:2020:EDC

[243] Priyan Malarvizhi Kumar and Usha Devi Gandhi. Enhanced DTLS with CoAPbased authentication scheme for the Internet of Things in healthcare application. The Journal of Supercomputing, 76(6):3963–3983, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alagarsamy:2020:IRB

[244] Ramachandran Alagarsamy and S. A. Sahaaya Arul Mary. Intelligent rulebased approach for effective information retrieval and dynamic storage in local repositories. The Journal of Supercomputing, 76(6):3984–3998, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Midhunchakkaravarthy:2020:FFA

[245] J. Midhunchakkaravarthy and S. SelvaBrunda. Feature fatigue analysis of product usability using hybrid ant colony optimization with artificial bee colony approach. The Journal of Supercomputing, 76(6):3999–4016, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pacha:2020:SAS

[246] Shobharani Pacha, Suresh Ramalingam Murugan, and R. Sethukarasi. Semantic annotation of summarized sensor data stream for effective query processing. The Journal of Supercomputing, 76(6):4017–4039, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pacha:2020:CSA

[247] Shobharani Pacha, Suresh Ramalingam Murugan, and R. Sethukarasi. Correction to: Semantic annotation of summarized sensor data stream for effective query processing. The Journal of Supercomputing, 76(6):4040, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        017-2212-6.pdf.

Sun:2020:DAS

[248] Jianguo Sun, Wenshan Wang, Liang Kou, Yun Lin, Liguo Zhang, Qingan Da, and Lei Chen. A data authentication scheme for UAV ad hoc network communication. The Journal of Supercomputing, 76(6):4041–4056, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bathla:2020:GBM

[249] Gourav Bathla, Himanshu Aggarwal, and Rinkle Rani. A graph-based model to improve social trust and influence for social recommendation. The Journal of Supercomputing, 76(6): 4057–4075, June 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hua:2020:OMP

[250] Weixin Hua, Dejun Mu, Zhigao Zheng, and Dawei Guo. Online multi-person tracking assist by high-performance detection. The Journal of Supercomputing, 76(6):4076–4094, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Di:2020:CRS

[251] Guohui Di, Fulin Su, and Xinbo Xu. Cross-range scaling of inverse synthetic aperture radar images with complex moving targets based on parameter estimation. The Journal of Supercomputing, 76(6):4095–4116, June 2020.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dong:2020:PCB

[252] Xiaoming Dong, Meng Qian, and Rong Jiang. Packet classification based on the decision tree with information entropy. The Journal of Supercomputing, 76(6):4117–4131, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sun:2020:GRO

[253] Ying Sun, Jiabing Hu, Gongfa Li, Guozhang Jiang, Hegen Xiong, Bo Tao, Zujia Zheng, and Du Jiang. Gear reducer optimal design based on computer multimedia simulation. The Journal of Supercomputing, 76(6): 4132–4148, June 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2020:RAF

[254] Wenxu Zhang, Lei Ma, Xiaonan Li, and W. Zhang. Research and application of fuzzy decision based on multiagent system. The Journal of Supercomputing, 76(6):4149–4168, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kishorekumar:2020:FSI

[255] R. Kishorekumar and P. Deepa. A framework for semantic image annotation using LEGION algorithm. The Journal of Supercomputing, 76(6): 4169–4183, June 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Joshua:2020:ANF

[256] K. Paul Joshua, J. Mohanalin, and S. T. Jaya Christa. Adaptive neurofuzzy inference system based underfrequency load shedding for Tamil nadu. The Journal of Supercomputing, 76(6):4184–4198, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Naveen:2020:GCS

[257] J. Naveen, P. J. A. Alphonse, and Sivaraj Chinnasamy. 3D grid clustering scheme for wireless sensor networks. The Journal of Supercomputing, 76(6): 4199–4211, June 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Han:2020:SAQ

[258] Jiawei Han, Yanheng Liu, Xin Sun, and Aiping Chen. A self-adjusting quantum key renewal management scheme in classical network symmetric cryptography. The Journal of Supercomputing, 76(6):4212–4230, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [259].

Han:2020:CSA

[259] Jiawei Han, Yanheng Liu, Xin Sun, and Aiping Chen. Correction to: A self-adjusting quantum key renewal management scheme in classical network symmetric cryptography. The Journal of Supercomputing, 76(6):4231, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2373-y.pdf. See [258].

Yin:2020:ICF

[260] Chunyong Yin, Lingfeng Shi, Ruxia Sun, and Jin Wang. Improved collaborative filtering recommendation algorithm based on differential privacy protection. The Journal of Supercomputing, 76(7):5161–5174, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tu:2020:OCS

[261] Li Tu, Shuai Liu, Yan Wang, Chi Zhang, and Ping Li. An optimized cluster storage method for realtime big data in Internet of Things. The Journal of Supercomputing, 76(7): 5175–5191, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bhattacherjee:2020:EEM

[262] Srimoyee Bhattacherjee, Rituparna Das, Sunirmal Khatua, and Sarbani Roy. Energy-efficient migration techniques for cloud environment: a step toward green computing. The Journal of Supercomputing, 76(7):5192–5220, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Huh:2020:RCM

[263] Jun-Ho Huh. Reefer container monitoring system using PLC-based communication technology for maritime edge computing. The Journal of Supercomputing, 76(7):5221–5243, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jung:2020:MCT

[264] Daeyong Jung, Daewon Lee, Myungil Kim, Hoyoon Kim, and Seung-Keun Park. Mesh convergence test system in integrated platform environment for finite element analysis. The Journal of Supercomputing, 76(7):5244–5258, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Son:2020:DII

[265] Yunsik Son, Junho Jeong, and Yang-Sun Lee. Design and implementation of an IoT–cloud converged virtual machine system. The Journal of Supercomputing, 76(7):5259–5275, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:EMP

[266] Songyuan Li, Hong Shen, Yingpeng Sang, and Hui Tian. An efficient method for privacy-preserving trajectory data publishing based on data partitioning. The Journal of Supercomputing, 76(7):5276–5300, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2020:NCS

[267] Yonggun Kim and Yoojae Won. A new cost-saving and efficient method for patch management using blockchain. The Journal of Supercomputing, 76(7): 5301–5319, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Salim:2020:DDS

[268] Mikail Mohammed Salim, Shailendra Rathore, and Jong Hyuk Park. Distributed denial of service attacks and its defenses in IoT: a survey. The Journal of Supercomputing, 76(7): 5320–5363, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2020:CTR

[269] Gun Ho Lee and Hee Seon Han. Clustering of tourist routes for individual tourists using sequential pattern mining. The Journal of Supercomputing, 76 (7):5364–5381, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2020:RHP

[270] Sul-Ho Kim, Seok-Woo Jang, Jin-Ho Park, and Gye-Young Kim. Robust hand pose estimation using visual sensor in IoT environment. The Journal of Supercomputing, 76(7):5382–5401, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Singh:2020:ICB

[271] Amanpreet Singh and Maninder Kaur. Intelligent content-based cybercrime detection in online social networks using cuckoo search metaheuristic approach. The Journal of Supercomputing, 76(7):5402–5424, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jung:2020:DNN

[272] YoungGiu Jung and Chang-Min Jeong. Deep neural network-based automatic unknown protocol classification system using histogram feature. The Journal of Supercomputing, 76(7):5425–5441, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pang:2020:EAM

[273] Beibei Pang, Fei Hao, Yixuan Yang, and Doo-Soon Park. An efficient approach for multi-user multi-cloud service composition in human-land sustainable computational systems. The Journal of Supercomputing, 76(7): 5442–5459, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Han:2020:DDL

[274] Seong-Soo Han, Yoon-Ki Kim, You-Boo Jeon, JinSoo Park, Doo-Soon Park, DuHyun Hwang, and Chang-Sung Jeong. Distributed deep learning platform for pedestrian detection on IT convergence environment. The Journal of Supercomputing, 76 (7):5460–5485, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-020-03195-0.pdf.

Ren:2020:OFO

[275] Fei Ren, Xiaoliang Chen, Fei Hao, Yajun Du, and Jianzhong Zheng. OFNE: a framework of opinion features regulated network embedding. The Journal of Supercomputing, 76(7):5486–5500, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:FFN

[276] Kun Li, Shigang Li, Shan Huang, Yifeng Chen, and Yunquan Zhang. FastNBL: fast neighbor lists establishment for molecular dynamics simulation based on bitwise operations. The Journal of Supercomputing, 76(7):

5501–5520, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [3].

Zuo:2020:PIN

[277] Xiaojun Zuo, Ze Chen, Limian Dong, Jie Chang, and Botao Hou. Power information network intrusion detection based on data mining algorithm. The Journal of Supercomputing, 76(7): 5521–5539, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Beigy:2020:ISA

[278] Hamid Beigy and Mohammad Reza Meybodi. An iterative stochastic algorithm based on distributed learning automata for finding the stochastic shortest path in stochastic graphs. The Journal of Supercomputing, 76(7): 5540–5562, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bai:2020:CPP

[279] Jianli Bai and Rong Hao. Comment on “Privacy-preserving public auditing for non-manager group shared data”. The Journal of Supercomputing, 76(7): 5563–5577, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See [2].

Khansari:2020:MWC

[280] Mina Emami Khansari and Saeed Sharifian. A modified water cycle evolutionary game theory algorithm to utilize QoS for IoT services in cloudassisted fog computing environments. The Journal of Supercomputing, 76(7): 5578–5608, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yue:2020:GOA

[281] Xiaofeng Yue and Hongbo Zhang. Grasshopper optimization algorithm with principal component analysis for global optimization. The Journal of Supercomputing, 76(7):5609–5635, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tinetti:2020:LCP

[282] Fernando G. Tinetti, Maximiliano J. Perez, Ariel Fraidenraich, and Adolfo E. Altenberg. Legacy code and parallel computing: updating and parallelizing a numerical model. The Journal of Supercomputing, 76(7):5636–5654, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chheang:2020:NEL

[283] Vuthea Chheang, Sangkwon Jeong, Gookhwan Lee, Jong-Sung Ha, and Kwan-Hee Yoo. Natural embedding of live actors and entities into 360^{∘} virtual reality scenes. The Journal of Supercomputing, 76(7):5655–5677, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:ENh

[284] Anonymous. Editor’s note. The Journal of Supercomputing, 76(7):4771, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        020-03333-8.pdf.

Srinivas:2020:PPB

[285] Jangirala Srinivas, Ashok Kumar Das, and Joel J. P. C. Rodrigues. 2PBDC: privacy-preserving bigdata collection in cloud environment. The Journal of Supercomputing, 76(7):4772–4801, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Adewole:2020:TSA

[286] Kayode Sakariyah Adewole, Tao Han, Wanqing Wu, Houbing Song, and Arun Kumar Sangaiah. Twitter spam account detection based on clustering and classification methods. The Journal of Supercomputing, 76(7):4802– 4837, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

He:2020:NPR

[287] Li He, Zhicheng Qian, and Fengjun Shang. A novel predicted replication strategy in cloud storage. The Journal of Supercomputing, 76(7):4838–4856, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fan:2020:ECS

[288] Kuan Fan, Mingxi Liu, Guofang Dong, and Wenbo Shi. Enhancing cloud storage security against a new replay attack with an efficient public auditing scheme. The Journal of Supercomputing, 76(7):4857–4883, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Guo:2020:ABE

[289] Rui Guo, Xiong Li, Dong Zheng, and Yinghui Zhang. An attribute-based encryption scheme with multiple authorities on hierarchical personal health record in cloud. The Journal of Supercomputing, 76(7):4884–4903, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Myung:2020:ESC

[290] Sein Myung and Jong-Hyouk Lee. Ethereum smart contract-based automated power trading algorithm in a microgrid environment. The Journal of Supercomputing, 76(7):4904–4914, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hashem:2020:MSA

[291] Ibrahim Abaker Targio Hashem, Nor Badrul Anuar, Mohsen Marjani, Ejaz Ahmed, Haruna Chiroma, Ahmad Firdaus, Muhamad Taufik Abdullah, Faiz Alotaibi, Waleed Kamaleldin Mahmoud Ali, Ibrar Yaqoob, and Abdullah Gani. MapReduce scheduling algorithms: a review. The Journal of Supercomputing, 76(7): 4915–4945, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2020:NNP

[292] Xingjun Zhang, Yi Cai, Yunfei Liu, Zhiwei Xu, and Xiaoshe Dong. NADE: nodes performance awareness and accurate distance evaluation for degraded read in heterogeneous distributed erasure code-based storage. The Journal of Supercomputing, 76(7):4946–4975, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:ENi

[293] Anonymous. Editor’s note. The Journal of Supercomputing, 76(7):4976, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
020-03332-9.pdf.

Bock:2020:PTB

[294] Alexander Asp Bock and Florian Biermann. Puncalc: task-based parallelism and speculative reevaluation in spreadsheets. The Journal of Supercomputing, 76(7):4977–4997, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-019-02823-8.pdf.

Emoto:2020:DGP

[295] Kento Emoto and Fumihisa Sadahira. A DSL for graph parallel programming with vertex subsets. The Journal of Supercomputing, 76(7):4998–5015, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Niculescu:2020:TPB

[296] Virginia Niculescu and Frédéric Loulergue. Transforming powerlist-based divide-and-conquer programs for an improved execution model. The Journal of Supercomputing, 76(7):5016– 5037, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ohberg:2020:HCG

[297] Tomas Öhberg, August Ernstsson, and Christoph Kessler. Hybrid CPU– GPU execution support in the skeleton programming framework SkePU. The Journal of Supercomputing, 76 (7):5038–5056, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-019-02824-7.pdf.

Lopez-Gomez:2020:DSV

[298] Javier López-Gómez, David del Rio Astorga, Manuel F. Dolz, Javier Fernández, and J. Daniel García. Detecting semantic violations of lockfree data structures through C++ contracts. The Journal of Supercomputing, 76(7):5057–5078, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Allombert:2020:PBM

[299] Victor Allombert and Frédéric Gava. Programming BSP and MULTI-BSP algorithms in ML. The Journal of Supercomputing, 76(7):5079–5097, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wrede:2020:GHP

[300] Fabian Wrede, Christoph Rieger, and Herbert Kuchen. Generation of highperformance code based on a domainspecific language for algorithmic skeletons. The Journal of Supercomputing, 76(7):5098–5116, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rasch:2020:DHL

[301] Ari Rasch, Julian Bigge, Martin Wrodarczyk, Richard Schulze, and Sergei Gorlatch. dOCAL: high-level distributed programming with OpenCL and CUDA. The Journal of Supercomputing, 76(7):5117–5138, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Garcia:2020:CAP

[302] J. Daniel Garcia, David del Rio, Marco Aldinucci, Fabio Tordini, Marco Danelutto, Gabriele Mencagli, and Massimo Torquati. Challenging the abstraction penalty in parallel patterns libraries. The Journal of Supercomputing, 76(7): 5139–5159, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:ENj

[303] Anonymous. Editor’s note. The Journal of Supercomputing, 76(7):5160, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
020-03276-0.pdf.

Raja:2020:CPB

[304] J. Raja and M. Ramakrishnan. Confidentiality-preserving based on attribute encryption using auditable access during encrypted records in cloud location. The Journal of Supercomputing, 76(8):6026–6039, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mythili:2020:ABL

[305] R. Mythili, Revathi Venkataraman, and T. Sai Raj. An attribute-based lightweight cloud data access control using hypergraph structure. The Journal of Supercomputing, 76(8):6040– 6064, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Preetha:2020:NDW

[306] C. A. S. Deiva Preetha and Subburaj Ramasamy. Nonparametric dynamically weighted combination model to determine when to stop testing. The Journal of Supercomputing, 76(8): 6065–6082, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mohan:2020:MBC

[307] N. Mohan. Multi-bearer coordinate grouping-based code division multiple access for data communication in wireless network. The Journal of Supercomputing, 76(8):6083–6093, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumaresan:2020:TVA

[308] S. Kumaresan and Vijayaragavan Shanmugam. Time-variant attributebased multitype encryption algorithm for improved cloud data security using user profile. The Journal of Supercomputing, 76(8):6094–6112, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cai:2020:CSL

[309] Wenwei Cai, Jiaxian Zhu, Weihua Bai, Weiwei Lin, Naqin Zhou, and Keqin Li. A cost saving and load balancing task scheduling model for computational biology in heterogeneous cloud datacenters. The Journal of Supercomputing, 76(8):6113–6139, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2020:ENk

[310] Anonymous. Editor’s note. The Journal of Supercomputing, 76(8):6140, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
020-03335-6.pdf.

Jeong:2020:ABQ

[311] Hwa-Young Jeong. ANP-based quantification method for the smart manufacturing system design decomposition. The Journal of Supercomputing, 76(8): 6141–6157, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2020:LDT

[312] Shih-Nung Chen, Jason Chi-Shun Hung, and Chun-Chia Wang. Large dynamic tiled display system implementation and application. The Journal of Supercomputing, 76(8):6158–6176, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shen:2020:SAI

[313] Jun-Hong Shen, Cheng-Jung Yu, Ching-Ta Lu, WenYen Lin, Neil Y. Yen, Tien-Chi Huang, and Hong-Ray Chu. Spatial air index with neighbor information for processing k-nearest neighbor searches in IoT mobile computing. The Journal of Supercomputing, 76(8):6177–6194, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chang:2020:MGB

[314] Min-Kuan Chang, Yung-Jen Mei, Yu-Wei Chan, Mei-Yu Wu, and Wun-Ren Chen. Matching game-based hierarchical spectrum sharing in cooperative cognitive radio networks. The Journal of Supercomputing, 76(8):6195– 6218, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Akhtar:2020:ELG

[315] Faheem Akhtar, Jianqiang Li, Muhammad Azeem, Shi Chen, Hui Pan, Qing Wang, and Ji-Jiang Yang. Effective large for gestational age prediction using machine learning techniques with monitoring biochemical indicators. The Journal of Supercomputing, 76(8):6219–6237, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2020:RTT

[316] Xianchuan Wang, Sulan Zhang, Shan Gao, Mian Zhang, Jie Zhang, Xianchao Wang, and Zheng Xu. Response time of a ternary optical computer that is based on queuing systems. The Journal of Supercomputing, 76(8): 6238–6257, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2020:ETS

[317] Rong-Chang Chen and Chen-Yi Lin. An efficient two-stage method for solving the order-picking problem. The Journal of Supercomputing, 76(8): 6258–6279, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yen:2020:DRA

[318] Chun-Tai Yen, Chen-Kun Tsung, and Wen-Fang Wu. Detecting removed attributes in the cyber system for smart manufacturing. The Journal of Supercomputing, 76(8):6280–6301, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kashikolaei:2020:ETS

[319] Seyedeh Monireh Ggasemnezhad Kashikolaei, Ali Asghar Rahmani Hosseinabadi, Behzad Saemi, Morteza Babazadeh Shareh, Arun Kumar Sangaiah, and Gui-Bin Bian. An enhancement of task scheduling in cloud computing based on imperialist competitive algorithm and firefly algorithm. The Journal of Supercomputing, 76(8):6302–6329, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yong:2020:RMC

[320] Binbin Yong, Liang Huang, Fucun Li, Jun Shen, Xin Wang, and Qingguo Zhou. A research of Monte Carlo optimized neural network for electricity load forecast. The Journal of Supercomputing, 76(8):6330–6343, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2020:CNL

[321] Chao-Tung Yang, Endah Kristiani, Yuan-Ting Wang, Geyong Min, Ching-Han Lai, and Wei-Je Jiang. On construction of a network log management system using ELK stack with Ceph. The Journal of Supercomputing, 76(8): 6344–6360, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kang:2020:IIB

[322] Tsan-Ching Kang, Chia-Hsien Wen, Shih-Wei Guo, Wei-Yueh Chang, and Chen-Lin Chang. The implementation of an IoT-based exercise improvement system. The Journal of Supercomputing, 76(8):6361–6375, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bok:2020:PCS

[323] Kyoungsoo Bok, Jieun Han, Jongtae Lim, and Jaesoo Yoo. Provenance compression scheme based on graph patterns for large RDF documents. The Journal of Supercomputing, 76(8): 6376–6398, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2020:ECL

[324] Qing Zhao, Dan Wang, Jianqiang Li, and Faheem Akhtar. Exploiting the concept level feature for enhanced name entity recognition in Chinese EMRs. The Journal of Supercomputing, 76(8):6399–6420, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2020:WIM

[325] Seongchul Park, Sanghyun Seo, Changhoon Jeong, and Juntae Kim. The weights initialization methodology of unsupervised neural networks to improve clustering stability. The Journal of Supercomputing, 76(8):6421–6437, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jeon:2020:LCQ

[326] Jun-Cheol Jeon. Low-complexity QCA universal shift register design using multiplexer and D flip-flop based on electronic correlations. The Journal of Supercomputing, 76(8):6438–6452, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ke:2020:OTS

[327] Chih-Kun Ke and Chiao-Min Chang. Optimizing target selection complexity of a recommendation system by skyline query and multi-criteria decision analysis. The Journal of Supercomputing, 76(8):6453–6474, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yen:2020:AIA

[328] Neil Y. Yen, Jia-Wei Chang, Jia-Yi Liao, and You-Ming Yong. Analysis of interpolation algorithms for the missing values in IoT time series: a case of air quality in Taiwan. The Journal of Supercomputing, 76(8): 6475–6500, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2020:ESP

[329] Hsiu-An Lee, Hsiao-Hsien Rau, Louis R. Chao, and Chien-Yeh Hsu. Establishing a survival probability prediction model for different lung cancer therapies. The Journal of Supercomputing, 76(8):6501–6514, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2020:OAM

[330] Zhong xun Wang, Yan fang Dong, and Zhan kai Bao. Optimal algorithm for multi-hop relay selection based on cognitive radio. The Journal of Supercomputing, 76(8):6515–6531, August 2020.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dhas:2020:NCA

[331] P. Edwin Dhas and B. Sankara Gomathi. A novel clustering algorithm by clubbing GHFCM and GWO for microarray gene data. The Journal of Supercomputing, 76(8):5679–5693, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Krishnamurthy:2020:PLI

[332] R. Krishnamurthy, T. Srinivas, and I. S. Amiri. Physical layer impairmentaware shared path protection in wavelength-routed optical networks. The Journal of Supercomputing, 76(8): 5694–5714, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Madhavi:2020:ELH

[333] Nachimuthu Madhavi and Muthusamy Madheswaran. Enhanced lifetime of heterogeneous wireless sensor network using stable election protocol with region-based energy-conscious sink movement. The Journal of Supercomputing, 76(8):5715–5731, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ponraj:2020:DEP

[334] Abraham Sudharson Ponraj and T. Vigneswaran. Daily evapotranspiration prediction using gradient boost regression model for irrigation planning. The Journal of Supercomputing, 76(8): 5732–5744, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Munirathinam:2020:NIF

[335] Deepak Raj Munirathinam and Mohanasundaram Ranganadhan. A new improved filter-based feature selection model for high-dimensional data. The Journal of Supercomputing, 76(8): 5745–5762, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Karthikeyan:2020:DIC

[336] S. Karthikeyan, K. Vimala Devi, and K. Valarmathi. Design and implementation of CfoTS networks for industrial fault detection and correction mechanism. The Journal of Supercomputing, 76(8):5763–5779, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sivakumar:2020:EHC

[337] R. Sivakumar and H. Mangalam. Ensemble hill climbing optimization in adaptive cruise control for safe automated vehicle transportation. The Journal of Supercomputing, 76(8): 5780–5800, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Deepa:2020:FSR

[338] P. Deepa and M. Suganthi. A fuzzy shape representation of a segmented vessel tree and kernel-induced random forest classifier for the efficient prediction of lung cancer. The Journal of Supercomputing, 76(8):5801– 5824, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Amulu:2020:CMH

[339] Lourdes Mary Amulu and Ravi Ramraj. Combinatorial meta-heuristics approaches for DVFS-enabled green clouds. The Journal of Supercomputing, 76(8):5825–5834, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumar:2020:ENC

[340] M. Rajeev Kumar and K. Arthi. An effective non-cooperative iris recognition system using hierarchical collaborative representation-based classification. The Journal of Supercomputing, 76(8): 5835–5848, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Satpathy:2020:NHD

[341] Sambit Satpathy, Prakash Mohan, Sanchali Das, and Swapan Debbarma. A new healthcare diagnosis system using an IoT-based fuzzy classifier with FPGA. The Journal of Supercomputing, 76(8):5849–5861, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Thirumoorthy:2020:TCE

[342] P. Thirumoorthy, P. Kalyanasundaram, R. Maheswar, P. Jayarajan, G. R. Kanagachidambaresan, and Iraj S. Amiri. Time-critical energy minimization protocol using PQM (TCEM-PQM) for wireless body sensor network. The Journal of Supercomputing, 76(8):5862–5872, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kalidindi:2020:DBB

[343] Kishore Raju Kalidindi, Pardha Saradhi Varma Gottumukkala, and Rajyalakshmi Davuluri. Derivative-based band clustering and multi-agent PSO optimization for optimal band selection of hyper-spectral images. The Journal of Supercomputing, 76(8):5873–5898, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mustafa:2020:DHS

[344] Mohammed Mustafa, G. M. T. Abdalla, S. Manimurugan, and Adel R. Alharbi. Detection of heartbeat sounds arrhythmia using automatic spectral methods and cardiac auscultatory. The Journal of Supercomputing, 76(8): 5899–5922, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mandhula:2020:PCO

[345] Trupthi Mandhula, Suresh Pabboju, and Narsimha Gugulotu. Predicting the customer’s opinion on Amazon products using selective memory architecture-based convolutional neural network. The Journal of Supercomputing, 76(8):5923–5947, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kolandaisamy:2020:ASR

[346] Raenu Kolandaisamy, Rafidah Md. Noor, Muhammad Reza Z’aba, Ismail Ahmedy, and Indraah Kolandaisamy. Adapted stream region for packet marking based on DDoS attack detection in vehicular ad hoc networks. The Journal of Supercomputing, 76(8):

5948–5970, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Saranya:2020:PPD

[347] K. Saranya and K. Premalatha. Privacy-preserving data publishing based on sanitized probability matrix using transactional graph for improving the security in medical environment. The Journal of Supercomputing, 76(8): 5971–5980, August 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Amuthadevi:2020:DFA

[348] C. Amuthadevi and Gayathri Monicka Subarnan. Development of fuzzy approach to predict the fetus safety and growth using AFI. The Journal of Supercomputing, 76(8):5981–5995, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Prakash:2020:RKD

[349] G. Prakash, Raja Krishnamoorthy, and P. T. Kalaivaani. Resource key distribution and allocation based on sensor vehicle nodes for energy harvesting in vehicular ad hoc networks for transport application. The Journal of Supercomputing, 76(8):5996–6009, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Malathy:2020:NEE

[350] S. Malathy, Ravi Rastogi, R. Maheswar, G. R. Kanagachidambaresan, T. V. P. Sundararajan, and D. Vigneswaran. A novel energy-efficient framework (NEEF) for the wireless body sensor network. The Journal of Supercomputing, 76(8):6010–6025, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Devi:2020:MLI

[351] R. Lakshmi Devi and V. Kalaivani. Machine learning and IoT-based cardiac arrhythmia diagnosis using statistical and dynamic features of ECG. The Journal of Supercomputing, 76(9): 6533–6544, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02873-y.

Ranjan:2020:NEC

[352] Joshua Arul Kumar Ranjan, Titus Sigamani, and Janet Barnabas. A novel and efficient classifier using spiking neural network. The Journal of Supercomputing, 76(9):6545– 6560, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02881-y.

Zhang:2020:FMS

[353] Lijuan Zhang. Feature mining simulation of video image information in multimedia learning environment based on BOW algorithm. The Journal of Supercomputing, 76(9):6561– 6578, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02890-x.

Zhang:2020:NCA

[354] Shuo Zhang, Yaping Liu, and Ning Zhou. A novel cost-aware algorithm for dynamic task placement problem in a heterogeneous Internetscale data center. The Journal of Supercomputing, 76(9):6579–6598, September 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02892-9.

Ramaswamy:2020:ECD

[355] Vanaja Ramaswamy and Saswati Mukherjee. An effective clinical decision support system using swarm intelligence. The Journal of Supercomputing, 76(9):6599– 6618, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02888-5.

Goswami:2020:GSI

[356] Sayan Goswami, Ayam Pokhrel, and Yang Zhou. GraphMap: scalable iterative graph processing using NoSQL. The Journal of Supercomputing, 76(9): 6619–6647, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03097-w.

Ray:2020:RTE

[357] Partha Pratim Ray, Dinesh Dash, and Debashis De. Real-time eventdriven sensor data analytics at the edge-Internet of Things for smart personal healthcare. The Journal of Supercomputing, 76(9):6648– 6668, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03089-w.

Barreda:2020:IFC

[358] María Barreda, José I. Aliaga, and Marc Casas. Iteration-fusing conjugate gradient for sparse linear systems with MPI + OmpSs. The Journal of Supercomputing, 76(9):6669– 6689, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03100-4.

Yin:2020:ENI

[359] Chuanlong Yin, Yuefei Zhu, and Hetong Zhang. Enhancing network intrusion detection classifiers using supervised adversarial training. The Journal of Supercomputing, 76(9):6690– 6719, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03092-1.

Mazumdar:2020:RMS

[360] Somnath Mazumdar and Alberto Scionti. Ring-mesh: a scalable and high-performance approach for manycore accelerators. The Journal of Supercomputing, 76(9):6720– 6752, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03072-5.

Arabnejad:2020:SSC

[361] Hamid Arabnejad, João Bispo, and Jorge G. Barbosa. Source-to-source compilation targeting OpenMP-based automatic parallelization of C applications. The Journal of Supercomputing, 76(9):6753–6785, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-019-03109-
                                        9.

Hatami:2020:ESM

[362] Elmira Hatami and Bahman Arasteh. An efficient and stable method to cluster software modules using ant colony optimization algorithm. The Journal of Supercomputing, 76(9):6786– 6808, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03112-0.

Zhou:2020:LTA

[363] Shuai Zhou and Xianfu Meng. A location and time-aware resource searching scheme in mobile P2P ad hoc networks. The Journal of Supercomputing, 76(9): 6809–6833, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03139-3.

Morinigo:2020:PDE

[364] José A. Moríñigo, Pablo García-Muller, and Rafael Mayo-García. Performance drop at executing communicationintensive parallel algorithms. The Journal of Supercomputing, 76(9):6834– 6859, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03142-8.

Fotohi:2020:NCT

[365] Reza Fotohi and Somayyeh Firoozi Bari. A novel countermeasure technique to protect WSN against denialof-sleep attacks using firefly and Hopfield neural network (HNN) algorithms. The Journal of Supercomputing, 76(9): 6860–6886, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03131-x.

Li:2020:UIG

[366] Ling Li, Sida Dai, and Kunmeng Yang. Using improved gradientboosted decision tree algorithm based on Kalman filter (GBDT-KF) in time series prediction. The Journal of Supercomputing, 76(9):6887– 6900, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03130-y.

Devaraj:2020:SDC

[367] Rajesh Devaraj. A solution to drawbacks in capturing execution requirements on heterogeneous platforms. The Journal of Supercomputing, 76(9): 6901–6916, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03145-w.

Seyedi:2020:NNI

[368] Behnam Seyedi and Reza Fotohi. NI-ASHPT: a novel intelligent agent-based strategy using hello packet table (HPT) function for trust Internet of Things. The Journal of Supercomputing, 76(9): 6917–6940, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03143-7.

Li:2020:ESO

[369] Chunlin Li, Chengyi Wang, and Youlong Luo. An efficient scheduling optimization strategy for improving consistency maintenance in edge cloud environment. The Journal of Supercomputing, 76(9):6941– 6968, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03133-9.

Javadpour:2020:DSM

[370] Amir Javadpour, Guojun Wang, and Kuan-Ching Li. Detecting straggler MapReduce tasks in big data processing infrastructure by neural network. The Journal of Supercomputing, 76(9): 6969–6993, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03136-6.

Li:2020:ERS

[371] Chunlin Li, Jingpan Bai, and Youlong Luo. Efficient resource scaling based on load fluctuation in edge-cloud computing environment. The Journal of Supercomputing, 76(9):6994– 7025, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03134-8.

Dokeroglu:2020:RPH

[372] Tansel Dokeroglu, Selen Pehlivan, and Bilgin Avenoglu. Robust parallel hybrid artificial bee colony algorithms for the multi-dimensional numerical optimization. The Journal of Supercomputing, 76(9):7026– 7046, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03127-7.

Gomez-Rodriguez:2020:CIE

[373] Mario A. Gomez-Rodriguez, Victor J. Sosa-Sosa, and Jose Luis Gonzalez. CloudBench: an integrated evaluation of VM placement algorithms in clouds. The Journal of Supercomputing, 76(9): 7047–7080, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03141-9.

Mabodi:2020:MLT

[374] Kobra Mabodi, Mehdi Yusefi, and Reza Fotohi. Multi-level trust-based intelligence schema for securing of Internet of Things (IoT) against security threats using cryptographic authentication. The Journal of Supercomputing, 76(9): 7081–7106, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03137-5.

Chuang:2020:TTC

[375] Yung-Ting Chuang and Feng-Wei Li. TCR: a trustworthy and churn-resilient academic distribution and retrieval system in P2P networks. The Journal of Supercomputing, 76(9):7107– 7139, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03146-9.

Ramezani:2020:PAS

[376] Reza Ramezani. A prefetch-aware scheduling for FPGA-based multitask graph systems. The Journal of Supercomputing, 76(9):7140– 7160, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03153-w.

Seo:2020:DAR

[377] Jung hyun Seo and HyeongOk Lee. Design and analysis of the rotational binary graph as an alternative to hypercube and torus. The Journal of Supercomputing, 76(9):7161– 7176, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03115-x.

Gandomi:2020:DMP

[378] Abolfazl Gandomi, Ali Movaghar, and Ahmad Khademzadeh. Designing a MapReduce performance model in distributed heterogeneous platforms based on benchmarking approach. The Journal of Supercomputing, 76(9):7177– 7203, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03162-9.

Kang:2020:EGC

[379] Minseo Kang and Jae-Gil Lee. Effect of garbage collection in iterative algorithms on Spark: an experimental analysis. The Journal of Supercomputing, 76(9):7204– 7218, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03150-z.

Abbes:2020:DRF

[380] Heithem Abbes, Thouraya Louati, and Christophe Cérin. Dynamic replication factor model for Linux containersbased cloud systems. The Journal of Supercomputing, 76(9):7219– 7241, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03158-5.

Javed:2020:DTF

[381] Ali Javed, Khalid Mahmood Malik, and Hafiz Malik. A decision tree framework for shot classification of field sports videos. The Journal of Supercomputing, 76(9):7242– 7267, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03155-8.

Karmakar:2020:BAC

[382] Kamalesh Karmakar, Rajib K. Das, and Sunirmal Khatua. Bandwidth allocation for communicating virtual machines in cloud data centers. The Journal of Supercomputing, 76(9): 7268–7289, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03128-6.

Liang:2020:LPT

[383] Bin Liang, Xiaoshe Dong, and Xingjun Zhang. A low-power task scheduling algorithm for heterogeneous cloud computing. The Journal of Supercomputing, 76(9):7290– 7314, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03163-8.

Nomura:2020:DTE

[384] Yukihiro Nomura, Issei Sato, and Osamu Abe. Development of training environment for deep learning with medical images on supercomputer system based on asynchronous parallel Bayesian optimization. The Journal of Supercomputing, 76(9):7315– 7332, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03164-7.

Priyadarshi:2020:DTW

[385] Rahul Priyadarshi, Bharat Gupta, and Amulya Anurag. Deployment techniques in wireless sensor networks: a survey, classification, challenges, and future research issues. The Journal of Supercomputing, 76(9):7333– 7373, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03166-5.

Mandal:2020:ATD

[386] Riman Mandal, Manash Kumar Mondal, and Utpal Biswas. An approach toward design and development of an energy-aware VM selection policy with improved SLA violation in the domain of green cloud computing. The Journal of Supercomputing, 76(9): 7374–7393, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03165-6.

Bir:2020:DIE

[387] Parth Bir, Shylaja Vinaykumar Karatangi, and Amrita Rai. Design and implementation of an elastic processor with hyperthreading technology and virtualization for elastic server models. The Journal of Supercomputing, 76(9): 7394–7415, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03174-5.

Dadfarnia:2020:ICF

[388] Mahila Dadfarnia, Fazlollah Adibnia, and Ali Dorri. Incremental collusive fraud detection in large-scale online auction networks. The Journal of Supercomputing, 76(9):7416– 7437, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03170-9.

Rajabzadeh:2020:NCM

[389] Mehdi Rajabzadeh, Abolfazl Toroghi Haghighat, and Amir Masoud Rahmani. New comprehensive model based on virtual clusters and absorbing Markov chains for energyefficient virtual machine management in cloud computing. The Journal of Supercomputing, 76(9):7438– 7457, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03169-2.

Chennupati:2020:DNN

[390] Gopinath Chennupati, Raviteja Vangara, and Boian Alexandrov. Distributed non-negative matrix factorization with determination of the number of latent features. The Journal of Supercomputing, 76(9):7458– 7488, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03181-6.

Chen:2020:TIT

[391] Liangchen Chen, Shu Gao, and Zhengwei Jiang. THS-IDPC: a three-stage hierarchical sampling method based on improved density peaks clustering algorithm for encrypted malicious traffic detection. The Journal of Supercomputing, 76(9):7489– 7518, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03372-1.

Wang:2020:CRP

[392] Feng Wang, Yanjun Wu, and Feiyue Huang. Correction to: Rio: a personal storage system in multi-device and cloud. The Journal of Supercomputing, 76(9):7519, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03309-8. See [148].

Bae:2020:FCF

[393] Han Jun Bae and Lynn Choi. Filter cache: filtering useless cache blocks for a small but efficient shared last-level cache. The Journal of Supercomputing, 76(10):7521– 7544, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03177-2.

Shirmarz:2020:PIS

[394] Alireza Shirmarz and Ali Ghaffari. Performance issues and solutions in SDNbased data center: a survey. The Journal of Supercomputing, 76(10): 7545–7593, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03180-7.

Chunlin:2020:DCC

[395] Li Chunlin and Jing Zhang. Dynamic cooperative caching strategy for delay-sensitive applications in edge computing environment. The Journal of Supercomputing, 76(10):7594– 7618, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03191-4.

Xiao:2020:SFI

[396] Wen Xiao and Juan Hu. SWEclat: a frequent itemset mining algorithm over streaming data using Spark Streaming. The Journal of Supercomputing, 76(10):7619– 7634, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03190-5.

Ebadifard:2020:SSW

[397] Fatemeh Ebadifard and Seyed Morteza Babamir. Scheduling scientific workflows on virtual machines using a Pareto and hypervolume based black hole optimization algorithm. The Journal of Supercomputing, 76(10): 7635–7688, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03183-4.

Fu:2020:NPM

[398] You Fu and Wei Zhou. A novel parallel Markov clustering method in biological interaction network analysis under multi-GPU computing environment. The Journal of Supercomputing, 76(10):7689– 7706, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03193-2.

Li:2020:EEC

[399] Chunlin Li, Jianhang Tang, and Youlong Luo. Elastic edge cloud resource management based on horizontal and vertical scaling. The Journal of Supercomputing, 76(10): 7707–7732, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03192-3.

Anonymous:2020:EN

[400] Anonymous. Editor’s note. The Journal of Supercomputing, 76(10): 7733, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03336-5.

Song:2020:RMH

[401] Justin JongSu Song and Wookey Lee. Relevance maximization for high-recall retrieval problem: finding all needles in a haystack. The Journal of Supercomputing, 76(10):7734–7757, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-016-1956-
8.

Chhaya:2020:KLO

[402] PhalPheaktra Chhaya, Chi-Hwan Choi, and Young-Sung Lee. KM-LOD: linked open data service for Korean medical database. The Journal of Supercomputing, 76(10):7758– 7776, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-017-1970-5.

Shang:2020:ABC

[403] Jingwen Shang, Chaokun Wang, and Jun Qian. An attribute-based community search method with graph refining. The Journal of Supercomputing, 76(10):7777–7804, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-017-1976-
z.

Lim:2020:ECR

[404] Jongtae Lim, Kyoungsoo Bok, and Jaesoo Yoo. An efficient continuous range query processing scheme in mobile P2P networks. The Journal of Supercomputing, 76(10):7805–7819, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-017-1986-
x.

Ryu:2020:III

[405] KyungSeok Ryu, JooSeok Park, and JaeHong Park. The influence of IT investment and IT governance on corporate performance of multibusiness firms. The Journal of Supercomputing, 76(10):7820–7846, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-017-1974-
1.

Kim:2020:MTF

[406] Svetlana Kim, Suan Lee, and Yong-Ik Yoon. MRTensorCube: tensor factorization with data reduction for context-aware recommendations. The Journal of Supercomputing, 76(10): 7847–7857, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-017-2002-1.

Jeong:2020:DTF

[407] Young-Seob Jeong, Sang-Hun Lee, and Ho-Jin Choi. Discovery of topic flows of authors. The Journal of Supercomputing, 76(10):7858–7882, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-017-2065-
z.

Gao:2020:ATA

[408] Wenhao Gao, Wenjian Luo, and Chenyang Bu. Adapting the TopLeaders algorithm for dynamic social networks. The Journal of Supercomputing, 76(10):7883–7905, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-017-2063-
1.

Seo:2020:EGS

[409] Hojin Seo, Kisung Park, and Young-Koo Lee. An effective graph summarization and compression technique for a large-scaled graph. The Journal of Supercomputing, 76(10):7906– 7920, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-018-2245-5.

Jung:2020:RPC

[410] Jihee Jung and Jang Hyun Baek. Reducing paging cost of tracking area list-based mobility management in LTE network. The Journal of Supercomputing, 76(10):7921–7935, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-018-2303-
                                        z.

Kim:2020:TGE

[411] Ho-Jun Kim, Eun-Jeong Ko, and Ki-Hoon Lee. Techniques and guidelines for effective migration from RDBMS to NoSQL. The Journal of Supercomputing, 76(10):7936–7950, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-018-2361-
                                        2.

Baek:2020:ESO

[412] Nakhoon Baek. An emulation scheme for OpenGL SC 2.0 over OpenGL. The Journal of Supercomputing, 76(10): 7951–7960, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-018-2399-1.

Heinrich:2020:HAU

[413] Dennis Heinrich, Stefan Werner, and Sven Groppe. Hardware-aided update acceleration in a hybrid Semantic Web database system. The Journal of Supercomputing, 76(10): 7961–7984, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-018-2462-y.

Jo:2020:EPR

[414] Yong-Yeon Jo, Myung-Hwan Jang, and Kyungsik Han. Efficient processing of recommendation algorithms on a single-machine-based graph engine. The Journal of Supercomputing, 76(10):7985–8002, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-018-2477-
4.

Kang:2020:TAR

[415] Jaewoong Kang, Jongmo Kim, and Mye Sohn. Transition activity recognition using fuzzy logic and overlapped sliding window-based convolutional neural networks. The Journal of Supercomputing, 76(10):8003–8020, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-018-2470-
y.

Zhao:2020:PSC

[416] Linjing Zhao, Xiaonan Zhang, and Kang G. Shin. Power saving with CoMP transmission for densely deployed small cell networks. The Journal of Supercomputing, 76(10): 8021–8039, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-018-2476-5.

Lee:2020:TBP

[417] Sang Il Lee and Seong Joon Yoo. Threshold-based portfolio: the role of the threshold and its applications. The Journal of Supercomputing, 76(10): 8040–8057, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-018-2577-1.

Chun:2020:DIK

[418] Sejin Chun, Jooik Jung, and Kyong-Ho Lee. Designing an integrated knowledge graph for smart energy services. The Journal of Supercomputing, 76(10):8058–8085, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-018-2672-
3.

Khan:2020:IFD

[419] Mohammad Azam Khan, Yong-Hwa Kim, and Jaegul Choo. Intelligent fault detection using raw vibration signals via dilated convolutional neural networks. The Journal of Supercomputing, 76(10):8086–8100, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-018-2711-
0.

Jeong:2020:PML

[420] Hyeok-June Jeong, Suh-Yong Choi, and Young-Guk Ha. Probability machine-learning-based communication and operation optimization for cloud-based UAVs. The Journal of Supercomputing, 76(10):8101– 8117, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-018-2728-4.

Kang:2020:DGC

[421] Seok Kang, Suan Lee, and Jinho Kim. Distributed graph cube generation using Spark framework. The Journal of Supercomputing, 76(10): 8118–8139, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02746-4.

Chen:2020:TCT

[422] Yangjun Chen, Bin Guo, and Xingyue Huang. δ-Transitive closures and triangle consistency checking: a new way to evaluate graph pattern queries in large graph databases. The Journal of Supercomputing, 76(10):8140– 8174, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02762-4.

Lee:2020:EAE

[423] Jae-Gil Lee, Donghwan Bae, and Mun Yong Yi. An effective approach to enhancing a focused crawler using Google. The Journal of Supercomputing, 76(10):8175– 8192, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02787-9.

Kim:2020:SMP

[424] Han-Gyu Kim, Gil-Jin Jang, and Ho-Jin Choi. Speech and music pitch trajectory classification using recurrent neural networks for monaural speech segregation. The Journal of Supercomputing, 76(10):8193– 8213, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02785-x.

Jung:2020:IPA

[425] Jihee Jung, Azizbek Marakhimov, and Jang Hyun Baek. Improvement and performance analysis of a power-saving mechanism considering traffic patterns. The Journal of Supercomputing, 76(10):8214– 8224, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02798-6.

Kim:2020:MSB

[426] Jinwoo Kim, Yonggeon Choi, and Sunghoon Kim. Multi-sensor-based detection and tracking of moving objects for relative position estimation in autonomous driving conditions. The Journal of Supercomputing, 76(10): 8225–8247, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02811-y.

Battulga:2020:HTP

[427] Lkhagvadorj Battulga, Sang-Hyun Lee, and Kwan-Hee Yoo. Hashtree PCA: accelerating PCA with hash-based grouping. The Journal of Supercomputing, 76(10):8248– 8264, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-02947-x.

Kuang:2020:LCS

[428] Sicong Kuang and Brian D. Davison. Learning class-specific word embeddings. The Journal of Supercomputing, 76(10):8265–8292, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-019-03024-

z. See correction [429].

Kuang:2020:CLC

[429] Sicong Kuang and Brian D. Davison. Correction to: Learning classspecific word embeddings. The Journal of Supercomputing, 76(10): 8293, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03059-2. See [428].

Lee:2020:MDL

[430] Sang Il Lee and Seong Joon Yoo. Multimodal deep learning for finance: integrating and forecasting international stock markets. The Journal of Supercomputing, 76(10):8294– 8312, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03101-3.

Baek:2020:ARS

[431] Nakhoon Baek and Kwan-Hee Yoo. An accelerated rendering scheme for massively large point cloud data. The Journal of Supercomputing, 76(10):

8313–8323, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03114-y.

Jeong:2020:PRN

[432] Young-Seob Jeong, Kyo-Joong Oh, and Ho-Jin Choi. Pseudo-random number generation using LSTMs. The Journal of Supercomputing, 76(10): 8324–8342, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03229-7.

Zhang:2020:PPA

[433] Yaming Zhang, Yaya H. Koura, and Yanyuan Su. Predator-prey approach in modeling users’ data packets forwarding. The Journal of Supercomputing, 76(10):8343– 8356, October 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03230-0.

Byun:2020:CSM

[434] Seokhyun Byun, Seunghyun Yoon, and Kyomin Jung. Comparative studies on machine learning for paralinguistic signal compression and classification. The Journal of Supercomputing, 76(10): 8357–8371, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03346-3.

Wang:2020:HSS

[435] Zhi-Hao Wang, Gwo-Jiun Horng, and Gwo-Jia Jong. Heart sound signal recovery based on time series signal prediction using a recurrent neural network in the long shortterm memory model. The Journal of Supercomputing, 76(11):8373– 8390, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03096-x.

Zhao:2020:SLD

[436] Ping Zhao, Guanglin Zhang, and Tariq Umer. A survey of local differential privacy for securing Internet of Vehicles. The Journal of Supercomputing, 76(11):8391– 8412, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03104-0.

Ozyurt:2020:EDF

[437] Fatih Özyurt. Efficient deep feature selection for remote sensing image recognition with fused deep learning architectures. The Journal of Supercomputing, 76(11):8413– 8431, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03106-y.

Zhao:2020:IDF

[438] Lingjun Zhao, Chunhua Su, and Zhaoyang Han. Indoor device-free passive localization with DCNN for location-based services. The Journal of Supercomputing, 76(11):8432– 8449, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03110-2.

Hou:2020:LCD

[439] Jun Hou, Qianmu Li, and Ye Tian. Low-cohesion differential privacy protection for industrial Internet. The Journal of Supercomputing, 76(11): 8450–8472, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03122-y.

Lin:2020:CBF

[440] Feng-Cheng Lin, Huu-Huy Ngo, and Chyi-Ren Dow. A cloudbased face video retrieval system with deep learning. The Journal of Supercomputing, 76(11):8473– 8493, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03123-x.

Orts:2020:SUP

[441] F. Orts, G. Ortega, and E. M. Garzón. On solving the unrelated parallel machine scheduling problem: active microrheology as a case study. The Journal of Supercomputing, 76(11): 8494–8509, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03121-z.

Shahbazian:2020:ODL

[442] Reza Shahbazian and Francesca Guerriero. Optimized distributed large-scale analytics over decentralized data sources with imperfect communication. The Journal of Supercomputing, 76(11):8510– 8528, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03129-5.

Duan:2020:DSS

[443] Yuntao Duan, Jiangdai Li, and Jyh-Haw Yeh. Data storage security for the Internet of Things. The Journal of Supercomputing, 76(11): 8529–8547, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03148-7.

Acharya:2020:DCN

[444] Smarika Acharya, Abeer Alsadoon, and Anand Deva. Deep convolutional network for breast cancer classification: enhanced loss function (ELF). The Journal of Supercomputing, 76(11): 8548–8565, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03157-6.

Alzubi:2020:CDB

[445] Omar A. Alzubi, Jafar A. Alzubi, and Mohammad Alsayyed. Cryptosystem design based on Hermitian curves for IoT security. The Journal of Supercomputing, 76(11): 8566–8589, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03144-x.

Khamparia:2020:IHT

[446] Aditya Khamparia, Deepak Gupta, and Rutvij H. Jhaveri. Internet of Health Things-driven deep learning system for detection and classification of cervical cells using transfer learning. The Journal of Supercomputing, 76(11):8590– 8608, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03159-4.

Safara:2020:PPB

[447] Fatemeh Safara, Alireza Souri, and Moayad Aloqaily. PriNergy: a priority-based energy-efficient routing method for IoT systems. The Journal of Supercomputing, 76(11): 8609–8626, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03147-8.

Li:2020:BDD

[448] Xingquan Li, Cong Cao, and Tao Zhang. Block diagonal dominancebased dynamic programming for detecting community. The Journal of Supercomputing, 76(11):8627– 8640, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03151-y.

Chang:2020:TSC

[449] Ray-I Chang, Yu-Hsuan Chiu, and Jeng-Wei Lin. Two-stage classification of tuberculosis culture diagnosis using convolutional neural network with transfer learning. The Journal of Supercomputing, 76(11): 8641–8656, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03152-x.

Vidhya:2020:MAN

[450] K. Vidhya and R. Shanmugalakshmi. Modified adaptive neurofuzzy inference system (M-ANFIS)- based multi-disease analysis of healthcare Big Data. The Journal of Supercomputing, 76(11):8657–8678, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03132-w.

Pantoja:2020:AMA

[451] Maria Pantoja, Maxence Weyrich, and Gerardo Fernández-Escribano. Acceleration of MRI analysis using multicore and manycore paradigms. The Journal of Supercomputing, 76(11): 8679–8690, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03154-9.

Dwivedi:2020:IEC

[452] Shubhra Dwivedi, Manu Vardhan, and Sarsij Tripathi. Incorporating evolutionary computation for securing wireless network against cyberthreats. The Journal of Supercomputing, 76(11): 8691–8728, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03161-w.

Ghahramani:2020:SBB

[453] Meysam Ghahramani, Reza Javidan, and Mohammad Shojafar. A secure biometric-based authentication protocol for global mobility networks in smart cities. The Journal of Supercomputing, 76(11):8729– 8755, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03160-x.

Xiong:2020:CDS

[454] Lu Xiong, Guanrong Tang, and Ruey-Shun Chen. Color disease spot image segmentation algorithm based on chaotic particle swarm optimization and FCM. The Journal of Supercomputing, 76(11):8756– 8770, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03171-8.

Tomas:2020:TSQ

[455] Andrés E. Tomás and Enrique S. Quintana-Ortí. Tall-and-skinny QR factorization with approximate Householder reflectors on graphics processors. The Journal of Supercomputing, 76(11):8771–8786, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-020-03176-
                                        3.

Stafford:2020:IUH

[456] Esteban Stafford and José Luis Bosque. Improving utilization of heterogeneous clusters. The Journal of Supercomputing, 76(11):8787– 8800, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03175-4.

Deng:2020:TEC

[457] Xiaoyi Deng, Yenchun Jim Wu, and Fuzhen Zhuang. Trust-embedded collaborative deep generative model for social recommendation. The Journal of Supercomputing, 76(11):8801– 8829, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03178-1.

Munoz-Montoro:2020:SIP

[458] A. J. Muñoz-Montoro, R. Cortina, and J. Ranilla. A score identification parallel system based on audio-to-score alignment. The Journal of Supercomputing, 76(11):8830– 8844, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03185-2.

Shaer:2020:ICT

[459] Bassam Shaer, Lesley Ray, and Danita Lee Marcum. Interactive capacitive touch music table with embedded microcontrollers. The Journal of Supercomputing, 76(11):8845– 8865, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03167-4.

Paoletti:2020:SRN

[460] Mercedes E. Paoletti, Juan M. Haut, and Antonio Plaza. Scalable recurrent neural network for hyperspectral image classification. The Journal of Supercomputing, 76(11):8866– 8882, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03187-0.

Barreda:2020:PMS

[461] Maria Barreda, Manuel F. Dolz, and Enrique S. Quintana-Ortí. Performance modeling of the sparse matrix-vector product via convolutional neural networks. The Journal of Supercomputing, 76(11):8883– 8900, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03186-1.

Taghizadeh:2020:CDC

[462] Ramin Gholami Taghizadeh, Reza Gholami Taghizadeh, and Seyyed Amir Asghari. CA-Dedupe: content-aware deduplication in SSDs. The Journal of Supercomputing, 76(11):8901– 8921, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03188-z.

Zhang:2020:EMD

[463] Yong Zhang, Shao Yong Ma, and Kai Leung Yung. EDGAN: motion deblurring algorithm based on enhanced generative adversarial networks. The Journal of Supercomputing, 76(11):

8922–8937, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03189-y.

Kumar:2020:MML

[464] Gulshan Kumar, Kutub Thakur, and Maruthi Rohit Ayyagari. MLEsIDSs: machine learning-based ensembles for intrusion detection systems — a review. The Journal of Supercomputing, 76(11):8938–8971, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-020-03196-
z.

Guha:2020:DPA

[465] Krishnendu Guha, Atanu Majumder, and Amlan Chakrabarti. Dynamic power-aware scheduling of real-time tasks for FPGA-based cyber physical systems against power draining hardware trojan attacks. The Journal of Supercomputing, 76(11): 8972–9009, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03184-3.

Yan:2020:TDR

[466] Shengzan Yan, Lijun Xu, and Shushan Wang. Three-dimensional rapid registration and reconstruction of multiview rigid objects based on end-to-end deep surface model. The Journal of Supercomputing, 76(11):9010– 9030, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03194-1.

Roshandel:2020:CPE

[467] Emad Roshandel, Faraj-Allah Dolatkhahi, and Hamid Davazdah-Emami. Cyberphysical energy system toward safety in smart grids. The Journal of Supercomputing, 76(11):9031–9062, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03204-2.

Moutafis:2020:DTS

[468] B. E. Moutafis, G. A. Gravvanis, and C. K. Filelis-Papadopoulos. On the design of two-stage multiprojection methods for distributed memory systems. The Journal of Supercomputing, 76(11):9063–9094, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03201-5.

Tarafdar:2020:EQS

[469] Anurina Tarafdar, Mukta Debnath, and Rajib K. Das. Energy and quality of service-aware virtual machine consolidation in a cloud data center. The Journal of Supercomputing, 76(11): 9095–9126, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03203-3.

Su:2020:NLC

[470] Yi-Jen Su, Wu-Chih Hu, and Ruei-Ye Su. A novel LMAEB-CNN model for Chinese microblog sentiment analysis. The Journal of Supercomputing, 76(11):9127–9141, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03198-x.

Tang:2020:EMF

[471] Qi Tang, Li-Hua Zhu, and Ji-Bo Wei. An efficient multi-functional duplication-based scheduling framework for multiprocessor systems. The Journal of Supercomputing, 76(11): 9142–9167, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03208-y.

Kayed:2020:NNM

[472] Mohammed Kayed and Ahmed A. Elngar. NestMSA: a new multiple sequence alignment algorithm. The Journal of Supercomputing, 76(11): 9168–9188, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03206-0.

Romero:2020:SSM

[473] Jose C. Romero, Antonio Vilches, and Rafael Asenjo. ScrimpCo: scalable matrix profile on commodity heterogeneous processors. The Journal of Supercomputing, 76(11):9189– 9210, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03199-w.

Kehrer:2020:EEC

[474] Stefan Kehrer and Wolfgang Blochinger. Equilibrium: an elasticity controller for parallel tree search in the cloud. The Journal of Supercomputing, 76(11): 9211–9245, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03197-y. See cor-

rection [932].

Raeisi-Varzaneh:2020:PNB

[475] Mostafa Raeisi-Varzaneh and Hossein Sabaghian-Bidgoli. A Petrinet-based communication-aware modeling for performance evaluation of NOC application mapping. The Journal of Supercomputing, 76(11):9246– 9269, November 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03207-z.

Roui:2020:ESS

[476] Mohamad Beheshti Roui, S. Kazem Shekofteh, and Ahad Harati. Efficient scheduling of streams on GPGPUs. The Journal of Supercomputing, 76(11): 9270–9302, November 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03209-x.

Yang:2020:ILI

[477] Chao-Tung Yang, Yuan-An Chen, and Po-Yu Liu. Influenza-like illness prediction using a long short-term memory deep learning model with multiple open data sources. The Journal of Supercomputing, 76(12):9303– 9329, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03182-5.

Jarrah:2020:NEH

[478] Muath Ibrahim Jarrah, A. S. M. Jaya, and Ahmed Ismail Abu-Khadrah. A novel explanatory hybrid artificial bee colony algorithm for numerical function optimization. The Journal of Supercomputing, 76(12): 9330–9354, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03083-2.

Li:2020:MTP

[479] HuiFang Li and Luan Su. Multimodal transport path optimization model and algorithm considering carbon emission multitask. The Journal of Supercomputing, 76(12):9355– 9373, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03103-1.

Ding:2020:PCL

[480] Pengxin Ding, Jianping Zhang, and Minghui Wang. Pyramid context learning for object detection. The Journal of Supercomputing, 76(12): 9374–9387, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03168-3.

Costero:2020:LKS

[481] Luis Costero, Francisco D. Igual, and Francisco Tirado. Leveraging knowledge-as-a-service (KaaS) for QoS-aware resource management in multi-user video transcoding. The Journal of Supercomputing, 76(12): 9388–9403, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-019-03117-9.

Han:2020:EHA

[482] Xiaoxia Han, Lin Yue, and Xinying Xu. Efficient hybrid algorithm based on moth search and fireworks algorithm for solving numerical and constrained engineering optimization problems. The Journal of Supercomputing, 76(12):9404– 9429, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03212-2.

Vegh:2020:FHM

[483] János Végh. Finally, how many efficiencies the supercomputers have? The Journal of Supercomputing, 76(12): 9430–9455, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03210-4.

Pakkaranang:2020:SMP

[484] Nuttapol Pakkaranang, Poom Kumam, and Yusuf I. Suleiman. Superiorization methodology and perturbation resilience of inertial proximal gradient algorithm with application to signal recovery. The Journal of Supercomputing, 76(12):9456– 9477, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03215-z.

Ordonez:2020:GAR

[485] Álvaro Ordóñez, Francisco Argüello, and Begüm Demir. GPU-accelerated registration of hyperspectral images using KAZE features. The Journal of Supercomputing, 76(12):9478– 9492, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03214-0.

Tabrizchi:2020:SSC

[486] Hamed Tabrizchi and Marjan Kuchaki Rafsanjani. A survey on security challenges in cloud computing: issues, threats, and solutions. The Journal of Supercomputing, 76(12): 9493–9532, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03213-1.

Pereira:2020:SPM

[487] Paulo Pereira, Jean Araujo, and Paulo Maciel. Stochastic performance model for web server capacity planning in fog computing. The Journal of Supercomputing, 76(12): 9533–9557, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03218-w.

Deng:2020:OLL

[488] Yuefan Deng, Meng Guo, and Weifeng Liu. Optimal low-latency network topologies for cluster performance enhancement. The Journal of Supercomputing, 76(12):9558– 9584, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03216-y.

Li:2020:HPR

[489] Ruitian Li, Liang Gong, and Minghai Xu. A heterogeneous parallel Red–Black SOR technique and the numerical study on SIMPLE. The Journal of Supercomputing, 76(12): 9585–9608, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03221-1.

Al-Sinayyid:2020:JSS

[490] Ali Al-Sinayyid and Michelle Zhu. Job scheduler for streaming applications in heterogeneous distributed processing systems. The Journal of Supercomputing, 76(12):9609– 9628, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03223-z.

Moreno:2020:SBW

[491] R. Moreno, E. Arias, and F. J. Tapiador. Seeking the best weather research and forecasting model performance: an empirical score approach. The Journal of Supercomputing, 76(12): 9629–9653, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03219-9.

Lu:2020:CBF

[492] Jiawei Lu and Yain-Whar Si. Clusteringbased force-directed algorithms for 3D graph visualization. The Journal of Supercomputing, 76(12):9654– 9715, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03226-w.

Deng:2020:DCM

[493] Xiaoheng Deng, Yuebin Xu, and Xi Zheng. Dynamic clustering method for imbalanced learning based on AdaBoost. The Journal of Supercomputing, 76(12):9716– 9738, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03211-3.

Moreno-Alvarez:2020:TDN

[494] Sergio Moreno-Álvarez, Juan M. Haut, and Javier Plaza. Training deep neural networks: a static load balancing approach. The Journal of Supercomputing, 76(12):9739– 9754, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03200-6.

Zhu:2020:SFM

[495] Anqing Zhu. Spatiotemporal feature mining algorithm based on multiple minimum supports of pattern growth in Internet of Things. The Journal of Supercomputing, 76(12): 9755–9771, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03217-x.

Kaity:2020:ISA

[496] Mohammed Kaity and Vimala Balakrishnan. An integrated semi-automated framework for domain-based polarity words extraction from an unannotated non-English corpus. The Journal of Supercomputing, 76(12): 9772–9799, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03222-0.

Jiang:2020:MGH

[497] Jincheng Jiang, Jinsong Chen, and Chisheng Wang. Multi-granularity hybrid parallel network simplex algorithm for minimum-cost flow problems. The Journal of Supercomputing, 76(12): 9800–9826, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03227-9.

Santander-Jimenez:2020:GAF

[498] Sergio Santander-Jiménez, Miguel A. Vega-Rodríguez, and Leonel Sousa. GPU acceleration of Fitch’s parsimony on protein data: from Kepler to Turing. The Journal of Supercomputing, 76(12):9827–9853, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03225-x.

Moodi:2020:SAP

[499] Mahdi Moodi, Mahdieh Ghazvini, and Behnam Ghavami. A smart adaptive particle swarm optimizationsupport vector machine: Android botnet detection application. The Journal of Supercomputing, 76(12): 9854–9881, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03233-x.

Bhattacharya:2020:TFB

[500] Subhayan Bhattacharya, Sankhamita Sinha, and Amarnath Gupta. Towards finding the best-fit distribution for OSN data. The Journal of Supercomputing, 76(12):9882– 9900, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03232-y.

Sharma:2020:MAP

[501] Kapil Sharma and Vivek Kumar Sehgal. Modern architecture for photonic networks-on-chip. The Journal of Supercomputing, 76(12):9901– 9921, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03220-2.

Camara:2020:ISH

[502] Jesús Cámara, Javier Cuenca, and Domingo Giménez. Integrating software and hardware hierarchies in an autotuning method for parallel routines in heterogeneous clusters. The Journal of Supercomputing, 76(12): 9922–9941, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03235-9.

Menouer:2020:OSR

[503] Tarek Menouer, Christophe Cérin, and Ching-Hsien Hsu. Opportunistic scheduling and resources consolidation system based on a new economic model. The Journal of Supercomputing, 76(12):9942– 9975, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03231-z.

Raca:2020:CCS

[504] Valon Raca and Eduard Mehofer. clusterCL: comprehensive support for multi-kernel data-parallel applications in heterogeneous asymmetric clusters. The Journal of Supercomputing, 76(12):9976–10008, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03234-w.

Ullah:2020:EDA

[505] Ihsan Ullah and Hee Yong Youn. Efficient data aggregation with node clustering and extreme learning machine for WSN. The Journal of Supercomputing, 76(12):10009–10035, December 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03236-8.

Chang:2020:ERC

[506] Yunseok Chang. An enhanced rerouting cost estimation algorithm towards Internet of Drone. The Journal of Supercomputing, 76(12):10036– 10049, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03243-9.

Gill:2020:TCS

[507] Sukhpal Singh Gill, Xue Ouyang, and Peter Garraghan. Tails in the cloud: a survey and taxonomy of straggler management within largescale cloud data centres. The Journal of Supercomputing, 76(12):10050– 10089, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03241-x.

Chan:2020:IBS

[508] Yung-Kuan Chan, Chun-Fu Hong, and Ping-Hsuan Sun. Image-based stroke rat brain atrophy volume and infarct volume computation. The Journal of Supercomputing, 76(12):10090– 10121, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03224-y.

Tanase:2020:DSI

[509] Cristian Andy Tanase. Dynamic scheduler implementation used for load distribution between hardware accelerators (RTL) and software tasks (CPU) in heterogeneous systems. The Journal of Supercomputing, 76(12):10122– 10139, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03242-w.

Kim:2020:SSS

[510] SungJin Kim, Jaeik Cho, and Taeshik Shon. Smart seed selectionbased effective black box fuzzing for IIoT protocol. The Journal of Supercomputing, 76(12):10140– 10154, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03245-7.

Ahmadpour:2020:DIR

[511] Seyed-Sajad Ahmadpour, Mohammad Mosleh, and Saeed Rasouli Heikalabad. The design and implementation of a robust single-layer QCA ALU using a novel fault-tolerant threeinput majority gate. The Journal of Supercomputing, 76(12):10155– 10185, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03249-3.

Khanbadr:2020:NMV

[512] Asal Khanbadr, Mohammadreza Binesh Marvasti, and Amir M. Rahmani. A novel method for victim block selection for NAND flash-based solid state drives based on scoring. The Journal of Supercomputing, 76(12): 10186–10211, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03250-w.

Wu:2020:TLU

[513] Ying-Jhih Wu, Shuo-Ting Yu, and Kuo-Chan Huang. Two-level utilizationbased processor allocation for scheduling moldable jobs. The Journal of Supercomputing, 76(12):10212– 10239, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03246-6.

Haghshenas:2020:PBU

[514] Kawsar Haghshenas and Siamak Mohammadi. Prediction-based underutilized and destination host selection approaches for energy-efficient dynamic VM consolidation in data centers. The Journal of Supercomputing, 76(12): 10240–10257, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03248-4.

Majumder:2020:EEA

[515] Atanu Majumder, Sangeet Saha, and Amlan Chakrabarti. EAAM: Energy-aware application management strategy for FPGA-based IoTcloud environments. The Journal of Supercomputing, 76(12):10258– 10287, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03240-y.

Sethi:2020:FHA

[516] Krishan Kumar Sethi and Dharavath Ramesh. A fast high averageutility itemset mining with efficient tighter upper bounds and novel list structure. The Journal of Supercomputing, 76(12):10288– 10318, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03247-5.

Abedini:2020:PSN

[517] Roshanak Abedini and Reza Ravanmehr. Parallel SEN: a new approach to improve the reliability of shuffle-exchange network. The Journal of Supercomputing, 76(12):10319– 10353, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03252-8.

Fadaeddini:2020:SDP

[518] Amin Fadaeddini, Babak Majidi, and Mohammad Eshghi. Secure decentralized peer-to-peer training of deep neural networks based on distributed ledger technology. The Journal of Supercomputing, 76(12):10354– 10368, December 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03251-9.

Javanmardi:2021:UBC

[519] Abdol Karim Javanmardi, S. Hadi Yaghoubyan, and Hamid Parvin. A unit-based, cost-efficient scheduler for heterogeneous Hadoop systems. The Journal of Supercomputing, 77(1):1– 22, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03256-4.

Bermejo:2021:CQS

[520] Belen Bermejo and Carlos Juiz. On the classification and quantification of server consolidation overheads. The Journal of Supercomputing, 77(1):23– 43, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03258-2.

Constantinescu:2021:EPD

[521] Denisa-Andreea Constantinescu, Angeles Navarro, and Rafael Asenjo. Efficiency and productivity for decision making on low-power heterogeneous CPU + GPU SoCs. The Journal of Supercomputing, 77(1):44– 65, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03257-3.

Moreno:2021:PRD

[522] J. J. Moreno, J. Miroforidis, and E. M. Garzón. Parallel radiation dose computations with GENOCOP III on GPUs. The Journal of Supercomputing, 77 (1):66–76, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03254-6.

Ko:2021:CET

[523] Hung-Jui Ko, Cheng-Ta Huang, and Shiuh-Jeng Wang. Cloud evidence tracks of storage service linking with iOS systems. The Journal of Supercomputing, 77(1):77– 94, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03255-5.

Laso:2021:IDH

[524] Ruben Laso, José C. Cabaleiro, and José A. Álvarez-Dios. IHP: a dynamic heterogeneous parallel scheme for iterative or time-step methodsimage denoising as case study. The Journal of Supercomputing, 77(1):95– 110, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03260-8.

Ye:2021:DTA

[525] Fang Ye, Jie Chen, and Tao Jiang. Decentralized task allocation for heterogeneous multi-UAV system with task coupling constraints. The Journal of Supercomputing, 77(1):111– 132, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03264-4.

Raj:2021:SBA

[526] Shashi Raj, Dharavath Ramesh, and Krishan Kumar Sethi. A Sparkbased apriori algorithm with reduced shuffle overhead. The Journal of Supercomputing, 77(1):133– 151, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03253-7.

Chen:2021:ANF

[527] Jizhi Chen and Junzhong Gu. ADOL: a novel framework for automatic domain ontology learning. The Journal of Supercomputing, 77(1):152– 169, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03261-7.

Shirmohammadi:2021:FTE

[528] Zahra Shirmohammadi and Marjan Asadinia. On-fly-TOD: an efficient mechanism for crosstalk fault reduction in WNoC. The Journal of Supercomputing, 77(1):170– 187, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03259-1.

Abadeh:2021:DED

[529] Maryam Nooraei Abadeh and Mansooreh Mirzaie. DiffPageRank: an efficient differential PageRank approach in MapReduce. The Journal of Supercomputing, 77(1):188– 211, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03265-3.

Yi:2021:DMM

[530] Sheng Yi, Xiaoyu She, and Jie Li. Distributed multi-munition cooperative guidance based on clock synchronization for switching and noisy networks. The Journal of Supercomputing, 77(1):212–243, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03244-8.

Chen:2021:CFR

[531] Yi-Cheng Chen, Lin Hui, and Tipajin Thaipisutikul. A collaborative filtering recommendation system with dynamic time decay. The Journal of Supercomputing, 77(1):244– 262, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03266-2.

Stachowski:2021:ABF

[532] Matthias Stachowski, Alexander Fiebig, and Thomas Rauber. Autotuning based on frequency scaling toward energy efficiency of blockchain algorithms on graphics processing units. The Journal of Supercomputing, 77 (1):263–291, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03263-5.

Wang:2021:GBE

[533] Hongzhi Wang, Ning Li, and Jianing Li. GPU-based efficient join algorithms on Hadoop. The Journal of Supercomputing, 77(1):292– 321, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03262-6.

Huang:2021:LBI

[534] Liang-Tsung Huang, Kai-Cheng Wei, and Jian-An Wang. A lightweight BLASTP and its implementation on CUDA GPUs. The Journal of Supercomputing, 77(1):322– 342, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03267-1.

Hung:2021:EWO

[535] Jason C. Hung and Chun-Chia Wang. Exploring the website object layout of responsive web design: results of eye tracking evaluations. The Journal of Supercomputing, 77(1): 343–365, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03283-1.

Gerami:2021:BMD

[536] Marzieh Gerami, Mohammad Eshghi, and Mehdi Hosseinzadeh. A biological multiplexer, designs, and simulations. The Journal of Supercomputing, 77(1):366–387, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-019-03138-4.

Dehraj:2021:RAM

[537] Pooja Dehraj and Arun Sharma. A review on architecture and models for autonomic software systems. The Journal of Supercomputing, 77 (1):388–417, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03268-0.

Gavagsaz:2021:PCP

[538] Elaheh Gavagsaz. Parallel computation of probabilistic skyline queries using MapReduce. The Journal of Supercomputing, 77(1):418– 444, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03279-x.

Wang:2021:SHA

[539] Ti Wang, Fangming Shao, and Kunping Zhu. Structural health analysis on cyber physical system based on reliability. The Journal of Supercomputing, 77 (1):445–470, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03280-4.

Choi:2021:HEB

[540] Seokwoo Choi, Taejoo Chang, and Yongsu Park. Hybrid emulation for bypassing anti-reversing techniques and analyzing malware. The Journal of Supercomputing, 77(1):471– 497, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03270-6.

Boroumand:2021:ECA

[541] Bahador Boroumand, Elham Yaghoubi, and Behrang Barekatain. An enhanced cost-aware mapping algorithm based on improved shuffled frog leaping in network on chips. The Journal of Supercomputing, 77(1):498– 522, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03271-5.

Tiwari:2021:GTC

[542] Rohit Kumar Tiwari and Rakesh Kumar. G-TOPSIS: a cloud service selection framework using Gaussian TOPSIS for rank reversal problem. The Journal of Supercomputing, 77 (1):523–562, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03284-0.

Ider:2021:EAT

[543] Masoud Ider and Behrang Barekatain. An enhanced AHP-TOPSIS-based load balancing algorithm for switch migration in software-defined networks. The Journal of Supercomputing, 77 (1):563–596, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03285-z.

Ramezani:2021:DST

[544] Reza Ramezani. Dynamic scheduling of task graphs in multi-FPGA systems using critical path. The Journal of Supercomputing, 77(1): 597–618, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03281-3.

Munoz-Montoro:2021:PMM

[545] A. J. Muñoz-Montoro, D. Suarez-Dou, and J. Ranilla. Parallel multichannel music source separation system. The Journal of Supercomputing, 77 (1):619–637, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03282-2.

Labinghisa:2021:NNB

[546] Boney A. Labinghisa and Dong Myung Lee. Neural network-based indoor localization system with enhanced virtual access points. The Journal of Supercomputing, 77(1):638– 651, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03272-4.

Zarei:2021:ILA

[547] Bagher Zarei and Mohammad Reza Meybodi. Improving learning ability of learning automata using chaos theory. The Journal of Supercomputing, 77(1):652–678, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03293-z.

Wu:2021:SLS

[548] Hao Wu, Xin Chen, and He Guo. Scheduling large-scale scientific workflow on virtual machines with different numbers of vCPUs. The Journal of Supercomputing, 77(1):679– 710, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03273-3.

Ghobaei-Arani:2021:ERP

[549] Mostafa Ghobaei-Arani and Ali Shahidinejad. An efficient resource provisioning approach for analyzing cloud workloads: a metaheuristicbased clustering approach. The Journal of Supercomputing, 77(1):711– 750, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03296-w.

Biswas:2021:FFR

[550] Amit Biswas, Ashish Kumar Maurya, and Samir Aknine. FRLLE: a failure rate and load-based leader election algorithm for a bidirectional ring in distributed systems. The Journal of Supercomputing, 77(1): 751–779, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03286-y.

Norouzi:2021:ECT

[551] Mojtaba Norouzi, Monireh Abdoos, and Ana L. C. Bazzan. Experience classification for transfer learning in traffic signal control. The Journal of Supercomputing, 77(1): 780–795, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03287-x.

Thakur:2021:BQI

[552] Abhijeet Singh Thakur, Tarun Biswas, and Pratyay Kuila. Binary quantuminspired gravitational search algorithmbased multi-criteria scheduling for multi-processor computing systems. The Journal of Supercomputing, 77 (1):796–817, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03292-0.

Guillen:2021:PEE

[553] Miguel A. Guillén, Antonio Llanes, and José M. Cecilia. Performance evaluation of edge-computing platforms for the prediction of low temperatures in agriculture using deep learning. The Journal of Supercomputing, 77(1):818–840, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03288-w.

Gadioli:2021:TAC

[554] Davide Gadioli, Gianluca Palermo, and Cristina Silvano. Tunable approximations to control time-to-solution in an HPC molecular docking Mini-App. The Journal of Supercomputing, 77 (1):841–869, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03295-x.

Silva:2021:RMW

[555] Matheus L. Silva, Joubert de Castro Lima, and Andre L. L. Aquino. Reproducibility model for wireless sensor networks parallel simulations.

The Journal of Supercomputing, 77 (1):870–889, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03298-8.

Sharma:2021:PBJ

[556] Rashmi Sharma, Nitin Nitin, and Deepak Dahiya. Priority-based joint EDF-RM scheduling algorithm for individual real-time task on distributed systems. The Journal of Supercomputing, 77(1):890– 908, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03306-x.

Mohammadhosseini:2021:REC

[557] MohammadReza Mohammadhosseini, Sara Najafzadeh, and Ebrahim Mahdipour. Reduce energy consumption in sensors using a smartphone, smartwatch, and the use of SFLA algorithms (REC-SSS). The Journal of Supercomputing, 77(1):909–935, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03202-4.

Aghamolaei:2021:WQU

[558] Sepideh Aghamolaei, Vahideh Keikha, and Ali Mohades. Windowing queries using Minkowski sum and their extension to MapReduce. The Journal of Supercomputing, 77(1):936– 972, January 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03299-7.

Burcak:2021:NDC

[559] Kadir Can Burçak, Ömer Kaan Baykan, and Harun Uguz. A new deep convolutional neural network model for classifying breast cancer histopathological images and the hyperparameter optimisation of the proposed model. The Journal of Supercomputing, 77 (1):973–989, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03321-y.

Pai:2021:CDC

[560] Kung-Jui Pai, Ruay-Shiung Chang, and Jou-Ming Chang. Constructing dual-CISTs of pancake graphs and performance assessment of protection routings on some Cayley networks. The Journal of Supercomputing, 77 (1):990–1014, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03297-9.

Kaliyar:2021:DIF

[561] Rohit Kumar Kaliyar, Anurag Goswami, and Pratik Narang. DeepFakE: improving fake news detection using tensor decomposition-based deep neural network. The Journal of Supercomputing, 77(2):1015–1037, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03294-y.

Abedi:2021:LBE

[562] Omid Abedi and Somayeh Razaghi Kariznoi. Load-balanced and energy-aware opportunistic routing with adaptive duty cycling for multi-channel WSNs. The Journal of Supercomputing, 77(2):1038– 1058, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03301-2.

Muhammad:2021:AST

[563] Asif Muhammad and Muhammad Aleem. A3-Storm: topology-, traffic- , and resource-aware storm scheduler for heterogeneous clusters. The Journal of Supercomputing, 77(2): 1059–1093, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03289-9.

Adeli:2021:MSP

[564] Morteza Adeli and Nasour Bagheri. MDSbSP: a search protocol based on MDS codes for RFID-based Internet of Vehicle. The Journal of Supercomputing, 77(2):1094– 1113, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03302-1.

Sadhukhan:2021:LRU

[565] Dipanwita Sadhukhan, Sangram Ray, and Mou Dasgupta. A lightweight remote user authentication scheme for IoT communication using elliptic curve cryptography. The Journal of Supercomputing, 77(2):1114– 1151, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03318-7.

Shrestha:2021:ARD

[566] Laghumee Shrestha, Abeer Alsadoon, and Sami Haddad. Augmented reality for dental implant surgery: enhanced ICP. The Journal of Supercomputing, 77(2):1152–1176, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-020-03322-
                                        x.

Aksoy:2021:RGS

[567] Sinan G. Aksoy, Paul Bruillard, and Mark Raugas. Ramanujan graphs and the spectral gap of supercomputing topologies. The Journal of Supercomputing, 77(2):1177– 1213, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03291-1.

Shirani:2021:DST

[568] Mohammad Reza Shirani and Faramarz Safi-Esfahani. Dynamic scheduling of tasks in cloud computing applying dragonfly algorithm, biogeographybased optimization algorithm and Mexican hat wavelet. The Journal of Supercomputing, 77(2):1214– 1272, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03317-8.

Mostafaeipour:2021:IPH

[569] Ali Mostafaeipour, Amir Jahangard Rafsanjani, and Joshuva Arockia Dhanraj. Investigating the performance of Hadoop and Spark platforms on machine learning algorithms. The Journal of Supercomputing, 77(2): 1273–1300, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03328-5.

Jo:2021:WBA

[570] Youngho Jo, Hyunwoo Lee, and Mincheol Whang. Web behavior analysis in social life logging. The Journal of Supercomputing, 77(2):1301– 1320, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03304-z.

Ardagna:2021:PPB

[571] D. Ardagna, E. Barbierato, and J. M. Almeida. Predicting the performance of big data applications on the cloud. The Journal of Supercomputing, 77(2): 1321–1353, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03307-w.

Wang:2021:PHA

[572] Zhichen Wang, Naisheng Guo, and Yang Xu. Prediction of highway asphalt pavement performance based on Markov chain and artificial neural network approach. The Journal of Supercomputing, 77(2):1354– 1376, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03329-4.

Ginny:2021:SPA

[573] Ginny, Chiranjeev Kumar, and Kshirasagar Naik. Smartphone processor architecture, operations, and functions: current state-of-the-art and future outlook: energy performance trade-off. The Journal of Supercomputing, 77(2): 1377–1454, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03312-z.

Goey:2021:ANT

[574] Jia-Zheng Goey, Wai-Kong Lee, and Wun-She Yap. Accelerating number theoretic transform in GPU platform for fully homomorphic encryption. The Journal of Supercomputing, 77(2): 1455–1474, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03156-7.

Alves:2021:IAF

[575] Marcelo M. Alves, Raul C. Almeida, Jr., and Karcius D. R. Assis. Impairment-aware fixed-alternate BSR routing heuristics applied to elastic optical networks. The Journal of Supercomputing, 77(2):1475– 1501, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03314-x.

He:2021:EPM

[576] Wei-Jia He, Ming-Lin Yang, and Xin-Qing Sheng. Efficient parallelization of multilevel fast multipole algorithm for electromagnetic simulation on many-core SW26010 processor. The Journal of Supercomputing, 77(2): 1502–1516, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03308-9.

Sharma:2021:TBR

[577] Abhilasha Sharma and R. G. Sangeetha. Terminal and broadcast reliability analysis of direct 2-D symmetric torus network. The Journal of Supercomputing, 77(2):1517– 1536, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03311-0.

Silva:2021:SMP

[578] Francisco Airton Silva, Iure Fé, and Glauber Gonçalves. Stochastic models for performance and cost analysis of a hybrid cloud and fog architecture. The Journal of Supercomputing, 77(2): 1537–1561, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03310-1.

Jeon:2021:DNQ

[579] Jun-Cheol Jeon. Designing nanotechnology QCA-multiplexer using majority function-based NAND for quantum computing. The Journal of Supercomputing, 77(2):1562– 1578, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03341-8.

BenHalima:2021:OBP

[580] Rania Ben Halima, Slim Kallel, and Walid Gaaloul. Optimal business process deployment cost in cloud resources. The Journal of Supercomputing, 77(2):1579– 1611, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03316-9.

Oliveira:2021:TNP

[581] Daniel Oliveira, Sean Blanchard, and Paolo Rech. Thermal neutrons: a possible threat for supercomputer reliability. The Journal of Supercomputing, 77(2):1612– 1634, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03324-9.

Lin:2021:EES

[582] Fang Lin, Yi Liu, and Depei Qian. ELS: Emulation system for debugging and tuning large-scale parallel programs on small clusters. The Journal of Supercomputing, 77(2): 1635–1666, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03319-6.

Atahary:2021:PPB

[583] Tanvir Atahary, Tarek M. Taha, and Scott Douglass. Parallelized pathbased search for constraint satisfaction in autonomous cognitive agents. The Journal of Supercomputing, 77(2): 1667–1692, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03339-2.

Hammami:2021:LAA

[584] Hamza Hammami, Sadok Ben Yahia, and Mohammad S. Obaidat. A lightweight anonymous authentication scheme for secure cloud computing services. The Journal of Supercomputing, 77(2):1693–1713, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03313-y.

Khan:2021:ODU

[585] Angshuman Khan and Rajeev Arya. Optimal demultiplexer unit design and energy estimation using quantum dot cellular automata. The Journal of Supercomputing, 77(2): 1714–1738, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03320-z.

Bharti:2021:OCB

[586] Monika Bharti and Himanshu Jindal. Optimized clustering-based discovery framework on Internet of Things. The Journal of Supercomputing, 77(2): 1739–1778, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03315-w.

Ngueilbaye:2021:AHM

[587] Alladoumbaye Ngueilbaye, Hongzhi Wang, and Daouda Ahmat Mahamat. Adoption of human metabolic processes as Data Quality Based Models. The Journal of Supercomputing, 77(2): 1779–1817, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03300-3.

Sole-Beteta:2021:ATS

[588] Xavier Solé-Beteta, Joan Navarro, and Alan Briones. Automatic tutoring system to support cross-disciplinary training in Big Data. The Journal of Supercomputing, 77(2):1818– 1852, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03330-x.

Li:2021:NSB

[589] Chunlin Li, YiHan Zhang, and Youlong Luo. Neighborhood search-based job scheduling for IoT big data realtime processing in distributed edgecloud computing environment. The Journal of Supercomputing, 77(2): 1853–1878, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03343-6.

Cores:2021:HTB

[590] Fernando Cores, Fernando Guirado, and Josep Lluis Lerida. High throughput BLAST algorithm using Spark and Cassandra. The Journal of Supercomputing, 77(2):1879– 1896, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03338-3.

Talib:2021:SLR

[591] Manar Abu Talib, Sohaib Majzoub, and Dina Jamal. A systematic literature review on hardware implementation of artificial intelligence algorithms. The Journal of Supercomputing, 77(2): 1897–1938, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03325-8.

Saadatmand:2021:TAT

[592] Faezeh Sadat Saadatmand, Nezam Rohbani, and Hamed Farbeh. TAMER: an adaptive task allocation method for aging reduction in multi-core embedded real-time systems. The Journal of Supercomputing, 77(2):1939– 1957, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03326-7.

Chen:2021:RNB

[593] Yu qiang Chen, Jian lan Guo, and Hong ling Liu. Research on navigation of bidirectional A∗ algorithm based on ant colony algorithm. The Journal of Supercomputing, 77(2):

1958–1975, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03303-0.

Magee:2021:ASR

[594] Daniel J. Magee, Anthony S. Walker, and Kyle E. Niemeyer. Applying the swept rule for solving explicit partial differential equations on heterogeneous computing systems. The Journal of Supercomputing, 77(2): 1976–1997, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03340-9.

Deepa:2021:ABI

[595] N. Deepa, B. Prabadevi, and Usman Tariq. An AI-based intelligent system for healthcare analysis using Ridge-Adaline Stochastic Gradient Descent Classifier. The Journal of Supercomputing, 77(2):1998– 2017, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03347-2.

Huang:2021:DPS

[596] Chenxi Huang, Hao Wang, and Fan Lin. A dynamic priority strategy for IoV data scheduling towards key data. The Journal of Supercomputing, 77(2): 2018–2032, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03350-7.

Saeidi:2021:HPD

[597] Mahmoud Saeidi and Ali Ahmadi. High-performance and deep pedestrian detection based on estimation of different parts. The Journal of Supercomputing, 77(2):2033– 2068, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03345-4.

Min-Allah:2021:DRT

[598] Nasro Min-Allah, Muhammad Bilal Qureshi, and Javid Taheri. Deployment of real-time systems in the cloud environment. The Journal of Supercomputing, 77(2):2069– 2090, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03334-7.

Miyata:2021:RTA

[599] Takafumi Miyata. A Riccati-type algorithm for solving generalized Hermitian eigenvalue problems. The Journal of Supercomputing, 77(2):2091– 2102, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03331-w.

Liu:2021:AFR

[600] Yong Liu, Huaxi Gu, and Yunhao Wang. An adaptive failure recovery mechanism based on asymmetric routing for data center networks. The Journal of Supercomputing, 77(2): 2103–2123, February 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03337-4.

Belloch:2021:ECP

[601] Jose A. Belloch, Germán León, and Enrique San Millan. Evaluating the computational performance of the Xilinx Ultrascale+ EG Heterogeneous MPSoC. The Journal of Supercomputing, 77(2):2124– 2137, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03342-7.

Chung:2021:PFA

[602] Moonyoung Chung and Woong-Kee Loh. α-Probabilistic flexible aggregate nearest neighbor search in road networks using landmark multidimensional scaling. The Journal of Supercomputing, 77(2):2138– 2153, February 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03521-6.

Huang:2021:CPS

[603] Anzhong Huang, Jie Cao, and Huimei Zhang. Construction of patient service system based on QFD in Internet of Things. The Journal of Supercomputing, 77(3):2155– 2171, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03359-y.

Chehelgerdi-Samani:2021:PAP

[604] Maryam Chehelgerdi-Samani and Faramarz Safi-Esfahani. PCVM.ARIMA: predictive consolidation of virtual machines applying ARIMA method. The Journal of Supercomputing, 77(3): 2172–2206, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03354-3.

Khomami:2021:CCB

[605] Mohammad Mehdi Daliri Khomami, Alireza Rezvanian, and Alireza Bagheri. CFIN: a community-based algorithm for finding influential nodes in complex social networks. The Journal of Supercomputing, 77(3):2207– 2236, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03355-2.

Hassan:2021:EED

[606] Mohammad Mehedi Hassan, Sana Ullah, and Abdulhameed Alelaiwi. An end-to-end deep learning model for human activity recognition from highly sparse body sensor data in Internet of Medical Things environment. The Journal of Supercomputing, 77 (3):2237–2250, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03361-4.

Sinha:2021:BFL

[607] Bam Bahadur Sinha and R. Dhanalakshmi. Building a fuzzy logic-based McCulloch–Pitts neuron recommendation model to uplift accuracy. The Journal of Supercomputing, 77(3): 2251–2267, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03344-5.

Wu:2021:OHB

[608] Hao Wu, Youlong Luo, and Chunlin Li. Optimization of heat-based cache replacement in edge computing system. The Journal of Supercomputing, 77(3):2268–2301, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03356-1.

Fernandez:2021:ISC

[609] Álvaro Fernández, Camino Fernández, and Miguel Á. Conde. Integrating supercomputing clusters into education: a case study in biotechnology. The Journal of Supercomputing, 77 (3):2302–2325, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03360-5.

Ghazvini:2021:NML

[610] Golnaz Aghaee Ghazvini, Mehran Mohsenzadeh, and Amir Masoud Rahmani. A new multi-level trust management framework (MLTM) for solving the invalidity and sparse problems of user feedback ratings in cloud environments. The Journal of Supercomputing, 77(3):2326–2354, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03348-1.

Balalavi:2021:ATI

[611] Maiya Balalavi, Hsiao-Ching Huang, and Yi-Maun Subeq. Applying Taiwanese indigenous health literacy for designing an elders’ prevention fall course: a statistical analysis and deep learning approach. The Journal of Supercomputing, 77(3):2355– 2382, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03358-z.

Dehkordi:2021:DAD

[612] Afsaneh Banitalebi Dehkordi, MohammadReza Soltanaghaei, and Farsad Zamani Boroujeni. The DDoS attacks detection through machine learning and statistical methods in SDN. The Journal of Supercomputing, 77 (3):2383–2415, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03323-w.

Aaker:2021:EFW

[613] Ole Edvard Aaker, Espen Birger Raknes, and Børge Arntsen. Elastodynamic full waveform inversion on GPUs with time-space tiling and wavefield reconstruction. The Journal of Supercomputing, 77(3):2416– 2457, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03352-5.

Meneses-Viveros:2021:ECM

[614] Amilcar Meneses-Viveros, Mireya Paredes-López, and Isidoro Gitler. Energy consumption model in multicore architectures with variable frequency. The Journal of Supercomputing, 77 (3):2458–2485, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03349-0.

Yang:2021:PBD

[615] Chao-Tung Yang, Jung-Chun Liu, and Chan-Fu Kuo. Performance benchmarking of deep learning framework on Intel Xeon Phi. The Journal of Supercomputing, 77(3):2486– 2510, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03362-3.

Eftekhari:2021:RSK

[616] Seyed Abdolreza Eftekhari, Morteza Nikooghadam, and Masoud Rafighi. Robust session key generation protocol for social Internet of Vehicles with enhanced security provision. The Journal of Supercomputing, 77 (3):2511–2544, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03363-2.

Mokarram:2021:CCF

[617] Marzieh Mokarram and Mohammad R. Khosravi. A cloud computing framework for analysis of agricultural big data based on Dempster–Shafer theory. The Journal of Supercomputing, 77(3):2545–2565, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03366-z.

Valentim:2021:MBS

[618] Thiago Valentim and Gustavo Callou. A model-based strategy for quantifying the impact of availability on the energy flow of data centers. The Journal of Supercomputing, 77(3):2566–2589, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03353-4.

Hosseinzadeh:2021:RDA

[619] Mehdi Hosseinzadeh, Jalil Koohpayehzadeh, and Mahdi Bohlouli. A review on diagnostic autism spectrum disorder approaches based on the Internet of Things and machine learning. The Journal of Supercomputing, 77(3):2590–2608, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03357-0.

Fernandez:2021:CSW

[620] Lisardo Fernández, Mariano Pérez, and Juan M. Orduña. A comparison study of wavelet transforms for the visualization of differentially methylated regions in DNA samples. The Journal of Supercomputing, 77 (3):2609–2623, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03269-z.

Aslinezhad:2021:ANF

[621] Mehdi Aslinezhad, Alireza Malekijavan, and Peyman Abbasi. Adaptive neuro-fuzzy modeling of a soft finger-like actuator for cyber-physical industrial systems. The Journal of Supercomputing, 77(3):2624– 2644, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03370-3.

Li:2021:CRC

[622] Pengzhi Li, Jianqiang Li, and Haihua Xie. Classification and recognition of computed tomography images using image reconstruction and information fusion methods. The Journal of Supercomputing, 77(3): 2645–2666, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03367-y.

Nalci:2021:IFH

[623] Yigitcan Nalci, Pinar Kullu, and Ozcan Ozturk. ILP formulation and heuristic method for energy-aware application mapping on 3D-NoCs. The Journal of Supercomputing, 77(3): 2667–2680, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03365-0.

Barina:2021:CVC

[624] David Barina. Convergence verification of the Collatz problem. The Journal of Supercomputing, 77(3): 2681–2688, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03368-x.

Oz:2021:SPI

[625] Dindar Öz and Isil Öz. Scalable parallel implementation of migrating birds optimization for the multi-objective task allocation problem. The Journal of Supercomputing, 77(3):2689– 2712, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03369-w.

Ashrafi-Payaman:2021:GGS

[626] Nosratali Ashrafi-Payaman, Mohammad Reza Kangavari, and Amir Mohammad Fander. GS4: Graph stream summarization based on both the structure and semantics. The Journal of Supercomputing, 77(3):2713– 2733, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03290-2.

Varshney:2021:RCU

[627] Shubham Varshney, Pankaj Charpe, and S. K. Pal. Relation collection using Pollard special-q sieving to solve integer factorization and discrete logarithm problem. The Journal of Supercomputing, 77(3):2734– 2769, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03351-6.

Carrascal:2021:FET

[628] Ginés Carrascal, Alberto A. del Barrio, and Guillermo Botella. First experiences of teaching quantum computing. The Journal of Supercomputing, 77(3):2770–2799, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03376-x.

Asghari:2021:TSR

[629] Ali Asghari, Mohammad Karim Sohrabi, and Farzin Yaghmaee. Task scheduling, resource provisioning, and load balancing on scientific workflows using parallel SARSA reinforcement learning agents and genetic algorithm. The Journal of Supercomputing, 77 (3):2800–2828, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03364-1.

He:2021:MSR

[630] Tongdi He and Shengxin Wang. Multispectral remote sensing land-cover classification based on deep learning methods. The Journal of Supercomputing, 77(3):2829–2843, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03377-w.

Too:2021:NFR

[631] Jingwei Too and Abdul Rahim Abdullah. A new and fast rival genetic algorithm for feature selection. The Journal of Supercomputing, 77 (3):2844–2874, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03378-9.

Du:2021:NAD

[632] Hongle Du and Yan Zhang. Network anomaly detection based on selective ensemble algorithm. The Journal of Supercomputing, 77(3):2875– 2896, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03374-z.

Yang:2021:SPA

[633] Yunchao Yang and S. Balachandar. A scalable parallel algorithm for direct-forcing immersed boundary method for multiphase flow simulation on spectral elements. The Journal of Supercomputing, 77(3): 2897–2927, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03371-2.

Wang:2021:TPM

[634] Jingjuan Wang and Qingkui Chen. A traffic prediction model based on multiple factors. The Journal of Supercomputing, 77(3):2928– 2960, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03373-0.

Chen:2021:CAB

[635] Huan Chen, Yue-Hsien Wang, and Chun-Hung Fan. A convolutional autoencoder-based approach with batch normalization for energy disaggregation. The Journal of Supercomputing, 77(3):2961–2978, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-020-03375-
y.

Lee:2021:SDA

[636] Chankyun Lee, Minseok Jang, and Woojin Seok. Scalable design and algorithm for science DMZ by considering the nature of research traffic. The Journal of Supercomputing, 77 (3):2979–2997, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03383-y.

Bui:2021:EEC

[637] Dinh-Mao Bui, Nguyen Anh Tu, and Eui-Nam Huh. Energy efficiency in cloud computing based on mixture power spectral density prediction. The Journal of Supercomputing, 77 (3):2998–3023, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03380-1.

Wang:2021:IMT

[638] Eric Ke Wang, Fan Wang, and Chien-Ming Chen. Intelligent monitor for typhoon in IoT system of smart city. The Journal of Supercomputing, 77 (3):3024–3043, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03381-0.

Kanwar:2021:DHB

[639] Vivek Kanwar and Ashok Kumar. DVhop-based range-free localization algorithm for wireless sensor network using runner-root optimization. The Journal of Supercomputing, 77(3): 3044–3061, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03385-w.

Park:2021:EOD

[640] Cheong Hee Park and Jiil Kim. An explainable outlier detection method using region-partition trees. The Journal of Supercomputing, 77(3): 3062–3076, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03384-x.

Hoorfar:2021:NCC

[641] Hamid Hoorfar and Alireza Bagheri. NP-completeness of chromatic orthogonal art gallery problem. The Journal of Supercomputing, 77(3):3077– 3109, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03379-8.

Fredj:2021:EBA

[642] Nissaf Fredj, Yessine Hadj Kacem, and Mohamed Abid. An eventbased approach for formally verifying runtime adaptive real-time systems. The Journal of Supercomputing, 77 (3):3110–3143, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03386-9.

Lu:2021:IEU

[643] Ching-Ta Lu, Ling-Ling Wang, and Jia-An Lin. Image enhancement using deep-learning fully connected neural network mean filter. The Journal of Supercomputing, 77(3):3144– 3164, March 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03389-6.

Sojoodi:2021:IGG

[644] Amir Hossein Sojoodi, Majid Salimi Beni, and Farshad Khunjush. Igniteg-GPU: a GPU-enabled in-memory computing architecture on clusters. The Journal of Supercomputing, 77(3): 3165–3192, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03390-z.

Nanehkaran:2021:ACM

[645] Y. A. Nanehkaran, Defu Zhang, and Najla Al-Nabhan. Analysis and comparison of machine learning classifiers and deep neural networks techniques for recognition of Farsi handwritten digits. The Journal of Supercomputing, 77(4):3193–3222, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03388-7.

Xia:2021:NMA

[646] Limin Xia and Zhenmin Li. A new method of abnormal behavior detection using LSTM network with temporal attention mechanism. The Journal of Supercomputing, 77(4):3223– 3241, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03391-y.

Vallathan:2021:SAD

[647] G. Vallathan, A. John, and Jerry Chun-Wei Lin. Suspicious activity detection using deep learning in secure assisted living IoT environments. The Journal of Supercomputing, 77 (4):3242–3260, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03387-8.

Corral-Garcia:2021:ECD

[648] Javier Corral-García, Felipe Lemus-Prieto, and Miguel-Ángel Pérez-Toledano. Efficient code development for improving execution performance in high-performance computing centers. The Journal of Supercomputing, 77(4):3261–3288, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03382-z.

Lee:2021:PIN

[649] Wai-Kong Lee, Sedat Akleylek, and Seong-Oun Hwang. Parallel implementation of Nussbaumer algorithm and number theoretic transform on a GPU platform: application to qTESLA. The Journal of Supercomputing, 77 (4):3289–3314, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03392-x.

Soria-Pardos:2021:UMC

[650] Víctor Soria-Pardos, Adrià Armejach, and Miquel Moretó. On the use of many-core Marvell ThunderX2 processor for HPC workloads. The Journal of Supercomputing, 77(4):3315– 3338, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03397-6.

Jeong:2021:EMS

[651] Yoon-Su Jeong and Byeong-Tae Ahn. An efficient management scheme of blockchain-based cloud user information using probabilistic weighting. The Journal of Supercomputing, 77 (4):3339–3358, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03396-7.

Chen:2021:STF

[652] Yulei Chen and Jianhua Chen. A secure three-factor-based authentication with key agreement protocol for e-health clouds. The Journal of Supercomputing, 77(4):3359– 3380, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03395-8.

Ujjwal:2021:SSM

[653] KC Ujjwal, Sudheer Kumar Battula, and Alexander Brown. SMOaaS: a scalable matrix operation as a service model in cloud. The Journal of Supercomputing, 77(4):3381– 3401, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03400-0.

Cheng:2021:GSF

[654] Chun Cheng, Yan Hu, and Michael Pecht. Generalized sparse filtering for rotating machinery fault diagnosis. The Journal of Supercomputing, 77(4):3402–3421, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03398-5.

Neeraj:2021:CAP

[655] Neeraj, Major Singh Goraya, and Damanpreet Singh. A comparative analysis of prominently used MCDM methods in cloud environment. The Journal of Supercomputing, 77(4): 3422–3449, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03393-w.

Zhu:2021:LRH

[656] Jiang Zhu, Lizan Wang, and Zhetao Li. A low redundancy and high time efficiency large-scale task assignment strategy for heterogeneous service-oriented cloud computing systems. The Journal of Supercomputing, 77(4):3450–3483, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03403-x.

Kalinnik:2021:PEO

[657] Natalia Kalinnik, Robert Kiesel, and Gudula Rünger. A performanceand energy-oriented extended tuning process for time-step-based scientific applications. The Journal of Supercomputing, 77(4):3484– 3515, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03402-y.

Karbasi:2021:SLS

[658] Amir Hassani Karbasi and Siyamak Shahpasand. SINGLETON: a lightweight and secure end-to-end encryption protocol for the sensor networks in the Internet of Things based on cryptographic ratchets. The Journal of Supercomputing, 77(4):3516– 3554, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03411-x.

Zhang:2021:CTD

[659] Jian Zhang, Feihu Feng, and Wanjuan Song. A compensation textures dehazing method for water alike area. The Journal of Supercomputing, 77 (4):3555–3570, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03406-8.

Atefinia:2021:NID

[660] Ramin Atefinia and Mahmood Ahmadi. Network intrusion detection using multi-architectural modular deep neural network. The Journal of Supercomputing, 77(4):3571– 3593, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03410-y.

Tseng:2021:DSS

[661] Kuo-Kun Tseng, Ran Zhang, and Mohammad Mehedi Hassan. DNetUnet: a semi-supervised CNN of medical image segmentation for supercomputing AI service. The Journal of Supercomputing, 77(4):3594– 3615, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03407-7.

Hosseinzadeh:2021:MMP

[662] Mehdi Hosseinzadeh, Omed Hassan Ahmed, and Hsiu-Sen Chiang. A multiple multilayer perceptron neural network with an adaptive learning algorithm for thyroid disease diagnosis in the Internet of Medical Things. The Journal of Supercomputing, 77 (4):3616–3637, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03404-w.

Nezarat:2021:IMP

[663] Amin Nezarat and N. Seifadini. An improved model for predicting trip mode distribution using convolution deep learning. The Journal of Supercomputing, 77(4):3638– 3652, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03394-9.

Elharrouss:2021:GRP

[664] Omar Elharrouss, Noor Almaadeed, and Ahmed Bouridane. Gait recognition for person re-identification. The Journal of Supercomputing, 77 (4):3653–3672, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03409-5.

Fan:2021:ABS

[665] Jiaqi Fan, Zhanyou Ma, and Changzhen Zhang. Analysis of blockchain system based on M/M_{1} /M_{2} vacation queueing model. The Journal of Supercomputing, 77(4):3673–3694, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03408-6.

Jang:2021:DDG

[666] Kyoung ae Jang and Woo-Je Kim. Development of data governance components using DEMATEL and content analysis. The Journal of Supercomputing, 77(4):3695–3709, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-020-03405-
                                        9.

Safara:2021:RRE

[667] Fatemeh Safara and Asri Ranga Abdullah Ramaiah. RenyiBS: Renyi entropy basis selection from wavelet packet decomposition tree for phonocardiogram classification. The Journal of Supercomputing, 77(4):3710– 3726, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03413-9.

Basumatary:2021:CBR

[668] Habila Basumatary, Arindam Debnath, and Bidyut Kumar Bhattacharyya. Centroid-based routing protocol with moving sink node for uniform and non-uniform distribution of wireless sensor nodes. The Journal of Supercomputing, 77(4): 3727–3751, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03414-8.

Zaghari:2021:ILS

[669] Nayereh Zaghari, Mahmood Fathy, and Mohammad Shahverdy. Improving the learning of self-driving vehicles based on real driving behavior using deep neural network techniques. The Journal of Supercomputing, 77 (4):3752–3794, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03399-4.

Balakrishnan:2021:SSA

[670] Vimala Balakrishnan, Pik Yin Lok, and Hajar Abdul Rahim. A semi-supervised approach in detecting sentiment and emotion based on digital payment reviews. The Journal of Supercomputing, 77(4):3795–3810, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03412-w.

Sharma:2021:BNN

[671] Himanshu Sharma and Elise Jennings. Bayesian neural networks at scale: a performance analysis and pruning study. The Journal of Supercomputing, 77(4):3811–3839, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03401-z.

Steno:2021:NER

[672] Priscilla Steno, Abeer Alsadoon, and Omar Hisham Alsadoon. A novel enhanced region proposal network and modified loss function: threat object detection in secure screening using deep learning. The Journal of Supercomputing, 77(4):3840– 3869, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03418-4.

Tang:2021:CVR

[673] Siyuan Tang and Feifei Yu. Construction and verification of retinal vessel segmentation algorithm for color fundus image under BP neural network model. The Journal of Supercomputing, 77(4):3870–3884, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03422-8.

Yang:2021:SFT

[674] Yuxing Yang, Xiaohui Li, and Jing Li. Structure fault tolerance of balanced hypercubes. The Journal of Supercomputing, 77(4):3885– 3898, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03419-3.

Guerrero-Higueras:2021:FLP

[675] Ángel Manuel Guerrero-Higueras, Lídia Sánchez-González, and Manuel Castejón-Limas. Facilitating the learning process in parallel computing by using instant messaging. The Journal of Supercomputing, 77(4):3899– 3913, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03416-6.

Zhang:2021:LSE

[676] Shasha Zhang, Dan Chen, and Albert Y. Zomaya. A lightweight solution to epileptic seizure prediction based on EEG synchronization measurement. The Journal of Supercomputing, 77 (4):3914–3932, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03426-4.

Toumi:2021:SID

[677] Lyazid Toumi and Ahmet Ugur. Static and incremental dynamic approaches for multi-objective bitmap join indexes selection in data warehouses. The Journal of Supercomputing, 77 (4):3933–3958, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03423-7.

Zakarya:2021:PTS

[678] Muhammad Zakarya, Lee Gillam, and Izaz Ur Rahman. PerficientCloudSim: a tool to simulate large-scale computation in heterogeneous clouds. The Journal of Supercomputing, 77(4): 3959–4013, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03425-5.

Flores-Contreras:2021:PPP

[679] Jesus Flores-Contreras, Hector A. Duran-Limon, and Sergio H. Almanza-Ruiz. Performance prediction of parallel applications: a systematic literature review. The Journal of Supercomputing, 77(4):4014– 4055, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03417-5.

Patonico:2021:TIE

[680] Simone Patonico, Thanh-Long Nguyen, and Kris Steenhaut. Toward the inclusion of end-to-end security in the OM2M platform. The Journal of Supercomputing, 77(4):4056– 4080, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03415-7.

Zuo:2021:HCR

[681] Haojia Zuo, Bo Cao, and Yan Huang. High-capacity ride-sharing via shortest path clustering on large road networks. The Journal of Supercomputing, 77(4):4081–4106, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03424-6.

Danesh:2021:EBC

[682] Malihe Danesh, Morteza Dorrigiv, and Farzin Yaghmaee. Ensemblebased clustering of large probabilistic graphs using neighborhood and distance metric learning. The Journal of Supercomputing, 77(4):4107– 4134, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03429-1.

Rajan:2021:LBD

[683] R. Sundara Rajan, Thomas Kalinowski, and T. M. Rajalaxmi. Lower bounds for dilation, wirelength, and edge congestion of embedding graphs into hypercubes. The Journal of Supercomputing, 77(4):4135– 4150, April 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03420-w.

Zhou:2021:UIA

[684] Hongzhen Zhou, Shuyuan Wang, and Kevin Yang. Ultrasound image analysis technology under deep belief networks in evaluation on the effects of diagnosis and chemotherapy of cervical cancer. The Journal of Supercomputing, 77(4):4151–4171, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03421-9.

Zameni:2021:TPN

[685] Mahsa Zameni, Aref Rezaei, and Leili Farzinvash. Two-phase node deployment for target coverage in rechargeable WSNs using genetic algorithm and integer linear programming. The Journal of Supercomputing, 77(4):4172–4200, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03431-7.

Rane:2021:NAF

[686] Chinmay Rane, Raj Mehrotra, and Mahua Bhattacharya. A novel attention fusion network-based framework to ensemble the predictions of CNNs for lymph node metastasis detection. The Journal of Supercomputing, 77 (4):4201–4220, April 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03432-6.

Wu:2021:FAP

[687] Xinfang Wu, Yong Xiang, and Xinzhi Zhou. Forecasting air passenger traffic flow based on the two-phase learning model. The Journal of Supercomputing, 77(5):4221–4243, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03428-2.

Lilian:2021:ANM

[688] J. Felicia Lilian, K. Sundarakantham, and S. Mercy Shalinie. Antinegation method for handling negation words in question answering system. The Journal of Supercomputing, 77(5):4244–4266, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03437-1.

Menouer:2021:KKC

[689] Tarek Menouer. KCSS: Kubernetes container scheduling strategy. The Journal of Supercomputing, 77 (5):4267–4293, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03427-3.

He:2021:NFU

[690] Qunfang He, Wenjie Chen, and Zhilei Chai. A novel framework for UAV returning based on FPGA. The Journal of Supercomputing, 77(5): 4294–4316, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03434-4.

Ferdinandy:2021:ESI

[691] Bence Ferdinandy, Ángel Manuel Guerrero-Higueras, and Ádám Miklósi. Exploratory study of introducing HPC to non-ICT researchers: institutional strategy is possibly needed for widespread adaption. The Journal of Supercomputing, 77(5):4317– 4331, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03438-0.

Sharma:2021:SDS

[692] Sumit Sharma and Sudip Roy. A survey on design and synthesis techniques for photonic integrated circuits. The Journal of Supercomputing, 77 (5):4332–4374, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03430-8.

Hashimoto:2021:ATP

[693] Takako Hashimoto, David Lawrence Shepard, and Takeaki Uno. Analyzing temporal patterns of topic diversity using graph clustering. The Journal of Supercomputing, 77(5): 4375–4388, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03433-5.

Louati:2021:DLC

[694] Ali Louati, Hassen Louati, and Zhaojian Li. Deep learning and case-based reasoning for predictive and adaptive traffic emergency management. The Journal of Supercomputing, 77 (5):4389–4418, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03435-3.

Wang:2021:AAM

[695] Suhua Wang, Lisa Zhang, and Yu Zhao. Attribute-aware multitask recommendation. The Journal of Supercomputing, 77(5):4419– 4437, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03440-6.

Asadinia:2021:SNB

[696] Sanaz Asadinia, Mahdi Mehrabi, and Elham Yaghoubi. Surix: Non-blocking and low insertion loss micro-ring resonator-based optical router for photonic network on chip. The Journal of Supercomputing, 77(5):4438– 4460, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03442-4.

Qin:2021:ITL

[697] Haoshu Qin and Huimei Zhang. Intelligent traffic light under fog computing platform in data control of real-time traffic flow. The Journal of Supercomputing, 77(5):4461– 4483, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03443-3.

Patel:2021:IGM

[698] Yashwant Singh Patel, Anjali Baheti, and Rajiv Misra. Interval graph multi-coloring-based resource reservation for energy-efficient containerized cloud data centers. The Journal of Supercomputing, 77(5):4484– 4532, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03439-z.

Han:2021:TET

[699] Qingchang Han, Hailong Yang, and Depei Qian. Towards efficient tile low-rank GEMM computation on Sunway many-core processors. The Journal of Supercomputing, 77(5):4533– 4564, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03444-2.

Tu:2021:SLM

[700] Yu-Ju Tu, Gaurav Kapoor, and Selwyn Piramuthu. Security of lightweight mutual authentication protocols. The Journal of Supercomputing, 77(5): 4565–4581, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03448-y.

Leon:2021:APA

[701] Betzabeth León, Daniel Franco, and Emilio Luque. Analysis of parallel application checkpoint storage for system configuration. The Journal of Supercomputing, 77(5):4582– 4617, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03445-1.

Chapaneri:2021:MLG

[702] Radhika Chapaneri and Seema Shah. Multi-level Gaussian mixture modeling for detection of malicious network traffic. The Journal of Supercomputing, 77(5):4618–4638, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03447-z.

Li:2021:UIT

[703] Bin Li, Ruey-Shun Chen, and C.- Y. Liu. Using intelligent technology and real-time feedback algorithm to improve manufacturing process in IoT semiconductor industry. The Journal of Supercomputing, 77 (5):4639–4658, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03457-x.

Gudivada:2021:NOT

[704] A. Arunkumar Gudivada and Gnanou Florence Sudha. Novel optimized treebased stack-type architecture for 2nbit comparator at nanoscale with energy dissipation analysis. The Journal of Supercomputing, 77(5):4659– 4680, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03453-1.

Chou:2021:PAD

[705] Chao-Lung Chou. Presentation attack detection based on score level fusion and challenge-response technique. The Journal of Supercomputing, 77 (5):4681–4697, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03461-1.

Wang:2021:GRH

[706] Yuzhu Wang, Mingxin Guo, and Jinrong Jiang. GPUs-RRTMG LW: high-efficient and scalable computing for a longwave radiative transfer model on multiple GPUs. The Journal of Supercomputing, 77(5): 4698–4717, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03451-3.

Talebi:2021:AAS

[707] Mohammadmehdi Talebi, Mohsen Sharifi, and Mohammadhesam Kalantari. ACEP: an adaptive strategy for proactive and elastic processing of complex events. The Journal of Supercomputing, 77(5):4718– 4753, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03454-0.

Hayouni:2021:NEE

[708] Haythem Hayouni and Mohamed Hamdi. A novel energy-efficient encryption algorithm for secure data in WSNs. The Journal of Supercomputing, 77 (5):4754–4777, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03465-x.

Sharma:2021:SER

[709] Saurabh Sharma and Vinod Kumar Verma. Security explorations for routing attacks in low power networks on Internet of Things. The Journal of Supercomputing, 77(5): 4778–4812, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03471-z.

Wong:2021:TFF

[710] Kok-Seng Wong, Yee Jian Chew, and Ying Han Pang. Toward forecasting future day air pollutant index in Malaysia. The Journal of Supercomputing, 77(5):4813– 4830, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03463-z.

Lee:2021:ESS

[711] Chun-Chia Lee, Hsiu-Sen Chiang, and Meng-Hsing Hsiao. Effects of screen size and visual presentation on visual fatigue based on regional brain wave activity. The Journal of Supercomputing, 77(5):4831–4851, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03458-w.

Park:2021:EIM

[712] Dong won Park, Seokhie Hong, and Sung Min Cho. Efficient implementation of modular multiplication over 192-bit NIST prime for 8-bit AVR-based sensor node. The Journal of Supercomputing, 77(5):4852– 4870, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03441-5.

Ni:2021:GSC

[713] Minghao Ni, Hongqiang Cheng, and Jinfeng Lai. GAN-SOM: a clustering framework with SOM-similar network based on deep learning. The Journal of Supercomputing, 77(5): 4871–4886, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03464-y.

Jazayeri:2021:LAE

[714] Fatemeh Jazayeri, Ali Shahidinejad, and Mostafa Ghobaei-Arani. A latency-aware and energy-efficient computation offloading in mobile fog computing: a hidden Markov model-based approach. The Journal of Supercomputing, 77(5):4887– 4916, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03476-8.

Mishra:2021:BTB

[715] Rahul Mishra, Dharavath Ramesh, and Damodar Reddy Edla. BB-tree based secure and dynamic public auditing convergence for cloud storage. The Journal of Supercomputing, 77 (5):4917–4956, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03456-y.

Wang:2021:ATD

[716] Farui Wang, Weizhe Zhang, and Zheng Wang. Automatic translation of data parallel programs for heterogeneous parallelism through OpenMP offloading. The Journal of Supercomputing, 77(5):4957–4987, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03452-2.

Zhu:2021:POT

[717] Zijie Zhu, Yongxian Wang, and Xinghua Cheng. Parallel optimization of three-dimensional wedge-shaped underwater acoustic propagation based on MPI + OpenMP hybrid programming model. The Journal of Supercomputing, 77(5):4988–5018, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03466-w.

Kommineni:2021:ACF

[718] Jenni Kommineni, Satria Mandala, and Parvathaneni Midhu Chakravarthy. Accurate computing of facial expression recognition using a hybrid feature extraction technique. The Journal of Supercomputing, 77(5):5019– 5044, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03468-8.

Hajian:2021:CCH

[719] R. Hajian and S. H. Erfani. CHESDA: continuous hybrid and energy-efficient secure data aggregation for WSN. The Journal of Supercomputing, 77 (5):5045–5075, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03455-z.

Faraji-Biregani:2021:SCB

[720] Maryam Faraji-Biregani and Reza Fotohi. Secure communication between UAVs using a method based on smart agents in unmanned aerial vehicles. The Journal of Supercomputing, 77 (5):5076–5103, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03462-0.

AlGhamdi:2021:DLM

[721] Rayed AlGhamdi, Azra Aziz, and Tarique Aziz. Deep learning model with ensemble techniques to compute the secondary structure of proteins. The Journal of Supercomputing, 77 (5):5104–5119, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03467-9.

Kiani:2021:SCS

[722] Mohsen Kiani and Amir Rajabzadeh. SDAM: a combined stack distanceanalytical modeling approach to estimate memory performance in GPUs. The Journal of Supercomputing, 77 (5):5120–5147, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03483-9.

Zhang:2021:FHS

[723] Hongjun Zhang, Heng Zhang, and Yanjun Wu. FastUDP: a highly scalable user-level UDP framework in multi-core systems for fast packet I/O. The Journal of Supercomputing, 77 (5):5148–5175, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03486-6.

Asadi:2021:TND

[724] Mohammad-Ali Asadi, Mohammad Mosleh, and Majid Haghparast. Toward novel designs of reversible ternary 6:2 compressor using efficient reversible ternary full-adders. The Journal of Supercomputing, 77(5): 5176–5197, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03485-7.

Jackins:2021:ABS

[725] V. Jackins, S. Vimal, and Mi Young Lee. AI-based smart prediction of clinical disease using random forest classifier and Naive Bayes. The Journal of Supercomputing, 77(5):5198– 5219, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03481-x.

Lin:2021:EET

[726] Yun Lin, Tundong Liu, and Yi Xie. An energy-efficient task migration scheme based on genetic algorithms for mobile applications in CloneCloud. The Journal of Supercomputing, 77 (5):5220–5236, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03470-0.

Khosravi:2021:IPC

[727] Mohammad R. Khosravi and Varun G. Menon. Intelligent and pervasive computing for cyber-physical systems. The Journal of Supercomputing, 77 (5):5237–5238, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03449-x.

Tavallali:2021:MTO

[728] Pooya Tavallali, Peyman Tavallali, and Mukesh Singhal. K-means tree: an optimal clustering tree for unsupervised learning. The Journal of Supercomputing, 77(5):5239– 5266, May 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03436-2. See cor-

rection [2870].

Bahig:2021:EPS

[729] Hazem M. Bahig and Khaled A. Fathy. An efficient parallel strategy for high-cost prefix operation. The Journal of Supercomputing, 77 (6):5267–5288, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03473-x.

Javanmardi:2021:ASC

[730] Abdol Karim Javanmardi, S. Hadi Yaghoubyan, and Hamid Parvin. An architecture for scheduling with the capability of minimum share to heterogeneous Hadoop systems. The Journal of Supercomputing, 77(6): 5289–5318, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03487-5.

Sheidaeian:2021:TGF

[731] Hamed Sheidaeian and Omid Fatemi. Toward a general framework for jointly processor-workload empirical modeling. The Journal of Supercomputing, 77(6):5319–5353, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03475-9.

Shekarriz:2021:QPO

[732] Mohsen Shekarriz, Seyed Morteza Babamir, and Meghdad Mirabi. Query processing optimization in broadcasting XML data in mobile communications. The Journal of Supercomputing, 77(6):5354–5380, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03479-5.

Lin:2021:FCF

[733] Hung-Yi Lin. Feature clustering and feature discretization assisting gene selection for molecular classification using fuzzy c-means and expectationmaximization algorithm. The Journal of Supercomputing, 77(6):5381– 5397, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03480-y.

Mahini:2021:EGA

[734] Hamidreza Mahini, Amir Masoud Rahmani, and Seyyedeh Mobarakeh Mousavirad. An evolutionary game approach to IoT task offloading in fog-cloud computing. The Journal of Supercomputing, 77(6):5398– 5425, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03484-8.

Chen:2021:CCR

[735] Genlang Chen, Jiajian Zhang, and Chaoyi Pang. CRState: checkpoint/restart of OpenCL program for in-kernel applications. The Journal of Supercomputing, 77(6):5426– 5467, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03460-2.

Mesquita:2021:MPT

[736] Leonardo A. J. Mesquita, Karcius D. R. Assis, and Raul C. Almeida Jr. Multi-period traffic on elastic optical networks planning: alleviating the capacity crunch. The Journal of Supercomputing, 77(6):5468– 5491, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03493-7.

Xia:2021:EDA

[737] Yuanyuan Xia, Shaozhong Guo, and Jinchen Xu. Error detection of arithmetic expressions. The Journal of Supercomputing, 77(6):5492– 5509, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03469-7.

Sun:2021:AAD

[738] Xiaoxin Sun, Lisa Zhang, and Bangzuo Zhang. Attribute-aware deep attentive recommendation. The Journal of Supercomputing, 77(6):5510– 5527, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03459-9.

Sharma:2021:CHR

[739] Monika Sharma and Mantosh Biswas. Classification of hyperspectral remote sensing image via rotation-invariant local binary pattern-based weighted generalized closest neighbor. The Journal of Supercomputing, 77(6): 5528–5561, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03474-w.

Braeken:2021:HEK

[740] An Braeken and Madhusanka Liyanage. Highly efficient key agreement for remote patient monitoring in MEC-enabled 5G networks. The Journal of Supercomputing, 77 (6):5562–5585, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03472-y.

Kristiani:2021:AQM

[741] Endah Kristiani, Ching-Fang Lee, and Wei-Cheng Chan. Air quality monitoring and analysis with dynamic training using deep learning. The Journal of Supercomputing, 77 (6):5586–5605, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03492-8.

Feizollahibarough:2021:SAV

[742] Sattar Feizollahibarough and Mehrdad Ashtiani. A security-aware virtual machine placement in the cloud using hesitant fuzzy decision-making processes. The Journal of Supercomputing, 77 (6):5606–5636, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03496-4.

Hjouji:2021:NSA

[743] Amal Hjouji, Belaid Bouikhalene, and Hassan Qjidaa. New set of adapted Gegenbauer–Chebyshev invariant moments for image recognition and classification. The Journal of Supercomputing, 77(6):5637– 5667, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03450-4.

Ali:2021:TSS

[744] Zainab H. Ali and Hesham A. Ali. Towards sustainable smart IoT applications architectural elements and design: opportunities, challenges, and open directions. The Journal of Supercomputing, 77(6):5668– 5725, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03477-7.

Tan:2021:NMS

[745] Zhiping Tan, Kangshun Li, and Najla Al-Nabhan. A novel mutation strategy selection mechanism for differential evolution based on local fitness landscape. The Journal of Supercomputing, 77(6):5726–5756, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03482-w.

Odajima:2021:PPC

[746] Tetsuya Odajima, Yuetsu Kodama, and Mitsuhisa Sato. Performance and power consumption analysis of ARM Scalable Vector Extension. The Journal of Supercomputing, 77(6): 5757–5778, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03495-5.

Kamrani:2021:DIM

[747] Hamed Kamrani and Saeed Rasouli Heikalabad. Design and implementation of multiplication algorithm in quantum-dot cellular automata with energy dissipation analysis. The Journal of Supercomputing, 77(6):5779– 5805, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03478-6.

Asgarnezhad:2021:AMO

[748] Razieh Asgarnezhad, S. Amirhassan Monadjemi, and Mohammadreza Soltanaghaei. An application of MOGW optimization for feature selection in text classification. The Journal of Supercomputing, 77(6): 5806–5839, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03490-w.

Mapetu:2021:DVC

[749] Jean Pepe Buanga Mapetu, Lingfu Kong, and Zhen Chen. A dynamic VM consolidation approach based on load balancing using Pearson correlation in cloud computing. The Journal of Supercomputing, 77 (6):5840–5881, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03494-6.

Mansouri:2021:CBA

[750] N. Mansouri, M. M. Javidi, and B. Mohammad Hasani Zade. A CSObased approach for secure data replication in cloud computing environment. The Journal of Supercomputing, 77(6):5882–5933, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03497-3.

Yang:2021:IDS

[751] Chao-Tung Yang, Tzu-Yang Chen, and Shyhtsun Felix Wu. The implementation of data storage and analytics platform for big data lake of electricity usage with Spark. The Journal of Supercomputing, 77(6): 5934–5959, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03505-6.

Li:2021:OIP

[752] Jingbo Li, Xingjun Zhang, and Chenglong Hu. OKCM: improving parallel task scheduling in high-performance computing systems using online learning. The Journal of Supercomputing, 77(6):5960–5983, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03506-5.

Rao:2021:MUU

[753] T. Ramalingeswara Rao, Soumya Kanti Ghosh, and Adrijit Goswami. Mining user-user communities for a weighted bipartite network using spark Graph-Frames and Flink Gelly. The Journal of Supercomputing, 77(6):5984– 6035, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03488-4.

Sharma:2021:POD

[754] Atul Sharma, Nitin Goyal, and Kalpna Guleria. Performance optimization in delay tolerant networks using backtracking algorithm for fully credits distribution to contrast selfish nodes. The Journal of Supercomputing, 77 (6):6036–6055, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03507-4.

Mohammadi:2021:ALC

[755] Maryam Mohammadi, Mahmood Fazlali, and Mehdi Hosseinzadeh. Accelerating Louvain community detection algorithm on graphic processing unit. The Journal of Supercomputing, 77(6):6056–6077, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03510-9.

Lu:2021:CTB

[756] JiaYu Lu, KongFa Hu, and TianShu Wang. A cluster-tree-based energyefficient routing protocol for wireless sensor networks with a mobile sink. The Journal of Supercomputing, 77 (6):6078–6104, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03501-w.

Jang:2021:ECT

[757] Seok-Woo Jang. Efficient covering of target areas using a location prediction-based algorithm. The Journal of Supercomputing, 77(6): 6105–6122, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03504-7.

Esmaeilyfard:2021:DCC

[758] Rasool Esmaeilyfard and Mahshid Naderi. Distributed composition of complex event services in IoT network. The Journal of Supercomputing, 77 (6):6123–6144, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03498-2.

Proficz:2021:ICR

[759] Jerzy Proficz and Krzysztof M. Ocetkiewicz. Improving Clairvoyant: reduction algorithm resilient to imbalanced process arrival patterns. The Journal of Supercomputing, 77 (6):6145–6177, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03499-1.

Kocak:2021:NDI

[760] Cem Kocak, Erol Egrioglu, and Eren Bas. A new deep intuitionistic fuzzy time series forecasting method based on long short-term memory. The Journal of Supercomputing, 77 (6):6178–6196, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03503-8.

Kholod:2021:PSO

[761] Ivan Kholod, Andrey Rukavitsyn, and Sergei Gorlatch. Parallelization of the self-organized maps algorithm for federated learning on distributed sources. The Journal of Supercomputing, 77 (6):6197–6213, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03509-2.

Gholizadeh:2021:KDI

[762] Nahid Gholizadeh, Hamid Saadatfar, and Nooshin Hanafi. K-DBSCAN: An improved DBSCAN algorithm for big data. The Journal of Supercomputing, 77(6):6214–6235, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03524-3.

Khan:2021:EMH

[763] Mohammad Ayoub Khan and Khaled Ali Abuhasel. An evolutionary multihidden Markov model for intelligent threat sensing in industrial Internet of Things. The Journal of Supercomputing, 77(6):6236– 6250, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03513-6.

Fei:2021:ADE

[764] Wenwei Fei, Hongyu Ni, and Wenxu Yan. Application of discrete eventtriggered H_{∞} control in wireless signal transmission of IOT. The Journal of Supercomputing, 77(6):6251– 6272, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03511-8.

Liu:2021:PCM

[765] Peini Liu and Jordi Guitart. Performance comparison of multi-container deployment schemes for HPC workloads: an empirical study. The Journal of Supercomputing, 77(6): 6273–6312, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03518-1.

Mohammed:2021:DND

[766] Thaha Mohammed, Aiiad Albeshri, and Rashid Mehmood. DIESEL: a novel deep learning-based tool for SpMV computations and solving sparse linear equation systems. The Journal of Supercomputing, 77(6):6313– 6355, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03489-3.

Alduailij:2021:FPE

[767] Mona A. Alduailij, Ioan Petri, and Abdulrahman S. Aldawood. Forecasting peak energy demand for smart buildings. The Journal of Supercomputing, 77(6):6356–6380, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03540-3.

Park:2021:KLM

[768] Joo Hyun Park, Jong-Woo Lee, and Soon-Bum Lim. Korean language math-to-speech rules for digital books for people with reading disabilities and their usability evaluation. The Journal of Supercomputing, 77(6): 6381–6407, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03519-0.

Shadi:2021:RTP

[769] Mahsa Shadi, Saeid Abrishami, and Behrooz Zolfaghari. Ready-time partitioning algorithm for computation offloading of workflow applications in mobile cloud computing. The Journal of Supercomputing, 77(6):6408– 6434, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03528-z.

Cheng:2021:VRC

[770] Sheng-Tzong Cheng, Chih-Wei Hsu, and Ci-Ruei Jiang. Video reasoning for conflict events through feature extraction. The Journal of Supercomputing, 77(6):6435–6455, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03514-5.

Bakshi:2021:EEC

[771] Mohana Bakshi, Chandreyee Chowdhury, and Ujjwal Maulik. Energyefficient cluster head selection algorithm for IoT using modified glowworm swarm optimization. The Journal of Supercomputing, 77(7):6457– 6475, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03536-z.

Tsuji:2021:NSS

[772] Miwako Tsuji, William T. C. Kramer, and Mitsuhisa Sato. A new sustained system performance metric for scientific performance evaluation. The Journal of Supercomputing, 77 (7):6476–6504, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03545-y.

Pasupathi:2021:TAU

[773] Subbulakshmi Pasupathi, Vimal Shanmuganathan, and Mucheol Kim. Trend analysis using agglomerative hierarchical clustering approach for time series big data. The Journal of Supercomputing, 77(7):6505– 6524, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03580-9.

Xu:2021:GOD

[774] Bin Xu and Tiefei Zhang. GPUoriented dynamic low-power data transmission method. The Journal of Supercomputing, 77(7):6525– 6539, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03515-4.

Omar:2021:SAS

[775] Mehnuma Tabassum Omar, K. M. Azharul Hasan, and Tatsuo Tsuji. A scalable array storage for efficient maintenance of future data. The Journal of Supercomputing, 77(7):6540– 6565, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03554-x.

Maghsoud:2021:PPE

[776] Zeinab Maghsoud, Hamid Noori, and Saadat Pour Mozaffari. PEPS: predictive energy-efficient parallel scheduler for multi-core processors. The Journal of Supercomputing, 77(7): 6566–6585, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03562-x.

Bhatia:2021:DCV

[777] Tarandeep Kaur Bhatia, Ramkumar Ketti Ramachandran, and Lei Pan. Data congestion in VANETs: research directions and new trends through a bibliometric analysis. The Journal of Supercomputing, 77(7): 6586–6628, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03520-7.

Pandey:2021:CAE

[778] Sarvesh Pandey and Udai Shanker. A contention aware EQS priority assignment heuristic for cohorts in DRT-DBS. The Journal of Supercomputing, 77(7):6629–6663, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03530-5.

Mirobi:2021:DDA

[779] G. Justy Mirobi and L. Arockiam. DAVmS: Distance Aware Virtual Machine Scheduling approach for reducing the response time in cloud computing. The Journal of Supercomputing, 77(7):6664–6675, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03563-w.

Tsai:2021:EIA

[780] Ming-Fong Tsai and Hung-Ju Tseng. Enhancing the identification accuracy of deep learning object detection using natural language processing. The Journal of Supercomputing, 77(7):6676–6691, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03525-2.

Shirmohammadi:2021:SCL

[781] Zahra Shirmohammadi, Ata Khorami, and Martin Eugenio Omana. ST-CAC: a low-cost crosstalk avoidance coding mechanism based on threevalued numerical system. The Journal of Supercomputing, 77(7):6692– 6713, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03527-0.

Tabrizi:2021:EPP

[782] Seyyed Parham Haghighate Pazhohe Tabrizi, Akram Reza, and Seyed Mahdi Jameii. Enhanced path planning for automated nanites drug delivery based on reinforcement learning and polymorphic improved ant colony optimization. The Journal of Supercomputing, 77(7):6714–6733, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03559-6.

Duan:2021:MTU

[783] Longzhen Duan, Shuqing Yang, and Dongbo Zhang. Multilevel thresholding using an improved cuckoo search algorithm for image segmentation. The Journal of Supercomputing, 77 (7):6734–6753, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03566-7.

Abadeh:2021:RES

[784] Maryam Nooraei Abadeh and Shohreh Ajoudanian. Reconfigurable edge as a service: enhancing edges using quality-based solutions. The Journal of Supercomputing, 77(7):6754– 6787, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03579-2.

Pandey:2021:CPV

[785] Vaibhav Pandey and Poonam Saini. Constraint programming versus heuristic approach to MapReduce scheduling problem in Hadoop YARN for energy minimization. The Journal of Supercomputing, 77(7):6788– 6816, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03516-3.

Shin:2021:ASF

[786] Mincheol Shin, Geunchul Park, and Mucheol Kim. Application-specific feature selection and clustering approach with HPC system profiling data. The Journal of Supercomputing, 77(7):6817–6831, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03533-2.

Lee:2021:EUI

[787] Yeon-Chang Lee, Jiwon Son, and Sang-Wook Kim. Exploiting uninteresting items for effective graphbased one-class collaborative filtering. The Journal of Supercomputing, 77 (7):6832–6851, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03573-8.

Pan:2021:RHB

[788] Husheng Pan, Yuzhen Li, and Dezhu Zhao. Recognizing human behaviors from surveillance videos using the SSD algorithm. The Journal of Supercomputing, 77(7):6852– 6870, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03578-3.

Park:2021:SCU

[789] Jung-Heum Park. A sufficient condition for the unpaired k-disjoint path coverability of interval graphs. The Journal of Supercomputing, 77 (7):6871–6888, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03512-7.

Kireeva:2021:PSD

[790] Anastasiya Kireeva, Karl K. Sabelfeld, and Sergey Kireev. Parallel simulation of drift-diffusion-recombination by cellular automata and global random walk algorithm. The Journal of Supercomputing, 77(7):6889– 6903, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03529-y.

Jiang:2021:DFA

[791] Yuan Jiang, Ru Qiao, and Guibao Wang. Data fusion of atmospheric ozone remote sensing lidar according to deep learning. The Journal of Supercomputing, 77(7):6904– 6919, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03537-y.

Belloch:2021:PGB

[792] Jose A. Belloch, José M. Badía, and Enrique S. Quintana-Ortí. On the performance of a GPU-based SoC in a distributed spatial audio system. The Journal of Supercomputing, 77 (7):6920–6935, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03577-4.

Kumar:2021:MCG

[793] Vivek Kumar, Dilip Kumar Sharma, and Vinay Kumar Mishra. Mille Cheval: a GPU-based in-memory high-performance computing framework for accelerated processing of bigdata streams. The Journal of Supercomputing, 77(7):6936–6960, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-020-03508-
3.

Ramakrishnan:2021:EAI

[794] Jayabrabu Ramakrishnan, Ramkumar Sivasakthivel, and Mythily R. Electrooculogram-aided intelligent sensing and high-performance communication control system for massive ALS individuals. The Journal of Supercomputing, 77(7):6961– 6978, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03517-2.

Jheng:2021:EPM

[795] Guan-Yi Jheng, Yi-Cheng Chen, and Hung-Ming Liang. Evolution pattern mining on dynamic social network. The Journal of Supercomputing, 77(7):6979–6991, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03534-1.

Wu:2021:PSA

[796] Tsu-Yang Wu, Zhiyuan Lee, and Raylin Tso. Provably secure authentication key exchange scheme using fog nodes in vehicular ad hoc networks. The Journal of Supercomputing, 77 (7):6992–7020, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03548-9.

Khan:2021:BLR

[797] Mehak Khan, Hongzhi Wang, and Sajida Karim. Bidirectional LSTM-RNN-based hybrid deep learning frameworks for univariate time series classification. The Journal of Supercomputing, 77(7):7021–7045, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-020-03560-
z.

Periola:2021:FCS

[798] A. A. Periola, O. A. Osanaiye, and A. T. Olusesi. Future cloud: spherical processors for realizing low-cost upgrade in underwater data centers. The Journal of Supercomputing, 77 (7):7046–7072, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03571-w. See cor-

rection [799].

Periola:2021:CFC

[799] A. A. Periola, O. A. Osanaiye, and A. T. Olusesi. Correction to: Future cloud: spherical processors for realizing low-cost upgrade in underwater data centers. The Journal of Supercomputing, 77(7):7073, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03866-6. See [798].

Ramanathan:2021:ACC

[800] Shalini Ramanathan and Mohan Ramasundaram. Accurate computation: COVID-19 rRT-PCR positive test dataset using stages classification through textual big data mining with machine learning. The Journal of Supercomputing, 77(7): 7074–7088, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03586-3. See re-

traction note [2039].

Safaei:2021:DMD

[801] AliAsghar Safaei and Saeede HabibiAsl. Diamond: multi-dimensional indexing technique for medical images retrieval using vertical fragmentation approach. The Journal of Supercomputing, 77(7):7089– 7148, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03522-5.

Xu:2021:EHF

[802] Wentao Xu, Xin Zhao, and Ge Nong. Enhancing HDFS with a full-text search system for massive small files. The Journal of Supercomputing, 77 (7):7149–7170, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03526-1.

Liu:2021:CVC

[803] Zhao Liu. Construction and verification of color fundus image retinal vessels segmentation algorithm under BP neural network. The Journal of Supercomputing, 77(7):7171– 7183, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03551-0.

Asghar:2021:AIR

[804] Hassan Asghar and Babar Nazir. Analysis and implementation of reactive fault tolerance techniques in Hadoop: a comparative study. The Journal of Supercomputing, 77(7): 7184–7210, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03491-9. See cor-

rection [805].

Asghar:2021:CAI

[805] Hassan Asghar and Babar Nazir. Correction to: Analysis and implementation of reactive fault tolerance techniques in Hadoop: a comparative study. The Journal of Supercomputing, 77(7):7211, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03651-5. See [804].

Sudarsan:2021:HPA

[806] M. V. Sudarsan, Ch. Sai Babu, and S. Satyanarayana. High-performance analysis of interleaved high-gain converter with active switched inductor using intelligent controller. The Journal of Supercomputing, 77(7): 7212–7235, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03546-x.

Polat:2021:CBT

[807] Özlem Polat and Cahfer Güngen. Classification of brain tumors from MR images using deep transfer learning. The Journal of Supercomputing, 77 (7):7236–7252, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03572-9.

Belhadi:2021:SSS

[808] Asma Belhadi, Youcef Djenouri, and Jerry Chun-Wei Lin. SS-ITS: secure scalable intelligent transportation systems. The Journal of Supercomputing, 77(7):7253–7269, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03582-7.

Jeon:2021:DDU

[809] MinHyuk Jeon, Odsuren Temuujin, and Dong-Hyuk Im. Distributed L-diversity using Spark-based algorithm for large resource description frameworks data. The Journal of Supercomputing, 77(7):7270– 7286, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03583-6.

Oh:2021:MDA

[810] Hyunyoung Oh, Yeongpil Cho, and Yunheung Paek. A metadata-driven approach to efficiently detect codereuse attacks on ARM multiprocessors. The Journal of Supercomputing, 77(7):7287–7314, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03542-1.

Sharma:2021:DAO

[811] Sumit Sharma and Sudip Roy. Design of all-optical parallel multipliers using semiconductor optical amplifierbased Mach–Zehnder interferometers. The Journal of Supercomputing, 77 (7):7315–7350, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03543-0.

Kim:2021:GRP

[812] Mingyu Kim and Nakhoon Baek. A 3D graphics rendering pipeline implementation based on the openCL massively parallel processing. The Journal of Supercomputing, 77(7):

7351–7367, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03581-8.

Umrao:2021:AFV

[813] Brajesh Kumar Umrao and Dharmendra Kumar Yadav. Algorithms for functionalities of virtual network: a survey. The Journal of Supercomputing, 77(7):7368–7439, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03502-9.

Cao:2021:DDO

[814] WenLiang Cao, LanLan Kang, and Zu-Wen Liu. Dual-drive oppositionbased non-inertial particle swarm optimization for deep learning in IoTs. The Journal of Supercomputing, 77 (7):7440–7454, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03531-4.

Kang:2021:BAM

[815] Shinjin Kang and Soo Kyun Kim. Behavior analysis method for indoor environment based on app usage mining. The Journal of Supercomputing, 77(7):7455–7475, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03532-3.

Nabi:2021:ORO

[816] Said Nabi and Masroor Ahmed. OG-RADL: overall performance-based resource-aware dynamic load-balancer for deadline constrained cloud tasks. The Journal of Supercomputing, 77 (7):7476–7508, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03544-z.

Zhan:2021:SAS

[817] Zhuoxin Zhan, Liulan Zhong, and Zhong Ming. Sequence-aware similarity learning for next-item recommendation. The Journal of Supercomputing, 77(7):7509–7534, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03555-w.

Zheng:2021:EEI

[818] Chunfeng Zheng, Lei Chen, and Zhonghui Gao. Efficacy evaluation of interventional therapy for primary liver cancer using magnetic resonance imaging and CT scanning under deep learning and treatment of vasovagal reflex. The Journal of Supercomputing, 77(7):7535–7548, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03539-w.

Li:2021:DSC

[819] Jinyan Li, Yaoyang Wu, and Kelvin K. L. Wong. Dynamic swarm class rebalancing for the process mining of rare events. The Journal of Supercomputing, 77(7):7549– 7583, July 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03500-x.

Zhu:2021:ACT

[820] Li Zhu and Jianbo Gao. Adoption of computerized tomography images in detection of lung nodules and analysis of neuropeptide correlative substances under deep learning algorithm. The Journal of Supercomputing, 77 (7):7584–7597, July 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03538-x.

Patidar:2021:EDI

[821] Mukesh Patidar and Namit Gupta. Efficient design and implementation of a robust coplanar crossover and multilayer hybrid full adder-subtractor using QCA technology. The Journal of Supercomputing, 77(8):7893– 7915, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03592-5.

Kumar:2021:TDI

[822] Randhir Kumar and Rakesh Tripathi. Towards design and implementation of security and privacy framework for Internet of Medical Things (IoMT) by leveraging blockchain and IPFS technology. The Journal of Supercomputing, 77(8): 7916–7955, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03570-x.

Akremi:2021:CHK

[823] Aymen Akremi and Mohsen Rouached. A comprehensive and holistic knowledge model for cloud privacy protection. The Journal of Supercomputing, 77(8):7956–7988, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-020-03594-
3.

Sukhoroslov:2021:TEE

[824] Oleg Sukhoroslov. Toward efficient execution of data-intensive workflows. The Journal of Supercomputing, 77(8): 7989–8012, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03612-4.

Basha:2021:SBV

[825] M. Suleman Basha, S. K. Mouleeswaran, and K. Rajendra Prasad. Samplingbased visual assessment computing techniques for an efficient social data clustering. The Journal of Supercomputing, 77(8):8013– 8037, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03618-6.

Punitha:2021:TCE

[826] V. Punitha and C. Mala. Traffic classification for efficient load balancing in server cluster using deep learning technique. The Journal of Supercomputing, 77(8):8038– 8062, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03613-3.

Combarro:2021:PIU

[827] Elías F. Combarro, Federico Carminati, and Ignacio F. Rúa. On protocols for increasing the uniformity of random bits generated with noisy quantum computers. The Journal of Supercomputing, 77(8):8063– 8081, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03574-7.

Abbasinezhad-Mood:2021:NCC

[828] Dariush Abbasinezhad-Mood, Arezou Ostad-Sharif, and Sayyed Majid Mazinani. Novel certificateless Chebyshev chaotic map-based key agreement protocol for advanced metering infrastructure. The Journal of Supercomputing, 77(8):8082– 8110, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03552-z.

Wang:2021:HAA

[829] Dongxia Wang, Yongmei Lei, and Guozheng Wang. HSAC-ALADMM: an asynchronous lazy ADMM algorithm based on hierarchical sparse allreduce communication. The Journal of Supercomputing, 77(8):8111– 8134, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03590-7.

Munoz-Montoro:2021:PSS

[830] A. J. Muñoz-Montoro, D. Suarez-Dou, and E. F. Combarro. Parallel source separation system for heart and lung sounds. The Journal of Supercomputing, 77(8):8135– 8150, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03616-0.

Jang:2021:MAB

[831] Kyoung ae Jang and Woo-Je Kim. A method of activity-based software maintenance cost estimation for package software. The Journal of Supercomputing, 77(8):8151– 8171, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03610-6.

Kumar:2021:RFT

[832] R. Dinesh Kumar, E. Golden Julie, and Sanghyun Seo. Recognition of food type and calorie estimation using neural network. The Journal of Supercomputing, 77(8):8172– 8193, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03622-w.

He:2021:ADL

[833] Jia He. Application of deep learning model under improved EMD in railway transportation investment benefits and national economic attribute analysis. The Journal of Supercomputing, 77 (8):8194–8208, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03609-z.

Mahindru:2021:HEA

[834] Arvind Mahindru and A. L. Sangal. HybriDroid: an empirical analysis on effective malware detection model developed using ensemble methods. The Journal of Supercomputing, 77 (8):8209–8251, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03569-4.

Devi:2021:MOH

[835] K. Lalitha Devi and S. Valli. Multiobjective heuristics algorithm for dynamic resource scheduling in the cloud computing environment. The Journal of Supercomputing, 77(8):8252– 8280, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03606-2.

Gaikwad:2021:PSL

[836] Vishesh P. Gaikwad, Jitendra V. Tembhurne, and Cheng-Chi Lee. Provably secure lightweight client authentication scheme with anonymity for TMIS using chaotic hash function. The Journal of Supercomputing, 77 (8):8281–8304, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03553-y.

Foroutan:2021:IMD

[837] Seyed Amir Hossein Foroutan, Reza Sabbaghi-Nadooshan, and Mohammad Bagher Tavakoli. Investigating multiple defects on a new fault-tolerant three-input QCA majority gate. The Journal of Supercomputing, 77(8): 8305–8325, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03567-6.

Malkovsky:2021:EPF

[838] Sergey I. Malkovsky, Aleksei A. Sorokin, and Vadim A. Kondrashev. Evaluating the performance of FFT library implementations on modern hybrid computing systems. The Journal of Supercomputing, 77(8):8326– 8354, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03591-6.

Lin:2021:ARE

[839] Wen-Yen Lin and Ching-Wen Huang. Absolute rotary encoder system based on optical sensor for angular measurement. The Journal of Supercomputing, 77(8):8355– 8373, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03585-4.

Kim:2021:GAP

[840] Jung-Yeon Kim, Suhwan Lee, and Yunyoung Nam. Gait analysis in patients with neurological disorders using ankle-worn accelerometers. The Journal of Supercomputing, 77(8): 8374–8390, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03587-2.

Kiani:2021:DVM

[841] Azaz Ahmed Kiani, Yaser Hafeez, and Sadia Ali. A dynamic variability management approach working with agile product line engineering practices for reusing features. The Journal of Supercomputing, 77 (8):8391–8432, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03627-5.

Orts:2021:OFT

[842] F. Orts, G. Ortega, and E. M. Garzón. Optimal fault-tolerant quantum comparators for image binarization. The Journal of Supercomputing, 77(8): 8433–8444, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03576-5.

Kumar:2021:SPN

[843] M. P. Pavan Kumar, B. Poornima, and C. Manjunath. Structure-preserving NPR framework for image abstraction and stylization. The Journal of Supercomputing, 77(8):8445– 8513, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03547-w.

Rundo:2021:CPM

[844] Leonardo Rundo, Andrea Tangherloni, and Marco S. Nobile. A CUDA-powered method for the feature extraction and unsupervised analysis of medical images. The Journal of Supercomputing, 77(8):8514– 8531, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03565-8.

Shariq:2021:NVS

[845] Mohd Shariq and Karan Singh. A novel vector-space-based lightweight privacy-preserving RFID authentication protocol for IoT environment. The Journal of Supercomputing, 77 (8):8532–8562, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03550-1.

Simsek:2021:DLB

[846] Mehmet Ulvi Simsek, Feyza Yildirim Okay, and Suat Ozdemir. A deep learning-based CEP rule extraction framework for IoT data. The Journal of Supercomputing, 77(8):8563– 8592, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03603-5.

Kim:2021:GBE

[847] JooHwan Kim, Shan Ullah, and Deok-Hwan Kim. GPU-based embedded edge server configuration and offloading for a neural network service. The Journal of Supercomputing, 77 (8):8593–8621, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03623-9.

Chen:2021:NBF

[848] Tai-Liang Chen, Yi-Ru Chen, and Jenq-Kuen Lee. NNBlocks: a Blockly framework for AI computing. The Journal of Supercomputing, 77(8): 8622–8652, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03631-9.

Maheshwari:2021:CBD

[849] Rashmi Maheshwari, Naveen Kumar, and Shailesh Tiwari. Consensusbased data replication protocol for distributed cloud. The Journal of Supercomputing, 77(8):8653–8673, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-021-03619-
                                        5.

Zhang:2021:AVD

[850] Jianhai Zhang, Jianhong Yu, and Xinhua Tian. Adoption value of deep learning and serological indicators in the screening of atrophic gastritis based on artificial intelligence. The Journal of Supercomputing, 77 (8):8674–8693, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03630-w.

Afanasyev:2021:VHP

[851] Ilya V. Afanasyev, Vladimir V. Voevodin, and Hiroaki Kobayashi. VGL: a high-performance graph processing framework for the NEC SX-Aurora TSUBASA vector architecture. The Journal of Supercomputing, 77(8): 8694–8715, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03564-9.

Shahzad:2021:ICV

[852] Atif Shahzad, Abdul Zubar Hameed, and Abdulrahman Basahel. Identification and control of the volatile organic compounds activity in confined environments (mosques). The Journal of Supercomputing, 77(8): 8716–8727, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03599-y.

Long:2021:URS

[853] Leijin Long, Feng He, and Hongjiang Liu. The use of remote sensing satellite using deep learning in emergency monitoring of high-level landslides disaster in Jinsha River. The Journal of Supercomputing, 77(8): 8728–8744, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03604-4. See cor-

rection [1491].

Savarapu:2021:HPB

[854] Sridhar Savarapu and Yadaiah Narri. High performance of brain emotional intelligent controller for DTC-SVM based sensorless induction motor drive. The Journal of Supercomputing, 77(8): 8745–8766, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03556-9.

Li:2021:TGC

[855] Xingquan Li and Hongxi Wu. Toward graph classification on structure property using adaptive motif based on graph convolutional network. The Journal of Supercomputing, 77(8): 8767–8786, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03628-4.

Negi:2021:CEL

[856] Sarita Negi, Man Mohan Singh Rauthan, and Neelam Panwar. CMODLB: an efficient load balancing approach in cloud computing environment. The Journal of Supercomputing, 77(8): 8787–8839, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03601-7.

Zhao:2021:SHO

[857] Ruxin Zhao, Yongli Wang, and Hao Li. A selfish herd optimization algorithm based on the simplex method for clustering analysis. The Journal of Supercomputing, 77(8):8840– 8910, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03597-0.

Hatami:2021:KFU

[858] Rashid Hatami and Hossein Bahramgiri. KDB: a fast update and high speed packet classifier in SDN. The Journal of Supercomputing, 77(8): 8911–8926, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03598-z.

Zhu:2021:RNT

[859] Xiaojun Zhu. Revisiting non-tree routing for maximum lifetime data gathering in wireless sensor networks. The Journal of Supercomputing, 77(8): 8927–8945, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03621-x.

Schembera:2021:LRD

[860] Björn Schembera. Like a rainbow in the dark: metadata annotation for HPC applications in the age of dark data. The Journal of Supercomputing, 77(8): 8946–8966, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03602-6.

Revathi:2021:MTP

[861] N. Revathi and G. Sumathi. Multistep temperature prediction for proactive thermal management on chip multiprocessors. The Journal of Supercomputing, 77(8):8967– 8994, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03611-5.

Bezerra:2021:ONP

[862] Diego de Freitas Bezerra, Guto Leoni Santos, and Patricia Takako Endo. Optimizing NFV placement for distributing micro-data centers in cellular networks. The Journal of Supercomputing, 77(8):8995–9019, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-021-03620-
                                        y.

Fang:2021:UBN

[863] Xiaodong Fang, Chan Chang, and Genggeng Liu. Using Bayesian network technology to predict the semiconductor manufacturing yield rate in IoT. The Journal of Supercomputing, 77 (8):9020–9045, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03649-z.

Chen:2021:SBB

[864] Chien-Ming Chen, Xiaoting Deng, and S. K. Hafizul Islam. A secure blockchain-based group key agreement protocol for IoT. The Journal of Supercomputing, 77(8):9046– 9068, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03561-y.

Hoffmann:2021:CAR

[865] Rolf Hoffmann, Dominique Désérable, and Franciszek Seredyński. A cellular automata rule placing a maximal number of dominoes in the square and diamond. The Journal of Supercomputing, 77(8):9069– 9087, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03549-8.

Dang:2021:MDP

[866] Chenpo Dang, Guirong Yi, and Shensong Li. MapReduce distributed parallel computing framework for diagnosis and treatment of knee joint Kashin–Beck disease. The Journal of Supercomputing, 77(8):9088– 9101, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03608-0.

Mohmmadzadeh:2021:EBC

[867] Hekmat Mohmmadzadeh and Farhad Soleimanian Gharehchopogh. An efficient binary chaotic symbiotic organisms search algorithm approaches for feature selection problems. The Journal of Supercomputing, 77(8):9102– 9144, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03626-6.

Pereira:2021:OWC

[868] Anil L. Pereira. Optical wireless communication using camera and RGB display. The Journal of Supercomputing, 77(8):9145– 9171, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03633-7.

Jang:2021:OTB

[869] Hee-Seon Jang and Jang Hyun Baek. Optimal TAL-based registration with cell-based central policy in mobile cellular networks: a semi-Markov process approach. The Journal of Supercomputing, 77(8): 9172–9189, August 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03624-8.

Paoletti:2021:DMP

[870] M. E. Paoletti, X. Tao, and A. Plaza. Deep mixed precision for hyperspectral image classification. The Journal of Supercomputing, 77(8):9190– 9201, August 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03638-2.

Bessonov:2021:PMW

[871] Oleg Bessonov and Sofiane Meradji. Parallel modeling of wildfires using efficient solvers for illconditioned linear systems. The Journal of Supercomputing, 77(9):9365– 9379, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03632-8.

Sood:2021:ETA

[872] Tanvi Sood and Kanika Sharma. Enhanced traffic-adaptive slotted MAC for IoT-based smart monitoring grid. The Journal of Supercomputing, 77(9): 9380–9410, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03588-1.

Pakdel:2021:FAP

[873] Hossein Pakdel and Reza Fotohi. A firefly algorithm for power management in wireless sensor networks (WSNs). The Journal of Supercomputing, 77(9): 9411–9432, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03639-1.

Hussain:2021:IHA

[874] Tassadaq Hussain, Saqib Amin, and Eduard Ayguadé. Implementation of a high-accuracy phase unwrapping algorithm using parallel-hybrid programming approach for displacement sensing using self-mixing interferometry. The Journal of Supercomputing, 77(9): 9433–9453, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03634-6.

Raman:2021:CWS

[875] Narayani Raman, Aisha Banu Wahab, and Sutherson Chandrasekaran. Computation of workflow scheduling using backpropagation neural network in cloud computing: a virtual machine placement approach. The Journal of Supercomputing, 77(9):9454– 9473, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03648-0.

Bai:2021:ELS

[876] Qingchun Bai, Kai Wei, and Liang He. Entity-level sentiment prediction in Danmaku video interaction. The Journal of Supercomputing, 77(9): 9474–9493, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03652-4.

Ohata:2021:NTL

[877] Elene Firmeza Ohata, João Victor Souza das Chagas, and Pedro Pedrosa Rebouças Filho. A novel transfer learning approach for the classification of histological images of colorectal cancer. The Journal of Supercomputing, 77(9):9494– 9519, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03575-6.

Mikavica:2021:BBS

[878] Branka Mikavica and Aleksandra Kostić-Ljubisavljević. Blockchainbased solutions for security, privacy, and trust management in vehicular networks: a survey. The Journal of Supercomputing, 77(9):9520– 9575, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03659-x.

Veeramakali:2021:IIT

[879] T. Veeramakali, R. Siva, and N. Krishnaraj. An intelligent Internet of Things-based secure healthcare framework using blockchain technology with an optimal deep learning model. The Journal of Supercomputing, 77(9): 9576–9596, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03637-3.

Kang:2021:HPS

[880] Ji-Hoon Kang, Jinyul Hwang, and Hoon Ryu. High-performance simulations of turbulent boundary layer flow using Intel Xeon Phi manycore processors. The Journal of Supercomputing, 77(9):9597–9614, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03642-6.

Melo:2021:DAP

[881] Carlos Melo, Jamilson Dantas, and Paulo Maciel. Distributed application provisioning over ethereum-based private and permissioned blockchain: availability modeling, capacity, and costs planning. The Journal of Supercomputing, 77(9):9615–9641, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://link.

springer.com/article/10.1007/s11227-
020-03617-z.

Liao:2021:MAO

[882] Yanping Liao and Xinyu Chen. Multiattribute overlapping radar working pattern recognition based on K-NN and SVM-BP. The Journal of Supercomputing, 77(9):9642– 9657, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03660-4.

Chuang:2021:CEL

[883] Po-Jen Chuang and Yen-Feng Tu. Constructing effective lightweight privacy protection in RFID-based systems. The Journal of Supercomputing, 77(9): 9658–9688, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03656-0.

Ramachandranpillai:2021:SDG

[884] Resmi Ramachandranpillai and Michael Arock. A solution to dynamic green vehicle routing problems with time windows using spiking neural P systems with modified rules and learning. The Journal of Supercomputing, 77(9): 9689–9720, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03635-5.

Toosi:2021:FFE

[885] Amir Bavafa Toosi and Mehdi Sedighi. FERMAT: FPGA energy reduction method by approximation theory. The Journal of Supercomputing, 77(9): 9721–9745, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03644-4.

Luccio:2021:CRG

[886] Fabrizio Luccio and Linda Pagli. Cops and robber on grids and tori: basic algorithms and their extension to a large number of cops. The Journal of Supercomputing, 77(9):9746– 9770, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03655-1.

Malyshkin:2021:PPE

[887] Victor Malyshkin, Darkhan Akhmed-Zaki, and Vladislav Perepelkin. Parallel programs execution optimization using behavior control in LuNA system. The Journal of Supercomputing, 77(9): 9771–9779, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03654-2.

Kim:2021:LOB

[888] Jongmo Kim, Junsik Kong, and Gyudong Park. Layered ontologybased multi-sourced information integration for situation awareness. The Journal of Supercomputing, 77(9): 9780–9809, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03629-3.

Jeong:2021:EDM

[889] Yoon-Su Jeong, Dong-Ryool Kim, and Seung-Soo Shin. Efficient data management techniques based on hierarchical IoT privacy using block chains in cloud environments. The Journal of Supercomputing, 77(9):9810– 9826, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03653-3.

Lee:2021:GTA

[890] Wookey Lee, Justin JongSu Song, and James J. H. Lee. Graph threshold algorithm. The Journal of Supercomputing, 77(9):9827– 9847, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03665-z.

Takahashi:2021:TTM

[891] Katsurou Takahashi and Hiroaki Ohshima. Transitivity of transformation matrices to bridge word vector spaces over 1000 years. The Journal of Supercomputing, 77(9):9848– 9878, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03584-5.

Einavipour:2021:IIB

[892] Sina Einavipour and Reza Javidan. An intelligent IoT-based positioning system for theme parks. The Journal of Supercomputing, 77(9):9879– 9904, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03669-9.

Pereira:2021:AMA

[893] Paulo Pereira, Jean Araujo, and Paulo Maciel. Analytical models for availability evaluation of edge and fog computing nodes. The Journal of Supercomputing, 77(9):9905– 9933, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03672-0.

Cano-Cano:2021:MEQ

[894] Javier Cano-Cano, Francisco J. Andújar, and José L. Sánchez. A methodology to enable QoS provision on InfiniBand hardware. The Journal of Supercomputing, 77(9):9934– 9946, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03667-x.

Lipare:2021:FRB

[895] Amruta Lipare, Damodar Reddy Edla, and Saidi Reddy Parne. Fuzzy rulebased system for energy efficiency in wireless sensor networks. The Journal of Supercomputing, 77(9):9947– 9970, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03668-w.

Barrachina:2021:PUF

[896] Sergio Barrachina, Adrián Castelló, and Jose I. Mestre. PyDTNN: a user-friendly and extensible framework for distributed deep learning. The Journal of Supercomputing, 77(9): 9971–9987, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03673-z.

Tang:2021:AMD

[897] Sai Tang, Hongrui Zhao, and Hongwei Wang. Analysis of the material distribution system of wise information technology of 120 under deep learning. The Journal of Supercomputing, 77(9): 9988–10002, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03646-2.

Kousika:2021:IPP

[898] N. Kousika and K. Premalatha. An improved privacy-preserving data mining technique using singular value decomposition with three-dimensional rotation data perturbation. The Journal of Supercomputing, 77(9):10003– 10011, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03643-5.

Gangadhar:2021:DLP

[899] A. Gangadhar and K. Babulu. Design of low-power and high-speed CNTFET-based TCAM cell for future generation networks. The Journal of Supercomputing, 77(9):10012– 10022, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03657-z.

Chien:2021:LMS

[900] Wei-Che Chien and Yueh-Min Huang. A lightweight model with spatialtemporal correlation for cellular traffic prediction in Internet of Things. The Journal of Supercomputing, 77 (9):10023–10039, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-021-03662-
2.

Quesada-Barriuso:2021:GAW

[901] Pablo Quesada-Barriuso, Dora Blanco Heras, and Francisco Argüello. GPU accelerated waterpixel algorithm for superpixel segmentation of hyperspectral images. The Journal of Supercomputing, 77(9):10040–10052, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03666-y.

Stafford:2021:PET

[902] Esteban Stafford and José Luis Bosque. Performance and energy task migration model for heterogeneous clusters. The Journal of Supercomputing, 77(9):10053–10064, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03663-1.

Chang:2021:SNE

[903] Yuan-Ming Chang, Chia-Yu Sung, and Jenq-Kuen Lee. Support NNEF execution model for NNAPI. The Journal of Supercomputing, 77(9):10065– 10096, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03625-7.

Huang:2021:RTE

[904] Jie-Fu Huang, Geng-Hua Zhang, and Sun-Yuan Hsieh. Real-time energy data compression strategy for reducing data traffic based on smart grid AMI networks. The Journal of Supercomputing, 77(9):10097– 10116, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-020-03557-8.

Hameed:2021:EHA

[905] Abdul Zubar Hameed, Balamurugan Ramasamy, and Ahmed Atef S. Bakhsh. Efficient hybrid algorithm based on genetic with weighted fuzzy rule for developing a decision support system in prediction of heart diseases. The Journal of Supercomputing, 77(9):10117–10137, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03677-9.

Rodriguez-de-Vera:2021:DDL

[906] Jesús M. Rodríguez de Vera, Josefa González-Carrillo, and Gregorio Bernabé. Deploying deep learning approaches to left ventricular non-compaction measurement. The Journal of Supercomputing, 77(9):10138–10151, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03664-0.

Benner:2021:FSG

[907] Peter Benner, Ernesto Dufrechou, and Enrique S. Quintana-Ortí. Factorized solution of generalized stable Sylvester equations using manycore GPU accelerators. The Journal of Supercomputing, 77(9):10152– 10164, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03658-y.

Singh:2021:HCR

[908] Harmanpreet Singh and Damanpreet Singh. Hierarchical clustering and routing protocol to ensure scalability and reliability in large-scale wireless sensor networks. The Journal of Supercomputing, 77(9):10165– 10183, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03671-1.

Chuang:2021:CBR

[909] Yung-Ting Chuang and Jia-Jun Tsai. cRedit-based and reputation retrieval system. The Journal of Supercomputing, 77(9):10184– 10225, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03675-x.

Herruzo:2021:EFE

[910] Jose M. Herruzo, Ivan Fernandez, and Oscar Plata. Enabling fast and energyefficient FM-index exact matching using processing-near-memory. The Journal of Supercomputing, 77(9):10226– 10251, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03661-3.

Sulaiman:2021:HLB

[911] Muhammad Sulaiman, Zahid Halim, and Dogan Aydin. A hybrid listbased task scheduling scheme for heterogeneous computing. The Journal of Supercomputing, 77(9):10252– 10288, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03685-9.

Prabha:2021:PCR

[912] P. Lakshmi Prabha, A. K. Jayanthy, and Balaji Ramraj. Prediction of cardiovascular risk by measuring carotid intima media thickness from an ultrasound image for type II diabetic mellitus subjects using machine learning and transfer learning techniques. The Journal of Supercomputing, 77(9):10289– 10306, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03676-w.

Jun-Feng:2021:POR

[913] Wang Jun-Feng, Ding Gang-Yi, and Zhang Fu-Quan. Parallel optimization of the ray-tracing algorithm based on the HPM model. The Journal of Supercomputing, 77(9):10307– 10332, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03680-0. See cor-

rection [914].

Jun-Feng:2021:CPO

[914] Wang Jun-Feng, Ding Gang-Yi, and Zhang Fu-Quan. Correction to: Parallel optimization of the ray-tracing algorithm based on the HPM model. The Journal of Supercomputing, 77 (9):10333, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03798-1. See [913].

Song:2021:IOS

[915] Yang Song, Helin Jin, and You Liu. IDCOS: optimization strategy for parallel complex expression computation on big data. The Journal of Supercomputing, 77(9):10334– 10356, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03674-y.

Tang:2021:OMM

[916] Hengliang Tang, Chunlin Li, and Youlong Luo. Optimal multilevel media stream caching in cloudedge environment. The Journal of Supercomputing, 77(9):10357–10376, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03683-x.

Mishra:2021:QIB

[917] Kaushik Mishra, Rosy Pradhan, and Santosh Kumar Majhi. Quantuminspired binary chaotic salp swarm algorithm (QBCSSA)-based dynamic task scheduling for multiprocessor cloud computing systems. The Journal of Supercomputing, 77(9):10377– 10423, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03695-7.

Solongontuya:2021:NSP

[918] Boldmaa Solongontuya, Kyung Joo Cheoi, and Mi-Hye Kim. Novel side pose classification model of stretching gestures using three-layer LSTM. The Journal of Supercomputing, 77(9):10424–10440, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-021-03684-
w.

Xing:2021:ACA

[919] Mengyan Xing, Zhonghua Ma, and Li Han. Analysis of complications after transcatheter arterial chemoembolization based on deep learning. The Journal of Supercomputing, 77(9):10441– 10462, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03687-7.

Son:2021:BCB

[920] Minjae Son, Seungwon Jung, and Eenjun Hwang. BCGAN: a CGAN-based over-sampling model using the boundary class for data balancing. The Journal of Supercomputing, 77(9):10463– 10487, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03688-6.

Gavel:2021:DID

[921] Shashank Gavel, Ajay Singh Raghuvanshi, and Sudarshan Tiwari. Distributed intrusion detection scheme using dual-axis dimensionality reduction for Internet of Things (IoT). The Journal of Supercomputing, 77(9):10488– 10511, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03697-5.

Sahni:2021:HAE

[922] Jyoti Sahni and Deo Prakash Vidyarthi. Heterogeneity-aware elastic scaling of streaming applications on cloud platforms. The Journal of Supercomputing, 77(9):10512– 10539, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03692-w.

Bakhsh:2021:HPC

[923] Ahmed A. Bakhsh. High-performance in classification of heart disease using advanced supercomputing technique with cluster-based enhanced deep genetic algorithm. The Journal of Supercomputing, 77(9):10540– 10561, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03689-5.

Baranwal:2021:FFO

[924] Gaurav Baranwal and Deo Prakash Vidyarthi. FONS: a fog orchestrator node selection model to improve application placement in fog computing. The Journal of Supercomputing, 77(9):10562–10589, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03702-x.

Wang:2021:AGS

[925] Yi Wang, Zhiping Tan, and Yeh-Cheng Chen. An adaptive gravitational search algorithm for multilevel image thresholding. The Journal of Supercomputing, 77(9):10590– 10607, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03706-7.

Oh:2021:AAL

[926] Amsuk Oh. Autofocus alignment of long distance near-infrared laser units for smoke sensing in wide closed area. The Journal of Supercomputing, 77(9):10608–10620, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03670-2.

Yoo:2021:SEI

[927] Seung Gyun Yoo and Byeongtae Ahn. A study for efficiency improvement of used car trading based on a public blockchain. The Journal of Supercomputing, 77(9):10621– 10635, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03681-z.

Banerjee:2021:ERU

[928] Sounak Banerjee, Sarbani Roy, and Sunirmal Khatua. Efficient resource utilization using multi-step-ahead workload prediction technique in cloud. The Journal of Supercomputing, 77(9):10636–10663, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03701-y.

Khanmohammadi:2021:EAT

[929] Elnaz Khanmohammadi, Behrang Barekatain, and Alfonso Ariza Quintana. An enhanced AHP-TOPSIS-based clustering algorithm for highquality live video streaming in flying ad hoc networks. The Journal of Supercomputing, 77(9):10664– 10698, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03645-3.

Duan:2021:ECD

[930] Li Duan, Yangyun Wang, and Ningming Zhou. Exploring the clinical diagnostic value of pelvic floor ultrasound images for pelvic organ prolapses through deep learning. The Journal of Supercomputing, 77(9):10699– 10720, September 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03682-y.

Acharya:2021:OCC

[931] Debasis Acharya and Dushmanta Kumar Das. Optimal coordination of over current relay using opposition learning-based gravitational search algorithm. The Journal of Supercomputing, 77(9):10721–10741, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03705-8.

Kehrer:2021:CEE

[932] Stefan Kehrer and Wolfgang Blochinger. Correction to: Equilibrium: an elasticity controller for parallel tree search in the cloud. The Journal of Supercomputing, 77(9):10742, September 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03993-0. See [474].

Ulabedin:2021:RDM

[933] Zain Ulabedin and Babar Nazir. Replication and data management-based workflow scheduling algorithm for multi-cloud data centre platform. The Journal of Supercomputing, 77(10): 10743–10772, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03541-2.

Do:2021:DNN

[934] Luu-Ngoc Do, Hyung-Jeong Yang, and In-Seop Na. Deep neural networkbased fusion model for emotion recognition using visual data. The Journal of Supercomputing, 77(10):10773– 10790, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03690-y.

Joo:2021:SVP

[935] Hae-Jong Joo and Hwa-Young Jeong. A study on VAL platform for 5G network for large-capacity data transmission. The Journal of Supercomputing, 77(10):10791–10803, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-021-03700-
z.

Kim:2021:UCI

[936] Kwanhee Kim, Jaehwan Lee, and Sangoh Park. A user-centric intelligent context-aware system for realizing Internet-of-Things environments. The Journal of Supercomputing, 77(10): 10804–10826, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03722-7.

Yi:2021:SMD

[937] Zhengming Yi and Yiping Yao. A stealing mechanism for delegation methods. The Journal of Supercomputing, 77(10):10827– 10849, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03719-2.

Ren:2021:NDB

[938] Hui Ren, Xiaochen Shen, and Xiaojun Jia. A novel dual-biological-community swarm intelligence algorithm with a commensal evolution strategy for multimodal problems. The Journal of Supercomputing, 77(10): 10850–10895, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03721-8.

Sharma:2021:DIT

[939] Shalini Sharma and Jerry Chou. Distributed and incremental travelling salesman algorithm on timeevolving graphs. The Journal of Supercomputing, 77(10):10896– 10920, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03716-5.

Wang:2021:RGP

[940] Yongzhen Wang, Xuefeng Yan, and Jun’an Zhang. Research on GPU parallel algorithm for direct numerical solution of two-dimensional compressible flows. The Journal of Supercomputing, 77(10):10921– 10941, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03704-9.

Gul:2021:MTE

[941] M. Junaid Gul, Gul Malik Urfa, and Eenjun Hwang. Mid-term electricity load prediction using CNN and Bi-LSTM. The Journal of Supercomputing, 77(10):10942–10958, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-021-03686-
8.

Ngueilbaye:2021:SSD

[942] Alladoumbaye Ngueilbaye, Hongzhi Wang, and Ibrahim A. Elgendy. SDLER: stacked dedupe learning for entity resolution in big data era. The Journal of Supercomputing, 77(10): 10959–10983, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03710-x.

Liu:2021:CDM

[943] Jung-Chun Liu, Chao-Tung Yang, and Wei-Je Jiang. Cyberattack detection model using deep learning in a network log system with data visualization. The Journal of Supercomputing, 77(10): 10984–11003, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03715-6.

Barbin:2021:NSI

[944] Javad Pashaei Barbin, Saleh Yousefi, and Behrooz Masoumi. Navigation in the social Internet-of-Things (SIoT) for discovering the influential service-providers using distributed learning automata. The Journal of Supercomputing, 77(10): 11004–11031, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03699-3.

Li:2021:DSP

[945] Pei Heng Li and Hee Yong Youn. Distributed stochastic principal component analysis using stabilized Barzilai– Borwein step-size for data compression with WSN. The Journal of Supercomputing, 77(10):11032– 11051, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03707-6.

Yazdanian:2021:EME

[946] Peyman Yazdanian and Saeed Sharifian. E2LG: a multiscale ensemble of LSTM/GAN deep learning architecture for multistep-ahead cloud workload prediction. The Journal of Supercomputing, 77(10):11052– 11082, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03723-6.

Qiu:2021:DLB

[947] Linrun Qiu, Dongbo Zhang, and Najla Al-Nabhan. Deep learning-based algorithm for vehicle detection in intelligent transportation systems. The Journal of Supercomputing, 77(10): 11083–11098, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03712-9.

Dhasarathan:2021:BIP

[948] Chandramohan Dhasarathan, Manish Kumar, and Manoj Kumar. A bio-inspired privacy-preserving framework for healthcare systems. The Journal of Supercomputing, 77(10): 11099–11134, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03720-9.

Xiao:2021:IEE

[949] Peng Xiao, Zhenyu Ni, and Zhigang Hu. Improving the energyefficiency of virtual machines by I/O compensation. The Journal of Supercomputing, 77(10):11135– 11159, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03678-8.

Kaity:2021:OAE

[950] Aishwarya Kaity and Sangeeta Singh. Optimized area efficient quantum dot cellular automata based reversible code converter circuits: design and energy performance estimation. The Journal of Supercomputing, 77(10): 11160–11186, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03693-9.

Elmeiligy:2021:EPI

[951] Manar A. Elmeiligy, Ali I. El Desouky, and Sally M. Elghamrawy. An efficient parallel indexing structure for multidimensional big data using spark. The Journal of Supercomputing, 77(10): 11187–11214, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03718-3.

Szaban:2021:SMS

[952] Miroslaw Szaban. Simple method of selecting totalistic rules for pseudorandom number generator based on nonuniform cellular automaton. The Journal of Supercomputing, 77(10):

11215–11227, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03698-4.

Singh:2021:BRT

[953] Tajinder Singh and Madhu Kumari. Burst: real-time events burst detection in social text stream. The Journal of Supercomputing, 77(10): 11228–11256, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03717-4.

SanJuan:2021:LPM

[954] Pablo San Juan, Rafael Rodríguez-Sánchez, and Enrique S. Quintana-Ortí. Low precision matrix multiplication for efficient deep learning in NVIDIA Carmel processors. The Journal of Supercomputing, 77(10): 11257–11269, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03636-4.

Zhang:2021:JPF

[955] Baoyin Zhang, Zeyao Mo, and Xiaolin Cao. JCOGIN: a programming framework for particle transport on combinatorial geometry. The Journal of Supercomputing, 77(10):11270– 11287, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03711-w.

Liu:2021:SGD

[956] Chong Liu, Wen-Ze Wu, and Wanli Xie. Study of the generalized discrete grey polynomial model based on the quantum genetic algorithm. The Journal of Supercomputing, 77(10): 11288–11309, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03713-8.

Zhou:2021:ABI

[957] Wenlong Zhou and Zhiyong Fu. Adoption of bio-image technology on rehabilitation intervention of sports injury of golf. The Journal of Supercomputing, 77(10):11310– 11327, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03732-5.

Nino:2021:HHP

[958] Alfonso Niño, Sebastián Reyes, and Ramón Carbó-Dorca. An HPC hybrid parallel approach to the experimental analysis of Fermat’s theorem extension to arbitrary dimensions on heterogeneous computer systems. The Journal of Supercomputing, 77(10): 11328–11352, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03727-2.

Son:2021:SDD

[959] Siwoon Son, Hyeonseung Im, and Yang-Sae Moon. Stochastic distributed data stream partitioning using task locality: design, implementation, and optimization. The Journal of Supercomputing, 77(10):11353– 11389, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03725-4.

Ghaffari:2021:STT

[960] Hamid Reza Ghaffari. Speeding up the testing and training time for the support vector machines with minimal effect on the performance. The Journal of Supercomputing, 77(10): 11390–11409, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03729-0.

Gulistan:2021:MVN

[961] Muhammad Gulistan, Ahmed Elmoasry, and Naveed Yaqoob. N - version of the neutrosophic cubic set: application in the negative influences of Internet. The Journal of Supercomputing, 77(10):11410– 11431, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-020-03615-1.

Jain:2021:EED

[962] Shubhra Jain, K. K. Pattanaik, and Anupam Shukla. EDVWDD: Eventdriven virtual wheel-based data dissemination for mobile sink-enabled wireless sensor networks. The Journal of Supercomputing, 77(10):11432– 11457, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.

1007/s11227-021-03714-7. See cor-

rection [963].

Jain:2021:CEE

[963] Shubhra Jain, K. K. Pattanaik, and Anupam Shukla. Correction to: EDVWDD: Event-Driven Virtual Wheel-based Data Dissemination for Mobile Sink-Enabled Wireless Sensor Networks. The Journal of Supercomputing, 77(10):11458– 11459, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03790-9. See [962].

Li:2021:CGB

[964] Pei Li, Shihao Zhou, and Jiageng Chen. A CPU–GPU-based parallel search algorithm for the best differential characteristics of block ciphers. The Journal of Supercomputing, 77(10): 11460–11480, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03703-w.

Taranto-Vera:2021:ASD

[965] Gilda Taranto-Vera, Purificación Galindo-Villardón, and Vanessa Salazar-Villalva. Algorithms and software for data mining and machine learning: a critical comparative view from a systematic review of the literature. The Journal of Supercomputing, 77(10): 11481–11513, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03708-5.

BenAlla:2021:NMP

[966] Hicham Ben Alla, Said Ben Alla, and Abdellah Touhafi. A novel multiclass priority algorithm for task scheduling in cloud computing. The Journal of Supercomputing, 77(10): 11514–11555, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03741-4.

Khosravi:2021:RAS

[967] Behnam Khosravi, Behrooz Khosravi, and Bahman Khosravi. Routing algorithms for the shuffle-exchange permutation network. The Journal of Supercomputing, 77(10):11556– 11574, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03694-8.

Li:2021:FAE

[968] Hongjian Li, Yaojun Wei, and Wenhong Tian. A frequency-aware and energy-saving strategy based on DVFS for Spark. The Journal of Supercomputing, 77(10):11575– 11596, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03740-5.

Jia:2021:NCW

[969] Zhao hong Jia, Lei Pan, and Xue jun Li. A novel cloud workflow scheduling algorithm based on stable matching game theory. The Journal of Supercomputing, 77(10): 11597–11624, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03742-3.

Condia:2021:DDR

[970] Josie E. Rodriguez Condia, Pierpaolo Narducci, and Luca Sterpone. DYRE: a DYnamic REconfigurable solution to increase GPGPU’s reliability. The Journal of Supercomputing, 77(10): 11625–11642, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03751-2.

Deng:2021:RAT

[971] Zexi Deng, Dunqian Cao, and Huimin Huang. Reliability-aware task scheduling for energy efficiency on heterogeneous multiprocessor systems. The Journal of Supercomputing, 77(10): 11643–11681, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03764-x.

Chaurasia:2021:CSE

[972] Nisha Chaurasia, Mohit Kumar, and Om Prakash Verma. Comprehensive survey on energy-aware server consolidation techniques in cloud computing. The Journal of Supercomputing, 77(10):11682– 11737, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03760-1.

Pradeepa:2021:EZC

[973] S. Pradeepa and K. R. Manjula. Epidemic zone of COVID-19 from social media using hypergraph with weighting factor (HWF). The Journal of Supercomputing, 77(10):11738– 11755, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03726-3.

Diaz:2021:NOR

[974] Javier Díaz, Pablo Ibáñez, and José M. Llabería. Near-optimal replacement policies for shared caches in multicore processors. The Journal of Supercomputing, 77(10):11756– 11785, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03736-1.

Jalali:2021:SCF

[975] Shiva Jalali and Monireh Hosseini. Social collaborative filtering using local dynamic overlapping community detection. The Journal of Supercomputing, 77(10):11786– 11806, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03734-3.

Dang:2021:EHR

[976] Tommy Dang, Ngan Nguyen, and Yong Chen. HiperView: real-time monitoring of dynamic behaviors of highperformance computing centers. The Journal of Supercomputing, 77(10): 11807–11826, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03724-5.

Zhang:2021:NST

[977] Yan Zhang, Xiaoxia Han, and Xinying Xu. A novel state transition simulated annealing algorithm for the multiple traveling salesmen problem. The Journal of Supercomputing, 77(10): 11827–11852, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03744-1.

Yang:2021:PFE

[978] Lei Yang, Li Feng, and Liwei Tian. Predicting freshmen enrollment based on machine learning. The Journal of Supercomputing, 77(10):11853– 11865, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03763-y.

Gorodnichev:2021:STD

[979] Maxim Gorodnichev and Danil Lebedev. Semantic tools for development of high-level interactive applications for supercomputers. The Journal of Supercomputing, 77(10): 11866–11880, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03731-6.

Jeong:2021:IRE

[980] SoonHyeong Jeong and Byeongtae Ahn. Implementation of real estate contract system using zero knowledge proof algorithm based blockchain. The Journal of Supercomputing, 77(10): 11881–11893, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03728-1.

Vimal:2021:MPD

[981] S. Vimal, Y. Harold Robinson, and Sanghyun Seo. A method of progression detection for glaucoma using Kmeans and the GLCM algorithm toward smart medical prediction. The Journal of Supercomputing, 77(10): 11894–11910, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03757-w.

Zakirov:2021:STR

[982] Andrey Zakirov, Anastasia Perepelkina, and Sergey Khilkov. Streaming techniques: revealing the natural concurrency of the lattice Boltzmann method. The Journal of Supercomputing, 77(10):11911– 11929, October 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03762-z.

Asgari:2021:PAP

[983] Samane Asgari, Shahram Jamali, and Mahdi Nooshyar. Performance-aware placement and chaining scheme for virtualized network functions: a particle swarm optimization approach. The Journal of Supercomputing, 77(11): 12209–12229, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03758-9.

Poongodi:2021:SHS

[984] M. Poongodi, Ashutosh Sharma, and Naveen Chilamkurti. Smart healthcare in smart cities: wireless patient monitoring system using IoT. The Journal of Supercomputing, 77(11): 12230–12255, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03765-w.

Semwal:2021:OHD

[985] Vijay Bhaskar Semwal, Anjali Gupta, and Praveen Lalwani. An optimized hybrid deep learning model using ensemble learning approach for human walking activities recognition. The Journal of Supercomputing, 77(11): 12256–12279, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03768-7.

Migallon:2021:MLP

[986] H. Migallón, A. Jimeno-Morenilla, and A. Belazi. Multi-level parallel chaotic Jaya optimization algorithms for solving constrained engineering design problems. The Journal of Supercomputing, 77(11):12280– 12319, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03737-0.

DiGennaro:2021:CAL

[987] Giovanni Di Gennaro, Amedeo Buonanno, and Francesco A. N. Palmieri. Considerations about learning Word2Vec. The Journal of Supercomputing, 77 (11):12320–12335, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-021-03743-
                                        2.

Lee:2021:EPM

[988] Yi-Hsuan Lee and Sheng-Jia Jian. Effective partitioning mechanisms for time-evolving graphs in the Flink system. The Journal of Supercomputing, 77(11):12336–12354, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03769-6.

Pradeepa:2021:IIP

[989] R. Pradeepa and M. Pushpalatha. IPR: Intelligent proactive routing model toward DDoS attack handling in SDN. The Journal of Supercomputing, 77(11):12355–12381, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03750-3.

Suplatov:2021:CDH

[990] Dmitry Suplatov, Maxim Shegay, and Vytas Svedas. Co-designing HPCsystems by computing capabilities and management flexibility to accommodate bioinformatic workflows at different complexity levels. The Journal of Supercomputing, 77(11):12382– 12398, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03691-x.

Balakrishnan:2021:ISS

[991] K. Balakrishnan, R. Dhanalakshmi, and Utkarsh Mahadeo Khaire. Improved Salp Swarm Algorithm based on the Levy flight for feature selection. The Journal of Supercomputing, 77(11):12399–12419, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03773-w.

Sedaghat:2021:RDR

[992] Shahrzad Sedaghat and Amir Hossein Jahangir. R2T-DSDN: reliable real-time distributed controller-based SDN. The Journal of Supercomputing, 77(11): 12420–12457, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03780-x.

Kulkarni:2021:GAR

[993] Ashwin Kumar Kulkarni and B. Annappa. GPU-aware resource management in heterogeneous cloud data centers. The Journal of Supercomputing, 77(11):12458–12485, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03779-4.

Li:2021:PEO

[994] Wei Li and Shunfu Jin. Performance evaluation and optimization of a task offloading strategy on the mobile edge computing with edge heterogeneity. The Journal of Supercomputing, 77(11):12486– 12507, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03781-w.

Seiz:2021:LPI

[995] Marco Seiz, Philipp Offenhäuser, and Michael Resch. Lustre I/O performance investigations on hazel hen: experiments and heuristics. The Journal of Supercomputing, 77(11):12508– 12536, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03730-7.

Latif:2021:NCM

[996] Rabia Latif, Syeda Hadia Afzaal, and Seemab Latif. A novel cloud management framework for trust establishment and evaluation in a federated cloud environment. The Journal of Supercomputing, 77(11):12537– 12560, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03775-8.

Lee:2021:EMF

[997] Junkyu Lee and Seog Park. Efficient methods for finding an optimal network location for travel planning. The Journal of Supercomputing, 77(11):

12561–12580, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03776-7. See cor-

rection [998].

Lee:2021:CEM

[998] Junkyu Lee and Seog Park. Correction to: Efficient methods for finding an optimal network location for travel planning. The Journal of Supercomputing, 77(11):12581, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03837-x. See [997].

Arulprakash:2021:PCC

[999] M. Arulprakash and R. Jebakumar. People-centric collective intelligence: decentralized and enhanced privacy mobile crowd sensing based on blockchain. The Journal of Supercomputing, 77(11):12582– 12608, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03756-x.

Kaur:2021:EAR

[1000] Mandeep Kaur, Amit Gupta, and Balwinder Singh Sohi. An enhanced architecture for route discovery and load balancing in WSN. The Journal of Supercomputing, 77(11):12609– 12629, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03777-6.

Gong:2021:ELS

[1001] Changqing Gong, Xinyao Wang, and Han Qi. Enhanced long shortterm memory with fireworks algorithm and mutation operator. The Journal of Supercomputing, 77(11):12630– 12646, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03679-7.

Chen:2021:PEC

[1002] Weiduo Chen, Xiaoshe Dong, and Xingjun Zhang. Performance evaluation of convolutional neural network on tianhe-3 prototype. The Journal of Supercomputing, 77(11): 12647–12665, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03759-8.

Hannah:2021:TCA

[1003] John Hannah, Robert Mills, and Douglas Hodson. Traffic collision avoidance system: false injection viability. The Journal of Supercomputing, 77(11): 12666–12689, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03766-9.

Liu:2021:SAI

[1004] Tao Liu, Yi Yuan, and Zhongyang Yu. The service architecture of Internet of Things terminal connection based on blockchain technology. The Journal of Supercomputing, 77(11): 12690–12710, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03774-9.

Santos:2021:AAE

[1005] Guto Leoni Santos, Theo Lynn, and Patricia Takako Endo. Availabilityaware and energy-aware dynamic SFC placement using reinforcement learning. The Journal of Supercomputing, 77(11):12711–12740, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03784-7.

Fu:2021:HEP

[1006] Hao Fu, Shanjiang Tang, and Jizhou Sun. HGP4CNN: an efficient parallelization framework for training convolutional neural networks on modern GPUs. The Journal of Supercomputing, 77(11):12741–12770, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03746-z.

Zhang:2021:AIC

[1007] Juan Zhang, Changsheng Wan, and Taochen Lu. Auditing images collected by sensors in ambient intelligence systems with privacy and high efficiency. The Journal of Supercomputing, 77(11):12771–12789, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03738-z.

Alqahtani:2021:PCI

[1008] Abdulrahman Saad Alqahtani. Performance computation and implementation of distributed controllers for reliable software-defined networks. The Journal of Supercomputing, 77(11): 12790–12800, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03785-6.

Kumar:2021:ISB

[1009] K. E. Naresh Kumar and V. Uma. Intelligent sentinet-based lexicon for context-aware sentiment analysis: optimized neural network for sentiment classification on social media. The Journal of Supercomputing, 77(11): 12801–12825, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03709-4.

Crespo:2021:UDF

[1010] Juan-José Crespo, José L. Sánchez, and José Duato. UPR: deadlockfree dynamic network reconfiguration by exploiting channel dependency graph compatibility. The Journal of Supercomputing, 77(11):12826– 12856, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03791-8.

Park:2021:SWB

[1011] Sungwoo Park, Seungmin Jung, and Eenjun Hwang. Sliding window-based LightGBM model for electric load forecasting using anomaly repair. The Journal of Supercomputing, 77(11): 12857–12878, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03787-4.

Kanemitsu:2021:CTC

[1012] Hidehiro Kanemitsu, Kenji Kanai, and Hidenori Nakazato. A containerized task clustering for scheduling workflows to utilize processors and containers on clouds. The Journal of Supercomputing, 77(11):12879– 12923, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03789-2.

Kim:2021:HTB

[1013] Jeong-Geun Kim, Yoon-Su Jo, and Shin-Dug Kim. History table-based linear analysis method for DRAM-PCM hybrid memory system. The Journal of Supercomputing, 77(11): 12924–12952, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03786-5.

Mokhayeri:2021:ZSS

[1014] Sara Mokhayeri and Mohammad Taghi Kheirabadi. Zone selection strategy in geocast routing algorithms in VANET: a review. The Journal of Supercomputing, 77(11):12953– 12986, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03783-8.

Yuan:2021:HEH

[1015] Zhu Yuan, Xindong You, and Ping Xie. HS6: An efficient H-code RAID-6 scaling by optimizing data migrating and parity updating. The Journal of Supercomputing, 77(11):12987– 13017, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03739-y.

Kumar:2021:EQO

[1016] Deepak Kumar and Vijay Kumar Jha. An efficient query optimization technique in big data using σ-ANFIS load balancer and CaM-BW optimizer. The Journal of Supercomputing, 77(11): 13018–13045, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03793-6.

Aguayo-Canela:2021:EMA

[1017] Francisco José Aguayo-Canela, Héctor Alaiz-Moretón, and Isaías García-Rodríguez. Enriched multi-agent middleware for building rule-based distributed security solutions for IoT environments. The Journal of Supercomputing, 77(11):13046–13068, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03797-2.

Wei:2021:ACS

[1018] Qianzhou Wei, Dongru Huang, and Yu Zhang. Artificial chicken swarm algorithm for multi-objective optimization with deep learning. The Journal of Supercomputing, 77(11):13069– 13089, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03770-z.

Zhang:2021:FTR

[1019] Yujie Zhang, Weibei Fan, and Ruchuan Wang. Fault-tolerant routing algorithm based on disjoint paths in 3-ary n-cube networks with structure faults. The Journal of Supercomputing, 77(11): 13090–13114, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03799-0.

Behnamian:2021:SSA

[1020] J. Behnamian, S. Memar Dezfooli, and H. Asgari. A scatter search algorithm with a novel solution representation for flexible open shop scheduling: a multi-objective optimization. The Journal of Supercomputing, 77(11): 13115–13138, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03794-5.

Li:2021:PHC

[1021] Huifang Li, Danjing Wang, and Orlando Bonilla Pineda. PSO+LOA: hybrid constrained optimization for scheduling scientific workflows in the cloud. The Journal of Supercomputing, 77(11):13139–13165, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03755-y.

Tayfour:2021:CDM

[1022] Omer Elsier Tayfour and Muhammad Nadzir Marsono. Collaborative detection and mitigation of DDoS in software-defined networks. The Journal of Supercomputing, 77(11):13166– 13190, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03782-9.

Tu:2021:DAC

[1023] Chia-Heng Tu, QiHui Sun, and Mu-Hsuan Cheng. On designing the adaptive computation framework of distributed deep learning models for Internet-of-Things applications. The Journal of Supercomputing, 77(11): 13191–13223, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03795-4.

Jeong:2021:MCB

[1024] Hwa-Young Jeong. Multi criteria based personalized recommendation service using analytical hierarchy process for Airbnb. The Journal of Supercomputing, 77(11):13224– 13242, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03812-6.

Tamilarasi:2021:ABR

[1025] R. Tamilarasi and S. Prabu. Automated building and road classifications from hyperspectral imagery through a fully convolutional network and support vector machine. The Journal of Supercomputing, 77(11): 13243–13261, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03954-7.

Cha:2021:FSG

[1026] Myung-Hoon Cha, Sang-Min Lee, and Kang-Ho Kim. Fast and secure global-heap for memory-centric computing. The Journal of Supercomputing, 77(11):13262–13291, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03806-4.

Biswas:2021:LTL

[1027] Amit Biswas, Anil Kumar Tripathi, and Samir Aknine. Lea-TN: leader election algorithm considering node and link failures in a torus network. The Journal of Supercomputing, 77(11): 13292–13329, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03803-7.

Salami:2021:EEC

[1028] Hamza Onoruoiza Salami, Abubakar Bala, and Idris Ismail. An energyefficient cuckoo search algorithm for virtual machine placement in cloud computing data centers. The Journal of Supercomputing, 77(11):13330– 13357, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03807-3.

Pasini:2021:SBT

[1029] Massimiliano Lupo Pasini, Vittorio Gabbi, and Nouamane Laanait. Scalable balanced training of conditional generative adversarial neural networks on image data. The Journal of Supercomputing, 77(11):13358– 13384, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03808-2.

Cheng:2021:SES

[1030] Huiwen Cheng, Bo Liu, and Ching-Hsien Hsu. A survey of energy-saving technologies in cloud data centers. The Journal of Supercomputing, 77(11): 13385–13420, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03805-5.

Zaghari:2021:IOD

[1031] Nayereh Zaghari, Mahmood Fathy, and Mohammad Shahverdy. The improvement in obstacle detection in autonomous vehicles using YOLO non-maximum suppression fuzzy algorithm. The Journal of Supercomputing, 77(11):13421–13446, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03813-5.

Seyedkolaei:2021:MCD

[1032] Ali Abdi Seyedkolaei and Seyed Amin Hosseini Seno. Multi-criteria decision-making for controller placement in software-defined wide-area networks. The Journal of Supercomputing, 77(11):13447–13473, November 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03815-3.

Roy:2021:DTR

[1033] Rupsa Roy, Swarup Sarkar, and Sourav Dhar. Design and testing of a reversible ALU by quantum cells automata electro-spin technology. The Journal of Supercomputing, 77(12): 13601–13628, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03767-8.

Shi:2021:ODH

[1034] Qinqin Shi, Chengji Wu, and Jianping Zhang. Optimization for DVhop type of localization scheme in wireless sensor networks. The Journal of Supercomputing, 77(12):13629– 13652, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03818-0.

Chen:2021:EMA

[1035] Yulei Chen and Jianhua Chen. An efficient mutual authentication and key agreement scheme without password for wireless sensor networks. The Journal of Supercomputing, 77(12): 13653–13675, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03820-6.

Yoon:2021:BCS

[1036] Daegun Yoon, Zhetao Li, and Sangyoon Oh. Balanced content space partitioning for pub/sub: a study on impact of varying partitioning granularity. The Journal of Supercomputing, 77(12): 13676–13702, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03821-5.

Bang:2021:NDS

[1037] A. O. Bang and Udai Pratap Rao. A novel decentralized security architecture against sybil attack in RPLbased IoT networks: a focus on smart home use case. The Journal of Supercomputing, 77(12):13703– 13738, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03816-2.

Hu:2021:DSM

[1038] Yichang Hu and Lu Lu. Design of a simulation model for high performance LINPACK in hybrid CPU–GPU systems. The Journal of Supercomputing, 77(12):13739– 13756, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03829-x.

Sharma:2021:AAI

[1039] Saurabh Sharma and Vinod Kumar Verma. AIEMLA: artificial intelligence enabled machine learning approach for routing attacks on Internet of Things. The Journal of Supercomputing, 77(12):13757– 13787, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03833-1.

Lee:2021:DCQ

[1040] Miyoung Lee and Byungjoo Park.

Design of CMOS QVCO with high-Q symmetric differential inductor for wireless LAN. The Journal of Supercomputing, 77(12):13788– 13805, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03796-3.

Kalantary:2021:RDI

[1041] Samira Kalantary, Javad Akbari Torkestani, and Ali Shahidinejad. Resource discovery in the Internet of Things integrated with fog computing using Markov learning model. The Journal of Supercomputing, 77(12): 13806–13827, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03824-2.

Kheirandish:2021:EDR

[1042] Davar Kheirandish, Majid Haghparast, and Mehdi Hosseinzadeh. Efficient designs of reversible sequential circuits. The Journal of Supercomputing, 77(12):13828–13862, December 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03735-2.

Fu:2021:ICC

[1043] Deqian Fu, Shunbo Hu, and Jianlong Qiu. An intelligent cloud computing of trunk logistics alliance based on blockchain and big data. The Journal of Supercomputing, 77(12): 13863–13878, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03800-w.

Cao:2021:AIR

[1044] Weiran Cao. Applying image registration algorithm combined with CNN model to video image stitching. The Journal of Supercomputing, 77(12): 13879–13896, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03840-2.

Qi:2021:SAK

[1045] Mingping Qi and Jianhua Chen. Secure authenticated key exchange for WSNs in IoT applications. The Journal of Supercomputing, 77(12):13897– 13910, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03836-y.

Priyadarshini:2021:NLC

[1046] Ishaani Priyadarshini and Chase Cotton. A novel LSTM-CNN-grid search-based deep neural network for sentiment analysis. The Journal of Supercomputing, 77(12):13911– 13932, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03838-w.

Li:2021:COS

[1047] Chunlin Li, Qianqian Cai, and Youlong Luo. Computation offloading and service allocation in mobile edge computing. The Journal of Supercomputing, 77(12):13933– 13962, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03749-w.

Liu:2021:TAO

[1048] Taotao Liu and Guijiang Duan.

Task allocation optimization model in mechanical product development based on Bayesian network and ant colony algorithm. The Journal of Supercomputing, 77(12):13963– 13991, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03831-3.

Olivera:2021:TPF

[1049] Vicente Matellán Olivera and José Luis González-Sánchez. Thinking in parallel: foreword. The Journal of Supercomputing, 77(12):13992–13994, December 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03848-8.

Rodriguez-Canal:2021:EHP

[1050] Gabriel Rodriguez-Canal, Yuri Torres, and Arturo Gonzalez-Escribano.

Efficient heterogeneous programming with FPGAs using the Controller model. The Journal of Supercomputing, 77(12):13995–14010, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-021-03792-
7.

Liang:2021:MDD

[1051] Bin Liang, Xiaoshe Dong, and Longxiang Wang. A multi-dimensional double descending maximum padding priority algorithm for cloud data centers. The Journal of Supercomputing, 77(12): 14011–14038, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03842-0.

Alqahtani:2021:IPI

[1052] Abdulrahman Saad Alqahtani, Khaled Ali Abuhasel, and Mohammed Alquraish. On implementing a powerful intrusion prevention system focused on big data. The Journal of Supercomputing, 77(12):14039–14052, December 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03856-8.

Touqeer:2021:SHS

[1053] Haseeb Touqeer, Shakir Zaman, and Muhammad Bilal. Smart home security: challenges, issues and solutions at different IoT layers. The Journal of Supercomputing, 77(12): 14053–14089, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03825-1.

Liu:2021:IHH

[1054] Chang Liu. An improved Harris hawks optimizer for job-shop scheduling problem. The Journal of Supercomputing, 77(12):14090– 14129, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03834-0.

Arslan:2021:ESR

[1055] Sanem Arslan and Osman Unsal. Efficient selective replication of critical code regions for SDC mitigation leveraging redundant multithreading. The Journal of Supercomputing, 77(12): 14130–14160, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03804-6.

Stojanovic:2021:OPA

[1056] Igor Stojanovic, Milica Jovanovic, and Goran Lj. Djordjevic. Optimal port allocation scheme for deflectionrouted networks-on-chip. The Journal of Supercomputing, 77(12):14161– 14179, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03850-0.

Mahajan:2021:SOR

[1057] Pratibha Mahajan and Pankaj Deep Kaur. Smart object recommendation (SORec) architecture using representation learning in Smart objects-Based Social Network (SBSN). The Journal of Supercomputing, 77(12): 14180–14206, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03828-y.

Rezaeiye:2021:APS

[1058] Payam Porkar Rezaeiye, Arash Sharifi, and Mehdi Dehghan. Access point selection in the network of Internet of Things (IoT) considering the strategic behavior of the things and users. The Journal of Supercomputing, 77(12): 14207–14229, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03788-3.

Xia:2021:HAR

[1059] Limin Xia and Wentao Ma. Human action recognition using highorder feature of optical flows. The Journal of Supercomputing, 77(12): 14230–14251, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03827-z.

Cheng:2021:ACO

[1060] Sheng-Tzong Cheng, Chih-Wei Hsu, and Sz-Yu Chen. Across-camera object tracking using a conditional random field model. The Journal of Supercomputing, 77(12):14252– 14279, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03862-w.

Dang:2021:DKA

[1061] Depeng Dang, Chuangxia Chen, and Xingjian Wang. Deep knowledgeaware framework for web service recommendation. The Journal of Supercomputing, 77(12):14280–14304, December 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03832-2.

Parihar:2021:TBA

[1062] Ashish Singh Parihar and Swarnendu Kumar Chakraborty. Tokenbased approach in distributed mutual exclusion algorithms: a review and direction to future research. The Journal of Supercomputing, 77(12): 14305–14355, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03802-8.

Westby:2021:FAM

[1063] Isaac Westby, Xiaokun Yang, and Hailu Xu. FPGA acceleration on a multi-layer perceptron neural network for digit recognition. The Journal of Supercomputing, 77(12):14356– 14373, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03849-7.

Tu:2021:TEI

[1064] Nguyen Anh Tu, Thien Huynh-The, and Young-Koo Lee. Toward efficient and intelligent video analytics with visual privacy protection for large-scale surveillance. The Journal of Supercomputing, 77(12):14374– 14404, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03865-7.

Combarro:2021:RTS

[1065] Elías F. Combarro, Sofia Vallecorsa, and Alberto Di Meglio. A report on teaching a series of online lectures on quantum computing from CERN. The Journal of Supercomputing, 77(12): 14405–14435, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03847-9. See cor-

rection [1066].

Combarro:2021:CRT

[1066] Elías F. Combarro, Sofia Vallecorsa, and Alberto Di Meglio. Correction to: A report on teaching a series of online lectures on quantum computing from CERN. The Journal of Supercomputing, 77(12):14436– 14437, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03923-0.  See

[1065].

BiTalebi:2021:LLA

[1067] Hossein BiTalebi and Farshad Safaei. LARA: Locality-aware resource allocation to improve GPU memoryaccess time. The Journal of Supercomputing, 77(12):14438–14460, December 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03854-w.

Hameed:2021:FVB

[1068] Khizar Hameed, Saurabh Garg, and Byeong Kang. A formally verified blockchain-based decentralised authentication scheme for the Internet of Things. The Journal of Supercomputing, 77(12):14461–14501, December 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03841-1.

Tao:2021:CDS

[1069] Xiaohan Tao, Jianmin Pang, and Yu Zhu. Compiler-directed scratchpad memory data transfer optimization for multithreaded applications on a heterogeneous many-core architecture. The Journal of Supercomputing, 77(12): 14502–14524, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03853-x.

Saeedizade:2021:DDD

[1070] Ehsan Saeedizade and Mehrdad Ashtiani. DDBWS: a dynamic deadline and budget-aware workflow scheduling algorithm in workflow-as-a-service environments. The Journal of Supercomputing, 77(12):14525– 14564, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03858-6.

Khosravi:2021:FRC

[1071] Mohammad R. Khosravi and Sadegh Samadi. Frame rate computing and aggregation measurement toward QoS/ QoE in video-SAR systems for UAVborne real-time remote sensing. The Journal of Supercomputing, 77(12): 14565–14582, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03869-3.

Lee:2021:TGM

[1072] Donggeon Lee, Kyung Seo Ki, and Gahgene Gweon. TM-generation model: a template-based method for automatically solving mathematical word problems. The Journal of Supercomputing, 77(12):14583– 14599, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03855-9.

Doshanlou:2021:EBQ

[1073] Abdollah Norouzi Doshanlou, Majid Haghparast, and Midia Reshadi. Efficient binary to quaternary and vice versa converters: embedding in quaternary arithmetic circuits. The Journal of Supercomputing, 77(12):14600– 14616, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03696-6.

Kotb:2021:RUA

[1074] Amer Kotb, Kyriakos E. Zoiros, and Wei Li. Realization of ultrafast alloptical NAND and XNOR logic functions using carrier reservoir semiconductor optical amplifiers. The Journal of Supercomputing, 77(12):14617– 14629, December 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03876-4.

Saravanan:2021:IAV

[1075] M. Saravanan and S. Manoj Kumar. Improved authentication in VANETs using a connected dominating set-based privacy preservation protocol. The Journal of Supercomputing, 77(12):14630–14651, December 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03911-4.

Liu:2022:DSE

[1076] Chia-Yin Liu, Hasitha Muthumala Waidyasooriya, and Masanori Hariyama. Design space exploration for an FPGAbased quantum annealing simulator with interaction-coefficient-generators. The Journal of Supercomputing, 78 (1):1–17, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03859-5.

Sefati:2022:LBC

[1077] SeyedSalar Sefati, Maryamsadat Mousavinasab, and Roya Zareh Farkhady. Load balancing in cloud computing environment using the grey wolf optimization algorithm based on the reliability: performance evaluation. The Journal of Supercomputing, 78 (1):18–42, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03810-8.

Lopez-Albelda:2022:FES

[1078] Bernabé López-Albelda, Francisco M. Castro, and Nicolás Guil. FlexSched: Efficient scheduling techniques for concurrent kernel execution on GPUs. The Journal of Supercomputing, 78 (1):43–71, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03819-z.

Esmaeilzade:2022:GCB

[1079] Saeid Esmaeilzade, Nasrollah Pakniat, and Ziba Eslami. A generic construction to build simple oblivious transfer protocols from homomorphic encryption schemes. The Journal of Supercomputing, 78(1):72– 92, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03826-0.

Memari:2022:LAT

[1080] Pedram Memari, Seyedeh Samira Mohammadi, and Reza Tavakkoli-Moghaddam. A latency-aware task scheduling algorithm for allocating virtual machines in a cost-effective and time-sensitive fog-cloud architecture. The Journal of Supercomputing, 78 (1):93–122, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03868-4.

Liu:2022:GRB

[1081] Ning Liu and Pai Liu. Goaling recognition based on intelligent analysis of real-time basketball image of Internet of Things. The Journal of Supercomputing, 78(1):123– 143, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03877-3.

Zhi-Bin:2022:NPH

[1082] Huang Zhi-Bin, Fu Guang-Tao, and Dai Zhi-Tao. Novel parallel hybrid genetic algorithms on the GPU for the generalized assignment problem. The Journal of Supercomputing, 78 (1):144–167, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03882-6.

Ozonbolagh:2022:EXD

[1083] Samaneh Fathi Ozonbolagh and Meghdad Mirabi. Efficient XML data placement schemes over multiple mobile wireless broadcast channels. The Journal of Supercomputing, 78(1): 168–199, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03884-4.

Yesil:2022:SHS

[1084] Serif Yesil and Ozcan Ozturk. Scheduling for heterogeneous systems in accelerator-rich environments. The Journal of Supercomputing, 78(1): 200–221, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03883-5.

Zhu:2022:OIW

[1085] Ting Zhu, Wen Qu, and Wenliang Cao. An optimized image watermarking algorithm based on SVD and IWT. The Journal of Supercomputing, 78 (1):222–237, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03886-2.

Reshadi:2022:MVN

[1086] Midia Reshadi and Seyedeh Yasaman Hosseini Mirmahaleh. Mapping and virtual neuron assignment algorithms for MAERI accelerator. The Journal of Supercomputing, 78(1):238– 257, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03893-3.

Jani:2022:HST

[1087] Kunal Jani, Ankit Kumar, and Ronak Nahata. Hpcfolder: a simple tool used to parallelize algorithms using the message passing interface (MPI). The Journal of Supercomputing, 78 (1):258–278, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03896-0.

Raji:2022:UUA

[1088] Mohsen Raji and Mohaddaseh Nikseresht. UMOTS: an uncertainty-aware multiobjective genetic algorithm-based static task scheduling for heterogeneous embedded systems. The Journal of Supercomputing, 78(1):279– 314, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03887-1.

Shahidi:2022:RMC

[1089] Seyed Mansour Shahidi and Shahram Etemadi Borujeni. Reversible multiplier with a column-wise structure and a reduced number of ancilla inputs and garbage outputs. The Journal of Supercomputing, 78(1):315– 342, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03870-w.

Jeong:2022:SAP

[1090] SoonHyeong Jeong and Byeongtae Ahn. A study of application platform for smart contract visualization based blockchain. The Journal of Supercomputing, 78(1):343– 360, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03879-1.

Prasanalakshmi:2022:IAC

[1091] B. Prasanalakshmi, K. Murugan, and Yu-Chen Hu. Improved authentication and computation of medical data transmission in the secure IoT using hyperelliptic curve cryptography. The Journal of Supercomputing, 78 (1):361–378, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03861-x.

Luo:2022:UDB

[1092] Ji Luo, Chuhao Zhao, and Guangqin Li. Using deep belief network to construct the agricultural information system based on Internet of Things. The Journal of Supercomputing, 78 (1):379–405, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03898-y.

Melo:2022:OEI

[1093] F. Melo, E. Andrade, and G. Callou. Optimization of electrical infrastructures at data centers through a DoE-based approach. The Journal of Supercomputing, 78(1):406– 439, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03874-6.

Li:2022:RST

[1094] Yingying Li, Huihui Sun, and Jianmin Pang. Revisiting split tiling for stencil computations in polyhedral compilation. The Journal of Supercomputing, 78(1):440–470, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03835-z.

Khan:2022:CEP

[1095] Minhaj Ahmad Khan. A cost-effective power-aware approach for scheduling cloudlets in cloud computing environments. The Journal of Supercomputing, 78(1):471–496, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03894-2.

Bouchaala:2022:ESE

[1096] Mariem Bouchaala, Cherif Ghazel, and Leila Azouz Saidane. Enhancing security and efficiency in cloud computing authentication and key agreement scheme based on smart card. The Journal of Supercomputing, 78 (1):497–522, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03857-7.

Khodadadi:2022:FPE

[1097] Elham Khodadadi, Behrang Barekatain, and Zahra Mogharrabi-Rad. FT-PDC: an enhanced hybrid congestionaware fault-tolerant routing technique based on path diversity for 3D NoC. The Journal of Supercomputing, 78 (1):523–558, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03906-1.

Guo:2022:MVF

[1098] Yi Guo. The microscopic visual forms in architectural art design following deep learning. The Journal of Supercomputing, 78(1):559– 577, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03888-0.

Hou:2022:ATD

[1099] Shengchao Hou. Auxiliary tumour diagnosis image with deep learning technology. The Journal of Supercomputing, 78(1):578– 595, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03881-7.

Nasrazadani:2022:SPS

[1100] Mina Nasrazadani, Afsaneh Fatemi, and Mohammadali Nematbakhsh. Sign prediction in sparse social networks using clustering and collaborative filtering. The Journal of Supercomputing, 78(1):596–615, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03902-5.

Li:2022:IDP

[1101] Jianxin Li, Yiting Chen, and Jinyu Huang. An improved DQN path planning algorithm. The Journal of Supercomputing, 78(1):616– 639, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03878-2.

Ean:2022:DPB

[1102] Sokchomrern Ean, Manas Bazarbaev, and Kwan-Hee Yoo. Dynamic programming-based computation of an optimal tap working pattern in nonferrous arc furnace. The Journal of Supercomputing, 78(1):640– 666, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03880-8.

Wadhwa:2022:TTR

[1103] Heena Wadhwa and Rajni Aron.

TRAM: Technique for resource allocation and management in fog computing environment. The Journal of Supercomputing, 78(1):667– 690, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03885-3.

Ou:2022:SSR

[1104] Zhilong Ou, YanMin Luo, and Geng Chen. SRFNet: selective receptive field network for human pose estimation. The Journal of Supercomputing, 78(1):691–711, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03889-z.

Chen:2022:FMI

[1105] Xianyao Chen, Kyung Tae Kim, and Hee Yong Youn. Feature matching and instance reweighting with transfer learning for human activity recognition using smartphone. The Journal of Supercomputing, 78(1): 712–739, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03844-y.

Abualigah:2022:AHM

[1106] Laith Abualigah and Muhammad Alkhrabsheh. Amended hybrid multiverse optimizer with genetic algorithm for solving task scheduling problem in cloud computing. The Journal of Supercomputing, 78(1):740– 765, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03915-0.

Doghri:2022:CPS

[1107] Wael Doghri, Ahlem Saddoud, and Lamia Chaari Fourati. Cyber-physical systems for structural health monitoring: sensing technologies and intelligent computing. The Journal of Supercomputing, 78(1):766– 809, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03875-5.

Chan:2022:TWP

[1108] Yu-Wei Chan, Tsan-Ching Kang, and Yin-Te Tsai. Tool wear prediction using convolutional bidirectional LSTM networks. The Journal of Supercomputing, 78(1):810– 832, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03903-4.

Kang:2022:SIM

[1109] Ji-Hoon Kang. Scalable implementation of multigrid methods using partial semi-aggregation of coarse grids. The Journal of Supercomputing, 78 (1):833–851, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03912-3.

Karthik:2022:FEI

[1110] P. C. Karthik, J. Sasikumar, and P. Kalyanasundaram. Field equations for incompressible non-viscous fluids using artificial intelligence. The Journal of Supercomputing, 78(1): 852–867, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03917-y.

Safaiezadeh:2022:NDS

[1111] Behrouz Safaiezadeh, Ebrahim Mahdipour, and Mehdi Hosseinzadeh. Novel design and simulation of reversible ALU in quantum dot cellular automata. The Journal of Supercomputing, 78 (1):868–882, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03860-y. See cor-

rection [1112].

Safaiezadeh:2022:CND

[1112] Behrouz Safaiezadeh, Ebrahim Mahdipour, and Mehdi Hosseinzadeh. Correction to: Novel design and simulation of reversible ALU in quantum dot cellular automata. The Journal of Supercomputing, 78(1): 883, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03933-y.  See

Keshavarz-Kohjerdi:2022:ELA

[1113] Fatemeh Keshavarz-Kohjerdi. Embedding linear arrays of the maximum length in O-shaped meshes. The Journal of Supercomputing, 78 (1):884–918, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03895-1.

Shimchenko:2022:ASP

[1114] Marina Shimchenko, Rubén Titos-Gil, and Alexandra Jimborean. Analysing software prefetching opportunities in hardware transactional memory. The Journal of Supercomputing, 78(1): 919–944, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03897-z.

Ritchie:2022:DPF

[1115] Robert Ritchie and Khodakhast Bibak. DOTMIX-Pro: faster and more efficient variants of DOTMIX for dynamicmultithreading platforms. The Journal of Supercomputing, 78(1):945– 961, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03904-3.

Mazinan:2022:MRB

[1116] Elham Mazinan, Hassan Naderi, and Saber Saati. Microblogs recommendations based on implicit similarity in content social networks. The Journal of Supercomputing, 78(1): 962–986, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03864-8.

Dhal:2022:CFG

[1117] K. Dhal, S. C. Rai, and S. Tripathy. CEMAR: a fine grained access control with revocation mechanism for centralized multi-authority cloud storage. The Journal of Supercomputing, 78 (1):987–1009, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03908-z.

Hu:2022:CFD

[1118] Shengzhou Hu, Xingfu Wang, and Tingting Zhong. Complex and flexible data access policy in attribute-based encryption. The Journal of Supercomputing, 78(1):1010–1029, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-021-03867-
5.

Sajedi:2022:FLF

[1119] Seyedeh Nafiseh Sajedi, Mohsen Maadani, and Meisam Nesari Moghadam. F-LEACH: a fuzzy-based data aggregation scheme for healthcare IoT systems. The Journal of Supercomputing, 78(1): 1030–1047, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03890-6.

Kuncan:2022:NAP

[1120] Fatma Kuncan, Yilmaz Kaya, and Melih Kuncan. A new approach for physical human activity recognition based on co-occurrence matrices. The Journal of Supercomputing, 78(1): 1048–1070, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03921-2.

Venkateswararao:2022:BPB

[1121] Kuna Venkateswararao and Pravati Swain. Binary-PSO-based energyefficient small cell deployment in 5G ultra-dense network. The Journal of Supercomputing, 78(1):1071– 1092, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03910-5.

Chen:2022:OSR

[1122] Che Chen, Rongzong Guo, and Chai Kiat Yeo. Optimal sequential relay-remote selection and computation offloading in mobile edge computing. The Journal of Supercomputing, 78 (1):1093–1116, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03919-w.

Rao:2022:SCQ

[1123] P. Venkateswara Rao and A. P. Siva Kumar. The societal communication of the Q&A community on topic modeling. The Journal of Supercomputing, 78 (1):1117–1143, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03852-y.

Bentaleb:2022:CTT

[1124] Ouafa Bentaleb, Adam S. Z. Belloum, and Aouaouche El-Maouhab. Containerization technologies: taxonomies, applications and challenges. The Journal of Supercomputing, 78(1): 1144–1181, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03914-1.

Nekooei-Joghdani:2022:DSI

[1125] Ahmad Nekooei-Joghdani and Faramarz Safi-Esfahani. Dynamic scheduling of independent tasks in cloud computing applying a new hybrid metaheuristic algorithm including Gabor filter, opposition-based learning, multiverse optimizer, and multi-tracker optimization algorithms. The Journal of Supercomputing, 78(1):1182– 1243, January 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03814-4.

Chaudhry:2022:MSO

[1126] S. R. Chaudhry, P. Liu, and M. Collier. A measurement study of offloading virtual network functions to the edge. The Journal of Supercomputing, 78(2): 1565–1582, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03907-0.

Ali:2022:PIT

[1127] Noureddine Ait Ali, Ahmed El abbassi, and Bouchaib Cherradi. The performances of iterative type-2 fuzzy Cmean on GPU for image segmentation. The Journal of Supercomputing, 78(2): 1583–1601, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03928-9.

Ali:2022:EUN

[1128] Zeeshan Ali, Aun Irtaza, and Muazzam Maqsood. An efficient U-Net framework for lung nodule detection using densely connected dilated convolutions. The Journal of Supercomputing, 78(2): 1602–1623, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03845-x.

Liu:2022:AID

[1129] Jun Liu. The analysis of innovative design and evaluation of energy storage system based on Internet of Things. The Journal of Supercomputing, 78(2): 1624–1641, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03931-0.

Bouhenni:2022:EPE

[1130] Sarra Bouhenni, Saïd Yahiaoui, and Hamamache Kheddouci. Efficient parallel edge-centric approach for relaxed graph pattern matching. The Journal of Supercomputing, 78(2):1642– 1671, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03938-7.

Ahmadpour:2022:EDQ

[1131] Seyed-Sajad Ahmadpour, Mohammad Mosleh, and Saeed Rasouli Heikalabad. Efficient designs of quantumdot cellular automata multiplexer and RAM with physical proof along with power analysis. The Journal of Supercomputing, 78(2):1672– 1695, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03913-2.

Du:2022:MPO

[1132] Qi Du and Hui Huang. MPI parameter optimization during debugging phase of HPC system. The Journal of Supercomputing, 78(2): 1696–1711, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03939-6.

Chen:2022:CBB

[1133] Jie Chen, Xianguo Qing, and Qian Sun. Consensus-based bundle algorithm with local replanning for heterogeneous multi-UAV system in the timesensitive and dynamic environment. The Journal of Supercomputing, 78(2): 1712–1740, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03940-z.

Wang:2022:HPB

[1134] Ruimin Wang, Zhiwei Yang, and Lu Lu. A high-performance batched matrix multiplication framework for GPUs under unbalanced input distribution. The Journal of Supercomputing, 78(2):1741–1758, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-021-03936-
9.

Yeon:2022:VAP

[1135] Hanbyul Yeon, Seongbum Seo, and Yun Jang. Visual analysis for panel data imputation with Bayesian network. The Journal of Supercomputing, 78(2): 1759–1782, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03934-x.

Ahanger:2022:NIF

[1136] Tariq Ahamed Ahanger, Usman Tariq, and Abdullah Sulman. A novel IoT-fog-cloud-based healthcare system for monitoring and predicting COVID-19 outspread. The Journal of Supercomputing, 78(2):1783– 1806, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03935-w.

Jiang:2022:DLT

[1137] Huiyan Jiang, Zhaoshuo Diao, and Yu-Dong Yao. Deep learning techniques for tumor segmentation: a review. The Journal of Supercomputing, 78(2): 1807–1851, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03901-6.

Wu:2022:TLC

[1138] Mingjie Wu, Qingkui Chen, and Jingjuan Wang. Toward low CPU usage and efficient DPDK communication in a cluster. The Journal of Supercomputing, 78(2):1852– 1884, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03942-x.

Zhang:2022:DRP

[1139] Mengzhao Zhang, Jeong-Geun Kim, and Shin-Dug Kim. Dynamic recognition prefetch engine for DRAM-PCM hybrid main memory. The Journal of Supercomputing, 78(2): 1885–1902, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03948-5.

Sheng:2022:TCM

[1140] Bao Sheng, Shi Wenzhong, and Xiang Haodong. A tight coupling mapping method to integrate the ESKF, g2o, and point cloud alignment. The Journal of Supercomputing, 78(2): 1903–1922, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03900-7.

Ahuja:2022:RTC

[1141] Nikita Ahuja, Harsha Bikkavilli, and Ariana Wilner. Real-time cellularlevel imaging and medical treatment with a swarm of wireless multifunctional robots. The Journal of Supercomputing, 78(2):1923– 1943, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03924-z.

Kirubanantham:2022:IWS

[1142] P. Kirubanantham, S. M. Udhaya Sankar, and P. C. Karthik. An intelligent web service group-based recommendation system for longterm composition. The Journal of Supercomputing, 78(2):1944– 1960, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03930-1.

Kim:2022:EED

[1143] Cheol jin Kim, Myung jae Lee, and Young guk Ha. End-to-end deep learning-based autonomous driving control for high-speed environment. The Journal of Supercomputing, 78(2): 1961–1982, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03929-8.

Gasmi:2022:SCO

[1144] Kaouther Gasmi, Selma Dilek, and Suat Ozdemir. A survey on computation offloading and service placement in fog computing-based IoT. The Journal of Supercomputing, 78(2): 1983–2014, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03941-y.

Safaei:2022:ESC

[1145] Alireza Safaei, Ramin Nassiri, and Amir Masoud Rahmani. Enterprise service composition models in IoT context: solutions comparison. The Journal of Supercomputing, 78(2): 2015–2042, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03873-7.

Wu:2022:PMD

[1146] Fu-Hsing Wu, Huey-Jen Lai, and Chih-Sheng Lin. Predictive models for detecting patients more likely to develop acute myocardial infarctions. The Journal of Supercomputing, 78(2): 2043–2071, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03916-z.

Idrees:2022:ESD

[1147] Ali Kadhum Idrees and Raphael Couturier. Energy-saving distributed monitoring-based firefly algorithm in wireless sensors networks. The Journal of Supercomputing, 78(2):2072– 2097, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03944-9.

Bahutair:2022:NFI

[1148] Mohammed Bahutair, Zaher Al Aghbari, and Ibrahim Kamel. NodeRank: Finding influential nodes in social networks based on interests. The Journal of Supercomputing, 78(2): 2098–2124, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03947-6.

Emami:2022:SET

[1149] Hojjat Emami. Stock exchange trading optimization algorithm: a humaninspired method for global optimization. The Journal of Supercomputing, 78(2):2125–2174, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-021-03943-
                                        w.

Rocco:2022:LFR

[1150] Roberto Rocco, Davide Gadioli, and Gianluca Palermo. Legio: fault resiliency for embarrassingly parallel MPI applications. The Journal of Supercomputing, 78(2):2175– 2195, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03951-w.

Hao:2022:DVP

[1151] Jiangwei Hao, Jinchen Xu, and YuanYuan Xia. Design of variable precision transcendental function automatic generator. The Journal of Supercomputing, 78(2):2196– 2218, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03937-8.

Bokharaie:2022:PSC

[1152] Vahhab Samadi Bokharaie and Ali Jahanian. Power side-channel leakage assessment and locating the exact sources of leakage at the early stages of ASIC design process. The Journal of Supercomputing, 78(2): 2219–2244, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03927-w.

Falcao:2022:AFU

[1153] Marcos Falcao, Caio Bruno Souza, and Kelvin Dias. An analytical framework for URLLC in hybrid MEC environments. The Journal of Supercomputing, 78(2):2245– 2264, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03945-8.

Li:2022:ALS

[1154] Baosheng Li, Baole Han, and Chuandong Qin. Application of large-scale L_{2}-SVM for microarray classification. The Journal of Supercomputing, 78(2): 2265–2286, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03962-7.

chandrika:2022:DAS

[1155] O. Mohana chandrika and M. Siva kumar. Design and analysis of SRAM cell using reversible logic gates towards smart computing. The Journal of Supercomputing, 78(2):2287– 2306, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03851-z.

Jalalian:2022:HMO

[1156] Zahra Jalalian and Mohsen Sharifi. A hierarchical multi-objective task scheduling approach for fast big data processing. The Journal of Supercomputing, 78(2):2307– 2336, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03960-9.

Khan:2022:HPN

[1157] Angshuman Khan and Rajeev Arya. High performance nanocomparator: a quantum dot cellular automatabased approach. The Journal of Supercomputing, 78(2):2337–2353, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03961-8.

Barai:2022:PMP

[1158] Atanu Barai, Yehia Arafa, and Stephan Eidenbenz. PPT-Multicore: performance prediction of OpenMP applications using reuse profiles and analytical modeling. The Journal of Supercomputing, 78(2):2354– 2385, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03949-4.

Yadav:2022:CCP

[1159] Vijay Kumar Yadav, Rakesh Kumar Yadav, and S. Venkatesan. CP2EH: a comprehensive privacy-preserving ehealth scheme over cloud. The Journal of Supercomputing, 78(2):2386– 2416, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03967-2.

Liao:2022:ODQ

[1160] Xia Liao, Yixian Shen, and Zhiguang Chen. Optimizing data query performance of bi-cluster for large-scale scientific data in supercomputers. The Journal of Supercomputing, 78(2): 2417–2441, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03965-4.

Renani:2022:NOH

[1161] Negin Bagheri Renani, Elham Yaghoubi, and Tofigh Abbasi. NLR-OP: a high-performance optical router based on North-Last turning model for multicore processors. The Journal of Supercomputing, 78(2):2442– 2476, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03920-3.

Vijayalakshmi:2022:AHP

[1162] T. Vijayalakshmi and R. Senthamarai. Application of homotopy perturbation and variational iteration methods for nonlinear imprecise preypredator model with stability analysis. The Journal of Supercomputing, 78(2): 2477–2502, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03956-5.

Shamsabadi:2022:CBM

[1163] Fatemeh Alidadi Shamsabadi and Shaghayegh Bakhtiari Chehelcheshmeh. A cloud-based mobile payment system using identity-based signature providing key revocation. The Journal of Supercomputing, 78(2): 2503–2527, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03830-4.

Aliakbarpour:2022:IRS

[1164] Hassan Aliakbarpour, Mohammad Taghi Manzuri, and Amir Masoud Rahmani. Improving the readability and saliency of abstractive text summarization using combination of deep neural networks equipped with auxiliary attention mechanism. The Journal of Supercomputing, 78(2):2528– 2555, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03950-x.

Shin:2022:CES

[1165] Hansub Shin, Kisung Lee, and Hyuk-Yoon Kwon. A comparative experimental study of distributed storage engines for big spatial data processing using GeoSpark. The Journal of Supercomputing, 78(2):2556– 2579, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03946-7.

Rekha:2022:DDA

[1166] D. Rekha, J. Sangeetha, and V. Ramaswamy. Digital document analytics using logistic regressive and deep transition-based dependency parsing. The Journal of Supercomputing, 78(2): 2580–2596, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03973-4.

Dehghani:2022:NAD

[1167] Abbas Dehghani, Ali Kavari, and Keyvan RahimiZadeh. A new approach for design of an efficient FPGA-based reconfigurable convolver for image processing. The Journal of Supercomputing, 78(2):2597– 2615, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03963-6.

Ashkouti:2022:DLD

[1168] Farough Ashkouti, Keyhan Khamforoosh, and Hana Khamfroush.

DHkmeans-ldiversity: distributed

hierarchical K-means for satisfaction of the l-diversity privacy model using Apache Spark. The Journal of Supercomputing, 78(2):2616– 2650, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03958-3.

Kanisha:2022:SCU

[1169] B. Kanisha, V. Mahalakshmi, and P. Kalyanasundaram. Smart communication using tri-spectral sign recognition for hearing-impaired people. The Journal of Supercomputing, 78(2): 2651–2664, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03968-1. See re-

traction note [1990].

Wu:2022:IVT

[1170] Hsin-Te Wu. The Internet-of-Vehicle traffic condition system developed by artificial intelligence of things. The Journal of Supercomputing, 78(2): 2665–2680, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03969-0.

Zhu:2022:EPE

[1171] Fushun Zhu and Hua Yan. An efficient parallel entropy coding method for JPEG compression based on GPU. The Journal of Supercomputing, 78(2): 2681–2708, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03971-6.

Zhang:2022:ICA

[1172] Tongpo Zhang, Xiaokai Nie, and Limin Yu. Improved camshift algorithm in AGV vision-based tracking with edge computing. The Journal of Supercomputing, 78(2):2709– 2723, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03974-3.

Kang:2022:FPO

[1173] Shinjin Kang and Soo Kyun Kim. Floor plan optimization for indoor environment based on multimodal data.

The Journal of Supercomputing, 78(2): 2724–2743, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03952-9.

Chen:2022:ASS

[1174] Ailin Chen, Pin Yang, and Pengsen Cheng. ACTSSD: social spammer detection based on active learning and co-training. The Journal of Supercomputing, 78(2):2744– 2771, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03966-3.

Ramakrishnan:2022:HPC

[1175] Jayabrabu Ramakrishnan, Rajesh Doss, and Karthik Srinivasan. High performance computation of human computer interface for neurodegenerative individuals using eye movements and deep learning technique. The Journal of Supercomputing, 78(2): 2772–2792, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03932-z.

Karmakar:2022:ABM

[1176] Kamalesh Karmakar, Rajib K. Das, and Sunirmal Khatua. An ACObased multi-objective optimization for cooperating VM placement in cloud data center. The Journal of Supercomputing, 78(3):3093– 3121, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03978-z.

ElBaz:2022:PBF

[1177] Didier El Baz, Bilal Fakih, and Vincent Boyer. Parallel best-first search algorithms for planning problems on multi-core processors. The Journal of Supercomputing, 78(3):3122– 3151, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03986-z.

Forsell:2022:PPC

[1178] Martti Forsell, Sara Nikula, and Jesper Larsson Träff. Performance and programmability comparison of the thick control flow architecture and current multicore processors. The Journal of Supercomputing, 78(3): 3152–3183, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03985-0.

Rahim:2022:ERS

[1179] Asma Rahim, Mehr Yahya Durrani, and Mucheol Kim. An efficient recommender system algorithm using trust data. The Journal of Supercomputing, 78(3):3184– 3204, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03991-2.

Zhang:2022:HEM

[1180] Yonghua Zhang, Hongxu Jiang, and Yu Du. High-efficient MPSoC-based CNNs accelerator with optimized storage and dataflow. The Journal of Supercomputing, 78(3):3205– 3225, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03909-y.

Shehab:2022:TFS

[1181] Noha Shehab, Mahmoud Badawy, and H. Arafat Ali. Toward feature selection in big data preprocessing based on hybrid cloud-based model. The Journal of Supercomputing, 78(3): 3226–3265, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03970-7.

Abdulhammed:2022:LBI

[1182] Omar Younis Abdulhammed. Load balancing of IoT tasks in the cloud computing by using sparrow search algorithm. The Journal of Supercomputing, 78(3):3266– 3287, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03989-w.

Chang:2022:EPS

[1183] Jia-Wei Chang. Enabling progressive system integration for AIoT and speech-based HCI through semanticaware computing. The Journal of Supercomputing, 78(3):3288–3324, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03996-x.

Korac:2022:MDI

[1184] Dragan Korać, Boris Damjanović, and Dejan Simić. A model of digital identity for better information security in e-learning systems. The Journal of Supercomputing, 78(3):3325– 3354, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03981-4. See cor-

rection [1185].

Korac:2022:CMD

[1185] Dragan Korać, Boris Damjanović, and Dejan Simić. Correction to: A model of digital identity for better information security in e-learning systems. The Journal of Supercomputing, 78 (3):3355, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04006-w.  See

[1184].

Kim:2022:TSP

[1186] Hyeonwoo Kim, Hyungjoon Kim, and Eenjun Hwang. Two-stage person re-identification scheme using crossinput neighborhood differences. The Journal of Supercomputing, 78(3): 3356–3373, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03994-z.

Khriji:2022:DIC

[1187] Sabrine Khriji, Yahia Benbelgacem, and Olfa Kanoun. Design and implementation of a cloud-based eventdriven architecture for real-time data processing in wireless sensor networks. The Journal of Supercomputing, 78(3): 3374–3401, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03955-6.

Venkateswarlu:2022:OCE

[1188] Chittetti Venkateswarlu and Nandanavanam Venkateswara Rao. Optimal channel estimation and interference cancellation in MIMO-OFDM system using MN-based improved AMO model. The Journal of Supercomputing, 78(3):3402– 3424, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03983-2.

Wang:2022:PAC

[1189] Yan Wang, Kenli Li, and Keqin Li. Performance-aware cache management for energy-harvesting nonvolatile processors. The Journal of Supercomputing, 78(3):3425– 3447, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03999-8.

Lu:2022:OMP

[1190] Jiawei Lu, Wei Zhao, and Gang Xiao. Optimal machine placement based on improved genetic algorithm in cloud computing. The Journal of Supercomputing, 78(3):3448– 3476, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03953-8.

Javadpour:2022:CIR

[1191] Amir Javadpour and Guojun Wang. cTMvSDN: improving resource management using combination of Markovprocess and TDMA in softwaredefined networking. The Journal of Supercomputing, 78(3):3477– 3499, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03871-9.

Xie:2022:ADL

[1192] Jun Xie, Jianwei Zhang, and Guohua Chen. Applying deep learning in football ankle injury for value of high-power magnetic resonance bioimaging evaluation. The Journal of Supercomputing, 78(3):3500– 3516, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04004-y.

Alawad:2022:DJR

[1193] Noor Aldeen Alawad and Bilal H. Abed-alguni. Discrete Jaya with refraction learning and three mutation methods for the permutation flow shop scheduling problem. The Journal of Supercomputing, 78(3):3517– 3538, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03998-9.

Sun:2022:CEU

[1194] Xiaoying Sun and Qiaoli Lu. Contrastenhanced ultrasound in optimization of treatment plans for diabetic nephropathy patients based on deep learning. The Journal of Supercomputing, 78(3): 3539–3560, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04002-0.

Li:2022:DBB

[1195] Chunlin Li, Qianqian Cai, and Youlong Luo. Data balancing-based intermediate data partitioning and check point-based cache recovery in Spark environment. The Journal of Supercomputing, 78(3):3561– 3604, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04000-2.

Atzori:2022:SVL

[1196] Marco Atzori, Wiebke Köpp, and Tino Weinkauf. In situ visualization of large-scale turbulence simulations in Nek5000 with ParaView Catalyst. The Journal of Supercomputing, 78(3): 3605–3620, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03990-3.

Lahouij:2022:OAC

[1197] Aida Lahouij, Lazhar Hamel, and Mohamed Graiet. An optimization approach for cloud composite services. The Journal of Supercomputing, 78(3): 3621–3645, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03995-y.

Mahmood:2022:IFD

[1198] Tariq Mahmood, Jianqiang Li, and Sirajuddin Qureshi. An intelligent fault detection approach based on reinforcement learning system in wireless sensor network. The Journal of Supercomputing, 78(3):3646– 3675, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04001-1.

Su:2022:SCI

[1199] Peng Su, Yuanyuan Chen, and Mengmeng Lu. Smart city information processing under Internet of Things and cloud computing. The Journal of Supercomputing, 78(3):3676– 3695, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03972-5.

Rana:2022:EDA

[1200] Saurabh Rana, Mohammad S. Obaidat, and Y. Sreenivasa Rao. Efficient design of an authenticated key agreement protocol for dewassisted IoT systems. The Journal of Supercomputing, 78(3):3696– 3714, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04003-z.

Alajlan:2022:ALM

[1201] Abrar Mohammed Alajlan and Marwah Mohammad Almasri. Automatic lane marking prediction using convolutional neural network and S-Shaped Binary Butterfly Optimization. The Journal of Supercomputing, 78(3): 3715–3745, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03988-x.

Toor:2022:ATC

[1202] Affan Ahmad Toor and Muhammad Usman. Adaptive telecom churn prediction for concept-sensitive imbalance data streams. The Journal of Supercomputing, 78(3):3746– 3774, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04021-x.

Zhang:2022:IPU

[1203] Jinlong Zhang and Shengye Huang.

Improving the performance of uplink visible light communication in urban streets. The Journal of Supercomputing, 78(3):3775– 3790, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04010-0.

Le:2022:FDP

[1204] Ye Le, Y. A. Nanehkaran, and Yimin Mao. FP-DCNN: a parallel optimization algorithm for deep convolutional neural network. The Journal of Supercomputing, 78(3):3791– 3813, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04012-y.

Bugingo:2022:CET

[1205] Emmanuel Bugingo, Zheng Wei, and Zhang Defu. Comparative evaluation of task priorities for processing and bandwidth capacities-based workflow scheduling for cloud environment. The Journal of Supercomputing, 78(3): 3814–3842, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-03979-y.

Lu:2022:SBA

[1206] Jing-You Lu, Hsu-Chao Lai, and Tian-Shyr Dai. Structural break-aware pairs trading strategy using deep reinforcement learning. The Journal of Supercomputing, 78(3):3843– 3882, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04013-x.

Abdennadher:2022:DDA

[1207] Imen Abdennadher. DAACS: a Decision Approach for Autonomic Computing Systems. The Journal of Supercomputing, 78(3):3883– 3904, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04011-z.

Chen:2022:RKS

[1208] Xiaozhong Chen, Rongli Chen, and Cairu Yang. Research to key success factors of intelligent logistics based on IoT technology. The Journal of Supercomputing, 78(3):3905– 3939, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04009-7.

Zhao:2022:NTS

[1209] Yuxuan Zhao, Ka Lok Man, and Sheng-Uei Guan. A novel two-stream structure for video anomaly detection in smart city management. The Journal of Supercomputing, 78(3):3940– 3954, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04007-9.

Chang:2022:ACD

[1210] Yigang Chang, Qian Wu, and Qiang Li. Adoption of combined detection technology of tumor markers via deep learning algorithm in diagnosis and prognosis of gallbladder carcinoma. The Journal of Supercomputing, 78(3): 3955–3975, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03843-z.

Wu:2022:TEK

[1211] Jimmy Ming-Tai Wu, Min Wei, and Shahab Tayeb. Top-k-dominating queries on incomplete large dataset. The Journal of Supercomputing, 78(3): 3976–3997, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04005-x.

Goldanloo:2022:HOB

[1212] Mina Javanmard Goldanloo and Farhad Soleimanian Gharehchopogh. A hybrid OBL-based firefly algorithm with symbiotic organisms search algorithm for solving continuous optimization problems. The Journal of Supercomputing, 78(3):3998– 4031, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04015-9.

Wang:2022:OES

[1213] Zhongmin Wang, Weiye Zhang, and Chen Lu. An optimal edge server placement approach for cost reduction and load balancing in intelligent manufacturing. The Journal of Supercomputing, 78(3):4032– 4056, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04017-7.

Kavitha:2022:EMD

[1214] D. Kavitha and C. H. Renumadhavi. An efficient multilayer deep detection perceptron (MLDDP) methodology for detecting testicular anomalies with or without congenital heart disease (TACHD). The Journal of Supercomputing, 78(3):4057– 4072, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04008-8.

Bai:2022:PRU

[1215] Qingchun Bai, Jie Zhou, and Liang He. PG-RNN: using position-gated recurrent neural networks for aspectbased sentiment classification. The Journal of Supercomputing, 78(3): 4073–4094, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04019-5.

Oz:2022:RSE

[1216] Isil Öz and Ömer Faruk Karadas.

Regional soft error vulnerability and error propagation analysis for GPGPU applications. The Journal of Supercomputing, 78(3):4095– 4130, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04026-6.

Aghdashi:2022:NDL

[1217] Arman Aghdashi and Seyedeh Leili Mirtaheri. Novel dynamic load balancing algorithm for cloud-based big data analytics. The Journal of Supercomputing, 78(3):4131– 4156, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04024-8.

Aghdashi:2022:CND

[1218] Arman Aghdashi and Seyedeh Leili Mirtaheri. Correction to: Novel dynamic load balancing algorithm for cloud-based big data analytics. The Journal of Supercomputing, 78 (3):4157, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04053-3.

Zhang:2022:NHA

[1219] Ting-Wei Zhang, Guang-Hui Xu, and Tao Han. A new hybrid algorithm for path planning of mobile robot. The Journal of Supercomputing, 78(3): 4158–4181, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04031-9.

Li:2022:IMD

[1220] Shanxi Li, Qingguo Zhou, and Qingquan Lv. Intelligent malware detection based on graph convolutional network. The Journal of Supercomputing, 78(3):4182– 4198, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04020-y.

Cai:2022:IOC

[1221] Weilin Cai, Heng Chen, and Xingjun Zhang. Implementation and optimization of ChaCha20 stream cipher on Sunway TaihuLight supercomputer. The Journal of Supercomputing, 78(3): 4199–4216, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04023-9.

Gonzalez-Dominguez:2022:MDP

[1222] Jorge González-Domínguez, José M.

Martín-Martínez, and Roberto R. Expósito. MPI-dot2dot: a parallel tool to find DNA tandem repeats on multicore clusters. The Journal of Supercomputing, 78(3):4217– 4235, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04025-7.

Yadav:2022:BOT

[1223] Ashish Mohan Yadav, Kuldeep Narayan Tripathi, and S. C. Sharma. A bi-objective task scheduling approach in fog computing using hybrid fireworks algorithm. The Journal of Supercomputing, 78(3):4236– 4260, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04018-6.

Grami:2022:EAS

[1224] Maziyar Grami. An energy-aware scheduling of dynamic workflows using big data similarity statistical analysis in cloud computing. The Journal of Supercomputing, 78(3):4261– 4289, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04016-8.

Huang:2022:FMD

[1225] Lingfeng Huang and Genggeng Liu.

Functional motion detection based on artificial intelligence. The Journal of Supercomputing, 78(3):4290– 4329, February 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04037-3.

Huang:2022:DID

[1226] Lingfeng Huang. Design of an IoT DDoS attack prediction system based on data mining technology. The Journal of Supercomputing, 78(4): 4601–4623, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04055-1.

Nabi:2022:PRP

[1227] Said Nabi and Masroor Ahmed. PSO-RDAL: particle swarm optimizationbased resource- and deadline-aware dynamic load balancer for deadline constrained cloud tasks. The Journal of Supercomputing, 78(4):4624– 4654, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04062-2.

Raheja:2022:SIF

[1228] Supriya Raheja, Mohammed Alshehri, and Thompson Stephan. A smart intuitionistic fuzzy-based framework for round-robin short-term scheduler. The Journal of Supercomputing, 78 (4):4655–4679, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04052-4. See cor-

rection [1229].

Raheja:2022:CSI

[1229] Supriya Raheja, Mohammed Alshehri, and Thompson Stephan. Correction to: A smart intuitionistic fuzzy-based framework for roundrobin short-term scheduler. The Journal of Supercomputing, 78(4): 4680, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04096-6.  See

[1228].

Zhao:2022:FER

[1230] Dezhu Zhao, Yufeng Qian, and Min Yang. The facial expression recognition technology under image processing and neural network. The Journal of Supercomputing, 78(4): 4681–4708, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04058-y.

Safa:2022:ADD

[1231] Ramin Safa, Peyman Bayat, and Leila Moghtader. Automatic detection of depression symptoms in Twitter using multimodal analysis. The Journal of Supercomputing, 78(4):4709– 4744, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04040-8.

Kumaran:2022:DAO

[1232] S. Senthil Kumaran, S. P. Balakannan, and Jun Li. A deep analysis of object capabilities for intelligence considering wireless IoT devices with the DNN approach. The Journal of Supercomputing, 78(4): 4745–4758, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04064-0.

Shi:2022:OEC

[1233] Jiawen Shi, Hong Li, and Chiyu Wang. Optimizing emotion-cause pair extraction task by using mutual assistance single-task model, clause position information and semantic features. The Journal of Supercomputing, 78 (4):4759–4778, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04067-x.

Li:2022:NSC

[1234] Jianjiang Li, Zhaochu Deng, and Jie Lin. A new software cache structure on Sunway TaihuLight. The Journal of Supercomputing, 78(4): 4779–4798, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04056-0.

Alkhateeb:2022:DHC

[1235] Faisal Alkhateeb, Bilal H. Abed-alguni, and Mohammad Hani Al-rousan. Discrete hybrid cuckoo search and simulated annealing algorithm for solving the job shop scheduling problem. The Journal of Supercomputing, 78 (4):4799–4826, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04050-6.

Cao:2022:SBP

[1236] Hao Cao, Shaozhong Guo, and Jinchen Xu. Superblock-based performance optimization for Sunway Math Library on SW26010 many-core processor. The Journal of Supercomputing, 78 (4):4827–4849, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03997-w.

Li:2022:SBS

[1237] Chunlin Li, Jing Zhang, and Xianmin Yang. Scalable blockchain storage mechanism based on two-layer structure and improved distributed consensus. The Journal of Supercomputing, 78 (4):4850–4881, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04061-3.

Moori:2022:LEL

[1238] Ayeh Moori, Behrang Barekatain, and Mehdi Akbari. LATOC: an enhanced load balancing algorithm based on hybrid AHP-TOPSIS and OPSO algorithms in cloud computing. The Journal of Supercomputing, 78(4): 4882–4910, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04042-6.

Shamsi:2022:RLT

[1239] Mahboubeh Shamsi, Abdolreza Rasouli Kenari, and Roghayeh Aghamohammadi. Reinforcement learning for traffic light control with emphasis on emergency vehicles. The Journal of Supercomputing, 78(4):4911– 4937, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04068-w.

Meshram:2022:EAK

[1240] Chandrashekhar Meshram, Rabha W. Ibrahim, and Agbotiname Lucky Imoize. An efficient authentication with key agreement procedure using Mittag-Leffler–Chebyshev summation chaotic map under the multiserver architecture. The Journal of Supercomputing, 78(4):4938– 4959, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04039-1.

Nain:2022:NNB

[1241] Sweety Nain and Prachi Chaudhary. A neural network-based approach for the performance evaluation of branch prediction in instructionlevel parallelism processors. The Journal of Supercomputing, 78(4):4960– 4976, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04045-3.

Smara:2022:RIC

[1242] Mounya Smara, Makhlouf Aliouat, and Al-Sakib Khan Pathan. Robustness improvement of componentbased cloud computing systems. The Journal of Supercomputing, 78(4): 4977–5009, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04054-2.

Liu:2022:DLI

[1243] Qingquan Liu, Yingli Wu, and Xin Li. Deep learning in the information service system of agricultural Internet of Things for innovation enterprise. The Journal of Supercomputing, 78(4):5010–5028, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04070-2.

Liu:2022:EIT

[1244] Xiao Liu, De yu Qi, and Hao tong Zhang. Exploring the Internet of Things sequence-structure detection and supertask network generation of temporal-spatial-based graph convolutional neural network. The Journal of Supercomputing, 78(4):5029– 5049, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04041-7.

Prateek:2022:RFL

[1245] Prateek and Rajeev Arya. Range free localization technique under erroneous estimation in wireless sensor networks. The Journal of Supercomputing, 78 (4):5050–5074, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04075-x.

Alazzam:2022:BPM

[1246] Hadeel Alazzam, Orieb AbuAlghanam, and Ahmad Sharieh. Best path in mountain environment based on parallel A∗ algorithm and Apache Spark. The Journal of Supercomputing, 78 (4):5075–5094, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04072-0.

Wang:2022:CBT

[1247] Zhongmin Wang, Gang Wang, and Jianwei Wang. Caching-based task scheduling for edge computing in intelligent manufacturing. The Journal of Supercomputing, 78(4):5095– 5117, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04071-1.

Gonzalez-Sanchez:2022:PMO

[1248] Belen Gonzalez-Sanchez, Miguel A. Vega-Rodríguez, and Sergio Santander-Jiménez. Parallel multi-objective optimization approaches for protein encoding. The Journal of Supercomputing, 78 (4):5118–5148, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04073-z.

Cao:2022:HCJ

[1249] Jian Cao and Yongjiang Guo. Handoff calls’ joining behavior and incentive mechanism in wireless cellular networks with retrial orbit. The Journal of Supercomputing, 78(4): 5149–5180, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04059-x.

Mao:2022:MBE

[1250] YiMin Mao, DeJin Gan, and XueYu Huang. A MapReduce-based Kmeans clustering algorithm. The Journal of Supercomputing, 78(4): 5181–5202, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04078-8.

Jain:2022:MLE

[1251] Praphula Kumar Jain, Rajendra Pamula, and Ephrem Admasu Yekun. A multi-label ensemble predicting model to service recommendation from social media contents. The Journal of Supercomputing, 78(4):5203– 5220, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04087-7.

Teimourian:2022:PWE

[1252] Hanifa Teimourian, Amir Teimourian, and Fadi Al-Turjman. The potential of wind energy via an intelligent IoT-oriented assessment. The Journal of Supercomputing, 78(4): 5221–5240, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04085-9.

Ishwarya:2022:PER

[1253] K. Ishwarya and A. Alice Nithya.

Performance-enhanced real-time lifestyle tracking model based on human activity recognition (PERT-HAR) model through smartphones. The Journal of Supercomputing, 78(4):5241– 5268, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04065-z.

Long:2022:DYP

[1254] Chhaihuoy Long, Eunhye Jo, and Yunyoung Nam. Development of a yoga posture coaching system using an interactive display based on transfer learning. The Journal of Supercomputing, 78 (4):5269–5284, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04076-w.

Pu:2022:AFR

[1255] Ying-Hung Pu, Po-Sheng Chiu, and Shih-Syun Lin. Aerial face recognition and absolute distance estimation using drone and deep learning. The Journal of Supercomputing, 78 (4):5285–5305, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04088-6.

Zeutouo:2022:HPC

[1256] Jerry Lacmou Zeutouo, Vianney Kengne Tchendji, and Jean Frédéric Myoupo. High-performance CGM-based parallel algorithms for minimum cost parenthesizing problem. The Journal of Supercomputing, 78(4):5306– 5332, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04069-9.

Karthik:2022:SUB

[1257] E. Karthik and T. Sethukarasi. Sarcastic user behavior classification and prediction from social media data using firebug swarm optimization-based long short-term memory. The Journal of Supercomputing, 78(4):5333– 5357, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04028-4.

Banaei:2022:EPS

[1258] Ali Banaei and Mohsen Sharifi.

ETAS: predictive scheduling of functions on worker nodes of Apache OpenWhisk platform. The Journal of Supercomputing, 78(4):5358– 5393, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04057-z.

Yan:2022:DWS

[1259] Zhang Yan, Du Hongle, and Yeh-Cheng Chen. Dynamic weighted selective ensemble learning algorithm for imbalanced data streams. The Journal of Supercomputing, 78(4): 5394–5419, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04084-w.

Chen:2022:PXQ

[1260] Rongxin Chen, Zhijin Wang, and Zongyue Wang. Parallel XPath query based on cost optimization. The Journal of Supercomputing, 78(4): 5420–5449, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04074-y.

Sinha:2022:ISO

[1261] Sankhamita Sinha, Subhayan Bhattacharya, and Sarbani Roy. Impact of second-order network motif on online social networks. The Journal of Supercomputing, 78(4): 5450–5478, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04079-7.

Shawkat:2022:OFG

[1262] Mai Shawkat, Mahmoud Badawi, and Ali El-desoky. An optimized FP-growth algorithm for discovery of association rules. The Journal of Supercomputing, 78(4):5479– 5506, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04066-y.

Hou:2022:ARN

[1263] Jiandong Hou and Zhisheng Tian. Application of recurrent neural network in predicting athletes’ sports achievement. The Journal of Supercomputing, 78(4):5507–5525, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04082-y.

Naskath:2022:CAV

[1264] J. Naskath, B. Paramasivan, and Hamza Aldabbas. Connectivity analysis of V2V communication with discretionary lane changing approach. The Journal of Supercomputing, 78 (4):5526–5546, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04086-8.

Huo:2022:ATE

[1265] Huaying Huo, Yigang Chang, and Yu Tang. Analysis of treatment effect of acupuncture on cervical spondylosis and neck pain with the data mining technology under deep learning. The Journal of Supercomputing, 78 (4):5547–5564, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-03959-2.

Xiao:2022:INB

[1266] Danyang Xiao, Xinxin Li, and Weigang Wu. Iteration number-based hierarchical gradient aggregation for distributed deep learning. The Journal of Supercomputing, 78(4):5565– 5587, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04083-x.

Zhang:2022:MAM

[1267] Ying Zhang and Haifeng Zhao. A multi-agent model for decision making on environmental regulation in urban agglomeration. The Journal of Supercomputing, 78(4):5588– 5609, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04094-8.

Refaee:2022:TEA

[1268] Eshrag A. Refaee and Shermin Shamsudheen. Trust- and energy-aware cluster head selection in a UAV-based wireless sensor network using Fit-FCM. The Journal of Supercomputing, 78 (4):5610–5625, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04092-w.

Kazemian:2022:CPS

[1269] Mohammad Mahdi Kazemian and Meghdad Mirabi. Controller placement in software defined networks using multi-objective antlion algorithm. The Journal of Supercomputing, 78 (4):5626–5649, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04109-4.

Luo:2022:MMB

[1270] Le Luo, Yi Liu, and Depei Qian. Magas: matrix-based asynchronous graph analytics on shared memory systems. The Journal of Supercomputing, 78 (4):5650–5680, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04091-x.

Ma:2022:CIC

[1271] Haizhou Ma and Aiping Ding. Construction and implementation of a college talent cultivation system under deep learning and data mining algorithms. The Journal of Supercomputing, 78(4):5681– 5696, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04036-4.

Yang:2022:ESP

[1272] Yunfei Yang, Jingwei Yang, and Xiantao Huang. Evaluation of sports public service under fuzzy integral and deep neural network. The Journal of Supercomputing, 78(4):5697– 5711, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04110-x.

Salehan:2022:CVO

[1273] Alireza Salehan and Arash Deldari.

Corona virus optimization (CVO): a novel optimization algorithm inspired from the Corona virus pandemic. The Journal of Supercomputing, 78 (4):5712–5743, March 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04100-z.

Rangwani:2022:FFM

[1274] Diksha Rangwani and Hari Om. Fourfactor mutual authentication scheme for health-care based on wireless body area network. The Journal of Supercomputing, 78(4):5744– 5778, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04099-3.

Nikmard:2022:CAD

[1275] Babak Nikmard, Naser Movahhedinia, and Mohammad Reza Khayyambashi. Congestion avoidance by dynamically cache placement method in named data networking. The Journal of Supercomputing, 78(4):5779– 5805, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04080-0.

Liu:2022:LZN

[1276] Qinghua Liu, Jiadui Chen, and Kai Yang. Linguistic Z-numbers and cloud model weighted ranking technology and its application in concept evaluation of information axiom. The Journal of Supercomputing, 78 (5):6061–6089, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04106-7.

Li:2022:WOA

[1277] Maodong Li, Guanghui Xu, and Xilin Zhao. Whale optimization algorithm based on dynamic pinhole imaging and adaptive strategy. The Journal of Supercomputing, 78(5):6090– 6120, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04116-5.

Suseendran:2022:MSI

[1278] G. Suseendran, D. Akila, and Anand Nayyar. Multi-sensor information fusion for efficient smart transport vehicle tracking and positioning based on deep learning technique. The Journal of Supercomputing, 78(5): 6121–6146, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04115-6.

Zhang:2022:SDP

[1279] Qing Zhang and Junhua Ren. Softwaredefect prediction within and across projects based on improved selforganizing data mining. The Journal of Supercomputing, 78(5):6147– 6173, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04113-8.

Hsu:2022:PBA

[1280] Ke-Jou Hsu and Jerry Chou. Performance benchmarking and autotuning for scientific applications on virtual cluster. The Journal of Supercomputing, 78(5):6174– 6206, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04103-w.

Sharma:2022:HLS

[1281] Gaurav Sharma, Lava Bhargava, and Vinod Kumar. Hybrid learning scenario path selection and abstraction framework for smart verification of complex SoCs. The Journal of Supercomputing, 78(5):6207– 6233, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04117-4.

Chauhan:2022:SAO

[1282] Sumika Chauhan, Govind Vashishtha, and Anil Kumar. A symbiosis of arithmetic optimizer with slime mould algorithm for improving global optimization and conventional design problem. The Journal of Supercomputing, 78 (5):6234–6274, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04105-8.

Hagras:2022:SEM

[1283] Tarek Hagras. Slack extender mechanism for greening dependenttasks scheduling on DVFS-enabled computing platforms. The Journal of Supercomputing, 78(5):6275– 6295, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04112-9.

Li:2022:CIT

[1284] Yongbin Li. Construction of Internet of Things English terms model and analysis of language features via deep learning. The Journal of Supercomputing, 78(5):6296–6317, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04130-7.

Cui:2022:ESB

[1285] Huanyu Cui, Nianbin Wang, and Yuezhu Xu. An effective SPMV based on block strategy and hybrid compression on GPU. The Journal of Supercomputing, 78(5):6318– 6339, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04123-6.

Esmaeilyfard:2022:IDW

[1286] Rasool Esmaeilyfard. Improving detection of web service antipatterns using crowdsourcing. The Journal of Supercomputing, 78(5):6340– 6370, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04134-3.

Qi:2022:AKE

[1287] Mingping Qi and Jianhua Chen. Authentication and key establishment protocol from supersingular isogeny for mobile environments. The Journal of Supercomputing, 78(5):6371– 6385, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04121-8.

Yang:2022:SPS

[1288] Canghai Yang, Han Wang, and Long Zuo. Semi-partitioned scheduling of mixed-criticality system on multiprocessor platforms. The Journal of Supercomputing, 78(5):6386– 6410, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04101-y.

Wankhade:2022:CCB

[1289] Mayur Wankhade, Chandra Sekhara Rao Annavarapu, and Mukul Kirti Verma.

CBVoSD: context based vectors over sentiment domain ensemble model for review classification. The Journal of Supercomputing, 78(5):6411– 6447, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04132-5.

Madera-Ramirez:2022:CGB

[1290] F. A. Madera-Ramirez, J. L. Lopez-Martinez, and J. Gomez-Montalvo. CPU-GPU buffer communication using compute shader to fill volumes with spheres. The Journal of Supercomputing, 78(5):6448– 6460, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04136-1.

Fakhouri:2022:SHI

[1291] Hussam N. Fakhouri, Faten Hamad, and Abedalsalam Alawamrah. Success history intelligent optimizer. The Journal of Supercomputing, 78(5): 6461–6502, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04093-9.

Wan:2022:FCS

[1292] Chang-Xuan Wan and Bo Li. Financial causal sentence recognition based on BERT-CNN text classification. The Journal of Supercomputing, 78(5):6503–6527, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04097-5.

Freire:2022:TSC

[1293] Daniela L. Freire, Rafael Z. Frantz, and Vitor Basto-Fernandes. Task scheduling characterisation in enterprise application integration. The Journal of Supercomputing, 78(5): 6528–6566, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04119-2.

Wen:2022:CSO

[1294] Heng Wen, Su Xin Wang, and Ma Cong Si. Colony search optimization algorithm using global optimization. The Journal of Supercomputing, 78 (5):6567–6611, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04127-2.

Kumar:2022:CSM

[1295] Sanjeev Kumar and Richa Agrawal.

A comprehensive survey on metaheuristic-based energy minimization routing techniques for wireless sensor network: classification and challenges. The Journal of Supercomputing, 78 (5):6612–6663, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04128-1.

Yadav:2022:PEN

[1296] Sonal Yadav and Ritu Raj. Power efficient network selector placement in control plane of multiple networks-on-chip. The Journal of Supercomputing, 78(5):6664–6695, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04098-4.

Liu:2022:LNN

[1297] Tianyi Liu, Shuoyuan Wang, and Lei Zhang. A lightweight neural network framework using linear grouped convolution for human activity recognition on mobile devices. The Journal of Supercomputing, 78(5):6696– 6716, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04140-5.

Sun:2022:IIS

[1298] Tao Sun, Zhengqiang Xiong, and Zhengxing Wang. Infrared image super-resolution method for edge computing based on adaptive nonlocal means. The Journal of Supercomputing, 78(5):6717– 6738, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04141-4.

Xu:2022:LEA

[1299] Liqiong Xu and Shuming Zhou. An O(log_{2} N) algorithm for reliability assessment of augmented cubes based on h-extra edge-connectivity. The Journal of Supercomputing, 78(5): 6739–6751, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04129-0.

Li:2022:AHB

[1300] Chunlin Li, Yong Zhang, and Youlong Luo. Adaptive handover based on traffic balancing and multi-dimensional collaborative resource management in MEC environment. The Journal of Supercomputing, 78(5):6752– 6787, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04120-9.

Palit:2022:ABB

[1301] Sudip Kumar Palit, Mohuya Chakraborty, and Subhalaxmi Chakraborty. AUGChain: blockchain-based mobile user authentication scheme in global mobility network. The Journal of Supercomputing, 78(5):6788–6816, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04139-y.

Nath:2022:CDM

[1302] Shubha Brata Nath, Subhrendu Chattopadhyay, and Soumya K. Ghosh. Containerized deployment of microservices in fog devices: a reinforcement learning-based approach. The Journal of Supercomputing, 78(5): 6817–6845, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04135-2.

Dehbozorgi:2022:RPM

[1303] Leila Dehbozorgi, Reza Sabbaghi-Nadooshan, and Alireza Kashaninia. Realization of processing-in-memory using binary and ternary quantumdot cellular automata. The Journal of Supercomputing, 78(5):6846– 6874, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04152-1.

Peng:2022:ALF

[1304] Ling Peng and Dongbo Zhang. An adaptive Lévy flight firefly algorithm for multilevel image thresholding based on Rényi entropy. The Journal of Supercomputing, 78(5): 6875–6896, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04150-3. See cor-

rection [1305].

Peng:2022:CAL

[1305] Ling Peng and Dongbo Zhang. Correction to: An adaptive Lévy flight firefly algorithm for multilevel image thresholding based on Rényi entropy. The Journal of Supercomputing, 78(5):6897, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04215-3.  See

[1304].

Khallouli:2022:CRS

[1306] Wael Khallouli and Jingwei Huang. Cluster resource scheduling in cloud computing: literature review and research challenges. The Journal of Supercomputing, 78(5):6898– 6943, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04138-z.

Gupta:2022:UHE

[1307] Vedika Gupta, Stuti Juyal, and Yu-Chen Hu. Understanding human emotions through speech spectrograms using deep neural network. The Journal of Supercomputing, 78(5):6944– 6973, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04124-5.

Kadry:2022:ASL

[1308] Seifedine Kadry, Venkatesan Rajinikanth, and Xiomara Patricia Blanco Valencia. Automated segmentation of leukocyte from hematological images — a study using various CNN schemes. The Journal of Supercomputing, 78 (5):6974–6994, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04125-4.

Wang:2022:EGE

[1309] Xinyang Wang, Lijuan Huang, and Keqin Li. The g-extra diagnosability of the balanced hypercube under the PMC and MM* model. The Journal of Supercomputing, 78(5): 6995–7015, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04126-3.

Wang:2022:NPD

[1310] Dongmei Wang, Yiwen Liang, and Xinmin Yang. NKA: a pathogen dosebased natural killer cell algorithm and its application to classification. The Journal of Supercomputing, 78 (5):7016–7037, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04133-4.

Xie:2022:PDA

[1311] Yuhao Xie, Jiarong Liang, and Changzhen Li. The properties and t/s-diagnosability of k-ary n-cube networks. The Journal of Supercomputing, 78(5):7038–7057, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04155-y. See cor-

rection [1312].

Xie:2022:CPE

[1312] Yuhao Xie, Jiarong Liang, and Changzhen Li. Correction to: The properties and t/s-diagnosability of k-ary n-cube networks. The Journal of Supercomputing, 78(5): 7058, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04228-y.  See

[1311].

Garcia-Molla:2022:PSD

[1313] Victor M. Garcia-Molla, M. Angeles Simarro, and Alberto Gonzalez. Parallel signal detection for generalized spatial modulation MIMO systems. The Journal of Supercomputing, 78 (5):7059–7077, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04163-y.

Kallel:2022:HBF

[1314] Ameni Kallel, Molka Rekik, and Mahdi Khemakhem. Hybrid-based framework for COVID-19 prediction via federated machine learning models. The Journal of Supercomputing, 78 (5):7078–7105, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04166-9.

Yang:2022:SCM

[1315] XianBen Yang and Wei Zhang. Simulation of cross-modal image-text retrieval algorithm under convolutional neural network structure and hash method. The Journal of Supercomputing, 78 (5):7106–7132, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04157-w.

Xu:2022:HQG

[1316] Saijuan Xu, Ling Wei, and Guolong Chen. A high-quality global routing algorithm based on hybrid topology optimization and heuristic search for data processing in MEC. The Journal of Supercomputing, 78 (5):7133–7157, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04147-y.

Morinigo:2022:ERT

[1317] José A. Moríñigo, Andrés Bustos, and Rafael Mayo-García. Error resilience of three GMRES implementations under fault injection. The Journal of Supercomputing, 78(5): 7158–7185, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04148-x.

Zhao:2022:OCB

[1318] Wen-Long Zhao, Wu Wang, and Qiao Wang. Optimization of cosmological N-body simulation with FMM-PM on SIMT accelerators. The Journal of Supercomputing, 78(5):7186– 7205, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04153-0.

Balakrishnan:2022:DLA

[1319] Vimala Balakrishnan, Zhongliang Shi, and Yue Fan. A deep learning approach in predicting products’ sentiment ratings: a comparative analysis. The Journal of Supercomputing, 78(5):7206–7226, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04169-6.

Chen:2022:ODA

[1320] Zhonglin Chen, Cuili Yang, and Junfei Qiao. The optimal design and application of LSTM neural network based on the hybrid coding PSO algorithm. The Journal of Supercomputing, 78 (5):7227–7259, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04142-3.

Trueman:2022:GBM

[1321] Tina Esther Trueman, P. Narayanasamy, and J. Ashok Kumar. A graphbased method for ranking of cloud service providers. The Journal of Supercomputing, 78(5):7260–7277, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-021-04156-
x.

Han:2022:SPL

[1322] Ling Bo Han, Bin Lao, and Ge Nong.

Succinct parallel Lempel–Ziv factorization on a multicore computer. The Journal of Supercomputing, 78 (5):7278–7303, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04165-w.

Zhu:2022:RRF

[1323] Xiaojun Zhu, Yu Wang, and Zhengrui Qin. RoV: receiving files from voice calls using dual-tone multi-frequency method. The Journal of Supercomputing, 78(5):7304– 7320, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04168-7.

Kadry:2022:ADA

[1324] Seifedine Kadry, V. Rajinikanth, and Elena Verdú. Automated detection of age-related macular degeneration using a pre-trained deep-learning scheme. The Journal of Supercomputing, 78 (5):7321–7340, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04181-w.

Durai:2022:CPP

[1325] C. Anand Deva Durai, Arshiya Begum, and Asfia Sabahath. COVID-19 pandemic, predictions and control in Saudi Arabia using SIR-F and age-structured SEIR model. The Journal of Supercomputing, 78 (5):7341–7353, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04149-w.

Bernal:2022:ECI

[1326] Adrián Bernal, M. Emilia Cambronero, and Valentín Valero. Evaluating cloud interactions with costs and SLAs. The Journal of Supercomputing, 78 (6):7529–7555, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04197-2.

Fuentes-Alventosa:2022:CEG

[1327] Antonio Fuentes-Alventosa, Juan Gómez-Luna, and R. Medina-Carnicer. CAVLCU: an efficient GPU-based implementation of CAVLC. The Journal of Supercomputing, 78(6):7556– 7590, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04183-8.

Beri:2022:NFC

[1328] Rydhm Beri, Mithilesh K. Dubey, and Aman Singh. A novel fogcomputing-assisted architecture of ehealthcare system for pregnant women. The Journal of Supercomputing, 78 (6):7591–7615, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04176-7.

Toshpulatov:2022:HPH

[1329] Mukhiddin Toshpulatov, Wookey Lee, and Arousha Haghighian Roudsari. Human pose, hand and mesh estimation using deep learning: a survey. The Journal of Supercomputing, 78 (6):7616–7654, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04184-7.

Hoffmann:2022:ORP

[1330] Renato B. Hoffmann, Júnior Löff, and Luiz G. Fernandes. OpenMP as runtime for providing high-level stream parallelism on multi-cores. The Journal of Supercomputing, 78 (6):7655–7676, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04182-9.

Durai:2022:NAE

[1331] Anand Deva Durai, Mythily Ganesh, and Dinesh Kumar Anguraj. A novel approach with an extensive case study and experiment for automatic code generation from the XMI schema of UML models. The Journal of Supercomputing, 78(6):7677– 7699, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04164-x.

Sharma:2022:BBC

[1332] Pratima Sharma, Rajni Jindal, and Malaya Dutta Borah. Blockchainbased cloud storage system with CP-ABE-based access control and revocation process. The Journal of Supercomputing, 78(6):7700– 7728, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04179-4.

Salami:2022:OEE

[1333] Bagher Salami, Hamid Noori, and Mahmoud Naghibzadeh. Online energy-efficient fair scheduling for heterogeneous multi-cores considering shared resource contention. The Journal of Supercomputing, 78(6): 7729–7748, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04159-8.

Chiang:2022:MCA

[1334] Chang-Po Chiang, Shiuh-Jeng Wang, and Yu-Shu Chen. Manipulating cyber army in pilot case forensics on social media. The Journal of Supercomputing, 78(6):7749–7767, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04172-x.

Mahdiani:2022:ZVN

[1335] H. R. Mahdiani. Z-Voter: a novel high-impedance voter for efficient realization of tristate logic in quantumdot cellular automata technology. The Journal of Supercomputing, 78(6): 7768–7787, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04167-8.

Cai:2022:AIR

[1336] Hang Cai. Application of intelligent real-time image processing in fitness motion detection under Internet of Things. The Journal of Supercomputing, 78(6):7788– 7804, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04145-0.

Liu:2022:RPO

[1337] Zongyun Liu and Jingqi Fu. Resource pricing and offloading decisions in mobile edge computing based on the Stackelberg game. The Journal of Supercomputing, 78(6):7805– 7824, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04246-w.

Belusso:2022:SSC

[1338] Cássio L. M. Belusso, Sandro Sawicki, and Fabricia Roos-Frantz. Selecting services in the cloud: a decision support methodology focused on infrastructure-as-a-service context. The Journal of Supercomputing, 78 (6):7825–7860, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04248-8.

Baranwal:2022:TTR

[1339] Gaurav Baranwal and Deo Prakash Vidyarthi. TRAPPY: a truthfulness and reliability aware application placement policy in fog computing. The Journal of Supercomputing, 78 (6):7861–7887, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04187-4.

Jin:2022:ODM

[1340] Xiaomin Jin, Feng Gao, and Yanping Chen. Optimal deployment of mobile cloudlets for mobile applications in edge computing. The Journal of Supercomputing, 78(6):7888– 7907, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04122-7.

Keshvari:2022:CBS

[1341] Roohollah Keshvari, Maryam Imani, and Mohsen Parsa Moghaddam. A clustering-based short-term load forecasting using independent component analysis and multi-scale decomposition transform. The Journal of Supercomputing, 78(6):7908– 7935, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04195-4.

Zhang:2022:UAR

[1342] Mingzu Zhang, Hongxi Liu, and Wenshui Lin. A unified approach to reliability and edge fault tolerance of cube-based interconnection networks under three hypotheses. The Journal of Supercomputing, 78(6):7936– 7947, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04185-6.

Nayak:2022:TST

[1343] Rajendra Prasad Nayak, Srinivas Sethi, and Deepak Puthal. TFMD-SDVN: a trust framework for misbehavior detection in the edge of softwaredefined vehicular network. The Journal of Supercomputing, 78(6):7948– 7981, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04227-z.

Li:2022:TSD

[1344] Xiaomei Li, Zhijiang Xie, and Yangjun Pi. Traffic sign detection based on improved faster r-CNN for autonomous driving. The Journal of Supercomputing, 78(6):7982– 8002, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04230-4.

Saxena:2022:OTO

[1345] Deepika Saxena and Ashutosh Kumar Singh. OFP-TM: an online VM failure prediction and tolerance model towards high availability of cloud computing environments. The Journal of Supercomputing, 78(6): 8003–8024, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04235-z.

Rajasekar:2022:FDD

[1346] P. Rajasekar and Yogesh Palanichamy. A flexible deadline-driven resource provisioning and scheduling algorithm for multiple workflows with VM sharing protocol on WaaS-cloud. The Journal of Supercomputing, 78(6): 8025–8055, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04225-1.

Bravo-Montes:2022:DIE

[1347] J. A. Bravo-Montes, A. Martín-Toledano, and F. Garcia-Herrero. Design and implementation of efficient QCA full-adders using faulttolerant majority gates. The Journal of Supercomputing, 78(6):8056– 8080, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04247-9.

Goncalves:2022:ELL

[1348] Marcio M. Goncalves, Josie E. Rodriguez Condia, and Jose Rodrigo Azambuja. Evaluating low-level software-based hardening techniques for configurable GPU architectures. The Journal of Supercomputing, 78 (6):8081–8105, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04154-z.

Najafimehr:2022:HML

[1349] Mohammad Najafimehr, Sajjad Zarifzadeh, and Seyedakbar Mostafavi. A hybrid machine learning approach for detecting unprecedented DDoS attacks. The Journal of Supercomputing, 78 (6):8106–8136, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04253-x.

Abdelghani:2022:DSM

[1350] Wafa Abdelghani, Ikram Amous, and Geoffrey Roman-Jimenez. Dynamic and scalable multi-level trust management model for Social Internet of Things. The Journal of Supercomputing, 78(6):8137– 8193, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04205-5.

Safari:2022:LTA

[1351] Maede Safari, Zahra Shirmohammadi, and Hamed Farbeh. LETHOR: a thermal-aware proactive routing algorithm for 3D NoCs with less entrance to hot regions. The Journal of Supercomputing, 78(6): 1–25, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04207-3.

Hsieh:2022:NSR

[1352] Sun-Yuan Hsieh, Chih-Wei Hsu, and Geng-Hua Zhang. Novel scheme for reducing communication data traffic in advanced metering infrastructure networks. The Journal of Supercomputing, 78(6):8219–8246, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04143-2.

Javed:2022:PCPa

[1353] Zeeshan Javed and Gon-Woo Kim. PanoVILD: a challenging panoramic vision, inertial and LiDAR dataset for simultaneous localization and mapping. The Journal of Supercomputing, 78 (6):8247–8267, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04198-1. See cor-

rection [1567].

Gou:2022:DAP

[1354] Chao Gou, Yuchen Zhou, and Dan Li. Driver attention prediction based on convolution and transformers. The Journal of Supercomputing, 78(6): 8268–8284, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04151-2.

Rajendran:2022:SAS

[1355] Dineshkumar Rajendran and S. Prasanna. Stealth assessment strategy in distributed systems using optimal deep learning with game based learning. The Journal of Supercomputing, 78 (6):8285–8301, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04236-y.

Amiri-Zarandi:2022:LLB

[1356] Mohammad Amiri-Zarandi, Rozita A.

Dara, and Evan Fraser. LBTM: a lightweight blockchain-based trust management system for Social Internet of Things. The Journal of Supercomputing, 78(6):8302– 8320, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04231-3.

Chen:2022:OMS

[1357] Lili Chen, Wensheng Gan, and Chien-Ming Chen. OHUQI: Mining onshelf high-utility quantitative itemsets. The Journal of Supercomputing, 78 (6):8321–8345, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04218-0.

Silva:2022:DBE

[1358] Jeydson Silva, Ronaldo Aquino, and Davidson Marques. Deep brain emotional learning-based intelligent controller applied to an inverted pendulum system. The Journal of Supercomputing, 78(6):8346– 8366, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04200-w.

Weng:2022:HFR

[1359] Chi-Yao Weng, Hao-Yu Weng, and Cheng-Ta Huang. High-fidelity reversible data hiding based on PVO and median preserving. The Journal of Supercomputing, 78(6):8367– 8388, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04226-0.

Razzaghi:2022:NST

[1360] Nasim Razzaghi and Shahram Babaie. A new selfish thing detection method based on Voronoi diagram for Internet of Things. The Journal of Supercomputing, 78(6):8389– 8408, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04202-8.

Ma:2022:ABF

[1361] Hualin Ma and Liyan Zhang. Attentionbased framework for weakly supervised video anomaly detection. The Journal of Supercomputing, 78(6): 8409–8429, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04190-9.

Khan:2022:DED

[1362] Angshuman Khan and Rajeev Arya. Design and energy dissipation analysis of simple QCA multiplexer for nanocomputing. The Journal of Supercomputing, 78(6):8430– 8444, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04191-8.

Tang:2022:OHD

[1363] Xin Tang, Wenjin Liu, and Baodan Chen. An optimized hardware design of a two-dimensional guide filter and its application in image denoising. The Journal of Supercomputing, 78(6):8445–8466, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04044-4.

Shan:2022:MPN

[1364] Chuanhui Shan, Jun Ou, and Xiumei Chen. Matrix-product neural network based on sequence block matrix product. The Journal of Supercomputing, 78(6):8467–8492, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04194-5.

Salunkhe:2022:EOC

[1365] Satish S. Salunkhe, Shelendra Pal, and Bos Mathew Jos. Energy optimization for CAN bus and media controls in electric vehicles using deep learning algorithms. The Journal of Supercomputing, 78(6): 8493–8508, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04186-5.

Yuan:2022:IBP

[1366] Yilin Yuan, Jianbiao Zhang, and Zheng Li. Identity-based public data integrity verification scheme in cloud storage system via blockchain. The Journal of Supercomputing, 78(6): 8509–8530, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04193-6.

Wu:2022:MMD

[1367] Yonghao Wu, Huyin Zhang, and Fei Yang. MVDLSTM: MultiView deep LSTM framework for online ridehailing order prediction. The Journal of Supercomputing, 78(6):8531– 8559, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04237-x.

Reddy:2022:NEH

[1368] K. Nageswara Reddy, M. Thillaikarasi, and T. Suresh. A novel elephant herd optimization model with a deep extreme learning machine for solar radiation prediction using weather forecasts. The Journal of Supercomputing, 78(6):8560–8576, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04244-y.

Shariq:2022:AAS

[1369] Mohd Shariq, Karan Singh, and David Taniar. AnonSURP: an anonymous and secure ultralightweight RFID protocol for deployment in Internet of Vehicles systems. The Journal of Supercomputing, 78(6):8577– 8602, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04232-2. See cor-

rection [1370].

Shariq:2022:CAA

[1370] Mohd Shariq, Karan Singh, and David Taniar. Correction to: Anon-SURP: an anonymous and secure ultralightweight RFID protocol for deployment in Internet of Vehicles systems. The Journal of Supercomputing, 78(6):8603, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04316-7.  See

[1369].

Ghuli:2022:ESE

[1371] Anand Ghuli, Damodar Reddy Edla, and João Manuel R. S. Tavares. Epileptic seizure endorsement technique using DWT power spectrum. The Journal of Supercomputing, 78 (6):8604–8624, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04196-3.

Alqahtani:2022:AQE

[1372] Abdullah Saleh Alqahtani, P. Saravanan, and Sami Alshmrany. An automatic query expansion based on hybrid CMO-COOT algorithm for optimized information retrieval. The Journal of Supercomputing, 78(6): 8625–8643, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04171-y.

Panahnejad:2022:ADK

[1373] Maryam Panahnejad and Meghdad Mirabi. APT-Dt-KC: advanced persistent threat detection based on kill-chain model. The Journal of Supercomputing, 78(6):8644– 8677, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04201-9.

Liu:2022:KDC

[1374] Qiang Liu and Qun Cong. Kinematic and dynamic control model of wheeled mobile robot under Internet of Things and neural network. The Journal of Supercomputing, 78 (6):8678–8707, April 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04160-1.

Naresh:2022:PSB

[1375] Vankamamidi S. Naresh, V. V. L. Divakar Allavarpu, and Sivaranjani Reddi. Provably secure blockchain privacy-preserving smart contract centric dynamic group key agreement for large WSN. The Journal of Supercomputing, 78(6):8708– 8732, April 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04175-8.

Chen:2022:DFF

[1376] Zhensen Chen, Wenyuan Yang, and Jingmin Yang. Deeply feature fused video super-resolution network using temporal grouping. The Journal of Supercomputing, 78(7):8999– 9016, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04299-x.

Fu:2022:HPI

[1377] You Fu and Wei Zhou. A heterogeneous parallel implementation of the Markov clustering algorithm for large-scale biological networks on distributed CPU-GPU clusters. The Journal of Supercomputing, 78(7): 9017–9037, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04204-6.

BenSaid:2022:MCS

[1378] Mouna Ben Said, Nader Ben Amor, and Khaled Lahbib. Multi-constraints self-adaptation for reconfigurable multimedia embedded systems. The Journal of Supercomputing, 78(7):9038– 9064, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04269-3.

Yadollahi:2022:MGI

[1379] Milad Yadollahi, Arash Shams Taleghani, and Vahid Esfahanian. Multi-grid ion thruster lifetime simulation using hybrid parallel processing. The Journal of Supercomputing, 78(7): 9065–9078, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04287-1.

Khodadad:2022:CEP

[1380] Farshid Samsami Khodadad and Hamid Noori. Characterizing energy and performance of soft-core-based homogeneous multiprocessor systems. The Journal of Supercomputing, 78 (7):9079–9101, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04273-7.

Nagendranth:2022:TIF

[1381] M. V. S. S. Nagendranth, M. Rajesh Khanna, and R. Surendran. Type II fuzzy-based clustering with improved ant colony optimizationbased routing (t2fcatr) protocol for secured data transmission in MANET. The Journal of Supercomputing, 78 (7):9102–9120, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04262-w.

Chhabra:2022:OBT

[1382] Amit Chhabra, Kuo-Chan Huang, and Tarik A. Rashid. Optimizing bag-of-tasks scheduling on cloud data centers using hybrid swarmintelligence meta-heuristic. The Journal of Supercomputing, 78(7):9121– 9183, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04199-0.

Morillo:2022:HPM

[1383] Julian Morillo, Maxime Vassaux, and Marta Garcia-Gasulla. Hybrid parallelization of molecular dynamics simulations to reduce load imbalance. The Journal of Supercomputing, 78 (7):9184–9215, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04214-4.

Poornima:2022:HPI

[1384] P. U. Poornima and R. Sridhar. Highperformance investigation of a snubber circuit-based split source inverter for nonlinear loads using PWM techniques. The Journal of Supercomputing, 78 (7):9216–9244, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04162-z.

Kim:2022:PPE

[1385] Hyeong-Jin Kim, Hyunjo Lee, and Jae-Woo Chang. Privacy-preserving k NN query processing algorithms via secure two-party computation over encrypted database in cloud computing. The Journal of Supercomputing, 78 (7):9245–9284, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04286-2.

Refaee:2022:CSI

[1386] Eshrag A. Refaee and Shermin Shamsudheen. A computing system that integrates deep learning and the Internet of Things for effective disease diagnosis in smart health care systems. The Journal of Supercomputing, 78(7):9285–9306, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04263-9.

Xiang:2022:CTM

[1387] Hua Xiang. The collection of theater music data and genre recognition under the Internet of Things and deep belief network. The Journal of Supercomputing, 78(7):9307– 9325, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04261-x.

Tian:2022:SEA

[1388] Junfeng Tian and Tianfeng Zhang. Secure and effective assured deletion scheme with orderly overwriting for cloud data. The Journal of Supercomputing, 78(7):9326– 9354, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04297-z.

Zhang:2022:CRS

[1389] Jiaqi Zhang, Ruijuan Zheng, and Qingtao Wu. A computational resources scheduling algorithm in edge cloud computing: from the energy efficiency of users’ perspective. The Journal of Supercomputing, 78(7): 9355–9376, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04146-z.

Manimurugan:2022:UBV

[1390] S. Manimurugan and Saad Almutairi. A user-based video recommendation approach using CAC filtering, PCA with LDOS-CoMoDa. The Journal of Supercomputing, 78 (7):9377–9391, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04213-5.

Wei:2022:UML

[1391] Yu-Chih Wei, Tzu-Yin Liao, and Wei-Chen Wu. Using machine learning to detect PII from attributes and supporting activities of information assets. The Journal of Supercomputing, 78(7):9392–9413, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04239-9.

Anand:2022:PPF

[1392] K. Anand, A. Vijayaraj, and M. Vijay Anand. Privacy preserving framework using Gaussian mutation based firebug optimization in cloud computing. The Journal of Supercomputing, 78(7):9414–9437, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04173-w.

Alqahtani:2022:FLI

[1393] Abdulrahman Saad Alqahtani. FSO-LSTM IDS: hybrid optimized and ensembled deep-learning networkbased intrusion detection system for smart networks. The Journal of Supercomputing, 78(7):9438– 9455, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04285-3.

Chen:2022:AAP

[1394] Zhengbo Chen, Di Wang, and Zuoning Chen. AMT: asynchronous inplace matrix transpose mechanism for Sunway many-core processor. The Journal of Supercomputing, 78(7):

9456–9474, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04282-6.

Jin:2022:PTT

[1395] Gang Jin. Player target tracking and detection in football game video using edge computing and deep learning. The Journal of Supercomputing, 78(7):9475–9491, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04274-6.

Ying:2022:EHO

[1396] Wei Ying, Ryu Fattah, and Xin Zhang. An efficient and high-order sliding mesh method for computational aeroacoustics. The Journal of Supercomputing, 78(7):9492– 9520, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04180-x.

Sharifian:2022:LNP

[1397] Zakieh Sharifian, Behrang Barekatain, and Faramarz Safi-Esfahani. LOADng-AT: a novel practical implementation of hybrid AHP-TOPSIS algorithm in reactive routing protocol for intelligent IoT-based networks. The Journal of Supercomputing, 78 (7):9521–9569, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04256-8.

Tian:2022:CEI

[1398] Zhenzhen Tian and Xinhua Wang.

Construction of enterprise innovation performance model using knowledge base and edge computing. The Journal of Supercomputing, 78(7):9570– 9594, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04211-7.

Liu:2022:UNN

[1399] Zhifang Liu. Using neural network to establish manufacture production performance forecasting in IoT environment. The Journal of Supercomputing, 78(7):9595–9618, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04210-8.

Narasimhulu:2022:NSB

[1400] K. Narasimhulu, K. T. Meena Abarna, and T. Suresh. A novel samplingbased visual topic models with computational intelligence for big social health data clustering. The Journal of Supercomputing, 78(7):9619– 9641, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04300-7.

Huang:2022:PHP

[1401] Zhi-Bin Huang, Guang-Tao Fu, and Wu-Bing Yang. A parallel highprecision critical point detection and location for large-scale 3D flow field on the GPU. The Journal of Supercomputing, 78(7):9642– 9667, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04220-6.

Kalsoom:2022:CAD

[1402] Anum Kalsoom, Muazzam Maqsood, and Seungmin Rho. A computer-aided diagnostic system for liver tumor detection using modified U-Net architecture. The Journal of Supercomputing, 78(7):9668–9690, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04266-6.

Yao:2022:USI

[1403] Yingbiao Yao, Xiaochong Kong, and Wei Feng. Uniform scheduling of interruptible garbage collection and request IO to improve performance and wear-leveling of SSDs. The Journal of Supercomputing, 78(7):9691– 9710, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04294-2.

Yang:2022:RAA

[1404] Cairu Yang and Xuzhong Zhang. Research into the application of AI robots in community home leisure interaction. The Journal of Supercomputing, 78(7):9711–9740, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04221-5.

Safaei:2022:EMC

[1405] F. Safaei, R. Akbar, and M. Moudi.

Efficient methods for computing the reliability polynomials of graphs and complex networks. The Journal of Supercomputing, 78(7):9741– 9781, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04216-2.

Lang:2022:PRB

[1406] Chen Lang, Ze Wang, and Shimin Sun.

POI recommendation based on a multiple bipartite graph network model. The Journal of Supercomputing, 78 (7):9782–9816, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04279-1.

Garcia-Molla:2022:PBT

[1407] Victor M. Garcia-Molla, Pedro Alonso-Jordá, and Ricardo García-Laguía. Parallel border tracking in binary images using GPUs. The Journal of Supercomputing, 78(7):9817– 9839, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04260-y.

Lou:2022:ADM

[1408] Wenzhong Lou, Jinkui Wang, and Dakui Wang. Application of distributed motion estimation for swarm MAVs in a GPS-restricted environment based on a wireless sensor network. The Journal of Supercomputing, 78(7):9840–9861, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04219-z.

Wan:2022:DRM

[1409] Xiaoji Wan, Hailin Li, and Yenchun Jim Wu. Dimensionality reduction for multivariate time-series data mining. The Journal of Supercomputing, 78 (7):9862–9878, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04303-4.

Wang:2022:IST

[1410] Fu-Hsing Wang and Shuo-I Wang. Independent spanning trees on WKrecursive networks and WK-recursive pyramids. The Journal of Supercomputing, 78(7):9879–9906, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-021-04296-
                                        0.

Zhou:2022:RAB

[1411] Kang Zhou, Chi Guo, and Huyin Zhang. Relational attention-based Markov logic network for visual navigation. The Journal of Supercomputing, 78(7):9907–9933, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04283-5.

Yuan:2022:EMR

[1412] Jianwen Yuan, Qinglin Zhao, and Yu-Teng Chang. Edge mining resources allocation among normal and gap blockchains using game theory.

The Journal of Supercomputing, 78 (7):9934–9951, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04249-7.

Andola:2022:SEC

[1413] Nitish Andola, Raghav Gahlot, and Shekhar Verma. Searchable encryption on the cloud: a survey. The Journal of Supercomputing, 78(7): 9952–9984, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04309-6.

Clemente:2022:AES

[1414] Danilo Clemente, Paulo Pereira, and Paulo Maciel. Availability evaluation of system service hosted in private cloud computing through hierarchical modeling process. The Journal of Supercomputing, 78(7):9985– 10024, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04217-1.

Yang:2022:RCD

[1415] Mi Yang, An Zhang, and Yunong Wang. A resource-constrained distributed task allocation method based on a two-stage coalition formation methodology for multi-UAVs. The Journal of Supercomputing, 78(7): 10025–10062, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04223-3.

Wang:2022:RDD

[1416] Chong Wang, Hao Jin, and Ke Zhou. Revocable, dynamic and decentralized data access control in cloud storage. The Journal of Supercomputing, 78 (7):10063–10087, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04277-3.

Liang:2022:TCT

[1417] Ningjing Liang, Xingjun Zhang, and Changjiang Zhang. Thou code: a triple-erasure-correcting horizontal code with optimal update complexity. The Journal of Supercomputing, 78 (7):10088–10117, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04271-9.

Perrot:2022:HSM

[1418] Gilles Perrot, Stéphane Domas, and Raphaël Couturier. How separable median filters can get better results than full 2D versions. The Journal of Supercomputing, 78(7): 10118–10148, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04233-1.

Sadr:2022:ATA

[1419] Hossein Sadr and Mojdeh Nazari Soleimandarabi. ACNN-TL: attentionbased convolutional neural network coupling with transfer learning and contextualized word representation for enhancing the performance of sentiment classification. The Journal of Supercomputing, 78(7):10149– 10175, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04208-2.

Kandan:2022:QOA

[1420] M. Kandan, Anbazhagan Krishnamurthy, and T. Tamilvizhi. Quasi oppositional Aquila optimizer-based task scheduling approach in an IoT enabled cloud environment. The Journal of Supercomputing, 78(7):10176– 10190, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04311-y.

Ashraf:2022:SOU

[1421] Asiya Ashraf, Zeshan Iqbal, and Sang oh Park. Scalable offloading using machine learning methods for distributed multi-controller architecture of SDN networks. The Journal of Supercomputing, 78(7):10191– 10210, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04313-w.

Mehdi:2022:FEM

[1422] Hamid Mehdi, Houman Zarrabi, and Amir Masoud Rahmani. A flexible energy management approach for smart healthcare on the internet of patients (IoP). The Journal of Supercomputing, 78(7):10211– 10249, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04240-2.

Hassan:2022:FGA

[1423] Md Rafiul Hassan, Walaa N. Ismail, and Mohammad Mehedi Hassan. A framework of genetic algorithm-based CNN on multi-access edge computing for automated detection of COVID-19. The Journal of Supercomputing, 78 (7):10250–10274, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04222-4. See cor-

rection [2175].

Sattar:2022:SDL

[1424] Naw Safrin Sattar and Shaikh Arifuzzaman. Scalable distributed Louvain algorithm for community detection in large graphs. The Journal of Supercomputing, 78(7):10275– 10309, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04224-2.

Tan:2022:SAM

[1425] Shih-Chin Tan, Chun-Liang Lai, and Jia-Jeng Hou. Scenario analysis on medical treatments of patients with knee osteoarthritis. The Journal of Supercomputing, 78(7):10310– 10325, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04206-4.

Patel:2022:HCL

[1426] Eva Patel and Dharmender Singh Kushwaha. A hybrid CNN-LSTM model for predicting server load in cloud computing. The Journal of Supercomputing, 78(8):1– 30, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04234-0.

Shenbagalakshmi:2022:AML

[1427] V. Shenbagalakshmi and T. Jaya.

Application of machine learning and IoT to enable child safety at home environment. The Journal of Supercomputing, 78(8):10357– 10384, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04310-z.

Yang:2022:IAB

[1428] Shengli Yang, Haoliang Wang, and Qiang Chen. An improved algorithm based on deep learning network for road image redundancy removal. The Journal of Supercomputing, 78 (8):10385–10404, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04302-5.

Mohanta:2022:ALA

[1429] Tapan Kumar Mohanta and Dushmanta Kumar Das. Advanced localization algorithm for wireless sensor networks using fractional order class topper optimization. The Journal of Supercomputing, 78(8):10405– 10433, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04278-2.

Martinez:2022:HOS

[1430] Millán A. Martínez, Basilio B. Fraguela, and José C. Cabaleiro. A highly optimized skeleton for unbalanced and deep divide-and-conquer algorithms on multi-core clusters. The Journal of Supercomputing, 78(8): 10434–10454, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04259-5.

Mahdavimoghadam:2022:IRL

[1431] Mahnoosh Mahdavimoghadam, Amin Nikanjam, and Monireh Abdoos. Improved reinforcement learning in cooperative multi-agent environments using knowledge transfer. The Journal of Supercomputing, 78(8):10455– 10479, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04305-w.

Senagi:2022:PCR

[1432] Kennedy Senagi and Nicolas Jouandeau. Parallel construction of Random Forest on GPU. The Journal of Supercomputing, 78(8):10480– 10500, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04290-6.

Wang:2022:EII

[1433] Tao Wang. Exploring intelligent image recognition technology of football robot using omnidirectional vision of Internet of Things. The Journal of Supercomputing, 78(8): 10501–10520, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04314-9.

Zhang:2022:ABC

[1434] Ruizhi Zhang. Analyzing body changes of high-level dance movements through biological image visualization technology by convolutional neural network. The Journal of Supercomputing, 78 (8):10521–10541, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04298-y.

Karaata:2022:RCM

[1435] Mehmet Hakan Karaata, Aysha Dabees, and Fawaz Alazemi. A reliable concurrent multicast algorithm for content distribution. The Journal of Supercomputing, 78(8):10542– 10574, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04291-5.

Vinoth:2022:IML

[1436] D. Vinoth and P. Prabhavathy. An intelligent machine learning-based sarcasm detection and classification model on social networks. The Journal of Supercomputing, 78(8):10575– 10594, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.

                                        1007/s11227-022-04312-x. See cor-

rection [2174].

Sharma:2022:EHE

[1437] Yanshul Sharma, Shounak Chakraborty, and Sanjay Moulik. ETA-HP: an energy and temperature-aware realtime scheduler for heterogeneous platforms. The Journal of Supercomputing, 78(8):1–25, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04257-7.

Noueihed:2022:KBV

[1438] Hamza Noueihed, Heba Harb, and Joe Tekli. Knowledge-based virtual outdoor weather event simulator using unity 3D. The Journal of Supercomputing, 78(8):10620– 10655, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04212-6.

Elawady:2022:TMR

[1439] Mohamed Elawady, Amany Sarhan, and Mahmoud A. M. Alshewimy. Toward a mixed reality domain model for time-sensitive applications using IoE infrastructure and edge computing (MRIoEF). The Journal of Supercomputing, 78(8):10656– 10689, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04307-8.

Alashti:2022:PAC

[1440] Amin Heydari Alashti, Ahmad Asgharian Rezaei, and Mohammad Ghodsi. Parsisanj: an automatic component-based approach toward search engine evaluation. The Journal of Supercomputing, 78(8):10690– 10711, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04306-9.

Baek:2022:DCB

[1441] Seungmin Baek, Eunhye Jo, and Yunyoung Nam. Development of a color block play device for child attentional capability test. The Journal of Supercomputing, 78(8):10712– 10724, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04301-6.

Ikram:2022:PIT

[1442] Sumaiya Thaseen Ikram, V. Priya, and Achyut Shankar. Prediction of IIoT traffic using a modified whale optimization approach integrated with random forest classifier. The Journal of Supercomputing, 78(8):10725– 10756, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04284-4.

Zhang:2022:MFI

[1443] Longqing Zhang, Liping Bai, and Yanghong Zhang. Multi-factor indicator of THIC intelligent lighting system with BP neural network. The Journal of Supercomputing, 78(8): 10757–10771, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04289-z.

Arasi:2022:AAB

[1444] V. Ezhil Arasi, K. Indra Gandhi, and K. Kulothungan. Auditable attribute-based data access control using blockchain in cloud storage. The Journal of Supercomputing, 78 (8):10772–10798, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04293-3.

Cai:2022:MCA

[1445] Weiwei Cai, Fazhi He, and Xiao Lv. Multi-core accelerated CRDT for large-scale and dynamic collaboration. The Journal of Supercomputing, 78 (8):10799–10828, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04308-7.

Zhang:2022:COO

[1446] Jianshan Zhang, Ming Li, and Bing Lin. Computation offloading for objectoriented applications in a UAV-based edge-cloud environment. The Journal of Supercomputing, 78(8):10829– 10853, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04288-0.

Sarrafzade:2022:GBA

[1447] Nazanin Sarrafzade, Reza Entezari-Maleki, and Leonel Sousa. A geneticbased approach for service placement in fog computing. The Journal of Supercomputing, 78(8):10854– 10875, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04254-w.

Xu:2022:AAE

[1448] Jia-Lang Xu and Ying-Lin Hsu. Analysis of agricultural exports based on deep learning and text mining. The Journal of Supercomputing, 78(8): 10876–10892, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04238-w.

Gao:2022:UNU

[1449] Ming Gao and YuBin Lu. URAP: a new ultra-lightweight RFID authentication protocol in passive RFID system. The Journal of Supercomputing, 78 (8):10893–10905, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04252-y.

Mahjoub:2022:MCD

[1450] Alireza Mahjoub, Fatemeh Vardi, and Roya Rad. A multi-criteria decision based on adaptive routing algorithms with discrete operators for on-chip networks. The Journal of Supercomputing, 78(8):10906– 10929, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04272-8.

Wang:2022:FWF

[1451] Yongwei Wang, Jingbo Wu, and Xin Jin. Flux-weakening fuzzy adaptive ST-SMO sensorless control algorithm for PMSM in EV. The Journal of Supercomputing, 78(8): 10930–10949, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04264-8.

Zhang:2022:CAS

[1452] Yuedong Zhang and Yuanbin Mo. Chaotic adaptive sailfish optimizer with genetic characteristics for global optimization. The Journal of Supercomputing, 78(8):10950–10996, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-021-04255-
                                        9.

Wang:2022:CPL

[1453] Min Wang and Zhihan Lv. Construction of personalized learning and knowledge system of chemistry specialty via the Internet of Things and clustering algorithm. The Journal of Supercomputing, 78(8):10997– 11014, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04315-8.

Banerjee:2022:CEM

[1454] Avishek Banerjee, Sudip Kumar De, and Samiran Chattopadhyay. Construction of energy minimized WSN using GA-SAMP-MWPSO and Kmean clustering algorithm with LDCF deployment strategy. The Journal of Supercomputing, 78(8):11015– 11050, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04265-7.

Tran:2022:BME

[1455] Duy Thanh Tran and Jun-Ho Huh.

Building a model to exploit association rules and analyze purchasing behavior based on rough set theory. The Journal of Supercomputing, 78 (8):11051–11091, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04275-5.

Ma:2022:STD

[1456] Chenyu Ma, Jinfang Jia, and Xiaoying Wang. Simulation of threedimensional phase field model with LBM method using OpenCL. The Journal of Supercomputing, 78(8): 11092–11110, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04321-w.

Bang:2022:DEN

[1457] A. O. Bang and Udai Pratap Rao. Design and evaluation of a novel whitebox encryption scheme for resourceconstrained IoT devices. The Journal of Supercomputing, 78(8):11111– 11137, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04322-9.

Mas:2022:QTM

[1458] Lluís Mas, Jordi Vilaplana, and Francesc Solsona. A queuing theory model for fog computing. The Journal of Supercomputing, 78(8): 11138–11155, May 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04328-3.

Albuquerque:2022:RST

[1459] Otavio de Paula Albuquerque, Marcelo Fantinato, and Muhammad Umair Shah. Recommendations for a smart toy parental control tool. The Journal of Supercomputing, 78(8):11156– 11194, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04319-4.

Jaiswal:2022:FHH

[1460] Kavita Jaiswal and Veena Anand. FAGWO-H: a hybrid method towards fault-tolerant cluster-based routing in wireless sensor network for IoT applications. The Journal of Supercomputing, 78(8):11195– 11227, May 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04333-6.

Xia:2022:DLB

[1461] Zhiyong Xia, Liping Zhang, and Jihong Tu. Deep learning-based hysteroscopic intelligent examination and ultrasound examination for diagnosis of endometrial carcinoma. The Journal of Supercomputing, 78(9):

11229–11244, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04046-2.

Zhang:2022:DLC

[1462] Xiaoyan Zhang, Xiaoyan Zhu, and Zhiqiang Xie. Deep learning in conebeam computed tomography image segmentation for the diagnosis and treatment of acute pulpitis. The Journal of Supercomputing, 78(9): 11245–11264, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04048-0.

Luo:2022:DSC

[1463] Pan Luo and Jinling Li. Dual-source computed tomography image information under deep learning algorithm in evaluation of coronary artery lesion in children with Kawasaki disease. The Journal of Supercomputing, 78 (9):11265–11282, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04077-9.

Chen:2022:DLM

[1464] Lieguang Chen, Ying Lu, and Xianxu Zhuang. Deep learning in molecular biology marker recognition of patients with acute myeloid leukemia. The Journal of Supercomputing, 78 (9):11283–11297, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04104-9.

Lv:2022:CEC

[1465] Ran Lv, Yinpu Tian, and Chenzhi Liu. Clinical efficacy of computed tomography-guided microwave ablation sequential transcatheter arterial chemoembolization under edge calculation for primary liver cancer over 5 cm in diameter. The Journal of Supercomputing, 78(9):11298– 11317, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04118-3.

Cao:2022:ODR

[1466] Lu Cao and Yuling Liu. Optimization design and research of simulation system for urban green ecological rainwater by genetic algorithm. The Journal of Supercomputing, 78 (9):11318–11344, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04192-7.

Bermejo:2022:GME

[1467] Belen Bermejo and Carlos Juiz. A general method for evaluating the overhead when consolidating servers: performance degradation in virtual machines and containers. The Journal of Supercomputing, 78(9):11345– 11372, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04318-5.

Calp:2022:IBD

[1468] Muhammed Hanefi Calp, Resul Butuner, and David Camacho. IoHTbased deep learning controlled robot vehicle for paralyzed patients of smart cities. The Journal of Supercomputing, 78(9):11373–11408, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-021-04292-
4.

Jude:2022:OBC

[1469] Dylan Jude, Jayanarayanan Sitaraman, and Andrew Wissink. An octree-based, cartesian Navier–Stokes solver for modern cluster architectures. The Journal of Supercomputing, 78 (9):11409–11440, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04324-7.

Tian:2022:SHS

[1470] Min Tian, Junjie Wang, and Tao Liu. swSuperLU: a highly scalable sparse direct solver on Sunway manycore architecture. The Journal of Supercomputing, 78(9):11441– 11463, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04270-w.

Liu:2022:AAM

[1471] Xiaoyan Liu, Yi Liu, and Depei Qian. Accelerating approximate matrix multiplication for near-sparse matrices on GPUs. The Journal of Supercomputing, 78(9):11464– 11491, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04334-5.

Jain:2022:SED

[1472] Shubhra Jain, Rahul Kumar Verma, and Anupam Shukla. A survey on event-driven and query-driven hierarchical routing protocols for mobile sink-based wireless sensor networks. The Journal of Supercomputing, 78 (9):11492–11538, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04327-4.

Stokfiszewski:2022:EIO

[1473] Kamil Stokfiszewski, Kamil Wieloch, and Mykhaylo Yatsymirskyy. An efficient implementation of onedimensional discrete wavelet transform algorithms for GPU architectures. The Journal of Supercomputing, 78 (9):11539–11563, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04331-8.

Majeed:2022:HPA

[1474] Ali Majeed and Esam Alkaldy. Highperformance adder using a new XOR gate in QCA technology. The Journal of Supercomputing, 78(9):11564– 11579, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04339-0.

Wang:2022:AFA

[1475] Yi Wang and Shuran Song. An adaptive firefly algorithm for multilevel image thresholding based on minimum cross-entropy. The Journal of Supercomputing, 78(9):11580– 11600, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04281-7.

Jiang:2022:ACO

[1476] Feng Jiang, Gang Wang, and Yue Wu.

Application of Canny operator threshold adaptive segmentation algorithm combined with digital image processing in tunnel face crevice extraction. The Journal of Supercomputing, 78 (9):11601–11620, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04330-9.

Zeng:2022:IME

[1477] Hui Zeng. Influences of mobile edge computing-based service preloading on the early-warning of financial risks. The Journal of Supercomputing, 78 (9):11621–11639, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04329-2.

Abitha:2022:EMO

[1478] R. Abitha, S. Mary Vennila, and I. Mohamed Zaheer. Evolutionary multi-objective optimization of artificial neural network for classification of autism spectrum disorder screening. The Journal of Supercomputing, 78 (9):11640–11656, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04268-4.

Mompean:2022:EED

[1479] Juan Mompeán, Juan L. Aragón, and Pablo Artal. Energy-efficient design of a presbyopia correction wearable powered by mobile GPUs and FPGAs. The Journal of Supercomputing, 78 (9):11657–11679, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04332-7.

Wang:2022:IIM

[1480] Dongmei Wang, Yiwen Liang, and Sai Liu. Innate immune memory and its application to artificial immune systems. The Journal of Supercomputing, 78(9):11680– 11701, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04295-1.

Ghose:2022:FFG

[1481] Anirban Ghose and Soumyajit Dey. FGFS: Feature Guided Frontier Scheduling for SIMT DAGs. The Journal of Supercomputing, 78(9): 11702–11743, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04323-8.

Fanfakh:2022:OGO

[1482] Ahmed Fanfakh, Hassan Noura, and Raphaël Couturier. ORSCA-GPU: one round stream cipher algorithm for GPU implementation. The Journal of Supercomputing, 78(9):11744– 11767, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04335-4.

Tian:2022:SPP

[1483] Liwei Tian, Li Feng, and Yuankai Guo. Stock price prediction based on LSTM and LightGBM hybrid model. The Journal of Supercomputing, 78 (9):11768–11793, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04326-5.

Makhadmeh:2022:HMV

[1484] Sharif Naser Makhadmeh, Ammar Kamal Abasi, and Mohammed Azmi Al-Betar. Hybrid multi-verse optimizer with grey wolf optimizer for power scheduling problem in smart home using IoT. The Journal of Supercomputing, 78(9):11794– 11829, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04325-6.

Mokhtari:2022:HCC

[1485] Sadaf Mokhtari, Hamid Barati, and Ali Barati. A hierarchical congestion control method in clustered Internet of Things. The Journal of Supercomputing, 78(9):11830– 11855, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04340-7.

Venkatesh:2022:RNP

[1486] K. Venkatesh and D. Narasimhan.

Revealing the novel precise subset identification and deduplication of audio substance over the shared public environment. The Journal of Supercomputing, 78(9):11856– 11872, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04317-6.

Jiang:2022:NFD

[1487] Fei Jiang, Xiao-Ya Ma, and Jin Tong. A new form of deep learning in smart logistics with IoT environment. The Journal of Supercomputing, 78 (9):11873–11894, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04343-4.

Zheng:2022:CSA

[1488] Huijun Zheng, Yong Peng, and Yeh-Cheng Chen. Course scheduling algorithm based on improved binary cuckoo search. The Journal of Supercomputing, 78(9):11895– 11920, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04341-6.

Padhy:2022:CCA

[1489] Satyajit Padhy, Ming-Han Tsai, and Jerry Chou. CAMIRA: a consolidationaware migration avoidance job scheduling strategy for virtualized parallel computing clusters. The Journal of Supercomputing, 78(9):11921– 11948, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04337-2.

Xia:2022:PAE

[1490] Xiaoyun Xia, Zhengxin Huang, and Yi Xiang. Performance analysis of evolutionary algorithm for the maximum internal spanning tree problem. The Journal of Supercomputing, 78 (9):11949–11973, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04342-5.

Long:2022:CUR

[1491] Leijin Long, Feng He, and Hongjiang Liu. Correction to: The use of remote sensing satellite using deep learning in emergency monitoring of high-level landslides disaster in Jinsha River. The Journal of Supercomputing, 78(9):11974, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04353-2. See [853].

Mohar:2022:LSN

[1492] Satinder Singh Mohar, Sonia Goyal, and Ranjit Kaur. Localization of sensor nodes in wireless sensor networks using bat optimization algorithm with enhanced exploration and exploitation characteristics. The Journal of Supercomputing, 78(9):11975– 12023, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04320-x.

Ahrabi:2022:EPD

[1493] Sima Sarv Ahrabi, Lorenzo Piazzo, and Enzo Baccarelli. Exploiting probability density function of deep convolutional autoencoders’ latent space for reliable COVID-19 detection on CT scans. The Journal of Supercomputing, 78 (9):12024–12045, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04349-y.

Wang:2022:ETL

[1494] Ziheng Wang, Heng Chen, and Xingjun Zhang. Extending τ-Lop to model MPI blocking primitives on shared memory. The Journal of Supercomputing, 78 (9):12046–12069, June 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04352-3.

Du:2022:IBR

[1495] Erle Du. Impact of bank research and development on total factor productivity and performance evaluation by RBF network. The Journal of Supercomputing, 78(9):12070– 12092, June 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04358-x.

Braeken:2022:AKA

[1496] An Braeken. Authenticated key agreement protocols for dew-assisted IoT systems. The Journal of Supercomputing, 78(10):12093–12113, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04364-
z.

Hussain:2022:IST

[1497] Zaid Hussain, Hosam AboElFotoh, and Bader AlBdaiwi. Independent spanning trees in Eisenstein– Jacobi networks. The Journal of Supercomputing, 78(10):12114– 12135, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04351-4.

Aldag:2022:AMQ

[1498] Mehmet Aldag, Yönal Kirsal, and Sadik Ülker. An analytical modelling and QoS evaluation of faulttolerant load balancer and web servers in fog computing. The Journal of Supercomputing, 78(10):12136– 12158, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04345-2.

Song:2022:IDC

[1499] Chi hoon Song and Young woo Sohn. The influence of dependability in cloud computing adoption. The Journal of Supercomputing, 78(10): 12159–12201, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04346-1.

Park:2022:QCP

[1500] Ju-Won Park, Min-Woo Kwon, and Taeyoung Hong. Queue congestion prediction for large-scale high performance computing systems using a hidden Markov model. The Journal of Supercomputing, 78(10): 12202–12223, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04356-z.

Lupion:2022:UMG

[1501] M. Lupión, J. F. Sanjuan, and P. M. Ortigosa. Using a multi-GPU node to accelerate the training of Pix2Pix neural networks. The Journal of Supercomputing, 78(10): 12224–12241, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04354-1.

Guo:2022:TAM

[1502] Ping Guo and Minliang Jiang. TSMSA: a 2DSPP algorithm with multistrategy rectangle selection. The Journal of Supercomputing, 78(10): 12242–12277, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04350-5.

Alweshah:2022:IDI

[1503] Mohammed Alweshah, Saleh Alkhalaileh, and Salwani Abdullah. Intrusion detection for IoT based on a hybrid shuffled shepherd optimization algorithm. The Journal of Supercomputing, 78(10):12278– 12309, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04357-y.

Cano-Cano:2022:PQS

[1504] Javier Cano-Cano, Francisco J.

Andújar, and Gaspar Mora. Providing quality of service in omni-path networks. The Journal of Supercomputing, 78(10):12310–12343, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04365-
y.

Dokulil:2022:OVE

[1505] Jiri Dokulil and Siegfried Benkner. The OCR-Vx experience: lessons learned from designing and implementing a task-based runtime system. The Journal of Supercomputing, 78(10): 12344–12379, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04355-0.

An:2022:DMB

[1506] MingShou An and Dae-Seong Kang.

The distance measurement based on corner detection for rebar spacing in engineering images. The Journal of Supercomputing, 78(10):12380– 12393, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04304-x.

Maniraj:2022:HDL

[1507] S. P. Maniraj and P. Sardar Maran.

A hybrid deep learning approach for skin cancer diagnosis using subband fusion of 3D wavelets. The Journal of Supercomputing, 78(10): 12394–12409, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04371-0.

Ye:2022:FFL

[1508] Qing Ye, Yuhao Zhou, and Jiancheng Lv. FLSGD: free local SGD with parallel synchronization. The Journal of Supercomputing, 78(10):12410– 12433, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-021-04267-5. See cor-

rection [1509].

Ye:2022:CFF

[1509] Qing Ye, Yuhao Zhou, and Jiancheng Lv. Correction to: FLSGD: free local SGD with parallel synchronization. The Journal of Supercomputing, 78(10):12434, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04440-4.  See

[1508].

Chang:2022:AIA

[1510] Yu-Teng Chang, Hui-Ru Yang, and Chien-Ming Chen. Analysis on improving the application of machine learning in product development. The Journal of Supercomputing, 78(10): 12435–12460, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04344-3.

Ye:2022:IBA

[1511] Yun Ye, Xiaojuan Zhao, and Lu Xiong. An improved bat algorithm with velocity weight and curve decreasing. The Journal of Supercomputing, 78 (10):12461–12475, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04368-9.

Zheng:2022:DAC

[1512] F. Yifeng Zheng, S. Lei Huang, and S. Chai Kiat Yeo. Distributed algorithm for computation offloading in mobile edge computing considering user mobility and task randomness. The Journal of Supercomputing, 78 (10):12476–12504, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04383-w.

Melo:2022:PAE

[1513] Carlos Melo, Felipe Oliveira, and Paulo Maciel. Performance and availability evaluation of the blockchain platform hyperledger fabric. The Journal of Supercomputing, 78(10): 12505–12527, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04361-2.

Chen:2022:BGP

[1514] Meihong Chen and Fuzhi Su. A basketball game prediction system based on artificial intelligence. The Journal of Supercomputing, 78(10): 12528–12552, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04375-w.

Praharenka:2022:VDC

[1515] Wyatt Praharenka, David Pankratz, and José Nelson Amaral. Vectorizing divergent control flow with active-lane consolidation on longvector architectures. The Journal of Supercomputing, 78(10):12553– 12588, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04359-w.

Packiaraj:2022:HFD

[1516] Muneeswaran Packiaraj and Sriram Kailasam. HyPar-FCA: a distributed framework based on hybrid partitioning for FCA. The Journal of Supercomputing, 78(10):12589– 12620, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04366-x.

Li:2022:WSR

[1517] Xiaoyu Li, Xiuguo Zhang, and Zhiying Cao. Web services recommendation based on Metapath-guided graph attention network. The Journal of Supercomputing, 78(10):12621– 12647, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04369-8.

Wang:2022:TTA

[1518] Yanni Wang, Rongchun Guo, and Suwen Zhao. Target tracking algorithm based on multiscale analysis and combinatorial matching. The Journal of Supercomputing, 78(10):

12648–12661, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04391-w.

Lakzaei:2022:JCR

[1519] Mahmood Lakzaei, Vahid Sattari-Naeini, and Amir Javadpour. A joint computational and resource allocation model for fast parallel data processing in fog computing. The Journal of Supercomputing, 78(10): 12662–12685, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04374-x.

Hu:2022:ABC

[1520] Xiaojian Hu, Tong Liu, and Chenxi Lin. Attention-based Conv-LSTM and Bi-LSTM networks for largescale traffic speed prediction. The Journal of Supercomputing, 78(10): 12686–12709, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04386-7.

Chaudhary:2022:SIA

[1521] Vidhi Chaudhary, Preetpal Kaur Buttar, and Manoj Kumar Sachan. Satellite imagery analysis for road segmentation using U-Net architecture. The Journal of Supercomputing, 78 (10):12710–12725, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04379-6.

Fani:2022:RHT

[1522] Reza Fani and Morteza Saheb Zamani. Runtime hardware Trojan detection by reconfigurable monitoring circuits. The Journal of Supercomputing, 78 (10):12726–12752, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04362-1.

Veni:2022:HPV

[1523] N. Veni and J. Manjula. Highperformance visual geometric group deep learning architectures for MRI brain tumor classification. The Journal of Supercomputing, 78(10):12753– 12764, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04384-9.

Farooq:2022:PPW

[1524] Muhammad Umar Farooq, Xingfu Wang, and Bushra Qureshi. POWER: probabilistic weight-based energyefficient cluster routing for largescale wireless sensor networks. The Journal of Supercomputing, 78(10): 12765–12791, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04372-z.

Meshram:2022:ERU

[1525] Chandrashekhar Meshram, Rabha W. Ibrahim, and Sharad Kumar Barve. An efficient remote user authentication with key agreement procedure based on convolution-Chebyshev chaotic maps using biometric. The Journal of Supercomputing, 78(10): 12792–12814, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-021-04280-8.

Qi:2022:PSP

[1526] Mingping Qi and Jianhua Chen.

Provably secure post-quantum authenticated key exchange from supersingular isogenies. The Journal of Supercomputing, 78(10):12815– 12833, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04378-7.

Ahanger:2022:CDM

[1527] Tariq Ahamed Ahanger, Abdullah Alqahtani, and Abdullah Algashami. Cognitive decision-making in smart police industry. The Journal of Supercomputing, 78(10):12834– 12860, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04392-9.

Liu:2022:IMI

[1528] Suping Liu, Jibo Zhang, and Wenlong Su. An improved method of identifying learner’s behaviors based on deep learning. The Journal of Supercomputing, 78(10):12861– 12872, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04402-w.

Balakrishna:2022:DAT

[1529] Sivadi Balakrishna. 4D-ACSM: a technique for dynamically assigning and adjusting cluster patterns for IoT data analysis. The Journal of Supercomputing, 78(10):12873– 12897, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04427-1.

Vinicius:2022:DPA

[1530] Lucas Vinícius, Laécio Rodrigues, and Francisco Airton Silva. Docker platform aging: a systematic performance evaluation and prediction of resource consumption. The Journal of Supercomputing, 78(10):12898– 12928, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04389-4.

Dolbeau:2022:CTP

[1531] Romain Dolbeau. Correction to: Theoretical peak FLOPS per instruction set: a tutorial. The Journal of Supercomputing, 78(10): 12929, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04443-1. See [1].

Hamid:2022:CAT

[1532] Laiba Hamid, Asmara Jadoon, and Hassan Asghar. Comparative analysis of task level heuristic scheduling algorithms in cloud computing. The Journal of Supercomputing, 78 (11):12931–12949, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04382-x.

Liang:2022:RLR

[1533] Hao Liang, Li Kang, and Jianjun Huang. A robust low-rank matrix completion based on truncated nuclear norm and Lp-norm. The Journal of Supercomputing, 78(11): 12950–12972, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04385-8.

Liu:2022:IPP

[1534] Tianbo Liu and Jindong Zhang. An improved path planning algorithm based on fuel consumption. The Journal of Supercomputing, 78(11): 12973–13003, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04395-6.

Mechouche:2022:CCA

[1535] Jeremy Mechouche, Roua Touihri, and Walid Gaaloul. Conformance checking for autonomous multi-cloud SLA management and adaptation. The Journal of Supercomputing, 78(11): 13004–13039, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04363-0.

Niu:2022:DOL

[1536] Yanbiao Niu, Xuefeng Yan, and Yanzhao Niu. Dynamic opposite learning enhanced artificial ecosystem optimizer for IIR system identification. The Journal of Supercomputing, 78 (11):13040–13085, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04367-w.

Srinivasavarma:2022:HBM

[1537] Vegesna S. M. Srinivasavarma, Shanmukha Rao Pydi, and S. Noor Mahammad. Hardware-based multimatch packet classification in NIDS: an overview and novel extensions for improving the energy efficiency of TCAM-based classifiers. The Journal of Supercomputing, 78(11):13086– 13121, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04377-8.

Li:2022:SSM

[1538] Jieling Li, Hao Zhang, and Zhihuang Liu. Semi-supervised machine learning framework for network intrusion detection. The Journal of Supercomputing, 78(11):13122– 13144, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04390-x.

Kansal:2022:CRM

[1539] Puneet Kansal, Manoj Kumar, and Om Prakash Verma. Classification of resource management approaches in fog/edge paradigm and future research prospects: a systematic review. The Journal of Supercomputing, 78 (11):13145–13204, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04338-1.

Mojez:2022:SCA

[1540] Hadi Mojez, Amir Massoud Bidgoli, and Hamid Haj Seyyed Javadi. Star capacity-aware latency-based next controller placement problem with considering single controller failure in software-defined wide-area networks. The Journal of Supercomputing, 78 (11):13205–13244, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04360-3.

Kapralos:2022:TPE

[1541] Michael Kapralos. Taylor polynomials as an estimator for certain Toeplitz matrices. The Journal of Supercomputing, 78(11):13245– 13275, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04373-y.

Li:2022:TOO

[1542] Wenhao Li, Zhan Zhang, and Tianming Liu. Toward optimal operator parallelism for stream processing topology with limited buffers. The Journal of Supercomputing, 78(11): 13276–13297, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04376-9.

Zhu:2022:CPS

[1543] Changpeng Zhu, Bo Han, and Yinliang Zhao. A comparative performance study of spark on Kubernetes. The Journal of Supercomputing, 78 (11):13298–13322, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04381-y.

Orts:2022:ITE

[1544] Francisco Orts, Gloria Ortega, and Ester M. Garzón. Implementation of three efficient 4-digit fault-tolerant quantum carry lookahead adders. The Journal of Supercomputing, 78(11): 13323–13341, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04401-x.

Mahdavi:2022:DBM

[1545] Nooshin Mahdavi, Farhad Razaghian, and Hamed Farbeh. Data block manipulation for error rate reduction in STT-MRAM based main memory. The Journal of Supercomputing, 78(11): 13342–13372, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04394-7.

Modi:2022:IRU

[1546] Tejas M. Modi and Pravati Swain. Intelligent routing using convolutional neural network in software-defined data center network. The Journal of Supercomputing, 78(11):13373– 13392, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04348-z.

Yang:2022:BGO

[1547] Zhiwei Yang, Lu Lu, and Ruimin Wang. A batched GEMM optimization framework for deep learning. The Journal of Supercomputing, 78(11): 13393–13408, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04336-3.

Rai:2022:EDF

[1548] Mritunjay Rai, Tanmoy Maity, and R. K. Yadav. Early detection of foot ulceration in type II diabetic patient using registration method in infrared images and descriptive comparison with deep learning methods. The Journal of Supercomputing, 78 (11):13409–13426, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04380-z.

Guzel:2022:FEA

[1549] Metehan Guzel and Suat Ozdemir. Fair and energy-aware IoT service composition under QoS constraints. The Journal of Supercomputing, 78(11): 13427–13454, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04398-3.

Moreno-Alvarez:2022:HGC

[1550] Sergio Moreno-Álvarez, Mercedes E.

Paoletti, and Juan M. Haut. Heterogeneous gradient computing optimization for scalable deep neural networks. The Journal of Supercomputing, 78 (11):13455–13469, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04399-2.

Xia:2022:AED

[1551] Limin Xia and Changhong Wei. Abnormal event detection in surveillance videos based on multi-scale feature and channel-wise attention mechanism. The Journal of Supercomputing, 78 (11):13470–13490, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04410-w.

Huang:2022:OCG

[1552] Xiaolong Huang, Alexandre F. Ramos, and Yuefan Deng. Optimal circulant graphs as low-latency network topologies. The Journal of Supercomputing, 78(11):13491– 13510, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04396-5.

Cheng:2022:SSS

[1553] Zongmao Cheng, Hao Tan, and Xiao Fu. A scheduling scheme for stochastic event capture based on Bayes statistical method. The Journal of Supercomputing, 78(11):13511– 13529, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04403-9.

Aghbari:2022:ACA

[1554] Zaher Al Aghbari, Ahmed M. Khedr, and Pravija P. V. Raj. An adaptive coverage aware data gathering scheme using KD-tree and ACO for WSNs with mobile sink. The Journal of Supercomputing, 78(11):13530– 13553, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04407-5.

Kim:2022:INB

[1555] HyunJin Kim and Taeshik Shon.

Industrial network-based behavioral anomaly detection in AI-enabled smart manufacturing. The Journal of Supercomputing, 78(11):13554– 13563, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04408-4.

Shuai:2022:VCA

[1556] Xin Shuai, Linbo Qing, and Xiaohai He. A video compression artifact reduction approach combined with quantization parameters estimation. The Journal of Supercomputing, 78 (11):13564–13582, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04412-8.

Hjouji:2022:HFO

[1557] Amal Hjouji and Jaouad El-Mekkaoui.

Helmet–Fourier orthogonal moments for image representation and recognition. The Journal of Supercomputing, 78(11):13583–13623, July 2022.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04414-
6.

Jain:2022:SSC

[1558] Khushboo Jain, Pawan Singh Mehra, and Arun Agarwal. SCADA: scalable cluster-based data aggregation technique for improving network lifetime of wireless sensor networks. The Journal of Supercomputing, 78(11): 13624–13652, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04419-1.

Huang:2022:OMC

[1559] En-Ming Huang and Jerry Chou. Optimization of multi-class 0/1 knapsack problem on GPUs by improving memory access efficiency. The Journal of Supercomputing, 78(11): 13653–13679, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04425-3.

Ahmed:2022:TTL

[1560] Ismail Hadj Ahmed, Abdelghani Djebbari, and Lotfi Senhadji. Telemedical transport layer security based platform for cardiac arrhythmia classification using quadratic time-frequency analysis of HRV signal. The Journal of Supercomputing, 78(11):13680– 13709, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04387-6.

Nimmagadda:2022:ERM

[1561] Ritvik Nimmagadda, Kritika Arora, and Miguel Vargas Martin. Emotion recognition models for companion robots. The Journal of Supercomputing, 78(11):13710–13727, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04416-
4.

Kjorveziroski:2022:KDE

[1562] Vojdan Kjorveziroski and Sonja Filiposka. Kubernetes distributions for the edge: serverless performance evaluation. The Journal of Supercomputing, 78(11):13728– 13755, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04430-6.

Ghaferi:2022:CML

[1563] Esmaeil Ghaferi, Razieh Malekhosseini, and Karamollah Bagherifard. A clustering method for locating services based on fog computing for the Internet of Things. The Journal of Supercomputing, 78(11):13756– 13779, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04393-8.

Miao:2022:HAI

[1564] Rui Miao, Hongxu Jiang, and Xiaobin Li. High anti-interference and FPGA-oriented method for realtime ship detection based on structured LBP features. The Journal of Supercomputing, 78(11):13780– 13813, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04400-y.

Martinez:2022:HCP

[1565] Pablo Antonio Martínez, Gregorio Bernabé, and José Manuel García. HDNN: a cross-platform MLIR dialect for deep neural networks. The Journal of Supercomputing, 78(11): 13814–13830, July 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04417-3.

Shanker:2022:CBM

[1566] Ravi Shanker and Mahua Bhattacharya. Classification of brain MR images using modified version of simplified pulse-coupled neural network and linear programming twin support vector machines. The Journal of Supercomputing, 78(11):13831– 13863, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04420-8.

Javed:2022:PCPb

[1567] Zeeshan Javed and Gon-Woo Kim. Publisher correction to: PanoVILD: a challenging panoramic vision, inertial and LiDAR dataset for simultaneous localization and mapping. The Journal of Supercomputing, 78(11): 13864, July 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04640-y.  See

[1353].

Zheng:2022:DOS

[1568] Meiguang Zheng, Dongbang Mao, and Zhigang Hu. DOSP: an optimal synchronization of parameter server for distributed machine learning. The Journal of Supercomputing, 78(12): 13865–13892, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04422-6.

Wu:2022:PSL

[1569] Tsu-Yang Wu, Qian Meng, and Saru Kumari. A provably secure lightweight authentication protocol in mobile edge computing environments. The Journal of Supercomputing, 78(12):13893– 13914, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04411-9.

Melentev:2022:FMS

[1570] V. A. Melent’ev. Formal method for the synthesis of optimal topologies of computing systems based on the projective description of graphs. The Journal of Supercomputing, 78(12): 13915–13941, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04421-7.

Pradhan:2022:DDB

[1571] Subrat Kumar Pradhan and Dushmanta Kumar Das. Delay-discretization-based sliding mode H_{∞} load frequency control scheme considering actuator saturation of windintegrated power system. The Journal of Supercomputing, 78(12):13942– 13987, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04397-4.

Zhang:2022:ECM

[1572] Jinrui Zhang. Exploration on coal mining-induced rockburst prediction using Internet of Things and deep neural network. The Journal of Supercomputing, 78(12):13988– 14008, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04424-4.

Ye:2022:FDA

[1573] Yingjun Ye, Yongdong Zhang, and Weicai Ye. Failure detection algorithm for fail-lagging model applied to HPC. The Journal of Supercomputing, 78(12): 14009–14033, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04347-0.

Noura:2022:DGB

[1574] Hassan N. Noura, Raphaël Couturier, and Kamel Mazouzi. DKEMA: GPUbased and dynamic key-dependent efficient message authentication algorithm. The Journal of Supercomputing, 78(12):14034–14071, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04433-
3.

Eddine:2022:DLB

[1575] Mekkaoui Djamel Eddine and Yanming Shen. A deep learning based approach for predicting the demand of electric vehicle charge. The Journal of Supercomputing, 78(12):14072– 14095, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04428-0.

Zeng:2022:FBA

[1576] Kai Zeng, Qian Ma, and Chenggang Yan. FPGA-based accelerator for object detection: a comprehensive survey. The Journal of Supercomputing, 78(12):14096– 14136, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04415-5.

Leiva:2022:FAA

[1577] Lucas Leiva, Martín Vázquez, and Jordina Torrents-Barrena. FPGA acceleration analysis of LibSVM predictors based on high-level synthesis. The Journal of Supercomputing, 78(12): 14137–14163, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04406-6.

Guo:2022:IAB

[1578] Ping Guo and Yicheng Jiang. An improved ant-based heuristic approach for solving the longest cycle problem in large-scale complex networks. The Journal of Supercomputing, 78(12): 14164–14190, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04409-3.

Liu:2022:CNN

[1579] Liying Liu and Yain-Whar Si. 1D convolutional neural networks for chart pattern classification in financial time series. The Journal of Supercomputing, 78(12):14191– 14214, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04431-5.

Li:2022:PEA

[1580] JianJiang Li, Jiali Li, and Ningming Nie. A parallel ETD algorithm for large-scale rate theory simulation. The Journal of Supercomputing, 78(12): 14215–14230, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04434-2.

Nabi:2022:RRD

[1581] Said Nabi, Muhammad Aleem, and Muhammad Azhar Iqbal. RADL: a resource and deadline-aware dynamic load-balancer for cloud tasks. The Journal of Supercomputing, 78(12): 14231–14265, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04426-2.

Jha:2022:EEW

[1582] Vivekanand Jha and Rashika Sharma.

An energy efficient weighted clustering algorithm in heterogeneous wireless sensor networks. The Journal of Supercomputing, 78(12):14266– 14293, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04429-z.

Wang:2022:AHL

[1583] Shan Wang and Shusheng Tong.

Analysis of high-level dance movements under deep learning and internet of things. The Journal of Supercomputing, 78(12):14294– 14316, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04454-y.

Zeng:2022:GSA

[1584] Guosun Zeng, Huanliang Xiong, and Canghai Wu. Game strategies among multiple cloud computing platforms for non-cooperative competing assignment user tasks. The Journal of Supercomputing, 78(12):14317– 14342, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04437-z.

Wang:2022:FCS

[1585] Yun Wang, Lu Huang, and Austin Lin Yee. Full-convolution Siamese network algorithm under deep learning used in tracking of facial video image in newborns. The Journal of Supercomputing, 78(12):14343– 14361, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04439-x.

Zhao:2022:MCU

[1586] Haoran Zhao, Tao Ren, and Yingyou Wen. Multi-context unsupervised domain adaption for HEp-2 cell classification using maximum partial classifier discrepancy. The Journal of Supercomputing, 78(12):14362– 14380, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04452-0.

Ansari:2022:DPD

[1587] Abdul Q. Ansari, Vipin Sharma, and Rajesh Mishra. A 3-disjoint path design of non-blocking shuffle exchange network by extra port alignment. The Journal of Supercomputing, 78(12): 14381–14401, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04450-2.

Eskandari:2022:MLF

[1588] Sadegh Eskandari. Multi-label feature selection using geometric series of relevance matrix. The Journal of Supercomputing, 78(12):14402– 14418, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04451-1.

Akbar:2022:SVB

[1589] Saeed Akbar and Ruixuan Li. A Shapley value-based thermal-efficient workload distribution in heterogeneous data centers. The Journal of Supercomputing, 78(12):14419– 14447, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04405-7.

Shi:2022:ECI

[1590] Hongyu Shi and Qiubo Li. Edge computing and the Internet of Things on agricultural green productivity. The Journal of Supercomputing, 78(12): 14448–14470, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04463-x.

Montazerolghaem:2022:SVV

[1591] Ahmadreza Montazerolghaem. Softwarization and virtualization of VoIP networks. The Journal of Supercomputing, 78(12):14471–14503, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04448-
w.

Xu:2022:SPS

[1592] Junwei Xu, Ruijuan Zheng, and Qingtao Wu. Service placement strategy for joint network selection and resource scheduling in edge computing. The Journal of Supercomputing, 78(12): 14504–14529, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04458-8.

She:2022:ELO

[1593] Chunyan She and Shaohua Zeng. An enhanced local outlier detection using random walk on grid information graph. The Journal of Supercomputing, 78(12):14530– 14547, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04459-7.

Mittal:2022:PEO

[1594] Payal Mittal, Akashdeep Sharma, and Arun Kumar Sangaiah. On the performance evaluation of object classification models in low altitude aerial data. The Journal of Supercomputing, 78(12):14548– 14570, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04469-5.

Zioviris:2022:CCF

[1595] Georgios Zioviris, Kostas Kolomvatsos, and George Stamoulis. Credit card fraud detection using a deep learning multistage model. The Journal of Supercomputing, 78(12):14571– 14596, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04465-9.

Yang:2022:ISC

[1596] Qifen Yang, Ziyang Li, and Yuhui Deng. An improvement of spectral clustering algorithm based on fast diffusion search for natural neighbor and affinity propagation. The Journal of Supercomputing, 78(12): 14597–14625, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04456-w. See cor-

rection [1674].

Yu:2022:SPC

[1597] Quan Yu, Jin-Yu Song, and Gang Chen. To solve the problems of combat mission predictions based on multiinstance genetic fuzzy systems. The Journal of Supercomputing, 78(12): 14626–14647, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04388-5.

Sabique:2022:SBF

[1598] P. V. Sabique, P. Ganesh, and R. Sivaramakrishnan. Stereovision based force estimation with stiffness mapping in surgical tool insertion using recurrent neural network. The Journal of Supercomputing, 78(12): 14648–14679, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04432-4.

Spurlock:2022:GMI

[1599] Kyle Spurlock and Heba Elgazzar. A genetic mixed-integer optimization of neural network hyper-parameters. The Journal of Supercomputing, 78(12): 14680–14702, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04475-7.

Guasque:2022:SAD

[1600] Ana Guasque, José María Aceituno, and Alfons Crespo. Schedulability analysis of dynamic priority realtime systems with contention. The Journal of Supercomputing, 78(12): 14703–14725, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04446-y.

Wang:2022:BBM

[1601] JunLu Wang, Qiang Liu, and Baoyan Song. Blockchain-based multi-malicious double-spending attack blacklist management model. The Journal of Supercomputing, 78(12): 14726–14755, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04370-1.

Aviles:2022:PAM

[1602] Pablo M. Aviles, Diego Lloria, and Maximo Cobos. Performance analysis of a millimeter wave MIMO channel estimation method in an embedded multi-core processor. The Journal of Supercomputing, 78(12): 14756–14767, August 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04479-3.

Kavezadeh:2022:SRS

[1603] Shirin Kavezadeh, Ashkan Farazin, and Alireza Hosseinzadeh. Supercomputing of reducing sequenced bases in de novo sequencing of the human genome. The Journal of Supercomputing, 78(13):14769– 14793, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04449-9.

Lee:2022:EGD

[1604] Sangmin Lee, Seongjoon Park, and Hwangnam Kim. Enhancing gas detection-based swarming through deep reinforcement learning. The Journal of Supercomputing, 78(13):14794– 14812, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04478-4.

Tripathi:2022:IAD

[1605] Dipty Tripathi, Amit Biswas, and Amrita Chaturvedi. An integrated approach of designing functionality with security for distributed cyberphysical systems. The Journal of Supercomputing, 78(13):14813–14845, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04481-9.

Li:2022:GCN

[1606] Xiaowen Li, Ran Lu, and Zhenfang Zhu. Graph convolutional networks with hierarchical multihead attention for aspect-level sentiment classification. The Journal of Supercomputing, 78(13):14846– 14865, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04480-w.

Nayak:2022:ELM

[1607] Janmenjoy Nayak, Saroj K. Meher, and S. Vimal. Extreme learning machine and Bayesian optimizationdriven intelligent framework for IoMT cyber-attack detection. The Journal of Supercomputing, 78(13):14866– 14891, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04453-z.

Mageshkumar:2022:ISF

[1608] N. Mageshkumar and L. Lakshmanan. An improved secure file deduplication avoidance using CKHO based deep learning model in a cloud environment. The Journal of Supercomputing, 78(13):14892–14918, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04436-0.

Ikram:2022:IEV

[1609] Muhammad Jawad Ikram, Mostafa Elsayed Saleh, and Osama Ahmed Abulnaja. Investigating the effect of varying block size on power and energy consumption of GPU kernels. The Journal of Supercomputing, 78(13):14919– 14939, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04473-9.

Anglada:2022:DSR

[1610] Martí Anglada, Enrique de Lucas, and Antonio González. Dynamic sampling rate: harnessing frame coherence in graphics applications for energy-efficient GPUs. The Journal of Supercomputing, 78(13):14940– 14964, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04413-7.

Song:2022:PDI

[1611] Kai Song, Wen Li, and Xianchao Wang.

Parallel design and implementation of Jacobi iterative algorithm based on ternary optical computer. The Journal of Supercomputing, 78(13):14965– 14990, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04471-x.

Yadav:2022:EEP

[1612] Vijay Kumar Yadav, Nitish Andola, and S. Venkatesan. EP2LBS: efficient privacy-preserving scheme for location-based services. The Journal of Supercomputing, 78(13):14991– 15013, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04486-4.

Sharma:2022:NSD

[1613] Rakhi Sharma and Shail Kumar Dinkar. A novel social deep autoencoder NMF incentive scheme to detect a selfish node in delay tolerant network. The Journal of Supercomputing, 78(13):15014– 15041, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04423-5.

Gasmi:2022:HDL

[1614] Karim Gasmi. Hybrid deep learning model for answering visual medical questions. The Journal of Supercomputing, 78(13):15042–15059, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04474-8.

Lin:2022:BSR

[1615] Pei-Chun Lin, Benjamin Yankson, and Manabu Tsukada. Building a speech recognition system with privacy identification information based on Google Voice for social robots. The Journal of Supercomputing, 78(13):15060– 15088, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04487-3.

Sarkar:2022:DLB

[1616] Indranil Sarkar and Sanjay Kumar. Deep learning-based energyefficient computational offloading strategy in heterogeneous fog computing networks. The Journal of Supercomputing, 78(13):15089–15106, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04461-z.

Mounica:2022:FSD

[1617] B. Mounica and K. Lavanya. Feature selection with a deep learning based high-performance computing model for traffic flow analysis of Twitter data. The Journal of Supercomputing, 78(13):15107– 15122, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04468-6.

Li:2022:MOS

[1618] Ming Li, Jianshan Zhang, and Xing Chen. MultiOff: offloading support and service deployment for multiple IoT applications in mobile edge computing. The Journal of Supercomputing, 78(13):15123– 15153, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04490-8.

Beranek:2022:AWS

[1619] Jakub Beránek, Stanislav Böhm, and Vojtech Cima. Analysis of workflow schedulers in simulated distributed environments. The Journal of Supercomputing, 78(13):15154– 15180, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04438-y.

Ameri:2022:CGG

[1620] Reyhaneh Ameri, Mohammad Reza Meybodi, and Mohammad Mehdi Daliri Khomami. Cellular Goore Game and its application to quality-of-service control in wireless sensor networks. The Journal of Supercomputing, 78(13):15181–15228, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04435-1.

Gao:2022:HEB

[1621] Xiue Gao, Bo Chen, and Yunming Wang. Hierarchy-entropy based method for command and control networks reconfiguration. The Journal of Supercomputing, 78(13):15229– 15249, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04445-z.

Heszberger:2022:HTE

[1622] Zalán Heszberger. Hyperbolic trees for efficient routing computation. The Journal of Supercomputing, 78 (13):15250–15268, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04485-
5.

Guo:2022:EPG

[1623] Bin Guo and Emil Sekerinski. Efficient parallel graph trimming by arc-consistency. The Journal of Supercomputing, 78(13):15269–15313, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04457-9.

Lo:2022:VRH

[1624] Chih-Min Lo, Jian-Hong Wang, and Hsing-Wen Wang. Virtual reality human-robot interaction technology acceptance model for learning direct current and alternating current. The Journal of Supercomputing, 78(13):15314–15337, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04455-x.

Ponte-Fernandez:2022:FPO

[1625] Christian Ponte-Fernández, Jorge González-Domínguez, and María J. Martín. Fiuncho: a program for any-order epistasis detection in CPU clusters. The Journal of Supercomputing, 78(13):15338–15357, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04477-5.

Kashi:2022:MAA

[1626] Somayeh Kashi, Ahmad Patooghy, and Mahdi Fazeli. A multi-application approach for synthesizing custom network-on-chips. The Journal of Supercomputing, 78(13):15358–15380, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04444-0.

Gong:2022:QAO

[1627] Changqing Gong, Ting Wang, and Han Qi. A quantum approximate optimization algorithm for solving Hamilton path problem. The Journal of Supercomputing, 78(13):15381– 15403, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04462-y.

Leon:2022:MCB

[1628] Betzabeth León, Sandra Méndez, and Emilio Luque. A model of checkpoint behavior for applications that have I/O. The Journal of Supercomputing, 78(13):15404–15436, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04482-8. See cor-

rection [1629].

Leon:2022:CMC

[1629] Betzabeth León, Sandra Méndez, and Emilio Luque. Correction to: A model of checkpoint behavior for applications that have I/O. The Journal of Supercomputing, 78(13):15437– 15438, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04571-8.  See

[1628].

Prakash:2022:HEM

[1630] P. N. Senthil Prakash and N. Rajkumar. HSVNN: an efficient medical data classification using dimensionality reduction combined with hybrid support vector neural network. The Journal of Supercomputing, 78(13):15439– 15462, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04500-9.

Sham:2022:ACR

[1631] Eht E. Sham and Deo Prakash Vidyarthi. Admission control and resource provisioning in fog-integrated cloud using modified fuzzy inference system. The Journal of Supercomputing, 78(13):15463–15503, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04483-7.

Liang:2022:RAP

[1632] Tengteng Liang, Mingzu Zhang, and Xing Yang. Reliability analysis of the pentanary n-cube based on hextra edge-connectivity with a concentration behavior. The Journal of Supercomputing, 78(13):15504– 15531, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04489-1.

Guo:2022:RMS

[1633] Jingjing Guo, Chunlin Li, and Youlong Luo. Resource management and switch migration in SDN-based multi-access edge computing environments. The Journal of Supercomputing, 78(13):15532–15566, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04493-5.

Kalaiarasi:2022:PEC

[1634] M. Kalaiarasi, V. R. Venkatasubramani, and S. Rajaram. A parallel elliptic curve crypto-processor architecture with reduced clock cycle for FPGA platforms. The Journal of Supercomputing, 78(13):15567– 15597, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04442-2.

Pashaei:2022:HBA

[1635] Elham Pashaei and Elnaz Pashaei. Hybrid binary arithmetic optimization algorithm with simulated annealing for feature selection in high-dimensional biomedical data. The Journal of Supercomputing, 78(13):15598–15637, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04507-2.

Shayegan:2022:EVS

[1636] Mohammad-Javad Shayegan and Mehdi Faizollahi-Samarin. An extended version of sectional MinHash method for near-duplicate detection. The Journal of Supercomputing, 78(13):15638– 15662, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04447-x.

Ji:2022:EAS

[1637] Zeyu Ji, Xingjun Zhang, and Zheng Wei. EP4DDL: addressing straggler problem in heterogeneous distributed deep learning. The Journal of Supercomputing, 78(13):15663– 15680, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04466-8.

Zhang:2022:EAH

[1638] Chuanzhu Zhang. Evaluation and analysis of human resource management mode and its talent screening factors based on decision tree algorithm. The Journal of Supercomputing, 78(13):15681–15713, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04499-z.

Belloch:2022:MIM

[1639] Jose A. Belloch, José M. Badía, and Vesa Välimäki. Multicore implementation of a multichannel parallel graphic equalizer. The Journal of Supercomputing, 78(14):15715– 15729, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04495-3.

Bahig:2022:SWF

[1640] Hazem M. Bahig, Dieaa I. Nassr, and Hatem M. Bahig. Speeding up wheel factoring method. The Journal of Supercomputing, 78(14):15730– 15748, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04470-y.

Rahmani:2022:NAS

[1641] Peyman Rahmani, Seyed Mostafa Fakhrahmad, and Mohammad Taheri. New attacks on secret sharing-based data outsourcing: toward a resistant scheme. The Journal of Supercomputing, 78(14):15749–15785, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04467-7.

Suwi:2022:RRP

[1642] Hanan Suwi, Laaziz Lahlou, and Claes Edstrom. RAFALE: Rethinking the provisioning of virtuAl network services using a fast and scAlable machine LEarning approach. The Journal of Supercomputing, 78(14):15786– 15819, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04492-6.

Souza:2022:SLR

[1643] Lubnnia Souza, Kádna Camboim, and Fernanda Alencar. A systematic literature review about integrating dependability attributes, performability and sustainability in the implantation of cooling subsystems in data center. The Journal of Supercomputing, 78(14):15820– 15856, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04515-2.

Kung:2022:EAS

[1644] Tzu-Liang Kung. Exact assessment of the super P_{k}-connectivity for the crossed cube interconnection network. The Journal of Supercomputing, 78(14):15857–15881, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04494-4.

Geng:2022:ABB

[1645] Tongtong Geng and Yueping Du. Applying the blockchain-based deep reinforcement consensus algorithm to the intelligent manufacturing model under Internet of Things. The Journal of Supercomputing, 78(14):15882– 15904, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04514-3.

Taherinia:2022:HPA

[1646] Mohsen Taherinia, Mahdi Esmaeili, and Behrouz Minaei-Bidgoli. A high-performance algorithm for finding influential nodes in large-scale social networks. The Journal of Supercomputing, 78(14):15905–15952, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04418-2.

Lin:2022:SQF

[1647] Hsiu-Hsia Lin, Tianyi Zhang, and Shih-Ku Kuang. A system for quantifying facial symmetry from 3D contour maps based on transfer learning and fast R-CNN. The Journal of Supercomputing, 78(14):15953– 15973, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04502-7.

Ruan:2022:ACD

[1648] Chunyan Ruan, Wen Qu, and Kuan-Han Lu. An algorithm for calculating the degree of similarity between English words through the different position and appearance coefficients of letters. The Journal of Supercomputing, 78(14):15974– 15994, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04511-6.

Marchang:2022:TRU

[1649] Ningrinla Marchang, Goldie M. Meitei, and Tejendra Thakur. Task reduction using regression-based missing data imputation in sparse mobile crowdsensing. The Journal of Supercomputing, 78(14):15995–16028, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04518-z.

Zhang:2022:SHI

[1650] Xiaomei Zhang, Pengming Zhang, and Chi hung Chi. sAuth: a hierarchical implicit authentication mechanism for service robots. The Journal of Supercomputing, 78(14):16029– 16055, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04472-w.

Alshawish:2022:EMA

[1651] Islam Alshawish and Ali Al-Haj.

An efficient mutual authentication scheme for IoT systems. The Journal of Supercomputing, 78(14):16056– 16087, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04520-5.

Song:2022:HPB

[1652] Yao Song, Liang Wang, and Jinquan Wang. Hypergraph-partitioning-based online joint scheduling of tasks and data. The Journal of Supercomputing, 78(14):16088–16117, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04460-0.

Devassy:2022:NNB

[1653] Deepa Devassy, J. Immanuel Johnraja, and Getzi Jeba Leelipushpam Paulraj. NBA: novel bio-inspired algorithm for energy optimization in WSN for IoT applications. The Journal of Supercomputing, 78(14):16118– 16135, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04505-4.

Kaviani:2022:CCQ

[1654] Farzaneh Kaviani and Mohammadreza Soltanaghaei. CQARPL: Congestion and QoS-aware RPL for IoT applications under heavy traffic. The Journal of Supercomputing, 78(14):16136– 16166, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04488-2.

Kumar:2022:RRA

[1655] Vinod Kumar, Mahmoud Shuker Mahmoud, and Adesh Kumari. RAPCHI: Robust authentication protocol for IoMT-based cloud-healthcare infrastructure. The Journal of Supercomputing, 78(14):16167–16196, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04513-4.

Ponnusamy:2022:DAT

[1656] Muruganantham Ponnusamy, Pradeep Bedi, and N. Yuvaraj. Design and analysis of text document clustering using salp swarm algorithm. The Journal of Supercomputing, 78(14):16197– 16213, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04525-0.

Cheng:2022:DSA

[1657] Qi Cheng, Yixin Chen, and Weidong Liu. A dual-stage attentionbased Bi-LSTM network for multivariate time series prediction. The Journal of Supercomputing, 78(14):16214– 16235, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04506-3.

Xiao:2022:ISE

[1658] Han Xiao, Shiyang Xiao, and Cailin Li.

Image Sobel edge extraction algorithm accelerated by OpenCL. The Journal of Supercomputing, 78(14):16236– 16265, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04404-8.

Lei:2022:GGE

[1659] Kai Lei, Hao Ye, and Jing Xiao.

GBRM: a graph embedding and blockchain-based resource management framework for 5G MEC. The Journal of Supercomputing, 78(14):16266– 16285, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04528-x.

Chung:2022:EEE

[1660] Moonyoung Chung, Soon J. Hyun, and Woong-Kee Loh. Efficient exact k-flexible aggregate nearest neighbor search in road networks using the M-tree. The Journal of Supercomputing, 78(14):16286–16302, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04496-2.

Ardakani:2022:DSA

[1661] Mahdi Maleknasab Ardakani, Mohammad Ali Tabarzad, and Mohammad Amin Shayegan. Detecting sybil attacks in vehicular ad hoc networks using fuzzy logic and arithmetic optimization algorithm. The Journal of Supercomputing, 78(14):16303– 16335, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04526-z.

Manimaran:2022:SSB

[1662] S. Manimaran, V. N. Sastry, and N. P. Gopalan. STMAD: sensor-based threat’s mitigation on smartphones using allowlist and denylist. The Journal of Supercomputing, 78(14):16336– 16363, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04523-2.

Salvakkam:2022:MLM

[1663] Dilli Babu Salvakkam and Rajendra Pamula. MESSB-LWE: multiextractable somewhere statistically binding and learning with errorbased integrity and authentication for cloud storage. The Journal of Supercomputing, 78(14):16364–16393, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04497-1.

Pal:2022:ESU

[1664] Gautam Pal. An efficient system using implicit feedback and lifelong learning approach to improve recommendation. The Journal of Supercomputing, 78(14):16394– 16424, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04484-6.

Wang:2022:ADL

[1665] Lei Wang. Application of deep learning to detect defects on the surface of steel balls in an IoT environment. The Journal of Supercomputing, 78(14):16425– 16452, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04516-1. See re-

traction note [2306].

Theerthagiri:2022:MPR

[1666] Prasannavenkatesan Theerthagiri. Mobility prediction for random walk mobility model using ARIMA in mobile ad hoc networks. The Journal of Supercomputing, 78(14):16453– 16484, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04503-6.

Li:2022:AIB

[1667] Kunmei Li and Nasser Fard. Analysis of impact of balanced level on MI-based and non-MI-based feature selection methods. The Journal of Supercomputing, 78(14):16485– 16497, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04504-5.

Ramachandran:2022:ANN

[1668] R. Ramachandran and K. Arutchelvan. ArRaNER: A novel named entity recognition model for biomedical literature documents. The Journal of Supercomputing, 78(14):16498– 16511, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04527-y.

Hu:2022:MRU

[1669] Yuxi Hu, Taimeng Fu, and Man-On Pun. 3D map reconstruction using a monocular camera for smart cities. The Journal of Supercomputing, 78(14):16512–16528, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04512-5.

Sun:2022:SAG

[1670] Kaili Sun, Yuan Li, and Quan Hu. Syntax-aware graph convolutional network for the recognition of Chinese implicit intersentence relations. The Journal of Supercomputing, 78(14):16529–16552, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04476-6.

Sultana:2022:ERM

[1671] Tangina Sultana and Young-Koo Lee. Efficient rule mining and compression for RDF style KB based on Horn rules. The Journal of Supercomputing, 78(14):16553–16580, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04519-y.

Li:2022:HDI

[1672] Xiao Li and Kewen Li. Highdimensional imbalanced biomedical data classification based on P-AdaBoost-PAUC algorithm. The Journal of Supercomputing, 78(14):16581– 16604, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04509-0.

Zhao:2022:ECS

[1673] Ying-Ze Zhao, Xiang-Jun Li, and Meijie Ma. Embedded connectivity of some BC networks. The Journal of Supercomputing, 78(14):16605– 16618, September 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04522-3.

Yang:2022:CIS

[1674] Qifen Yang, Ziyang Li, and Yuhui Deng. Correction to: An improvement of spectral clustering algorithm based on fast diffusion search for natural neighbor and affinity propagation. The Journal of Supercomputing, 78(14):16619–16620, September 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04743-6.  See

[1596].

Liu:2022:EAA

[1675] Xi Liu, Jun Liu, and Hong Wu.

Energy-aware allocation for delaysensitive multitask in mobile edge computing. The Journal of Supercomputing, 78(15):16621–16646, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04550-
z.

Battula:2022:BBF

[1676] Sudheer Kumar Battula, Saurabh Garg, and Erfan Aghasian. A blockchain-based framework for automatic SLA management in fog computing environments. The Journal of Supercomputing, 78(15):16647– 16677, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04545-w.

Hajian:2022:LAK

[1677] R. Hajian, S. H. Erfani, and S. Kumari. A lightweight authentication and key agreement protocol for heterogeneous IoT with special attention to sensing devices and gateway. The Journal of Supercomputing, 78(15): 16678–16720, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04464-w.

Wu:2022:ONR

[1678] Guanqaun Wu and Desheng Zeng. Optimization of network resource management based on software-defined networking in the 5G environment. The Journal of Supercomputing, 78(15): 16721–16744, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04547-8.

Priya:2022:DDB

[1679] J. Sathya Priya, A. Bhagyalakshmi, and N. Deepa. DBAHHO: Deep belief network-based adaptive Harris hawks optimization for adaptive offloading strategy in mobile edge computing. The Journal of Supercomputing, 78(15): 16745–16769, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04501-8.

Patil:2022:DLM

[1680] Jitendra Patil, Vrinda Tokekar, and Anil Rawat. Discriminate, locate and mitigate DDoS traffic in presence of Flash Crowd in Software Defined Network. The Journal of Supercomputing, 78(15):16770– 16793, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04538-9.

Liu:2022:DDS

[1681] Jianhang Liu, Jiebing Wang, and Xuerong Cui. A data distribution scheme for VANET based on fountain code. The Journal of Supercomputing, 78(15):16794– 16819, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04536-x.

Barriga:2022:ADM

[1682] R. Barriga, M. Romero, and H. Hassan. Advanced data modeling for industrial drying machine energy optimization. The Journal of Supercomputing, 78(15):16820– 16840, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04498-0.

Huang:2022:PPP

[1683] Po-Hsuan Huang, Ting-Wei Chang, and Shen-Ming Chung. POPS: an offpeak precomputing scheme for privacypreserving computing. The Journal of Supercomputing, 78(15):16841– 16860, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04552-x.

Gupta:2022:LDA

[1684] Sakshi Gupta and Itu Snigdh. Leveraging data aggregation algorithm in LoRa networks. The Journal of Supercomputing, 78(15):16861– 16875, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04534-z.

Gong:2022:NAD

[1685] Changqing Gong, Weiqi Guan, and Han Qi. Network attack detection scheme based on variational quantum neural network. The Journal of Supercomputing, 78(15):16876– 16897, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04542-z.

Ghosh:2022:RRA

[1686] Subha Ghosh and Debashis De. RAM:

resource allocation in MIMO-MISO cognitive IoT for 5G wireless networks using two-level weighted majority cooperative game. The Journal of Supercomputing, 78(15):16898– 16950, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04546-9.

Bala:2022:TCP

[1687] R. Bala and R. Manoharan. Trusted consensus protocol for blockchain networks based on fuzzy inference system. The Journal of Supercomputing, 78(15):16951–16974, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04510-
7.

Hariri:2022:HAS

[1688] Malihe Hariri, Mostafa Nouri-Baygi, and Saeid Abrishami. A hybrid algorithm for scheduling scientific workflows in IaaS cloud with deadline constraint. The Journal of Supercomputing, 78(15):16975– 16996, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04563-8.

Faraji-Mehmandar:2022:SLA

[1689] Mohammad Faraji-Mehmandar, Sam Jabbehdari, and Hamid Haj Seyyed Javadi. A self-learning approach for proactive resource and service provisioning in fog environment. The Journal of Supercomputing, 78(15): 16997–17026, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04521-4.

Deldari:2022:CNM

[1690] Arash Deldari, Abolghasem Yousofi, and Alireza Salehan. CDA: a novel multicore scheduling for cost-aware deadline-constrained scientific workflows on the IaaS cloud. The Journal of Supercomputing, 78(15):17027– 17054, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04551-y.

Zhou:2022:EDF

[1691] Binbin Zhou and Lu Lu. An effective 3-D Fast Fourier Transform framework for multi-GPU accelerated distributed-memory systems. The Journal of Supercomputing, 78(15): 17055–17073, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04491-7.

Dupak:2022:HTW

[1692] Lucindia Dupak and Subhasish Banerjee. Hybrid trust and weight evaluation-based trust assessment using ECK-ANFIS and AOMDV-REPO-based optimal routing in MANET environment. The Journal of Supercomputing, 78(15):17074– 17094, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04530-3.

Zhang:2022:DMM

[1693] Chongchong Zhang, Yannan Shi, and Yiying Wang. A denoising method of mine microseismic signal based on NAEEMD and frequencyconstrained SVD. The Journal of Supercomputing, 78(15):17095– 17113, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04554-9.

Cheng:2022:CSF

[1694] Xiaochun Cheng, Seifedine Kadry, and Rubén González Crespo. CNN supported framework for automatic extraction and evaluation of dermoscopy images. The Journal of Supercomputing, 78(15):17114– 17131, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04561-w.

Shivhare:2022:SSB

[1695] Anubhav Shivhare, Manish Kumar Maurya, and Manish Kumar. A secret sharing-based scheme for secure and energy efficient data transfer in sensor-based IoT. The Journal of Supercomputing, 78(15):17132– 17149, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04533-0.

Raca:2022:REC

[1696] Valon Raca, Seeun William Umboh, and Bernhard Scholz. Runtime and energy constrained work scheduling for heterogeneous systems. The Journal of Supercomputing, 78(15):17150– 17177, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04556-7.

Goswami:2022:MCD

[1697] Veena Goswami and Gopinath Panda. Multimedia content delivery services in the cloud with partial sleep and abandonment. The Journal of Supercomputing, 78(15):17178– 17201, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04532-1.

Goudarzi:2022:RDA

[1698] Parisa Goudarzi, Amir Masoud Rahmani, and Mohammad Mosleh. Resource discovery approaches in cloudIoT: a systematic review. The Journal of Supercomputing, 78(15): 17202–17230, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04541-0.

Camara:2022:PPC

[1699] Jesús Cámara, José-Carlos Cano, and Mariano Saura-Sánchez. PARCSIM: a parallel computing simulator for scalable software optimization. The Journal of Supercomputing, 78(15): 17231–17246, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04537-w.

Yin:2022:HDR

[1700] Shanshan Yin and Liqiong Xu. Hybrid diagnosis of regular networks under the HPMC fault model. The Journal of Supercomputing, 78(15): 17247–17260, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04531-2.

Li:2022:BBM

[1701] Yue Li, Mingcheng Xu, and Gaojian Xu. Blockchain-based mutual authentication protocol without CA. The Journal of Supercomputing, 78(15): 17261–17283, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04558-5.

Rani:2022:BBI

[1702] Poonam Rani, Preeti Kaur, and Sweety Nain. Blockchain-based IoT enabled health monitoring system. The Journal of Supercomputing, 78(15):17284– 17308, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04584-3.

Ahmadi:2022:DSA

[1703] Kourosh Dadashtabar Ahmadi, Ali Jabar Rashidi, and Ali Massomi Moghri. Design and simulation of autonomous military vehicle control system based on machine vision and ensemble movement approach. The Journal of Supercomputing, 78(15):17309– 17347, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04565-6.

Masoumi:2022:SAA

[1704] Maryam Masoumi and Hassan Motallebi. A structure-aware algorithm for fault-tolerant scheduling of scientific workflows. The Journal of Supercomputing, 78(15):17348– 17377, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04529-w.

Li:2022:CRF

[1705] Fei Li, Yuzhu Wang, and He Zhang. CC-RRTMG SW++: Further optimizing a shortwave radiative transfer scheme on GPU. The Journal of Supercomputing, 78(15):17378– 17402, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04566-5.

Binbusayyis:2022:ICM

[1706] Adel Binbusayyis, Haya Alaskar, and M. Dinesh. An investigation and comparison of machine learning approaches for intrusion detection in IoMT network. The Journal of Supercomputing, 78(15):17403– 17422, October 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04568-3.

Pirozmand:2022:GHA

[1707] Poria Pirozmand, Amir Javadpour, and Forough Ja’fari. GSAGA: a hybrid algorithm for task scheduling in cloud infrastructure. The Journal of Supercomputing, 78(15): 17423–17449, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04539-8.

Yu:2022:NIR

[1708] Anqi Yu and Nuo Wang. Nodeimportance ranking in scale-free networks: a network metric response model and its solution algorithm. The Journal of Supercomputing, 78(15): 17450–17469, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04544-x.

Gao:2022:FSP

[1709] Zheng Gao, Chenxiang Zhang, and Zhengyin Li. Financial sequence prediction based on swarm intelligence algorithms and Internet of Things. The Journal of Supercomputing, 78(15): 17470–17490, October 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04572-7.

Liu:2022:OMT

[1710] Xi Liu and Jun Liu. An online mechanism for task allocation and pricing in crowd sensing systems. The Journal of Supercomputing, 78(16): 17594–17618, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04564-7.

Li:2022:CLB

[1711] Wei Li, Liangqilin Ni, and Lei Wang. Cumulative learning-based competitive swarm optimizer for large-scale optimization. The Journal of Supercomputing, 78(16):17619– 17656, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04553-w.

Zhu:2022:MOO

[1712] Sifeng Zhu, Mingyang Zhao, and Qinghua Zhang. Multi-objective optimal offloading decision for multiuser structured tasks in intelligent transportation edge computing scenario. The Journal of Supercomputing, 78(16):17797–17825, November 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04549-6.

Xu:2022:DCN

[1713] Yan Xu, Yuwen Wang, and Yuyue Du. Deep convolutional neural networks for bias field correction of brain magnetic resonance images. The Journal of Supercomputing, 78(16):17943– 17968, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04575-4.

Xu:2022:DIC

[1714] Tong Xu, Changrong Zhao, and Zhaoyang You. Design of intelligent control system for printing and dyeing wastewater treatment under internet of things and deep learning. The Journal of Supercomputing, 78(16): 18023–18050, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04524-1.

Wu:2022:MOO

[1715] Xian Wu and Zhaocai Wang. Multiobjective optimal allocation of regional water resources based on slime mould algorithm. The Journal of Supercomputing, 78(16):18288– 18317, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04599-w.

Raamesh:2022:TCM

[1716] Lilly Raamesh, S. Jothi, and S. Radhika. Test case minimization and prioritization for regression testing using SBLA-based AdaBoost convolutional neural network. The Journal of Supercomputing, 78(16):18379– 18403, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04540-1.

Zhong:2022:SSD

[1717] Yong Zhong, Liang Chen, Changlin Dan, and Amin Rezaeipanah. A systematic survey of data mining and big data analysis in Internet of Things. The Journal of Supercomputing, 78(17): 18405–18453, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04594-1.

Ghadamgahi:2022:NTA

[1718] Seyed Mahdi Ghadamgahi, Reza Sabbaghi-Nadooshan, and Keivan Navi. Novel ternary adders and subtractors in quantum cellular automata. The Journal of Supercomputing, 78(17):18454–18496, November 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04593-2.

Paliwal:2022:MAB

[1719] Priyanka Paliwal, Julian L. Webber, Abolfazl Mehbodniya, Mohd Anul Haq, Anil Kumar, and Prem Kumar Chaurasiya. Multi-agent-based approach for generation expansion planning in isolated micro-grid with renewable energy sources and battery storage. The Journal of Supercomputing, 78(17):18497–18523, November 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04609-x. See cor-

rection [1810].

Huan:2022:RSI

[1720] Hai Huan, Nan Zou, Yi Zhang, Yaqin Xie, and Chao Wang. Remote sensing image reconstruction using an asymmetric multi-scale superresolution network. The Journal of Supercomputing, 78(17):18524– 18550, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04617-x.

Wang:2022:PJF

[1721] Qiqi Wang, Hongjie Zhang, Jing Li, Yu Shen, and Xiaohui Liu. Predicting job finish time based on parameter features and running logs in supercomputing system. The Journal of Supercomputing, 78(17):18551– 18577, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04582-5.

Bukhari:2022:EVC

[1722] Maryam Bukhari, Mehr Yahya Durrani, Saira Gillani, Sadaf Yasmin, Seungmin Rho, and Sang-Soo Yeo. Exploiting vulnerability of convolutional neural network-based gait recognition system. The Journal of Supercomputing, 78(17):18578– 18597, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04611-3.

Fang:2022:APM

[1723] Zheng Fang, Bichao Ye, Bingan Yuan, Tingjun Wang, Shuo Zhong, Shunren Li, and Jianyi Zheng. Angle prediction model when the imaging plane is tilted about z-axis. The Journal of Supercomputing, 78(17): 18598–18615, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04595-0.

Guan:2022:TPC

[1724] Yingrong Guan, Yaoyu Qiu, and Cheng Tian. Trajectory planning in college football training using deep learning and the Internet of Things. The Journal of Supercomputing, 78(17): 18616–18635, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04619-9.

Manshad:2022:NMW

[1725] Mozhdeh Khaksar Manshad, Mohammad Reza Meybodi, and Afshin Salajegheh. A new multi-wave continuous action-set cellular learning automata for link prediction problem in weighted multi-layer social networks. The Journal of Supercomputing, 78(17):18636– 18665, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04615-z.

Kandasamy:2022:NOU

[1726] Nehru Kandasamy, Vaishali Dhare, and Nagarjuna Telagam. Novel optimized ultra-dense 1-bit magnitude comparator design in quantumdot cellular automata technology based on MV32 gate. The Journal of Supercomputing, 78(17):18666– 18690, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04604-2.

Ozturk:2022:SEP

[1727] Zuhal Ozturk, Haluk Rahmi Topcuoglu, and Mahmut Taylan Kandemir. Studying error propagation on application data structure and hardware. The Journal of Supercomputing, 78(17): 18691–18724, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04625-x.

BenHafaiedh:2022:DFB

[1728] Imene Ben Hafaiedh and Maroua Ben Slimane. A distributed formal-based model for self-healing behaviors in autonomous systems: from failure detection to self-recovery. The Journal of Supercomputing, 78(17):18725– 18753, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04614-0.

Li:2022:GCB

[1729] Ying Li, Xueting Cui, Jiahao Fan, and Tan Wang. Global chaotic bat algorithm for feature selection. The Journal of Supercomputing, 78(17): 18754–18776, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04606-0.

Alagheband:2022:AES

[1730] Mahdi R. Alagheband and Atefeh Mashatan. Advanced encryption schemes in multi-tier heterogeneous Internet of Things: taxonomy, capabilities, and objectives. The Journal of Supercomputing, 78(17):18777– 18824, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04586-1.

Sindhu:2022:WCS

[1731] Korrapati Sindhu, Karthick Seshadri, and Chidambaran Kollengode. Workload characterization and synthesis for cloud using generative stochastic processes. The Journal of Supercomputing, 78(17):18825–18855, November 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04597-y.

Wang:2022:VNS

[1732] Yun Wang, Xingquan Zuo, Zhiqiang Wu, Hui Wang, and Xinchao Zhao. Variable neighborhood search based multiobjective ACO-list scheduling for cloud workflows. The Journal of Supercomputing, 78(17):18856– 18886, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04616-y.

Gao:2022:UCT

[1733] Long Gao, Fen Zheng, and Jian Yong Bian. Using computer theory to detect PCB defects in an IoT environment. The Journal of Supercomputing, 78(17):18887–18914, November 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04610-4.

Lin:2022:ADT

[1734] Chuen-Horng Lin, Chia-Ching Yu, and Huan-Yu Chen. Augmentation dataset of a two-dimensional neural network model for use in the car parts segmentation and car classification of three dimensions. The Journal of Supercomputing, 78(17):18915– 18958, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04630-0.

He:2022:RDI

[1735] Qinlu He, Genqing Bian, Weiqi Zhang, and Zhen Li. RTFTL: design and implementation of real-time FTL algorithm for flash memory. The Journal of Supercomputing, 78(17):18959– 18993, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04602-4.

Tian:2022:EAS

[1736] Junfeng Tian and Qianqian Song. An efficient auditing scheme with a novel structure for multiple replicas. The Journal of Supercomputing, 78(17): 18994–19019, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04598-x.

Chen:2022:NNN

[1737] Chong Chen, Yixuan Dou, Jie Chen, and Yaru Xue. A novel neural network training framework with data assimilation. The Journal of Supercomputing, 78(17):19020– 19045, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04629-7.

Qamar:2022:CSA

[1738] Shamimul Qamar, Abdul Azeem, Tanweer Alam, and Izhar Ahmad. A crow search algorithm integrated with dynamic awareness probability for cellular network cost management. The Journal of Supercomputing, 78(17): 19046–19069, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04623-z.

Pandey:2022:PIP

[1739] Shubhang Pandey and T. G. Venkatesh. Performance investigation of packetbased communication in 3D-memories. The Journal of Supercomputing, 78 (17):19070–19096, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04605-
1.

Moreno:2022:HEE

[1740] J. J. Moreno, E. M. Garzón, J. J. Fernández, and A. Martínez-Sánchez. HPC enables efficient 3D membrane segmentation in electron tomography. The Journal of Supercomputing, 78(17):19097–19113, November 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04607-z.

Khair:2022:DEM

[1741] Younes Khair, Abdeslem Dennai, and Youssef Elmir. Dynamic and elastic monitoring of VMs in cloud environment. The Journal of Supercomputing, 78(17):19114–19137, November 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04624-y.

Su:2022:BPS

[1742] Fuzhi Su and Meihong Chen. Basketball players’ score prediction using artificial intelligence technology via the Internet of Things. The Journal of Supercomputing, 78(17):19138– 19166, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04573-6.

Chen:2022:SFA

[1743] Heng Chen, Ziheng Wang, Xi Xiao, Jingbo Li, Xiaoshe Dong, and Xingjun Zhang. SunwayURANS: 3D fullannulus URANS simulations of transonic axial compressors on Sunway TaihuLight. The Journal of Supercomputing, 78(17):19167–19187, November 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04628-8.

Masoudi:2022:SSM

[1744] Sepide Masoudi and Faramarz Safi-Esfahani. SM@RMFFOG: sensor mining at resource management framework of fog computing. The Journal of Supercomputing, 78(17):19188– 19227, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04592-3.

Lim:2022:ERR

[1745] Chhayly Lim, Jungyeon Kim, Jeongseok Kim, Byeong-Gwon Kang, and Yunyoung Nam. Estimation of respiratory rate in various environments using microphones embedded in face masks. The Journal of Supercomputing, 78(17):19228– 19245, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04622-0.

Ullah:2022:EAI

[1746] Farhan Ullah, Jihoon Moon, Hamad Naeem, and Sohail Jabbar. Explainable artificial intelligence approach in combating real-time surveillance of COVID19 pandemic from CT scan and X-ray images using ensemble model. The Journal of Supercomputing, 78(17):19246– 19271, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04631-z.

Zhai:2022:GPS

[1747] Xiaomeng Zhai, Hong Zhang, Xu Huang, and Shouhua Zhang. Graph partitioning strategies: one size does not fit all. The Journal of Supercomputing, 78(17):19272–19295, November 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04620-2.

Choudhary:2022:OCC

[1748] Pankaj Kumar Choudhary and Dushmanta Kumar Das. Optimal coordination of over-current relay in a power distribution network using aggrandized class topper optimization (A-CTO) algorithm. The Journal of Supercomputing, 78(17):19296– 19321, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04632-y.

Wang:2022:LSI

[1749] SuHua Wang, ZhiQiang Ma, XiaoXin Sun, HuiNan Zhao, XiuZhuo Wei, Rui Ma, and Bo Tang. Leveraging side information as adjusting embedding to improve user representation for recommendations. The Journal of Supercomputing, 78(17):19322– 19345, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04635-9.

Tirado:2022:SPA

[1750] Felipe Tirado, Alvaro Wong, Dolores Rexachs, and Emilio Luque. Scalable performance analysis method for SPMD applications. The Journal of Supercomputing, 78(17):19346– 19371, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04588-z.

Bai:2022:IBT

[1751] Jing Bai, Junfeng Zhou, Ming Du, and Ziyang Chen. Index-based top kα-maximal-clique enumeration over uncertain graphs. The Journal of Supercomputing, 78(17):19372– 19400, November 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04613-1.

Gan:2022:CNN

[1752] Baiqiang Gan, Yuqiang Chen, Qiuping Dong, Jianlan Guo, and Rongxia Wang. A convolutional neural network intrusion detection method based on data imbalance. The Journal of Supercomputing, 78(18):19401– 19434, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04633-x.

Chuang:2022:PAE

[1753] Yung-Ting Chuang and Chiu-Shun Hsiang. A popularity-aware and energy-efficient offloading mechanism in fog computing. The Journal of Supercomputing, 78(18):19435– 19458, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04626-w.

Kang:2022:PAO

[1754] Shinjin Kang and Soo Kyun Kim. Pleasure-arousal-outlier model for quantitative evaluation of game experiences. The Journal of Supercomputing, 78(18):19459–19477, December 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04636-8.

Zhou:2022:IOI

[1755] Wen Zhou, Kaixin Zhang, Zhe Ming, Jingliang Chen, and Yiwen Liang. Immune optimization inspired artificial natural killer cell earthquake prediction method. The Journal of Supercomputing, 78(18):19478– 19500, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04618-w.

Jiang:2022:AIE

[1756] Yanxu Jiang, Linfei Han, and Yifang Gao. Artificial intelligence-enabled smart city construction. The Journal of Supercomputing, 78(18):19501– 19521, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04638-6.

Teijeiro:2022:LCT

[1757] Diego Teijeiro, Margarita Amor, Ramón Doallo, Eduardo Corbelle, Juan Porta, and Jorge Parapar. Land consolidation through parcel exchange among landowners using a distributed Spark-based genetic algorithm. The Journal of Supercomputing, 78(18): 19522–19544, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04627-9.

Qu:2022:TSL

[1758] Wen Qu, Ting Zhu, Jie Liu, and Jianxin Li. A time sequence location method of long video violence based on improved C3D network. The Journal of Supercomputing, 78(18): 19545–19565, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04649-3. See cor-

rection [1811].

Yang:2022:EDO

[1759] Yang Yang, Chen Qian, Haomiao Li, Yuchao Gao, Jinran Wu, Chan-Juan Liu, and Shangrui Zhao. An efficient DBSCAN optimized by arithmetic optimization algorithm with opposition-based learning. The Journal of Supercomputing, 78(18):19566– 19604, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04634-w.

Patidar:2022:EAB

[1760] Nilesh Patidar and Namit Gupta. An extensible architecture of 32bit ALU for high-speed computing in QCA technology. The Journal of Supercomputing, 78(18):19605– 19627, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04608-y.

Tavakoli-Zaniani:2022:IHM

[1761] Maryam Tavakoli-Zaniani, Mohammad Reza Gholamian, and Seyyed Alireza Hashemi-Golpayegani. Improving heuristics miners for healthcare applications by discovering optimal dependency graphs. The Journal of Supercomputing, 78(18):19628–19661, December 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04637-7.

Kuranga:2022:PSO

[1762] Cry Kuranga, Njodzi Ranganai, and Tendai S. Muwani. Particle swarm optimization-based empirical mode decomposition predictive technique for nonstationary data. The Journal of Supercomputing, 78(18):19662– 19683, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04646-6.

You:2022:TRI

[1763] Sangseok You and Lionel P. Robert, Jr. Team robot identification theory (TRIT): robot attractiveness and team identification on performance and viability in human-robot teams. The Journal of Supercomputing, 78(18): 19684–19706, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04645-7.

Han:2022:SMM

[1764] Changnian Han, Peng Zhang, Yicong Zhu, Guojing Cong, James R. Kozloski, Chih Chieh Yang, Leili Zhang, and Yuefan Deng. Scalable multiscale modeling of platelets with 100 million particles. The Journal of Supercomputing, 78(18):19707– 19724, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04648-4.

Abasabadi:2022:HFS

[1765] Sedighe Abasabadi, Hossein Nematzadeh, Homayun Motameni, and Ebrahim Akbari. Hybrid feature selection based on SLI and genetic algorithm for microarray datasets. The Journal of Supercomputing, 78(18): 19725–19753, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04650-w.

Dhinakaran:2022:CPD

[1766] D. Dhinakaran and P. M. Joe Prathap. Correction to: Protection of data privacy from vulnerability using two-fish technique with apriori algorithm in data mining. The Journal of Supercomputing, 78(18): 19754, December 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04652-8.

Zhao:2023:KKG

[1767] Jing zhuan Zhao, Xuan Zhang, Chen Gao, Zhu dong Li, and Bao lei Wang. KG2Lib: knowledge-graph-based convolutional network for thirdparty library recommendation. The Journal of Supercomputing, 79(1):1– 26, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04603-3.

Alam:2023:EFE

[1768] Md. Golam Rabiul Alam, Abde Musavvir Khan, Myesha Farid Shejuty, Syed Ibna Zubayear, Md. Nafis Shariar, Meteb Altaf, Mohammad Mehedi Hassan, Salman A. AlQahtani, and Ahmed Alsanad. Ejection fraction estimation using deep semantic segmentation neural network. The Journal of Supercomputing, 79(1):27– 50, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04642-w.

Markovic:2023:VST

[1769] Ivica Marković and Suzana Stojković.

4-valued spectral transforms implementation on GPU with tensor cores. The Journal of Supercomputing, 79 (1):51–74, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04651-9.

Malekijou:2023:LBA

[1770] Hanieh Malekijou, Vesal Hakami, Nastooh Taheri Javan, and Amirhossein Malekijoo. Q-learning-based algorithms for dynamic transmission control in IoT equipment. The Journal of Supercomputing, 79(1):75– 108, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04643-9.

Li:2023:BBS

[1771] Dun Li, Dezhi Han, Noel Crespi, Roberto Minerva, and Kuan-Ching Li. A blockchain-based secure storage and access control scheme for supply chain finance. The Journal of Supercomputing, 79(1):109– 138, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04655-5.

Gupta:2023:FIS

[1772] Supriya Gupta, Aakanksha Sharaff, and Naresh Kumar Nagwani. Frequent item-set mining and clustering based ranked biomedical text summarization. The Journal of Supercomputing, 79 (1):139–159, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04578-1.

Bao:2023:APC

[1773] Liang Bao, Jin Yang, Zhengtong Zhang, Wenjing Liu, Junhao Chen, and Chase Wu. On accurate prediction of cloud workloads with adaptive pattern mining. The Journal of Supercomputing, 79(1):160– 187, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04647-5.

Bitalebi:2023:CAP

[1774] Hossein Bitalebi and Farshad Safaei. Criticality-aware priority to accelerate GPU memory access. The Journal of Supercomputing, 79(1):188– 213, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04657-3.

Su:2023:CIN

[1775] Yuerong Su and Weiwei Sun. Classification and interaction of new media instant music video based on deep learning under the background of artificial intelligence. The Journal of Supercomputing, 79(1):214– 242, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04672-4.

Kamanga:2023:MCD

[1776] Célestin Tshimanga Kamanga, Emmanuel Bugingo, Simon Ntumba Badibanga, and Eugène Mbuyi Mukendi. A multi-criteria decision making heuristic for workflow scheduling in cloud computing environment. The Journal of Supercomputing, 79 (1):243–264, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04677-z.

Kaur:2023:EEP

[1777] Kiranbir Kaur, Salil Bharany, Sumit Badotra, Karan Aggarwal, Anand Nayyar, and Sandeep Sharma. Energyefficient polyglot persistence database live migration among heterogeneous clouds. The Journal of Supercomputing, 79(1):265–294, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04662-

Kaur:2023:NAB

[1778] Jasleen Kaur, Rajeev Kumar, Alka Agrawal, and Raees Ahmad Khan. A neutrosophic AHP-based computational technique for security management in a fog computing network. The Journal of Supercomputing, 79 (1):295–320, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04674-2.

Yang:2023:LBD

[1779] Qinglin Yang, Taiyu Wang, Kaiming Zhu, Junbo Wang, Yu Han, and Chunhua Su. Loss-based differentiation strategy for privacy preserving of social robots. The Journal of Supercomputing, 79(1):321– 348, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04660-8. See cor-

rection [1812].

Wang:2023:ERM

[1780] Chunqiu Wang and Young Chun Ko. Emotional representation of music in multi-source data by the Internet of things and deep learning. The Journal of Supercomputing, 79 (1):349–366, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04665-3.

Ibrahim:2023:NFT

[1781] Osman Ali Sadek Ibrahim, Belal A. Hamed, and Tarek Abd El-Hafeez. A new fast technique for pattern matching in biological sequences. The Journal of Supercomputing, 79(1): 367–388, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04673-3.

Zheng:2023:IFL

[1782] Qing Zheng and Fangming Shao. Intelligent failure localization and maintenance of network based on reliability. The Journal of Supercomputing, 79 (1):389–418, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04653-7.

Xiang:2023:DFF

[1783] Lingyun Xiang, Huiqing You, Guoqing Guo, and Qian Li. Deep feature fusion for cold-start spam review detection. The Journal of Supercomputing, 79(1):419–434, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04685-z.

Imani:2023:FBW

[1784] Maryam Imani. Fuzzy-based weighting long short-term memory network for demand forecasting. The Journal of Supercomputing, 79(1):435– 460, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04659-1.

Hadian:2023:ETA

[1785] Hamid Hadian, Mohammadreza Farrokh, Mohsen Sharifi, and Ali Jafari. An elastic and traffic-aware scheduler for distributed data stream processing in heterogeneous clusters. The Journal of Supercomputing, 79 (1):461–498, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04669-z.

Yannam:2023:EAG

[1786] V Ramanjaneyulu Yannam, Jitendra Kumar, Korra Sathya Babu, and Bidyut Kumar Patra. Enhancing the accuracy of group recommendation using slope one. The Journal of Supercomputing, 79(1):499– 540, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04664-4.

Ram:2023:GNG

[1787] Pintu Kumar Ram and Pratyay Kuila. GAAE: a novel genetic algorithm based on autoencoder with ensemble classifiers for imbalanced healthcare data. The Journal of Supercomputing, 79 (1):541–572, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04679-x.

Vahdat-Nejad:2023:CAS

[1788] Hamed Vahdat-Nejad. CAMID: architectural support of middleware for multiple-domain ubiquitous computing and IoT. The Journal of Supercomputing, 79(1):573– 590, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04696-w.

Shen:2023:KES

[1789] Wenfeng Shen, Zhengsen Liu, Yunjie Tan, Zhaokai Luo, and Zhou Lei. KubeGPU: efficient sharing and isolation mechanisms for GPU resource management in container cloud. The Journal of Supercomputing, 79(1): 591–625, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04682-2.

Abdi:2023:FEF

[1790] Athena Abdi and Sina Shahoveisi.

FT-EALU: fault-tolerant arithmetic and logic unit for critical embedded and real-time systems. The Journal of Supercomputing, 79(1): 626–649, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04698-8.

Bennai:2023:EHA

[1791] Soufia Bennai, Kamal Amroun, Samir Loudni, and Abdelkader Ouali. An efficient heuristic approach combining maximal itemsets and area measure for compressing voluminous table constraints. The Journal of Supercomputing, 79(1):650– 676, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04667-1.

Zhao:2023:ISR

[1792] Weikang Zhao, KinTak U, and Huibin Luo. An image super-resolution method based on polynomial exponential function and non-uniform rectangular partition. The Journal of Supercomputing, 79(1):677– 701, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04691-1.

Yang:2023:IAM

[1793] Liwei Yang, Furong Zhu, Wencong Lai, Qiulian Zhang, Xinlai Liu, and Wenjie Zhang. Interference analysis for MIMO–OFDM based indoor visible light communication. The Journal of Supercomputing, 79(1): 702–724, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04656-4.

Truong:2023:PUA

[1794] Toan-Thinh Truong, Minh-Triet Tran, Anh-Duc Duong, Phuong-Nam Nguyen-Pham, Hoang-Anh Nguyen, and Trong-Nguyen Nguyen. Provable user authentication scheme on ECC in multi-server environment. The Journal of Supercomputing, 79(1):725– 761, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04641-x.

Segura:2023:IAR

[1795] Albert Segura, Jose Maria Arnau, and Antonio Gonzalez. Irregular accesses reorder unit: improving GPGPU memory coalescing for graph-based workloads. The Journal of Supercomputing, 79(1):762–787, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04621-1.

Imine:2023:DFD

[1796] Youcef Imine, Hicham Lakhlef, Michel Raynal, and François Taïani. DM-CSC: a fully distributed multi-coloring approach for scalable communication in synchronous broadcast networks. The Journal of Supercomputing, 79 (1):788–813, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04700-3.

Andujar:2023:EVT

[1797] Francisco J. Andújar, Miguel Sánchez de la Rosa, Jesus Escudero-Sahuquillo, and José L. Sánchez. Extending the VEF traces framework to model data center network workloads. The Journal of Supercomputing, 79(1): 814–831, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04692-0.

Singh:2023:MON

[1798] Manish Kumar Singh, Amit Choudhary, Sandeep Gulia, and Anurag Verma. Multi-objective NSGA–II optimization framework for UAV path planning in an UAV-assisted WSN. The Journal of Supercomputing, 79 (1):832–866, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04701-2.

Lu:2023:RIG

[1799] Rong Lu and Quanyu Song. Research on the improved gesture tracking algorithm in sign language synthesis. The Journal of Supercomputing, 79 (1):867–879, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04705-y.

Shan:2023:DDR

[1800] Chuanhui Shan, Ao Li, and Xiumei Chen. Deep delay rectified neural networks. The Journal of Supercomputing, 79(1):880–896, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04704-z.

Zhang:2023:DSS

[1801] Qiuyu Zhang and Zhenyu Zhao. Distributed storage scheme for encryption speech data based on blockchain and IPFS. The Journal of Supercomputing, 79(1):897–923, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04702-1.

Sirisha:2023:CVQ

[1802] D. Sirisha. Complexity versus quality: a trade-off for scheduling workflows in heterogeneous computing environments. The Journal of Supercomputing, 79(1):924–946, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04687-x.

Yu:2023:NWO

[1803] Bengong Yu and Shuwen Zhang.

A novel weight-oriented graph convolutional network for aspect-based sentiment analysis. The Journal of Supercomputing, 79(1):947– 972, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04689-9.

Kim:2023:DFA

[1804] Minju Kim, Yeonghun Shin, Wooyeon Jo, and Taeshik Shon. Digital forensic analysis of intelligent and smart IoT devices. The Journal of Supercomputing, 79(1):973–997, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04639-5.

Mazzonetto:2023:TDH

[1805] Angela Mazzonetto, Rafael Z. Frantz, Sandro Sawicki, Fabricia Roos-Frantz, and Gerson Battisti. Towards a dynamic heuristic for task scheduling in application integration platforms to handle large volumes of data. The Journal of Supercomputing, 79 (1):998–1031, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04695-x.

Seo:2023:EEU

[1806] Jiwon Seo, Inyoung Bang, Yungi Cho, Jangseop Shin, Dongil Hwang, Donghyun Kwon, Yeongpil Cho, and Yunheung Paek. Exploring effective uses of the tagged memory for reducing bounds checking overheads. The Journal of Supercomputing, 79(1): 1032–1064, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04694-y.

Nozal:2023:MLB

[1807] Raúl Nozal and Jose Luis Bosque. Mashing load balancing algorithm to boost hybrid kernels in molecular dynamics simulations. The Journal of Supercomputing, 79(1):1065– 1080, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04671-5.

El-Feshawy:2023:IFB

[1808] Somaya A. El-Feshawy, Waleed Saad, Mona Shokair, and Moawad Dessouky. IoT framework for brain tumor detection based on optimized modified ResNet 18 (OMRES). The Journal of Supercomputing, 79(1):1081– 1110, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04678-y.

Saif:2023:CICa

[1809] Mufeed Ahmed Naji Saif, S. K. Niranjan, Belal Abdullah Hezam Murshed, Fahd A. Ghanem, and Ammar Abdullah Qasem Ahmed. CSO-ILB: chicken swarm optimized intercloud load balancer for elastic containerized multi-cloud environment. The Journal of Supercomputing, 79(1): 1111–1155, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04688-w.

Paliwal:2023:CMA

[1810] Priyanka Paliwal, Julian L. Webber, Abolfazl Mehbodniya, Mohd Anul Haq, Anil Kumar, and Prem Kumar Chaurasiya. Correction to: Multiagent-based approach for generation expansion planning in isolated microgrid with renewable energy sources and battery storage. The Journal of Supercomputing, 79(1):1156– 1157, January 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04683-1.  See

[1719].

Qu:2023:CTS

[1811] Wen Qu, Ting Zhu, Jie Liu, and Jianxin Li. Correction to: A time sequence location method of long video violence based on improved C3D network. The Journal of Supercomputing, 79(1):1158, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04699-7.  See

[1758].

Yang:2023:CLB

[1812] Qinglin Yang, Taiyu Wang, Kaiming Zhu, Junbo Wang, Yu Han, and Chunhua Su. Correction to: Lossbased differentiation strategy for privacy preserving of social robots. The Journal of Supercomputing, 79(1): 1159–1160, January 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04716-9.  See

[1779].

Ouyang:2023:TRB

[1813] Yiming Ouyang, Chenglong Sun, Ruifeng Li, Qi Wang, and Jianhua Li. Transit ring: bubble flow control for eliminating inter-ring communication congestion. The Journal of Supercomputing, 79(2):1161– 1181, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04712-z.

Ren:2023:MBM

[1814] Fujia Ren, Chenhui Yang, and Y. A. Nanehkaran. MRI-based model for MCI conversion using deep zeroshot transfer learning. The Journal of Supercomputing, 79(2):1182– 1200, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04668-0.

Almanza-Ruiz:2023:PPM

[1815] Sergio H. Almanza-Ruiz, Arturo Chavoya, and Hector A. Duran-Limon. Parallel protein multiple sequence alignment approaches: a systematic literature review. The Journal of Supercomputing, 79(2):1201– 1234, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04697-9.

Chang:2023:NAM

[1816] Yu-Teng Chang and Neng-Hsun Fan.

A novel approach to market segmentation selection using artificial intelligence techniques. The Journal of Supercomputing, 79(2):1235– 1262, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04666-2.

Wang:2023:MOP

[1817] Jun Wang, Yadan Zhang, Xichao Wang, Pengjun Mao, and Bo Liu. A multi-objective parameter optimization approach to maximize lifetime of wireless sensor networks inspired by spider web. The Journal of Supercomputing, 79(2):1263– 1288, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04676-0.

Jeong:2023:CRF

[1818] Chang-Hoo Jeong and Mun Yong Yi.

Correcting rainfall forecasts of a numerical weather prediction model using generative adversarial networks. The Journal of Supercomputing, 79(2): 1289–1317, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04686-y.

Zhang:2023:SML

[1819] Ziyu Zhang, Yuelin Gao, and Eryang Guo. A supercomputing method for large-scale optimization: a feedback biogeography-based optimization with steepest descent method. The Journal of Supercomputing, 79(2):1318– 1373, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04644-8.

Narmatha:2023:OCC

[1820] C. Narmatha, P. Manimegalai, J. Krishnadass, Prajoona Valsalan, S. Manimurugan, and Mohammed Mustafa. Ovarian cysts classification using novel deep reinforcement learning with Harris hawks optimization method. The Journal of Supercomputing, 79(2): 1374–1397, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04709-8.

Nehra:2023:ERA

[1821] P. Nehra and Nishtha Kesswani. Efficient resource allocation and management by using load balanced multidimensional bin packing heuristic in cloud data centers. The Journal of Supercomputing, 79(2):1398– 1425, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04707-w.

Beiki:2023:GSD

[1822] Zohre Beiki and Ali Jahanian. Generic and scalable DNA-based logic design methodology for massive parallel computation. The Journal of Supercomputing, 79(2):1426– 1450, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04693-z.

Alaie:2023:HBO

[1823] Yeganeh Asghari Alaie, Mirsaeid Hosseini Shirvani, and Amir Masoud Rahmani. A hybrid bi-objective scheduling algorithm for execution of scientific workflows on cloud platforms with execution time and reliability approach. The Journal of Supercomputing, 79(2): 1451–1503, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04703-0.

Song:2023:MMO

[1824] Yixin Song, Junyang Yu, JinJiang Wang, and Xin He. Memory management optimization strategy in Spark framework based on less contention. The Journal of Supercomputing, 79(2): 1504–1525, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04663-5.

Chen:2023:AEC

[1825] Min Chen and Lili Zhang. Application of edge computing combined with deep learning model in the dynamic evolution of network public opinion in emergencies. The Journal of Supercomputing, 79(2):1526– 1543, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04733-8.

Peng:2023:UAI

[1826] Yanhua Peng, Yipu Yan, Guoyu Chen, Biao Feng, and Xingyu Gao. An underwater attenuation image enhancement method with adaptive color compensation and detail optimization. The Journal of Supercomputing, 79(2): 1544–1570, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04720-z.

Munoz-Montoro:2023:ADM

[1827] Antonio J. Muñoz-Montoro, Pablo Revuelta-Sanz, Damian Martínez-Muñoz, Juan Torre-Cruz, and José Ranilla. An ambient denoising method based on multi-channel nonnegative matrix factorization for wheezing detection. The Journal of Supercomputing, 79(2):1571– 1591, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04706-x.

Etefaghi:2023:AAI

[1828] Amir Etefaghi and Saeed Sharifian. AdaInNet: an adaptive inference engine for distributed deep neural networks offloading in IoT–FOG applications based on reinforcement learning. The Journal of Supercomputing, 79(2): 1592–1621, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04728-5.

Du:2023:HPH

[1829] Xiaoyu Du, Cheng Cheng, Zhijie Han, Weibei Fan, and Shuai Ding. Hamiltonian properties of HCN and BCN networks. The Journal of Supercomputing, 79(2):1622– 1653, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04723-w.

Bozorgvar:2023:PDG

[1830] Niloofar Bozorgvar, Abbas Rasoolzadegan, and Ahad Harati. Probabilistic detection of GoF design patterns. The Journal of Supercomputing, 79(2): 1654–1682, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04718-7.

Cavdar:2023:DMA

[1831] Tuğrul Çavdar, Nader Ebrahimpour, Muhammet Talha Kakız, and Faruk Baturalp Günay. Decisionmaking for the anomalies in IIoTs based on 1D convolutional neural networks and Dempster–Shafer theory (DS-1DCNN). The Journal of Supercomputing, 79(2):1683– 1704, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04739-2.

Kouchak:2023:GEG

[1832] M. M. Emadi Kouchak, F. Safaei, and M. Reshadi. Graph entropiesgraph energies indices for quantifying network structural irregularity.

The Journal of Supercomputing, 79(2): 1705–1749, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04724-9.

Jahanshahi:2023:USB

[1833] Saeedeh Jahanshahi, Amir Sabbagh Molahosseini, and Azadeh Alsadat Emrani Zarandi. uLog: a software-based approximate logarithmic number system for computations on SIMD processors. The Journal of Supercomputing, 79(2):1750– 1783, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04713-y.

Nematpour:2023:EGA

[1834] Mohammad Nematpour, Habib Izadkhah, and Farnaz Mahan. Enhanced genetic algorithm with some heuristic principles for task graph scheduling. The Journal of Supercomputing, 79(2): 1784–1813, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04684-0.

Xia:2023:CMA

[1835] Yuanqing Xia, Yufeng Zhan, Li Dai, and Yuehong Chen. A cost and makespan aware scheduling algorithm for dynamic multi-workflow in cloud environment. The Journal of Supercomputing, 79(2):1814– 1833, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04681-3.

Khan:2023:DLB

[1836] Umair Khan, Muazzam Maqsood, Saira Gillani, Mehr Yahya Durrani, Irfan Mehmood, and Sanghyun Seo. A deep learning-based framework for accurate identification and crop estimation of olive trees. The Journal of Supercomputing, 79(2):1834– 1855, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04738-3.

Pasini:2023:SPT

[1837] Massimiliano Lupo Pasini and Junqi Yin. Stable parallel training of Wasserstein conditional generative adversarial neural networks. The Journal of Supercomputing, 79(2):1856– 1876, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04721-y.

Rahmanian:2023:TCG

[1838] Mehdi Rahmanian, Ramin Nassiri, Mehran Mohsenzadeh, and Reza Ravanmehr. Test case generation for enterprise business services based on enterprise architecture design. The Journal of Supercomputing, 79(2): 1877–1907, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04742-7.

Venna:2023:DNA

[1839] Rama Krishna Reddy Venna and G. Durga Jayakumar. Design of novel area-efficient coplanar reversible arithmetic and logic unit with an energy estimation in quantumdot cellular automata. The Journal of Supercomputing, 79(2):1908– 1925, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04740-9.

Malekzadeh:2023:EER

[1840] Mina Malekzadeh. Enabling efficient and reliable IoT deployment in 5G and LTE cellular areas for optimized service provisioning. The Journal of Supercomputing, 79(2): 1926–1955, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04722-x.

Rajak:2023:NTO

[1841] Ranjit Rajak, Shrawan Kumar, Shiv Prakash, Nidhi Rajak, and Pratibha Dixit. A novel technique to optimize quality of service for directed acyclic graph (DAG) scheduling in cloud computing environment using heuristic approach. The Journal of Supercomputing, 79(2):1956– 1979, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04729-4.

Lin:2023:EDF

[1842] Pei-Chun Lin, Patrick C. K. Hung, Ying Jiang, Carolina Padilla Velasco, and Marco Antonio Martínez Cano. An experimental design for facial and color emotion expression of a social robot. The Journal of Supercomputing, 79(2): 1980–2009, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04734-7.

Xinquan:2023:SHS

[1843] Wu Xinquan, Yan Xuefeng, Li Xingchan, and Wang Yongzhen. Simulating hybrid SysML models: a model transformation approach under the DEVS framework. The Journal of Supercomputing, 79(2):2010– 2030, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04654-6.

Nuhu:2023:MLB

[1844] Abubakar Abdussalam Nuhu, Qasim Zeeshan, Babak Safaei, and Muhammad Atif Shahzad. Machine learningbased techniques for fault diagnosis in the semiconductor manufacturing process: a comparative study. The Journal of Supercomputing, 79(2): 2031–2081, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04730-x.

Ramya:2023:TAD

[1845] G. Ramya and R. Manoharan. Trafficaware dynamic controller placement in SDN using NFV. The Journal of Supercomputing, 79(2):2082– 2107, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04717-8.

Guan:2023:MCM

[1846] Runwei Guan, Ka Lok Man, Haocheng Zhao, Ruixiao Zhang, Shanliang Yao, Jeremy Smith, Eng Gee Lim, and Yutao Yue. MAN and CAT: mix attention to nn and concatenate attention to YOLO. The Journal of Supercomputing, 79(2):2108– 2136, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04726-7.

Zhang:2023:TSD

[1847] Yongliang Zhang, Yang Lu, Wuqiang Zhu, Xing Wei, and Zhen Wei. Traffic sign detection based on multiscale feature extraction and cascade feature fusion. The Journal of Supercomputing, 79(2):2137– 2152, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04670-6.

Piramuthu:2023:VAP

[1848] Otto B. Piramuthu and Matthew Caesar. VANET authentication protocols: security analysis and a proposal. The Journal of Supercomputing, 79(2): 2153–2179, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04736-5.

Baouya:2023:TCD

[1849] Abdelhakim Baouya, Otmane Ait Mohamed, and Samir Ouchani. Toward a context-driven deployment optimization for embedded systems: a product line approach. The Journal of Supercomputing, 79(2):2180– 2211, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04741-8.

Wadhwa:2023:OTS

[1850] Heena Wadhwa and Rajni Aron. Optimized task scheduling and preemption for distributed resource management in fog-assisted IoT environment. The Journal of Supercomputing, 79(2): 2212–2250, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04747-2.

Kumar:2023:HOO

[1851] Ram Kumar and S. C. Sharma. Hybrid optimization and ontology-based semantic model for efficient text-based information retrieval. The Journal of Supercomputing, 79(2):2251– 2280, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04708-9.

Bolton:2023:ABC

[1852] Sarah Bolton, Richard Dill, Michael R.

Grimaila, and Douglas Hodson. ADS-B classification using multivariate long short-term memory-fully convolutional networks and data reduction techniques. The Journal of Supercomputing, 79(2):2281– 2307, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04737-4.

Salih:2023:EBL

[1853] Shalaw Faraj Salih and Alan Anwer Abdulla. An effective bi-layer contentbased image retrieval technique. The Journal of Supercomputing, 79(2): 2308–2331, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04748-1.

Wang:2023:PSS

[1854] Ziheng Wang, Xiaoshe Dong, Yan Kang, and Heng Chen. Parallel SHA-256 on SW26010 many-core processor for hashing of multiple messages. The Journal of Supercomputing, 79(2): 2332–2355, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04750-7.

Gharib:2023:SCS

[1855] Mohammed Gharib and MohammadAmin Fazli. Secure cloud storage with anonymous deduplication using ID-based key management. The Journal of Supercomputing, 79(2): 2356–2382, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04751-6.

Xie:2023:NER

[1856] Ping Xie, Zhu Yuan, and Yu Hu.

Nscale: an efficient RAID-6 online scaling via optimizing data migration. The Journal of Supercomputing, 79(3): 2383–2403, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04752-5.

Asad:2023:GGC

[1857] Arghavan Asad and Farah Mohammadi. Godiva: green on-chip interconnection for DNNs. The Journal of Supercomputing, 79(3):2404– 2430, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04749-0.

Thilakavathy:2023:IQE

[1858] P. Thilakavathy and B. Diwan. Intelligent quotient estimation from MRI images using optimal light gradient boosting machine. The Journal of Supercomputing, 79(3):2431– 2450, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04711-0.

Cui:2023:DEC

[1859] Yu Cui, Yang Zhang, Xia Li, and Shunfu Jin. A dynamic energy conservation scheme with dual-rate adjustment and semi-sleep mode in cloud system. The Journal of Supercomputing, 79(3):2451– 2487, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04715-w.

Gowthami:2023:NAT

[1860] V. Gowthami, K. Bhoopathy Bagan, and S. Ewins Pon Pushpa. A novel approach towards high-performance image compression using multilevel wavelet transformation for heterogeneous datasets. The Journal of Supercomputing, 79(3):2488–2518, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04744-5.

NasehiMoghaddam:2023:RAI

[1861] Saeed NasehiMoghaddam, Mohammad Fathian, and Babak Amiri. Riskaverse influence maximization. The Journal of Supercomputing, 79(3): 2519–2569, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04731-w.

Bian:2023:CNC

[1862] Genqing Bian, Mingxuan Song, and Bilin Shao. Certificateless network coding scheme from certificateless public auditing protocol. The Journal of Supercomputing, 79(3):2570– 2602, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04719-6.

Xu:2023:FTQ

[1863] Heyang Xu, Sen Xu, Wei Wei, and Naixuan Guo. Fault tolerance and quality of service aware virtual machine scheduling algorithm in cloud data centers. The Journal of Supercomputing, 79(3):2603– 2625, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04760-5.

Albishari:2023:DLB

[1864] Mohammed Albishari, Mingchu Li, Runfa Zhang, and Esmail Almosharea. Deep learning-based early stage detection (DL-ESD) for routing attacks in Internet of Things networks. The Journal of Supercomputing, 79(3): 2626–2653, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04753-4.

Wen:2023:PEM

[1865] Yean-Fu Wen and Chen-Min Hsu. A performance evaluation of modular functions and state databases for Hyperledger Fabric blockchain systems. The Journal of Supercomputing, 79(3): 2654–2690, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04762-3.

Yu:2023:GWO

[1866] Xianrui Yu, Qiuhong Zhao, Qi Lin, and Tongyu Wang. A grey wolf optimizerbased chaotic gravitational search algorithm for global optimization. The Journal of Supercomputing, 79(3): 2691–2739, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04754-3.

Sepahvand:2023:SSP

[1867] Abdolah Sepahvand and Mohammadreza Razzazi. Spanning simple path inside a simple polygon. The Journal of Supercomputing, 79(3): 2740–2766, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04765-0.

Tran:2023:AAC

[1868] Nghi C. Tran, Jian-Hong Wang, Toan H. Vu, Tzu-Chiang Tai, and Jia-Ching Wang. Anti-aliasing convolution neural network of finger vein recognition for virtual reality (VR) human-robot equipment of metaverse. The Journal of Supercomputing, 79(3): 2767–2782, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04680-4.

Park:2023:CIG

[1869] Jung-Heum Park and Hyeong-Seok Lim. Characterization of interval graphs that are paired 2disjoint path coverable. The Journal of Supercomputing, 79(3):2783– 2800, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04768-x.

Shirmohammadi:2023:RNS

[1870] Zahra Shirmohammadi, Yassin Allivand, Fereshte Mozafari, Ahmad Patooghy, Mona Jalal, and Sanaz Kazemi Abharian. ReNo: novel switch architecture for reliability improvement of NoCs. The Journal of Supercomputing, 79(3):2801–2818, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04732-9. See cor-

rection [1942].

Liu:2023:DED

[1871] Zizhen Liu, Si Chen, Jing Ye, Junfeng Fan, Huawei Li, and Xiaowei Li. DHSA: efficient doubly homomorphic secure aggregation for crosssilo federated learning. The Journal of Supercomputing, 79(3):2819– 2849, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04745-4.

Ahrabi:2023:HMB

[1872] Sima Sarv Ahrabi, Alireza Momenzadeh, Enzo Baccarelli, Michele Scarpiniti, and Lorenzo Piazzo. How much BiGAN and CycleGAN-learned hidden features are effective for COVID-19 detection from CT images? A comparative study. The Journal of Supercomputing, 79(3): 2850–2881, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04775-y.

Ragesh:2023:TBS

[1873] G. K. Ragesh and Ajay Kumar.

Trust-based secure routing and message delivery protocol for signal processing attacks in IoT applications. The Journal of Supercomputing, 79(3): 2882–2909, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04766-z.

Wu:2023:LTA

[1874] Han Wu, Jiahao Nie, Ziming Zhu, Zhiwei He, and Mingyu Gao. Leveraging temporal-aware fine-grained features for robust multiple object tracking. The Journal of Supercomputing, 79(3): 2910–2931, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04776-x.

Xiao:2023:EES

[1875] Ling Xiao, Beiji Zou, Chengzhang Zhu, and Fanbo Nie. ESDedup: An efficient and secure deduplication scheme based on data similarity and blockchain for cloud-assisted medical storage systems. The Journal of Supercomputing, 79(3): 2932–2960, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04746-3.

Dizaji:2023:WGA

[1876] S. Haleh S. Dizaji, Saeid Pashazadeh, and Javad Musevi Niya. Wasserstein generative adversarial networks for modeling marked events. The Journal of Supercomputing, 79(3): 2961–2983, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04781-0.

Bevara:2023:DEQ

[1877] Vasudeva Bevara, Syed Alihussain, P. N. S. B. S. V. Prasad, and Pradyut K. Sanki. Design of an efficient QCA-based median filter with energy dissipation analysis. The Journal of Supercomputing, 79(3):2984– 3004, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04780-1.

Liu:2023:PPE

[1878] Zhenpeng Liu, Jingyi Wang, Zilin Gao, and Jianhang Wei. Privacy-preserving edge computing offloading scheme based on whale optimization algorithm. The Journal of Supercomputing, 79(3): 3005–3023, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04756-1.

Rajasekar:2023:LRG

[1879] P. Rajasekar, H. Mangalam, and C. S. Subash Kumar. Logic realization of Galois field for AES SBOX using quantum dot cellular automata. The Journal of Supercomputing, 79(3): 3024–3054, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04779-8.

Dinachali:2023:PAO

[1880] Bijan Pourghorbani Dinachali, Sam Jabbehdari, and Hamid Haj Seyyed Javadi. A pricing approach for optimal use of computing resources in cloud federation. The Journal of Supercomputing, 79(3):3055– 3094, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04725-8.

Li:2023:IMM

[1881] Peng Li and Rui Hou. Int-Monitor: a model triggered hardware trojan in deep learning accelerators. The Journal of Supercomputing, 79(3): 3095–3111, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04759-y.

Baysal:2023:BTA

[1882] Merve Vildan Baysal, Özden Özcan-Top, and Aysu Betin-Can. Blockchain technology applications in the health domain: a multivocal literature review. The Journal of Supercomputing, 79(3): 3112–3156, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04772-1.

Xie:2023:NIM

[1883] Shanshan Xie, Yan Zhang, Danjv Lv, Xu Chen, Jing Lu, and Jiang Liu. A new improved maximal relevance and minimal redundancy method based on feature subset. The Journal of Supercomputing, 79(3):3157– 3180, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04763-2.

Kumar:2023:HRH

[1884] P. Hari Kumar and G. S. AnandhaMala. HMAC-R: Hash-based message authentication code and Rijndaelbased multilevel security model for data storage in cloud environment. The Journal of Supercomputing, 79(3): 3181–3209, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04714-x.

Chen:2023:PLL

[1885] Junde Chen, Adnan Zeb, Yuandong Sun, and Defu Zhang. A power line loss analysis method based on boost clustering. The Journal of Supercomputing, 79(3):3210– 3226, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04777-w.

Zhang:2023:NII

[1886] Xi Zhang, Yongxiong Liu, Ting Chen, and Tao Gao. Noise-insensitive image representation via multiple extended LDB and class supervised intelligent coordination feature selection. The Journal of Supercomputing, 79(3): 3227–3256, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04658-2.

Ba:2023:SSC

[1887] Lina Ba, Hailun Wu, and Heping Zhang. Star-structure connectivity of folded hypercubes and augmented cubes. The Journal of Supercomputing, 79(3):3257– 3276, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04758-z. See cor-

rection [1986].

allahMottaki:2023:EHG

[1888] Nemat allah Mottaki, Homayun Motameni, and Hosein Mohamadi. An effective hybrid genetic algorithm and tabu search for maximizing network lifetime using coverage sets scheduling in wireless sensor networks. The Journal of Supercomputing, 79(3): 3277–3297, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04710-1.

ElMansoum:2023:VST

[1889] Imane El Mansoum, Laaziz Lahlou, Fawaz A. Khasawneh, Nadjia Kara, and Claes Edstrom. VALKYRIE: a suite of topology-aware clustering approaches for cloud-based virtual network services. The Journal of Supercomputing, 79(3):3298– 3328, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04786-9.

Kumar:2023:CDC

[1890] Sanjay Kumar, Abhishek Mallik, and Sandeep Singh Sengar. Community detection in complex networks using stacked autoencoders and crow search algorithm. The Journal of Supercomputing, 79(3):3329– 3356, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04767-y.

Abad:2023:HAT

[1891] Sudeh Shirkavand Saleh Abad and Mohammad Hossein Moaiyeri. Hardwareaccuracy trade-offs for error-resilient applications using an ultra-efficient hybrid approximate multiplier. The Journal of Supercomputing, 79(3): 3357–3372, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04789-6.

Tao:2023:EII

[1892] Zhifu Tao, Bingxin Yao, and Jiaming Zhu. Exploring interval implicitization in real-valued time series classification and its applications. The Journal of Supercomputing, 79(3): 3373–3391, February 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04792-x.

Douiba:2023:IAD

[1893] Maryam Douiba, Said Benkirane, Azidine Guezzaz, and Mourade Azrour. An improved anomaly detection model for IoT security using decision tree and gradient boosting. The Journal of Supercomputing, 79(3):3392– 3411, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04783-y.

Chang:2023:MFM

[1894] Wenbing Chang, Ruowen Li, Yu Fu, Yiyong Xiao, and Shenghan Zhou. A multistep forecasting method for online car-hailing demand based on wavelet decomposition and deep Gaussian process regression. The Journal of Supercomputing, 79(3):3412– 3436, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04773-0.

Dogani:2023:MWR

[1895] Javad Dogani, Farshad Khunjush, Mohammad Reza Mahmoudi, and Mehdi Seydali. Multivariate workload and resource prediction in cloud computing using CNN and GRU by attention mechanism. The Journal of Supercomputing, 79(3):3437– 3470, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04782-z.

Park:2023:DDC

[1896] Andrew T. Park, Nathaniel Peck, Richard Dill, Douglas D. Hodson, Michael R. Grimaila, and Wayne C. Henry. Distribution of DDS-cerberus authenticated facial recognition streams. The Journal of Supercomputing, 79(3):3471– 3488, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04771-2.

Park:2023:ISB

[1897] Sungwoo Park, Jaeuk Moon, Seongkuk Cho, and Eenjun Hwang. Instance segmentation-based review photo validation scheme. The Journal of Supercomputing, 79(3):3489– 3510, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04784-x.

Safiri:2023:LBO

[1898] Saadat Safiri and Amirhossein Nikoofard. Ladybug Beetle Optimization algorithm: application for real-world problems. The Journal of Supercomputing, 79(3):3511– 3560, February 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04755-2.

Zhai:2023:MCA

[1899] Rui Zhai, Libo Zhang, Yingqi Wang, Yalin Song, and Junyang Yu. A multichannel attention graph convolutional neural network for node classification. The Journal of Supercomputing, 79 (4):3561–3579, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04778-9.

Mefgouda:2023:NNI

[1900] Brahim Mefgouda and Hanen Idoudi.

New network interface selection based on MADM and multi-objective whale optimization algorithm in heterogeneous wireless networks. The Journal of Supercomputing, 79(4):3580– 3615, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04791-y.

Park:2023:QDC

[1901] Andrew T. Park, Nathaniel Peck, Richard Dill, Douglas D. Hodson, Michael R. Grimaila, and Wayne C. Henry. Quantifying DDS-cerberus network control overhead. The Journal of Supercomputing, 79(4):3616– 3642, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04770-3.

Adarsh:2023:LLH

[1902] Abhinav Adarsh, Shashwat Pathak, Digvijay Singh Chauhan, and Basant Kumar. Low-latency and high–reliability FBMC modulation scheme using optimized filter design for enabling NextG real-time smart healthcare applications. The Journal of Supercomputing, 79(4):3643– 3665, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04799-4.

Rani:2023:RAW

[1903] Radha Rani and Dharmendra Prasad Mahato. A randomized algorithm for the wait-free consensus problem. The Journal of Supercomputing, 79 (4):3666–3690, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04774-z.

Dhal:2023:AOB

[1904] Krishna Gopal Dhal, Arunita Das, Swarnajit Ray, Rebika Rai, and Tarun Kumar Ghosh. Archimedes optimizer-based fast and robust fuzzy clustering for noisy image segmentation. The Journal of Supercomputing, 79(4):3691–3730, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04769-w.

Verma:2023:HLD

[1905] Richa Verma and Shalini Chandra.

HBI-LB: A dependable fault-tolerant load balancing approach for fog based Internet-of-Things environment. The Journal of Supercomputing, 79(4): 3731–3749, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04797-6.

Singh:2023:PPM

[1906] Ashutosh Kumar Singh and Jatinder Kumar. A privacy-preserving multidimensional data aggregation scheme with secure query processing for smart grid. The Journal of Supercomputing, 79(4):3750–3770, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04794-9.

Richa:2023:HLP

[1907] Majdi Richa, Jean-Christophe Prévotet, Mickaël Dardaillon, Mohamad Mroué, and Abed Ellatif Samhat. High-level power estimation techniques in embedded systems hardware: an overview.

The Journal of Supercomputing, 79 (4):3771–3790, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04798-5.

Gui:2023:FPI

[1908] Weixia Gui, Fulai Pan, Dandan Zhu, and Feng Li. Faulty processor identification for a multiprocessor system under the malek model using an improved binary bat algorithm. The Journal of Supercomputing, 79 (4):3791–3820, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04790-z.

Chen:2023:RUC

[1909] Yanping Chen, Shengsheng He, Xiaomin Jin, Zhongmin Wang, Fengwei Wang, and Lei Chen. Resource utilization and cost optimization oriented container placement for edge computing in industrial internet. The Journal of Supercomputing, 79 (4):3821–3849, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04801-z.

Pramanik:2023:CES

[1910] Amit Kumar Pramanik, Mahabub Hasan Mahalat, Jayanta Pal, Seyed-Sajad Ahmadpour, and Bibhash Sen. Costeffective synthesis of QCA logic circuit using genetic algorithm. The Journal of Supercomputing, 79(4): 3850–3877, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04757-0.

Pan:2023:DRA

[1911] Wenjun Pan and Lin Miao. Dynamics and risk assessment of a remanufacturing closed-loop supply chain system using the Internet of Things and neural network approach. The Journal of Supercomputing, 79(4): 3878–3901, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04727-6.

Palit:2023:PAL

[1912] Sudip Kumar Palit, Mohuya Chakraborty, and Subhalaxmi Chakraborty. Performance analysis of 5GMAKA: lightweight mutual authentication and key agreement scheme for 5G network. The Journal of Supercomputing, 79 (4):3902–3935, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04807-7.

Singh:2023:NPC

[1913] Manisha Singh, Gaurav Baranwal, and Anil Kumar Tripathi. A novel 2phase consensus with customized feedback based group decision-making involving heterogeneous decision-makers. The Journal of Supercomputing, 79 (4):3936–3973, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04796-7.

DiRocco:2023:USV

[1914] Lorenzo Di Rocco, Umberto Ferraro Petrillo, and Francesco Palini. Using software visualization to support the teaching of distributed programming. The Journal of Supercomputing, 79(4):3974–3998, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04805-9.

Ahmad:2023:IFB

[1915] Israr Ahmad, Saima Abdullah, and Adeel Ahmed. IoT-fog-based healthcare 4.0 system using blockchain technology. The Journal of Supercomputing, 79(4):3999–4020, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-022-04788-
                                        7.

Tian:2023:DDC

[1916] Junfeng Tian, Hongwei Xu, and Jin Tian. DCCP: a dependable committee consensus protocol for permissionless blockchain. The Journal of Supercomputing, 79(4):4021– 4047, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04811-x.

Tran:2023:BLH

[1917] Duy Thanh Tran, Jun-Ho Huh, and Jae-Hwan Kim. Building a Lucy hybrid model for grocery sales forecasting based on time series. The Journal of Supercomputing, 79(4): 4048–4083, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04824-6.

Qin:2023:FCF

[1918] Yugang Qin, Yurong Qian, Hongyang Wei, Yingying Fan, and Peiyun Feng. FE-CSP: a fast and efficient pedestrian detector with center and scale prediction. The Journal of Supercomputing, 79(4):4084–4104, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04815-7.

Singh:2023:SAU

[1919] Pankaj Singh and Rupali Singh. A sliced architecture using novel configurable logic modules in quantum dot cellular automata for application of field-programmable gate arrays. The Journal of Supercomputing, 79(4):4105–4125, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04812-w.

Pai:2023:TED

[1920] Kung-Jui Pai, Ro-Yu Wu, Sheng-Lung Peng, and Jou-Ming Chang. Three edge-disjoint Hamiltonian cycles in crossed cubes with applications to fault-tolerant data broadcasting. The Journal of Supercomputing, 79 (4):4126–4145, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04825-5.

Jeong:2023:ELC

[1921] Young-Sang Jeong and Nam-Wook Cho. Evaluation of e-learners’ concentration using recurrent neural networks. The Journal of Supercomputing, 79(4):4146–4163, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04804-w.

Zhang:2023:DAU

[1922] Jing Zhang, Qihan Huang, Jian-Yu Hu, and Xiu-Cai Ye. Dimension-aware under spatiotemporal constraints: an efficient privacy-preserving framework with peak density clustering. The Journal of Supercomputing, 79(4): 4164–4191, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04826-4.

Gong:2023:SOG

[1923] Hao Gong, Ying Liu, Xiaoying Chen, and Cheng Wang. Scene optimization of GPU-based back-projection algorithm. The Journal of Supercomputing, 79(4):4192–4214, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04785-w.

Aghamohammadpour:2023:ATH

[1924] Ali Aghamohammadpour, Ebrahim Mahdipour, and Iman Attarzadeh. Architecting threat hunting system based on the DODAF framework. The Journal of Supercomputing, 79 (4):4215–4242, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04808-6.

Zhao:2023:ICD

[1925] Yanjun Zhao, Saeid Belkasim, and Guillaume Aubry. Image compression and denoising using multiresolution region-based image description scheme. The Journal of Supercomputing, 79 (4):4243–4265, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04806-8.

Amnyieh:2023:CRN

[1926] Zahra Heydaran Daroogheh Amnyieh, Seyed Mohammad Jalal Rastegar Fatemi, Maryam Rastgarpour, and Golnaz Aghaee Ghazvini. CNN-RDM: a new image processing model for improving the structure of deep learning based on representational dissimilarity matrix. The Journal of Supercomputing, 79(4):4266– 4290, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04661-7.

Wang:2023:TOE

[1927] Zhongmin Wang, Yurong Ding, Xiaomin Jin, Yanping Chen, and Cong Gao. Task offloading for edge computing in industrial Internet with joint data compression and security protection. The Journal of Supercomputing, 79(4):4291–4317, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04821-9.

Elsway:2023:EAD

[1928] Ahmed A. Elsway, Ahmed M. Khedr, Oruba Alfawaz, and Walid Osamy. Energy-aware disjoint dominating setsbased whale optimization algorithm for data collection in WSNs. The Journal of Supercomputing, 79(4): 4318–4350, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04814-8.

Fu:2023:DSP

[1929] Wenlong Fu, Xing Liu, Fanwu Chu, Bailin Li, and Jiahao Gu. A disassembly sequence planning method with improved discrete grey wolf optimizer for equipment maintenance in hydropower station. The Journal of Supercomputing, 79(4):4351– 4382, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04822-8.

Liang:2023:CDC

[1930] Zongnan Liang, Jiawei Nian, Hongjin Liu, Xuru Wang, and Mengfei Yang. C-DMR: a cache-based fault-tolerant protection method for register file. The Journal of Supercomputing, 79 (4):4383–4397, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04836-2.

Pan:2023:SSC

[1931] Kaige Pan and Dongqin Cheng. Star structure connectivity of Cayley graphs generated by transposition trees. The Journal of Supercomputing, 79(4): 4398–4411, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04837-1.

Huang:2023:PMS

[1932] Jianying Huang, Seunghyeok Yang, Jinhui Li, Jeill Oh, and Hoon Kang. Prediction model of sparse autoencoder-based bidirectional LSTM for wastewater flow rate. The Journal of Supercomputing, 79(4):4412– 4435, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04827-3.

Wu:2023:SLE

[1933] Bin Wu and Sean He. Self-learning and explainable deep learning network toward the security of artificial intelligence of things. The Journal of Supercomputing, 79(4):4436– 4467, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04818-4.

Cheng:2023:CVB

[1934] Min-Yuan Cheng, Minh-Tu Cao, and Christian Kentaro Nuralim. Computer vision-based deep learning for supervising excavator operations and measuring real-time earthwork productivity. The Journal of Supercomputing, 79 (4):4468–4492, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04803-x.

Sait:2023:OFB

[1935] Sadiq M. Sait, Aiman El-Maleh, Mohammad Altakrouri, and Ahmad Shawahna. Optimization of FPGA-based CNN accelerators using metaheuristics. The Journal of Supercomputing, 79(4):4493– 4533, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04787-8.

Kalsoom:2023:SMF

[1936] Nadia Kalsoom, Asad Waqar Malik, Anis U. Rahman, Muhammad Ahsan, and Arsalan Ahmad. SUDV: Malicious fog node management framework for software update dissemination in connected vehicles. The Journal of Supercomputing, 79(4): 4534–4555, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04829-1.

Siegel:2023:RGR

[1937] Pierre Siegel, Andrei Doncescu, Vincent Risch, and Sylvain Sené. Representation of gene regulation networks by hypothesis logic-based Boolean systems. The Journal of Supercomputing, 79(4):4556–4581, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04809-5.

Zhao:2023:EPM

[1938] Kaiyi Zhao, Li Li, Zeqiu Chen, Jiayao Li, Ruizhi Sun, and Gang Yuan. An efficient parallelization method of Dempster–Shafer evidence theory based on CUDA. The Journal of Supercomputing, 79(4):4582– 4601, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04810-y.

Liu:2023:AAI

[1939] Qiuyan Liu, Xuan Zhao, and Kaihan Shi. The analysis of agricultural Internet of things product marketing by deep learning. The Journal of Supercomputing, 79(4):4602– 4621, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04817-5.

Dezhkam:2023:BBC

[1940] Arsalan Dezhkam, Mohammad Taghi Manzuri, Ahmad Aghapour, Afshin Karimi, Ali Rabiee, and Shervin Manzuri Shalmani. A Bayesian-based classification framework for financial time series trend prediction. The Journal of Supercomputing, 79(4):4622– 4659, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04834-4.

Lima:2023:PMA

[1941] André Luis da Cunha Dantas Lima, Vítor Moraes Aranha, and Erick Giovani Sperandio Nascimento. Predictive maintenance applied to mission critical supercomputing environments: remaining useful life estimation of a hydraulic cooling system using deep learning. The Journal of Supercomputing, 79(4):4660– 4684, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04833-5.

Shirmohammadi:2023:CRN

[1942] Zahra Shirmohammadi, Yassin Allivand, Fereshte Mozafari, Ahmad Patooghy, Mona Jalal, and Sanaz Kazemi Abharian. Correction to: ReNo: novel switch architecture for reliability improvement of NoCs. The Journal of Supercomputing, 79(4): 4685, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04800-0.  See

[1870].

Justus:2023:RAS

[1943] J. Jean Justus and M. Anuradha. Resource allocation scheme for CCRN using hybrid Giza Pyramids constructionbased complex-valued satin bowerbird optimization. The Journal of Supercomputing, 79(5):4687– 4712, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04761-4.

Shanker:2023:RRE

[1944] Ravi Shanker, Heet Sankesara, Surendra Nagar, and Mahua Bhattacharya. RESPNet: resource-efficient and structure-preserving network for deformable image registration. The Journal of Supercomputing, 79(5):

4713–4736, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04840-6.

Gaudiani:2023:CMA

[1945] Adriana Gaudiani, Alvaro Wong, Emilio Luque, and Dolores Rexachs. A computational methodology applied to optimize the performance of a river model under uncertainty conditions. The Journal of Supercomputing, 79 (5):4737–4759, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04816-6.

Wang:2023:FGF

[1946] Zhehe Wang, Yunfu Shen, Shuang Li, and Shuxin Wang. A fine-grained fast parallel genetic algorithm based on a ternary optical computer for solving traveling salesman problem. The Journal of Supercomputing, 79 (5):4760–4790, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04813-9.

Benachour:2023:FPA

[1947] Amira Benachour, Saïd Yahiaoui, Didier El Baz, Nadia Nouali-Taboudjemat, and Hamamache Kheddouci. Fast parallel algorithms for finding elementary circuits of a directed graph: a GPU-based approach. The Journal of Supercomputing, 79(5):4791– 4819, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04835-3.

Tsmots:2023:DGM

[1948] Ivan Tsmots, Vasyl Teslyuk, Natalia Kryvinska, Oleksa Skorokhoda, and Iryna Kazymyra. Development of a generalized model for parallelstreaming neural element and structures for scalar product calculation devices. The Journal of Supercomputing, 79(5):4820–4846, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04838-0.

Priyadarshi:2023:DCO

[1949] Rahul Priyadarshi and Bharat Gupta. 2-D coverage optimization in obstaclebased FOI in WSN using modified PSO. The Journal of Supercomputing, 79(5):4847–4869, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04832-6.

Wang:2023:MGA

[1950] Jingjing Wang, Hongji Meng, Jian Yang, and Zhi Xie. Multi-GPU accelerated cellular automaton model for simulating the solidification structure of continuous casting bloom. The Journal of Supercomputing, 79 (5):4870–4894, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04839-z.

Elshahed:2023:PST

[1951] Eman M. Elshahed, Randa M. Abdelmoneem, Eman Shaaban, Hayam A.

Elzahed, and Shahinaz M. Al-Tabbakh. Prioritized scheduling technique for healthcare tasks in cloud computing. The Journal of Supercomputing, 79 (5):4895–4916, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04823-7.

BenHalima:2023:MKP

[1952] Riadh Ben Halima, Marwa Hachicha, Ahmed Jemal, and Ahmed Hadj Kacem. MAPE-K patterns for selfadaptation in cyber-physical systems. The Journal of Supercomputing, 79 (5):4917–4943, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04828-2.

Cao:2023:CCD

[1953] Lifeng Cao, Shoucai Zhao, Zhen-Sheng Gao, and Xuehui Du. Crosschain data traceability mechanism for cross-domain access. The Journal of Supercomputing, 79(5):4944– 4961, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04793-w.

Yadav:2023:PPS

[1954] Vijay Kumar Yadav, Nitish Andola, Shekhar Verma, and S. Venkatesan. PSCLS: provably secure certificateless signature scheme for IoT device on cloud. The Journal of Supercomputing, 79(5):4962–4982, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04795-8.

Shih:2023:IVN

[1955] Wen-Chung Shih, Chao-Tung Yang, Cheng-Tian Jiang, and Endah Kristiani. Implementation and visualization of a netflow log data lake system for cyberattack detection using distributed deep learning. The Journal of Supercomputing, 79(5): 4983–5012, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04802-y.

Gudakahriz:2023:OTS

[1956] Sajjad Jahanbakhsh Gudakahriz, Amir Masoud Eftekhari Moghadam, and Fariborz Mahmoudi. Opinion texts summarization based on texts concepts with multi-objective pruning approach. The Journal of Supercomputing, 79 (5):5013–5036, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04842-4.

Cheng:2023:ESM

[1957] Dun-Wei Cheng, Jo-Yi Chang, Chen-Yen Lin, Limei Lin, Yanze Huang, Krishnaiyan Thulasiraman, and Sun-Yuan Hsieh. Efficient survivable mapping algorithm for logical topology in IP-over-WDM optical networks against node failure. The Journal of Supercomputing, 79(5): 5037–5063, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04841-5.

Huang:2023:SMM

[1958] Hai Huang and Haoran Zong. Secure matrix multiplication based on fully homomorphic encryption. The Journal of Supercomputing, 79(5): 5064–5085, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04850-4.

Yu:2023:AGA

[1959] Xue Yu. The appeal of green advertisements on consumers’ consumption intention based on low-resource machine translation. The Journal of Supercomputing, 79(5):5086– 5108, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04846-0.

Gupta:2023:RON

[1960] Rohit Kumar Gupta, Shashwat Mahajan, and Rajiv Misra. Resource orchestration in network slicing using GAN-based distributional deep Qnetwork for industrial applications. The Journal of Supercomputing, 79 (5):5109–5138, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04867-9.

Chehreghani:2023:UAS

[1961] Mostafa Haghir Chehreghani. On using affine sketches for multipleresponse dynamic graph regression. The Journal of Supercomputing, 79 (5):5139–5153, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04865-x.

Chen:2023:MLW

[1962] Zhiqiang Chen, Meng Han, Hongxin Wu, Muhang Li, and Xilong Zhang. A multi-level weighted concept drift detection method. The Journal of Supercomputing, 79(5):5154– 5180, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04864-y.

Zabin:2023:HDT

[1963] Mahe Zabin, Ho-Jin Choi, and Jia Uddin. Hybrid deep transfer learning architecture for industrial fault diagnosis using Hilbert transform and DCNN–LSTM. The Journal of Supercomputing, 79(5):5181– 5200, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04830-8.

Yang:2023:RDD

[1964] Peng Yang, Juncheng Leng, Guangzhen Zhao, Wenjun Li, and Haisheng Fang. Rumor detection driven by graph attention capsule network on dynamic propagation structures. The Journal of Supercomputing, 79(5):5201– 5222, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04831-7.

Lyu:2023:PRI

[1965] Chunyan Lyu, Tong Xu, Kejun Wang, and Jing Chen. Person re-identification based on human semantic parsing and message passing. The Journal of Supercomputing, 79(5):5223– 5247, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04866-w.

Esmaeilyfard:2023:IMD

[1966] Rasool Esmaeilyfard and Mahsa Moghisi. An incentive mechanism design for multitask and multipublisher mobile crowdsensing environment. The Journal of Supercomputing, 79(5):5248–5275, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04852-2.

Chang:2023:SIR

[1967] Shih-Hao Chang, Chih-Hsien Hsia, and Wei-Zhi Hong. A secured Internet of Robotic Things (IoRT) for longterm care services in a smart building. The Journal of Supercomputing, 79(5):5276–5290, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04845-1. See cor-

rection [2037].

Madani:2023:KOT

[1968] Mirmorsal Madani, Homayun Motameni, and Hosein Mohamadi. KN-NGAN: an oversampling technique for textual imbalanced datasets. The Journal of Supercomputing, 79(5):

5291–5326, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04851-3.

Tsai:2023:UIP

[1969] Ming-Fong Tsai and Yi-Hong Wu. User intent prediction search engine system based on query analysis and image recognition technologies. The Journal of Supercomputing, 79(5): 5327–5359, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04874-w.

Cheng:2023:MBE

[1970] Jing-Ru C. Cheng, Corwin Stanford, Steven R. Glandon, Anthony L. Lam, and Warren R. Williams. Macro benchmarking edge devices using enhanced super-resolution generative adversarial networks (ESRGANs). The Journal of Supercomputing, 79(5): 5360–5373, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04819-3.

He:2023:DAC

[1971] Dun He, Jue Chen, and Xihe Qiu. A density algorithm for controller placement problem in software defined wide area networks. The Journal of Supercomputing, 79(5):5374– 5402, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04873-x.

Bai:2023:WSK

[1972] Qingchun Bai, Jun Xiao, and Jie Zhou. A weakly supervised knowledge attentive network for aspectlevel sentiment classification. The Journal of Supercomputing, 79(5): 5403–5420, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04820-w.

Do:2023:AGC

[1973] Cong Thuan Do, Cheol Hong Kim, and Sung Woo Chung. Aggressive GPU cache bypassing with monolithic 3D-based NoC. The Journal of Supercomputing, 79(5):5421– 5442, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04878-6. See cor-

rection [2036].

Amanov:2023:AEM

[1974] Anuarbek Amanov, Akbar Majidi, Nazila Jahnabakhsh, and Aydın Çetin. Adjusting ECN marking threshold in multi-queue DCNs with deep learning. The Journal of Supercomputing, 79(5):5443–5468, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04893-7.

Ding:2023:CNC

[1975] Xiaoyao Ding, Gang Zhou, Jicang Lu, and Taojie Zhu. CLGR-Net: a collaborative local-global reasoning network for document-level relation extraction. The Journal of Supercomputing, 79 (5):5469–5485, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04875-9.

Najjar-Ghabel:2023:DHW

[1976] Samad Najjar-Ghabel, Leili Farzinvash, and Seyed Naser Razavi. Data harvesting in wireless sensor networks using mobile sinks under realworld circumstances. The Journal of Supercomputing, 79(5):5486– 5515, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04888-4.

Mewada:2023:SAH

[1977] Arvind Mewada and Rupesh Kumar Dewang. SA-ASBA: a hybrid model for aspect-based sentiment analysis using synthetic attention in pre-trained language BERT model with extreme gradient boosting. The Journal of Supercomputing, 79(5):5516–5551, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04881-x.

Kristiani:2023:FSR

[1978] Endah Kristiani, Yi-Chun Chen, Chao-Tung Yang, and Chia-Hsin Li. Flame and smoke recognition on smart edge using deep learning. The Journal of Supercomputing, 79(5):5552– 5575, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04884-8.

Chen:2023:HHO

[1979] Lei Chen, Na Song, and Yunpeng Ma.

Harris hawks optimization based on global cross-variation and tent mapping. The Journal of Supercomputing, 79(5):5576–5614, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04869-7.

Achroufene:2023:RBG

[1980] Achour Achroufene. RSSI-based geometric localization in wireless sensor networks. The Journal of Supercomputing, 79(5):5615–5642, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04887-
5.

Brisaboa:2023:STE

[1981] Nieves R. Brisaboa, Ana Cerdeira-Pena, Guillermo de Bernardo, Antonio Fariña, and Gonzalo Navarro. Space/time-efficient RDF stores based on circular suffix sorting. The Journal of Supercomputing, 79(5):5643– 5683, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04890-w.

Byun:2023:TSF

[1982] Jin Wook Byun. Towards serverless fast one round authentication with two mobile end devices. The Journal of Supercomputing, 79(5): 5684–5704, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04891-9.

Chen:2023:CCH

[1983] Bo Chen, Hang Tao, Xuehuan Jiang, Yufeng Chen, Xiue Gao, Panling Jiang, and Rui Tong. A command-and-control hypernetwork modeling approach based on hierarchy-betweenness edge-linking strategy. The Journal of Supercomputing, 79(5):5705– 5729, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04876-8.

Horng:2023:ISO

[1984] Shih-Cheng Horng and Shieh-Shing Lin. Incorporate seagull optimization into ordinal optimization for solving the constrained binary simulation optimization problems. The Journal of Supercomputing, 79(5):5730– 5758, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04880-y.

Tesser:2023:CHS

[1985] Rafael Keller Tesser and Edson Borin. Containers in HPC: a survey. The Journal of Supercomputing, 79(5): 5759–5827, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04848-y.

Ba:2023:CSS

[1986] Lina Ba, Hailun Wu, and Heping Zhang. Correction to: Starstructure connectivity of folded hypercubes and augmented cubes. The Journal of Supercomputing, 79(5): 5828, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04843-3.  See

[1887].

Khanna:2023:PCDa

[1987] Gaurav Khanna, Sanjay K. Chaturvedi, and Mohamed Othman. Publisher correction: On design and performance analysis of improved shuffle exchange gamma interconnection network layouts. The Journal of Supercomputing, 79(5):5829–5830, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04844-2.

AlZubi:2023:RNB

[1988] Ahmad Ali AlZubi. Retraction note:

Big data analytic diabetics using map reduce and classification techniques. The Journal of Supercomputing, 79(5):5831, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04858-w. See [217].

Karthikeyan:2023:RNE

[1989] K. Karthikeyan, R. Sunder, K. Shankar, S. K. Lakshmanaprabu, V. Vijayakumar, Mohamed Elhoseny, and Gunasekaran Manogaran. Retraction note: Energy consumption analysis of virtual machine migration in cloud using hybrid swarm optimization (ABC-BA). The Journal of Supercomputing, 79(5):5832–5833, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04860-2. See [208].

Kanisha:2023:RNS

[1990] B. Kanisha, V. Mahalakshmi, M. Baskar, K. Vijaya, and P. Kalyanasundaram. Retraction note: Smart communication using tri-spectral sign recognition for hearing-impaired people. The Journal of Supercomputing, 79 (5):5834–5835, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04859-9.  See

[1169].

Hui:2023:RNM

[1991] Qiuli Hui. Retraction note: Motion video tracking technology in sports training based on mean–shift algorithm. The Journal of Supercomputing, 79(5):5836, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04863-z. See [5].

Kim:2023:RNS

[1992] Hyun-Woo Kim, Jungho Kang, and Young-Sik Jeong. Retraction note: Simulator considering modeling and performance evaluation for high-performance computing of collaborative-based mobile cloud infrastructure. The Journal of Supercomputing, 79(5): 5837, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04857-x.

Cho:2023:GCF

[1993] Hojin Cho and Myungsun Kim. gCFS:

completely fair scheduling on multiple GPUs for improved multi-DNN execution in terms of performance isolation. The Journal of Supercomputing, 79(6):5851–5877, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04901-w.

Yan:2023:IHM

[1994] Zheping Yan, Jinyu Yan, Yifan Wu, and Chao Zhang. An improved hybrid mayfly algorithm for global optimization. The Journal of Supercomputing, 79(6):5878–5919, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04883-9.

Xu:2023:FSF

[1995] Jiuyun Xu and Xingru Zhao. FPSA-SMS: first price sealed auction-based service migration strategy in mobile edge computing. The Journal of Supercomputing, 79(6):5920– 5938, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04879-5.

Huang:2023:CAR

[1996] Fengyi Huang, Hengzhou Ye, and Wei Hao. Cost-aware resource management based on market pricing mechanisms in edge federation environments.

The Journal of Supercomputing, 79 (6):5939–5961, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04870-0.

Osamy:2023:TTA

[1997] Walid Osamy, Ahmed M. Khedr, Dilna Vijayan, and Ahmed Salim. TAC-TIRSO: trust aware clustering technique based on improved rat swarm optimizer for WSN-enabled intelligent transportation system. The Journal of Supercomputing, 79(6):5962– 6016, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04889-3.

Ali:2023:DUE

[1998] Zainab H. Ali, John F. Zaki, and Nora El-Rashidy. Dynamic urban evaluation routing protocol for enhanced vehicle ad hoc networks. The Journal of Supercomputing, 79(6): 6017–6039, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04877-7.

Lee:2023:DAM

[1999] Chang-Eun Lee, Jaeuk Baek, Jeany Son, and Young-Guk Ha. Deep AI military staff: cooperative battlefield situation awareness for commander’s decision making. The Journal of Supercomputing, 79(6):6040– 6069, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04882-w.

Qiao:2023:EFT

[2000] Hongwei Qiao, Jixiang Meng, and Eminjan Sabir. The edge faulttolerant spanning laceability of the enhanced hypercube networks. The Journal of Supercomputing, 79(6): 6070–6086, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04896-4.

Elleuch:2023:CCI

[2001] Islam Elleuch, Achraf Makni, and Rafik Bouaziz. CICAPS: a cooperative intersection collision avoidance persistent system for cooperative intersection ADAS. The Journal of Supercomputing, 79(6):6087– 6114, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04849-x.

Yang:2023:RPT

[2002] Hao Yang, Dongming Zhou, Jinde Cao, Qian Zhao, and Miao Li. Rain-Former: a pyramid transformer for single image deraining. The Journal of Supercomputing, 79(6):6115– 6140, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04895-5.

Shabir:2023:FMA

[2003] Balawal Shabir, Anis U. Rahman, Asad Waqar Malik, Rajkumar Buyya, and Muazzam A. Khan. A federated multi-agent deep reinforcement learning for vehicular fog computing. The Journal of Supercomputing, 79 (6):6141–6167, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04911-8.

Li:2023:MBS

[2004] Wei Li, Haonan Luo, and Lei Wang. Multifactorial brain storm optimization algorithm based on direct search transfer mechanism and concave lens imaging learning strategy. The Journal of Supercomputing, 79(6):6168– 6202, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04916-3.

Muniz:2023:RTM

[2005] Rubén Muñiz, Fernando Nuño, Juan Díaz, María González, Miguel J. Prieto, and Óliver Menéndez. Realtime monitoring solution with vibration analysis for industry 4.0 ventilation systems. The Journal of Supercomputing, 79(6):6203– 6227, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04897-3.

Wei:2023:SFD

[2006] Xing Wei, Ting Bai, Yan Zhai, Lei Chen, Hui Luo, Chong Zhao, and Yang Lu. Source-free domain adaptive object detection based on pseudosupervised mean teacher. The Journal of Supercomputing, 79(6):6228– 6251, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04915-4.

Pal:2023:CBP

[2007] Debranjan Pal, Md Rasid Ali, Abhijit Das, and Dipanwita Roy Chowdhury. A cluster-based practical key recovery attack on reduced-round AES using impossible-differential cryptanalysis. The Journal of Supercomputing, 79(6):6252–6289, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04872-y.

Zhang:2023:EDL

[2008] Maoyuan Zhang, Lisha Liu, Jiaxin Mi, and Xianqi Yuan. Enhanced duallevel dependency parsing for aspectbased sentiment analysis. The Journal of Supercomputing, 79(6):6290– 6308, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04898-2.

Cheng:2023:QWL

[2009] Sheng-Tzong Cheng, Chih-Wei Hsu, Gwo-Jiun Horng, and Kuan-Ting Tsai. Quadro-W learning for human behavior prediction in an evolving environment: a case study of the intelligent butler technology. The Journal of Supercomputing, 79(6):6309– 6346, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04899-1.

Han:2023:CBC

[2010] Ju Hee Han, Jee In Kim, Jang Won Suh, and Hyungseok Kim. Customizing blendshapes to capture facial details. The Journal of Supercomputing, 79(6):6347–6372, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04885-7.

Giannoula:2023:HPB

[2011] Christina Giannoula, Athanasios Peppas, Georgios Goumas, and Nectarios Koziris. High-performance and balanced parallel graph coloring on multicore platforms. The Journal of Supercomputing, 79(6): 6373–6421, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04894-6.

Liao:2023:IPM

[2012] Yanping Liao, Fan Jiang, and Jinli Wang. Intra-pulse modulation recognition of radar signals based on multifeature random matching fusion network. The Journal of Supercomputing, 79(6):6422–6451, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04902-9.

Chen:2023:PPS

[2013] Min Chen and Wenhu Du. The predicting public sentiment evolution on public emergencies under deep learning and Internet of Things. The Journal of Supercomputing, 79(6): 6452–6470, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04900-x.

Mukhopadhyay:2023:DCE

[2014] Nirmalya Mukhopadhyay and Babul P.

Tewari. Dynamic cost effective solution for efficient cloud infrastructure. The Journal of Supercomputing, 79 (6):6471–6506, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04913-6.

Wang:2023:ACO

[2015] Yiwen Wang, Hao Liu, Guiyan Ding, and Liangping Tu. Adaptive chimp optimization algorithm with chaotic map for global numerical optimization problems. The Journal of Supercomputing, 79(6):6507–6537, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04886-6.

Alagarsamy:2023:SOE

[2016] Aravindhan Alagarsamy, Sundarakannan Mahilmaran, Lakshminarayanan Gopalakrishnan, and Seok-Bum Ko. SaHNoC: an optimal energy efficient hybrid networks-on-chip architecture. The Journal of Supercomputing, 79 (6):6538–6559, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04910-9.

Liu:2023:LSM

[2017] Weiming Liu. Literature survey of multi-track music generation model based on generative confrontation network in intelligent composition. The Journal of Supercomputing, 79(6): 6560–6582, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04914-5.

Singh:2023:IMC

[2018] Tinku Singh, Riya Khanna, Satakshi, and Manish Kumar. Improved multi-class classification approach for imbalanced big data on spark. The Journal of Supercomputing, 79(6): 6583–6611, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04908-3.

Chen:2023:NMP

[2019] Tao Chen, Yong Chen, Zhicheng He, Eric Li, Chenglin Zhang, and Yuanyi Huang. A novel marine predators algorithm with adaptive update strategy. The Journal of Supercomputing, 79(6):6612–6645, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04903-8.

Khalili:2023:SSS

[2020] Ebrahim Khalili, Razieh Malekhosseini, S. Hadi Yaghoubyan, Karamollah Bagherifard, and Hamid Parvin. Sequential semi-supervised active learning model in extremely low training set (SSSAL). The Journal of Supercomputing, 79(6):6646– 6673, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04847-z.

Funika:2023:ACR

[2021] Włodzimierz Funika, Paweł Koperek, and Jacek Kitowski. Automated cloud resources provisioning with the use of the proximal policy optimization. The Journal of Supercomputing, 79 (6):6674–6704, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04924-3.

Kumar:2023:OLL

[2022] Sumit Kumar, Neera Batra, and Shrawan Kumar. Optimized localization in large-scale heterogeneous WSN. The Journal of Supercomputing, 79 (6):6705–6729, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04922-5.

Haimour:2023:ACA

[2023] Jumana Haimour and Ali Al-Haj.

An adaptive centralized authentication control method to reduce association delay in the IoT 802.11ah protocol. The Journal of Supercomputing, 79(6):6730–6755, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04919-0.

Tran:2023:NML

[2024] Duy Thanh Tran and Jun-Ho Huh.

New machine learning model based on the time factor for e-commerce recommendation systems. The Journal of Supercomputing, 79(6):6756– 6801, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04909-2.

Aldabbas:2023:EBA

[2025] Hamza Aldabbas. Efficient bandwidth allocation in SDN-based peerto-peer data streaming using machine learning algorithm. The Journal of Supercomputing, 79(6):6802– 6824, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04929-y.

Nandhini:2023:ERA

[2026] P. S. Nandhini, S. Kuppuswami, S. Malliga, and R. DeviPriya. Enhanced Rank Attack Detection Algorithm (E-RAD) for securing RPLbased IoT networks by early detection and isolation of rank attackers. The Journal of Supercomputing, 79 (6):6825–6848, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04921-6.

Qiu:2023:VNF

[2027] Rixuan Qiu, Jiawen Bao, Yuancheng Li, Xin Zhou, Liang Liang, Hui Tian, Yanting Zeng, and Jie Shi. Virtual network function deployment algorithm based on graph convolution deep reinforcement learning. The Journal of Supercomputing, 79(6): 6849–6870, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04947-w.

Sadeghi:2023:ICC

[2028] Maryam Sadeghi, Mohammad Naderi Dehkordi, Behrang Barekatain, and Naser Khani. Improve customer churn prediction through the proposed PCA-PSO-K means algorithm in the communication industry. The Journal of Supercomputing, 79(6): 6871–6888, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04907-4. See cor-

rection [2351].

Yoon:2023:WDH

[2029] Daegun Yoon, Minjoong Jeong, and Sangyoon Oh. WAVE: designing a heuristics-based three-way breadthfirst search on GPUs. The Journal of Supercomputing, 79(6):6889– 6917, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04934-1.

Khanal:2023:SLQ

[2030] Bikram Khanal, Javier Orduz, Pablo Rivas, and Erich Baker. Supercomputing leverages quantum machine learning and Grover’s algorithm. The Journal of Supercomputing, 79 (6):6918–6940, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04923-4.

Wang:2023:AIN

[2031] Wanwan Wang, Yu Duan, Longhan Cao, and Zhenghong Jiang. Application of improved Naive Bayes classification algorithm in 5G signaling analysis. The Journal of Supercomputing, 79(6):6941–6964, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04946-x.

Topcu:2023:SEV

[2032] Burak Topçu and Isil Öz. Soft error vulnerability prediction of GPGPU applications. The Journal of Supercomputing, 79(6):6965– 6990, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04933-2.

Chang:2023:TCH

[2033] Guanghui Chang, Shiyang Hu, and Haihui Huang. Two-channel hierarchical attention mechanism model for short text classification. The Journal of Supercomputing, 79(6): 6991–7013, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04950-1.

Souaidia:2023:IVA

[2034] Chouaib Souaidia, Tawfik Thelaidjia, and Salah Chenikher. Independent vector analysis based on binary grey wolf feature selection and extreme learning machine for bearing fault diagnosis. The Journal of Supercomputing, 79(6):7014–7036, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04931-4.

Safaiezadeh:2023:NHP

[2035] Behrouz Safaiezadeh, Lauri Kettunen, and Majid Haghparast. Novel highperformance QCA Fredkin gate and designing scalable QCA binary to gray and vice versa. The Journal of Supercomputing, 79(6):7037– 7060, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04939-w.

Do:2023:CAG

[2036] Cong Thuan Do, Cheol Hong Kim, and Sung Woo Chung. Correction to: Aggressive GPU cache bypassing with monolithic 3D-based NoC. The Journal of Supercomputing, 79 (6):7061, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04917-2. See cor-

rection [1973].

Chang:2023:CSI

[2037] Shih-Hao Chang, Chih-Hsien Hsia, and Wei-Zhi Hong. Correction to: A secured Internet of Robotic Things (IoRT) for long-term care services in a smart building. The Journal of Supercomputing, 79(6): 7062, April 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04927-0.  See

[1967].

Shankar:2023:RNO

[2038] K. Shankar, S. K. Lakshmanaprabu, Deepak Gupta, Andino Maseleno, and Victor Hugo C. de Albuquerque. Retraction note to: Optimal featurebased multi-kernel SVM approach for thyroid disease classification. The Journal of Supercomputing, 79(6): 7063–7064, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04941-2. See [40].

Ramanathan:2023:RNA

[2039] Shalini Ramanathan and Mohan Ramasundaram. Retraction note to: Accurate computation: COVID-19 rRT-PCR positive test dataset using stages classification through textual big data mining with machine learning. The Journal of Supercomputing, 79(6):7065, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04937-y. See [800].

Feng:2023:GSC

[2040] Junhong Feng, Jie Zhang, Xiaoshu Zhu, and Jian-Hong Wang. Gene selection and clustering of single-cell data based on Fisher score and genetic algorithm. The Journal of Supercomputing, 79(7):7067–7093, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04920-7.

Garg:2023:PEL

[2041] Aakansha Garg, Rajeev Arya, and Maheshwari Prasad Singh. Price elasticity log-log model for cost optimization in D2D underlay mobile edge computing system. The Journal of Supercomputing, 79(7):7094– 7131, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04928-z.

Jang:2023:GLB

[2042] Sung su Jang, Cheol jin Kim, Seong yeon Hwang, Myung jae Lee, and Young guk Ha. L-GAN: landmarkbased generative adversarial network for efficient face de-identification. The Journal of Supercomputing, 79(7): 7132–7159, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04954-x.

Guyeux:2023:UDS

[2043] Christophe Guyeux, Gaby Bou Tayeh, Abdallah Makhoul, Stéphane Chrétien, Julien Bourgeois, and Jacques M. Bahi. Using data science to predict firemen interventions: a case study. The Journal of Supercomputing, 79 (7):7160–7175, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04956-9.

Rodrigues:2023:FRA

[2044] Iago Richard Rodrigues, Marrone Dantas, Assis T. de Oliveira Filho, Gibson Barbosa, Daniel Bezerra, Ricardo Souza, Maria Valéria Marquezini, Patricia Takako Endo, Judith Kelner, and Djamel Sadok. A framework for robotic arm pose estimation and movement prediction based on deep and extreme learning models. The Journal of Supercomputing, 79(7):7176–7205, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04936-z.

Gao:2023:OPT

[2045] Jinsheng Gao, Xiaomin Zhu, and Runtong Zhang. Optimization of parallel test task scheduling with constraint satisfaction. The Journal of Supercomputing, 79(7):7206– 7227, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04943-0.

Zhou:2023:UNL

[2046] Jingyi Zhou, Qingfang He, Guang Cheng, and Zhiying Lin. Unionnet: lightweight deep neural network model suitable for small data sets. The Journal of Supercomputing, 79 (7):7228–7243, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04963-w.

Xu:2023:ETS

[2047] Jinyi Xu, Hao Shi, and Yixiang Chen. Efficient tasks scheduling in multicore systems integrated with hardware accelerators. The Journal of Supercomputing, 79(7):7244– 7271, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04955-w.

Sardroud:2023:GND

[2048] Asghar A. Asgharian Sardroud and Mohsen Ghasemi. The g-good-neighbor diagnosability of triangle-free graphs. The Journal of Supercomputing, 79 (7):7272–7285, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04942-1.

Egrioglu:2023:NGA

[2049] Erol Egrioglu, Crina Grosan, and Eren Bas. A new genetic algorithm method based on statistical-based replacement for the training of multiplicative neuron model artificial neural networks. The Journal of Supercomputing, 79 (7):7286–7304, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04935-0.

Xue:2023:DBO

[2050] Jiankai Xue and Bo Shen. Dung beetle optimizer: a new meta-heuristic algorithm for global optimization. The Journal of Supercomputing, 79 (7):7305–7336, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04959-6.

Aghazadeh:2023:AMB

[2051] Nasser Aghazadeh, Paria Moradi, Giovanna Castellano, and Parisa Noras. An automatic MRI brain image segmentation technique using edgeregion-based level set. The Journal of Supercomputing, 79(7):7337– 7359, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04948-9.

Hammami:2023:NEL

[2052] Hamza Hammami, Sadok Ben Yahia, and Mohammad S. Obaidat. A novel efficient and lightweight authentication scheme for secure smart grid communication systems. The Journal of Supercomputing, 79(7):7360– 7376, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04944-z.

Ahmad:2023:HCA

[2053] Shahnawaz Ahmad, Shabana Mehfuz, and Javed Beg. Hybrid cryptographic approach to enhance the mode of key management system in cloud environment. The Journal of Supercomputing, 79(7):7377–7413, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04964-9.

Cheng:2023:PFP

[2054] Yuzhu Cheng and Qiuying Shi. PCMIgr: a fast packet classification method based on information gain ratio. The Journal of Supercomputing, 79(7):7414–7437, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04951-0.

Zheng:2023:AIA

[2055] Huijun Zheng, Jianlan Guo, Qin Zhou, Yong Peng, and Yuqiang Chen. Application of improved ant colony algorithm in load balancing of software-defined networks. The Journal of Supercomputing, 79(7): 7438–7460, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04957-8.

Savio:2023:ALA

[2056] Paolo Savio, Alberto Scionti, Giacomo Vitali, Paolo Viviani, Chiara Vercellino, Olivier Terzo, Huy-Nam Nguyen, Donato Magarielli, Ennio Spano, Michele Marconcini, and Francesco Poli. Accelerating legacy applications with spatial computing devices. The Journal of Supercomputing, 79(7):7461–7483, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04925-2.

Li:2023:CEA

[2057] Zengpeng Li, Huiqun Yu, and Guisheng Fan. Cost-effective approaches for deadline-constrained workflow scheduling in clouds. The Journal of Supercomputing, 79(7):7484– 7512, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04962-x.

Garea:2023:HCO

[2058] Alberto S. Garea, Dora B. Heras, Francisco Argüello, and Begüm Demir. A hybrid CUDA, OpenMP, and MPI parallel TCA-based domain adaptation for classification of very highresolution remote sensing images.

The Journal of Supercomputing, 79 (7):7513–7532, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04961-y.

Singh:2023:RTT

[2059] Tinku Singh, Vinarm Rajput, Satakshi, Umesh Prasad, and Manish Kumar. Real-time traffic light violations using distributed streaming. The Journal of Supercomputing, 79 (7):7533–7559, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04977-4.

Wan:2023:ECS

[2060] YunBo Wan, Lei He, Yong Zhang, Zhong Zhao, Jie Liu, and HaoYuan Zhang. An efficient communication strategy for massively parallel computation in CFD. The Journal of Supercomputing, 79(7):7560– 7583, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04940-3.

Chen:2023:DFM

[2061] Meirun Chen, Michel Habib, and Cheng-Kuan Lin. Diagnosability for a family of matching composition networks. The Journal of Supercomputing, 79(7):7584–7608, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04949-8.

Lupion:2023:ANN

[2062] Marcos Lupión, N. C. Cruz, Juan F.

Sanjuan, B. Paechter, and Pilar M. Ortigosa. Accelerating neural network architecture search using multi-GPU high-performance computing. The Journal of Supercomputing, 79 (7):7609–7625, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04960-z.

Nezamdoust:2023:OPC

[2063] Seyed Soroush Nezamdoust, Mohammad Ali Pourmina, and Farbod Razzazi. Optimal prediction of cloud spot instance price utilizing deep learning. The Journal of Supercomputing, 79 (7):7626–7647, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04970-x.

Badia:2023:SPF

[2064] Jose M. Badia, Adrian Amor-Martin, Jose A. Belloch, and Luis Emilio Garcia-Castillo. Strategies to parallelize a finite element mesh truncation technique on multi-core and many-core architectures. The Journal of Supercomputing, 79(7):7648– 7664, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04975-6.

Bang:2023:APZ

[2065] Jiwoo Bang, Chungyong Kim, Eun-Kyu Byun, Hanul Sung, Jaehwan Lee, and Hyeonsang Eom. Accelerating I/O performance of ZFS-based Lustre file system in HPC environment. The Journal of Supercomputing, 79 (7):7665–7691, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04966-7.

Shilpi:2023:OSN

[2066] Shilpi, Prateek Raj Gautam, Sunil Kumar, and Arvind Kumar. An optimized sensor node localization approach for wireless sensor networks using RSSI. The Journal of Supercomputing, 79 (7):7692–7716, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04971-w.

Jing:2023:RID

[2067] Chen Jing, Huang Jincheng, Xin Chen, and Liu Mingxin. Research on information dissemination model based on heat transfer in online social network. The Journal of Supercomputing, 79(7):7717–7735, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04968-5.

Varma:2023:ISP

[2068] Pothuri Surendra Varma and Veena Anand. Intelligent scanning period dilation based Wi–Fi fingerprinting for energy efficient indoor positioning in IoT applications. The Journal of Supercomputing, 79(7):7736– 7761, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04980-9.

Pinheiro:2023:PMF

[2069] Thiago Felipe da Silva Pinheiro, Paulo Pereira, Bruno Silva, and Paulo Maciel. A performance modeling framework for microservices-based cloud infrastructures. The Journal of Supercomputing, 79(7):7762–7803, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04967-6.

Rocher-Gonzalez:2023:CMH

[2070] Jose Rocher-Gonzalez, Jesus Escudero-Sahuquillo, Pedro J. Garcia, and Francisco J. Quiles. Congestion management in high-performance interconnection networks using adaptive routing notifications. The Journal of Supercomputing, 79(7):7804– 7834, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04926-1.

Tirandazi:2023:RSI

[2071] Peyman Tirandazi, Seyed Mojtaba Hosseini Bamakan, and Aref Toghroljerdi. A review of studies on Internet of Everything as an enabler of neuromarketing methods and techniques. The Journal of Supercomputing, 79 (7):7835–7876, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04988-1.

Magdy:2023:CIE

[2072] Maged Magdy, Fayed F. M. Ghaleb, Dawlat A. El A. Mohamed, and Wael Zakaria. CC-IFIM: an efficient approach for incremental frequent itemset mining based on closed candidates. The Journal of Supercomputing, 79 (7):7877–7899, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04976-5.

Jaiswal:2023:NAD

[2073] Vineet Jaiswal and Trailokya Nath Sasamal. Novel approach for the design of efficient full adder in MQCA. The Journal of Supercomputing, 79 (7):7900–7915, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04989-0.

Kayani:2023:MMT

[2074] Maemoona Kayani, Abdul Ghafoor, and M. Mohsin Riaz. Multi-modal text recognition and encryption in scanned document images. The Journal of Supercomputing, 79(7):7916– 7936, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04912-7.

Agrawal:2023:ASP

[2075] Supriya Agrawal and Prachi Natu. ABGS segmenter: pixel wise adaptive background subtraction and intensity ratio based shadow removal approach for moving object detection. The Journal of Supercomputing, 79 (7):7937–7969, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04972-9.

Li:2023:MEC

[2076] Yijun Li, Ziqiong Lin, Wenjie Zhang, Yifeng Zheng, and Jingmin Yang. Mobile edge computingenabled blockchain: contract-guided computation offloading. The Journal of Supercomputing, 79(7):7970– 7996, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04990-7.

Hu:2023:PIP

[2077] Junying Hu, Xiaofei Qian, Changchun Tan, and Xinbao Liu. Point and interval prediction of aircraft engine maintenance cost by bootstrapped SVR and improved RFE. The Journal of Supercomputing, 79(7):7997– 8025, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04986-3.

Zhou:2023:SBA

[2078] Mingqiang Zhou, Qianqian Zeng, and Ping Guo. SMC-BRB: an algorithm for the maximum clique problem over large and sparse graphs with the upper bound via s^{+}-index. The Journal of Supercomputing, 79 (7):8026–8047, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04982-7.

Oza:2023:DET

[2079] Parita Oza, Paawan Sharma, and Samir Patel. Deep ensemble transfer learning-based framework for mammographic image classification. The Journal of Supercomputing, 79(7): 8048–8069, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04992-5.

Kumar:2023:AAV

[2080] Pankaj Kumar and Hari Om. An anonymous and authenticated V2I communication with a simplified user revocation and re-registration strategy. The Journal of Supercomputing, 79(7):8070–8096, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04978-3.

Alagheband:2023:FUA

[2081] Mahdi R. Alagheband and Atefeh Mashatan. Formal unlinkability analysis of message authentication protocols based on authentic channels. The Journal of Supercomputing, 79 (7):8097–8123, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04993-4.

Alaejos:2023:MKP

[2082] Guillermo Alaejos, Adrián Castelló, Héctor Martínez, Pedro Alonso-Jordá, Francisco D. Igual, and Enrique S. Quintana-Ortí. Micro-kernels for portable and efficient matrix multiplication in deep learning. The Journal of Supercomputing, 79(7): 8124–8147, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05003-3.

Galiano:2023:CUD

[2083] Vicente Galiano, Héctor Migallón, Miguel Martínez-Rach, Otoniel López-Granado, and Manuel P. Malumbres. Correction to: On the use of deep learning and parallelism techniques to significantly reduce the HEVC intra-coding time. The Journal of Supercomputing, 79(7):8148–8149, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04918-1.

Khanna:2023:PCD

[2084] Gaurav Khanna, Sanjay K. Chaturvedi, and Mohamed Othman. Publisher correction to: On design and performance analysis of improved shuffle exchange gamma interconnection network layouts. The Journal of Supercomputing, 79(7):8150–8151, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04892-8.

Fakhfakh:2023:MOA

[2085] Fairouz Fakhfakh, Saoussen Cheikhrouhou, Bouthaina Dammak, Monia Hamdi, and Mouna Rekik. Multi-objective approach for scheduling time-aware business processes in cloud-fog environment. The Journal of Supercomputing, 79(8):8153–8177, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04690-2. See cor-

rection [2128].

Amini:2023:EEA

[2086] Vahid Amini, Mahmoud Momtazpour, and Morteza Saheb Zamani. An energy-efficient and accuracy-aware edge computing framework for heart arrhythmia detection: a joint model selection and task offloading approach. The Journal of Supercomputing, 79 (8):8178–8204, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04987-2.

Asorey:2023:CHE

[2087] Hernán Asorey and Rafael Mayo-García. Calculation of the highenergy neutron flux for anticipating errors and recovery techniques in exascale supercomputer centres. The Journal of Supercomputing, 79 (8):8205–8235, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04981-8.

Wang:2023:ANM

[2088] Qihang Wang, Decheng Zuo, Zhan Zhang, Siyuan Chen, and Tianming Liu. An adaptive nonmigrating load-balanced distributed stream window join system. The Journal of Supercomputing, 79(8): 8236–8264, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04991-6.

Patidar:2023:UAE

[2089] Mukesh Patidar, Upendra Singh, Surendra Kumar Shukla, Giriraj Kumar Prajapati, and Namit Gupta. An ultra-area-efficient ALU design in QCA technology using synchronized clock zone scheme. The Journal of Supercomputing, 79(8):8265– 8294, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05012-2.

Fu:2023:MAB

[2090] Yiyang Fu, Xiaojun Xie, and Tao Zhang. MRAN: a attention-based approach for social recommendation. The Journal of Supercomputing, 79 (8):8295–8321, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04985-4.

Liao:2023:ABB

[2091] Xiaoxin Liao, Jingyi Yuan, Zemin Cai, and Jian huang Lai. An attentionbased bidirectional GRU network for temporal action proposals generation. The Journal of Supercomputing, 79 (8):8322–8339, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04973-8.

Zou:2023:MRP

[2092] Awei Zou, Lei Wang, Weimin Li, Jingcao Cai, Hai Wang, and Tielong Tan. Mobile robot path planning using improved mayfly optimization algorithm and dynamic window approach.

The Journal of Supercomputing, 79 (8):8340–8367, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04998-z.

Kumar:2023:SSD

[2093] Rohit Kumar and Neha Agrawal. A survey on software-defined vehicular networks (SDVNs): a security perspective. The Journal of Supercomputing, 79(8):8368–8400, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05008-y.

Tian:2023:CCC

[2094] Junfeng Tian, Haoyi Jia, and Wenqing Bai. CCECGP: causal consistency model of edge-cloud collaborative based on grouping protocol. The Journal of Supercomputing, 79 (8):8401–8424, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04997-0.

UelintondaSilva:2023:POO

[2095] Hígor Uélinton da Silva, Natiele Lucca, Claudio Schepke, Dalmo Paim de Oliveira, and César Flaubiano da Cruz Cristaldo. Parallel OpenMP and OpenACC porous media simulation. The Journal of Supercomputing, 79 (8):8425–8446, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05004-2.

Fuentes-Alventosa:2023:GHO

[2096] Antonio Fuentes-Alventosa, Juan Gómez-Luna, and R. Medina-Carnicer. GVLE: a highly optimized GPU-based implementation of variable-length encoding. The Journal of Supercomputing, 79(8):8447–8474, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-022-04994-
3.

Garcia-Nava:2023:FTT

[2097] J.-Luis García-Nava, Juan J. Flores, Victor M. Tellez, and Felix Calderon. Fast training of a transformer for global multi-horizon time series forecasting on tensor processing units. The Journal of Supercomputing, 79 (8):8475–8498, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05009-x.

Yu:2023:RSC

[2098] Cui Yu, Fei Gao, Boyong Gao, and Yawen Chen. h-restricted Hstructure connectivity and h-restricted H-substructure connectivity of hypercube. The Journal of Supercomputing, 79(8):8499–8512, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04983-6.

Paramasivam:2023:CEC

[2099] Suguna Paramasivam and R. Leela Velusamy. Cor-ENTC: correlation with ensembled approach for network traffic classification using SDN technology for future networks. The Journal of Supercomputing, 79(8): 8513–8537, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04969-4.

Zhang:2023:SEM

[2100] Wenbo Zhang, Xiaotong Huo, and Zhenshan Bao. A secure and efficient multi-domain data sharing model on consortium chain. The Journal of Supercomputing, 79(8):8538– 8582, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05010-4.

Abdollahizad:2023:SEA

[2101] Solmaz Abdollahizad, Mohammad Ali Balafar, Bakhtiar Feizizadeh, Amin Babazadeh Sangar, and Karim Samadzamini. Stacking ensemble approach in data mining methods for landslide prediction. The Journal of Supercomputing, 79(8):8583–8610, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05006-0.

Yi:2023:DME

[2102] Yufan Yi, Yan Tian, Cong He, Yajing Fan, Xinli Hu, and Yiping Xu. DBT: multimodal emotion recognition based on dual-branch transformer. The Journal of Supercomputing, 79 (8):8611–8633, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05001-5.

Yang:2023:TMB

[2103] Xin Yang, Abla Smahi, Hui Li, Ping Lu, Huayu Zhang, and Shuo-Yen Robert Li. Triple methods-based empirical assessment of the effectiveness of adaptive cyber defenses in the cloud. The Journal of Supercomputing, 79(8):8634–8667, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04984-5.

Guo:2023:NOA

[2104] Ming Guo and Jia Lu. ND-S: an oversampling algorithm based on natural neighbor and density peaks clustering. The Journal of Supercomputing, 79(8):8668–8698, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04965-8.

Perez-Wohlfeil:2023:IAA

[2105] Esteban Perez-Wohlfeil, Oswaldo Trelles, and Nicolás Guil. Irregular alignment of arbitrarily long DNA sequences on GPU. The Journal of Supercomputing, 79(8):8699– 8728, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05007-z.

Liang:2023:PEM

[2106] Chi-Hsiu Liang, Chun-Ho Cheng, Hong-Lin Wu, Chao-Chin Li, Po-Lin Huang, and Chi-Chuan Hwang. Performance evaluation of multi-exaflops machines using Equality network topology. The Journal of Supercomputing, 79(8):8729–8753, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05005-1.

Saha:2023:ECC

[2107] Munmun Saha, Sanjaya Kumar Panda, Suvasini Panigrahi, and David Taniar. An efficient composite cloud service model using multi-criteria decisionmaking techniques. The Journal of Supercomputing, 79(8):8754– 8788, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05013-1.

Zhang:2023:IHA

[2108] Jianhui Zhang, Xuezhen Cheng, Meng Zhao, and Jiming Li. ISSWOA: hybrid algorithm for function optimization and engineering problems. The Journal of Supercomputing, 79 (8):8789–8842, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04996-1.

Brahim:2023:NFI

[2109] A. Hadj Brahim, A. Ali Pacha, and N. Hadj Said. A new fast image compression-encryption scheme based on compressive sensing and parallel blocks. The Journal of Supercomputing, 79(8):8843–8889, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-022-04999-
y.

DiDomenico:2023:NPB

[2110] Daniel Di Domenico, João V. F.

Lima, and Gerson G. H. Cavalheiro. NAS Parallel Benchmarks with Python: a performance and programming effort analysis focusing on GPUs. The Journal of Supercomputing, 79 (8):8890–8911, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04932-3.

Vani:2023:PAN

[2111] K. Vani and G. Sudha Sadasivam.

Privacy-aware network embeddingbased ensemble for social recommendation. The Journal of Supercomputing, 79(8):8912–8939, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04958-7.

Padulano:2023:LOS

[2112] Vincenzo Eduardo Padulano, Pablo Oliver Cortés, Pedro Alonso-Jordá, Enric Tejedor Saavedra, Sebastián Risco, and Germán Moltó. Leveraging an open source serverless framework for high energy physics computing. The Journal of Supercomputing, 79 (8):8940–8965, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05016-y.

Pham:2023:IIY

[2113] Thi-Ngot Pham, Viet-Hoan Nguyen, and Jun-Ho Huh. Integration of improved YOLOv5 for face mask detector and auto-labeling to generate dataset for fighting against COVID-19. The Journal of Supercomputing, 79(8):8966–8992, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04979-2.

Yu:2023:WSR

[2114] Ting Yu, Dongjin Yu, Dongjing Wang, and Xueyou Hu. Web service recommendation for mashup creation based on graph network. The Journal of Supercomputing, 79(8):8993– 9020, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05011-3.

Luo:2023:MTL

[2115] Jie Luo, Shubo Liu, Zhaohui Cai, Chang Xiong, and Guoqing Tu. A multi-task learning model for nonintrusive load monitoring based on discrete wavelet transform. The Journal of Supercomputing, 79(8):9021– 9046, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05000-6.

Idrees:2023:ECT

[2116] Ali Kadhum Idrees and Marwa Saieed Khlief. Efficient compression technique for reducing transmitted EEG data without loss in IoMT networks based on fog computing. The Journal of Supercomputing, 79(8):9047– 9072, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05027-9.

Liao:2023:TKE

[2117] Shengbin Liao, Zongkai Yang, Qingzhou Liao, and Zhangxiong Zheng. TopicLPRank: a keyphrase extraction method based on improved TopicRank. The Journal of Supercomputing, 79 (8):9073–9092, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05022-0.

Sadiq:2023:SEC

[2118] Muhammad Umair Sadiq and Muhammad Murtaza Yousaf. Space-efficient computation of parallel approximate string matching. The Journal of Supercomputing, 79(8):9093– 9126, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05038-6.

Barve:2023:DCO

[2119] Yashoda Barve and Jatinderkumar R. Saini. Detecting and classifying online health misinformation with ‘Content Similarity Measure (CSM)’ algorithm: an automated fact-checking-based approach. The Journal of Supercomputing, 79 (8):9127–9156, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05032-y.

Yang:2023:SSF

[2120] Lulu Yang, Xiaohui Hua, and Yuxing Yang. On structure and substructure fault tolerance of star networks. The Journal of Supercomputing, 79 (8):9157–9179, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-05036-8.

Li:2023:NID

[2121] Jieling Li, Hao Zhang, Zhihuang Liu, and Yanhua Liu. Network intrusion detection via tri-broad learning system based on spatial-temporal granularity. The Journal of Supercomputing, 79(8):9180–9205, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-05025-x.

Garcia:2023:MBD

[2122] Adriano Marques Garcia, Dalvan Griebler, Claudio Schepke, and Luiz Gustavo Fernandes. Micro-batch and data frequency for stream processing on multi-cores. The Journal of Supercomputing, 79(8):9206– 9244, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-05024-y.

Thomadakis:2023:TRS

[2123] Polykarpos Thomadakis and Nikos Chrisochoides. Toward runtime support for unstructured and dynamic exascale-era applications. The Journal of Supercomputing, 79(8):9245– 9272, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05023-z.

DeRango:2023:TLB

[2124] Alessio De Rango, Andrea Giordano, Giuseppe Mendicino, Rocco Rongo, and William Spataro. Tailoring load balancing of cellular automata parallel execution to the case of a two-dimensional partitioned domain. The Journal of Supercomputing, 79 (8):9273–9287, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05043-3.

Reddy:2023:DBT

[2125] K. Rasool Reddy and Ravindra Dhuli.

Detection of brain tumors from MR images using fuzzy thresholding and texture feature descriptor. The Journal of Supercomputing, 79(8):9288– 9319, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05033-x.

Moin:2023:ERF

[2126] Anam Moin, Farhan Aadil, Zeeshan Ali, and Dongwann Kang. Emotion recognition framework using multiple modalities for an effective humancomputer interaction. The Journal of Supercomputing, 79(8):9320– 9349, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05026-w.

Chehreghani:2023:IEE

[2127] Mostafa Haghir Chehreghani and Zahra Yaghoobi. Improving empirical efficiency of CUR decomposition. The Journal of Supercomputing, 79 (8):9350–9366, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-05039-5.

Fakhfakh:2023:CMO

[2128] Fairouz Fakhfakh, Saoussen Cheikhrouhou, Bouthaina Dammak, Monia Hamdi, and Mouna Rekik. Correction to: Multi-objective approach for scheduling time-aware business processes in cloud-fog environment. The Journal of Supercomputing, 79(8): 9367, May 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04905-6.  See

[2085].

Khabou:2023:ISI

[2129] Nesrine Khabou, Ismael Bouassida Rodriguez, and Mohamed Jmaiel. Introduction to the special issue on new trends in autonomous systems engineering. The Journal of Supercomputing, 79(8):9368–9370, May 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-023-05157-
                                        8.

Lloria:2023:HCG

[2130] Diego Lloria, Pablo M. Aviles, Jose A. Belloch, Sandra Roger, Carmen Botella-Mascarell, and Almudena Lindoso. Hybrid CPU–GPU implementation of the transformed spatial domain channel estimation algorithm for mmWave MIMO systems. The Journal of Supercomputing, 79 (9):9371–9382, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05018-w.

Mukherjee:2023:ULG

[2131] Aradhita Mukherjee, Rituparna Chaki, and Nabendu Chaki. An unsupervised learning-guided multi-node failure-recovery model for distributed graph processing systems. The Journal of Supercomputing, 79(9):9383– 9408, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05028-8.

deCastro:2023:EEP

[2132] Manuel de Castro, Inmaculada Santamaria-Valenzuela, Yuri Torres, Arturo Gonzalez-Escribano, and Diego R. Llanos. EPSILOD: efficient parallel skeleton for generic iterative stencil computations in distributed GPUs. The Journal of Supercomputing, 79 (9):9409–9442, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05040-y.

Abadeh:2023:DML

[2133] Maryam Nooraei Abadeh and Mansooreh Mirzaie. A differential machine learning approach for trust prediction in signed social networks.

The Journal of Supercomputing, 79 (9):9443–9466, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05044-2.

Seo:2023:DFI

[2134] Seunghee Seo, Byoungjin Seok, and Changhoon Lee. Digital forensic investigation framework for the metaverse. The Journal of Supercomputing, 79(9):9467–9485, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05045-1.

Belgacem:2023:MLM

[2135] Ali Belgacem, Saïd Mahmoudi, and Mohamed Amine Ferrag. A machine learning model for improving virtual machine migration in cloud computing. The Journal of Supercomputing, 79(9):9486–9508, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-05031-z.

Shaji:2023:HPF

[2136] B. Shaji, R. Lal Raja Singh, and K. L. Nisha. High-performance fuzzy optimized deep convolutional neural network model for big data classification based on the Social Internet of Things. The Journal of Supercomputing, 79 (9):9509–9537, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04974-7.

Garcia:2023:IDD

[2137] Jerónimo S. García, Savíns Puertas-Martín, Juana L. Redondo, Juan José Moreno, and Pilar M. Ortigosa. Improving drug discovery through parallelism. The Journal of Supercomputing, 79(9):9538–9557, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05014-0.

Wei:2023:LSC

[2138] Wei Wei, Xiaowan Li, Beibei Zhang, Linfeng Li, Robertas Damaševičius, and Rafal Scherer. LSTM-SN: complex text classifying with LSTM fusion social network. The Journal of Supercomputing, 79(9):9558– 9583, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05034-w.

He:2023:DDM

[2139] Qinlu He, Fan Zhang, Genqing Bian, Weiqi Zhang, Zhen Li, and Chen Chen. Dynamic decision-making strategy of replica number based on data hot. The Journal of Supercomputing, 79 (9):9584–9603, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05029-7.

Wei:2023:LCG

[2140] Junlin Wei, Jinrong Jiang, Hailong Liu, Feng Zhang, Pengfei Lin, Pengfei Wang, Yongqiang Yu, Xuebin Chi, Lian Zhao, Mengrong Ding, Yiwen Li, Zipeng Yu, Weipeng Zheng, and Yuzhu Wang. LICOM3-CUDA: a GPU version of LASG/IAP climate system ocean model version 3 based on CUDA. The Journal of Supercomputing, 79 (9):9604–9634, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05020-2.

Carratala-Saez:2023:ULF

[2141] Rocío Carratalá-Sáez, Yuri Torres, José Sierra-Pallares, Sergio López-Huguet, and Diego R. Llanos. UVaF-TLE: Lagrangian finite time Lyapunov exponent extraction for fluid dynamic applications. The Journal of Supercomputing, 79(9):9635– 9665, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05017-x.

Jiang:2023:ALS

[2142] Baoxing Jiang, Guangtao Xu, and Peiyu Liu. Aspect-level sentiment classification via location enhanced aspectmerged graph convolutional networks. The Journal of Supercomputing, 79 (9):9666–9691, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05002-4.

Rashid:2023:ARP

[2143] Tarik A. Rashid, Bryar A. Hassan, Abeer Alsadoon, Shko Qader, S. Vimal, Amit Chhabra, and Zaher Mundher Yaseen. Awareness requirement and performance management for adaptive systems: a survey. The Journal of Supercomputing, 79 (9):9692–9714, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05021-1.

Kaveh:2023:TNT

[2144] Mehrdad Kaveh, Mohammad Saadi Mesgari, Diego Martín, and Masoud Kaveh. TDMBBO: a novel three-dimensional migration model of biogeography-based optimization (case study: facility planning and benchmark problems). The Journal of Supercomputing, 79(9):9715– 9770, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05047-z.

Modi:2023:HDL

[2145] Tejas M. Modi and Pravati Swain.

Hybrid deep learning models and link probability based routing in software defined-DCN. The Journal of Supercomputing, 79(9):9771– 9794, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04995-2.

Yang:2023:RBH

[2146] Peng Yang, Wenjun Li, Guangzhen Zhao, and Xianyu Zha. Row-based hierarchical graph network for multihop question answering over textual and tabular data. The Journal of Supercomputing, 79(9):9795– 9818, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-05035-9.

Dolz:2023:PET

[2147] Manuel F. Dolz, Sergio Barrachina, Héctor Martínez, Adrián Castelló, Antonio Maciá, Germán Fabregat, and Andrés E. Tomás. Performance-energy trade-offs of deep learning convolution algorithms on ARM processors. The Journal of Supercomputing, 79 (9):9819–9836, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05050-4.

Itoo:2023:RRK

[2148] Samiulla Itoo, Musheer Ahmad, Vinod Kumar, and Ahmed Alkhayyat. RK-MIS: robust key management protocol for industrial sensor network system. The Journal of Supercomputing, 79(9):9837–9865, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-05041-x.

deCastro:2023:IME

[2149] Manuel de Castro, Roberto R. Osorio, David L. Vilariño, Arturo Gonzalez-Escribano, and Diego R. Llanos. Implementation of a motion estimation algorithm for Intel FPGAs using OpenCL. The Journal of Supercomputing, 79 (9):9866–9888, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05051-3.

Cai:2023:CSN

[2150] Gang Cai and Lingyan Wu. Cycle sampling neural network algorithms and applications. The Journal of Supercomputing, 79(9):9889– 9914, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05019-9.

Garcia-Molla:2023:PBT

[2151] Victor M. Garcia-Molla and Pedro Alonso-Jordá. Parallel border tracking in binary images for multicore computers. The Journal of Supercomputing, 79(9):9915–9931, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05052-2.

Zhou:2023:EGD

[2152] Peng Zhou, Longxin Lin, Tengjiao He, and Zhen Zhang. GHDC: a dual-centric data center network architecture by using multi-port servers with greater incremental scalability. The Journal of Supercomputing, 79 (9):9932–9963, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05046-0.

Liang:2023:EFD

[2153] Jiarong Liang, Yang Guo, Yuhao Xie, and Xinyu Liang. An evolutionary fault diagnosis algorithm for interconnection networks under the PMC model. The Journal of Supercomputing, 79 (9):9964–9984, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05054-0.

Ye:2023:SCL

[2154] Wei Ye, Minggang Dong, Yan Wang, Guojun Gan, and Deao Liu. Similar classes latent distribution modellingbased oversampling method for imbalanced image classification. The Journal of Supercomputing, 79(9): 9985–10019, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-05037-7.

Uzer:2023:NFS

[2155] Mustafa Serter Uzer and Onur Inan. A novel feature selection using binary hybrid improved whale optimization algorithm. The Journal of Supercomputing, 79(9):10020– 10045, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05067-9.

Wang:2023:SGS

[2156] Jia Wang and Yanmin Luo. SA-GCN: structure-aware graph convolutional networks for crowd pose estimation. The Journal of Supercomputing, 79 (9):10046–10062, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05055-z.

Radaideh:2023:PCB

[2157] Moh’d A. Radaideh, Nadil Iyad Mohammad, and Maya Mohammad Mukbil. A proposed cloud-based platform for facilitating donation services in support to needy-students. The Journal of Supercomputing, 79 (9):10063–10092, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05062-0.

Paul:2023:RSS

[2158] Aditi Paul and Somnath Sinha. Received signal strength-based location verification technique in wireless sensor network using spline curve. The Journal of Supercomputing, 79(9): 10093–10116, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05069-7.

Chen:2023:LLO

[2159] Naiyuan Chen, Yan Li, Zhuomin Yang, Zhensong Lu, Sai Wang, and Junang Wang. LODNU: lightweight object detection network in UAV vision. The Journal of Supercomputing, 79 (9):10117–10138, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05065-x.

Zhao:2023:PRT

[2160] Hao Zhao, Jingwei Geng, and Shunfu Jin. Performance research on a task offloading strategy in a two-tier edge structure-based MEC system. The Journal of Supercomputing, 79(9): 10139–10177, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05059-9.

Quislant:2023:TSA

[2161] Ricardo Quislant, Ivan Fernandez, Eladio Gutierrez, and Oscar Plata. Time series analysis acceleration with advanced vectorization extensions. The Journal of Supercomputing, 79(9): 10178–10207, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05060-2.

Song:2023:DTS

[2162] Yao Song, Liang Wang, Limin Xiao, Runnan Shen, Jinquan Wang, and Chenhao Zhang. Dynamic twoside matching of tasks and resources in wide-area distributed computing environments. The Journal of Supercomputing, 79(9):10208– 10231, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05056-y.

Roshani:2023:DMI

[2163] Reza Roshani, Homayon Motameni, and Hosein Mohamadi. A decentralized method for initial populations of genetic algorithms. The Journal of Supercomputing, 79(9): 10232–10251, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05066-w.

Navarro-Torres:2023:BBA

[2164] Agustín Navarro-Torres, Jesús Alastruey-Benedé, Pablo Ibáñez, and Víctor Viñals-Yúfera. BALANCER: bandwidth allocation and cache partitioning for multicore processors. The Journal of Supercomputing, 79(9): 10252–10276, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05070-0.

Morales-Garcia:2023:ELP

[2165] Juan Morales-García, Andrés Bueno-Crespo, Raquel Martínez-España, Juan-Luis Posadas, Pietro Manzoni, and José M. Cecilia. Evaluation of low-power devices for smart greenhouse development. The Journal of Supercomputing, 79(9):10277– 10299, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05076-8.

Li:2023:CES

[2166] Hongjian Li, Jie Shen, Lei Zheng, Yuzheng Cui, and Zhi Mao. Costefficient scheduling algorithms based on beetle antennae search for containerized applications in Kubernetes clouds. The Journal of Supercomputing, 79 (9):10300–10334, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05077-7.

Rico:2023:KRA

[2167] Noelia Rico, Pedro Alonso, and Irene Díaz. Kemeny ranking aggregation meets the GPU. The Journal of Supercomputing, 79(9):10335– 10352, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05058-w.

Shalini:2023:EDM

[2168] P. V. Shalini, V. Radha, and Sriram G. Sanjeevi. Early detection and mitigation of TCP SYN flood attacks in SDN using chi-square test. The Journal of Supercomputing, 79 (9):10353–10385, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05057-x.

Chauhan:2023:ECD

[2169] Pinkey Chauhan and Mithilesh Atulkar. An efficient centralized DDoS attack detection approach for Software Defined Internet of Things. The Journal of Supercomputing, 79(9): 10386–10422, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05072-y.

Quiroz-Fabian:2023:PEP

[2170] J. L. Quiroz-Fabián, G. Román-Alonso, M. A. Castro-García, and M. Aguilar-Cornejo. PAARes: an efficient process allocation based on the available resources of cluster nodes. The Journal of Supercomputing, 79(9): 10423–10441, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05085-7.

Rahmani:2023:SDO

[2171] Peyman Rahmani, Seyed Mostafa Fakhrahmad, and Mohammad Taheri.

Secure data outsourcing based on seedresidual shares and order-shuffling encryption. The Journal of Supercomputing, 79(9):10442–10480, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05048-
y.

Saravanan:2023:MMB

[2172] R. Saravanan, K. Suresh, and S. S.

Arumugam. A modified k-means-based cluster head selection and Philippine eagle optimization-based secure routing for MANET. The Journal of Supercomputing, 79(9):10481– 10504, June 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05053-1.

Sadeghi:2023:CICa

[2173] Maryam Sadeghi, Mohammad Naderi Dehkordi, Behrang Barekatain, and Naser Khani. Correction to: Improve customer churn prediction through the proposed PCA–PSO-K means algorithm in the communication industry. The Journal of Supercomputing, 79(9):10505, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05015-z.  See

[2028].

Vinoth:2023:CIM

[2174] D. Vinoth and P. Prabhavathy. Correction to: An intelligent machine learning-based sarcasm detection and classification model on social networks. The Journal of Supercomputing, 79(9):10506, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05071-z. See cor-

rigendum [1436].

Hassan:2023:CFG

[2175] Md Raful Hassan, Walaa N. Ismail, Ahmad Chowdhury, Sharara Hossain, Shamsul Huda, and Mohammad Mehedi Hassan. Correction to: A framework of genetic algorithm-based CNN on multi-access edge computing for automated detection of COVID-19. The Journal of Supercomputing, 79 (9):10507–10508, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05063-z.  See

[1423].

Pu:2023:RNA

[2176] Zhenzi Pu, Qijie Jiang, Hongzhu Yue, and Maria Tsaptsinos. Retraction note: Agent-based supply chain allocation model and its application in smart manufacturing enterprises. The Journal of Supercomputing, 79(9):10509, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05230-2. See [194].

Moayed:2023:EPM

[2177] Hojjat Moayed, Eghbal G. Mansoori, and Mohammad R. Moosavi. An efficient pruning method for subgraph matching in large-scale graphs. The Journal of Supercomputing, 79(10):

10511–10532, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05061-1.

Alarcon:2023:CGB

[2178] Sonia Lopez Alarcon and Federico Rueda. Compilation of Gaussian boson samplers for quantum computing. The Journal of Supercomputing, 79 (10):10533–10554, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05075-9.

Sabelfeld:2023:PIR

[2179] Karl K. Sabelfeld, Sergey Kireev, and Anastasiya Kireeva. Parallel implementations of randomized vector algorithm for solving large systems of linear equations. The Journal of Supercomputing, 79(10):10555– 10569, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05079-5.

Fang:2023:PCS

[2180] Juan Fang, Yixiang Xu, Han Kong, and Min Cai. A prefetch control strategy based on improved hill-climbing method in asymmetric multi-core architecture. The Journal of Supercomputing, 79(10):10570– 10588, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05078-6.

Dolz:2023:EPW

[2181] Manuel F. Dolz, Héctor Martínez, Adrián Castelló, Pedro Alonso-Jordá, and Enrique S. Quintana-Ortí. Efficient and portable Winograd convolutions for multi-core processors. The Journal of Supercomputing, 79(10): 10589–10610, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05088-4.

Abdelkhalek:2023:ACI

[2182] Ahmed Abdelkhalek and Maggie Mashaly. Addressing the class imbalance problem in network intrusion detection systems using data resampling and deep learning. The Journal of Supercomputing, 79(10): 10611–10644, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05073-x.

Seghir:2023:IDF

[2183] Fateh Seghir and Ghizlane Khababa. An improved discrete flower pollination algorithm for fuzzy QoSaware IoT services composition based on skyline operator. The Journal of Supercomputing, 79(10):10645– 10676, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05074-w.

Li:2023:RAD

[2184] John Li, Maria Pantoja, and Gerardo Fernández-Escribano. Reliable adaptive distributed hyperparameter optimization (RadHPO) for deep learning training and uncertainty estimation. The Journal of Supercomputing, 79 (10):10677–10690, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05081-x.

Karozis:2023:MNH

[2185] Stelios Karozis, Xenia Ziouvelou, and Vangelis Karkaletsis. Mapping the national HPC ecosystem and training needs: The Greek paradigm. The Journal of Supercomputing, 79(10): 10691–10705, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05080-y.

Dang:2023:IYR

[2186] Thi Phuc Dang, Ngoc Trinh Tran, Van Hau To, and Minh Khoa Tran Thi. Improved YOLOv5 for realtime traffic signs recognition in bad weather conditions. The Journal of Supercomputing, 79(10):10706– 10724, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05097-3.

Dai:2023:CPM

[2187] Fei Dai, Yawen Chen, Zhiyi Huang, Haibo Zhang, Hao Zhang, and Chengpeng Xia. Comparing the performance of multi-layer perceptron training on electrical and optical network-on-chips. The Journal of Supercomputing, 79 (10):10725–10746, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04945-y.

Alrashidi:2023:SIB

[2188] Malek Alrashidi. Synergistic integration between internet of things and augmented reality technologies for deaf persons in e-learning platform. The Journal of Supercomputing, 79(10): 10747–10773, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04952-z.

Jeannot:2023:IML

[2189] Emmanuel Jeannot and Richard Sartori. An introspection monitoring library to improve MPI communication time. The Journal of Supercomputing, 79(10):10774– 10795, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05084-8.

AbdulKader:2023:PPD

[2190] M. Mohideen AbdulKader and S. Ganesh Kumar. A privacy-preserving data transfer in a blockchain-based commercial real estate platform using random address generation mechanism. The Journal of Supercomputing, 79 (10):10796–10822, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04906-5.

Zhang:2023:AII

[2191] Jinzhong Zhang, Gang Zhang, Min Kong, and Tan Zhang. Adaptive infinite impulse response system identification using an enhanced golden jackal optimization. The Journal of Supercomputing, 79(10): 10823–10848, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05086-6.

Lee:2023:ACC

[2192] Hee-Jin Lee and Bum Yong Park.

Adaptive CVCF controller design for an energy storage system considering an operation mode change in a standalone microgrid. The Journal of Supercomputing, 79(10):10849– 10863, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04953-y.

Cui:2023:ITL

[2193] Jingkai Cui, Tianyu Liu, Mingchao Zhu, and Zhenbang Xu. Improved team learning-based grey wolf optimizer for optimization tasks and engineering problems. The Journal of Supercomputing, 79(10):10864– 10914, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04930-5.

AlAghbari:2023:FFT

[2194] Zaher Al Aghbari, P. V. Pravija Raj, and Ahmed M. Khedr. FtCFt: a fault-tolerant coverage preserving strategy for face topology-based wireless sensor networks. The Journal of Supercomputing, 79(10):10915– 10940, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05092-8.

Aryai:2023:SSI

[2195] Pouya Aryai, Ahmad Khademzadeh, Somayyeh Jafarali Jassbi, and Mehdi Hosseinzadeh. SIMOF: swarm intelligence multi-objective fuzzy thermalaware routing protocol for WBANs. The Journal of Supercomputing, 79 (10):10941–10976, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05102-9.

Malik:2023:EES

[2196] Annu Malik and Rashmi Kushwah. Energy-efficient scheduling in IoT using Wi-Fi and ZigBee cross-technology. The Journal of Supercomputing, 79 (10):10977–11006, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05093-7.

Kumar:2023:PCF

[2197] Jatinder Kumar, Rishabh Gupta, Deepika Saxena, and Ashutosh Kumar Singh. Power consumption forecast model using ensemble learning for smart grid. The Journal of Supercomputing, 79(10):11007– 11028, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05096-4.

Qorich:2023:TSC

[2198] Mohammed Qorich and Rajae El Ouazzani. Text sentiment classification of Amazon reviews using word embeddings and convolutional neural networks. The Journal of Supercomputing, 79(10):11029– 11054, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05094-6.

Shen:2023:CBM

[2199] Jingcheng Shen, Linbo Long, Xin Deng, Masao Okita, and Fumihiko Ino. A compression-based memoryefficient optimization for out-of-core GPU stencil computation. The Journal of Supercomputing, 79(10):11055– 11077, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05103-8.

Fernandez-Gomez:2023:NAS

[2200] Antonio M. Fernández-Gómez, David Gutiérrez-Avilés, Alicia Troncoso, and Francisco Martínez-Álvarez. A new Apache Spark-based framework for big data streaming forecasting in IoT networks. The Journal of Supercomputing, 79(10):11078– 11100, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05100-x.

Singh:2023:TFB

[2201] Indu Singh and Rajni Jindal. Trust factor-based analysis of user behavior using sequential pattern mining for detecting intrusive transactions in databases. The Journal of Supercomputing, 79(10):11101– 11133, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05090-w.

Tewari:2023:EER

[2202] Peeyush Tewari and Sandesh Tripathi. An energy efficient routing scheme in internet of things enabled WSN: neuro-fuzzy approach. The Journal of Supercomputing, 79(10): 11134–11158, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05091-9.

Kwak:2023:EPM

[2203] Soobin Kwak, Seokjun Ham, Youngjin Hwang, and Junseok Kim. Estimation and prediction of the multiply exponentially decaying daily case fatality rate of COVID-19. The Journal of Supercomputing, 79(10): 11159–11169, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05119-0.

Zhao:2023:ISO

[2204] Pengjun Zhao and Sanyang Liu. An improved symbiotic organisms search algorithm with good point set and memory mechanism. The Journal of Supercomputing, 79(10):11170– 11197, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05083-9.

Dong:2023:AAL

[2205] Yongchuan Dong, Qiaosha Zou, and Chuanjin Richard Shi. Augmenting aspect-level sentiment classification with distance-related local context input. The Journal of Supercomputing, 79(10):11198– 11217, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05108-3.

Shukla:2023:MMO

[2206] Prashant Shukla and Sudhakar Pandey.

MAA: multi-objective artificial algae algorithm for workflow scheduling in heterogeneous fog-cloud environment. The Journal of Supercomputing, 79 (10):11218–11260, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05110-9.

Yadav:2023:SKB

[2207] Awaneesh Kumar Yadav, An Braeken, and Manoj Misra. Symmetric key-based authentication and key agreement scheme resistant against semi-trusted third party for fog and dew computing. The Journal of Supercomputing, 79(10):11261– 11299, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05064-y.

Guo:2023:EHF

[2208] Ruyan Guo, Yan Wang, Jianxi Fan, and Weibei Fan. Embedding hierarchical folded cubes into linear arrays and complete binary trees with minimum wirelength. The Journal of Supercomputing, 79(10):11300– 11327, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05095-5.

Usharani:2023:ISD

[2209] S. Usharani and K. Dhanalakshmi. An image storage duplication detection method using recurrent learning for smart application services. The Journal of Supercomputing, 79(10): 11328–11354, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05042-4.

Ahmed:2023:RAU

[2210] Usman Ahmed and Jerry Chun-Wei Lin. Robust adversarial uncertainty quantification for deep learning fine-tuning. The Journal of Supercomputing, 79(10):11355– 11386, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05087-5.

Yoon:2023:STF

[2211] Daegun Yoon, Minjoong Jeong, and Sangyoon Oh. SAGE: toward on-the-fly gradient compression ratio scaling. The Journal of Supercomputing, 79 (10):11387–11409, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05120-7.

Chen:2023:AMM

[2212] Ruibo Chen, Yanjun Pu, Bowen Shi, and Wenjun Wu. An automatic model management system and its implementation for AIOps on microservice platforms. The Journal of Supercomputing, 79(10):11410– 11426, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05123-4.

Pan:2023:SHE

[2213] Jingshan Pan, Lei Xiao, Min Tian, Tao Liu, and Yinglong Wang. sw-ParaFEM: a highly efficient parallel finite element solver on Sunway many-core architecture. The Journal of Supercomputing, 79(10):11427– 11451, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05114-5.

Wankhade:2023:MBB

[2214] Mayur Wankhade, Chandra Sekhara Rao Annavarapu, and Ajith Abraham. MAPA BiLSTM–BERT: multi-aspects position aware attention for aspect level sentiment analysis. The Journal of Supercomputing, 79(10):11452– 11477, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05112-7.

Kim:2023:NSA

[2215] Eunsook Kim, Kyungwoon Lee, and Chuck Yoo. Network SLO-aware container scheduling in Kubernetes. The Journal of Supercomputing, 79(10): 11478–11494, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05122-5.

Li:2023:GGS

[2216] David Chunhu Li, Pin-Hao Chen, and Li-Der Chou. GAP4NSH: a genetic service function chaining with network service header for P4-based software-defined networks. The Journal of Supercomputing, 79(10):11495– 11529, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05121-6.

Alharbi:2023:NUE

[2217] Mohammed Alharbi, Gerard Edwards, and Richard Stocker. Novel ultraenergy-efficient reversible designs of sequential logic quantum-dot cellular automata flip-flop circuits. The Journal of Supercomputing, 79(10): 11530–11557, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05134-1.

Guo:2023:TLC

[2218] Peng Guo, Qi Liu, Shui Yu, Jianyu Xiong, Xiang Tan, and Chao Guo. A transformer with layer-cross decoding for remaining useful life prediction. The Journal of Supercomputing, 79 (10):11558–11584, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05126-1.

Sethi:2023:SCT

[2219] Biswajeet Sethi, Sourav Kanti Addya, Jay Bhutada, and Soumya K. Ghosh. Shipping code towards data in an inter-region serverless environment to leverage latency. The Journal of Supercomputing, 79(10): 11585–11610, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05104-7.

Khanna:2023:DPA

[2220] Gaurav Khanna, Sanjay K. Chaturvedi, and Mohamed Othman. On design and performance analysis of improved shuffle exchange gamma interconnection network layouts. The Journal of Supercomputing, 79(11): 11611–11640, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04735-6.

Galiano:2023:UDL

[2221] Vicente Galiano, Héctor Migallón, Miguel Martínez-Rach, Otoniel López-Granado, and Manuel P. Malumbres. On the use of deep learning and parallelism techniques to significantly reduce the HEVC intra-coding time. The Journal of Supercomputing, 79 (11):11641–11659, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-022-04764-1.

Folino:2023:EBF

[2222] Gianluigi Folino, Carla Otranto Godano, and Francesco Sergio Pisani. An ensemble-based framework for user behaviour anomaly detection and classification for cybersecurity. The Journal of Supercomputing, 79(11): 11660–11683, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05049-x.

Makmuang:2023:SSV

[2223] Dawrawee Makmuang, Wachiraphong Ratiphaphongthon, and Rabian Wangkeeree. Smooth support vector machine with generalized pinball loss for Pattern Classification. The Journal of Supercomputing, 79(11):11684– 11706, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05082-w.

Choi:2023:DEI

[2224] Yurim Choi and Hangbae Chang. Design of evaluation items of the security levels for suppliers in the manufacturing industry. The Journal of Supercomputing, 79(11):11707– 11742, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05101-w.

Chen:2023:IPP

[2225] Yen-Chun Chen and Chin-Feng Lai. An intuitive pre-processing method based on human-robot interactions: zeroshot learning semantic segmentation based on synthetic semantic template. The Journal of Supercomputing, 79 (11):11743–11766, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05068-8.

Packiaraj:2023:HFI

[2226] Muneeswaran Packiaraj and Sriram Kailasam. HyPar-FCA+: an improved workload-aware elastic framework for FCA. The Journal of Supercomputing, 79(11):11767– 11796, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05116-3.

Ay:2023:CAM

[2227] Sevket Ay, Ekin Ekinci, and Zeynep Garip. A comparative analysis of metaheuristic optimization algorithms for feature selection on ML-based classification of heart-related diseases. The Journal of Supercomputing, 79(11): 11797–11826, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05132-3.

Ibanez:2023:PDM

[2228] Mario Ibáñez, Manuel Luna, Jose Luis Bosque, and Ramón Beivide. Parallelisation of decision-making techniques in aquaculture enterprises. The Journal of Supercomputing, 79(11): 11827–11843, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05124-3.

Arkan:2023:UHI

[2229] AhmadShahab Arkan and Mahmood Ahmadi. An unsupervised and hierarchical intrusion detection system for software-defined wireless sensor networks. The Journal of Supercomputing, 79(11):11844– 11870, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05117-2.

Siagh:2023:ISC

[2230] Asma Siagh, Fatima Zohra Laallam, Okba Kazar, and Hajer Salem. An improved sentiment classification model based on data quality and word embeddings. The Journal of Supercomputing, 79(11):11871– 11894, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05099-1.

Ruiz-Villafranca:2023:MMA

[2231] Sergio Ruiz-Villafranca, Javier Carrillo-Mondéjar, Juan Manuel Castelo Gómez, and José Roldán-Gómez. ME-CInOT: a multi-access edge computing and industrial internet of things emulator for the modelling and study of cybersecurity threats. The Journal of Supercomputing, 79(11):11895– 11933, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05098-2.

Yan:2023:MGD

[2232] Surong Yan, Haosen Wang, Yixiao Li, Chunqi Wu, Long Han, Chenglong Shi, and Ruilin Guo. Metapathguided dual semantic-aware filtering for HIN-based recommendation. The Journal of Supercomputing, 79(11): 11934–11964, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05113-6.

Bera:2023:CDE

[2233] Somnath Bera, Tanushree Dey, Anwesha Mukherjee, and Rajkumar Buyya. E-CropReco: a dew-edge-based multi-parametric crop recommendation framework for internet of agricultural things. The Journal of Supercomputing, 79(11):11965– 11999, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05131-4.

Afiya:2023:MOL

[2234] Syeda Afiya and M. Rajesh. MinLA of (K_{9} − C_{9})^{n} and its optimal layout into certain trees. The Journal of Supercomputing, 79(11):12000– 12012, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05140-3.

Kumar:2023:LBB

[2235] Ashish Kumar and Kakali Chatterjee. A lightweight blockchainbased framework for medical cyberphysical system. The Journal of Supercomputing, 79(11):12013– 12041, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05133-2.

Mirhosseini:2023:FMS

[2236] Mina Mirhosseini, Mahmood Fazlali, Mohammad K. Fallah, and Jeong-A Lee. A fast MILP solver for high-level synthesis based on heuristic model reduction and enhanced branch and bound algorithm. The Journal of Supercomputing, 79(11): 12042–12073, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05109-2.

Merdassi:2023:NLA

[2237] Imen Merdassi, Cherif Ghazel, and Leila Saidane. A new LTMA–ABE location and time access security control scheme for mobile cloud. The Journal of Supercomputing, 79(11): 12074–12105, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05107-4.

Meng:2023:GRB

[2238] Chunyun Meng, Xiaobing He, Zhen Tan, and Li Luan. Gait recognition based on 3D human body reconstruction and multi-granular feature fusion. The Journal of Supercomputing, 79 (11):12106–12125, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05143-0.

Elbaghazaoui:2023:PNW

[2239] Bahaa Eddine Elbaghazaoui, Mohamed Amnai, and Youssef Fakhri. Predicting the next word using the Markov chain model according to profiling personality. The Journal of Supercomputing, 79(11):12126– 12141, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05125-2.

Ebrahimpour:2023:HBP

[2240] Hossein Ebrahimpour, Mehrdad Ashtiani, Fatemeh Bakhshi, and Ghazaleh Bakhtiariazad. A heuristic-based package-aware function scheduling approach for creating a trade-off between cold start time and cost in FaaS computing environments. The Journal of Supercomputing, 79(11): 12142–12190, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05128-z.

Selvi:2023:ERS

[2241] K. Selvi, K. Muthumanickam, P. Vijayalakshmi, and P. C. Senthil Mahesh. ECC-reliant secure authentication protocol for cloud server and smart devices in IoT. The Journal of Supercomputing, 79(11):12191– 12218, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05135-0.

Saini:2023:CNF

[2242] A. Saini, A. Tsokanos, and R. Kirner.

CryptoQNRG: a new framework for evaluation of cryptographic strength in quantum and pseudorandom number generation for key-scheduling algorithms. The Journal of Supercomputing, 79(11):12219–12237, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05115-
4.

Hemavathi:2023:FFR

[2243] S. Hemavathi and B. Latha. FRHO: Fuzzy rule-based hybrid optimization for optimal cluster head selection and enhancing quality of service in wireless sensor network. The Journal of Supercomputing, 79(11): 12238–12265, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05106-5.

Xu:2023:CVI

[2244] Saijuan Xu, Canyang Guo, Yuhan Zhu, Genggeng Liu, and Neal Xiong. CNN-VAE: An intelligent text representation algorithm. The Journal of Supercomputing, 79(11):12266– 12291, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05139-w.

Kaur:2023:SSR

[2245] Prabhjot Kaur, Nitin Kumar, and Maheep Singh. SURFBCS: speeded up robust features based fuzzy vault scheme in biometric cryptosystem. The Journal of Supercomputing, 79(11): 12292–12316, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05142-1.

Fang:2023:WWS

[2246] Juan Fang, Li’ang Zhao, Min Cai, and Huijing Yang. WSMP: a warp scheduling strategy based on MFQ and PPF. The Journal of Supercomputing, 79 (11):12317–12340, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05127-0.

Li:2023:EDB

[2247] Wei Li, Junqing Yuan, and Lei Wang. An enhanced decompositionbased multiobjective evolutionary algorithm with adaptive neighborhood operator and extended distance-based environmental selection. The Journal of Supercomputing, 79(11):12341– 12393, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05118-1.

Dal:2023:IID

[2248] Deniz Dal and Esra Celik. Investigation of the impact of different versions of GCC on various metaheuristic-based solvers for traveling salesman problem. The Journal of Supercomputing, 79 (11):12394–12440, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05152-z.

Zagan:2023:FIH

[2249] Ionel Zagan and Vasile Gheorghita Gaitan. FPGA implementation of hardware accelerated RTOS based on real-time event handling. The Journal of Supercomputing, 79(11): 12441–12471, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05151-0. See cor-

rection [2352].

Kayali:2023:MLB

[2250] Devrim Kayali, Nemah Abu Shama, Suleyman Asir, and Kamil Dimililer. Machine learning-based models for the qualitative classification of potassium ferrocyanide using electrochemical methods. The Journal of Supercomputing, 79(11):12472– 12491, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05137-y.

Trivedi:2023:SAK

[2251] Chandan Trivedi and Udai Pratap Rao. Secrecy aware key management scheme for Internet of Healthcare Things. The Journal of Supercomputing, 79(11): 12492–12522, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05144-z.

Thungon:2023:LCB

[2252] Leki Chom Thungon, Subhas Chandra Sahana, and Md. Iftekhar Hussain. A lightweight certificate-based authentication scheme for 6LoWPANbased Internet of Things. The Journal of Supercomputing, 79(11):12523– 12548, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05089-3.

Lopez:2023:FTQ

[2253] Luis O. López, Francisco Orts, Gloria Ortega, Vicente González-Ruiz, and Ester M. Garzón. Fault-tolerant quantum algorithm for dual-threshold image segmentation. The Journal of Supercomputing, 79(11):12549– 12562, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05148-9.

Liu:2023:TAV

[2254] Bo Liu, Rui Chen, Weiwei Lin, Wentai Wu, Jianpeng Lin, and Keqin Li. Thermal-aware virtual machine placement based on multiobjective optimization. The Journal of Supercomputing, 79(11):12563– 12590, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05136-z. See pub-

lisher correction [2441].

Nekouie:2023:MFS

[2255] Nadia Nekouie, Morteza Romoozi, and Mahdi Esmaeili. Multimodal feature selection from microarray data based on Dempster–Shafer evidence fusion. The Journal of Supercomputing, 79 (11):12591–12621, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05138-x. See cor-

rection [2439].

Santhadevi:2023:SDL

[2256] D. Santhadevi and B. Janet. Stacked deep learning framework for edge– based intelligent threat detection in IoT network. The Journal of Supercomputing, 79(11):12622– 12655, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05153-y.

Orts:2023:EDQ

[2257] Francisco Orts, Gloria Ortega, Elías F. Combarro, Ignacio F. Rúa, Antonio M. Puertas, and Ester M. Garzón. Efficient design of a quantum absolutevalue circuit using Clifford+T gates. The Journal of Supercomputing, 79 (11):12656–12670, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05162-x.

Zare:2023:APM

[2258] Mansoureh Zare, Yasser Elmi Sola, and Hesam Hasanpour. An autonomous planning model for solving IoT service placement problem using the imperialist competitive algorithm. The Journal of Supercomputing, 79(11): 12671–12690, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05172-9.

Tran:2023:FSC

[2259] Duy Thanh Tran and Jun-Ho Huh.

Forecast of seasonal consumption behavior of consumers and privacypreserving data mining with new S-Apriori algorithm. The Journal of Supercomputing, 79(11):12691– 12736, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05105-6.

Palo:2023:FDS

[2260] Patitapaban Palo, Aurobinda Routray, Rahul Mahadik, and Sanjai Singh. Fault detection in seismic data using graph convolutional network. The Journal of Supercomputing, 79(11): 12737–12765, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05173-8.

Al-Bashabsheh:2023:ICD

[2261] Emran Al-Bashabsheh, Ahmad Alaiad, Mahmoud Al-Ayyoub, Othman Beni-Yonis, Raed Abu Zitar, and Laith Abualigah. Improving clinical documentation: automatic inference of ICD-10 codes from patient notes using BERT model. The Journal of Supercomputing, 79(11): 12766–12790, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05160-z.

Mahjoub:2023:OUN

[2262] Shabnam Mahjoub, Mehdi Golsorkhtabaramiri, and Seyed Sadegh Salehi Amiri. Optimal uniformization for non-uniform two-level loops using a hybrid method. The Journal of Supercomputing, 79(11):12791– 12814, July 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05194-3.

Yousef:2023:MBM

[2263] Reem N. Yousef, Abeer T. Khalil, Ahmed S. Samra, and Mohamed Maher Ata. Model-based and modelfree deep features fusion for high performed human gait recognition. The Journal of Supercomputing, 79(12): 12815–12852, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05156-9.

Jesus:2023:GPA

[2264] Lauê Jesus, Peterson Nogueira, João Speglich, and Murilo Boratto. GPU performance analysis for viscoacoustic wave equations using fast stencil computation from the symbolic specification. The Journal of Supercomputing, 79(12):12853– 12868, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05178-3.

Ukken:2023:SAB

[2265] Amal Francis V. Ukken, Arjun Bindu Jayachandran, Jaideep Kumar Punnath Malayathodi, and Pranesh Das. Statistically aided Binary Multi–Objective Grey Wolf Optimizer: a new feature selection approach for classification. The Journal of Supercomputing, 79(12):12869– 12901, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05145-y.

Wang:2023:PAQ

[2266] Shunzhi Wang, Zhanyou Ma, Rong Wang, and Wenming Fang. Performance analysis of a queueing system based on working vacation with repairable fault in the P2P network. The Journal of Supercomputing, 79(12): 12902–12923, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05154-x.

Gao:2023:CCM

[2267] Shuang Gao, Kaiwei Ma, Yang Gao, Zhenjiang Jiang, Xin Shen, Mingxing Yang, and Fengyu Xu. Combined compensation method of robot kinematics error based on MRIPN–IMA. The Journal of Supercomputing, 79(12): 12924–12948, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05177-4.

Gong:2023:SNS

[2268] Shuai Gong, Zhenfang Zhu, Jiangtao Qi, Wenqing Wu, and Chunling Tong. SeburSum: a novel set-based summary ranking strategy for summarylevel extractive summarization. The Journal of Supercomputing, 79(12): 12949–12977, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05165-8.

Lages:2023:CBM

[2269] Diogo Lages, Eric Borba, Eduardo Tavares, Andson Balieiro, and Erica Souza. A CPN-based model for assessing energy consumption of IoT networks. The Journal of Supercomputing, 79(12):12978– 13000, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05185-4.

Mohar:2023:EDT

[2270] Satinder Singh Mohar, Sonia Goyal, and Ranjit Kaur. Exploration of different topologies for optimal sensor nodes deployment in wireless sensor networks using Jaya-sine cosine optimization algorithm. The Journal of Supercomputing, 79(12):13001– 13030, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05147-w.

Nasersharif:2023:MLM

[2271] Babak Nasersharif, Manije Ebrahimpour, and Navid Naderi. Multilayer maximum mean discrepancy in auto-encoders for cross-corpus speech emotion recognition. The Journal of Supercomputing, 79(12):13031– 13049, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05161-y.

Liang:2023:CNL

[2272] Jiarong Liang, Weijian Zeng, and Xiaojiang Du. Construction of nodeand link-fault-tolerant virtual backbones in wireless networks. The Journal of Supercomputing, 79(12): 13050–13074, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05180-9.

Mozhdehi:2023:TED

[2273] Mahsa Hadikhah Mozhdehi and Amir-Masoud Eftekhari Moghadam. Textual emotion detection utilizing a transfer learning approach. The Journal of Supercomputing, 79(12):13075– 13089, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05168-5.

Batool:2023:PHD

[2274] Eliza Batool, Saira Gillani, Sheneela Naz, Maryam Bukhari, Muazzam Maqsood, Sang-Soo Yeo, and Seungmin Rho. POSNet: a hybrid deep learning model for efficient person re-identification. The Journal of Supercomputing, 79(12): 13090–13118, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05169-4.

Gunish:2023:HIP

[2275] Gunish Gunish, Sheema Madhusudhanan, and Arun Cyril Jose. Hybrid image processing model: a base for smart emergency applications. The Journal of Supercomputing, 79(12): 13119–13141, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05174-7.

Tiwari:2023:NSI

[2276] Anurag Tiwari and Vinay Kumar Srivastava. Novel schemes for the improvement of lifting wavelet transformbased image watermarking using Schur decomposition. The Journal of Supercomputing, 79(12):13142– 13179, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05167-6.

Tripathy:2023:AEF

[2277] Gyananjaya Tripathy and Aakanksha Sharaff. AEGA: enhanced feature selection based on ANOVA and extended genetic algorithm for online customer review analysis. The Journal of Supercomputing, 79(12):13180– 13209, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05179-2.

He:2023:ATS

[2278] Suna He, Jigang Wu, Bing Wei, and Jiaxin Wu. Algorithms for treeshaped task partition and allocation on heterogeneous multiprocessors. The Journal of Supercomputing, 79(12): 13210–13240, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05186-3.

Elnakib:2023:EDL

[2279] Omar Elnakib, Eman Shaaban, Mohamed Mahmoud, and Karim Emara. EIDM: deep learning model for IoT intrusion detection systems. The Journal of Supercomputing, 79(12): 13241–13261, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05197-0.

Chuang:2023:BVN

[2280] Po-Jen Chuang and Pang-Yu Huang.

B-VAE: a new dataset balancing approach using batched Variational AutoEncoders to enhance network intrusion detection. The Journal of Supercomputing, 79(12):13262– 13286, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05171-w.

Sawant:2023:ABP

[2281] Shrutika S. Sawant, F. X. Erick, St.

Göb, Nina Holzer, Elmar W. Lang, and Theresa Götz. An adaptive binary particle swarm optimization for solving multi-objective convolutional filter pruning problem. The Journal of Supercomputing, 79(12): 13287–13306, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05150-1.

Sun:2023:ASC

[2282] Keshuo Sun and Haiying Gao. Adaptively secure CP–ABE for circuits with fan-in n and fan-out 1. The Journal of Supercomputing, 79(12): 13307–13340, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05130-5.

Kia:2023:HHA

[2283] Keihaneh Kia and Amir Rajabzadeh. HDSAP: heterogeneity-aware dynamic scheduling algorithm to improve performance of nanoscale many-core processors for unknown workloads. The Journal of Supercomputing, 79(12): 13341–13369, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05159-6.

Vecil:2023:EGI

[2284] Francesco Vecil, José Miguel Mantas, and Pedro Alonso-Jordá. Efficient GPU implementation of a Boltzmann–Schrödinger–Poisson solver for the simulation of nanoscale DG MOSFETs. The Journal of Supercomputing, 79(12):13370– 13401, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05189-0.

daSilva:2023:ERN

[2285] Lucas Ferreira da Silva and João V. F. Lima. An evaluation of relational and NoSQL distributed databases on a low-power cluster. The Journal of Supercomputing, 79(12):13402– 13420, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05166-7.

Gong:2023:QQM

[2286] Changqing Gong, Hongsheng Zhu, Abdullah Gani, and Han Qi. QGA-QGCNN: a model of quantum gate circuit neural network optimized by quantum genetic algorithm. The Journal of Supercomputing, 79(12): 13421–13441, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05158-7.

Timucin:2023:CSC

[2287] Tunahan Timuçin and Serdar Biroğul.

Collaborative Smart Contracts (CoSC): example of real estate purchase and sale(s). The Journal of Supercomputing, 79(12):13442–13461, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05196-
1.

Zhang:2023:MMC

[2288] Chi Zhang, Shiqiang Nie, Jinyu Wang, Song Liu, and Weiguo Wu. MCB: a multidevice cooperative buffer management strategy for boosting the write performance of the SSD– SMR hybrid storage. The Journal of Supercomputing, 79(12):13462– 13489, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05201-7.

Li:2023:CCO

[2289] Hongjian Li, Dongjun Li, Xue Zhang, and Hu Sun. Cooperative computation offloading combined with data compression in mobile edge computing system. The Journal of Supercomputing, 79(12):13490– 13518, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05200-8.

Reddy:2023:EER

[2290] M. Venkata Krishna Reddy, P. V. S. Srinivas, and M. Chandra Mohan. Energy efficient routing with secure and adaptive trust threshold approach in mobile ad hoc networks. The Journal of Supercomputing, 79(12): 13519–13544, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05187-2.

Eskandari:2023:BTW

[2291] Hosein Eskandari, Maryam Imani, and Mohsen Parsa Moghaddam. Best-tree wavelet packet transform bidirectional GRU for short-term load forecasting. The Journal of Supercomputing, 79(12):13545– 13577, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05193-4.

BenAmor:2023:SLM

[2292] Arij Ben Amor, Sarra Jebri, Mohamed Abid, and Aref Meddeb. A secure lightweight mutual authentication scheme in Social Industrial IoT environment. The Journal of Supercomputing, 79(12):13578– 13600, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05176-5.

Xu:2023:RSP

[2293] Xiaojie Xu and Yun Zhang. Regional steel price index forecasts with neural networks: evidence from east, south, north, central south, northeast, southwest, and northwest China. The Journal of Supercomputing, 79(12): 13601–13619, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05207-1.

Kasturi:2023:OOS

[2294] Anirudh Kasturi and Chittaranjan Hota. OSGAN: One-shot distributed learning using generative adversarial networks. The Journal of Supercomputing, 79(12):13620– 13640, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05182-7.

Sun:2023:RAE

[2295] Xueli Sun, Jianxi Fan, Eminjan Sabir, Baolei Cheng, and Jia Yu. Reliability of augmented k-ary n-cubes under the extra connectivity condition. The Journal of Supercomputing, 79(12): 13641–13669, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05141-2.

Zhen:2023:FWS

[2296] Yan Zhen, Yuexian Li, Puning Zhang, Zhigang Yang, and Rongjian Zhao. Frequent words and syntactic context integrated biomedical discontinuous named entity recognition method. The Journal of Supercomputing, 79(12): 13670–13695, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05224-0.

Cai:2023:PTM

[2297] Linqin Cai, Haodu Fang, and Zhiqing Li. Pre-trained multilevel fuse network based on vision-conditioned reasoning and bilinear attentions for medical image visual question answering. The Journal of Supercomputing, 79(12): 13696–13723, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05195-2.

Zhang:2023:MAA

[2298] Maoyuan Zhang, Xiang Li, and Fei Wu. Moka-ADA: adversarial domain adaptation with modeloriented knowledge adaptation for cross-domain sentiment analysis. The Journal of Supercomputing, 79(12): 13724–13743, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05191-6.

Rafiee:2023:FMC

[2299] Mojtaba Rafiee. Flexible multi-client functional encryption for set intersection. The Journal of Supercomputing, 79(12):13744–13765, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05129-
y.

Guaman:2023:EEC

[2300] Daniel Guamán, Jennifer Pérez, and Priscila Valdiviezo-Diaz. Estimating the energy consumption of model-view-controller applications. The Journal of Supercomputing, 79(12):13766– 13793, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05202-6.

Zhu:2023:CEL

[2301] Feng Zhu, Jingjing Zhang, and Xin Wang. Communication-efficient local SGD with age-based worker selection. The Journal of Supercomputing, 79(12):13794–13816, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05190-
7.

Bustos:2023:RHH

[2302] Andrés Bustos, Antonio Juan Rubio-Montero, Roberto Méndez, Sergio Rivera, Francisco González, Xandra Campo, Hernán Asorey, and Rafael Mayo-García. Response of HPC hardware to neutron radiation at the dawn of exascale. The Journal of Supercomputing, 79(12):13817– 13838, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05199-y.

Fu:2023:PEM

[2303] Weihong Fu and Yue Chen. Parameters estimation method and ISAR imaging of multi-target with complex maneuvering. The Journal of Supercomputing, 79(12):13839– 13863, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05188-1.

Machovec:2023:SMS

[2304] Dylan Machovec, Howard Jay Siegel, James A. Crowder, Sudeep Pasricha, Anthony A. Maciejewski, and Ryan D. Friese. Surveillance mission scheduling with unmanned aerial vehicles in dynamic heterogeneous environments. The Journal of Supercomputing, 79(12):13864– 13888, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05225-z.

Zhou:2023:SUH

[2305] Jian Zhou, Weixin Wang, Jian Lu, and Lingzhe Liu. Small unmanned helicopter modeling method based on a hybrid kernel function PSO–LSSVM. The Journal of Supercomputing, 79(12): 13889–13906, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05211-5.

Wang:2023:RNA

[2306] Lei Wang. Retraction note: Application of deep learning to detect defects on the surface of steel balls in an IoT environment. The Journal of Supercomputing, 79(12): 13907, August 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05274-4.  See

[1665].

Yue:2023:RNR

[2307] Hongzhu Yue, Qijie Jiang, Chuanbin Yin, and Jonny Wilson. Retraction note: Research on data aggregation and transmission planning with Internet of Things technology in WSN multi-channel aware network. The Journal of Supercomputing, 79 (12):13908, August 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05353-6. See [202].

Ye:2023:AFA

[2308] Zongli Ye, Chin-Ling Chen, Wei Weng, Hongyu Sun, Woei-Jiunn Tsaur, and Yong-Yuan Deng. An anonymous and fair auction system based on blockchain. The Journal of Supercomputing, 79(13):13909– 13951, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05155-w.

Jha:2023:VTD

[2309] Pranava K. Jha. Vertex transitivity and distance metric of the quad-cube. The Journal of Supercomputing, 79(13):13952–13970, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05181-8.

Kuyu:2023:CIE

[2310] Yigit Çagatay Kuyu and Fahri Vatansever. A conceptual investigation of the effect of random numbers over the performance of metaheuristic algorithms. The Journal of Supercomputing, 79(13):13971– 14038, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05111-8.

Precioso:2023:TMN

[2311] Daniel Precioso and David Gómez-Ullate. Thresholding methods in nonintrusive load monitoring. The Journal of Supercomputing, 79(13):14039– 14062, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05149-8.

Singh:2023:CPN

[2312] Nikita Singh. CPU power and network bandwidth-aware optimal block size computation for blockchain-based applications using meta-heuristic algorithms. The Journal of Supercomputing, 79(13):14063–14078, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05210-6.

Yu:2023:GTE

[2313] Changyong Yu, Huimin Liu, Fazal Wahab, Zihan Ling, Tianmei Ren, Haitao Ma, and Yuhai Zhao. Global triangle estimation based on first edge sampling in large graph streams. The Journal of Supercomputing, 79(13):14079– 14116, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05205-3.

Raghav:2023:PTP

[2314] Raghav, Nitish Andola, Katyayani Verma, S. Venkatesan, and Shekhar Verma. Proactive threshold-proxy re-encryption scheme for secure data sharing on cloud. The Journal of Supercomputing, 79(13):14117–14145, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05221-3. See cor-

rection [2440].

Zhang:2023:PPP

[2315] Jing Zhang and Chuanwen Li. A practical privacy-preserving nearest neighbor searching method over encrypted spatial data. The Journal of Supercomputing, 79(13):14146– 14171, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05170-x.

Zhang:2023:CLC

[2316] Qi Zhang, Yi Liu, Tao Liu, and Depei Qian. CoFB: latency-constrained co-scheduling of flows and batches for deep learning inference service on the CPU–GPU system. The Journal of Supercomputing, 79(13):14172– 14199, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05183-6.

Golrasan:2023:PCM

[2317] Elham Golrasan and Marzieh Varposhti. Probabilistic coverage in mobile directional sensor networks: a game theoretical approach. The Journal of Supercomputing, 79(13):14200– 14220, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05231-1.

Liang:2023:MMB

[2318] Xiufang Liang, Yingzheng Zhu, Huajuan Duan, Fuyong Xu, Peiyu Liu, and Ran Lu. MISR: a multiple behavior interactive enhanced learning model for social-aware recommendation. The Journal of Supercomputing, 79(13):14221–14244, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05175-6. See pub-

lisher correction [2442].

Abdi:2023:IIS

[2319] Yousef Abdi and Mohammad Asadpour. On the impact of informationsharing model between subpopulations in the Island-based evolutionary algorithms: search manager framework as a case study. The Journal of Supercomputing, 79(13):14245– 14286, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05218-y.

Maronas:2023:MNE

[2320] Marcos Maroñas, Antoni Navarro, Eduard Ayguadé, and Vicenç Beltran. Mitigating the NUMA effect on taskbased runtime systems. The Journal of Supercomputing, 79(13):14287– 14312, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05164-9.

Zhang:2023:LLO

[2321] Yonghua Zhang, Hongxu Jiang, Yuting Zhu, Runhua Zhang, Yongxiang Cao, Chenhui Zhu, Wei Wang, Dong Dong, and Xiaobin Li. LOCP: Latency-optimized channel pruning for CNN inference acceleration on GPUs. The Journal of Supercomputing, 79(13):14313–14341, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05212-
4.

Cabanas-Molero:2023:MDM

[2322] Pablo Cabañas-Molero, Antonio J.

Muñoz-Montoro, Pedro Vera-Candeas, and José Ranilla. The music demixing machine: toward real-time remixing of classical music. The Journal of Supercomputing, 79(13):14342– 14357, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05192-5.

Jaberi:2023:PPM

[2323] Mehrad Jaberi and Hamid Mala. Privacy-preserving multi-party PCA computation on horizontally and vertically partitioned data based on outsourced QR decomposition. The Journal of Supercomputing, 79(13):14358– 14387, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05206-2.

Zhao:2023:EDT

[2324] Xiongfei Zhao, Gerui Zhang, and Yain-Whar Si. An efficient dynamic transaction storage mechanism for sustainable high-throughput Bitcoin. The Journal of Supercomputing, 79(13):14388– 14426, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05237-9.

Zadehbagheri:2023:ECO

[2325] Mahmoud Zadehbagheri and Ali Reza Abbasi. Energy cost optimization in distribution network considering hybrid electric vehicle and photovoltaic using modified whale optimization algorithm. The Journal of Supercomputing, 79(13):14427– 14456, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05214-2.

Javidi:2023:NNM

[2326] Mohammad Javidi and Mahdi Saedshoar Heris. New numerical methods for solving the partial fractional differential equations with uniform and non-uniform meshes. The Journal of Supercomputing, 79(13):14457– 14488, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05198-z.

Braik:2023:EWO

[2327] Malik Braik, Mohammed Awadallah, Mohammed Azmi Al-Betar, and Heba Al-Hiary. Enhanced whale optimization algorithm-based modeling and simulation analysis for industrial system parameter identification. The Journal of Supercomputing, 79(13):14489–14544, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05215-1.

Xu:2023:EIP

[2328] Fuyong Xu, Zhaoxin Ding, Zhenfang Zhu, and Peiyu Liu. Exploring implicit persona knowledge for personalized dialogue generation. The Journal of Supercomputing, 79(13):14545– 14570, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05209-z.

Wang:2023:CWD

[2329] Guan Wang, Jie-Sheng Wang, Hong-Yu Wang, and Jia-Xu Liu. Componentwise design method of fuzzy Cmeans clustering validity function based on CRITIC combination weighting. The Journal of Supercomputing, 79(13):14571–14601, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-023-05234-
                                        y.

Raza:2023:UAU

[2330] Ali Raza, Zeshan Iqbal, and Farhan Aadil. UAV-assisted ubiquitous communication architecture for urban VANET environment. The Journal of Supercomputing, 79(13):14602– 14632, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05223-1.

Ghosh:2023:TWM

[2331] Subha Ghosh and Debashis De. TARA: weighted majority cooperative game theory-based task assignment and resource allocation in 5G heterogeneous fog network for IoT. The Journal of Supercomputing, 79(13):14633– 14683, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05228-w.

Kuhn:2023:SVS

[2332] Martin J. Kühn, Johannes Holke, Annette Lutz, Jonas Thies, Melven Röhrig-Zöllner, Alexander Bleh, Jan Backhaus, and Achim Basermann. SIMD vectorization for simultaneous solution of locally varying linear systems with multiple right-hand sides. The Journal of Supercomputing, 79(13):14684–14706, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05220-4.

Bayraktar:2023:RTA

[2333] Cihan Bayraktar, Ziya Karakaya, and Hadi Gökçen. Real-time anomaly detection system within the scope of smart factories. The Journal of Supercomputing, 79(13):14707– 14742, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05236-w.

Zhang:2023:TRM

[2334] Xiaohong Zhang, Wenqi Du, and Ata Jahangir Moshayedi. A traceable and revocable multi-authority attribute-based access control scheme for mineral industry data secure storage in blockchain. The Journal of Supercomputing, 79(13):14743– 14779, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05222-2.

Ma:2023:QAF

[2335] Zongmin Ma, Xiaowen Zhang, and Yuhan Zhao. Queries with aggregate functions over fuzzy RDF data. The Journal of Supercomputing, 79(13):14780–14807, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-023-05235-
                                        x.

Dewangan:2023:TBS

[2336] Narendra K. Dewangan and Preeti Chandrakar. TempChain: a blockchain scheme for telehealth data sharing between two blockchains using property mapping function. The Journal of Supercomputing, 79(13):14808– 14826, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05259-3.

Ouyang:2023:UUR

[2337] Yiming Ouyang, Jiaxin Wang, Chenglong Sun, Qi Wang, and Huaguo Liang. URMP: using reconfigurable multicast path for NoC-based deep neural network accelerators. The Journal of Supercomputing, 79(13):14827– 14847, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05255-7.

Cheng:2023:PCB

[2338] Dongqin Cheng. Path covers of bubble-sort star graphs. The Journal of Supercomputing, 79(13):14848– 14868, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05256-6.

Bai:2023:YAO

[2339] Yuxuan Bai, Mingshuai Dong, Shimin Wei, Jian Li, and Xiuli Yu. YOLOOD: an arbitrary-oriented flexible flat cable detection method in robotic assembly. The Journal of Supercomputing, 79(13):14869–14893, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05254-8.

Tyagi:2023:GWH

[2340] Vikas Tyagi and Samayveer Singh.

GM-WOA: a hybrid energy efficient cluster routing technique for SDN-enabled WSNs. The Journal of Supercomputing, 79(13):14894– 14922, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05263-7.

Sachan:2023:EHA

[2341] Anuj Sachan and Neetesh Kumar.

S-Edge: heterogeneity-aware, lightweighted, and edge computing integrated adaptive traffic light control framework. The Journal of Supercomputing, 79(13):14923–14953, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05216-0.

Asghari:2023:EAE

[2342] Ali Asghari, Marjan Sayadi, and Hossein Azgomi. Energy-aware edge server placement using the improved butterfly optimization algorithm. The Journal of Supercomputing, 79(13):14954– 14980, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05271-7.

Ayyildiz:2023:SSS

[2343] Tülin Erçelebi Ayyildiz and Elmas Burcu Mamak Ekinci. Selection of Six Sigma projects based on integrated multi-criteria decision-making methods: the case of the software development industry. The Journal of Supercomputing, 79(13):14981– 15003, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05250-y.

Hong:2023:CCP

[2344] Hanshu Hong and Zhixin Sun. Constructing conditional PKEET with verification mechanism for data privacy protection in intelligent systems. The Journal of Supercomputing, 79(13):15004–15022, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05253-
9.

Jokandan:2023:PPA

[2345] Seyed Mohsen Ebadi Jokandan, Peyman Bayat, and Mehdi Farrokhbakht Foumani. Predicting product advertisement links using hybrid learning within social networks. The Journal of Supercomputing, 79(13):15023– 15050, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05213-3.

Freire:2023:GMA

[2346] Manuel Freire, Juan Ferrand, Franco Seveso, Ernesto Dufrechou, and Pablo Ezzatti. A GPU method for the analysis stage of the SP-TRSV kernel. The Journal of Supercomputing, 79(13):15051–15078, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05238-8.

Ali:2023:SUF

[2347] Shallaw Mohammed Ali and Gabor Kecskemeti. SeQual: an unsupervised feature selection method for cloud workload traces. The Journal of Supercomputing, 79(13):15079– 15097, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05163-w.

Doan:2023:SIH

[2348] Thi Van Thao Doan, Mohamed-Lamine Messai, Gérald Gavin, and Jérôme Darmont. A survey on implementations of homomorphic encryption schemes. The Journal of Supercomputing, 79(13):15098–15139, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05233-z.

Sudha:2023:HPG

[2349] V. Sudha and K. S. Easwarakumar. Hexagonal picture generation in DNA computing. The Journal of Supercomputing, 79(13):15140– 15184, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05247-7.

Qu:2023:JOS

[2350] Xiaofeng Qu and Huiqiang Wang. A joint optimization scheme of task caching and offloading for smart factories. The Journal of Supercomputing, 79(13):15185–15211, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05257-5.

Sadeghi:2023:CICb

[2351] Maryam Sadeghi, Mohammad Naderi Dehkordi, Behrang Barekatain, and Naser Khani. Correction to: Improve customer churn prediction through the proposed PCA–PSO-K means algorithm in the communication industry. The Journal of Supercomputing, 79 (13):15212, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05204-4.  See

[2028].

Zagan:2023:CFI

[2352] Ionel Zagan and Vasile Gheorghita Gaitan. Correction to: FPGA implementation of hardware accelerated RTOS based on real-time event handling.

The Journal of Supercomputing, 79 (13):15213, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05229-9.  See

[2249].

Firouznia:2023:ACC

[2353] Marjan Firouznia, Pietro Ruiu, and Giuseppe A. Trunfio. Adaptive cooperative coevolutionary differential evolution for parallel feature selection in high-dimensional datasets. The Journal of Supercomputing, 79(14):15215– 15244, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05226-y.

Liu:2023:ATF

[2354] Tianbo Liu and Jindong Zhang.

An adaptive traffic flow prediction model based on spatiotemporal graph neural network. The Journal of Supercomputing, 79(14):15245– 15269, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05261-9.

Zeng:2023:MHH

[2355] Liang Zeng, Yanyan Li, Hao Zhang, Ming Li, and Shanshan Wang. A mixed Harris hawks optimization algorithm based on the pinhole imaging strategy for solving numerical optimization problems. The Journal of Supercomputing, 79(14):15270– 15323, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05260-w.

Soltani:2023:FNF

[2356] Mohammad Soltani, Behrang Barekatain, Faramarz Hendesi, and Zahra Beheshti. FSCN: a novel forwarding method based on Shannon entropy and COPRAS decision process in named data networking. The Journal of Supercomputing, 79(14):15324– 15357, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05265-5.

Zhao:2023:NCP

[2357] Lingxiao Zhao, Zhiyang Li, Yue Ma, and Leilei Qu. A novel cryptocurrency price time series hybrid prediction model via machine learning with MATLAB/Simulink. The Journal of Supercomputing, 79(14):15358– 15389, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05242-y.

Feng:2023:BBB

[2358] Pengbin Feng, Li Yang, Di Lu, Ning Xi, and Jianfeng Ma. BejaGNN: behavior-based Java malware detection via graph neural network. The Journal of Supercomputing, 79(14):15390– 15414, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05243-x.

He:2023:TPD

[2359] Xinfeng He and Shuchao Sun. A two-phase detection method against APT attack on flow table management in SDN. The Journal of Supercomputing, 79(14):15415–15434, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05281-5.

Sharma:2023:RFL

[2360] Mradula Sharma and Parmeet Kaur.

Reliable federated learning in a cloudfog-IoT environment. The Journal of Supercomputing, 79(14):15435– 15458, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05252-w.

Ajmera:2023:SPS

[2361] Kashav Ajmera and Tribhuwan Kumar Tewari. SR-PSO: server residual efficiency-aware particle swarm optimization for dynamic virtual machine scheduling. The Journal of Supercomputing, 79(14):15459–15495, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05270-8.

Jiang:2023:DND

[2362] Luyu Jiang, Dantong Ouyang, Huisi Zhou, Naiyu Tian, and Liming Zhang. DPAHMA: a novel dual-population adaptive hybrid memetic algorithm for non-slicing VLSI floorplans. The Journal of Supercomputing, 79(14):15496– 15534, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05277-1.

Bergil:2023:CAM

[2363] Erhan Bergil, Canan Oral, and Engin Ufuk Ergül. Classification of arithmetic mental task performances using EEG and ECG signals. The Journal of Supercomputing, 79(14):15535– 15547, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05294-0.

Thaherbasha:2023:UAE

[2364] Shaik Thaherbasha, Ravindra Dhuli, S. D. Nageena Parveen, and A. Hareesh. A unified approach to evaluate the precise outage probability of NOMA over various fading scenarios. The Journal of Supercomputing, 79(14):15548–15578, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05275-3.

Zhi:2023:ATP

[2365] ChunYu Zhi, HuaiJiang Sun, and Tian Xu. Adaptive trajectory prediction without catastrophic forgetting. The Journal of Supercomputing, 79(14):15579–15596, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05241-z.

HamaKarim:2023:MBS

[2366] Bakhtyar Rafeeq HamaKarim, Rojiar Pir Mohammadiani, Amir Sheikhahmadi, Bryar Rafiq Hamakarim, and Mehri Bahrami. A method based on k-shell decomposition to identify influential nodes in complex networks. The Journal of Supercomputing, 79(14):15597–15622, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05296-y.

Medina:2023:SHS

[2367] Marie Chantelle Cruz Medina and João Fausto L. de Oliveira. A selective hybrid system for stateof-charge forecasting of lithium-ion batteries. The Journal of Supercomputing, 79(14):15623–15642, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05285-1.

Mandloi:2023:QLB

[2368] Dilip Mandloi and Rajeev Arya. Qlearning-based UAV-mounted base station positioning in a disaster scenario for connectivity to the users located at unknown positions. The Journal of Supercomputing, 79(14):15643– 15674, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05292-2.

Gupta:2023:BEH

[2369] Aditya Gupta, Monu Bhagat, and Vibha Jain. Blockchain-enabled healthcare monitoring system for early Monkeypox detection. The Journal of Supercomputing, 79(14):15675– 15699, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05288-y.

Hayat:2023:MLB

[2370] Asad Hayat, Yasir Noman Khalid, Muhammad Siraj Rathore, and Muhammad Nadeem Nadir. A machine learning-based resource-efficient task scheduler for heterogeneous computer systems. The Journal of Supercomputing, 79(14):15700–15728, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05266-4.

Liu:2023:MMC

[2371] Yifan Liu, Jincai Chen, Ping Lu, Chuanbo Zhu, Yugen Jian, Chao Sun, and Han Liang. MOON-LIT: momentum-contrast and largekernel for multi-fine-grained deraining. The Journal of Supercomputing, 79(14):15729–15759, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05286-0.

Wu:2023:ELS

[2372] Jiayang Wu, Shihong Huang, Yanxu Yang, and Bingzhi Zhang. Evaluation of 3D LiDAR SLAM algorithms based on the KITTI dataset. The Journal of Supercomputing, 79(14):15760– 15772, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05267-3.

Guenifi:2023:OTC

[2373] Naima Guenifi, Shiromani Balmukund Rahi, Faiza Benmahdi, and Houda Chabane. Optimization of tunneling current in ferroelectric tunnel FET using genetic algorithm. The Journal of Supercomputing, 79(14):15773– 15789, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05240-0.

Thao:2023:PDL

[2374] Le Quang Thao, Nguyen Duy Thien, Ngo Chi Bach, Duong Duc Cuong, Le Duc Anh, Dang Gia Khanh, Nguyen Ha Minh Hieu, and Nguyen Trieu Hoang Minh. PesViT: a deep learning approach for detecting misuse of pesticides on farm. The Journal of Supercomputing, 79(14):15790– 15813, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05302-3.

Dehkordi:2023:ESC

[2375] Afsaneh Banitalebi Dehkordi. Examining the status of CPU working load, processing load and controller bandwidth under the influence of packet-in buffer status located in openflow switches in SDNbased IoT framework. The Journal of Supercomputing, 79(14):15814– 15834, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05258-4.

Cheng:2023:NIC

[2376] Xiaoyu Cheng, Peng Han, Wei He, and Guohui Zhou. A new interval constructed belief rule base with rule reliability. The Journal of Supercomputing, 79(14):15835– 15867, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05284-2.

Talaat:2023:EAV

[2377] Amira Samy Talaat and Shaker El-Sappagh. Enhanced aerial vehicle system techniques for detection and tracking in fog, sandstorm, and snow conditions. The Journal of Supercomputing, 79(14):15868– 15893, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05245-9.

Wang:2023:DSG

[2378] Haoyu Wang and Jianwei An. Dynamic stochastic game-based security of edge computing based on blockchain. The Journal of Supercomputing, 79(14):15894–15926, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05289-x.

Li:2023:MFA

[2379] Xingxing Li, Weidong Li, and Xuejie Zhang. Multiresource fair allocation with time window constraints. The Journal of Supercomputing, 79(14):15927–15954, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05248-6.

Zhang:2023:RAM

[2380] Yaofang Zhang, Zibo Wang, Yingzhou Wang, Kuan Lin, Tongtong Li, Hongri Liu, Chao Li, and Bailing Wang. A risk assessment model for similar attack scenarios in industrial control system. The Journal of Supercomputing, 79(14):15955– 15979, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05269-1.

Liu:2023:FEC

[2381] Xuemei Liu, Jixiang Meng, and Yingzhi Tian. On forcibly k-edge-connected and forcibly super edgeconnected uniform hypergraphic sequences. The Journal of Supercomputing, 79(14):15980–15996, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05287-z.

Zhou:2023:HMA

[2382] Xinrong Zhou, Fang Wang, Chao Zhou, and Rui Shan. The HSGWO– MPIO algorithm based on improved search capability. The Journal of Supercomputing, 79(14):15997–16016, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05246-8.

Al-hajjar:2023:OML

[2383] Alaa Lateef Noor Al-hajjar and Ali Kadhum M. Al-Qurabat. An overview of machine learning methods in enabling IoMT-based epileptic seizure detection. The Journal of Supercomputing, 79(14):16017– 16064, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05299-9.

Mars:2023:SAA

[2384] Rawya Mars, Saoussen Cheikhrouhou, Slim Kallel, and Ahmed Hadj Kacem. A survey on automation approaches of smart contract generation. The Journal of Supercomputing, 79(14):16065– 16097, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05262-8.

Xiong:2023:IOA

[2385] Yijin Xiong, Xin Gao, and Guoying Zhang. Interactive object annotation based on one-click guidance. The Journal of Supercomputing, 79(14):16098– 16117, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05279-z.

Louati:2023:ECP

[2386] Hassen Louati, Ali Louati, Slim Bechikh, and Elham Kariri. Embedding channel pruning within the CNN architecture design using a bilevel evolutionary approach. The Journal of Supercomputing, 79(14):16118– 16151, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05273-5.

Bakhtiari:2023:TTW

[2387] Niloofar Barati Bakhtiari, Masood Rafighi, and Reza Ahsan. TTLA: two-way trust between clients and fog servers using Bayesian learning automata. The Journal of Supercomputing, 79(14):16152–16180, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05264-6.

Ashraf:2023:LAS

[2388] Zeeshan Ashraf, Adnan Sohail, and Muhammad Yousaf. Lightweight and authentic symmetric session key cryptosystem for client-server mobile communication. The Journal of Supercomputing, 79(14):16181– 16205, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05239-7.

St-Onge:2023:MOF

[2389] Cédric St-Onge, Nadjia Kara, and Claes Edstrom. Multivariate outlier filtering for A-NFVLearn: an advanced deep VNF resource usage forecasting technique. The Journal of Supercomputing, 79(14):16206– 16232, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05283-3.

Bhan:2023:BES

[2390] Rati Bhan, Rajendra Pamula, Parvez Faruki, and Jyoti Gajrani. Blockchainenabled secure and efficient data sharing scheme for trust management in healthcare smartphone network. The Journal of Supercomputing, 79(14):16233–16274, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05272-
6.

Chen:2023:POM

[2391] Renjiang Chen, Tao Liu, Zhaoyuan Liu, Li Wang, Min Tian, Ying Guo, Jingshan Pan, Xiaoming Wu, and Meihong Yang. Parallel optimization of method of characteristics based on Sunway Bluelight II supercomputer. The Journal of Supercomputing, 79(14):16275–16299, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05313-0.

Wang:2023:EDP

[2392] Xi Wang, Xinzhi Hu, Weibei Fan, and Ruchuan Wang. Efficient data persistence and data division for distributed computing in cloud data center networks. The Journal of Supercomputing, 79(14):16300–16327, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05276-2.

Karotia:2023:CUT

[2393] Akanksha Karotia and Seba Susan. CovSumm: an unsupervised transformer-cum-graph-based hybrid document summarization model for CORD-19. The Journal of Supercomputing, 79(14):16328–16350, September 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05291-3.

Jayanthan:2023:ACT

[2394] K. S. Jayanthan and S. Domnic. An attentive convolutional transformer-based network for road safety. The Journal of Supercomputing, 79(14):16351–16377, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05293-
1.

Chauhan:2023:NCM

[2395] Parul Chauhan, Anjana Gupta, and Tanya Malhotra. A novel cloud model based on multiplicative unbalanced linguistic term set. The Journal of Supercomputing, 79(14):16378– 16408, September 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05295-z.

Das:2023:ICF

[2396] Nayan Ranjan Das, Imon Mukherjee, Anubhav D. Patel, and Goutam Paul. An intelligent clustering framework for substitute recommendation and player selection. The Journal of Supercomputing, 79(15):16409– 16441, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05314-z.

Bahramlou:2023:ECF

[2397] Ainaz Bahramlou, Massoud Reza Hashemi, and Zeinab Zali. Ensemble clustering and feature weighting in time series data. The Journal of Supercomputing, 79(15):16442– 16478, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05290-4.

Roopa:2023:MMP

[2398] Vuppula Roopa and Himansu Shekhar Pradhan. Mathematical modeling and performance evaluation of Be-RAN for 6G wireless networks. The Journal of Supercomputing, 79(15): 16479–16528, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05321-0.

DiTullio:2023:PQK

[2399] Daniele Di Tullio and Manoj Gyawali. A post-quantum key exchange protocol from the intersection of quadric surfaces. The Journal of Supercomputing, 79(15):16529–16558, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05146-
x.

Zhang:2023:TQM

[2400] Junbo Zhang and Cheng Zhang. Teaching quality monitoring and evaluation of physical education teaching in ordinary college based on edge computing optimization model. The Journal of Supercomputing, 79(15): 16559–16579, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05324-x.

Al-Omaisi:2023:IID

[2401] Hussein Al-Omaisi, Elankovan A. Sundararajan, Raed Alsaqour, Nor Fadzilah Abdullah, Khairul Azmi Abu Bakar, and Maha Abdelhaq. ID-TracS: an interest–data-flow trackingbased forwarding scheme for vehicular named data networks. The Journal of Supercomputing, 79(15): 16580–16615, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05268-2.

Albahrani:2023:NSE

[2402] Ekhlas Abbas Albahrani, Tayseer Karam Alshekly, and Sadeq H. Lafta. New secure and efficient substitution and permutation method for audio encryption algorithm. The Journal of Supercomputing, 79(15):16616– 16646, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05249-5.

Liang:2023:FFG

[2403] Hong Liang, Shaoshuai Han, Mingwen Shao, and Qian Zhang. FGP-GAN: a finer-grained CNN pruning via generative adversarial network. The Journal of Supercomputing, 79(15): 16647–16663, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05320-1.

Lai:2023:OHS

[2404] Chun-Ming Lai, Ting-Wei Chang, and Chao-Tung Yang. An online and highly-scalable streaming platform for filtering trolls with transfer learning. The Journal of Supercomputing, 79(15):16664– 16687, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05312-1.

Roy:2023:ACT

[2405] Chitrangada Roy and Dushmanta Kumar Das. Ameliorated class topper optimizer for cost optimization using demand side management program in a day-ahead energy market. The Journal of Supercomputing, 79(15): 16688–16719, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05297-x.

Kyriakopoulos:2023:HDW

[2406] Christos Kyriakopoulos, Efstratios Gallopoulos, and Ioannis E. Venetis. Hierarchical dynamic workload scheduling on heterogeneous clusters for grid search of inverse problems. The Journal of Supercomputing, 79(15): 16720–16772, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05306-z.

Deb:2023:IBD

[2407] Priti Deb, Anwesha Mukherjee, Debashis De, and Soumya K. Ghosh. IoBT: beamforming design in internet of things. The Journal of Supercomputing, 79(15):16773– 16792, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05307-y.

Sheibani:2023:FDT

[2408] Samaneh Sheibani, Hassan Shakeri, and Reza Sheibani. Fourdimensional trust propagation model for improving the accuracy of recommender systems. The Journal of Supercomputing, 79(15):16793– 16820, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05278-0.

Ghosh:2023:DDC

[2409] Subha Ghosh and Debashis De.

DewGame: D2D communication enabled dew computing for 5G IoT using coalition formation game. The Journal of Supercomputing, 79(15): 16821–16858, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05311-2.

Huang:2023:TVC

[2410] Haijun Huang, Hong Zhang, Zidi Zhao, and Shouhua Zhang. A traceable and verifiable CP–ABE scheme with blockchain in VANET. The Journal of Supercomputing, 79(15): 16859–16883, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05322-z.

Singh:2023:FTP

[2411] Raushan Kumar Singh and Mukesh Kumar. Future trends of path planning framework considering accident attributes for smart cities. The Journal of Supercomputing, 79(15): 16884–16913, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05305-0. See cor-

rections [2573, 2672].

Meshram:2023:ECG

[2412] Chandrashekhar Meshram, Rabha W. Ibrahim, Preecha Yupapin, Ismail Bahkali, Agbotiname Lucky Imoize, and Sarita Gajbhiye Meshram. An efficient certificateless group signcryption scheme using Quantum Chebyshev Chaotic Maps in HC– IoT environments. The Journal of Supercomputing, 79(15):16914– 16939, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05303-2.

Khedr:2023:HDJ

[2413] Ahmed M. Khedr, S. Sheeja Rani, and Mohamed Saad. Hybridized Dragonfly and Jaya algorithm for optimal sensor node location identification in mobile wireless sensor networks. The Journal of Supercomputing, 79(15): 16940–16962, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05326-9.

Hussein:2023:DPC

[2414] Ahmed Mohammed Hussein, Ali Kadhum Idrees, and Raphaël Couturier. A distributed prediction-compression-based mechanism for energy saving in IoT networks. The Journal of Supercomputing, 79(15):16963– 16999, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05317-w.

Emil:2023:DEA

[2415] Demyana Emil, Mohammed Hamdy, and Gihan Nagib. Development an efficient AXI-interconnect unit between set of customized peripheral devices and an implemented dualcore RISC-V processor. The Journal of Supercomputing, 79(15):17000– 17019, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05304-1.

Yan:2023:EAD

[2416] Kaixiang Yan, Jiatian Mei, Dongming Zhou, and Lifen Zhou. Externalattention dual-modality fusion network for RGBT tracking. The Journal of Supercomputing, 79(15):17020– 17041, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05329-6.

Xuan:2023:NAS

[2417] Cho Do Xuan, Dao Hoang Mai, Ma Cong Thanh, and Bui Van Cong. A novel approach for software vulnerability detection based on intelligent cognitive computing. The Journal of Supercomputing, 79(15):17042– 17078, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05282-4.

Shilpa:2023:LDD

[2418] B. Shilpa, Puranam Revanth Kumar, and Rajesh Kumar Jha. LoRa DL: a deep learning model for enhancing the data transmission over LoRa using autoencoder. The Journal of Supercomputing, 79(15):17079– 17097, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05355-4.

Wu:2023:COH

[2419] Ruizhi Wu and Bo Li. Computation offloading and heterogeneous task caching in MEC-enabled vehicular networks. The Journal of Supercomputing, 79(15):17098– 17122, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05328-7.

Labidi:2023:VPT

[2420] Taher Labidi and Zaineb Sakhrawi. On the value of parameter tuning in stacking ensemble model for software regression test effort estimation. The Journal of Supercomputing, 79(15): 17123–17145, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05334-9.

Yadav:2023:HKH

[2421] Pooja Yadav and S. C. Sharma.

HFBO-KSELM: Hybrid Flash Butterfly Optimization-based Kernel Softplus Extreme Learning Machine for classification of chronic kidney disease. The Journal of Supercomputing, 79(15): 17146–17169, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05337-6.

Yao:2023:LOT

[2422] Liang Yao, Huaguo Liang, and Yingchun Lu. Low-overhead TRNG based on MUX for cryptographic protection using multiphase sampling. The Journal of Supercomputing, 79(15): 17170–17186, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05349-2.

Gong:2023:DRA

[2423] Changqing Gong, Wanying He, Ting Wang, Abdullah Gani, and Han Qi. Dynamic resource allocation scheme for mobile edge computing. The Journal of Supercomputing, 79(15): 17187–17207, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05323-y.

Yadav:2023:ADC

[2424] Sonal Yadav, Vijay Laxmi, Hemangee Kapoor, Manoj Singh Gaur, and Amit Kumar. Adaptive distribution of control messages for improving bandwidth utilization in multiple NoC. The Journal of Supercomputing, 79(15): 17208–17246, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05208-0.

Casas-Ordaz:2023:IOB

[2425] Angel Casas-Ordaz, Diego Oliva, Mario A. Navarro, Alfonso Ramos-Michel, and Marco Pérez-Cisneros. An improved opposition-based Runge Kutta optimizer for multilevel image thresholding. The Journal of Supercomputing, 79(15):17247– 17354, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05227-x.

Umair:2023:VSA

[2426] Areeba Umair, Elio Masciari, and Muhammad Habib Ullah. Vaccine sentiment analysis using BERT + NB-SVM and geo-spatial approaches. The Journal of Supercomputing, 79(15): 17355–17385, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05319-8.

Oztekin:2023:UFB

[2427] Halit Öztekin, Abdelkader Lazzem, and Ihsan Pehlivan. Using FPGAbased content-addressable memory for mnemonics instruction searching in assembler design. The Journal of Supercomputing, 79(15):17386– 17418, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05357-2.

Dong:2023:PPC

[2428] Dong Dong, Hongxu Jiang, Hanqun Lin, and Yanfei Song. PCGC: a performance compact graph compiler based on multilevel fusion-splitting rules. The Journal of Supercomputing, 79(15): 17419–17444, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05298-w.

Tang:2023:PBC

[2429] Minghua Tang, Enrico Russo, and Maurizio Palesi. The positionbased compression techniques for DNN model. The Journal of Supercomputing, 79(15):17445–17474, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05339-

4. See correction [2623].

Mohammadzadeh:2023:STP

[2430] Mojtaba Mohammadzadeh, Abdoul-Ahad Choupani, and Farshid Afshar. The short-term prediction of daily traffic volume for rural roads using shallow and deep learning networks: ANN and LSTM. The Journal of Supercomputing, 79(15):17475– 17494, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05333-w.

Mukherjee:2023:MSM

[2431] Anwesha Mukherjee, Shreya Ghosh, Soumya K. Ghosh, and Rajkumar Buyya. Mobi-sense: mobilityaware sensor-fog paradigm for missioncritical applications using network coding and steganography. The Journal of Supercomputing, 79(15):17495– 17518, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05300-5.

Mora:2023:AAI

[2432] Manuel Mora, Olayele Adelakun, Paola Yuritzy Reyes-Delgado, and Oswaldo Diaz. AVS FD MVITS: an agile IT service design workflow for small data centers. The Journal of Supercomputing, 79(15):17519– 17561, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05244-w.

Yan:2023:TCI

[2433] Fengru Yan, Guanghua Zhang, Dongwen Zhang, Xinghua Sun, Botao Hou, and Naiwen Yu. TL-CNN-IDS: transfer learning-based intrusion detection system using convolutional neural network. The Journal of Supercomputing, 79(15):17562– 17584, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05347-4.

Naim:2023:NCS

[2434] M. Naim and A. Ali Pacha. A new chaotic satellite image encryption algorithm based on a 2D filter and Fisher–Yates shuffling. The Journal of Supercomputing, 79(15): 17585–17618, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05346-5.

Jazaeri:2023:CCC

[2435] Seyedeh Shabnam Jazaeri, Parvaneh Asghari, Sam Jabbehdari, and Hamid Haj Seyyed Javadi. Composition of caching and classification in edge computing based on quality optimization for SDN-based IoT healthcare solutions. The Journal of Supercomputing, 79(15):17619– 17669, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05332-x.

Alonso-Mencia:2023:ACV

[2436] Javier Alonso-Mencía, Marta Castro-Rodríguez, Beatriz Herrero-Pinilla, Juan M. Alonso-Weber, Leocadio Rodríguez-Mañas, and Rodrigo Pérez-Rodríguez. ADELA: a conversational virtual assistant to prevent delirium in hospitalized older persons. The Journal of Supercomputing, 79(15): 17670–17690, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05352-7.

Liu:2023:ANA

[2437] Yuanxuan Liu and Dequan Li. AdaXod: a new adaptive and momental bound algorithm for training deep neural networks. The Journal of Supercomputing, 79(15):17691– 17715, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05338-5.

Karaja:2023:DBT

[2438] Mouna Karaja, Abir Chaabani, Ameni Azzouz, and Lamjed Ben Said. Dynamic bag-of-tasks scheduling problem in a heterogeneous multi-cloud environment: a taxonomy and a new bi-level multi-follower modeling. The Journal of Supercomputing, 79(15): 17716–17753, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05341-w.

Nekouie:2023:CMF

[2439] Nadia Nekouie, Morteza Romoozi, and Mahdi Esmaeili. Correction to: Multimodal feature selection from microarray data based on Dempster–Shafer evidence fusion. The Journal of Supercomputing, 79(15):

17754, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05344-7.  See

[2255].

Raghav:2023:CPT

[2440] Raghav, Nitish Andola, Katyayani Verma, S. Venkatesan, and Shekhar Verma. Correction to: Proactive threshold-proxy re-encryption scheme for secure data sharing on cloud. The Journal of Supercomputing, 79 (15):17755, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05343-8.  See

[2314].

Liu:2023:PCT

[2441] Bo Liu, Rui Chen, Weiwei Lin, Wentai Wu, Jianpeng Lin, and Keqin Li. Publisher correction to: Thermalaware virtual machine placement based on multi-objective optimization. The Journal of Supercomputing, 79(15): 17756–17757, October 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05345-6.  See

[2254].

Liang:2023:PCM

[2442] Xiufang Liang, Yingzheng Zhu, Huajuan Duan, Fuyong Xu, Peiyu Liu, and Ran Lu. Publisher correction to: MISR: a multiple behavior interactive enhanced learning model for social-aware recommendation. The Journal of Supercomputing, 79(15):

17758, October 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05301-4.  See

[2318].

Xiong:2023:UOS

[2443] Yu Xiong, Minghe Yan, Xiang Hu, Chaohui Ren, and Hang Tian. An unsupervised opinion summarization model fused joint attention and dictionary learning. The Journal of Supercomputing, 79(16):17759–17783, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05316-x.

Hong:2023:GEA

[2444] Yan Hong, Yuzhu Wang, Xuanying Zhang, Xiaocong Wang, He Zhang, and Jinrong Jiang. A GPU-enabled acceleration algorithm for the CAM5 cloud microphysics scheme. The Journal of Supercomputing, 79(16):17784– 17809, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05360-7.

Shi:2023:TIM

[2445] Xiayang Shi, Zhenqiang Yu, Xuhui Wang, Yijun Li, and Yufeng Niu. Text-image matching for multi-model machine translation. The Journal of Supercomputing, 79(16):17810– 17823, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05318-9.

Khalil:2023:EFS

[2446] Ali Ahmed Khalil, Mostafa I. El Sayeid, Fatma E. Ibrahim, Ashraf A. M. Khalaf, Entessar Gemeay, Hossam Kasem, Salah Eldeen A. Khamis, Ghada M. El-Banby, Walid El-Shafai, El-Sayed M. El-Rabaie, Adel S. El-Fishawy, Moawad I. Dessouky, Ibrahim M. El-Dokany, Turkey Alotaiby, Saleh A. Alshebeili, and Fathi E. Abd El-Samie. Efficient frameworks for statistical seizure detection and prediction. The Journal of Supercomputing, 79(16):17824– 17858, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-05030-0.

Seyhan:2023:NPA

[2447] Kübra Seyhan and Sedat Akleylek.

A new password-authenticated module learning with rounding-based key exchange protocol: Saber.PAKE. The Journal of Supercomputing, 79(16): 17859–17896, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05251-x.

Riki:2023:DOB

[2448] Samira Riki and Fatemeh Serajeh Hassani. Designing a one-bit coplanar QCA ALU using a novel robust area-efficient three-input majority gate design. The Journal of Supercomputing, 79(16):17897– 17918, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05280-6.

Correia:2023:BSE

[2449] Leonel Feitosa Correia, Jamilson Ramalho Dantas, and Francisco Airton Silva. Blockchain as a service environment: a dependability evaluation. The Journal of Supercomputing, 79(16): 17919–17943, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05342-9.

Tighkhorshid:2023:CDE

[2450] Amirhossein Tighkhorshid, Seyed Mohamad Ali Tousi, and Amirhossein Nikoofard. Car depth estimation within a monocular image using a light CNN. The Journal of Supercomputing, 79(16):17944– 17961, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05359-0.

Cho:2023:ECE

[2451] Joong-Yeon Cho, Pu-Rum Seo, and Hyun-Wook Jin. Exploiting copy engines for intra-node MPI collective communication. The Journal of Supercomputing, 79(16):17962– 17982, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05340-x.

Mazumdar:2023:NBH

[2452] Somnath Mazumdar, Alberto Scionti, Stéphane Zuckerman, and Antoni Portero. NoC-based hardware software co-design framework for dataflow thread management. The Journal of Supercomputing, 79(16):17983– 18020, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05335-8.

Mahdy:2023:AEO

[2453] Araby Mahdy, Abdullah Shaheen, Ragab El-Sehiemy, and Ahmed Ginidi. Artificial ecosystem optimization by means of fitness distance balance model for engineering design optimization. The Journal of Supercomputing, 79(16):18021–18052, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05331-y.

Kabaou:2023:PEO

[2454] Mohamed Ouwais Kabaou, Zoghlami Nesrine, Hamouda Hassen, and Baabou Fatma. Performance evaluation of opportunistic schedulers based on fairness and throughput in newgeneration mobile networks. The Journal of Supercomputing, 79(16):18053– 18088, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05308-x.

Reyes-Delgado:2023:AER

[2455] Paola Yuritzy Reyes-Delgado, Manuel Mora, Fen Wang, and Jorge Marx Gómez. AHP evaluation of rigorous and agile IT service design-building phases-workflows in data centers. The Journal of Supercomputing, 79(16): 18089–18166, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05219-x.

Das:2023:XRA

[2456] Surajit Das, Mahamuda Sultana, Suman Bhattacharya, Diganta Sengupta, and Debashis De. XAI-reduct: accuracy preservation despite dimensionality reduction for heart disease classification using explainable AI. The Journal of Supercomputing, 79(16): 18167–18197, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05356-3.

Qu:2023:SSD

[2457] Hua Qu, Ke Wang, and Jihong Zhao. Survivable SFC deployment method based on federated learning in multi-domain network. The Journal of Supercomputing, 79(16):18198– 18226, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05382-1.

Bakshi:2023:CSO

[2458] Mohana Bakshi, Chandreyee Chowdhury, and Ujjwal Maulik. Cuckoo search optimization-based energy efficient job scheduling approach for IoT-edge environment. The Journal of Supercomputing, 79(16):18227– 18255, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05358-1.

Calheno:2023:ICC

[2459] Rui Calheno, Paulo Carvalho, Solange Rito Lima, Pedro Rangel Henriques, and Mateo Ramos Merino. Improving conformance checking in process modelling: a multiperspective algorithm. The Journal of Supercomputing, 79(16):18256–18292, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05315-y.

Su:2023:STG

[2460] Ziyi Su, Tong Liu, Xiatong Hao, and Xiaojian Hu. Spatial-temporal graph convolutional networks for traffic flow prediction considering multiple traffic parameters. The Journal of Supercomputing, 79(16):18293– 18312, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05383-0.

Xuebin:2023:FAP

[2461] Jin Xuebin, Chen Yewang, Fan Wentao, Zhang Yong, and Du Jixiang. Fast algorithm for parallel solving inversion of large scale small matrices based on GPU. The Journal of Supercomputing, 79(16):18313– 18339, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05336-7.

Bagies:2023:RBD

[2462] Taghreed Bagies, Wei Le, Jeremy Sheaffer, and Ali Jannesari. Reducing branch divergence to speed up parallel execution of unit testing on GPUs. The Journal of Supercomputing, 79(16): 18340–18374, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05375-0.

Dong:2023:LST

[2463] Yumin Dong, Yongfu Zha, Yongjian Zhang, and Xinji Zha. Long- and shortterm collaborative attention networks for sequential recommendation. The Journal of Supercomputing, 79(16): 18375–18393, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05348-3.

Horvath:2023:TRB

[2464] Gábor Horváth, András Mészáros, and Péter Szilágyi. TeleDAL: a regression-based template-less unsupervised method for finding anomalies in log sequences. The Journal of Supercomputing, 79(16):18394– 18416, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05379-w.

fu:2023:KME

[2465] Lifang fu, Huanxin Peng, and Shuai Liu. KG-MFEND: an efficient knowledge graph-based model for multidomain fake news detection. The Journal of Supercomputing, 79(16):

18417–18444, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05381-2.

Chuah:2023:ESM

[2466] Edward Chuah, Arshad Jhumka, and Sai Narasimhamurthy. An empirical study of major page faults for failure diagnosis in cluster systems. The Journal of Supercomputing, 79(16): 18445–18479, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05366-1.

Faqir-Rhazoui:2023:EPP

[2467] Youssef Faqir-Rhazoui and Carlos García. Exploring the performance and portability of the k-means algorithm on SYCL across CPU and GPU architectures. The Journal of Supercomputing, 79(16):18480– 18506, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05373-2.

Mallik:2023:CAF

[2468] Manjarini Mallik and Chandreyee Chowdhury. Characteristic analysis of fingerprint datasets from a pragmatic view of indoor localization using machine learning approaches. The Journal of Supercomputing, 79(16): 18507–18546, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05386-x.

Li:2023:ESC

[2469] Qianyu Li, Wensheng Zhang, and Mengxing Huang. Efficient slot correlation learning network for multidomain dialogue state tracking. The Journal of Supercomputing, 79(16): 18547–18568, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05217-z.

Mohammadzadeh:2023:EAW

[2470] Ali Mohammadzadeh, Mahdi Akbari Zarkesh, Pouria Haji Shahmohamd, Javid Akhavan, and Amit Chhabra. Energy-aware workflow scheduling in fog computing using a hybrid chaotic algorithm. The Journal of Supercomputing, 79(16):18569– 18604, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05330-z.

Zhang:2023:CFA

[2471] Mengxing Zhang, Lin Qi, and Yulong Guo. Construction of feature analysis model for demeanor evidence investigation based on data mining algorithm. The Journal of Supercomputing, 79(16):18605–18626, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05370-5.

Ye:2023:GBA

[2472] Hengzhou Ye, Bochao Feng, and Xinxiao Li. A game-based approach for cloudlet resource pricing for cloudlet federation. The Journal of Supercomputing, 79(16):18627– 18647, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05374-1.

Ran:2023:IPR

[2473] Ruisheng Ran, Qianghui Zeng, Xiaopeng Jiang, and Bin Fang. Isometric projection with reconstruction. The Journal of Supercomputing, 79(16): 18648–18666, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05354-5.

Liu:2023:SIS

[2474] Ying-Ho Liu and Chia-Yu Kuo.

SiMAIM: identifying sockpuppets and puppetmasters on a single forumoriented social media site. The Journal of Supercomputing, 79(16):18667– 18698, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05376-z.

Chang:2023:TAN

[2475] Zhuoqing Chang, Shubo Liu, Run Qiu, Song Song, Zhaohui Cai, and Guoqing Tu. Time-aware neural ordinary differential equations for incomplete time series modeling. The Journal of Supercomputing, 79(16): 18699–18727, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05327-8.

Karaca:2023:NAP

[2476] Osman Volkan Karaca, Kayhan M. Imre, and Ali Ziya Alkar. Network accelerator for parallel discrete event simulations. The Journal of Supercomputing, 79(16):18728– 18747, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05365-2.

Wang:2023:SKS

[2477] Yifei Wang and Mert Pilanci. Sketching the Krylov subspace: faster computation of the entire ridge regularization path. The Journal of Supercomputing, 79(16):18748–18776, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05309-w.

Kalita:2023:LKB

[2478] Dhruba Jyoti Kalita, Vibhav Prakash Singh, and Vinay Kumar. A lightweight knowledge-based PSO for SVM hyper-parameters tuning in a dynamic environment. The Journal of Supercomputing, 79(16):18777– 18799, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05385-y.

kumar:2023:DBA

[2479] Raman kumar and Anuj Jain. Driving behavior analysis and classification by vehicle OBD data using machine learning. The Journal of Supercomputing, 79(16):18800– 18819, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05364-3.

Mu:2023:SLC

[2480] Tong Mu, Qiaochuan Ren, BiLin Shao, Genqing Bian, and Jing Song. A secure and lightweight cloud-centric intelligent medical system based on Internet of Medical Things. The Journal of Supercomputing, 79(16):18820– 18848, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05363-4.

Ghazaan:2023:FCE

[2481] Majid Ilchi Ghazaan, Pedram Ghaderi, and Amirali Rezaeizadeh. A fast convergence EO-based multi-objective optimization algorithm using archive evolution path and its application to engineering design problems. The Journal of Supercomputing, 79(16):18849– 18885, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05362-5.

Zheke:2023:IST

[2482] Yuan Zheke, Niu Jun, Lu Xurong, and Yang Fangmeng. Identify spatiotemporal properties of network traffic by model checking. The Journal of Supercomputing, 79(16):18886– 18909, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05388-9.

Ramezanzad:2023:RTA

[2483] Ali Ramezanzad, Mehran Rezaei, Hooman Nikmehr, and Mahdi Kalbasi. Real-time approximate and combined 2D convolvers for FPGAbased image processing. The Journal of Supercomputing, 79(16):18910– 18946, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05377-y.

Wang:2023:CLY

[2484] Huilin Wang, Huaming Qian, Shuai Feng, and Shuya Yan. CALYOLOv4: lightweight YOLOv4 target detection based on coordinated attention. The Journal of Supercomputing, 79(16): 18947–18969, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05380-3.

Lee:2023:CCO

[2485] Gun Ho Lee. Clustering customer orders in a smart factory using sequential pattern mining. The Journal of Supercomputing, 79(16):18970– 18992, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05351-8.

Gadhamsetty:2023:ASA

[2486] Venkata Datta Adithya Gadhamsetty and Rohit Kumar Das. ASM-SDN: an automated station migration system in cluster-based heterogeneous software-defined network. The Journal of Supercomputing, 79(17):18993– 19018, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05392-z.

Mohammadi:2023:MPA

[2487] Somayeh Mohammadi, Latif PourKarimi, Felix Droop, Ninon De Mecquenem, Ulf Leser, and Knut Reinert. A mathematical programming approach for resource allocation of data analysis workflows on heterogeneous clusters. The Journal of Supercomputing, 79(17): 19019–19048, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05325-w.

Macedo:2023:MAF

[2488] Daniel Macedo, Danilo Santos, Angelo Perkusich, and Dalton Valadares. A mobility-aware federated learning coordination algorithm. The Journal of Supercomputing, 79(17):19049– 19063, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05372-3.

Tang:2023:MRS

[2489] Ji’ao Tang, Xiaojun Zhu, Lu Lin, Chao Dong, and Lei Zhang. Monitoring routing status of UAV networks with NB–IoT. The Journal of Supercomputing, 79(17):19064– 19094, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05411-z.

Wang:2023:NST

[2490] Tao Wang, Xianghong Tang, Jianguang Lu, and Fangjie Liu. A novel spatio-temporal hybrid neural network for remaining useful life prediction. The Journal of Supercomputing, 79(17): 19095–19117, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05404-y.

Chen:2023:BTF

[2491] Fengmei Chen, Bin Zhao, Yilong Gao, and Wenyin Zhang. BTDA: Two-factor dynamic identity authentication scheme for data trading based on alliance chain. The Journal of Supercomputing, 79(17):19118– 19137, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05393-y.

Rawat:2023:NSB

[2492] Arjun Singh Rawat, Maroti Deshmukh, and Maheep Singh. Natural sharebased lightweight (n, n) single secret image sharing scheme using LSB stuffing for medical images. The Journal of Supercomputing, 79(17):19138– 19167, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05396-9.

Absardi:2023:QDS

[2493] Zeinab Nazemi Absardi and Reza Javidan. A QoE-driven SDN traffic management for IoT-enabled surveillance systems using deep learning based on edge cloud computing. The Journal of Supercomputing, 79(17):19168– 19193, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05412-y.

Kavianpour:2023:CBM

[2494] Parisa Kavianpour, Mohammadreza Kavianpour, Ehsan Jahani, and Amin Ramezani. A CNN–BiLSTM model with attention mechanism for earthquake prediction. The Journal of Supercomputing, 79(17):19194– 19226, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05369-y. See cor-

rection [2673].

Qin:2023:NCT

[2495] Yuhua Qin, Haoran Liu, Rongrong Yin, Shiwei Zhao, and Mingru Dong. Non-cooperative target tracking method based on underwater acoustic sensor networks. The Journal of Supercomputing, 79(17):19227– 19253, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05367-0.

Khan:2023:AMG

[2496] Muhammad Hassan Khan, Hiba Azam, and Muhammad Shahid Farid. Automatic multi-gait recognition using pedestrian’s spatiotemporal features. The Journal of Supercomputing, 79(17):19254–19276, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05391-0.

Luo:2023:KSA

[2497] Jiang-Yao Luo and Jian-Hua Yuan. A kernel search algorithm for virtual machine consolidation problem in cloud computing. The Journal of Supercomputing, 79(17):19277– 19296, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05406-w.

Tian:2023:HFG

[2498] Ting Tian, Shumin Zhang, and Yalan Li. Hybrid fault g-good-neighbor conditional diagnosability of star graphs. The Journal of Supercomputing, 79(17):19297–19311, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05368-z.

Yang:2023:ICS

[2499] Chao Yang, Zijing Huang, Bibo Jiang, Menglian Zhu, Aoran Luo, and Jianfeng He. Improved clonal selection algorithm based on the directional update strategy. The Journal of Supercomputing, 79(17):19312– 19331, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05405-x.

Chen:2023:ESA

[2500] Long Chen, Fangyi Xu, Kezhong Jin, and Zhenzhou Tang. Energy-saving access point configurations in WLANs: a swarm intelligent approach. The Journal of Supercomputing, 79(17): 19332–19364, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05402-0.

Zhao:2023:LDP

[2501] Jianzhe Zhao, Chenxi Huang, Wenji Wang, Rulin Xie, Rongrong Dong, and Stan Matwin. Local differentially private federated learning with homomorphic encryption. The Journal of Supercomputing, 79(17):19365– 19395, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05378-x.

Fan:2023:DCB

[2502] Wenguang Fan, Bahman Arasteh, Asgarali Bouyer, and Vahid Majidnezhad. A divide and conquer based development of gray wolf optimizer and its application in data replication problem in distributed systems. The Journal of Supercomputing, 79(17):19396– 19430, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05401-1.

Toprak:2023:ENE

[2503] Ahmet Toprak and Metin Turan.

Enhanced Named Entity Recognition algorithm for financial document verification. The Journal of Supercomputing, 79(17):19431–19451, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05371-4.

Kumari:2023:LIE

[2504] Punam Kumari and Bhaskar Mondal. Lightweight image encryption algorithm using NLFSR and CBC mode. The Journal of Supercomputing, 79(17):19452–19472, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05415-9.

Andalib:2023:ADP

[2505] Azam Andalib and Seyed Morteza Babamir. Anomaly detection of policies in distributed firewalls using data log analysis. The Journal of Supercomputing, 79(17):19473– 19514, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05417-7.

Kar:2023:MHS

[2506] Purbani Kar and Swapan Debbarma. Multilingual hate speech detection sentimental analysis on social media platforms using optimal feature extraction and hybrid diagonal gated recurrent neural network. The Journal of Supercomputing, 79(17):19515– 19546, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05361-6.

Guo:2023:NAM

[2507] Yijie Guo, Lu Lu, and Songxiang Zhu. Novel accelerated methods for convolution neural network with matrix core. The Journal of Supercomputing, 79(17):19547–19573, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05399-6.

Ngueilbaye:2023:DQM

[2508] Alladoumbaye Ngueilbaye, Joshua Zhexue Huang, Mehak Khan, and Hongzhi Wang. Data quality model for assessing public COVID-19 big datasets. The Journal of Supercomputing, 79(17): 19574–19606, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05410-0.

Souza:2023:DES

[2509] Lubnnia Souza, Kádna Camboim, Jean Araujo, Fernanda Alencar, Paulo Maciel, and Joao Ferreira. Dependability evaluation and sensitivity analysis of data center cooling systems. The Journal of Supercomputing, 79(17): 19607–19635, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05419-5.

Wen:2023:DPM

[2510] Yean-Fu Wen and Chia-Pei Wang.

Data privacy mechanisms development and performance evaluation for personal and ubiquitous blockchainbased storage. The Journal of Supercomputing, 79(17):19636–19670, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05425-7.

Feng:2023:FML

[2511] Jinxiang Feng, Jie Xu, Yizhi Deng, and Jun Gao. A Fechner multiscale local descriptor for face recognition. The Journal of Supercomputing, 79(17): 19671–19698, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05421-x.

Zhao:2023:CCB

[2512] Yanpu Zhao, Changsheng Huang, Mengjie Zhang, and Cheng Lv. COLMA: a chaos-based mayfly algorithm with opposition-based learning and Levy flight for numerical optimization and engineering design. The Journal of Supercomputing, 79(17): 19699–19745, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05400-2.

Yakut:2023:NRA

[2513] Selman Yakut, Furkan Öztemiz, and Ali Karci. A new robust approach to solve minimum vertex cover problem: Malatya vertex-cover algorithm. The Journal of Supercomputing, 79(17): 19746–19769, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05397-8.

Megzari:2023:ACS

[2514] Abdelmoujib Megzari, P. V. Pravija Raj, Walid Osamy, and Ahmed M. Khedr. Applications, challenges, and solutions to single- and multiobjective critical node detection problems: a survey. The Journal of Supercomputing, 79(17):19770–19808, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05429-3.

Seilsepour:2023:TSA

[2515] Azam Seilsepour, Reza Ravanmehr, and Ramin Nassiri. Topic sentiment analysis based on deep neural network using document embedding technique. The Journal of Supercomputing, 79(17):19809–19847, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05423-9.

Kirisci:2023:NKP

[2516] Murat Kirisci and Necip Simsek. A novel kernel principal component analysis with application disaster preparedness of hospital: interval-valued Fermatean fuzzy set approach. The Journal of Supercomputing, 79(17):19848– 19878, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05395-w.

Zhou:2023:MEA

[2517] Guanqun Zhou, Shusen Yuan, Hongbo Xing, Youjun Jiang, Pinyong Geng, Yewen Cao, and Xianye Ben. Microexpression action unit recognition based on dynamic image and spatial pyramid. The Journal of Supercomputing, 79(17):19879–19902, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05409-7.

Li:2023:BAS

[2518] Hongzhi Li and Dezhi Han. Blockchainassisted secure message authentication with reputation management for VANETs. The Journal of Supercomputing, 79(17):19903–19933, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05394-x.

Metin:2023:TFT

[2519] Ahmet Metin, Ahmet Kasif, and Cagatay Catal. Temporal fusion transformer-based prediction in aquaponics. The Journal of Supercomputing, 79(17):19934–19958, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05389-8.

Wang:2023:DSU

[2520] Fengxian Wang, Shaozhi Feng, Youmei Pan, Huanlong Zhang, Senlin Bi, and Jiaxiang Zhang. Dynamic spiral updating whale optimization algorithm for solving optimal power flow problem. The Journal of Supercomputing, 79(17):19959–20000, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05427-5.

Wang:2023:CTM

[2521] Fei Wang, Huawei Li, Wanli Yang, Shidong Jin, and Peng Gao. Cell tracking with multifeature fusion. The Journal of Supercomputing, 79(17): 20001–20018, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05384-z.

Nolp:2023:SNC

[2522] Klaus Nölp and Lena Oden. Simplifying non-contiguous data transfer with MPI for Python. The Journal of Supercomputing, 79(17):20019– 20040, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05398-7.

Shen:2023:GNE

[2523] Tao Shen, Tianyu Li, Zhuo Yu, Fenhua Bai, and Chi Zhang. GT-NRSM: efficient and scalable sharding consensus mechanism for consortium blockchain. The Journal of Supercomputing, 79(17):20041– 20075, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05414-w.

ElMenbawy:2023:EEC

[2524] Noha El Menbawy, Hesham A. Ali, Mohamed S. Saraya, Amr M. T. Ali-Eldin, and Mohamed M. Abdelsalam. Energy-efficient computation offloading using hybrid GA with PSO in internet of robotic things environment. The Journal of Supercomputing, 79(17): 20076–20115, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05387-w.

Zhang:2023:AHP

[2525] Jian Zhang, Liang Deng, Ruitian Li, Ming Li, Jie Liu, and Zhe Dai. Achieving high performance and portable parallel GMRES algorithm for compressible flow simulations on unstructured grids. The Journal of Supercomputing, 79(17): 20116–20140, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05430-w.

Singh:2023:DGD

[2526] Manisha Singh, Gaurav Baranwal, and Anil Kumar Tripathi. Decentralized group decision making using blockchain. The Journal of Supercomputing, 79(17):20141–20178, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05426-6.

Singh:2023:ALB

[2527] Tinku Singh, Shivam Gupta, Satakshi, and Manish Kumar. Adaptive load balancing in cluster computing environment. The Journal of Supercomputing, 79(17):20179– 20207, November 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05434-6.

Chen:2023:JRP

[2528] Fengxian Chen. Job runtime prediction of HPC cluster based on PC–Transformer. The Journal of Supercomputing, 79(17):20208–20234, November 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05470-2.

Wen:2023:EMT

[2529] Xiaoqiang Wen, Zhibin Wu, Mengchong Zhou, Jianguo Wang, and Lifeng Wu. Economic mining of thermal power plant based on improved Hadoop-based framework and Spark-based algorithms. The Journal of Supercomputing, 79(18):20235– 20262, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05443-5.

Hamian:2023:NLA

[2530] Melika Hamian, Karim Faez, Soheila Nazari, and Malihe Sabeti. A novel learning approach in deep spiking neural networks with multi-objective optimization algorithms for automatic digit speech recognition. The Journal of Supercomputing, 79(18):20263– 20288, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05420-y.

Sun:2023:IRF

[2531] Haibin Sun and Meng Tian. Improved range-free localization algorithm based on reliable node optimization and enhanced sand cat optimization algorithm. The Journal of Supercomputing, 79(18):20289– 20323, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05436-4.

Bera:2023:VBK

[2532] Sourav Bera, Suryakant Prasad, and Y. Sreenivasa Rao. Verifiable and Boolean keyword searchable attributebased signcryption for electronic medical record storage and retrieval in cloud computing environment. The Journal of Supercomputing, 79(18): 20324–20382, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05416-8.

Tasyurek:2023:NDL

[2533] Murat Tasyurek and Ertugrul Gul. A new deep learning approach based on grayscale conversion and DWT for object detection on adversarial attacked images. The Journal of Supercomputing, 79(18):20383– 20416, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05456-0.

Li:2023:GEM

[2534] Bing Li, Peng Yang, Zhongjian Hu, Yuankang Sun, and Meng Yi. Graphenhanced multi-answer summarization under question-driven guidance. The Journal of Supercomputing, 79(18): 20417–20444, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05457-z.

Zhang:2023:NAD

[2535] Hao Zhang, Zude Xiao, Jason Gu, and Yanhua Liu. A network anomaly detection algorithm based on semisupervised learning and adaptive multiclass balancing. The Journal of Supercomputing, 79(18):20445– 20480, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05474-y.

Ding:2023:DIN

[2536] Junyuan Ding, Xiaoliang Chen, Peng Lu, Zaiyan Yang, Xianyong Li, and Yajun Du. DialogueINAB: an interaction neural network based on attitudes and behaviors of interlocutors for dialogue emotion recognition. The Journal of Supercomputing, 79(18): 20481–20514, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05439-1.

Deepthi:2023:DNA

[2537] M. Deepthi, G. N. V. R. Vikram, and P. Venkatappareddy. Development of a novel activation function based on Chebyshev polynomials: an aid for classification and denoising of images. The Journal of Supercomputing, 79(18):20515– 20531, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05466-y.

Jeng:2023:IQC

[2538] Mingyoung Jeng, SM Ishraq Ul Islam, David Levy, Andrew Riachi, Manu Chaudhary, Md. Alvir Islam Nobel, Dylan Kneidel, Vinayak Jha, Jack Bauer, Anshul Maurya, Naveed Mahmud, and Esam El-Araby. Improving quantum-to-classical data decoding using optimized quantum wavelet transform. The Journal of Supercomputing, 79(18):20532– 20561, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05433-7.

Benrazek:2023:TBI

[2539] Ala-Eddine Benrazek, Zineddine Kouahla, Brahim Farou, Hamid Seridi, Imane Allele, and Mohamed Amine Ferrag. Tree-based indexing technique for efficient and real-time label retrieval in the object tracking system. The Journal of Supercomputing, 79(18): 20562–20599, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05478-8.

Liu:2023:IDM

[2540] Yaodi Liu, Kun Zhang, and Zhendong Wang. Intrusion detection of manifold regularized broad learning system based on LU decomposition. The Journal of Supercomputing, 79(18): 20600–20648, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05403-z.

Munoz-Montoro:2023:EPK

[2541] Antonio J. Muñoz-Montoro, Julio J.

Carabias-Orti, Daniele Salvati, and Raquel Cortina. Efficient parallel kernel based on Cholesky decomposition to accelerate multichannel nonnegative matrix factorization. The Journal of Supercomputing, 79(18): 20649–20664, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05471-1.

Singh:2023:ADF

[2542] Ashutosh Kumar Singh and Pooja Rani. An AI-driven fault-tolerant aggregation model for smart grid. The Journal of Supercomputing, 79(18): 20665–20683, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05461-3.

Wu:2023:MHA

[2543] Huafeng Wu, Yuxuan Zhang, Linian Liang, Xiaojun Mei, Dezhi Han, Bing Han, Tien-Hsiung Weng, and Kuan-Ching Li. Multi-head attention-based model for reconstructing continuous missing time series data. The Journal of Supercomputing, 79(18):20684– 20711, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05465-z. See cor-

rection [2622].

Sharma:2023:TPY

[2544] Sonia Sharma and Rajendra Kumar Bharti. Two phased Yarn scheduling with novel binary hypergraphed PSO. The Journal of Supercomputing, 79(18):20712–20746, December 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05310-3.

Yin:2023:EPT

[2545] Junqi Yin, Sajal Dash, John Gounley, Feiyi Wang, and Georgia Tourassi. Evaluation of pre-training large language models on leadership-class supercomputers. The Journal of Supercomputing, 79(18):20747–20768, December 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05479-7.

Alseid:2023:MIM

[2546] Marya Alseid, Ali A. El-Moursy, Oruba Alfawaz, and Ahmed M. Khedr. MSSAMTO-IoV: modified sparrow search algorithm for multihop task offloading for IoV. The Journal of Supercomputing, 79(18): 20769–20789, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05446-2.

Roy:2023:DSS

[2547] Aditi Roy, J. Kokila, N. Ramasubramanian, and B. Shameedha Begum. Device-specific security challenges and solution in IoT edge computing: a review. The Journal of Supercomputing, 79(18):20790– 20825, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05450-6.

Liao:2023:PRF

[2548] Jing Liao, Xiande Su, Lei Jiang, Kuan-Ching Li, Tien-Hsiung Weng, and Subhash Bhalla. Performance of representation fusion model for entity and relationship extraction within unstructured text. The Journal of Supercomputing, 79(18):20826– 20844, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05442-6.

Chowdhuri:2023:NPE

[2549] Rajib Chowdhuri and Mrinal Kanti Deb Barma. Node position estimation based on optimal clustering and detection of coverage hole in wireless sensor networks using hybrid deep reinforcement learning. The Journal of Supercomputing, 79(18):20845– 20877, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05494-8.

Feng:2023:DFM

[2550] Shuai Feng, Huaming Qian, and Huilin Wang. A deep feature matching pipeline with triple search strategy. The Journal of Supercomputing, 79(18):20878–20898, December 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05418-6.

Morabbi:2023:IGD

[2551] Sajedeh Morabbi, Hadi Soltanizadeh, Saeed Mozaffari, and Mohammad Javad Fadaeieslam. Improving generalization in deep neural network using knowledge transformation based on Fisher criterion. The Journal of Supercomputing, 79(18):20899– 20922, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05448-0.

Youssef:2023:SGS

[2552] Taki Youssef, Elmoukhtar Zemmouri, and Anas Bouzid. STM-GCN: a spatiotemporal multi-graph convolutional network for pedestrian trajectory prediction. The Journal of Supercomputing, 79(18):20923–20937, December 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05467-x.

Reddy:2023:OAB

[2553] Posham Bhargava Reddy and Chapram Sudhakar. An osmotic approachbased dynamic deadline-aware task offloading in edge-fog-cloud computing environment. The Journal of Supercomputing, 79(18):20938– 20960, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05440-8.

Sivasubramani:2023:GBA

[2554] Santhosh Sivasubramani, Sanghamitra Debroy, Swati Ghosh Acharyya, and Amit Acharyya. Graphene-based area efficient power planning architecture design methodology for nanomagnetic logic implementation. The Journal of Supercomputing, 79(18): 20961–20983, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05449-z.

Tang:2023:TFC

[2555] Mengmeng Tang, Yuanbo Guo, Qingchun Bai, and Han Zhang. Trigger-free cybersecurity event detection based on contrastive learning. The Journal of Supercomputing, 79(18): 20984–21007, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05454-2.

Sun:2023:SAS

[2556] Kai Sun, HuaJie Jiang, Yongli Hu, and BaoCai Yin. Substructureaware subgraph reasoning for inductive relation prediction. The Journal of Supercomputing, 79(18):21008– 21027, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05493-9.

Sun:2023:TTC

[2557] Dawei Sun, Minghui Wu, Zhihong Yang, Atul Sajjanhar, and Rajkumar Buyya. A two-tier coordinated load balancing strategy over skewed data streams. The Journal of Supercomputing, 79(18):21028– 21056, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05473-z.

Saini:2023:FIF

[2558] Kanika Saini, Sheetal Kalra, and Sandeep K. Sood. Fog-inspired framework for emergency rescue operations in post-disaster scenario. The Journal of Supercomputing, 79(18):21057– 21088, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05475-x.

Du:2023:CCS

[2559] Ming Du, Wanting Ma, Yuting Tan, and Junfeng Zhou. Continuous community search with attribute constraints in temporal graphs. The Journal of Supercomputing, 79(18): 21089–21115, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05451-5.

Sagharichian:2023:IDP

[2560] Masoud Sagharichian and Morteza Alipour Langouri. iPartition: a distributed partitioning algorithm for block-centric graph processing systems. The Journal of Supercomputing, 79(18):21116– 21143, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05492-w.

Ali:2023:CBR

[2561] Muhammad Ali, Asad Waqar Malik, and Anis Ur Rahman. Clusteringbased re-routing framework for network traffic congestion avoidance on urban vehicular roads. The Journal of Supercomputing, 79(18):21144– 21165, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05455-1.

Kudkelwar:2023:AMA

[2562] Sagar B. Kudkelwar, Bam Bahadur Sinha, and Sravan Kumar Gunturi. An Archimedes metaheuristic algorithm based optimum relay coordination in microgrid and combined overhead/cable distribution network. The Journal of Supercomputing, 79(18): 21166–21184, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05486-8.

Ahmadluei:2023:CNN

[2563] Saeed Ahmadluei, Karim Faez, and Behrooz Masoumi. Convolutional neural network pruning based on misclassification cost. The Journal of Supercomputing, 79(18):21185– 21234, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05487-7.

Joshi:2023:SAL

[2564] Nandkishor Joshi and S. C. Sharma. Switching algorithm in listen-and-talk-based MAC protocols for fullduplex cognitive radio networks with type 2 fuzzy cooperative spectrum sensing. The Journal of Supercomputing, 79(18):21235–21264, December 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05477-9.

Barhoush:2023:IDS

[2565] Malek Barhoush, Bilal H. Abed-alguni, and Nour Elhuda A. Al-qudah. Improved discrete salp swarm algorithm using exploration and exploitation techniques for feature selection in intrusion detection systems. The Journal of Supercomputing, 79(18):21265– 21309, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05444-4.

Ghunaim:2023:SKQ

[2566] Tasneem Ghunaim, Ibrahim Kamel, and Zaher Al Aghbari. Secure kNN query of outsourced spatial data using two-cloud architecture. The Journal of Supercomputing, 79(18):21310– 21345, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05495-7.

Liu:2023:SEG

[2567] Zhuohua Liu, Bin Yang, Jingrui An, and Caijuan Huang. Similarity evaluation of graphic design based on deep visual saliency features. The Journal of Supercomputing, 79(18): 21346–21367, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05468-w.

Chraibi:2023:NDM

[2568] Amine Chraibi, Said Ben Alla, Abdellah Touhafi, and Abdellah Ezzati. A novel dynamic multi-objective task scheduling optimization based on Dueling DQN and PER. The Journal of Supercomputing, 79(18):21368– 21423, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05489-5.

Gomes:2023:NBS

[2569] Carlos Gomes, Meuse Nogueira de O. Junior, Bruno Nogueira, Paulo Maciel, and Eduardo Tavares. NoSQLbased storage systems: influence of consistency on performance, availability and energy consumption. The Journal of Supercomputing, 79(18): 21424–21448, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05488-6.

Sheikhi:2023:ADC

[2570] Farnaz Sheikhi, Aliakbar Taghdiri, Danial Moradisabzevar, Hanieh Rezakhani, Hasti Daneshkia, and Mobina Goodarzi. Automatic detection of COVID-19 and pneumonia from chest X-ray images using texture features. The Journal of Supercomputing, 79(18):21449–21473, December 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05452-4.

Kumar:2023:PLA

[2571] E. Praveen Kumar and S. Priyanka. A password less authentication protocol for multi-server environment using physical unclonable function. The Journal of Supercomputing, 79(18): 21474–21506, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05437-3.

Ozkan:2023:NPG

[2572] Kürsad Özkan, Ahmet Mert, and Serkan Özdemir. A new proposed GLCM texture feature: modified Rényi Deng entropy. The Journal of Supercomputing, 79(18):21507– 21527, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05627-z.

Singh:2023:CFT

[2573] Raushan Kumar Singh and Mukesh Kumar. Correction to: Future trends of path planning framework considering accident attributes for smart cities. The Journal of Supercomputing, 79 (18):21528, December 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05431-9.  See

[2411].

Zhou:2024:LIP

[2574] Zhiyong Zhou, Yuanning Liu, Xiaodong Zhu, Shuai Liu, Shaoqiang Zhang, Yuanfeng Li, and Zhen Liu. Lifelong iris presentation attack detection without forgetting. The Journal of Supercomputing, 80(1):1– 19, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05445-3.

Xia:2024:ASR

[2575] Dahai Xia, Xinyun Wu, Meng Yan, and Caiquan Xiong. An adaptive stochastic ranking-based tournament selection method for differential evolution. The Journal of Supercomputing, 80(1):20–49, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05390-1.

Javeed:2024:CEF

[2576] Khalid Javeed, Ali El-Moursy, and David Gregg. E^{2}CSM: efficient FPGA implementation of elliptic curve scalar multiplication over generic prime field GF (p). The Journal of Supercomputing, 80(1):50– 74, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05428-4.

Kumar:2024:GAC

[2577] Chandan Kumar, Sudhanshu Kumar Jha, Dilip Kumar Yadav, Shiv Prakash, and Mukesh Prasad. A generalized approach to construct node probability table for Bayesian belief network using fuzzy logic. The Journal of Supercomputing, 80(1):75– 97, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05458-y.

Katal:2024:EOC

[2578] Avita Katal, Tanupriya Choudhury, and Susheela Dahiya. Energy optimized container placement for cloud data centers: a metaheuristic approach. The Journal of Supercomputing, 80(1):98– 140, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05462-2.

Sameh:2024:ECM

[2579] Sally Mohamed Sameh, Hossam El-Din Moustafa, Ehab H. AbdelHay, and Mohamed Maher Ata. An effective chaotic maps image encryption based on metaheuristic optimizers. The Journal of Supercomputing, 80(1):141–201, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05413-x.

Wu:2024:NFC

[2580] Huafeng Wu, Feng Wang, Xiaojun Mei, Linian Liang, Bing Han, Dezhi Han, Tien-Hsiung Weng, and Kuan-Ching Li. A novel fuzzy control path planning algorithm for intelligent ship based on scale factors. The Journal of Supercomputing, 80 (1):202–225, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05438-2.

Abdulhammed:2024:RIS

[2581] Omar Younis Abdulhammed. A robust image steganography based on a novel technique by using improved DNA and modified chaotic approach. The Journal of Supercomputing, 80 (1):226–248, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05459-x.

Dong:2024:WBC

[2582] Na Dong, Qingyue Feng, Jianfang Chang, and Xiaoming Mai. White blood cell classification based on a novel ensemble convolutional neural network framework. The Journal of Supercomputing, 80(1):249– 270, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05490-y.

Rajalakshmi:2024:NDL

[2583] R. Rajalakshmi, P. Sivakumar, L. Krishna Kumari, and M. Chengathir Selvi. A novel deep learning model for diabetes melliuts prediction in IoTbased healthcare environment with effective feature selection mechanism. The Journal of Supercomputing, 80 (1):271–291, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05496-6.

Emanuel:2024:EAF

[2584] Rebecca H. K. Emanuel, Paul D. Docherty, Helen Lunt, and Knut Möller. The effect of activation functions on accuracy, convergence speed, and misclassification confidence in CNN text classification: a comprehensive exploration. The Journal of Supercomputing, 80(1):292– 312, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05441-7.

Lai:2024:CEC

[2585] Yonghao Lai and Xiaohui Hua. Component edge connectivity and extra edge connectivity of alternating group networks. The Journal of Supercomputing, 80(1):313–330, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05464-0.

Yang:2024:PAD

[2586] Jin Yang, Wangdong Yang, Ruixuan Qi, Qinyun Tsai, Shengle Lin, Fengkun Dong, Kenli Li, and Keqin Li. Parallel algorithm design and optimization of geodynamic numerical simulation application on the Tianhe new-generation high-performance computer. The Journal of Supercomputing, 80(1):331–362, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05469-9.

Chang:2024:JJC

[2587] Chih-Yung Chang, Syu-Jhih Jhang, Shih-Jung Wu, and Diptendu Sinha Roy. JCF: joint coarse- and finegrained similarity comparison for plagiarism detection based on NLP. The Journal of Supercomputing, 80(1): 363–394, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05472-0.

Navimipour:2024:NSA

[2588] Nima Jafari Navimipour, Seyed-Sajad Ahmadpour, and Senay Yalcin. A nano-scale arithmetic and logic unit using a reversible logic and quantumdots. The Journal of Supercomputing, 80(1):395–412, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05491-x.

Ran:2024:PLD

[2589] Ruisheng Ran, Ting Wang, Zheng Li, and Bin Fang. Polynomial linear discriminant analysis. The Journal of Supercomputing, 80(1):413– 434, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05485-9.

Park:2024:APJ

[2590] Ju-Won Park, Xin Huang, and Chul-Ho Lee. Analyzing and predicting job failures from HPC system log. The Journal of Supercomputing, 80 (1):435–462, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05482-y.

Ahishakiye:2024:DGC

[2591] Emmanuel Ahishakiye, Waweru Mwangi, Petronilla Muriithi, Fredrick Kanobe, Godliver Owomugisha, Danison Taremwa, and Lenard Nkalubo. Deep Gaussian convolutional neural network model in classification of cassava diseases using spectral data. The Journal of Supercomputing, 80(1):463– 485, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05498-4.

Ustubioglu:2024:AFD

[2592] Beste Ustubioglu, Gul Tahaoglu, Guzin Ulutas, Arda Ustubioglu, and Muhammed Kilic. Audio forgery detection and localization with superresolution spectrogram and keypointbased clustering approach. The Journal of Supercomputing, 80(1):486– 518, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05504-9.

Tian:2024:PPT

[2593] Tian Tian, Anshi Wang, Xiuting Yang, Dunwei Gong, Tie Hou, and Xiangjuan Yao. Parallel program testing based on critical communication and branch transformation. The Journal of Supercomputing, 80 (1):519–548, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05460-4.

Piontek:2024:CEA

[2594] Tobias Piontek, Kawsar Haghshenas, and Marco Aiello. Carbon emissionaware job scheduling for Kubernetes deployments. The Journal of Supercomputing, 80(1):549– 569, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05506-7.

Salami:2024:SFS

[2595] Yashar Salami, Vahid Khajehvand, and Esmaeil Zeinali. SOS-FCI: a secure offloading scheme in fog-cloud-based IoT. The Journal of Supercomputing, 80(1):570–600, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05499-3.

Jalilian:2024:TVM

[2596] Hamed Jalilian. Total variation method based on modified Barzilai– Borwein algorithm to noise reduction in MRI images. The Journal of Supercomputing, 80(1):601– 619, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05500-z.

Li:2024:USB

[2597] Qun Li, Bei Tang, Jianxin Li, and Siguang Chen. User satisfactionbased energy-saving computation offloading in fog computing networks. The Journal of Supercomputing, 80 (1):620–641, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05484-w.

Gao:2024:HGT

[2598] Rong Gao, Wei He, Lingyu Yan, Donghua Liu, Yonghong Yu, and Zhiwei Ye. Hybrid graph transformer networks for multivariate time series anomaly detection. The Journal of Supercomputing, 80(1):642– 669, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05503-w.

Vardi:2024:HMA

[2599] Fatemeh Vardi and Alireza Mahjoub. A hot-module-aware mapping approach in network-on-chip. The Journal of Supercomputing, 80(1):670– 702, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05424-8.

Sahoo:2024:PLB

[2600] Sujit Sangram Sahoo and Vijay Kumar Chaurasiya. Proof of location based delivery system using multi-party virtual state channel: a blockchain model. The Journal of Supercomputing, 80 (1):703–733, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05510-x.

Chiang:2024:DWB

[2601] Mao-Lun Chiang, Hui-Ching Hsieh, Tzu-Ling Lin, Tsui-Ping Chang, and Hong-Wei Chen. Dynamic weightbased connectivity recovery in wireless sensor and actor networks. The Journal of Supercomputing, 80(1): 734–760, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05518-3.

Dalirinia:2024:LEO

[2602] Elham Dalirinia, Mehrdad Jalali, Mahdi Yaghoobi, and Hamid Tabatabaee. Lotus effect optimization algorithm (LEA): a lotus nature-inspired algorithm for engineering design optimization. The Journal of Supercomputing, 80(1):761–799, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05513-8.

Bahrani:2024:NIK

[2603] Payam Bahrani, Behrouz Minaei-Bidgoli, Hamid Parvin, Mitra Mirzarezaee, and Ahmad Keshavarz. A new improved KNN-based recommender system. The Journal of Supercomputing, 80(1):800–834, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05447-1.

Chauhan:2024:SAD

[2604] Dikshit Chauhan and Anupam Yadav. Stability and agent dynamics of artificial electric field algorithm. The Journal of Supercomputing, 80 (1):835–864, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05502-x.

Singh:2024:TTF

[2605] Namrata Singh and Ayan Kumar Das. TFAS: two factor authentication scheme for blockchain enabled IoMT using PUF and fuzzy extractor. The Journal of Supercomputing, 80(1):865–914, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05507-6.

Mahmoodabadi:2024:AAV

[2606] Zahra Mahmoodabadi and Mostafa Nouri-Baygi. An approximation algorithm for virtual machine placement in cloud data centers. The Journal of Supercomputing, 80(1): 915–941, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05505-8.

Kumar:2024:UUP

[2607] Rakesh Kumar, Sunil K. Singh, and D. K. Lobiyal. UPSRVNet: Ultralightweight, privacy preserved, and secure RFID-based authentication protocol for VIoT networks. The Journal of Supercomputing, 80(1):942– 969, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05463-1.

Cheng:2024:PUN

[2608] Sheng-Tzong Cheng, Gwo-Jiun Horng, Chih-Wei Hsu, and Z.-Yu Su. Per-user network access control kernel module with secure multifactor authentication. The Journal of Supercomputing, 80 (1):970–1008, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05480-0.

Khalifavi:2024:FLR

[2609] Maedeh Khalifavi, Zahra Shirmohammadi, and Sahar Kianian. FASR-LED: reducing energy consumption in wireless body area networks by an efficient smart method. The Journal of Supercomputing, 80(1):1009– 1036, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05481-z.

Shah:2024:CNL

[2610] Tariq Shah and Tanveer ul Haq. Construction of 24-by-24 nonlinear layer for symmetric algorithm and its application to data encryption in parallel with DNA transform. The Journal of Supercomputing, 80(1):1037– 1058, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05512-9.

Bozdal:2024:CAD

[2611] Mehmet Bozdal, Kadir Ileri, and Ali Ozkahraman. Comparative analysis of dimensionality reduction techniques for cybersecurity in the SWaT dataset. The Journal of Supercomputing, 80(1): 1059–1079, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05511-w.

Indushree:2024:NBB

[2612] M. Indushree and Manish Raj. A novel blockchain-based authentication scheme for telecare medical information system. The Journal of Supercomputing, 80(1):1080– 1108, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05526-3.

Azizi:2024:SFP

[2613] Sadoon Azizi, Pedram Farzin, Mohammad Shojafar, and Omer Rana. A scalable and flexible platform for service placement in multi-fog and multi-cloud environments. The Journal of Supercomputing, 80(1):1109– 1136, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05520-9.

Liu:2024:RLB

[2614] Tundong Liu, Fengqiang Gao, Xin Chen, Guifang Shao, and Chenhan Wang. Research on load-balancing scheduling optimization for wafer surface defect detection. The Journal of Supercomputing, 80(1):1137– 1159, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05521-8.

Latrech:2024:CDH

[2615] Jihene Latrech, Zahra Kodia, and Nadia Ben Azzouna. CoDFi-DL: a hybrid recommender system combining enhanced collaborative and demographic filtering based on deep learning. The Journal of Supercomputing, 80 (1):1160–1182, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05519-2.

Zhang:2024:HSA

[2616] Zuoyan Zhang, Jinchen Xu, Jiangwei Hao, Yang Qu, Haotian He, and Bei Zhou. Hierarchical search algorithm for error detection in floatingpoint arithmetic expressions. The Journal of Supercomputing, 80(1): 1183–1205, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05523-6.

Baouya:2024:DWR

[2617] Abdelhakim Baouya, Salim Chehida, Saddek Bensalem, Levent Gürgen, Richard Nicholson, Miquel Cantero, Mario Diaznava, and Enrico Ferrera. Deploying warehouse robots with confidence: the BRAIN–IoT framework’s functional assurance. The Journal of Supercomputing, 80(1): 1206–1237, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05483-x.

Sharma:2024:COC

[2618] Sumit Sharma and Sarika Jain.

CovidO: an ontology for COVID-19 metadata. The Journal of Supercomputing, 80(1):1238–1267, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05509-
4.

Qi:2024:EGB

[2619] Sibo Qi, Juan Chen, Peng Chen, Peian Wen, Xianhua Niu, and Lei Xu. An efficient GAN-based predictive framework for multivariate time series anomaly prediction in cloud data centers. The Journal of Supercomputing, 80(1): 1268–1293, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05534-3.

Li:2024:PPI

[2620] Xiaomei Li, Ye Lu, Xiaoyu Zhao, Xiong Deng, and Zhijiang Xie. Path planning for intelligent vehicles based on improved d* Lite. The Journal of Supercomputing, 80(1):1294– 1330, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05528-1.

Pabitha:2024:DMR

[2621] C. Pabitha and B. Vanathi. Dense mesh RCNN: assessment of human skin burn and burn depth severity. The Journal of Supercomputing, 80(1): 1331–1362, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05660-y.

Wu:2024:CMH

[2622] Huafeng Wu, Yuxuan Zhang, Linian Liang, Xiaojun Mei, Dezhi Han, Bing Han, Tien˘2011Hsiung Weng, and Kuan˘2011Ching Li. Correction to: Multi-head attention-based model for reconstructing continuous missing time series data. The Journal of Supercomputing, 80(1):1363– 1364, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05515-6.  See

[2543].

Tang:2024:CPB

[2623] Minghua Tang, Enrico Russo, and Maurizio Palesi. Correction to: The position-based compression techniques for DNN model. The Journal of Supercomputing, 80(1): 1365, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05514-7.  See

[2429].

Buttar:2024:SSS

[2624] Preetpal Kaur Buttar and Manoj Kumar Sachan. Semantic segmentation of satellite images for crop type identification in smallholder farms. The Journal of Supercomputing, 80(2): 1367–1395, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05533-4.

Zhonghua:2024:SCA

[2625] Chen Zhonghua, S. B. Goyal, and Anand Singh Rajawat. Smart contracts attribute-based access control model for security & privacy of IoT system using blockchain and edge computing. The Journal of Supercomputing, 80(2): 1396–1425, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05517-4.

Gao:2024:FNS

[2626] Xiaoxin Gao, Fazhi He, Songwei Zhang, Jinkun Luo, and Bo Fan. A fast nondominated sorting-based MOEA with convergence and diversity adjusted adaptively. The Journal of Supercomputing, 80(2):1426– 1463, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05516-5.

Rashidi:2024:FTE

[2627] Bahram Rashidi. Fault-tolerant and error-correcting 4-bit S-boxes for cryptography applications with multiple errors detection. The Journal of Supercomputing, 80(2):1464– 1490, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05530-7.

Alavizadeh:2024:EDS

[2628] Alavieh Sadat Alavizadeh, Seyed Hossein Erfani, Meghdad Mirabi, and Amir Sahafi. An efficient distributed and secure algorithm for transaction confirmation in IOTA using cloud computing. The Journal of Supercomputing, 80 (2):1491–1521, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05525-4.

Uppada:2024:NEA

[2629] Santosh Kumar Uppada, B. S. Ashwin, and B. Sivaselvan. A novel evolutionary approach-based multimodal model to detect fake news in OSNs using text and metadata. The Journal of Supercomputing, 80(2): 1522–1553, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05531-6.

Amiri:2024:DIC

[2630] Melika Amiri, Massoud Dousti, and Majid Mohammadi. Design and implementation of carry-save adder using quantum-dot cellular automata. The Journal of Supercomputing, 80(2): 1554–1567, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05532-5.

Yeasin:2024:OOB

[2631] Md. Yeasin and Ranjit Kumar Paul.

OptiSembleForecasting: optimizationbased ensemble forecasting using MCS algorithm and PCA-based error index. The Journal of Supercomputing, 80(2): 1568–1597, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05542-3.

Carrion:2024:RSO

[2632] Carmen Carrión. Research streams and open challenges in the metaverse. The Journal of Supercomputing, 80(2): 1598–1639, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05544-1.

Tian:2024:CPA

[2633] Zhaoxia Tian, Mingzu Zhang, and Xing Feng. Concentration phenomenon about h-extra edge-connectivity of the n-th cartesian product of complete graph K_{4} with large-scale faulty links. The Journal of Supercomputing, 80(2): 1640–1667, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05350-9.

Rostampour:2024:UPE

[2634] Samad Rostampour, Nasour Bagheri, Behnam Ghavami, Ygal Bendavid, Saru Kumari, Honorio Martin, and Carmen Camara. Using a privacyenhanced authentication process to secure IoT-based smart grid infrastructures. The Journal of Supercomputing, 80(2):1668–1693, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-023-05535-
                                        2.

Yang:2024:TNA

[2635] Yi Yang, Jiaxuan Wei, Zhixuan Yu, and Ruisheng Zhang. A trustworthy neural architecture search framework for pneumonia image classification utilizing blockchain technology. The Journal of Supercomputing, 80(2): 1694–1727, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05541-4. See cor-

rection [2869].

Qtaish:2024:OKM

[2636] Amjad Qtaish, Malik Braik, Dheeb Albashish, Mohammad T. Alshammari, Abdulrahman Alreshidi, and Eissa Jaber Alreshidi. Optimization of K-means clustering method using hybrid capuchin search algorithm. The Journal of Supercomputing, 80(2): 1728–1787, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05540-5.

Sood:2024:SDS

[2637] Shaurya Sood and Neminath Hubballi. SlowTrack: detecting slow rate Denial of Service attacks against HTTP with behavioral parameters. The Journal of Supercomputing, 80(2): 1788–1817, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05453-3.

Mewada:2024:NSG

[2638] Arvind Mewada and Rupesh Kumar Dewang. NRWalk2Vec-HIN: spammer group detection based on heterogeneous information network embedding over social media. The Journal of Supercomputing, 80(2):1818– 1851, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05537-0.

Magdy:2024:DMS

[2639] Mirna Magdy, Mohamed Grida, and Gawaher Hussein. Disruption mitigation in the semiconductors supply chain by using public blockchains. The Journal of Supercomputing, 80(2): 1852–1906, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05543-2.

Zhang:2024:CED

[2640] Huanwen Zhang, Yan Wang, Jianxi Fan, Yuejuan Han, and Baolei Cheng. Constructing edge-disjoint spanning trees in several cube-based networks with applications to edge faulttolerant communication. The Journal of Supercomputing, 80(2):1907– 1934, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05546-z.

Aikyn:2024:EFE

[2641] Nartay Aikyn, Ardan Zhanegizov, Temirlan Aidarov, Dinh-Mao Bui, and Nguyen Anh Tu. Efficient facial expression recognition framework based on edge computing. The Journal of Supercomputing, 80(2):1935– 1972, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05548-x.

Paikaray:2024:OCN

[2642] Divya Paikaray, Ashok Kumar Mehta, and Danish Ali Khan. Optimized convolutional neural network for the classification of lung cancer. The Journal of Supercomputing, 80(2): 1973–1989, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05550-3.

Haraty:2024:DMS

[2643] Ramzi A. Haraty and Ali Assaf.

DG-means: a superior greedy algorithm for clustering distributed data. The Journal of Supercomputing, 80(2): 1990–2024, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05508-5.

Han:2024:YSS

[2644] Yanjiang Han, Fengping Wang, Wei Wang, Xiangyu Li, and Jianyang Zhang. YOLO-SG: Small traffic signs detection method in complex scene. The Journal of Supercomputing, 80(2): 2025–2046, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05547-y.

Wang:2024:IEA

[2645] Yu-Ju Wang, Liang-Hong Wu, and Liang-Chuan Wu. An integrative extraction approach for index-tracking portfolio construction and forecasting under a deep learning framework. The Journal of Supercomputing, 80(2): 2047–2066, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05538-z.

Esfahani:2024:BBE

[2646] Maryam Nasr Esfahani, Behrouz Shahgholi Ghahfarokhi, and Shahram Etemadi Borujeni. Blockchain-based end-to-end privacy-preserving scheme for IoTbased healthcare systems. The Journal of Supercomputing, 80(2):2067– 2127, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05522-7.

Seifpour:2024:SMO

[2647] Massoud Seifpour, Seyyed Amir Asghari, and Mostafa Ghobaei-Arani. A stochastic multi-objective optimization method for railways scheduling: a NSGA–II-based hybrid approach. The Journal of Supercomputing, 80(2): 2128–2163, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05529-0.

Mathivanan:2024:CIE

[2648] P. Mathivanan and Ponnambalam Maran. Color image encryption based on novel kolam scrambling and modified 2D logistic cascade map (2D LCM). The Journal of Supercomputing, 80(2):2164– 2195, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05539-y. See cor-

rection [2722].

Li:2024:DGN

[2649] Xiaohong Li, Qixuan Peng, Ruihong Li, and Huifang Ma. Dual graph neural network for overlapping community detection. The Journal of Supercomputing, 80(2):2196– 2222, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05435-5.

Pramanik:2024:MCC

[2650] Pijush Kanti Dutta Pramanik, Saurabh Pal, and Prasenjit Choudhury. Mobile crowd computing: potential, architecture, requirements, challenges, and applications. The Journal of Supercomputing, 80(2):2223–2318, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05545-
0.

Liu:2024:ERL

[2651] Haiying Liu, ZhiHao Li, Kuihua Huang, Rui Wang, Guangquan Cheng, and Tiexiang Li. Evolutionary reinforcement learning algorithm for largescale multi-agent cooperation and confrontation applications. The Journal of Supercomputing, 80(2):2319– 2346, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05551-2.

S:2024:UAD

[2652] Saumiya S. and S. Wilfred Franklin.

Unified automated deep learning framework for segmentation and classification of liver tumors. The Journal of Supercomputing, 80(2):2347– 2380, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05524-5.

Guo:2024:OOD

[2653] Hengliang Guo, Long Zhang, Yi Zhang, Jianan Li, Xiaoyue Xu, Lu Liu, Kuangsheng Cai, Dan Wu, Shuxin Yang, Lingbo Kong, and Xu Gao. OpenMP offloading data transfer optimization for DCUs. The Journal of Supercomputing, 80(2):2381– 2402, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05422-w.

Nassar:2024:RMB

[2654] Shaimaa E. Nassar, Ibrahim Yasser, Hanan M. Amer, and Mohamed A. Mohamed. A robust MRI-based brain tumor classification via a hybrid deep learning technique. The Journal of Supercomputing, 80(2): 2403–2427, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05549-w.

Sivanandam:2024:NLG

[2655] Chakaravarthi Sivanandam, Vaishnnave Mala Perumal, and Jagadeesh Mohan. A novel light GBM-optimized long short-term memory for enhancing quality and security in web service recommendation system. The Journal of Supercomputing, 80(2):

2428–2460, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05552-1.

Patel:2024:ALB

[2656] Usha Patel and Vibha Patel. Active learning-based hyperspectral image classification: a reinforcement learning approach. The Journal of Supercomputing, 80(2):2461– 2486, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05568-7.

Xia:2024:MML

[2657] Chenxing Xia, Difeng Chen, Xiuju Gao, Bin Ge, Kuan-Ching Li, Xianjin Fang, Yan Zhang, and Ke Yang. MFCINet: multi-level feature and context information fusion network for RGB-D salient object detection. The Journal of Supercomputing, 80(2): 2487–2513, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05561-0.

Meshram:2024:EPP

[2658] Chandrashekhar Meshram, Mohammad S. Obaidat, Rabha W. Ibrahim, Sarita Gajbhiye Meshram, and Arpit Vijay Raikwar. An efficient privacypreserved authentication technique based on conformable fractional chaotic map for TMIS under smart homes environments. The Journal of Supercomputing, 80(2):2514– 2537, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05536-1.

Singh:2024:FFA

[2659] Tinka Singh and Ramagopal V. S. Uppaluri. Feed-forward ANN and traditional machine learning-based prediction of biogas generation rate from meteorological and organic waste parameters. The Journal of Supercomputing, 80(2):2538– 2571, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05569-6.

Darabkh:2024:NRP

[2660] Khalid A. Darabkh, Thara A. Alzboun, and Marwa H. Al-Tahaineh. A novel routing protocol for stable route selection in IoT networks. The Journal of Supercomputing, 80(2):2572– 2600, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05570-z.

Ghasemi:2024:NGN

[2661] Bahar Ghasemi, Mahmood Ahmadi, and Hamed Alimohammadi. A neural gas network-based scheme for SDN many-field packet classification. The Journal of Supercomputing, 80(2): 2601–2632, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05564-x.

Azzouzi:2024:NAE

[2662] Oussama Azzouzi, Mohamed Anane, Mouloud Koudil, Mohamed Issad, and Yassine Himeur. Novel area-efficient and flexible architectures for optimal ate pairing on FPGA. The Journal of Supercomputing, 80(2):2633– 2659, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05578-5.

Kamboj:2024:RLB

[2663] Anil Kumar Kamboj, Poonam Jindal, and Pankaj Verma. Reinforcement learning-based secure joint relay and jammer selection in dualhop wireless networks. The Journal of Supercomputing, 80(2):2660– 2680, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05575-8.

Rani:2024:DBB

[2664] Divya Rani and Sachin Tripathi. Design of blockchain-based authentication and key agreement protocol for health data sharing in cooperative hospital network. The Journal of Supercomputing, 80(2):2681– 2717, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05577-6.

Arepalli:2024:DLE

[2665] Peda Gopi Arepalli and K. Jairam Naik. A deep learning-enabled IoT framework for early hypoxia detection in aqua water using light weight spatially shared attention-LSTM network. The Journal of Supercomputing, 80(2): 2718–2747, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05580-x.

Zhou:2024:NMD

[2666] Bo Zhou, Hai Huang, Jun Xia, and Donghai Tian. A novel malware detection method based on API embedding and API parameters. The Journal of Supercomputing, 80(2): 2748–2766, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05556-x.

Qi:2024:RBP

[2667] Lijun Qi and Jincheng Zhuang. RLWEbased public key searchable encryption: securer, faster, and lower endto-end delay for cloud computing. The Journal of Supercomputing, 80(2): 2767–2798, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05574-9.

Wang:2024:DBC

[2668] Youwei Wang, Lizhou Feng, Ao Liu, Weiqi Wang, and Yudong Hou. Dual BiGRU–CNN-based sentiment classification method combining global and local attention. The Journal of Supercomputing, 80(2):2799– 2837, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05558-9.

Cheng:2024:BGC

[2669] Tengteng Cheng, Guosun Zeng, and Zhipeng Sun. A big graph clustering method to support parallel processing by perceiving graph’s application algorithm semantics. The Journal of Supercomputing, 80(2):2838– 2861, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05572-x.

Sivapriya:2024:CAD

[2670] G. Sivapriya, R. Manjula Devi, and P. Keerthika. Computer-aided diagnostic model for retinal vascular diseases using graph-based attention mechanism. The Journal of Supercomputing, 80(2):2862– 2888, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05581-w.

Tang:2024:SSE

[2671] Yuxin Tang, Yudi Qiu, Yanwei Liu, Jie Jiao, Peng Zhang, and Yibo Fan. Scalable short-entry dualgrain coherence directories with flexible region granularity. The Journal of Supercomputing, 80(2):2889– 2911, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05559-8.

Singh:2024:CFT

[2672] Raushan Kumar Singh and Mukesh Kumar. Correction to: Future trends of path planning framework considering accident attributes for smart cities. The Journal of Supercomputing, 80(2):2912, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05554-z.  See

[2411].

Kavianpour:2024:CCB

[2673] Parisa Kavianpour, Mohammadreza Kavianpour, Ehsan Jahani, and Amin Ramezani. Correction to: A CNN– BiLSTM model with attention mechanism for earthquake prediction. The Journal of Supercomputing, 80(2): 2913, January 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05497-5.  See

[2494].

Biri:2024:FFP

[2674] Ismail Biri, Ulku Tuncer Kucuktas, Fatih Uysal, and Firat Hardalac. Forecasting the future popularity of the anti-vax narrative on Twitter with machine learning. The Journal of Supercomputing, 80(3):2917– 2947, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05567-8.

Chen:2024:IYR

[2675] Jiaxin Chen, Xuguang Zhang, Yinggan Tang, and Hui Yu. ICE-YoloX: research on face mask detection algorithm based on improved YoloX network. The Journal of Supercomputing, 80(3):2948– 2969, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05585-6.

Awasthi:2024:IMT

[2676] Divyanshu Awasthi, Priyank Khare, and Vinay Kumar Srivastava. Internet of medical things-based authentication for an optimized watermarking of encrypted EEG. The Journal of Supercomputing, 80(3):2970– 3004, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05566-9.

Al-Haddad:2024:FDA

[2677] Luttfi A. Al-Haddad, Alaa Abdulhady Jaber, Sinan A. Al-Haddad, and Yousif M. Al-Muslim. Fault diagnosis of actuator damage in UAVs using embedded recorded data and stacked machine learning models. The Journal of Supercomputing, 80(3): 3005–3024, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05584-7.

Wang:2024:SSD

[2678] Guangjin Wang, Yuanying Wang, Fuyong Xu, Yongsheng Zhang, and Peiyu Liu. Syntactic and semantic dualenhanced bidirectional network for aspect sentiment triplet extraction. The Journal of Supercomputing, 80(3):

3025–3041, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05573-w.

Zhang:2024:AAE

[2679] Hongxia Zhang, Mengde Wu, Qiqi Feng, and Hao Li. AERQP: adaptive embedding representationbased QoS prediction for web service recommendation. The Journal of Supercomputing, 80(3):3042– 3065, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05582-9.

Ahmadi:2024:EBF

[2680] Arash Ahmadi and Amanj Khorramian. Efficient brute-force state space search for Yin–Yang puzzle. The Journal of Supercomputing, 80(3): 3066–3088, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05565-w.

Maheswari:2024:CBD

[2681] K. Uma Maheswari, S. Mary Saira Bhanu, and Nickolas Savarimuthu. Clustering-based data integrity verification approach for multi-replica in a fog environment. The Journal of Supercomputing, 80(3):3089– 3113, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05576-7.

Chen:2024:HCR

[2682] Yilu Chen, Miaodi Su, Hongzhi Ding, Shaohong Weng, Zhifeng Lin, and Xiqiong Bai. High-correlation 3D routability estimation for congestionguided global routing. The Journal of Supercomputing, 80(3):3114– 3141, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05553-0.

Wang:2024:PVD

[2683] Xuesong Wang, Zheng Zhang, Dongsheng Zhang, and Limin Ding. Pose visual detection method for cellphone dropping process incorporating prior information. The Journal of Supercomputing, 80(3):3142– 3161, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05527-2.

Xu:2024:YAL

[2684] Ruihang Xu, Peisen Wang, Xuanjing Li, and Ruifeng Nie. YOLO-ARGhost: a lightweight face mask detection model. The Journal of Supercomputing, 80(3):3162–3182, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05588-3.

Bai:2024:LBS

[2685] Mei Bai, Qibo Wang, Shihan Chang, Xite Wang, and Guanyu Li. Locationbased skyline query processing technology in road networks. The Journal of Supercomputing, 80(3):3183– 3211, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05563-y.

Yang:2024:ATS

[2686] Yanni Yang, Yiting Tan, Jintao Min, and Zhengwei Huang. Automatic text summarization for government news reports based on multiple features. The Journal of Supercomputing, 80(3): 3212–3228, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05599-0.

Khouni:2024:NOA

[2687] Saif Eddine Khouni and Tidjani Menacer. Nizar optimization algorithm: a novel metaheuristic algorithm for global optimization and engineering applications. The Journal of Supercomputing, 80(3):3229– 3281, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05579-4.

Zhang:2024:CCI

[2688] Lin Zhang, Fang’ai Liu, Hongchen Wu, Xuqiang Zhuang, and Yaoyao Yan. CFF: combining interactive features and user interest features for click-through rate prediction. The Journal of Supercomputing, 80(3): 3282–3309, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05598-1.

Liu:2024:ELR

[2689] Xiao Liu, Junlong Chen, Deyu Qi, and Tong Zhang. Exploration of low-resource language-oriented machine translation system of genetic algorithm-optimized hyper-task network under cloud platform technology. The Journal of Supercomputing, 80(3): 3310–3333, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05604-6.

Tang:2024:SSP

[2690] Jianxin Tang, Shihui Song, Hongyu Zhu, Qian Du, and Jitao Qu. Sequential seeding policy on social influence maximization: a Qlearning-driven discrete differential evolution optimization. The Journal of Supercomputing, 80(3):3334– 3359, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05601-9.

Wang:2024:LVM

[2691] Junfan Wang, Yi Chen, Yeting Gu, Yunfeng Yan, Qi Li, Mingyu Gao, and Zhekang Dong. A lightweight vehicle mounted multi-scale traffic sign detector using attention fusion pyramid. The Journal of Supercomputing, 80(3): 3360–3381, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05594-5.

Gu:2024:SSN

[2692] Xu Gu, Xiaoliang Chen, Peng Lu, Xiang Lan, Xianyong Li, and Yajun Du. SiMaLSTM-SNP: novel semantic relatedness learning model preserving both Siamese networks and membrane computing. The Journal of Supercomputing, 80(3):3382– 3411, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05592-7.

Ding:2024:SMS

[2693] Xinnan Ding, Shan Du, Yu Zhang, and Kejun Wang. Spatiotemporal multi-scale bilateral motion network for gait recognition. The Journal of Supercomputing, 80(3):3412– 3440, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05607-3.

Belkhiri-Brahmi:2024:CCB

[2694] Louiza Belkhiri-Brahmi, Samira Yessad, and Fouzi Semchedine. Congestion control-based sink MOBility pattern for data gathering optimization in WSN. The Journal of Supercomputing, 80(3):3441– 3479, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05596-3.

Li:2024:MPS

[2695] Wei Li, Yetong Gao, and Lei Wang. A multipopulation particle swarm optimization based on divergent guidance and knowledge transfer for multimodal multiobjective problems. The Journal of Supercomputing, 80(3):

3480–3527, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05624-2.

Khobdeh:2024:BAR

[2696] Soroush Babaee Khobdeh, Mohammad Reza Yamaghani, and Siavash Khodaparast Sareshkeh. Basketball action recognition based on the combination of YOLO and a deep fuzzy LSTM network. The Journal of Supercomputing, 80(3):3528– 3553, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05611-7.

Dwivedi:2024:NAS

[2697] Rajesh Dwivedi, Aruna Tiwari, Neha Bharill, Milind Ratnaparkhe, Parul Mogre, Pranjal Gadge, and Kethavath Jagadeesh. A novel Apache Spark-based 14-dimensional scalable feature extraction approach for the clustering of genomics data. The Journal of Supercomputing, 80(3): 3554–3588, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05602-8.

Yuan:2024:SCC

[2698] Quan Yuan, Shanshan Wang, Mai Hu, and Liang Zeng. SLDChOA: a comprehensive and competitive multi-strategy-enhanced chimp algorithm for global optimization and engineering design. The Journal of Supercomputing, 80(3):3589– 3643, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05617-1.

Liu:2024:CLF

[2699] Runsheng Liu, Wencong Cai, Junyang Zhang, Xiaoling Wu, Lilin Yang, and Kaiqing Luo. CF-lines: a fusing contour features optimization method for line segment detector. The Journal of Supercomputing, 80(3):3644– 3662, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05615-3.

Gao:2024:MSR

[2700] Nan Gao, Bowei Yang, Peng Chen, and Liping Qian. A multi-stage recognizer for nested named entity with weakly labeled data. The Journal of Supercomputing, 80(3):3663– 3693, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05619-z.

Farahani:2024:NDE

[2701] Samira Shirinabadi Farahani, Mohammad Reza Reshadinezhad, and Seyed Erfan Fatemieh. New design for error-resilient approximate multipliers used in image processing in CNTFET technology. The Journal of Supercomputing, 80(3):3694– 3712, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05623-3.

Ahila:2024:EDL

[2702] A. Ahila, V. Prema, S. Ayyasamy, and M. Sivasubramanian. An enhanced deep learning model for highspeed classification of plant diseases with bioinspired algorithm. The Journal of Supercomputing, 80(3):3713– 3737, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05622-4.

Cherbal:2024:SIT

[2703] Sarra Cherbal, Abdelhak Zier, Sara Hebal, Lemia Louail, and Boubakeur Annane. Security in Internet of Things: a review on approaches based on blockchain, machine learning, cryptography, and quantum computing. The Journal of Supercomputing, 80(3): 3738–3816, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05616-2.

Tang:2024:CDC

[2704] Yaqian Tang, Baolei Cheng, Yan Wang, Yuejuan Han, Jia Yu, and Jianxi Fan. Connectivity and diagnosability of a class of recursive networks. The Journal of Supercomputing, 80(3): 3817–3848, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05589-2.

Wang:2024:MTS

[2705] Zhenghong Wang, Fanhua Yu, Dan Wang, Taihui Liu, and Rongjun Hu. Multi-threshold segmentation of breast cancer images based on improved dandelion optimization algorithm. The Journal of Supercomputing, 80(3): 3849–3874, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05605-5.

Qi:2024:SBS

[2706] Liuling Qi, Junfeng Tian, Mengjia Chai, and Hongyun Cai. STFM: a blockchain sharding algorithm based on trust field model for heterogeneous Internet of Things. The Journal of Supercomputing, 80(3):3875– 3901, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05610-8.

Arrar:2024:CSL

[2707] Djihad Arrar, Nadjet Kamel, and Abdelaziz Lakhfif. A comprehensive survey of link prediction methods. The Journal of Supercomputing, 80(3): 3902–3942, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05591-8.

Qasem:2024:NHB

[2708] Sultan Noman Qasem. A novel honey badger algorithm with multilayer perceptron for predicting COVID-19 time series data. The Journal of Supercomputing, 80(3):3943– 3969, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05560-1.

Boureima:2024:DMN

[2709] Ismael Boureima, Manish Bhattarai, Maksim Eren, Erik Skau, Philip Romero, Stephan Eidenbenz, and Boian Alexandrov. Distributed out-of-memory NMF on CPU/GPU architectures. The Journal of Supercomputing, 80(3):3970– 3999, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05587-4. See cor-

rection [2771].

Bansal:2024:SGD

[2710] Neha Bansal, Shelly Sachdeva, and Lalit K. Awasthi. Schema generation for document stores using workload-driven approach. The Journal of Supercomputing, 80(3):4000– 4048, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05613-5.

Leem:2024:EMA

[2711] Subeen Leem, Jisong Oh, Jihoon Moon, Mucheol Kim, and Seungmin Rho. Enhancing multistep-ahead bikesharing demand prediction with a two-stage online learning-based timeseries model: insight from Seoul. The Journal of Supercomputing, 80(3): 4049–4082, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05593-6.

Aydemir:2024:ISC

[2712] Salih Berkan Aydemir. Ideal solution candidate search for starling murmuration optimizer and its applications on global optimization and engineering problems. The Journal of Supercomputing, 80(3):4083– 4156, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05618-0.

Abidi:2024:SMA

[2713] Rihab Abidi, Nadia Ben Azzouna, Wassim Trojet, Ghaleb Hoblos, and Nabil Sahli. A study of mechanisms and approaches for IoV trust models requirements achievement. The Journal of Supercomputing, 80(3): 4157–4201, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05620-6.

Zheng:2024:MGM

[2714] Xiaoyu Zheng, Dewang Chen, Zhiming Lin, Liping Zhuang, and Wendi Zhao. Method on generating massive virtual driving curves for highspeed trains of the Cross–Taiwan Strait Railway and its statistical analysis. The Journal of Supercomputing, 80(3): 4202–4225, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05621-5.

Khani:2024:ALB

[2715] Mohsen Khani, Shahram Jamali, and Mohammad Karim Sohrabi. Approximate Q-learning-based (AQL) network slicing in mobile edge-cloud for delay-sensitive services. The Journal of Supercomputing, 80(3):4226– 4247, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05614-4.

Yin:2024:ORS

[2716] Chao Yin, Qin Fang, Hongyi Li, Yingjian Peng, Xiaogang Xu, and Dan Tang. An optimized resource scheduling algorithm based on GA and ACO algorithm in fog computing. The Journal of Supercomputing, 80(3): 4248–4285, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05571-y.

Ketu:2024:IHC

[2717] Shwet Ketu and Pramod Kumar Mishra. An intelligent hybrid classification model for heart disease detection using imbalanced electrocardiogram signals. The Journal of Supercomputing, 80(3):4286– 4308, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05583-8.

Jiao:2024:MKC

[2718] Junjun Jiao, Zhisheng Chen, and Tao Zhou. Minimal Kapur cross-entropy-based image segmentation for distribution grid inspection using improved INFO optimization algorithm. The Journal of Supercomputing, 80(3): 4309–4352, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05628-y.

Liang:2024:MTC

[2719] Shengbin Liang, Fuqi Sun, Haoran Sun, Tingting Chen, and Wencai Du. A medical text classification approach with ZEN and capsule network. The Journal of Supercomputing, 80(3): 4353–4377, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05612-6.

Yadav:2024:PEC

[2720] Ashok Kumar Yadav, Karan Singh, and Pramod Kumar Srivastava. Probabilistic estimation of comprehensive utility based on user preference and network condition for network selection in future in HetNet. The Journal of Supercomputing, 80(3):4378– 4400, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05595-4.

Zhou:2024:DVG

[2721] Yuan Zhou, Lei Xiang, Fan Liu, Haoran Duan, and Yang Long. Dynamic visual-guided selection for zero-shot learning. The Journal of Supercomputing, 80(3):4401– 4419, February 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05625-1.

Mathivanan:2024:CCI

[2722] P. Mathivanan and Ponnambalam Maran. Correction to: Color image encryption based on novel kolam scrambling and modified 2D logistic cascade map (2D LCM). The Journal of Supercomputing, 80(3): 4420–4422, February 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05590-9.  See

[2648].

Li:2024:SCA

[2723] Di Li, Lang Li, and Yu Ou. Sidechannel analysis based on Siamese neural network. The Journal of Supercomputing, 80(4):4423–4450, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-023-05631-
3.

Zhang:2024:CPF

[2724] Hongwei Zhang, Shuaibo Liu, Chi Wang, Shuai Lu, and Wenbo Xiong. Color-patterned fabric defect detection algorithm based on triplet attention multi-scale u-shape denoising convolutional auto-encoder. The Journal of Supercomputing, 80(4): 4451–4476, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05639-9.

Reddy:2024:DLB

[2725] K. Hemant Kumar Reddy, Rajat Shubhra Goswami, and Diptendu Sinha Roy. A deep learningbased smart service model for contextaware intelligent transportation system. The Journal of Supercomputing, 80(4):4477–4499, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05597-2.

Liu:2024:AII

[2726] Yanru Liu, Jiali Xu, and Austin Lin Yee. The analysis of Iris image acquisition and real-time detection system using convolutional neural network. The Journal of Supercomputing, 80(4):4500–4532, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05629-x.

Bulbul:2024:OAN

[2727] Mehmet Akif Bülbül. Optimization of artificial neural network structure and hyperparameters in hybrid model by genetic algorithm: iOS-Android application for breast cancer diagnosis/prediction. The Journal of Supercomputing, 80(4):4533– 4553, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05635-z.

Kuwano:2024:PRU

[2728] Masashi Kuwano, Mio Hosoe, and Taku Moriyama. Prediction of railroad user count using number of route searches via bivariate state-space modeling. The Journal of Supercomputing, 80(4):4554–4576, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05642-0.

Song:2024:MSF

[2729] Yuqin Song, Jitao Zhao, and Chunliang Shang. A multi-stage feature fusion defogging network based on the attention mechanism. The Journal of Supercomputing, 80(4):4577– 4599, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05644-y.

Afzali:2024:ERA

[2730] Mahboubeh Afzali, Amin Mohammad Vali Samani, and Hamid Reza Naji. An efficient resource allocation of IoT requests in hybrid fog-cloud environment. The Journal of Supercomputing, 80(4):4600– 4624, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05586-5.

Deng:2024:HSP

[2731] Shao Deng, Shanzhu Xiao, Qiuqun Deng, and Huanzhang Lu. A hovering swarm particle swarm optimization algorithm based on node resource attributes for hardware/ software partitioning. The Journal of Supercomputing, 80(4):4625– 4647, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05603-7.

Ghosal:2024:DFA

[2732] Anit Kumar Ghosal, Amit Sardar, and Dipanwita Roy Chowdhury. Differential fault analysis attack-tolerant hardware implementation of AES. The Journal of Supercomputing, 80 (4):4648–4681, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05632-2.

Tian:2024:SEP

[2733] Min Tian, Qi Liu, Jingshan Pan, Ying Gou, and Zanjun Zhang. sw-PTS: an efficient parallel Thomas split algorithm for tridiagonal systems on Sunway manycore processors. The Journal of Supercomputing, 80 (4):4682–4706, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05641-1.

Jafari:2024:ARU

[2734] Sadiqa Jafari and Yung Cheol Byun. Accurate remaining useful life estimation of lithium-ion batteries in electric vehicles based on a measurable featurebased approach with explainable AI. The Journal of Supercomputing, 80 (4):4707–4732, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05648-8.

Sham:2024:MFS

[2735] Eht E Sham and Deo Prakash Vidyarthi. A modified fuzzy similarity measure for trapezoidal fuzzy number with their applications. The Journal of Supercomputing, 80(4): 4733–4759, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05608-2.

Alatrash:2024:CFI

[2736] Rawaa Alatrash, Rojalina Priyadarshini, and Hadi Ezaldeen. Collaborative filtering integrated fine-grained sentiment for hybrid recommender system. The Journal of Supercomputing, 80 (4):4760–4807, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05600-w.

Dabba:2024:NGW

[2737] Ali Dabba, Abdelkamel Tari, and Samy Meftali. A novel grey wolf optimization algorithm based on geometric transformations for gene selection and cancer classification. The Journal of Supercomputing, 80(4): 4808–4840, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05643-z.

Borode:2024:OAI

[2738] Adeola Borode and Peter Olubambi.

Optimisation of artificial intelligence models and response surface methodology for predicting viscosity and relative viscosity of GNP-alumina hybrid nanofluid: incorporating the effects of mixing ratio and temperature. The Journal of Supercomputing, 80 (4):4841–4869, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05652-y.

Sharma:2024:MAB

[2739] Neetu Sharma and Rajesh Rohilla. A multilevel authentication-based blockchain powered medicine anticounterfeiting for reliable IoT supply chain management. The Journal of Supercomputing, 80(4):4870– 4913, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05654-w.

Cao:2024:MMS

[2740] Ruyi Cao, Wanghao Mo, and Wendong Zhang. MFMDet: multiscale face mask detection using improved Cascade rcnn. The Journal of Supercomputing, 80(4):4914– 4942, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05653-x.

Sun:2024:DTB

[2741] Xiaoxiao Sun, Siqing Yang, Chenying Zhao, and Dongjin Yu. Designtime business process compliance assessment based on multi-granularity semantic information. The Journal of Supercomputing, 80(4):4943– 4971, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05626-0.

Lu:2024:GGT

[2742] Jingan Lu, Zhenfang Zhu, Kefeng Li, Shuai Gong, Hongli Pei, and Wenling Wang. Gtpsum: guided tensor product framework for abstractive summarization. The Journal of Supercomputing, 80(4):4972–4995, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05630-4.

Zhang:2024:SBD

[2743] Guifang Zhang and Hon-Cheng Wong.

Saliency-based dual-attention network for unsupervised video object segmentation. The Journal of Supercomputing, 80(4):4996–5010, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05637-x.

Landolsi:2024:HMN

[2744] Mohamed Yassine Landolsi, Lotfi Ben Romdhane, and Lobna Hlaoua. Hybrid medical named entity recognition using document structure and surrounding context. The Journal of Supercomputing, 80(4):5011– 5041, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05647-9.

Kang:2024:PIP

[2745] Yan Kang, Xiaoshe Dong, Ziheng Wang, Heng Chen, and Qiang Wang. Parallel implementations of postquantum Leighton–Micali signature on multiple nodes. The Journal of Supercomputing, 80(4):5042– 5072, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05662-w.

Jawaddi:2024:ICA

[2746] Siti Nuraishah Agos Jawaddi, Azlan Ismail, Muhammad Nur Haziq Mohammad Hatta, and Anis Faqihah Kamarulzaman. Insights into cloud autoscaling: a unique perspective through MDP and DTMC formal models. The Journal of Supercomputing, 80 (4):5073–5107, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05665-7.

Shree:2024:DPI

[2747] Shreyshi Shree, Chen Zhou, and Masoud Barati. Data protection in Internet of Medical Things using blockchain and secret sharing method. The Journal of Supercomputing, 80(4): 5108–5135, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05657-7.

Xiao:2024:LBS

[2748] Junbi Xiao, Xingjian Pan, Jianhang Liu, Jian Wang, Peiying Zhang, and Laith Abualigah. Load balancing strategy for SDN multi-controller clusters based on load prediction. The Journal of Supercomputing, 80(4): 5136–5162, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.

1007/s11227-023-05658-6.  See

[2820].

Gupta:2024:UWS

[2749] Swati Gupta and Niraj Pratap Singh.

Underwater wireless sensor networks: a review of routing protocols, taxonomy, and future directions. The Journal of Supercomputing, 80(4): 5163–5196, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05646-w.

Xiao:2024:TOA

[2750] Chunjing Xiao, Shenkai Lv, Wei Fan, and W. H. Ip. Temporal-order association-based dynamic graph evolution for recommendation. The Journal of Supercomputing, 80(4):5197– 5223, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05645-x.

Sheena:2024:EEM

[2751] N. Sheena, Shelbi Joseph, S. Shailesh, and Bharat Bhushan. EMCP: evolutionary multi-objective clustering protocol for prolonged lifetime of heterogeneous IoT networks. The Journal of Supercomputing, 80(4):5224– 5254, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05651-z.

Jafari:2024:OTU

[2752] Ehsan Jafari, Ardeshir Dolati, and Kamran Layeghi. Object tracking using local structural information and energy minimization. The Journal of Supercomputing, 80(4): 5255–5278, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05650-0.

Tong:2024:RSC

[2753] Yinsheng Tong, Zuoyong Li, Hui Huang, Libin Gao, Minghai Xu, and Zhongyi Hu. Research of spatial context convolutional neural networks for early diagnosis of Alzheimer’s disease. The Journal of Supercomputing, 80 (4):5279–5297, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05655-9.

Akan:2024:BRO

[2754] Taymaz Akan, Diego Oliva, Ali-Reza Feizi-Derakhshi, Amir-Reza Feizi-Derakhshi, Marco Pérez-Cisneros, and Mohammad Alfrad Nobel Bhuiyan. Battle royale optimizer for multilevel image thresholding. The Journal of Supercomputing, 80(4):5298– 5340, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05664-8.

Song:2024:HCS

[2755] Liqiang Song, Huaiguang Wang, and Baojian Yang. Hybrid compression scheme based on VMD optimization algorithm application to mechanical equipment monitoring. The Journal of Supercomputing, 80(4):5341– 5362, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05663-9.

Zhao:2024:SSD

[2756] Yue Zhao, Yuan Yao, Tong He, Xingshe Zhou, and Bo Shen. Sl4u: a scenario description language for unmanned swarm. The Journal of Supercomputing, 80(4):5363– 5389, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05636-y.

Sun:2024:ECP

[2757] Jingbo Sun, Weiming Peng, Tianbao Song, Haitao Liu, Shuqin Zhu, and Jihua Song. Enhanced crossprompt trait scoring via syntactic feature fusion and contrastive learning. The Journal of Supercomputing, 80 (4):5390–5407, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05640-2.

Sun:2024:HPB

[2758] Ru Sun and Ting Yang. Hybrid parameter-based PSO flexible needle percutaneous puncture path planning. The Journal of Supercomputing, 80 (4):5408–5427, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05661-x.

Yu:2024:CLU

[2759] Zhiyi Yu, Hong Li, and Jialin Feng.

Contrastive learning for unsupervised sentence embeddings using negative samples with diminished semantics. The Journal of Supercomputing, 80 (4):5428–5445, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05682-6.

Xu:2024:AES

[2760] Jin Xu, Huiqun Yu, Guisheng Fan, Jiayin Zhang, Zengpeng Li, and Qifeng Tang. Adaptive edge service deployment in burst load scenarios using deep reinforcement learning. The Journal of Supercomputing, 80(4): 5446–5471, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05656-8.

Liu:2024:UVA

[2761] Jing-Wei Liu, Ming-Hua Hsu, Chun-Liang Lai, and Sheng-K Wu. Using video analysis and artificial neural network to explore association rules and influence scenarios in elite table tennis matches. The Journal of Supercomputing, 80(4):5472– 5489, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05684-4.

Jie:2024:PGM

[2762] Chang Jie, Chen Jiming, Shao Ying, Tong Yanchun, and Ren Haodong. A pyramid GNN model for CXRbased COVID-19 classification. The Journal of Supercomputing, 80(4): 5490–5508, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05633-1.

Yang:2024:SHP

[2763] Fei Yang, Huyin Zhang, Shiming Tao, and Xiying Fan. Simple hierarchical PageRank graph neural networks. The Journal of Supercomputing, 80 (4):5509–5539, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05666-6.

Vaish:2024:AHC

[2764] Prashant Vaish, Niharika Anand, Vishal Krishna Singh, and Gaurav Sharma. Applications hosting over cloud-assisted IOT: a productivity model and method defining accessibility of data security. The Journal of Supercomputing, 80(4):5540– 5564, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05668-4.

Singh:2024:HOE

[2765] Raghuraj Singh and Kuldeep Kumar. Hybrid optimization-enabled deep Q network for fault prediction in service-oriented architecture. The Journal of Supercomputing, 80 (4):5565–5592, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05659-5.

Nie:2024:CLS

[2766] ZiXiang Nie, Jin Li, FengHui Duan, and Yueming Lu. A collaborative ledger storing model for lightweight blockchains based on Chord Ring. The Journal of Supercomputing, 80 (4):5593–5615, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05667-5.

Li:2024:PMQ

[2767] Zhen Li, Zining Cao, and Chao Xing. Performance modeling and quantitative evaluation for cyber-physical systems based on LTS. The Journal of Supercomputing, 80(4):5616– 5653, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05669-3.

Fan:2024:HCE

[2768] Weibei Fan, Xuanli Liu, and Mengjie Lv. Hamiltonian cycle embedding with fault-tolerant edges and adaptive diagnosis in half hypercube. The Journal of Supercomputing, 80(4): 5654–5674, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05674-6.

Johari:2024:PMA

[2769] Azad Johari and Azlan Ismail. PM-LAP: a methodology for annotating SSML elements into HTML5. The Journal of Supercomputing, 80(4): 5675–5707, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05673-7.

Reddy:2024:DPD

[2770] G. Viswanatha Reddy, B. S. N. V.

Chaitanya, P. Prathyush, M. Sumanth, C. Mrinalini, P. Dileep Kumar, and Snehasis Mukherjee. DFW-PP: dynamic feature weighting-based popularity prediction for social media content. The Journal of Supercomputing, 80(4):5708–5730, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05672-8.

Boureima:2024:CDM

[2771] Ismael Boureima, Manish Bhattarai, Maksim Eren, Erik Skau, Philip Romero, Stephan Eidenbenz, and Boian Alexandrov. Correction to: Distributed out-of-memory NMF on CPU/GPU architectures. The Journal of Supercomputing, 80(4):5731– 5732, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05675-5.  See

[2709].

Wang:2024:REM

[2772] Zhaojun Wang, Xindong You, and Xueqiang Lv. A relation enhanced model for temporal knowledge graph alignment. The Journal of Supercomputing, 80(5):5733– 5755, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05670-w.

Al-Janabi:2024:IRE

[2773] Samaher Al-Janabi and Ghada Mohammed. An intelligent returned energy model of cell and grid using a gain sharing knowledge enhanced long short-term memory neural network. The Journal of Supercomputing, 80 (5):5756–5814, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05609-1.

Li:2024:NGL

[2774] Peng Li, Cheng Che, and Rui Hou. Nacc-Guard: a lightweight DNN accelerator architecture for secure deep learning. The Journal of Supercomputing, 80(5):5815– 5831, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05671-9.

Huang:2024:ICP

[2775] Kangjin Huang, Yonggang Che, Chuanfu Xu, Zhe Dai, and Jian Zhang. Improving CUDA performance of an unstructured high-order CFD application under OP2 framework. The Journal of Supercomputing, 80(5): 5832–5846, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05679-1.

Xiang:2024:ECA

[2776] Yueting Xiang, Yangfen Chen, Wenjuan Fan, and Hui Ye. Enhancing computer-aided translation system with BiLSTM and convolutional neural network using a knowledge graph approach. The Journal of Supercomputing, 80(5):5847– 5869, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05686-2.

Ullah:2024:PID

[2777] Inam Ullah, Asra Noor, Shah Nazir, Farhad Ali, Yazeed Yasin Ghadi, and Nida Aslam. Protecting IoT devices from security attacks using effective decision-making strategy of appropriate features. The Journal of Supercomputing, 80(5):5870– 5899, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05685-3.

Muthulakshmi:2024:IFS

[2778] S. Muthulakshmi and R. Chitra. Interplanetary file system and blockchain for secured smart grid networks. The Journal of Supercomputing, 80(5): 5900–5922, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05680-8.

Rammal:2024:UAO

[2779] Abbas Rammal, Kenneth Ezukwoke, Anis Hoayek, and Mireille Batton-Hubert. Unsupervised approach for an optimal representation of the latent space of a failure analysis dataset. The Journal of Supercomputing, 80 (5):5923–5949, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05634-0.

Bushra:2024:NJB

[2780] S. Nikkath Bushra, Nalini Subramanian, G. Shobana, and S. Radhika. A novel Jarratt butterfly Ebola optimization-based attentional random forest for data anonymization in cloud environment. The Journal of Supercomputing, 80(5):5950– 5978, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05606-4.

Alp:2024:MMO

[2781] Sait Alp, Rahim Dehkharghani, Taymaz Akan, and Mohammad A. N. Bhuiyan. MOBRO: multi-objective battle royale optimizer. The Journal of Supercomputing, 80(5):5979– 6016, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05676-4.

Shilpi:2024:AJA

[2782] Shilpi and Arvind Kumar. Application of Jaya algorithm for solving localization problem in a distributed Wireless Sensor Network. The Journal of Supercomputing, 80 (5):6017–6041, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05683-5.

Chang:2024:MMI

[2783] Rui yun Chang, Xiu fang Feng, Cheng-Ye Zou, and Hao Zhang. Multimodal medical image protection algorithm based on 3D-CICCM and Fresnel transform. The Journal of Supercomputing, 80(5):6042–6067, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05703-4.

Li:2024:CBM

[2784] Fanfan Li. Chord-based music generation using long short-term memory neural networks in the context of artificial intelligence. The Journal of Supercomputing, 80(5):6068– 6092, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05704-3.

Ma:2024:TAN

[2785] Xiaoxuan Ma, Boyang Sun, Weifeng Liu, Dong Sui, Sihan Shan, Jing Chen, and Zhaofeng Tian. Tnseg: adversarial networks with multi-scale joint loss for thyroid nodule segmentation. The Journal of Supercomputing, 80 (5):6093–6118, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05689-z.

Geng:2024:NPP

[2786] Hongjie Geng and Mingming Zhou.

Novel post-photographic technique based on deep convolutional neural network and blockchain technology. The Journal of Supercomputing, 80 (5):6119–6139, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05705-2.

Xiong:2024:DPC

[2787] Jingwen Xiong, Wenke Zang, Yuzhen Zhao, and Xiyu Liu. Density peaks clustering algorithm with connected local density and punished relative distance. The Journal of Supercomputing, 80(5):6140–6168, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05688-0.

Liu:2024:ILW

[2788] Huan Liu, Jian Wu, and Rui He. IDP-Net: a light-weight network and its variants for human pose estimation. The Journal of Supercomputing, 80 (5):6169–6191, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05691-5.

Wang:2024:LSI

[2789] Qianqiu Wang, Ge Nong, and Wenbo Wu. Linear structure index for network-constrained moving objects. The Journal of Supercomputing, 80 (5):6192–6220, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05690-6.

Chen:2024:FGB

[2790] Tianhua Chen, Yanyue Li, and Qinghua Qiao. Fine-grained bird image classification based on counterfactual method of vision transformer model. The Journal of Supercomputing, 80(5):6221–6239, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05701-6.

Wang:2024:TOM

[2791] Jubin Wang, Yuan Zhuang, and Yunhui Zeng. A transmission optimization method for MPI communications. The Journal of Supercomputing, 80 (5):6240–6263, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05699-x.

Bakry:2024:ASF

[2792] Ahmed N. Bakry, Almohammady S.

Alsharkawy, Mohamed S. Farag, and K. R. Raslan. Automatic suppression of false positive alerts in anti-money laundering systems using machine learning. The Journal of Supercomputing, 80(5):6264–6284, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05708-z.

Sun:2024:EAT

[2793] Yifei Sun, Jigang Wu, Yalan Wu, Long Chen, and Weijun Sun. Efficient approaches for task offloading in pointof-interest based vehicular fog computing. The Journal of Supercomputing, 80 (5):6285–6310, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05698-y.

Fang:2024:TTT

[2794] Juan Fang, Zhichao Wei, Yaqi Liu, and Yumin Hou. TB-TBP: a taskbased adaptive routing algorithm for network-on-chip in heterogeneous CPU–GPU architectures. The Journal of Supercomputing, 80(5):6311– 6335, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05700-7.

Wu:2024:PAK

[2795] Huarui Wu, Chang Liu, and Chunjiang Zhao. Personalized agricultural knowledge services: a framework for privacy-protected user portraits and efficient recommendation. The Journal of Supercomputing, 80 (5):6336–6355, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05557-w.

Abbas:2024:MAB

[2796] Ali Abbas and Raza Hasan. A multi-attribute-based data forwarding scheme for delay tolerant networks. The Journal of Supercomputing, 80 (5):6356–6381, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05702-5.

Asghari:2024:EAS

[2797] Ali Asghari, Hossein Azgomi, Ali Abbas Zoraghchian, and Abbas Barzegarinezhad. Energy-aware server placement in mobile edge computing using trees social relations optimization algorithm. The Journal of Supercomputing, 80(5):6382–6410, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05692-4.

Shukla:2024:UTD

[2798] Neeraj Kumar Shukla, Abdulilah M.

Mayet, M. Ramkumar Raja, Muneer Parayangat, Mohammed Usman, Rajesh Verma, and Javed Khan Bhutto. A unified test data volume compression scheme for circular scan architecture using hosted cuckoo optimization. The Journal of Supercomputing, 80(5):6411–6434, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05638-w.

Naz:2024:ATL

[2799] Sumera Naz, Aqsa Shafiq, and Maheen Abbas. An approach for 2tuple linguistic q-rung orthopair fuzzy MAGDM for the evaluation of historical sites with power Heronian mean. The Journal of Supercomputing, 80 (5):6435–6485, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05678-2.

Kadum:2024:AEO

[2800] Muhanad Mohammed Kadum and Xiaoheng Deng. Allocating energyobjective aware workflow in distributed edge micro data centres. The Journal of Supercomputing, 80(5):6486– 6521, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05681-7.

Li:2024:SGI

[2801] Haiyan Li, Shaolin Peng, Xun Lang, Shuhua Ye, and Hongsong Li. Spatialguided informative semantic joint transformer for single-image deraining. The Journal of Supercomputing, 80(5):6522–6551, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05697-z.

Khojand:2024:CPS

[2802] Mahnaz Khojand, Kambiz Majidzadeh, Mohammad Masdari, and Yousef Farhang. Controller placement in SDN using game theory and a discrete hybrid metaheuristic algorithm. The Journal of Supercomputing, 80 (5):6552–6600, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05709-y.

Varposhti:2024:DHBa

[2803] Marzieh Varposhti. Distributed homology-based sensor selection and scheduling in wireless sensor networks. The Journal of Supercomputing, 80 (5):6601–6621, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05716-z.

Idrissi:2024:ABB

[2804] Hind Idrissi and Paolo Palmieri. Agentbased blockchain model for robust authentication and authorization in IoT-based healthcare systems. The Journal of Supercomputing, 80(5): 6622–6660, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05649-7.

Huang:2024:RCS

[2805] Yibo Huang, Bo Wang, Xiangrong Pu, Yian Li, and Qiuyu Zhang. Research on ciphertext speech biohashing authentication based on chaotic system and improved public chain. The Journal of Supercomputing, 80 (5):6661–6698, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05693-3.

Valadanzoj:2024:HSY

[2806] Zahra Valadanzoj, Hassan Daryanavard, and Abbas Harifi. High-speed YOLOv4-tiny hardware accelerator for self-driving automotive. The Journal of Supercomputing, 80(5):6699– 6724, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05713-2.

Ma:2024:DST

[2807] Hui Ma. Development of a smart tourism service system based on the Internet of Things and machine learning. The Journal of Supercomputing, 80 (5):6725–6745, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05719-w.

Hashim:2024:NNA

[2808] Ali Jameel Hashim, M. A. Balafar, and Jafar Tanha. NEAE: NeuroEvolution AutoEncoder for anomaly detection in internet traffic data. The Journal of Supercomputing, 80(5): 6746–6777, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05715-0.

Zhang:2024:BBP

[2809] Shiqiang Zhang and Dongzhi Cao. A blockchain-based provably secure anonymous authentication for edge computing-enabled IoT. The Journal of Supercomputing, 80(5):6778– 6808, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05696-0.

Ghafouri:2024:IMI

[2810] Saeid Ghafouri, Seyed Hossein Khasteh, and Seyed Omid Azarkasb. Influence maximization (IM) in complex networks with limited visibility using statistical methods. The Journal of Supercomputing, 80(5):6809– 6854, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05695-1.

Gowda:2024:TFT

[2811] Naveen Chandra Gowda, Sunilkumar S. Manvi, A. Bharathi Malakreddy, and Rajkumar Buyya. TAKM-FC: Two-way authentication with efficient Key Management in Fog Computing Environments. The Journal of Supercomputing, 80(5):6855– 6890, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05712-3.

Benarmas:2024:DLB

[2812] Redouane Benabdallah Benarmas and Kadda Beghdad Bey. A deep learningbased framework for road traffic prediction. The Journal of Supercomputing, 80(5):6891–6916, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05718-x.

Cheng:2024:MOD

[2813] Yuqing Cheng, Zhiying Cao, Xiuguo Zhang, Qilei Cao, and Dezhen Zhang. Multi objective dynamic task scheduling optimization algorithm based on deep reinforcement learning. The Journal of Supercomputing, 80(5): 6917–6945, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05714-1.

Varposhti:2024:DHBb

[2814] Marzieh Varposhti. Distributed homology-based algorithm for solving set k-cover problem in heterogeneous directional sensor networks. The Journal of Supercomputing, 80 (5):6946–6964, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05721-2.

Niu:2024:SSR

[2815] Yong Niu, Xing Xing, Zhichun Jia, Mindong Xin, and Junye Xing.

SMIGNN: social recommendation with multi-intention knowledge distillation based on graph neural network. The Journal of Supercomputing, 80(5): 6965–6988, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05720-3.

Patidar:2024:UDC

[2816] Mukesh Patidar, Ankit Jain, Keshav Patidar, Surendra Kumar Shukla, Ali H. Majeed, Namit Gupta, and Nilesh Patidar. An ultra-dense and cost-efficient coplanar RAM cell design in quantum-dot cellular automata technology. The Journal of Supercomputing, 80(5):6989– 7027, March 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05722-1.

Nedjah:2024:DHD

[2817] Nadia Nedjah, Sérgio Raposo, and Luiza de Macedo Mourelle. Dedicated hardware design for efficient quantum computations using classical logic gates. The Journal of Supercomputing, 80(5):7028–7070, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05687-1.

Yang:2024:RRT

[2818] Kun Yang, Lin Jiang, Rui Shan, Kangle Li, and Xinyue Cui. RMSRM: real-time monitoring-based self-reconfiguration mechanism in reconfigurable PE array. The Journal of Supercomputing, 80(5):7071–7101, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05707-0.

Garcia:2024:ECM

[2819] Donaldo Garcia, José de Jesús Rubio, Humberto Sossa, Jaime Pacheco, Guadalupe Juliana Gutierrez, and Carlos Aguilar-Ibañez. Electricity consumption modeling by a chaotic convolutional radial basis function network. The Journal of Supercomputing, 80(5):7102–7119, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05733-y.

Xiao:2024:CLB

[2820] Junbi Xiao, Xingjian Pan, Jianhang Liu, Jian Wang, Peiying Zhang, and Laith Abualigah. Correction to: Load balancing strategy for SDN multicontroller clusters based on load prediction. The Journal of Supercomputing, 80(5):7120–7121, March 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05710-5.  See

[2748].

Wang:2024:UFA

[2821] Junfeng Wang. The use of fuzzy authentication integrated with convolutional neural networks in digital content protection. The Journal of Supercomputing, 80(6):7123– 7146, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05738-7.

Wang:2024:MSA

[2822] Xiaohong Wang, Xu Zhao, Kun Xu, and Shihao Xu. Multi-scale adaptive atrous graph convolution for point cloud analysis. The Journal of Supercomputing, 80(6):7147– 7170, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05694-2.

Wu:2024:HAC

[2823] Lisheng Wu, Xiaoming You, and Sheng Liu. Heterogeneous ant colony algorithm based on selective evolution mechanism and game strategy. The Journal of Supercomputing, 80 (6):7171–7206, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05706-1.

Almaspoor:2024:DIV

[2824] Mohammad Hassan Almaspoor, Ali A. Safaei, Afshin Salajegheh, and Behrouz Minaei-Bidgoli. Distributed independent vector machine for big data classification problems. The Journal of Supercomputing, 80(6):7207– 7244, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05711-4.

He:2024:MDC

[2825] Xinfeng He and Riyang Li. Malware detection for container runtime based on virtual machine introspection. The Journal of Supercomputing, 80(6):7245–7268, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05727-w.

Song:2024:RRO

[2826] Kai Song, Huaqiong Ma, Haiming Zhang, and Liping Yan. Research of ReLU output device in ternary optical computer based on parallel fully connected layer. The Journal of Supercomputing, 80(6):7269– 7292, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05737-8.

Song:2024:LLB

[2827] Yuqin Song, Chunliang Shang, and Jitao Zhao. LBCNet: A lightweight bilateral cascaded feature fusion network for real-time semantic segmentation. The Journal of Supercomputing, 80(6):7293–7315, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05740-z.

Khan:2024:GGL

[2828] Saif Ur Rehman Khan, Ming Zhao, Sohaib Asif, Xuehan Chen, and Yusen Zhu. GLNET: global-local CNN’sbased informed model for detection of breast cancer categories from histopathological slides. The Journal of Supercomputing, 80(6):7316– 7348, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05742-x.

Wang:2024:TOL

[2829] Chenghao Wang, Zhongqiang Luo, and Ziyuan Qi. Transformer oil leakage detection with sampling-WIoU module. The Journal of Supercomputing, 80(6):7349–7368, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05748-5.

Chen:2024:UPN

[2830] Yen-Liang Chen, Chia-Chi Wu, and Po-Cheng Shih. Using personalized next session to improve session-based recommender systems. The Journal of Supercomputing, 80(6):7369– 7398, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05743-w.

Zhang:2024:EDD

[2831] Jing Zhang, Zuan yang Zeng, Kun liang Si, and Xiu cai Ye. Entropy-driven differential privacy protection scheme based on social graphlet attributes. The Journal of Supercomputing, 80 (6):7399–7432, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05751-w.

Antony:2024:NMS

[2832] Blessy Antony and S. Revathy. A novel model for Sybil attack detection in online social network using optimal three-stream double attention network. The Journal of Supercomputing, 80(6):7433–7482, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05677-3.

Shi:2024:CVP

[2833] Jun Shi, Peiyi Zhang, Sihan Du, Wanyong Liang, Weifeng Cao, Qingbo Li, and Hechao Hou. Chipping value prediction for dicing saw based on sparrow search algorithm and neural networks. The Journal of Supercomputing, 80(6):7483–7506, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05728-9.

Emami:2024:BBP

[2834] Ashkan Emami, Ghazaleh Keshavarz Kalhori, Sheyda Mirzakhani, and Mohammad Ali Akhaee. A blockchainbased privacy-preserving anti-collusion data auction mechanism with an off-chain approach. The Journal of Supercomputing, 80(6):7507– 7556, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05736-9.

Li:2024:AAR

[2835] Dan Li, Wenbo Shi, Ning Lu, Sang-Su Lee, and Sokjoon Lee. ARdetector: Android ransomware detection framework. The Journal of Supercomputing, 80(6):7557–7584, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05741-y.

Wang:2024:PPE

[2836] Yuanxiang Wang, Zhen Xu, and Lei Yang. Plus-profile energy harvested prediction and adaptive energy management for solar-powered wireless sensor networks. The Journal of Supercomputing, 80(6):7585– 7603, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05755-6.

Zhu:2024:TCG

[2837] Chao Zhu, Benshun Yi, and Laigan Luo. Triple-channel graph attention network for improving aspectlevel sentiment analysis. The Journal of Supercomputing, 80(6):7604– 7623, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05745-8.

Chen:2024:HFS

[2838] You-Shyang Chen, Jieh-Ren Chang, Yaswanth P. K. Thotakura, and Ashraf Mohammad. A hybrid four-stage detection model to pre-identify the sustainable manufacturing process of Li-ion battery pack. The Journal of Supercomputing, 80(6):7624– 7661, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05757-4.

Guan:2024:VAE

[2839] Yaonan Guan, Yunwen Xu, Yugeng Xi, and Dewei Li. Variational auto encoder fused with Gaussian process for unsupervised anomaly detection. The Journal of Supercomputing, 80 (6):7662–7690, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05747-6.

Zahedy:2024:RRN

[2840] Niloofar Zahedy, Behrang Barekatain, and Alfonso Ariza Quintana. RI-RPL: a new high-quality RPL-based routing protocol using Q-learning algorithm. The Journal of Supercomputing, 80 (6):7691–7749, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05724-z.

Shobeiri:2024:PAH

[2841] Peyman Shobeiri, Mehdi Akbarian Rastaghi, Saeid Abrishami, and Behnam Shobiri. PCP-ACO: a hybrid deadline-constrained workflow scheduling algorithm for cloud environment. The Journal of Supercomputing, 80 (6):7750–7780, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05753-8.

Angbera:2024:AXB

[2842] Ature Angbera and Huah Yong Chan.

An adaptive XGBoost-based optimized sliding window for concept drift handling in non-stationary spatiotemporal data streams classifications. The Journal of Supercomputing, 80(6): 7781–7811, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05729-8.

Rostami:2024:EET

[2843] Safdar Rostami, Ali Broumandnia, and Ahmad Khademzadeh. An energyefficient task scheduling method for heterogeneous cloud computing systems using capuchin search and inverted ant colony optimization algorithm. The Journal of Supercomputing, 80(6):7812–7848, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05725-y.

AlShahrani:2024:BEF

[2844] Ali M. Al Shahrani, Ali Rizwan, Manuel Sánchez-Chero, Lilia Lucy Campos Cornejo, and Mohammad Shabaz. Blockchain-enabled federated learning for prevention of power terminals threats in IoT environment using edge zero-trust model. The Journal of Supercomputing, 80 (6):7849–7875, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05763-6.

Ali:2024:IIA

[2845] Bakht Sher Ali, Inam Ullah, Tamara Al Shloul, Izhar Ahmed Khan, Ijaz Khan, Yazeed Yasin Ghadi, Akmalbek Abdusalomov, Rashid Nasimov, Khmaies Ouahada, and Habib Hamam. ICS-IDS: application of big data analysis in AI-based intrusion detection systems to identify cyberattacks in ICS networks. The Journal of Supercomputing, 80(6):7876–7905, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05764-5.

Moradikashkooli:2024:EOA

[2846] A. Moradikashkooli, H. Haj Seyyed Javadi, and S. Jabbehdari. An efficient optimization algorithm for nonlinear 2D fractional optimal control problems. The Journal of Supercomputing, 80(6):7906–7930, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05732-z.

Liu:2024:DCD

[2847] Yang Liu. Depression clinical detection model based on social media: a federated deep learning approach. The Journal of Supercomputing, 80 (6):7931–7954, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05754-7.

Mittal:2024:TPF

[2848] Shweta Mittal and Mohona Ghosh. A three-phase framework for secure storage and sharing of healthcare data based on blockchain, IPFS, proxy reencryption and group communication. The Journal of Supercomputing, 80 (6):7955–7992, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05735-w.

Singh:2024:RID

[2849] Tajinder Singh, Madhu Kumari, and Daya Sagar Gupta. Rumor identification and diffusion impact analysis in real-time text stream using deep learning. The Journal of Supercomputing, 80(6):7993–8037, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05726-x.

Han:2024:SYL

[2850] Jingjing Han, Guangqi Yang, Hongyang Wei, Weijun Gong, and Yurong Qian. ST-YOLOX: a lightweight and accurate object detection network based on Swin Transformer. The Journal of Supercomputing, 80(6):8038– 8059, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05744-9.

Deng:2024:FDC

[2851] Xiangyu Deng, Xikai Huang, and Haiyue Yu. Frequency-domain characteristic analysis of PCNN. The Journal of Supercomputing, 80(6): 8060–8093, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05750-x.

Sidhom:2024:TPH

[2852] Ones Sidhom, Haythem Ghazouani, and Walid Barhoumi. Three-phases hybrid feature selection for facial expression recognition. The Journal of Supercomputing, 80(6):8094– 8128, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05758-3.

Roodbali:2024:HEL

[2853] Khadijeh Moeini Roodbali, Ebrahim Abiri, and Kourosh Hassanli. Highly efficient low-area gate-diffusion-input-based approximate full adders for image processing computing. The Journal of Supercomputing, 80(6):8129– 8155, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05768-1.

Kanwal:2024:HFD

[2854] Ayesha Kanwal, Kashif Javed, Sara Ali, Saddaf Rubab, Muhammad Attique Khan, Areej Alasiry, Mehrez Marzougui, and Mohammad Shabaz. A hybrid framework for detection of autism using ConvNeXt-T and embedding clusters. The Journal of Supercomputing, 80(6):8156– 8178, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05761-8.

Khalouta:2024:NDT

[2855] Ali Khalouta. A new decomposition transform method for solving nonlinear fractional logistic differential equation. The Journal of Supercomputing, 80(6):8179–8201, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05730-1.

Verma:2024:OFC

[2856] Nilesh Kumar Verma and K. Jairam Naik. Optimized fog community framework with advanced genetic algorithm for enhanced performance dynamics. The Journal of Supercomputing, 80 (6):8202–8235, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05769-0.

Feizabadi:2024:PNH

[2857] Reza Feizabadi, Mehri Bagherian, Hamidreza Vaziri, and Maziar Salahi. PLEACH: a new heuristic algorithm for pure parsimony haplotyping problem. The Journal of Supercomputing, 80(6):8236–8258, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05746-7.

Ferraris:2024:SIT

[2858] Davide Ferraris, Carmen Fernandez-Gago, Rodrigo Roman, and Javier Lopez. A survey on IoT trust model frameworks. The Journal of Supercomputing, 80(6):8259– 8296, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05765-4.

Shu:2024:ELP

[2859] Yiming Shu and Yiru Dai. An effective link prediction method for industrial knowledge graphs by incorporating entity description and neighborhood structure information. The Journal of Supercomputing, 80 (6):8297–8329, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05767-2.

Rohani:2024:EGB

[2860] Faezeh Rohani, Kamrad Khoshhal Roudposhti, Hamidreza Taheri, Ali Mashhadi, and Andreas Mueller. Extracting gait and balance pattern features from skeleton data to diagnose attention deficit/hyperactivity disorder in children. The Journal of Supercomputing, 80(6):8330– 8356, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05731-0.

Jayachitra:2024:TVB

[2861] J. Jayachitra, K. Suganya Devi, S. V. Manisekaran, and Satish Kumar Satti. Terahertz video-based hidden object detection using YOLOv5m and mutation-enabled salp swarm algorithm for enhanced accuracy and faster recognition. The Journal of Supercomputing, 80(6):8357– 8382, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05717-y.

Mahmud:2024:ASB

[2862] Minhaz Mahmud, Md. Soharab Hossain Sohan, Saha Reno, M. A. Baten Sikder, and Fakir Sharif Hossain. Advancements in scalability of blockchain infrastructure through IPFS and dual blockchain methodology. The Journal of Supercomputing, 80(6):8383– 8405, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05734-x.

Gupta:2024:IHW

[2863] Arti Gupta and Vijay Kumar Chaurasiya. Integrating Healthcare 4.0 and WBAN: efficient redundancy reduction and adaptive packet scheduling using AR– DRL. The Journal of Supercomputing, 80(6):8406–8433, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05752-9.

Nikooghadam:2024:RRE

[2864] Mahdi Nikooghadam, Haleh Amintoosi, and Hamid Reza Shahriari. REACH: Robust Efficient Authentication for Crowdsensing-based Healthcare. The Journal of Supercomputing, 80(6):8434–8468, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05749-4.

Li:2024:ABV

[2865] Shibao Li, Chenxu Ma, Yunwu Zhang, Longfei Li, Chengzhi Wang, Xuerong Cui, and Jianhang Liu. Attentionbased variable-size feature compression module for edge inference. The Journal of Supercomputing, 80(6): 8469–8484, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05779-y.

Jang:2024:HBI

[2866] Joonhyouk Jang, Minho Park, and Jiman Hong. Hybrid booting with incremental hibernation for the baseboard management controllers. The Journal of Supercomputing, 80(6): 8485–8504, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05762-7.

Deng:2024:OMD

[2867] Kaiyi Deng and Guanen Wang. Online mode development of Korean art learning in the post-epidemic era based on artificial intelligence and deep learning. The Journal of Supercomputing, 80(6):8505–8528, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05776-1.

Zhang:2024:GAN

[2868] Wei Zhang, Ping He, Chuntian Qin, Fan Yang, and Ying Liu. A graph attention network-based model for anomaly detection in multivariate time series. The Journal of Supercomputing, 80(6):8529–8549, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05772-5.

Yang:2024:CTN

[2869] Yi Yang, Jiaxuan Wei, Zhixuan Yu, and Ruisheng Zhang. Correction to: A trustworthy neural architecture search framework for pneumonia image classification utilizing blockchain technology. The Journal of Supercomputing, 80(6):8550, April 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05760-9.  See

[2635].

Tavallali:2024:CMT

[2870] Pooya Tavallali, Peyman Tavallali, and Mukesh Singhal. Correction: K-means tree: an optimal clustering tree for unsupervised learning. The Journal of Supercomputing, 80 (6):8551, April 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05723-0. See [728].

Guan:2024:PPC

[2871] Bowen Guan, Chencheng Zhao, Xianghui Yuan, Jun Long, and Xiang Li. Price prediction in China stock market: an integrated method based on time series clustering and image feature extraction. The Journal of Supercomputing, 80(7):8553– 8591, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05562-z.

Khelil:2024:TEW

[2872] Hathem Khelil and Mahmoud Brahimi. Toward an efficient web service composition based on an improved BTLBO algorithm. The Journal of Supercomputing, 80(7):8592– 8613, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05777-0.

Sun:2024:ACK

[2873] Xiaowen Sun, Zhenfang Zhu, Jiangtao Qi, Zhen Zhao, and Hongli Pei. Affective Commonsense Knowledge Enhanced Dependency Graph for aspect sentiment triplet extraction. The Journal of Supercomputing, 80 (7):8614–8636, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05778-z.

Shao:2024:CCC

[2874] Chao Shao and Hairui Zhang. Climate change characteristics and population health impact factors using deep neural network and hyperautomation mechanism. The Journal of Supercomputing, 80(7):8637–8667, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05795-y.

Gupta:2024:KAE

[2875] Khushboo Gupta and Vinod Kumar.

KMS-AMI: an efficient and scalable key management scheme for secure two-way communications in advanced metering infrastructure of smart grid. The Journal of Supercomputing, 80 (7):8668–8701, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05756-5.

Liu:2024:MPM

[2876] Xin-Duo Liu, Wei-Jia He, Ming-Lin Yang, and Xin-Qing Sheng. Massive parallelization of multilevel fast multipole algorithm for 3-D electromagnetic scattering problems on SW26010 many-core cluster. The Journal of Supercomputing, 80(7): 8702–8718, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05759-2.

Rubio:2024:ORB

[2877] José de Jesús Rubio, Marco Antonio Islas, Donaldo Garcia, Jaime Pacheco, Alejandro Zacarias, and Carlos Aguilar-Ibañez. Optimized radial basis function network for the fatigue driving modeling. The Journal of Supercomputing, 80(7):8719– 8741, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05775-2.

Li:2024:CIE

[2878] Youlin Li. Constructing the intelligent expressway traffic monitoring system using the Internet of Things and inspection robot. The Journal of Supercomputing, 80(7):8742– 8766, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05794-z.

Zhang:2024:DAC

[2879] Jie Zhang, Jingshun Bi, Jinyou Hou, and Qinggang Xie. Dynamical analysis, circuit realization, and applications of 4D hyperchaotic systems with bursty oscillations and infinite attractor coexistence. The Journal of Supercomputing, 80(7):8767– 8800, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05781-4.

Yang:2024:DPA

[2880] Yumeng Yang, Li Wang, and Zizhen Cheng. Density peaks algorithm based on information entropy and merging strategy for power load curve clustering. The Journal of Supercomputing, 80(7):8801–8832, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05793-0.

Yang:2024:VLC

[2881] Liwei Yang, Boyu Jia, Xue Liang, Fang Wang, and Xiangyuan Peng. Visible light communication and WiFi hybrid networks based on dynamic resource allocation algorithm. The Journal of Supercomputing, 80(7): 8833–8856, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05791-2.

Deng:2024:NHA

[2882] Xuzhen Deng, Dengxu He, and Liangdong Qu. A novel hybrid algorithm based on arithmetic optimization algorithm and particle swarm optimization for global optimization problems. The Journal of Supercomputing, 80 (7):8857–8897, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05773-4.

Puthiyidam:2024:EAS

[2883] Jiby J. Puthiyidam, Shelbi Joseph, and Bharat Bhushan. Enhanced authentication security for IoT client nodes through T-ECDSA integrated into MQTT broker. The Journal of Supercomputing, 80(7):8898– 8932, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05789-w.

Miao:2024:ADL

[2884] Yue Miao, Siyuan Tang, Zhuqiang Zhang, Jukun Song, Zhi Liu, Qiang Chen, and Miao Zhang. Application of deep learning and XG-Boost in predicting pathological staging of breast cancer MR images. The Journal of Supercomputing, 80 (7):8933–8953, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05797-w.

Sadeghi:2024:IHS

[2885] Mersedeh Sadeghi, Alessio Carenini, Oscar Corcho, Matteo Rossi, Riccardo Santoro, and Andreas Vogelsang. Interoperability of heterogeneous systems of systems: from requirements to a reference architecture. The Journal of Supercomputing, 80 (7):8954–8987, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05774-3.

Abbasi:2024:ADI

[2886] Negar Abbasi, Mohammadreza Soltanaghaei, and Farsad Zamani Boroujeni. Anomaly detection in IOT edge computing using deep learning and instance-level horizontal reduction. The Journal of Supercomputing, 80(7):8988– 9018, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05771-6.

Ansari:2024:NHT

[2887] Manzoor Ansari and Mansaf Alam. A novel hybrid time-series approach for IoT-cloud-enabled environment monitoring. The Journal of Supercomputing, 80(7):9019–9053, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05782-3.

Pan:2024:SBU

[2888] Dewei Pan, Zhijie Liu, Qiang Zhang, Ying Liu, Wencai Feng, Shiyu Ji, Wei Zhang, and Zhen Min. Simulationbased uprighting of a capsized ship in wave-induced environments. The Journal of Supercomputing, 80(7): 9054–9072, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05798-9.

Ouyang:2024:MMT

[2889] Fan Ouyang and Bo Shen. A mutual mean teacher framework for crossdomain aspect-based sentiment analysis. The Journal of Supercomputing, 80(7):9073–9095, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05792-1.

Lin:2024:AAR

[2890] Yu e Lin, Houguo Li, Xingzhu Liang, Mengfan Li, and Huilin Liu. AAR: Attention remodulation for weakly supervised semantic segmentation. The Journal of Supercomputing, 80(7): 9096–9114, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05786-z.

Wang:2024:CQE

[2891] Junhao Wang, Jiarong Liang, and Qingnian Li. On the construction of quality extended virtual backbones in wireless sensor networks using cooperative communication. The Journal of Supercomputing, 80(7):9115– 9139, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05799-8.

Ren:2024:GUL

[2892] Juanjuan Ren and Siti Salwa Salleh. Green urban logistics path planning design based on physical network system in the context of artificial intelligence . The Journal of Supercomputing, 80(7):9140–9161, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-023-05796-x.

Hou:2024:TBN

[2893] Ke Hou, Jianping Sun, Mingcheng Guo, Ming Pang, and Na Wang. TentISSA-BPNN: a novel evaluation model for cloud service providers for petroleum enterprises. The Journal of Supercomputing, 80(7):9162– 9193, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05803-1.

Zhu:2024:TTM

[2894] Fangyin Zhu, Wei Xu, Duanyang Liu, and Haiyan Shi. TSANET: transportation mode recognition model with global and local spatiotemporal features. The Journal of Supercomputing, 80(7):9194–9219, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05785-0.

Bomnale:2024:NUI

[2895] Archana Bomnale and Avinash More.

Node utilization index-based data routing and aggregation protocol for energy-efficient wireless sensor networks. The Journal of Supercomputing, 80(7):9220–9252, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05800-4.

Kumari:2024:ERS

[2896] Apurva Kumari and Subhendu Kumar Sahoo. An effective and robust single-image dehazing method based on gamma correction and adaptive Gaussian notch filtering. The Journal of Supercomputing, 80(7): 9253–9276, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05805-z.

Wang:2024:CMN

[2897] Xiaolong Wang, Ling Cai, and Yunhao Xu. Creation mechanism of new media art combining artificial intelligence and internet of things technology in a metaverse environment. The Journal of Supercomputing, 80 (7):9277–9297, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05819-7.

Tosini:2024:AEI

[2898] Marcelo Tosini, Martín Vázquez, and Lucas Leiva. Analysis and efficient implementation of IEEE-754 decimal floating point adders/subtractors in FPGAs for DPD and BID encoding. The Journal of Supercomputing, 80 (7):9298–9326, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05808-w.

She:2024:IMH

[2899] Lina She, Hongfang Gong, and Siyu Zhang. An interactive multi-head self-attention capsule network model for aspect sentiment classification. The Journal of Supercomputing, 80 (7):9327–9352, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05813-z.

Zhu:2024:CPT

[2900] Yi’an Zhu, Jie Chen, and Lian Li.

A critical path task scheduling algorithm based on sequential failure factor. The Journal of Supercomputing, 80(7):9353–9383, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05788-x.

Shirvani:2024:SST

[2901] Mirsaeid Hosseini Shirvani. A survey study on task scheduling schemes for workflow executions in cloud computing environment: classification and challenges. The Journal of Supercomputing, 80(7):9384– 9437, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05806-y.

Deepanjali:2024:TBI

[2902] S. Deepanjali and S. K. Noor Mahammad. A twofold bio-inspired system for mitigating SEUs in the controllers of digital system deployed on FPGA. The Journal of Supercomputing, 80 (7):9438–9470, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05804-0.

Tang:2024:SSO

[2903] Tao Tang, Kai Lu, Lin Peng, Yingbo Cui, Jianbin Fang, Chun Huang, Ruibo Wang, Canqun Yang, and Yifei Guo. SNCL: a supernode OpenCL implementation for hybrid computing arrays. The Journal of Supercomputing, 80(7):9471–9493, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05766-3.

Elsakaan:2024:NPA

[2904] Nadim Elsakaan and Kamal Amroun. A novel privacy-aware global infrastructure for ecological footprint calculator based on the Internet of Things and blockchain. The Journal of Supercomputing, 80(7):9494– 9531, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05821-z.

Ferreira:2024:EAP

[2905] Martha Dais Ferreira, Zahra Sadeghi, and Stan Matwin. Exploring autoregression patterns for automatic vessel type classification. The Journal of Supercomputing, 80(7):9532– 9553, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05826-8.

Wang:2024:UPA

[2906] Sheng Wang, Shiping Chen, and Yumei Shi. Utilization-prediction-aware energy optimization approach for heterogeneous GPU clusters. The Journal of Supercomputing, 80(7):9554– 9578, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05807-x.

Khan:2024:RSD

[2907] Habib Ullah Khan, Waseem Afsar, Shah Nazir, Asra Noor, Mahwish Kundi, Mashael Maashi, and Haya Mesfer Alshahrani. Revolutionizing software developmental processes by utilizing continuous software approaches. The Journal of Supercomputing, 80(7):9579– 9608, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05818-8.

Sabuwala:2024:AES

[2908] Noshin A. Sabuwala and Rohin D.

Daruwala. An approach to enhance the security of unmanned aerial vehicles (UAVs). The Journal of Supercomputing, 80(7):9609– 9639, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05811-1.

Fan:2024:FBD

[2909] Chenguang Fan, Muyao Li, Wenzhong Liu, and Jingjing Cheng. FPGA-based downhole real-time inversion of petrophysical information for NMR–LWD tools with periodic thermal management. The Journal of Supercomputing, 80(7):9640–9662, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05827-7.

Chowdary:2024:LPB

[2910] S. Dayasagar Chowdary and M. S.

Sudhakar. Linear programming-based multi-objective floorplanning optimization for system-on-chip. The Journal of Supercomputing, 80(7):9663– 9686, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05812-0.

Sheikhattar:2024:TMB

[2911] Mohammadreza Sheikhattar and Alireza Mansouri. A topic mapping-based framework to analyze textual risk reports from social media big data contents. The Journal of Supercomputing, 80(7):9687–9712, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05783-2.

Zhou:2024:SAM

[2912] Zhiheng Zhou, Xiaomei Huang, Naixue Xiong, Guoqiong Liao, and Xiaobin Deng. A self-attention model with contrastive learning for online group recommendation in event-based social networks. The Journal of Supercomputing, 80(7):9713–9741, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05801-3.

Mora:2024:ANR

[2913] H. Mora, M. T. Signes-Pont, F. A. Pujol López, J. Mora-Pascual, and J. M. García Chamizo. Advancements in number representation for high-precision computing. The Journal of Supercomputing, 80(7):9742– 9761, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05814-y.

Roy:2024:FMS

[2914] Sanjib Roy and Ayan Kumar Das.

Fuzzy miner selection toward blockchainbased secure communication using multifactor authentication. The Journal of Supercomputing, 80(7):9762– 9811, May 2024. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05844-6.

Kim:2024:CBA

[2915] JongHyuk Kim, Yong Moon, and Hoon Ko. Correlation-based advanced feature analysis for wireless sensor networks. The Journal of Supercomputing, 80(7):9812–9828, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05739-6.

Kim:2024:GTT

[2916] Hyunjun Kim and Hwansoo Han.

GPU thread throttling for page-level thrashing reduction via static analysis. The Journal of Supercomputing, 80(7):9829–9847, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05787-y.

Bano:2024:SPE

[2917] Shahzadi Bano, Weimei Zhi, Baozhi Qiu, Muhammad Raza, Nabila Sehito, Mian Muhammad Kamal, Ghadah Aldehim, and Nuha Alruwais. Selfpaced ensemble and big data identification: a classification of substantial imbalance computational analysis. The Journal of Supercomputing, 80(7):9848–9869, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05828-6.

Chen:2024:LEF

[2918] Yinsheng Chen, Jiahao Li, Kun Sun, and Ying Zhang. A lightweight early forest fire and smoke detection method. The Journal of Supercomputing, 80 (7):9870–9893, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05835-7.

Bai:2024:EER

[2919] Qingchun Bai, Mengmeng Tang, Yang Song, and Jun Xiao. Extracting entity and relationship interactions from danmaku-video comments using a neural bootstrapping framework. The Journal of Supercomputing, 80 (7):9894–9916, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05817-9.

Fu:2024:IPG

[2920] Weihong Fu and Haoyi Li. An improved phase gradient autofocus algorithm for ISAR phase autofocus problem. The Journal of Supercomputing, 80(7):9917–9934, May 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-023-05838-4.